A Complete Bibliography of Publications in the ICES Journal of Marine Science (2010–2019)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

15 October 2024 Version 1.13

Title word cross-reference

> [GUN^{+}19]. ^{13} [DKB14]. ^{14} [BTA^{+}18]. ^{15} [DKB14]. ^{86} [PBW15]. ^{87} [PBW15]. _{2} [BTB^{+}17, BIKP^{+}17, BWM^{+}16, BSHC17, BWH16, BMM^{+}19, CCE17, CPP17, DI11, DKPM16, EMW^{+}16, FGMC^{+}13, GQX^{+}19, GM17, HTRF17, HWR^{+}16, HBG^{+}16, HLMT16, IZP^{+}16, KPGH16, KTLB16, KPD^{+}16, LK17, LBG18, LSF16, LEE17, McE17, MWP^{+}16, MFB17, NMF^{+}16, PWCS17, PPS17, RCG^{+}18, RCF^{+}17, RFT^{+}16, SSH^{+}17, SLF^{+}16a, SMB^{+}16a, SY16, SJ16, SLF16b, TNFC16, VSB^{+}16, WWMF17]. a [ALd12, KYK^{+}12]. B_{MEY} [PSS^{+}14]. B_{MSY} [PSS^{+}14]. β [DCHMHQ12]. Δ [BTA^{+}18, DKB14]. F_{max} [LKS14a]. L_{inf} [FWC^{+}19b].

-dependent [TCBD10]. -driven [KTLB16]. -independent [PCLQ^{+}16]. -induced [GQX^{+}19, IZP^{+}16, LBG18].

/K [FWC^{+}19b].

0-group [EISJ12, HMM12]. ’09 [Rod10].

150-year [KOK^{+}14]. 18-kHz [SOK^{+}12]. 1838 [BF17, Man17a]. 1970s [CFT^{+}19, RYS11]. 1990s [GSP12]. 1991/1992 [MYI12]. 19th [JCS16, Pin17].

200 [GUN^{+}19]. 2000s [RYS11]. 2007/2008 [MYI12]. 2010 [GGK^{+}11]. 2010s [CFT^{+}19]. 2011 [WRC^{+}12]. 20th [DK18, Eer12, Pin17]. 21st [FKH^{+}16, KRG^{+}12]. 24^{o} [PG10].

3D [SKM18]. 3NO [NPGT15].

40-year [Ger17]. 40-year-long [CCA^{+}18]. 4D [HHR17]. 4th [LKG^{+}19].

5-MHz [KLN^{+}10]. 5-year [BUDM15, DD10]. 50-year [AFP12, Ósk18]. 500-kHz [Mel16]. 57^{o} [ACA19]. 5th [MNG15].

67 [MAP^{+}11].

71-8 [Ano14a]. 72 suppl 1 [Ano15]. 73 [BF17, Man17a]. 73-3 [Ano16a]. 73 suppl 1 [Ano16b]. 74-4 [Ano17].

8-fjords [BLJ^{+}17].

a4a [JMM^{+}15a]. abalone [BLM^{+}17]. Abdulrazzak [GGT^{+}14]. ability [FS14, RMJF14]. Abiotic [OBG12, GFSC17, LEB^{+}14, RLP^{+}13]. ABNJ [SJ18]. abnormal [THSM^{+}18]. abnormalities [OTAK15]. abrupt [KKAV^{+}19]. absence [PSH^{+}11]. Absolute [GML13]. absorption [HSFiS12]. absorption-based [HSFiS12]. Abundance [Che10, GS12, LSCL18, MONS14, AFLvDH15, BHC^{+}15, BKMdMS13, BHMR14, BRH^{+}15, BNP19, BMC^{+}13, BBD^{+}13, CBBM14, CN16, EDP14, ESD^{+}16, FGMC^{+}13, FCB17, FFPL14, aFKA11, GPG14, GHG^{+}10, HTM11, HAF^{+}16, HHJB14, HAH^{+}16, HKS12, IaF11, ISL19, JMF12, Kan11, KdS14, KIP14, KKLM13, LLM12, LGvPH15, LGR^{+}19, LDT^{+}11, LLF18, LBTS^{+}19, LHH^{+}13, MHLW12, MOG^{+}14a, MOG^{+}14b, MARD13, MBA^{+}19, NUÓ^{+}16, ØFNH12, OIMP18, OFB^{+}17, PSS^{+}14, RDB^{+}18a, RJ12, STKT16, SOK^{+}12, SGM^{+}17, SBCF16, SPS^{+}10, SS17, SBKA16, SRKV15, TSP12, TSWS15, TB17, TNS13, TID^{+}10, WBW^{+}18, WKL13, WWI^{+}16, YCZY19, ZKD^{+}14]. abundance-at-age [WWI^{+}16]. abundances [SFD^{+}16]. Abyssal [HLL^{+}12]. acanthias [AJM19, DED13]. Acanthocybium [ZGT13, ZGTL13]. Acanthopagrus [BBOE10]. Acartia [LTQ14, MAB12]. accelerator [KHF14]. accelerometer [SLS15]. accepted [DG11]. access [WJM16, YS15]. accessibility [WJM16]. accessible [LSJ10]. accompanying [PCBO^{+}18]. according [PKK13]. account [PSGY^{+}12].

Accounting [BN16, KPL^{+}19, Cad13, DDR10, LSCL18, OIMP18, RSV^{+}17, RTS19]. accounts [DBA^{+}19]. accumulation [COS11]. Accuracy [HJT12, SJUE11, BSF^{+}19, McB15, MBA^{+}19, MBP19]. achieve [WW16b]. Achieving [UVD^{+}17, Por11]. acicula [RRM^{+}12]. Acid [LLLB^{+}10, LHK18, PFPG12, YPN^{+}18]. acidic [ELBS17]. acidification [BMM^{+}19, Bro16, Bro17b, BHM13, CGAD16, CRC^{+}16, CH16, CPLH16, DCS^{+}11, DDF^{+}12, DRVV^{+}17, FAP^{+}17, GQX^{+}19, HSS^{+}16, Hum17, HFM13, Jok16, LBG18, LFH^{+}19, LVSF17, LPSF19, LAD^{+}17, Mac17b, MGL^{+}15, McE17, NF16, NMF^{+}16, PSH13, PFD^{+}16, RGAS^{+}18, RAH^{+}16, SAB^{+}16, SFCS^{+}17, SD17, SFD^{+}16, VGB^{+}17, WZZ^{+}16, ZSC16]. acidified [GRF^{+}16]. acidifying [CSJ16a]. acids [BDF^{+}14, GSMRM^{+}16, ÖPPM19]. Acoupa [FBL^{+}16, FBL^{+}16]. Acoustic [BMC^{+}13, DKK15, HRC^{+}13, MMG^{+}17, STB^{+}11, BHD13, BDHC19, CCD^{+}19, CJP^{+}14, CD14, Dav11, DW11, DZC^{+}13, DMT10b, DRB10, FHOK14, GCBI17, GHB^{+}10, JLL17, JVPM12, KNH11, KPJC14, KLN^{+}10, KRKG16, Kor10, KHPI15, LCM10, LCS^{+}15, MDS13, OKDJ18, RK16, SG17, SOK^{+}12, SLS15, SR16, TMS^{+}15, TBC15, TGDG15, VBA10, WJM16, WCK^{+}10, WWI^{+}16, XMM12, ZDD^{+}19, vdKKS11, vdKFS^{+}16]. acoustic- [JVPM12]. acoustic-backscatter [KLN^{+}10]. acoustic-optical [RK16]. acoustic-trawl [DZC^{+}13, WWI^{+}16]. acoustical [WLL^{+}13]. acoustics [FCG^{+}16, GHB^{+}14, KBB13, LAG^{+}16, QFM^{+}10, SWvSR11, TOA^{+}16]. Acropora [BIKP^{+}17, dCVB14]. across [HBB^{+}16, KDF^{+}19a, LMG^{+}14, MRP^{+}16, MFB17, PWCS17, PSO^{+}14, RSB^{+}15, RTAT17, RST18, SHK^{+}15, SBNL12, SCY^{+}10, VCH16]. Act [HMN^{+}12]. action [HCR13]. Active [Bak14, HUPBL18, LAG^{+}16, PPHK19, SKA15, SKA22]. activities [CMC^{+}12, GBJ^{+}15, KJW^{+}18, KPJ^{+}15, LJH^{+}12, MSR^{+}19, MMAS11, TTG^{+}15]. activity [BLD17, BVC15, DI11, GDS15, KKLM13, LLLB^{+}10, MRMK11, PE18, RCF^{+}18, SLS15, SJ16]. aculeatus [OJK^{+}19]. Acute [TWB^{+}19, AS14]. acutus [GIPS12]. adapt [SH14b, DDR10]. Adaptation [Cri12, MS12, YFP^{+}19, COL^{+}13, FB19, VP19]. adaptations [BCL^{+}11, DL12]. Adapting [TWHB19]. Adaptive [PTvOR13, AHK^{+}18, CAC^{+}18, DC14, GVA19, GMKS11, OHD^{+}16, PEN^{+}19b]. Additional [FCG^{+}16]. additive [LPT10]. addressing [SWRC^{+}18]. adequacy [PWS^{+}11]. adjacent [NC12, OAM^{+}10]. Adriatic [BCC^{+}14, BPPP17, CMU^{+}17, PDZ^{+}13, ŠBGT^{+}17]. adult [ALWH18, CMU^{+}17, HAF^{+}16, HHM^{+}16, MTP^{+}12, MVB^{+}18, ØU13, REFJ12, SSF^{+}18a, SCC^{+}13, SBSE14, UFJ^{+}18]. Advanced [LRJ18]. Advances [HHRW12, HBG^{+}19a, HBG^{+}19b, Paf18]. Advancing [HKK^{+}17, SMOH18]. advantages [SMMK11]. advection [WAH^{+}16]. adverse [FWRM12, KCMS19]. Advice [DAP^{+}12, BCRM^{+}18, CC11, GKJ^{+}15, Jen13, KPL^{+}19, KHMS19, LPFH13, MLB^{+}14, Mor18, PDD^{+}10b, SPG^{+}17, SH14a, URV^{+}11, vGA18a]. advisory [LKS14a]. Advocacy [Ric11]. Aegean [MKD^{+}18]. aeglefinus [BGM^{+}14, CPVS14, LSY^{+}14]. Aerial [CSR11, BVC15, CBK18, LDT^{+}11]. Aerosol [MSL^{+}11]. aff [HMHL19]. affect [GÖN^{+}12, HHT13, LKS^{+}14b, LBK17, MIM^{+}19, OBJA^{+}18, SLF17, TNH^{+}10, VOM11]. affected [DSA13, HWR^{+}16, ZSC16]. affecting [BBB^{+}19, CAB^{+}10, HBB^{+}16, HvDH^{+}10, HAH^{+}16, KHPI15, LYW^{+}18, SDH^{+}18, THKP11]. affects [KPGH16, KFL^{+}15, MRC^{+}18a, MRC^{+}18b, NF16]. Africa [APK11, CST^{+}19, MRC^{+}10, dGGW^{+}11, vdHBSM10]. African [GKC^{+}15, RDS12, RBP15, dMBD11]. aft [URE^{+}18]. after [ABJ^{+}19, BF17, BBHC11, CMH19, CFHC10, JDFN12, KKC19, Man17a, MCSL16, MBBC11, SMBH16, SWBJ13, SBS^{+}19c, VQGDAMdL14]. Ag [LLLB^{+}10]. against [COL^{+}13, GMM^{+}12, PMOH13]. Agardy [Aga18a, Day18a]. Age [Ada17, AEC11, Bru10, DWPS11, FMC10, GMH^{+}14, GOPG10, GFML10, HRF^{+}10, HCL^{+}18, HCF^{+}12a, JÓT^{+}12, KMF13, KDBP11, MHK19, Pow14, WNM10, AST^{+}15a, ALWH19, AH19, BMR^{+}11, BGM^{+}11, BTA^{+}18, BN16, BHS^{+}11b, BMR10a, BPC10, BP13, CKH19, DGC12, FFH^{+}11, FHH^{+}13a, FHH^{+}13b, HLCC16, HJS14, HAH^{+}16, HHQ11, HHQ12, HFSV^{+}15, Hüs10, HCF^{+}12b, HRP^{+}16, HGH^{+}16, JHB^{+}12, JSOO12, JOSO16, KAT11, LKR16, LCSI17, LB11, MTP^{+}12, McB15, Nel19, OYIN18, OSJ^{+}16, OLM^{+}15, PDAC^{+}16, PHM13, SHW^{+}15, SLH^{+}17, SBH^{+}14, SWA^{+}16, WMN^{+}11, WNW^{+}15, WMJ13, WWI^{+}16, ZGT13]. age-0 [HAH^{+}16, LKR16, PDAC^{+}16, WMJ13]. Age-based [WNM10, BN16]. age-dependent [MTP^{+}12]. Age-specific [Ada17, Pow14]. Age-structure-dependent [Bru10]. age-structured [BPC10, DGC12, HHQ12, HFSV^{+}15, OYIN18, OLM^{+}15, SBH^{+}14]. Ageing [BA10, WOW^{+}17, ALWH19, GBM^{+}19, HLCC16]. agent [TSP12]. agent-based [TSP12]. ages [SLM^{+}11]. agglomerative [SHS11]. aggregate [DPL10]. aggregated [WWCO15, ZKD^{+}14]. Aggregating [MKB^{+}17, Bjö18, BMS^{+}18, BMO^{+}19, CAL14, FS14, SO16]. aggregation [BS13, GRH^{+}14, HFGT18, HAR19, MPB11, TCBD10, vDHvKR15]. Aggregations [FDJ18, FDJ19, CWRB11, GKR14, GR15, LDD^{+}10, LBTS^{+}19, OdJN^{+}12, RB10, VPH18]. Aggressive [BFDGLÁ15]. agreement [FCB17, McB15, SPS^{+}10, WLL^{+}13]. Agulhas [MRC^{+}10]. AHD [TGDG15]. ahead [DCPTN14]. aim [PHV^{+}15]. air [PHO13]. Airborne [KGP^{+}15]. aircraft [CBK18]. Airgun [FHD^{+}19]. AIS [SHS^{+}18]. Al [GGT^{+}14, ABJ^{+}19, CMP^{+}11, CSJ16b, FWP^{+}16b, FWC^{+}19b, HKB^{+}18, HLL18, HR14, HPCW19, MAP^{+}11]. Al-Abdulrazzak [GGT^{+}14]. al. [AAP14a, CMH19, CN14, HS10, RPM14b, RR10]. alalunga [CAB^{+}10, WKK^{+}15, WLN^{+}13]. Alaska [BDW18, Fau11, IS15, LKR16, OSL^{+}15, OIMP18, PMH^{+}13, PH17, PSGY^{+}12, RWG^{+}18, SZM^{+}10, SRRZ16, SMK15, SOL^{+}15, WMJ13, WSW^{+}19, ZHAG17, ZGK^{+}17]. Alaskan [CBL19, Cri12, FB11, SOL^{+}15]. albacares [LLL^{+}11, MMSS18, WLN^{+}13]. albacore [CAB^{+}10, PCB^{+}14, WKK^{+}15, WLN^{+}13]. albatross [MRP^{+}16]. albatrosses [RB10]. albida [HLS^{+}15, MMRG18, VBW^{+}18]. Alcyonacea [CSBHS^{+}11]. Aleutian [DMO10, RWG^{+}18]. Alewife [TMR^{+}16]. Alexandrium [MMH17]. alga [GQX^{+}19]. algae [BGM^{+}19, CCE17]. algal [DKPM16, MLMS15, TTJ^{+}18, TTJ^{+}19]. Algarve [SPE14]. algorithm [DCG10, DGC12]. algorithms [SHS11, Sub18a, Sub18b]. aliased [RD16]. alien [FPRA11]. Alitta [SCD^{+}15]. alive [WS13]. alkalization [LFH^{+}19]. Allee [GDHFS^{+}18, Hut14, RLD^{+}12]. Allelic [NNY17]. allis [RLD^{+}12, RDB^{+}18a]. allocation [CSFS10, CK18, HLD10, KOK^{+}14, LHV^{+}16, MLT11, PBQ^{+}10, STM19]. Allometric [BI13, SV10]. allometry [OG16]. allowable [SRGF15]. Allowing [HFHS10]. allows [MTO^{+}18]. allozyme [GMKS11]. Almejas [NdlPIR18]. along [ALd12, BPPP17, FGMC^{+}13, FC19, FSC^{+}12, GHD^{+}10, HTRF17, HGRAR10, HFSB^{+}14, JDFN12, KMF13, KBT14, KdS14, LGB^{+}18, MSL^{+}11, MSS^{+}19, MDMC11, NC12, ØFNH12, RSBC13, TIS^{+}19, WKW^{+}10, WKL13]. Alosa [RLD^{+}12]. alosine [BSC14]. alter [DFB^{+}10, HSS^{+}16, SiTiM^{+}11, SSA^{+}18]. altered [ALMFFBP19]. altering [LGB^{+}18]. alternate [Pep16, TDR^{+}16]. alternative [CECL16, CSF^{+}17, HABV19, HMP^{+}15, LJH^{+}12, LFGW13, LBB15, MHHT10, MPB11, PO14, RCH^{+}15, WPeEA16, vH10]. alters [SLF16b]. always [DW11, GR15]. ambivalence [Aga18c]. America [KBT14, ME11, SRCR12, dGGW^{+}11]. American [BM16, HdBR^{+}16, LCG^{+}18, QR15, STR16, TCS^{+}19, THSM^{+}18, VOS10, WDOJ15, WWMF17, WABZ16, XS12, ZCW11, dlTQMUE10]. americana [BHDB12]. americanus [THSM^{+}18, WWMF17, ZCW11]. Ammodytes [KO12, WCR^{+}19]. Among [PSS11, FCB17, GSL^{+}11, MJS^{+}15b, MJAS14, PQR10, SBG10, SJRB18, SOL^{+}15, WOW^{+}17]. Among-stock [PSS11]. amongst [BHM^{+}18, CK18]. Amphibalanus [PSH13]. Amphipoda [PCFR13]. amphitrite [CYLT16]. amplification [JFJ^{+}17]. Amplifier [DBA^{+}19]. Amsterdam [GRG^{+}10]. anadromous [PBW15]. analogues [RGAS^{+}18]. analyse [AHC15, MRP^{+}16]. analyses [APHC15, BWR^{+}15, GKJ^{+}15, HvDH^{+}10, PBW15, PA18, SPE14, YPN^{+}18]. Analysing [CLMP16, Pay10]. Analysis [ALd12, Loh11, MIJ^{+}17, PDS15a, AB14, APB^{+}10, ABJ^{+}19, ASB^{+}19, BJH^{+}14, BF17, BPKH13, BCT^{+}10, BNB^{+}14, BHS^{+}16, Bru10, BYQ^{+}19, CSS^{+}12, CMH19, CMLVGLP14, CMG^{+}17, CAB^{+}10, CAOG15, DBDP10, DMT^{+}10a, DWPS11, ETG^{+}15, EMW^{+}16, FPSF11, FA10, GRC16a, GFMO11, GSMRM^{+}16, HA15, HGF^{+}18, HRF^{+}10, HHZ^{+}18, HBF14, IEL^{+}15, JHB^{+}12, JQD^{+}17, JvdM15, KNH11, KKC19, KAT11, LJH^{+}12, LMPP10, LGR^{+}19, MIM^{+}19, Man16, Man17a, MNV^{+}11, MYI12, May14, MSE12, MAS19, MMSS18, OYI17, OCG10, OHD^{+}16, PWCS17, PFPG12, PE18, PSF12, QG12, RHB^{+}12, RHB^{+}13, RHHH^{+}12, SSP^{+}13, SBS^{+}19c, TVJ^{+}14, TSPL14, VPSV^{+}10, VSP^{+}14, WZZ^{+}16, WW16b, Win15, XMM12, ZMF12]. Analytical [BPC10, Ham15]. Anarhichas [PSDG12]. anchialine [CAMM12]. anchoita [MMG^{+}17]. anchor [NMM^{+}17]. anchovy [AIA^{+}15, AFQ^{+}11, BJH^{+}14, BCBI13, BPPB12, BMC^{+}13, CAC^{+}18, DBL^{+}16, IFU11, LMPP10, LLPV^{+}12, MMG^{+}17, MHHT10, MOVU10, PDMG11, PDH^{+}14, RBAB17, RML^{+}16, TID^{+}10, VSP^{+}14, dMBD11]. ancient [CGM^{+}19]. Andersen [FWP^{+}16b]. angle [CD14, FCG^{+}16, FA10, KO12, TKA^{+}16]. angled [BBH^{+}10, BBHC11, RBBC11]. angled-and-released [BBH^{+}10, RBBC11]. anglerfish [CSS^{+}12]. angling [BCC^{+}19, FCG^{+}16, LSF^{+}18, ODSC10, PVCB17]. Angola [VBA10]. Anguilla [ACM^{+}16, AFP12, BM16, BLB^{+}18, CMC^{+}12, DHKE16, HdBR^{+}16, JTM^{+}16, LØM^{+}18, MM12, MMWC16, PR16, PMOH13, SSP12]. Anguillicoloides [HdBR^{+}16]. anguillid [CD16, MFT16]. animal [CJC^{+}15, RIBB^{+}19]. animals [BCC^{+}19, FHH^{+}13a, FHH^{+}13b]. Annual [TCH^{+}16, BJR17, Ben13, BTTV^{+}17, BG18, DMO10, SHS15, VGWJ11, WLN^{+}13, ZGT13]. annual- [WLN^{+}13]. annually [GUN^{+}19]. anomalies [Kam14]. Antarctic [BS13, BGM^{+}19, CK18, CWRB11, FWT^{+}14, MDD^{+}14, RB10, RT19, SSM^{+}16, STF^{+}16]. Antarctica [PBG14, LTT^{+}18, TRHB13]. Anthropocene [CSB18, RIBB^{+}19]. anthropogenic [MFT16, PPS17]. anticyclonic [KYK^{+}12]. antillarum [RBMAOCL18]. aper [ECW^{+}17, FHC^{+}12, FOJ13, HCF^{+}12a, HCF^{+}12b, WMN^{+}11]. aperture [KGP^{+}15]. Aphanopus [MMM11]. Apocalypse [DGPR11]. Apparent [RBMAOCL18, SWS13]. appear [Cla18d]. Application [GW16, ALBR14, BSW^{+}11a, BSW^{+}11b, BMS^{+}19, BPC10, BMC^{+}17, CLdR^{+}15, DCG10, DBL^{+}16, EPH15, FCPJ10, Fle15, FWD15, GVA19, GLC15, HRB15a, HPN^{+}17, IFU11, Lap11, LTGP18, LSR^{+}13, MGPC^{+}14, ME11, PDMG11, PKR^{+}19, PSGY^{+}12, QG12, RFMR10, SOB^{+}11, SPS^{+}10, SR16, TSP12, THC16]. applications [HKS17, KM16, LSD^{+}11, MKHK18, NDB^{+}15]. applied [Bru10, DUM^{+}12, MLT11, Nel19, SBCF16, UBvH11]. apply [HHA^{+}11]. Applying [Bro16, LLSJ18, vH13]. approach [Abl16, ACCA19, AH15, BCRM^{+}18, BOP^{+}19, BPMR17, BLJ^{+}17, CMM^{+}15, CPVS14, CLS^{+}18, CBR^{+}11, CCL^{+}14, DC14, FWP^{+}16a, FWC^{+}18, FWC^{+}19a, FWC^{+}19b, GLH14, GLK^{+}10, GGP11, GBJ^{+}15, GD15, HS10, HP17, HHA^{+}11, HSFiS12, HPCW19, JLH12, JOSO16, JMM^{+}15b, JLS^{+}15, KJW^{+}18, KVA18, KPJ^{+}15, KRG^{+}12, LHS^{+}18, LGL^{+}12, LKS14a, LLCJ10, LH19, Llo17, LLSJ18, LBTS^{+}19, MPR12, MLT11, MNV^{+}11, MLS^{+}17, Mat11, NSO^{+}19, OD17, ÖHSNV17, OSSL16, PDZ^{+}13, PDD^{+}11, PSS11, PMBM18, RCF^{+}18, RBBC11, RJP^{+}17, RGPG10, ŠBGT^{+}17, SGV18, SRDC^{+}14, SRB^{+}14, SMMK11, SWRC^{+}18, TWHB19, TCS^{+}19, TLF16, UVD^{+}17, UFLH^{+}13, WGCL^{+}19, WFS^{+}11, YDT14, ZMKC14, ZHLK11, dMBD11]. approaches [Ben15, BBD^{+}13, GF14, HJS^{+}17, KHC^{+}17, LBLJ17, LRJ18, LBB15, PFH16, RMF^{+}15, RD14, SML14, SFG^{+}15]. appropriate [LCMR^{+}18]. approximation [SRW13]. Aquaculture [RSY11, ATLC16, BJH^{+}16, BCC^{+}19, FRM^{+}16, GKC^{+}16, HSV^{+}17, Ham14, JQD^{+}17, LHJ15, LSR^{+}13, PCBO^{+}18, RMF^{+}15, SMR^{+}11]. aquarium [BCL^{+}11, CBC^{+}11b, Pin17, STD11]. aquatic [BCC^{+}19, HBB^{+}18, HKB^{+}18, HLL18, HHT13]. Arabian [BFB^{+}11, LGR^{+}19, MDP^{+}15, MSL^{+}11, RLQ^{+}15]. aragonite [RRM^{+}12]. Arbacia [VGB^{+}17]. archipelago [AV15, MB14, MRMK11]. architecture [Sca18]. Architeuthis [GRNG^{+}10]. archival [AGG^{+}18, CAM^{+}15, CAB^{+}10, HLA^{+}11, LFM^{+}18, Loh11, STC^{+}17]. Arctic [BTB^{+}17, BMFSH12, BA12, CPVS14, CBBM14, CSF^{+}17, DMT^{+}10a, DMS18, EKR^{+}19, GDBM^{+}17, HMLKR18, HAH^{+}14, HSK^{+}14, KT18, KOK^{+}14, KLCC18, KCL18, LHJ12, LSY^{+}14, LCSI17, LCSI18, MDD^{+}14, MBA^{+}19, PCFR13, RDB^{+}18b, SHG^{+}18, SAS^{+}18, VDK15, WPGW12, WSW^{+}19, Yar10]. Arctica [TJR^{+}10]. Area [DOB^{+}17, AV15, AIA13, BGM^{+}14, BLJ^{+}17, CBW15, DMO10, FVSL^{+}10, FMM^{+}12, FDC^{+}16, GKC^{+}16, HWK^{+}15, HHHE16, JDH^{+}14, LHHK12, LNWS18, MS12, MBLS15, MR14a, MJS^{+}16, OCS14, OAM^{+}10, SWS13, TK18, THR16, VHH10, VRF^{+}16, YS15, dJBLH15, LLM12]. Areas [Ano18p, AV15, AFP12, BHC^{+}15, Ber18b, BSP15, CMA^{+}16, CDAN^{+}14, Fri19, GGTW17, GRH^{+}14, JKvdK^{+}15, JRWM11, KKC19, LMPP10, MHM^{+}11, Mos18, NS17, OE16, PMP^{+}16, PHM^{+}14, PAB^{+}18, PDAC^{+}16, PKHG14, RYS11, SG18b, Sca18, SDSA15, SSRT10, SG16a, SJUE11, SJRB18, SMS^{+}17, STB^{+}11]. Argentina [BHDB12, DAB^{+}12, MDPH12]. Argentine [MMG^{+}17]. argentinus [CAOG15]. Argonauta [RS10]. argonauts [RS10]. Argopecten [RFN^{+}19]. argus [BMG15, CHM15, GBM^{+}19, GBB15]. Arkona [HHE^{+}16]. armoured [HGS^{+}19]. arrangement [TMW19]. array [GZS^{+}19]. arrival [BWR^{+}15]. art [Sca18]. artefacts [DCHN^{+}15, PA18]. articles [Ano18p, Sei14]. Artificial [HUPBL18, BTL17, GS14, GS15, KNH11, LGB^{+}18, PS13, SLT^{+}14, SLM^{+}11, SPS15, SLT^{+}17]. artificial-reef [KNH11]. artisanal [BMSC16, DMT^{+}10a, FVSL^{+}10, KKAV^{+}19, OCG10, OCWG16, PG10, PLCC18, dlTQMUE10]. Asia [dGGW^{+}11]. Asian [HZZ^{+}15]. aspects [ESD^{+}16]. assemblage [AIA13, FDC^{+}16, GS15, MST^{+}18, MJS^{+}15b, PCBO^{+}18, PGN^{+}11]. assemblages [BFH13, CD11, EEP^{+}11, HGS^{+}19, HAF^{+}16, JDM^{+}16, KTMV16, MCO^{+}17, NC12, NPGT15, NGTT16, RFHG19, RTAT17, SHS11, tHR11, vdKKS11]. assess [BNK10, BBOE10, EPH15, GG13, GRC^{+}14, GHG^{+}10, GJWS19, JVPM12, KGP^{+}15, MIJ^{+}17, PSF12, RMJF14, RBBC11, SHS15, SWBJ13]. assessed [KAT11]. Assessing [ABRL12, APHC15, FWRM12, GSMRM^{+}16, HABV19, JRG^{+}16, JLH12, KJW^{+}18, LHS^{+}19, LBTS^{+}19, NSS15, PRBF18, PEN^{+}19b, PWS^{+}11, RBM15, SRDC^{+}14, SMK15, SSTB15, TDR^{+}16, TLPS15, ZCR15, ASB^{+}19, CJP^{+}14, DID^{+}16, DSA13, DWPS11, HMN^{+}12, HP17, HBiI^{+}11, JHB^{+}17, MRP^{+}16, MBP19, SOB^{+}11, SR16, ZHLK11, ZDF^{+}19]. Assessment [LJB16, LAD^{+}17, POIM12, Par13, TCD^{+}16, ANT18, ACCA19, APOG11, ABRL12, AH15, Ast15b, BN16, CDC15, CTC14, CAM^{+}15, CECL16, CR11, CIJ18, CBS15, CGGH^{+}17, DG11, DDR10, DED13, DBM^{+}15, DCVC15, DK19, EHB^{+}15, EHD^{+}15, FCPJ10, FCB17, Fle15, FPLB19, GCJL16, GPP17, GB15c, GLC15, HPDB19, HKS17, HLCC16, HHQ12, HCE^{+}11, HFSV^{+}15, JMM^{+}15a, JOSO16, JMM^{+}15b, KDH^{+}10, KDF^{+}19a, LGL^{+}12, LTGP18, LBB15, LWN^{+}13, LB14, LB11, LDT^{+}11, MWP^{+}15, MWJ11, MGVH^{+}10, MGPC^{+}14, MP15, MG15, MRD19, MJS^{+}15a, MBA^{+}19, MLB^{+}14, Nee15, OJK^{+}19, PHV^{+}15, PKK13, PDD^{+}10b, PHM13, RGRL11, RBB^{+}16, RMKG11, RTS19, SSS^{+}12, ŠBGT^{+}17, SDDA11, SH14a, SSF^{+}12, SM15, SSZH12, Sub18a, Sub18b, SC16, SIP18, TBG16, TCS^{+}19, TKB^{+}15, TRHB13, Tho11, Tho14, TJH15, Tow14, TFC^{+}19a, TFC^{+}19b]. assessment [TPS^{+}11, TCF^{+}13, VCF^{+}12, WW16b, WNW^{+}15, ZMF12, ZPS^{+}18]. assessments [ASV^{+}19, BPD^{+}18a, BPD^{+}18b, BG18, Cad14, DGL^{+}17, DC14, GB15a, GKJ^{+}15, HMR15, JTE19, KET^{+}17, LKS^{+}14c, MKL^{+}19, MBT^{+}19, MBP19, PA18, Pow14, PSS11, SHL^{+}14, SFG^{+}15, WM14, XPC11, ZHAG17, ZE17]. assign [BHL^{+}15]. assignment [CMG^{+}17, ETG^{+}15, FHH^{+}13a, FHH^{+}13b, SLKS10]. assignments [SLKS10]. Assimilating [BGA^{+}19, WSGD^{+}17]. assistance [DVV^{+}11]. associated [BTL17, BHL^{+}15, BMS^{+}18, BMO^{+}19, CBD^{+}10a, CMP^{+}11, CSBHS^{+}11, DKC12, DD10, FTM^{+}17, FS14, GCJL16, GvRKB19, Ham14, HHM^{+}19, LLL^{+}11, LWS^{+}18, MAP^{+}11, MVP^{+}14, PH17, RLW^{+}15, SLM^{+}11, SG16b, SWA^{+}16]. association [BHCSP^{+}11, TMR^{+}16, WMJ13]. associations [FDJ18, FDJ19, GGV^{+}18, PCB^{+}14]. assume [GBB16]. Assumptions [ABF^{+}16, FWP^{+}16b, Jok16, SCL^{+}19]. assurance [HBF14]. asterias [FMC10]. Asturias [RGGFA16]. asymmetry [DGPC^{+}15, MIM^{+}19]. asymptotic [HTS^{+}10, HHT13]. asynchrony [JOSO16]. at-sea [BSC14]. Atlantic [AWS^{+}13, BMH^{+}16, COM11, CAM^{+}15, DK19, FMC10, HLS^{+}15, MDP^{+}15, MMAS11, PDS15a, PDS15b, PDCP14, PHB^{+}16, PGN^{+}11, RHB^{+}13, SRHJ12, VSP^{+}14, ALWH19, AV15, AIA13, ABHT^{+}16, AGG^{+}18, BJH^{+}14, BGM^{+}11, BJH^{+}16, BR12, BKMdMS13, BHL^{+}15, BMFSH12, BSC14, BCS^{+}17, BFC^{+}16, BKT12, Bow14, BPF19c, BHS^{+}16, BGA^{+}19, BFH^{+}18, BT10, Bru10, BDHC19, BMH^{+}16, BYQ^{+}19, BCOS11, BFH13, CNBK11, CSS^{+}12, CTC14, CMH^{+}16, CMA^{+}16, COS11, CMH^{+}11, CSC^{+}13, CB12, Cha12, CCD^{+}19, CDSP11, CCC^{+}14, CKD^{+}17, CCL^{+}14, CHMY15, CCSG18, COHdP19, CAB^{+}10, DMZC17, DED13, DHZA14, DEH^{+}19, DBDP10, Der18, DFS^{+}19, DPD^{+}12, DK18, Elv15, EGC^{+}17, EPH15, FHC^{+}12, FGMC^{+}13, FWH^{+}15, FFK^{+}11, FBM^{+}18, FHK14, FBF^{+}17, FAR15, FLP^{+}13, FPLB19, FMM^{+}12, GMN^{+}17, GMS^{+}18, GMD^{+}11, GCS^{+}18, GCW^{+}18, GDJ11a, GSL^{+}11, GUN^{+}19, GDBM^{+}17]. Atlantic [GSMA15, GLP^{+}11, GIW16, GKJ^{+}15, GIR^{+}19, GCS^{+}17, HSSB14, Han13, HG10, HLL^{+}12, HUT^{+}11, HdBR^{+}16, HPS^{+}11, HRC^{+}13, HWK^{+}15, íHHJ^{+}13, HFH10, HAH^{+}16, HP11, HCF^{+}12a, HCF^{+}12b, JHB^{+}12, JKvdK^{+}15, JÓT^{+}12, JWS^{+}18, JOvdM^{+}14, Kam14, KHH^{+}17a, KMF13, KDF^{+}16, KUM^{+}12, KDF^{+}19a, KCK14, KLC13, KWF11, KdS14, Kor10, KR12, KN17, KRO^{+}16, LLL^{+}11, Lap11, LHHK12, LLL^{+}13, LFGW13, LFM^{+}18, LSWD12, LJB16, LYW^{+}18, LBK^{+}19, Llo17, LOOO16, LSCL18, MMM11, MWS^{+}17, MAA^{+}19a, MAA^{+}19b, MVB^{+}18, MSOM10, MALM17, MGPC^{+}14, MKB^{+}17, MCSL16, MBC11, MPJ12, MSK^{+}11, Mel16, ML12, MJSB14, MKC^{+}11, MSR^{+}12, MMDSP15, MVP^{+}14, MSR14, MGH^{+}12, MBL^{+}17, NSS15, NSO^{+}19, NGTT16, NUÓ^{+}16, OSJ^{+}16, OEG^{+}16, OBLP^{+}19, OEK^{+}18, OAT^{+}12, OFYS^{+}11, OBL12, ÖPPM19, OVW^{+}17, PSDG12, PSO^{+}14, PHM^{+}14, PRKASR12, PKHG14, PKR^{+}19, QG12]. Atlantic [RSBC13, RHB^{+}12, RBA^{+}19, RC12, RHOJ10, RPDF14, RJ12, RFHS14, RFT^{+}16, RAB^{+}12, RCH^{+}19, SHK^{+}15, SRC11, SOK^{+}12, SSRT10, SLS18, SGM^{+}17, SZOL19, SMNE14, SBNL12, SFNdH16, SPM^{+}18, SLKS10, SRCR12, SHW^{+}15, SPS15, SPS^{+}10, SCS^{+}11, SBL^{+}17, SWA^{+}16, SKA^{+}12, SS10b, SHAM10, SHS15, SMK^{+}11, SHST18, SCC^{+}13, SBD^{+}15, SO10, SBKA16, STC^{+}17, SFN^{+}19, TCBD10, TKB^{+}15, TMR^{+}17, TFM^{+}12, TMP12, THC16, TB16, TPS^{+}11, TMR^{+}16, UBV^{+}11, UFJ^{+}18, VF17, VCF^{+}12, VDK^{+}18, WPGW12, WS13, WSG^{+}17, WMN^{+}11, WÓG^{+}16, WMG^{+}17, WDG^{+}17, WRC^{+}12, WVG^{+}13, WRDF10, WMG11, ZMKC14, ZHD^{+}14, ZLC^{+}17, ZDD^{+}19, dHFF^{+}16, vdKFS^{+}16]. Atlantis [SFD15]. Atlanto [BSW^{+}11a, BSW^{+}11b]. atlas [Ott10]. atolls [GVA19]. attain [RKJ14]. attenuation [RDKK15]. attitudes [GAB^{+}14]. attracting [HUPBL18]. attributes [ACS15, GS15, MFS^{+}14, PS13]. attribution [vPFF^{+}16]. audax [KDBP11, SSP^{+}13]. audit [SMMK11]. aurata [FSC^{+}12, GSMRM^{+}16]. auratus [WPH^{+}15, WPH^{+}17]. Australia [KS17, BTL17, CGG^{+}11, Ful11, GMH^{+}14, GD15, HD11, HBB^{+}13a, JRG^{+}16, Lav15, LMG^{+}14, LWT^{+}11, MWJ11, MTCF^{+}17, NPB^{+}15, OTM^{+}16, PDD^{+}10b, RHP^{+}12, SSH^{+}14, SFA^{+}17, TMG16, WNM10, WWCO15, ZCH19, dL14, dLCF^{+}15]. Australian [MFS^{+}14, AFK15, BBHC11, DPD^{+}15, HMHL19, MBBC11, OCBJ10, PPD17, PCdL15, PCS^{+}11, TBD^{+}18]. australis [BBH^{+}10, BBOE10, OOD13, TDN^{+}19]. autocorrelation [SBG10]. autodiametric [DPRG^{+}18]. Automatic [MBAGVG^{+}18, SSM^{+}18, SHS^{+}17]. automatically [VPH18]. automating [SHS^{+}17]. autonomous [SKH^{+}12]. Autumn [LF14, MO18, OTAK15]. autumn-spawning [MO18, OTAK15]. AUV [GCBI17, SKH^{+}12]. AUV-based [GCBI17]. Auxiliary [Tho11]. availability [CST^{+}19, KLCC18, KCL18, KHPI15, LLPV^{+}12, MOVU10, SBSE14]. available [HBB^{+}13a, VKS^{+}12, XCRP13]. Aveiro [LTQ14]. averaging [MJS^{+}15a]. avoid [Cam18, CZBFLÁ^{+}15, DMO10, Mac12, Moo19]. Avoidance [SOL^{+}15, DH13, OSSL16, Peñ19a, PDR19, SJ16, WLL^{+}13]. Avoiding [Bar19, RD16]. away [LFMJ19a, LFMJ19b]. Azorean [CJP^{+}14]. Azores [BHCSP^{+}11, MMM11, PCD^{+}13, PDCP14]. azoricus [BCL^{+}11, CBC^{+}11b]. Azov [VG15].

back [DB16, HMS^{+}13, JMM19, Omm18]. backed [TSBV^{+}15]. background [MWS^{+}17]. backscatter [Dav11, DW11, DLSJ^{+}19, KLN^{+}10, MBQ10, SG17]. backscattering [CD14, DKK15, FHOK14, JLL17, TK18]. backtracking [BGS^{+}14]. bacterial [HTRF17]. Bacterioplankton [WZZ^{+}16, DRVV^{+}17]. bad [BSR^{+}10, Mur10]. Bahamas [ODSC10]. Bahía [MDPH12, NdlPIR18]. bairdi [HCR13, PFD^{+}16]. bait [HBHR18, HUPBL18]. baited [BLB^{+}12, COR15, HUPBL18]. Baja [dÁMNGS^{+}11]. balance [GSC10]. Balanced [CGR16, GRC16a, HHBSM16, KGZH16, Pla17, BCN16, DK19, FPLB19, FWP^{+}16a, LPK12, RGS^{+}16]. Balancing [EHT17, WLB11, GRC16b, WHMP15]. Balanus [CYLT16, EWH16]. Balearic [MNGA^{+}17]. baleen [LHJ12]. ballan [DMF13]. Ballast [DHNL12, CBBM14]. Baltic [CTH^{+}19a, CTH^{+}19b, HKB^{+}18, HLL18, HSHÖ14, MO18, MHK19, AL10a, BMR^{+}11, BNK10, BNE^{+}15, BGS^{+}14, BBD^{+}13, CECL16, DLL12, Eer12, EHHH14, EHB^{+}15, EKR^{+}19, FCB17, FOC^{+}18, FDR^{+}17, GNB^{+}16, GOS^{+}13, HBB^{+}18, HMS^{+}13, HHM^{+}11, HHH12a, HPN^{+}17, HNAK12, HBG^{+}16, Hüs10, Hüs11, HHH12b, HRP^{+}16, HGH^{+}16, HHE^{+}16, HHHE16, JCLO17, JMM19, KUM^{+}12, KOT^{+}17, KFL^{+}15, KHH^{+}17b, KVFK19, LKHK11, LHK18, LSB^{+}10, LHLK10, MRD^{+}12, MGBvD14, MKC^{+}11, MLB^{+}14, MJS^{+}15b, NKSE14, OTAK15, OLR^{+}18, OND^{+}18, OBG12, OTO^{+}15, OJK^{+}19, OEK^{+}18, OFB^{+}17, OHD^{+}16, PVQS12, PHM^{+}14, PKHM12, PNC15, PKHG14, RHV14, RHHH^{+}12, RTJS14, RLP^{+}13, SSR^{+}16, SOGT^{+}19, SR17, Str10, SSZH12, SH15, SC16, TVJ^{+}14, VHS^{+}11, VHQ^{+}11, WW16b, WKP^{+}17]. balticus [Eer12]. ban [BUDM15, BCN16, CEP^{+}18, CCG14, PMP^{+}16, Pas14]. banana [ZBE^{+}15]. band [GBM^{+}19]. bandwidth [FHOK14]. Bangladesh [FKH^{+}16]. Bank [MRC^{+}10, BWP10, DMZC17, FLLG15, GF11, KHH^{+}17a, LF14, Mag11, TCS^{+}19]. Banks [Lev11, MMAS11, NPGT15, HBB^{+}13a]. barely [IHL19]. Barents [JIB^{+}14, ASDSM18, AMS^{+}12, BPF^{+}19a, BPF^{+}19b, BKT12, BMEBM^{+}16, CP15, DIS^{+}12, DJHO10, DD10, DD13, EISJ12, EDP14, ESD^{+}16, Eri16, FPRA11, GHM^{+}16, GDBM^{+}17, HBD^{+}19a, HBD^{+}19b, HB10, HF14, HHBSM16, IKF17, ITT19, JIB^{+}12, JLS^{+}15, JPTC16, NHL^{+}19, OAM^{+}10, OBY^{+}14, PJDN12, PT19, Rey12, SJUE11, SSF18b, SEDO19, SFN^{+}19]. barium [PDD^{+}10a]. barnacle [CYLT16, EWH16, JCSC10, PSH13, POIM12, Par13, RGGFA16]. barotrauma [FWH^{+}15]. barracuda [ODSC10, ODSC10]. barrel [EHT17]. Barrier [Day18c, HBB^{+}13a, PEV^{+}16, HZZ^{+}15]. bartramii [Che10, YCZY19]. base [HB19b]. based [ANU11, ACT^{+}18, ABF^{+}16, ABRL12, AHK^{+}18, ASB^{+}19, APHC15, BSHE18, BNK10, BUNB10, BRH^{+}15, BN16, BCL^{+}11, BVC15, BMR10b, CC17, CLT^{+}14, CK18, CKA^{+}17, CGGH^{+}17, CMD17, CSD16, DGC12, DHS19, DHZA14, DOB^{+}17, DCER^{+}14, DMS16, DPL16, Eid18, EBB^{+}16a, EBB^{+}16b, ES18, EP14, FGRR11, Fle15, FFH^{+}11, FHH^{+}13a, FHH^{+}13b, FWP^{+}16b, GNFM11, GLH14, GRC16b, GGP11, GNDC11, GCBI17, GWS17, GBJ^{+}15, GRF^{+}12, GOS^{+}13, GTBT18, GFF^{+}19, HKS17, HSHÖ14, HSFiS12, HOV^{+}15, HBHR18, HP11, HHT13, JHB^{+}12, KIA^{+}18, KS14, KVA18, KRG^{+}12, LÁG^{+}19, LSF^{+}18, LB14, LB17, LDCR^{+}19, LWT^{+}11, LBK^{+}19, Loh11, LDT^{+}11, LBCOJ19, MIM^{+}19, MKL^{+}19, MSMW18, MHM^{+}11, MBQ10, MDdPMQ10, MTLG14, MR14a, MRD19, MNWB18, MLB^{+}14, MUEO17, NPB^{+}15, NJFH19, ØHS10, OFB^{+}17, POIM12, Par13, PNE^{+}14, PCdL15, PVD14, PKR^{+}19, Pow14]. based [PKK13, RSY11, RCH^{+}15, RHB^{+}12, RHB^{+}13, RFF^{+}14, RDB^{+}18b, SHL^{+}14, Sas19, SCY^{+}10, SLKS10, SFW^{+}12, SMOH18, SKH^{+}12, SFA^{+}17, SR17, SPE14, SSP12, SWT^{+}19, Sub18a, Sub18b, SWRC^{+}18, SP13, TLR^{+}17, TCBD10, TCS^{+}19, TBD^{+}18, TLF16, TSP12, TED16, TSPL14, UBV^{+}11, VF17, VE13, VKN^{+}11, WNM10, WLL^{+}13, WFM^{+}15, ZGT13, dCVB14]. baseline [CP15, DHCE13, GCW^{+}18, JLS^{+}15, OVW^{+}17]. basic [BBD^{+}10, RR10]. Basin [CHMY15, KYK^{+}12, HHE^{+}16]. Basin-wide [CHMY15]. basis [EDBMB14, FPSF11]. Basque [ZRU^{+}15]. bass [BSB^{+}13, BvdMA^{+}18, KS18b, LSF^{+}18, PPS17, SHK17, SBCF16, SCC^{+}13, VWMS19, ZRU^{+}15, OTM^{+}16]. Basurto [Obu18a, Bas18a]. batch [GMS^{+}18, KS18b, dBWMD^{+}14]. bathymetric [CBD10b, VCH16]. bathymetry [CBD10b, EDBMB14]. Bathymodiolus [BCL^{+}11, CBC^{+}11b]. Bathynomus [TBG16]. bathypelagic [RGC^{+}19]. Bathyraja [AEC11]. baxteri [HD11]. Bay [ADB16, KS17, LHJ12, LSI^{+}15, DD13, AIA^{+}15, AFQ^{+}11, ASV^{+}19, ALB^{+}19, BI13, BCBI13, BMC^{+}13, BT10, CSR11, DKB14, FSN^{+}11, FFK^{+}11, FC19, FAR15, GSP12, HRC^{+}13, HAH^{+}16, IFU11, Jon14, KZS10, LAD^{+}19, LGL^{+}12, LPB^{+}14, LMPP10, LLPV^{+}12, Lor11, MTDP11, MOG^{+}14a, MOG^{+}14b, NDF^{+}15, PHGK^{+}16, PDMG11, PDH^{+}14, Pre19, PSGY^{+}12, SMSR10, TMW19, Tow14, UHB12, WWCO15, dPJGB13]. Bayesian [APB^{+}10, AFQ^{+}11, BHC^{+}15, HKS12, IFU11, JOSO16, JVPM12, LKHK11, MWP^{+}15, MGPC^{+}14, MS15, MBT^{+}19, NSO^{+}19, RDB^{+}18a, RFMR10, ŠBGT^{+}17, SPS^{+}10, SR16, TMP^{+}17, WFS^{+}11]. be [Abl16, ALBR14, BKTS11, Bjö18, DCHN^{+}15, EHG^{+}14, GNFM11, GR15, LGR^{+}14, MHM^{+}11, SPG^{+}17, SBS^{+}10, SMB^{+}16a, TMR^{+}16, XCRP13, ZE17]. beaked [CBD^{+}10a, CMP^{+}11, FCCLA10, MAP^{+}11, PJDN12]. beaks [GRNG^{+}10, XPC11]. Beam [BM12, CD14, DID^{+}16, EDBMB14, FA10, PTvOR13, QFM^{+}10, SHJI10, SLV16, THKP11, TK18, UTA^{+}16, VPO17]. beam-trawled [UTA^{+}16]. bearing [SMB^{+}18]. bears [Qui18]. Beaufort [GRKF16, FRF14, WSW^{+}19]. become [XCRP13]. becoming [CHM15]. bed [KKMS14, RD14]. beds [COL^{+}13, MQ11]. been [Bro14, Fra15, Ric19]. before [KKC19, MCSL16]. Beggiatoa [Ham14]. beginning [Pas14]. Behaviour [HUT^{+}11, AFH^{+}17, AUEC12, BPH16, BvdMA^{+}18, BFDGLÁ15, BLS^{+}18, BBOE10, CFM^{+}14, Cla18a, DHZA14, DRB10, EPVC14, FHD^{+}19, FMK10, GKV^{+}15, GSMA15, GDS15, GCL^{+}18, GTGD15, HWR^{+}16, HLA^{+}11, KTLB16, KRD^{+}18, LK17, LTT^{+}18, Mac12, MFP^{+}19, OSC^{+}17, OMTY14, OS14, PvdKE^{+}12, PST14, SRC11, SPS15, STD11, TTR^{+}19, VRF^{+}16, dHFF^{+}16]. Behavioural [Dil17, ACCA19, CTH^{+}19a, CTH^{+}19b, HHR17, JOF14, Loh11, MKHK18, RS10, SJ16, dPLF^{+}19, vPGFT13]. Behaviours [TMR^{+}17, CGAD16, LFGW13, PDR19]. behind [ABF^{+}16, FWP^{+}16b, OSP^{+}13, QG12]. being [Obu18c, Obu18d, PDT^{+}18]. Belgian [VVP^{+}14]. belief [LKHK11]. Belize [BHC^{+}15]. Bell [Man17a]. benefit [BMM^{+}19]. benefits [EHD^{+}15, KP13, LTGP18, TJD17, WMJ13, WHP16]. Benguela [GKV^{+}15, GES^{+}12, KYK^{+}13]. benthic [Alf10, BUDM15, BMF14, BBD^{+}13, BMNKA^{+}16, BFH13, CSS^{+}15, CAS^{+}16, CSFS10, DHCE13, FMM^{+}12, FDC^{+}16, GGTW17, GNB^{+}16, JLS^{+}15, KJW^{+}18, MS19, MBS^{+}15, MOG^{+}14a, MOG^{+}14b, MGS^{+}15, PG16, PCS^{+}11, RBB^{+}16, RKW^{+}17, SHJI10, SLS18, STR16, SKH^{+}12, SCD^{+}15, TWB^{+}19, TSP^{+}19, WCT^{+}16, WHAA^{+}18, vDHvKR15]. benthic-mapping [CSFS10]. Benthos [RBB^{+}15, GDD^{+}18, MS14, SLV16, GIDP^{+}18]. Benthosema [Sas19]. Bergeseth [Cla18c]. bergylta [DMF13]. Bering [AEC11, BA12, CDAN^{+}14, CN16, DW11, DMO10, EPK^{+}18, FSN^{+}11, GMKS11, HGC^{+}19a, HGC^{+}19b, HSFiS12, HCE^{+}11, HMM12, HCDAF15, IHH^{+}11, MS14, MBIH11, PBW15, PH12a, PCFR13, SDAM^{+}18, SRRZ16, SRW13, SHZ^{+}16, SP13, Urb15, WKL13, WWI^{+}16]. Berlengas [JCSC10]. Bernard [Ano12e]. Bernhard [BF17, Man17a, Man16]. BESITO [GIDP^{+}18]. best [HR14, LB13, MBP19, PMP^{+}16]. better [BKSB12, BT15, DFG^{+}17, SM15, TWHB19]. Between [BDCvD^{+}10, BLD17, BHS11a, BR12, BÁOMRV13, BHCSP^{+}11, CLT^{+}14, CGAD16, Coc17, CLV^{+}14, DTL^{+}11, DHCE13, ECFL15, ET15, Fau11, GS15, Har13, HLCC16, HKK^{+}17, HFH10, HTS^{+}10, HHGtIWGoOH12, HHHE16, JLH12, JRWM11, JWS^{+}18, KCK14, KHC^{+}17, KEM^{+}17, LD17, LAG^{+}16, LBG18, Loh11, LTT^{+}18, MMC19, MKF^{+}15, MKC^{+}11, Mos18, MMSS18, MBP19, OKDJ18, PRB^{+}15, PVD14, PDAC^{+}16, Pin17, PG10, PRO14, PSF12, RBM15, RDB^{+}18a, SPS^{+}10, SHST18, SFD^{+}16, SRKV15, SPV^{+}16, TDN^{+}19, TKA^{+}16, TMP12, VBGG^{+}11, VPSV^{+}10, VBO^{+}15, ZHD10, dJBLH15, tHR11]. Between-year [BDCvD^{+}10]. Beverton [Pla17, PHV^{+}15]. Beyond [BV19b, BV19a, Fri19, GM18, HD18, LKS14a, Mos18, Rid18, Sca18]. Bias [Nel19, HFSV^{+}15, OKDJ18, RDKK15, WOW^{+}17]. biased [HSD^{+}16, KBH^{+}18, OBL12, PRF^{+}17]. biases [TLPS15]. bibliometric [AB14]. big [RGR^{+}11]. bigeye [BMO^{+}19, WLN^{+}13]. bigger [BG18]. Bight [HFSB^{+}14, LGBA11, BFH^{+}18, GRP10, SLS18]. bilateral [MIM^{+}19]. Billfish [ZCH19, SPM^{+}18]. Billfishes [KS18a]. binomial [HKS12]. Bio [GGTW17, LBKTW19, SLT^{+}17]. bio-economic [SLT^{+}17]. bio-logging [LBKTW19]. Bio-physical [GGTW17]. bioaccumulation [BWM^{+}16]. Biochemical [SFD^{+}16, LFH^{+}19]. biodegradable [GHV^{+}18]. Biodiversity [CP15, Ast15b, FOC^{+}18, GRC16b, GM18, GRF^{+}12, HD18, JC15, KYK^{+}13, MST^{+}18, Mos18, RSGM17, RG11, Rid18, TED16, UFLH^{+}13, Van11]. Bioeconomic [GPS^{+}17, TDHC18, DRGCGJ17, DCG10, DGC12, MHHT10, MLT11, Pla17, PDD^{+}10b]. Bioeconomics [GDHFS^{+}18]. Bioerosion [SFCS^{+}17, SD17, STF^{+}17]. biofilms [HTRF17]. biogenic [TB17]. biogeochemical [BFC^{+}16, GDJ11b, LRJ18]. biogeochemistry [Pac18, RRCT^{+}17]. Biogeographic [BFB^{+}11, RBCH10, RHOL11]. biogeographical [CPSCCE^{+}17, GDD^{+}18]. bioindicator [GIPS12]. Biological [Can16, Yar10, CAC^{+}18, CMM^{+}15, CECL16, DPL10, DGPC^{+}15, EHB^{+}15, KJW^{+}18, KCK14, KHC^{+}17, KTMV16, LÁG^{+}19, LKHK11, MTDP11, MGVH^{+}10, MMSS18, NJFH19, NC12, OAM^{+}10, PDCP14, PS13, QCA^{+}18a, QCA^{+}18b, RHB^{+}12, RHB^{+}13, RAB^{+}12, TWHB19, WÓG^{+}16, WMG^{+}17, ZCW11]. biology [AH15, HKKS10, KSZ10, Mat19, Mor18, TM15, dPJGB13]. Biomass [DMJ^{+}14, HMHL19, MYI12, ASDSM18, BS13, DKK15, DZC^{+}13, EDP14, ESD^{+}16, FC19, GRC^{+}14, HKRM18, IFU11, JTJ18, MHHK13, NPB^{+}15, NUÓ^{+}16, OJK^{+}19, OSBG16, PGOM11, PDT^{+}18, PO14, PHK^{+}19, RCH^{+}15, RK16, Sas19, SRGF15, TCH^{+}16, WWI^{+}16, ZBE^{+}15]. biomass- [WWI^{+}16]. biome [PDWH11]. biomes [PDWH11]. biomineralization [VSB^{+}16]. biophotonics [LAD^{+}17]. biophysical [BSHE18, BPST^{+}16, EPH15, HGC^{+}19a, HGC^{+}19b, HDCH^{+}11, HPN^{+}17, HSK^{+}14]. bioregion [APHC15]. biosecurity [APHC15]. biota [ELBS17]. biotelemetry [BBOE10]. Biotic [LEB^{+}14, LHS^{+}19]. biotope [EDBMB14]. bioturbation [COS11]. birds [GSP12, GGV^{+}18, HBB^{+}18, HLL18, SSF^{+}12, HKB^{+}18]. birth [Rot15]. Biscay [AIA^{+}15, AFQ^{+}11, ADB16, ASV^{+}19, ALB^{+}19, BI13, BCBI13, BMC^{+}13, FC19, FAR15, IFU11, LAD^{+}19, LGL^{+}12, LPB^{+}14, LMPP10, LLPV^{+}12, Lor11, MTDP11, NDF^{+}15, PDMG11, PDH^{+}14, Pre19, SMSR10, dPJGB13]. bites [RLW^{+}15]. bitter [MDMC11, SZM^{+}10]. bivalve [FAP^{+}17, OCG13, RMF^{+}15, VMAMAG11]. black [BSB^{+}13, CSB18, KKMS14, KS18b, MMM11, PVCB17, SBCF16, TSBV^{+}15, SG17]. black-backed [TSBV^{+}15]. blackspot [Lor11, PDCP14]. bladder [KRKG16]. Blanca [MDPH12]. blasts [FHD^{+}19]. bleaching [NF16]. bleekeri [TNS13]. blind [FFH^{+}11, FHH^{+}13a, FHH^{+}13b, SSR^{+}16]. blind-reading [FFH^{+}11, FHH^{+}13a, FHH^{+}13b]. bloom [BL15, CHMY15, LF14, LSF^{+}15, MDP^{+}15, SLTL15, TRSM15, VKT^{+}12, VKS^{+}12, WAH^{+}16]. blooms [ET15, KLN^{+}10, LLST15, MMH17, QCA^{+}18a, QCA^{+}18b, TTJ^{+}18, TTJ^{+}19]. blue [CJFS16, CMH^{+}11, CAM^{+}15, DJHO10, GM17, KRML11, LDD^{+}10, LVSF17, LPSF19, LPT10, MARD13, OOD13, PGOM11, SSP^{+}11]. blue-footed [MARD13]. Blueback [TMR^{+}16]. Bluefin [MBAGVG^{+}18, ALWH19, ABHT^{+}16, AGG^{+}18, BWR^{+}15, FAR15, FL14, LBM19, MALM17, Mel16, MLLL11, MBL^{+}17, MTO^{+}18, RBA^{+}19, RCH^{+}19, SLH^{+}17, SWA^{+}16, WLN^{+}13]. bluefish [SMR12]. board [HMP^{+}15]. boards [MBS^{+}15]. Boarfish [ECW^{+}17, FOJ13, FHC^{+}12, HCF^{+}12a, HCF^{+}12b, WMN^{+}11]. boat [MNGA^{+}17]. boats [RGR^{+}11]. bodied [TPBL17]. body [AFK15, DJW^{+}15, GNKS18, KPGH16, KKMS14, SSA^{+}18, TKA^{+}16, WVG^{+}13]. BOFFFFs [HJS14]. bogaraveo [Lor11, PDCP14]. bomb [BTA^{+}18]. bonitos [PKR^{+}19]. boobies [ZHD10]. booby [MARD13]. boom [dMBD11]. boreal [LPD^{+}16a, SRKV15]. borealis [EMP11, JTJ18, KJG^{+}15, OFYS^{+}11]. Boreogadus [KLCC18, KCL18, LCSI17, LCSI18, WSW^{+}19]. Born [SRW13]. Both [MRD^{+}12, AV15]. Bothnia [BMR^{+}11]. Bothnian [GÖN^{+}12, PHGK^{+}16, RHV14]. bottleneck [GMKS11]. bottlenecks [PKHG14]. Bottom [OCR14, AFH^{+}17, CB12, EBH^{+}17, ÉSMGB15, GDJ11b, GLP^{+}11, IZP^{+}16, KIP14, KHPI15, KFN15, MS14, OKDJ18, QHG^{+}12, RFF^{+}14, SKM18, SMZ^{+}14, TWB^{+}19, UBvH11, vDHvKR15, vHCP15, vHCP17]. bottom-dwelling [SMZ^{+}14]. bottom-set [GLP^{+}11, SKM18]. bottom-trawl [KIP14, KFN15, OKDJ18, UBvH11, vHCP15]. Bottom-up [OCR14, ÉSMGB15, GDJ11b, IZP^{+}16, RFF^{+}14]. Boundaries [GM18, DMS18, KKZ^{+}12]. box [BNB19, CSB18]. brachiopod [CPLH16]. brachydactyla [VBGG^{+}11]. brackish [PR16]. brake [KHF14]. branches [GDBM^{+}17]. branching [BIKP^{+}17]. branchline [RBM15]. brasiliensis [dSMGKP12]. brassica [FTM^{+}17]. brassica-florida [FTM^{+}17]. Brazil [GRP10, BNB^{+}14, MAB12, PG10, VHH10]. Brazilian [KGP^{+}15, dSMGKP12]. break [AIA^{+}15]. Breaking [Ove19]. bream [BBOE10, FSC^{+}12, Lor11, PVCB17, PDCP14]. breeding [DMJ^{+}14, JMF12, MRC^{+}18a, MRC^{+}18b, SH14b]. brevicaudata [LLM12]. Brevoortia [ALWH18, BMH^{+}16, HAH^{+}16, RRH^{+}15]. bridge [CLT^{+}14, Dil17]. Bridging [BLD17, GRC16b]. bridles [HRB15a]. brief [KDF^{+}19a]. Bringing [WM14]. Bristol [FSN^{+}11, KZS10, PSGY^{+}12]. British [SQ14, CCS10, GNDC11, LDD^{+}10, LDT^{+}11, MSE12, SKKM15]. brittlestars [CGAD16]. Broad [FHOK14, GF11, BNP19, Qui18]. Broad-scale [GF11]. Broadband [BDW18, BAFRJ18, LCM10, LBLJ17, SCJI10]. broadcast [ROK^{+}18]. broadcast-spawning [ROK^{+}18]. broader [Bro17b]. brood [PRF^{+}17]. brooding [CPB19a, CPB19b]. broods [VBGG^{+}11]. Brosme [HMN^{+}12]. Broughton [MRMK11]. brown [HD11, KS17, RGFT14, SDH^{+}18, SCD^{+}15, STSP17, SRKV15, TH15, TCH^{+}16]. Brunswick [RC12]. Buccinum [HPS^{+}15, HCL^{+}18]. budget [BGM^{+}19, TWHB19]. buffer [BHLS18]. buffers [PSGY^{+}12]. Building [CFA^{+}16, HSD^{+}16, LS10]. built [BTL17]. Bullimore [HR14]. Bump [GS11]. buoy [DW10]. buoyancy [HHH12a]. burden [FGRR11]. burrowed [JH13]. bust [dMBD11]. butterfish [Ada17, JLL17]. Butterworth [RR10]. By-catch [PCBO^{+}18]. Bycatch [APK11, RCC14a, AMS^{+}12, ACC^{+}12, BMR^{+}11, BSK^{+}11, BSC14, FBCD^{+}17, FDC13, FCCLA10, GD15, HCM^{+}15, HB19a, HRB15a, HBHR18, IS15, KFN15, LJB16, LGB^{+}18, MKHK18, OCS14, OSSL16, PDS15a, SLV16, SSF^{+}12, SCC^{+}13, SBR11, SWT^{+}19, SI15, VG15, WSGD^{+}17, WNPY15, ZDF^{+}19, dQCD^{+}10]. bycatches [EGC^{+}17].

C [BTA^{+}18, DKB14]. C. [HCR13]. Cadiz [Llo17]. Cadrin [CMP^{+}11, MAP^{+}11]. Caenogastropoda [BHCSP^{+}11]. cage [LHJ15]. calanoid [HSS^{+}16, IZP^{+}16]. Calanoida [LTQ14]. Calanus [ASDSM18, BTB^{+}17, CCL^{+}14, CSF^{+}17, FHD^{+}19, GPG14, HSS^{+}16, JFJ^{+}17, KGH^{+}16, RFT^{+}16, SG17, YLLG18]. calcification [BJR17, CSJ16a, FAP^{+}17, LAD^{+}17, SLF^{+}16a, SLF16b]. calcified [CCE17, KPD^{+}16, MNG15]. calcifying [NMF^{+}16]. Calcium [CSJ16b, WHH16]. calculating [BPMR17, Ham15]. calibration [HKRM18, Mac11]. California [CS15, dÁMNGS^{+}11, BWR^{+}15, DZ14, FBD^{+}19, GHB^{+}10, KPGH16, KAH^{+}11, KEM^{+}17, KDF^{+}19b, LAD^{+}17, MPBH11, SRS11, SPRL18]. californianus [KPGH16]. Caligus [GKC^{+}16]. call [HJS^{+}17, NDB^{+}16]. Callinectes [GM17]. Callista [VMAMAG11]. Callogorgia [CSBHS^{+}11]. Calloria [CPLH16]. cameras [FCG^{+}16, TBD^{+}18]. campechanus [BTA^{+}18, SBG10]. camtschaticus [DD13, FPRA11, GMKS11, Hje14, HKS12, LPSF19, SLF17, WHNS14, Win15]. Can [BSP15, DFB^{+}10, HBB^{+}13b, Jon14, JOF14, LB13, MHM^{+}11, SBS^{+}10, SH14b, ZE17, ALBR14, BHL^{+}15, DMS16, EHG^{+}14, GMN^{+}17, GDJ11b, Jen13, Mat11, Moo19, NCH17, SMZ^{+}14, SiTiM^{+}11, TMR^{+}16, WHMP15, XCRP13]. Canada [BCOS11, HS10, OSM18, BT10, CC17, CB12, CCS10, FOT^{+}17, Ham14, LLCJ10, MPL^{+}14, MS10, OSBG16, PMM^{+}14, RC12, RBCH10, SBFC10, SSTB15, STC^{+}17, THSM^{+}18, VWMS19]. Canadian [CJFS16, DHNL12, GRKF16, Hub14]. Canal [TBC^{+}14, TPBL17, LTQ14]. Canary [CAMM12, Pac18]. Cancer [BMVPN10, BHS11a, CR11, HBB^{+}16, SQ14]. candidate [HMN^{+}12]. candy [PH13]. candy-fish [PH13]. cannibalism [SMR^{+}10]. Cantabrian [PHA^{+}10]. cantharus [PVCB17]. Cap [MMAS11, PRKASR12]. capacity [MNV^{+}11, PEN^{+}19b]. Cape [GKV^{+}15, GKC^{+}15]. capelin [CMD17, DPBM12, GHM^{+}16, HGRAR10, HF14, LPD^{+}16a, MR14b, OCR14, OAM^{+}10, PGG^{+}12, SSF18b]. capensis [GKV^{+}15, GES^{+}12, WRMJ13]. capitata [BJR17]. Capros [ECW^{+}17, FHC^{+}12, FOJ13, HCF^{+}12a, HCF^{+}12b, WMN^{+}11]. caps [OCS14]. captain [FS14]. Capture [TJR^{+}10, BGM^{+}11, BPH16, CPT^{+}10, ECW^{+}17, FDC13, FWH^{+}15, HJT12, HMP^{+}15, HHZ^{+}18, JGCH18, MMA^{+}10, RB10, SBCF16, TPHdJ19, URE^{+}18, WESS18]. captured [SMS^{+}16]. capturing [OBL12]. Carangidae [NPB^{+}15]. carapace [PWCS17]. carbo [MMM11]. carbon [AML^{+}19, BLMW17, BMM^{+}19, LHJ15, MBBC11, NDM^{+}16, PPHK19, SFD^{+}16]. carbon-steel [MBBC11]. carbonate [CSJ16b, RAH^{+}16, RRCT^{+}17, RST18, WHH16]. carcass [TIS^{+}19]. Carcharhinus [PHB^{+}16]. Carcinus [FLCQ19]. Cardinale [ABJ^{+}19, WW16a]. care [PVCB17]. career [And15, Cam18, DKC^{+}17, MN18, Ove19, Qui18]. Caretta [KCP^{+}11]. Caribbean [BIKP^{+}17, BM16, BMG15, CZBFLÁ^{+}15, CHM15, EMW^{+}16, FPSF11, GBM^{+}19, GBB15, JVRT10, KPBB15]. carinatus [Lap11]. Carlin [MSR^{+}12]. Carlo [LH19]. carneipes [Lav15]. carryover [LSF16]. cartilaginous [CBS15]. case [ALWH19, ACC^{+}12, AFQ^{+}11, AFK15, BMC^{+}18, BFC^{+}16, CAM^{+}15, CLS^{+}18, CGG^{+}11, CSR11, CLMP16, DFB^{+}10, EHT17, FB19, FBM^{+}18, HR14, HMS^{+}13, HJS^{+}17, HHBSM16, JPPY11, KRML11, KPBB15, LGL^{+}12, LHV^{+}16, LBB15, LCG^{+}18, MCAM14, MHH^{+}18, MBC11, MR14a, MHMP^{+}18, NSS15, PCdL15, Pin17, PDR19, RGGFA16, Sas19, SA14, SGM^{+}17, SBH^{+}14, SDT15, SWRC^{+}18, TCS^{+}19, TMW19, TLPS15, TMG16, VF17, WPvBS15, ZH13, dMBD11]. cases [TCF^{+}13]. Casitas [GBB15]. CAST [JFJ^{+}17]. Catalonia [DCVC15]. catastrophic [Hol10, PEN^{+}19b]. Catch [LB11, ZCH19, ANU11, ABRL12, AH15, BCK^{+}15, BKSB12, BT15, BDCvD^{+}10, BHFH14, BTTV^{+}17, BMC^{+}17, CPT^{+}10, CMH^{+}11, CDSP11, CCG14, DPD^{+}15, DMO10, ESAV15, Fau11, FKH^{+}16, FWS^{+}13, GL11, HLCC16, HUPBL18, IFU11, JCA10, JSOO12, JOSO16, KLKD11, KdS14, KVFK19, MONS14, Nel19, NHL^{+}19, OCBJ10, ODSC10, OCG13, PDS15b, PMH^{+}13, PCD^{+}13, PVCB17, PCBO^{+}18, RYS11, SA14, SDH^{+}18, SDH12, SPS^{+}10, STD11, TMP^{+}17, VKN^{+}11, WMLJ17, Wel11, Win15, WHMP15, ZPS^{+}18]. catch-and-release [CPT^{+}10, FWS^{+}13, ODSC10, PVCB17]. catch-and-releases [BCK^{+}15]. Catch-at-age [LB11, HLCC16, JSOO12, JOSO16]. catch-based [ABRL12]. catch-curve [Nel19]. catch-only [ZPS^{+}18]. catch-per-unit-effort [BKSB12]. catch-pooling [SDH12]. catch-ratios [WMLJ17]. Catchability [HPS^{+}11, AFH^{+}17, MTDP11, MVH^{+}11, SLS15, THH18, Tho11, TPHC15, ZBE^{+}15]. catcher [FB11]. catcher-vessel [FB11]. catches [ACA19, AAP14a, AAP14b, Bar19, Ben13, CRAPMN12, CTH^{+}19a, CTH^{+}19b, ECFL15, EGC^{+}17, GGT^{+}14, IHL19, MPR17, NdlPIR18, SSZH12, TPHdJ19, TMR^{+}16, ZRU^{+}15, dlTQMUE10, vHCP17, vdHdGL16]. catching [JH13, TJD17, VOM11]. catchments [RDB^{+}18a]. catenella [MMH17]. cathepsin [LLLB^{+}10]. catshark [PHH^{+}16]. caught [BCKA11, LHK18, SPS15, SLS15]. Causal [SSF18b, PRBF18]. cause [MO18, MFT16]. caused [AS14, BPMR17, ETG^{+}15, NKSE14, SZOL19, VMAMAG11, dLCF^{+}15]. causes [DVV^{+}11]. cave [DFB^{+}10]. caveat [Hut14]. caveats [Sub18a, Sub18b]. Cd [LLLB^{+}10]. cease [GMN^{+}17]. Celtic [MPR17, SGK^{+}11, SRG11, SFH^{+}12, VBO^{+}15]. censored [JCA10]. censoring [TGH^{+}12]. census [SMMK11]. Centers [OdS19]. central [ASI^{+}16b, DCHMHQ12, GCG^{+}10, GHB^{+}10, GFML10, GOS^{+}13, HPN^{+}17, KDF^{+}19b, SRS11, VHS^{+}11, dGGW^{+}11]. centre [AIA^{+}15]. Centrophoridae [DHS19]. Centrophorus [DHS19]. Centropristis [BSB^{+}13, KS18b, SBCF16]. Centroscymnus [VMG11]. Centrostephanus [DDF^{+}12]. century [BJMG19, DK18, Eer12, FKH^{+}16, JCS16, Kam14, KRG^{+}12, Mäl12, Pin17, RFT^{+}16]. Cephalopod [JVRT10, MAS19, Rod10, RGPG10]. Cephalopoda [RYS11, SLF^{+}16a]. Cephalopods [XPC11, PDD^{+}10a, VHH10, XCRP13]. certification [AGSPH14]. certifications [KK19]. cervicornis [BIKP^{+}17]. Cetacean [GRS^{+}14, VHGTFR10, WSGD^{+}17]. cetaceans [LAG^{+}16, MKF^{+}15, SBR11]. Chaetognath [NC12, KN17]. chain [BPH16, DDE^{+}19, HHZ^{+}18]. chaining [CC11]. chalcogramma [BHD13, HFM13, IHH^{+}11, MBIH11, WPWH11, WMJ13]. chalcogrammus [GSC10, KLCC18, KCL18, LCSI18, PDAC^{+}16]. challenge [CBC^{+}11a, Hin15, Hub14, LNH^{+}15, TH15]. challenged [GNKS18]. challengeri [TPHC15]. Challenges [WBW^{+}18, vGA18a, ACT^{+}18, DG17, DWPS11, EP19, EHB^{+}15, GRC16b, HKK^{+}17, KK19, MAGK18, RGS^{+}16, SML14]. Challenging [HRP^{+}16]. chamaeleonticeps [FFPL14]. Change [CFT^{+}19, LKG^{+}19, ASB^{+}11, AFQ^{+}11, BJR17, BKVT16, Bar19, BR12, BJMG19, Ber18b, BNB19, BHS^{+}11b, CAC^{+}18, CTT14, CDSP11, CPS13, CFA^{+}16, ECG15, EP19, EEP^{+}11, FKH^{+}16, Ful11, GLH14, GSP12, GDD^{+}18, GKC^{+}15, HABV19, HMN^{+}12, HP12, HBiI^{+}11, HBB^{+}13c, HBG^{+}19a, HBG^{+}19b, IaF11, iIOKW13, JC15, JTJ18, KK18, KKAV^{+}19, KK19, KCMS19, MS12, MMH17, MSA19, MBC11, MNWB18, MBIH11, MLLL11, Mur11, NRTK^{+}19, OBG12, PBC^{+}16, PEN^{+}19b, RGAS^{+}18, RG11, RDCPvH11, Rod10, RLD^{+}12, RCC^{+}14b, SMR12, SPdJFMS17, SPRL18, Sch14, SH14b, SH16, TTJ^{+}18, TTJ^{+}19, TMG16, VHQ^{+}11, WDOJ15, ZHLK11, ZRN18, vPFF^{+}16]. Changes [BCB^{+}18, CFHC10, JIB^{+}12, JIB^{+}14, LGvPH15, LBK17, MGR^{+}13, NPGT15, NGTT16, OSJ^{+}16, PRKASR12, RIBB^{+}19, SMR12, SLH^{+}17, SS10b, VBGG^{+}11, VCH16, VEA^{+}10, WPGW12, AFK15, BNE^{+}17, BHMR14, BCT^{+}10, BÖBL12, DL11, EHB^{+}15, FC19, FOC^{+}18, GPG14, HDF^{+}19, HFSB^{+}14, Hje14, HF14, JBSD11, JMF12, JDFN12, KYK^{+}13, KHMS19, KGW^{+}12, LPB^{+}14, MRD^{+}12, MWJ11, MKC^{+}11, MPBH11, MFT16, NUO^{+}10, NUÓ^{+}16, NRTK^{+}19, PB11, PGG^{+}12, PHGK^{+}16, RBAB17, RvDW10, RLP^{+}13, SV10, SHW^{+}18, SRHJ12, SiTiM^{+}11, SWA^{+}16, SHST18, SH14a, SLS15, SPM^{+}17, TNY^{+}13, TRPH16, TFM^{+}12, TTJ^{+}18, TTJ^{+}19, TMP^{+}17, VAP12, VQGDAMdL14, WDG^{+}17, WVG^{+}13]. Changing [GAB^{+}14, BFLÁ15, CFT^{+}19, DG17, FB19, FDR^{+}17, IHH^{+}11, KT18, KPL^{+}19, KS18a, PWS^{+}11, RDB^{+}18b, RJP^{+}17, RPB16, SMOH18, SBKA16, TGG^{+}15, TFM^{+}12]. Channel [BPBD15, BCSG13, DMS^{+}12, DPL10, GRC^{+}14, HFSB^{+}14, LVM^{+}11, LGBA11, MLT11, MLC16, NMO^{+}17, YDT14]. Characteristics [GES^{+}12, Alf10, BMG15, CBC^{+}11b, CSD16, DPOL13, HCF^{+}12b, MS15, MHE^{+}16, OFYS^{+}11, RHB^{+}12, RHB^{+}13, RG11, RAB^{+}12, SRHJ12, WNM^{+}13, WPH^{+}15, WMJ13, dHFF^{+}16, dJBLH15]. Characterization [GNDC11, SLTL15, BMVPN10, CAMM12, KCP^{+}11, PVQS12, ŠBGT^{+}17]. characterized [BFH13]. Characterizing [BHD13, BL15, CLT^{+}14, DPD^{+}12, FSS^{+}17, PSKM^{+}16, PBG14, RJ12, KYK^{+}13]. Charleston [GS11]. Charybdis [HZZ^{+}15]. chemical [TNH^{+}10]. chemistry [BIKP^{+}17, KSJ14, MIJ^{+}17, RAH^{+}16, RRCT^{+}17, RST18, TJMC11]. cheni [KKMS14]. Chesapeake [CSR11, DKB14, HAH^{+}16, Jon14, Tow14]. Chile [ACA19, CAC^{+}18, DCHMHQ12, NDM^{+}16, VHGTFR10, VOS10]. Chilean [FWD15, HLCC16, MMH17, RFN^{+}19, dlTQMUE10]. chimaeras [HFNH13]. China [KCP^{+}11, MUW^{+}19, STKT16, Sas19, SLT^{+}17]. chinensis [PWPW18]. Chinook [HTT^{+}17, IS15, KBT14, ME11]. chinstrap [LTT^{+}18]. Chioggia [Can16]. chione [VMAMAG11]. Chionoecetes [HCR13, MDMC11, OSM18, PFD^{+}16, Urb15]. chlorophyll [ALd12, KYK^{+}12, WFS^{+}11]. choice [AUEC12, BBM15, CRAM^{+}14, MUEO17]. choices [FR16]. chokka [DRB10, MRC^{+}10, RDS12, SHL10]. chondrichthyan [WSGD^{+}17]. chondrichthyans [FDJ18, FDJ19]. Chondrus [FLCQ19]. choruses [PSKM^{+}16, PWPW18]. chronic [BMEBM^{+}16]. Chronobiology [Mat19]. chronometer [BTA^{+}18]. Chrysaora [QCA^{+}18a, QCA^{+}18b]. Chrysophrys [WPH^{+}15, WPH^{+}17]. Chukchi [HSFiS12, MYI12, TOA^{+}16]. chum [aFKA11]. CIAC [Rod10]. circle [GPP17]. circulation [JV12a, Kam14, RCH^{+}19, WAH^{+}16]. circuli [UBV^{+}11]. cisco [PBW15]. Citizen [BLS^{+}18, Gro11]. clam [BPBD15, CGM^{+}19, GVA19, MBA^{+}19, TJR^{+}10, VMAMAG11]. clams [LK17]. Clarifying [KK18, LDCR^{+}19]. clarity [Aks15]. class [GHM^{+}16, PJDN12, RMJF14, TLF16]. classes [CSD16]. Classification [FFF16, XMM12, BSR^{+}10, CSS^{+}15, CD14, DMS16, FHC^{+}12, HR14, MBQ10, MDS13, PCS^{+}11, SSM^{+}18, SCJI10, SCL^{+}19]. Claudet [Ber18a, Cla18b]. claw [PWCS17]. cleaner [HSV^{+}17]. Climate [AVGÓ12, DIS^{+}12, HHM^{+}11, Hum17, HCE^{+}11, KAH^{+}11, LBM19, LKG^{+}19, SH16, YCY^{+}13, ASB^{+}11, AFQ^{+}11, BJR17, BKVT16, Bar19, BRH^{+}15, BKT12, BHS^{+}11b, CDSP11, CPS13, CFA^{+}16, DHAH12, EEP^{+}11, EISJ12, aFKA11, Ful11, GLH14, GDD^{+}18, GHF^{+}11, GKR10, GF11, HABV19, HMN^{+}12, HTKB19, HP12, HGGJ15, HBiI^{+}11, HBB^{+}13c, HBG^{+}19a, HBG^{+}19b, HHH^{+}19, HF14, HP11, IaF11, iIOKW13, JIB^{+}12, JIB^{+}14, JC15, KK18, KBC^{+}12, KGN^{+}10, Kim12, KOT^{+}17, KGH^{+}16, KCMS19, LTGP18, LSR^{+}13, MS12, MMH17, MDD^{+}14, MBC11, MLC16, MW19, MBIH11, MLLL11, Mur11, NRTK^{+}19, PBC^{+}16, PHGK^{+}16, PCB^{+}14, PEN^{+}19b, PWS^{+}11, PAB^{+}14, QOH^{+}13, QCA^{+}18a, QCA^{+}18b, RGAS^{+}18, RBAB17, RG11, RCC^{+}14b, SMR12, dÁMNGS^{+}11, SPdJFMS17, SPRL18, SBNL12, SHZ^{+}16, SH14b, SBKA16, TNS13, TFM^{+}12, TTJ^{+}18, TTJ^{+}19, TMP12, TSY^{+}11, WDOJ15, WPGW12]. climate [YLLG18, YYCC16, ZHLK11, tHR11, vPFF^{+}16, HHGtIWGoOH12]. climate-change [MS12, PBC^{+}16]. Climate-driven [HHM^{+}11, HGGJ15, HF14, HP11, LTGP18, RBAB17, TSY^{+}11]. Climate-related [AVGÓ12]. Climatic [AASD12]. Climbing [DB16]. clinically [CZBFLÁ^{+}15]. close [Elv15, ØHS10]. closed [FMM^{+}12, KHH^{+}17a, SSRT10, SG16a]. closure [CKD^{+}17, CBW15, PSM^{+}10, RH17]. closures [BOP^{+}19, BSP15, ETG^{+}15, MBLS15, NC11, OCS14, WESS18]. Clupea [AFLvDH15, GKR10, GOS^{+}13, HNAK12, HP11, KWF11, LHK18, LGR^{+}14, MSK^{+}11, OTAK15, OSJ^{+}16, PHM^{+}14, PKHM12, PKHG14, RHV14, RHOJ10, RTJS14, SOK^{+}12, SKA^{+}12]. clupeid [BCSG13]. clupeoid [YDT14]. clustering [KBB13, SHS11]. clutches [THSM^{+}18]. Cnidaria [CSBHS^{+}11]. Co [vH10, BHB^{+}19, DGB^{+}15, DPE^{+}17, FGRR11, HBC^{+}11, MNGA^{+}17, NPB^{+}15, PLSD16, RAH^{+}16, TMW19, BWM^{+}16, BWH16, BMM^{+}19, DI11, GQX^{+}19, HTRF17, HWR^{+}16, HBG^{+}16, HLMT16, IZP^{+}16, KTLB16, LK17, LBG18, McE17, MWP^{+}16, MFB17, NMF^{+}16, PPS17, RCG^{+}18, RFT^{+}16, SSH^{+}17, SY16, TNFC16]. co-create [BHB^{+}19]. co-located [DPE^{+}17]. Co-management [vH10, DGB^{+}15, FGRR11, HBC^{+}11, MNGA^{+}17, TMW19]. co-occurrence [PLSD16]. co-occurring [NPB^{+}15]. co-variation [RAH^{+}16]. coalescent [GÁE16]. coarse [MONS14]. Coast [CMH19, dPLF^{+}19, ALd12, ABJ^{+}19, BPPP17, BLJ^{+}17, CC17, FSC^{+}12, HFSB^{+}14, JDFN12, KMF13, KBT14, KCP^{+}11, KGP^{+}15, MWJ11, MS10, MDMC11, ØFNH12, RSBC13, SBS^{+}19c, TPL11, WNM10, WKW^{+}10, HABV19, JW18, KHF14, MMC19, NC12, RH17, TSWS15]. Coastal [Aks15, DLL12, JFJ^{+}17, MOG^{+}14a, MOG^{+}14b, SFN^{+}19, ACM^{+}16, AASD12, AFP12, ALMFFBP19, BGM^{+}14, BUDM15, BVS^{+}16, BMS^{+}19, BKTS11, BLM^{+}17, BÁOMRV13, BVW^{+}18, CLT^{+}14, DCHMHQ12, DI11, DRVV^{+}17, DHCE13, ECFL15, FBCD^{+}17, FRM^{+}16, FW10, GDJ11a, GSMA15, HCDAF15, JGCH18, JWS^{+}18, JCS16, KPJC14, KLN^{+}10, KOK^{+}14, KOT^{+}17, KKAV^{+}19, LKR16, LLM12, LGRC14, LSB^{+}10, LTT^{+}18, MQ11, MCSL16, MSL^{+}11, MFB17, MJS^{+}15b, MJS^{+}16, MSAS14, MJAS14, OMTY14, OBG12, OTO^{+}15, PKHG14, PHB^{+}16, RFN^{+}19, RSB^{+}15, ROBC11, RC15, RLP^{+}13, SWB^{+}14, Sei14, SMH^{+}19, SRB^{+}14, SBSE14, TBC15, TMG16, VKN^{+}11, YFP^{+}19, dJBLH15]. coasts [DPE^{+}17, GTM12, TIS^{+}19]. Cochlodinium [KLN^{+}10]. Cod [KFL^{+}15, OAM^{+}10, SSZH12, ABJ^{+}19, BNK10, BGM^{+}14, BKTS11, Bjö18, BHLS18, BCOS11, CTC14, CMH^{+}16, CMH19, CECL16, CTH^{+}19a, CTH^{+}19b, CKD^{+}17, CCSG18, Coo19, DL11, DMZC17, DHZA14, DEH^{+}19, EHHH14, EHB^{+}15, ES18, FRF14, FWH^{+}15, FCB17, GHF^{+}11, GDJ11a, GIW16, GHV^{+}18, GF11, GCS^{+}17, HSSB14, HCC^{+}14, HUT^{+}11, HSNO12, HRC^{+}13, HMS^{+}13, HWK^{+}15, HHM^{+}11, HHH12a, HSK^{+}14, HBC^{+}11, HBD^{+}19a, HBD^{+}19b, HF14, HUPBL18, HMM12, HCDAF15, Hüs10, Hüs11, HHH12b, HRP^{+}16, HGH^{+}16, HHE^{+}16, HHHE16, IHL19, IKF17, JWS^{+}18, JKS^{+}18, JOvdM^{+}14, KDH^{+}10, KCK14, KOK^{+}14, KLCC18, KCL18, KHH^{+}17b, KVSV14, KR12, KRO^{+}16, KRD^{+}18, LSY^{+}14, LCSI17, LCSI18, LFGW13, LYW^{+}18, LWN^{+}13, LK15, LBK17, MV10, Mag11, MSOM10, MCSL16, MM17, MHK19, MJSB14, MO13, MSR14, MKAR^{+}18, MR14b, MSAS14, MJAS14, NUO^{+}10, NC11]. cod [NKSE14, NRTK^{+}19, OFB^{+}17, OAT^{+}12, OBY^{+}14, OBL12, PB11, RHHH^{+}12, REFJ12, SSRT10, SHW^{+}15, SG10, SG16a, SHG^{+}18, SMK^{+}11, SSP12, SBD^{+}15, SO10, SMR^{+}10, Str10, SH15, SBS^{+}19c, TCBD10, TNH^{+}10, TJMC11, UBV^{+}11, UOB^{+}15, WS13, WSW^{+}19, WMG11, Yar10, ZMKC14, ZHD^{+}14, ZLC^{+}17, ZDD^{+}19, dHFF^{+}16]. cod-capelin [HF14]. codend [BHLS18, Mil10]. codends [FMK10]. coefficient [THH18]. coelolepis [VMG11]. cognitive [vPFF^{+}16]. Coherent [NBS^{+}10, ASV^{+}19, GZS^{+}19]. Coho [CHD^{+}18]. Cohort [BHFH14, AST^{+}15a, YCZY19]. coincident [LSB^{+}10]. coioides [LGR^{+}19]. cold [CD11, DK18, EPK^{+}18, JDH^{+}14, PDAC^{+}16, SDAM^{+}18]. collaboration [DGB^{+}15, SOL^{+}15]. collaborative [GD15]. Collapse [RLD^{+}12, BWP10, Coo19, DCNB^{+}10, FLP^{+}13, Mac12, MCSL16, Ósk18, VQGDAMdL14]. collapses [PSM^{+}10]. collect [LÓGR11]. collected [BMR^{+}11, FC19]. collection [DGB^{+}15, ECG15, HGS11, RR11, VQB^{+}11, dGGW^{+}11]. collector [EF15]. Cololabis [iIOKW13, TSY^{+}11, TSS^{+}13]. colonies [JMF12]. colony [RBP15]. colour [SG10, Wil11]. Columbia [CCS10, LEB^{+}14, SQ14, SKKM15]. column [LRJ18]. Combination [HCC^{+}14, SHG^{+}18, AFQ^{+}11]. Combined [BSHC17, CYLT16, EPH15, PPS17, RFHS14, CGK^{+}17, DMT10b, GBJ^{+}15, HB10, QOH^{+}13, RFN^{+}19]. Combining [BWR^{+}15, BHB^{+}19, GFF^{+}17, GHG^{+}10, LPB^{+}14, LMPP10, OHD^{+}16, AH19, HHZ^{+}18, KRG^{+}12, RDB^{+}18a]. come [Abl16, MHE^{+}16]. comes [SH15]. coming [HMLKR18]. Commemorating [Bro14]. Comment [BF17, CN14, GGT^{+}14, HS10, HLL18, HPCW19, MAP^{+}11, Man17a, RR10, SD17, SC16, AAP14a, CMP^{+}11, DK19, FPLB19, FWC^{+}19b, RPM14b, STF^{+}17]. Comments [CMH19, ABJ^{+}19, HKB^{+}18, WW16a]. Commercial [CS15, Lev11, RKCP18, BHS11a, BMC^{+}18, BCK^{+}15, BTTV^{+}17, BCC^{+}19, CR19, ESAV15, EPPL^{+}11, FMK10, GKC^{+}15, HS10, HSHÖ14, IFU11, JWS^{+}18, JKS^{+}18, JBE14, JOvdM^{+}14, LLCJ10, Loh11, MO18, MARD13, NBS^{+}10, PMH^{+}13, PKK13, REFJ12, RK16, SCJI10, TCS^{+}19, WMLJ17, WLB11]. commercially [BVW^{+}18, CSTJ19, HCL^{+}18, OHLK19a, OHLK19b, RCC^{+}14b, SWB^{+}14, SPFM^{+}10, ZGK^{+}17]. commitments [Ric14]. Common [ALMFFBP19, HTKB19, SBNL12, VQB^{+}11, WFM^{+}15, BLB^{+}12, ELBS17, FCCLA10, FB19, HPS^{+}15, HRF^{+}10, JMF12, KGN^{+}10, MCAM14, MAA^{+}19a, MAA^{+}19b, Nel19, PDZ^{+}13, RSB^{+}15, RYS11, SPE14, SCC^{+}13, SMS^{+}17]. commonly [TBG16]. commons [STSP17]. communicating [SCY^{+}10, SS10a, SSB^{+}10a]. communication [BA17]. communications [OCS14]. communities [ATK^{+}10, Alf10, BMC^{+}18, BTL17, BJMG19, BCB^{+}18, BOP^{+}19, BÖBL12, CAS^{+}16, DKK15, DRVV^{+}17, FJSJ15, FOC^{+}18, GLK^{+}10, GNB^{+}16, MCSL16, RDD^{+}10, RDGP13, SLT^{+}14, SHW^{+}18, SRDC^{+}14, Str10, TNFC16, TUUC14, TMG16, VKT^{+}12, YFP^{+}19, ZGK^{+}17]. Community [FGMC^{+}13, PHH^{+}16, dCVB14, AR10, BUDM15, BFB^{+}11, CNBK11, CAC^{+}16, CJC^{+}15, CGM^{+}19, DSN16, DCHMHQ12, DKPM16, DAB^{+}12, EPK^{+}18, FC19, FMM^{+}12, GLH14, Gar11, GHD^{+}10, GGV^{+}18, GRR^{+}11, GFR^{+}12, HFSB^{+}14, HBG^{+}16, HDG^{+}17, JDFN12, KBC^{+}12, LEB^{+}14, LLPV^{+}12, MCO^{+}17, MS19, MGL^{+}15, MVP^{+}14, OBG12, PRKASR12, PVD14, PCS^{+}11, PLSD16, QOH^{+}13, RSGM17, SLM^{+}11, SSH^{+}17, SPdJFMS17, SGK^{+}11, SDAM^{+}18, SMB^{+}16b, TNY^{+}13, TRSM15, TD19, TMP12, WZZ^{+}16, WDG^{+}17, ZHL^{+}19]. Community-based [dCVB14]. community-led [HDG^{+}17]. community-level [PVD14]. comparable [SCY^{+}10]. Comparative [BUNB10, DHAH12, LCSI18, LGR^{+}19, Bak14, BPKH13, BBB^{+}19, DUM^{+}12, GB15a, HTB^{+}16, PWCS17, PNC15, SSS^{+}12]. compare [MST^{+}10, SBS^{+}10]. compared [SV10]. Comparing [LS10, SCY^{+}10, TD19, TB17, Tow14, ZRU^{+}15, tHR11, DMS^{+}12, SS10a, SSB^{+}10a]. Comparison [CR11, DDE^{+}19, FJSJ15, KKZ^{+}12, OYIN18, RRM^{+}12, SBS19a, SBS19b, TUUC14, WLN^{+}13, BCT^{+}10, BMC^{+}17, CBD10b, DW11, DLSJ^{+}19, Fau11, HGS11, LLL^{+}13, MSK^{+}11, MOVU10, OSBG16, RHP^{+}12, SHJI10, SHS^{+}17, SHS^{+}18, SG16a, SHST18, SR17, ZHD10]. Comparisons [MSOF18, CBBM14, DWW^{+}10, PRO14, PSS11, WOW^{+}17, WMLJ17]. compensate [SSM^{+}18]. compete [FRF14]. Competition [BNE^{+}15, HBB^{+}18, HLL18, ET15, HKB^{+}18, SFW^{+}12]. Competitive [SPV^{+}16, HFH10]. competitors [BWH16]. compiled [EDBMB14]. complementary [BTTV^{+}17]. Complete [Ano14a, Ano15, Ano16a, Ano16b, Ano17]. Complex [GB15a, Bak14, DC14, GF14, GCS^{+}17, HRB^{+}15b, NMF^{+}16, Pun17, SBH^{+}14]. complexity [MCAM14]. compliance [Ber18b, HSHÖ14, PMOH13]. component [BHFH14, OD17, PA18, SRRZ16, ZLC^{+}17]. components [BDCvD^{+}10, MTDP11, Pay10, TCF^{+}13]. composition [ALWH19, AH19, BDCvD^{+}10, BHS^{+}16, BMEBM^{+}16, EPK^{+}18, FS14, GS12, GFR^{+}12, HBD^{+}19a, HBD^{+}19b, HOS^{+}15, HHQ11, JDM^{+}16, LHK18, LHLK10, MHE^{+}16, OLM^{+}15, PBW15, PWCS17, RRCT^{+}17, SFH^{+}12, SLH^{+}17, SDAM^{+}18, SMSR10, SBD^{+}15, SMB^{+}16b, TNH^{+}10]. compositional [Tho14]. compositions [AST^{+}15a, Fau11, WNW^{+}15]. Comprehensive [OVW^{+}17, Bow14, EHD^{+}15, GTBT18, PMBM18]. comprise [HGF^{+}18]. compromise [LØM^{+}18, MCO^{+}17]. computer [SKM18]. Conant [KN17]. concensus [BV19a]. concentration [ALd12, ES18]. concentrations [BDF^{+}14, CCC^{+}14, HHM^{+}11, KDF^{+}19b, Rey12, RFT^{+}16]. concept [Ano19, DCMGB^{+}17, EEP^{+}11, KK18, PHV^{+}15, Tho11, WFS^{+}11]. conceptual [PGV^{+}17, TD19]. concerns [JdFBB^{+}10]. concise [SLG^{+}15]. conclusions [GRR^{+}12]. Concurrent [DW10, JDFN12, GM17]. Condition [OBL12, ALWH18, CPT^{+}10, CECL16, FMLM^{+}14, FCB17, GKV^{+}15, GNKS18, GES^{+}12, HS17, KCL18, MSK^{+}11, MKAR^{+}18, RPDF14, WPvBS15, WVG^{+}13, YST^{+}10]. condition-corrected [CECL16]. conditions [BvdMA^{+}18, BGM^{+}19, BLM^{+}17, Cla18d, DL12, EBSW15, ESR^{+}10, FDR^{+}17, GKV^{+}15, GPG14, HGC^{+}19a, HGC^{+}19b, HMM12, KVFK19, LLL^{+}11, NJFH19, OBJA^{+}18, PFD^{+}16, QHG^{+}12, RHV14, RPB16, SAB^{+}16, SDW19, VHS^{+}10, VG15, WCT^{+}16, WVG^{+}13]. conducive [NJFH19]. conduct [CBK18]. confidence [CFA^{+}16]. configurations [LKS^{+}14b]. confinement [BBHC11]. confirm [STB^{+}11]. conflict [LFMJ19a, LFMJ19b, SSF^{+}12]. conflicting [NCH17, SWRC^{+}18]. Conflicts [BVW^{+}18]. Confronting [LVPK10]. congeneric [KN17]. Connecting [ANT18]. connections [KBC^{+}12, Pin17]. Connectivity [HHHE16, SJRB18, BGA^{+}19, CMG^{+}17, CKD^{+}17, FSS^{+}17, GGTW17, HKK^{+}17, JQD^{+}17, KCF^{+}17, KCS10, KPBB15, LBK17, MRSG^{+}17, MJAS14, Ósk18, PG16, PDD^{+}10a, PDAC^{+}16, RTAT17, TVJ^{+}14, ZLC^{+}17]. consecutive [VBGG^{+}11, dLCF^{+}15]. consensus [BV19b]. Consequences [KCK14, MFB17, ODSC10, PVCB17, BI13, EGC^{+}17, GSP12, GVA19, GTGD15, HHM^{+}11, HHM^{+}19, HFH10, KCS10, KHMS19, KGW^{+}12, RTAT17, TDR^{+}16]. Conservation [GRH^{+}14, LLM12, BMS^{+}19, CMA^{+}16, DJW^{+}15, DMS^{+}12, Dil17, EHT17, EPH15, GGTW17, GRC16b, GGIC^{+}15, HMLKR18, HSGL18, HRC^{+}13, HFNH13, KOC^{+}16, Lev11, LNH^{+}15, MBLS15, MAS19, Mer18, SDSA15, SSF^{+}12, XS12, YS15]. Conserving [PMBM18, HJS14]. considerations [CGR16, HHM^{+}19, KWF11, LPFH13, PDCP14, RDT^{+}17, RKCP18, ZHAG17]. considering [BG18]. consistency [GVA19, SSR^{+}16]. consistent [PCLQ^{+}16]. conspecific [KSZ10]. conspecifics [CZBFLÁ^{+}15]. constrained [LPD^{+}16b]. constraints [CBO15, LBJ15]. Construction [LSD^{+}11, RCF^{+}18]. consumers [JvdM15, MB14]. consumption [DMJ^{+}14, DMO^{+}15b, GM17, HBHR18, PLCC18, SRH^{+}14, VRF^{+}16, vDKR14]. containing [SLS18]. contamination [FBL^{+}16, SDW19]. Contemporary [MP15, MAS19]. content [GSMRM^{+}16, KUM^{+}12, SM12, SMSR10]. contents [Der18, MDV^{+}15, XCRP13, YPN^{+}18]. context [ETM^{+}16, FDR^{+}17, GOH^{+}19, PKR^{+}19, RCH^{+}15, RAH^{+}16, SFG^{+}15, SWRC^{+}18]. contexts [BFDGLÁ15]. contiguous [NBS^{+}10]. continental [BNP19, BYQ^{+}19, CNBK11, LLL^{+}13, LPB^{+}14, LK15, MFT16, Mos18, NST^{+}14, RHOJ10, SPFM^{+}10, WKL13]. Continuous [Pet19, BLJ^{+}17, HLD10, MR10, TWHB19]. continuum [DPL16]. contours [HA15]. contraction [PDWH11]. contradiction [BCN16]. contrast [SFD^{+}16]. contrasted [TCH^{+}16]. Contrasting [BA12, HGS^{+}19, HCDAF15, SAB^{+}16, WNM^{+}13, BGM^{+}19, BBM15, HSNO12, HMFW14, LYS^{+}10, MDD^{+}14, tHR11, Coo19]. contrasts [UvHS^{+}14]. contributes [Paf18]. contributing [BPH16]. Contribution [ASDSM18, BGM^{+}19, Gar11, Hol14, MCAM14]. Contributions [SSF^{+}18a, JW18]. Control [Eid18, HCE^{+}11, BPvHR10, CCS10, DHH^{+}14, Dek16, GDJ11b, KKC19, KCMS19, LWT^{+}11, MHM^{+}11, PAB^{+}14, RFF^{+}14, RC15, VPSV^{+}10, VSP^{+}14, ZCW11, ZD14]. control-rule-based [MHM^{+}11]. controlled [CBC^{+}11b, DBL^{+}16]. controlling [aFKA11]. controls [LVPK10, ME11]. convection [WPvBS15]. conventional [DW11, GHV^{+}18, MBBC11, MBS^{+}15, MDdPMQ10]. Convergence [AL10b]. converging [BSB^{+}13]. conversation [GFF^{+}17]. conversion [MSE12]. Convictfish [MSS^{+}19]. convolutional [AHR^{+}19]. cool [MHSG19]. cool-water [MHSG19]. cooling [PCFR13]. cooperation [Fri19, Paf18]. cooperative [DGB^{+}15]. cooperatives [ETG^{+}15, SDH12]. Copepod [VKT^{+}12, BTB^{+}17, BHDB12, EKR^{+}19, FHD^{+}19, MPR12, SG17, SSH^{+}17, TIS^{+}19, vDKR14]. Copepoda [LTQ14]. copepods [HSS^{+}16, IZP^{+}16, KOT^{+}17, OLR^{+}18]. Coping [KKAV^{+}19]. Coral [BFH13, MA13, SLT^{+}14, BJR17, BHCSP^{+}11, BWH16, HBB^{+}13a, JDH^{+}14, Jok16, KK18, LGR^{+}19, MCO^{+}17, MS15, PS13, RBMAOCL18, RTAT17, RWG^{+}18, SPdJFMS17, SMK15, TBHK17, WLB11, BFH13, ZGT13]. Coral-characterized [BFH13]. coralline [DKPM16]. corals [BIKP^{+}17, CCE17, MMAS11, NF16, SFD^{+}16, dCVB14]. core [BNP19]. Coregonus [LHV^{+}16, PBW15]. coriacea [GHB^{+}10]. corkwing [HSD^{+}16, HSV^{+}17]. Cormorant [BÖBL12]. cormorants [HLL18, VEA^{+}10]. corrected [CECL16]. Correcting [KIP14]. correction [EBB^{+}16a, FFN^{+}16, FHH^{+}13a, JIB^{+}14, OKDJ18, Par13, RHB^{+}13]. Corrections [TK18]. correlated [ANT18, BN16]. correlates [Ham15, STF^{+}16]. correlation [OIMP18]. correlations [GJWS19, PSF12]. correlative [MBL^{+}17]. Corrigendum [Ano14b, Ano14c, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano18j]. Corroboration [DZ14]. Coryphaena [MAA^{+}19a, MAA^{+}19b]. cosmopolitan [KN17]. cost [FAP^{+}17, KE14, LSD^{+}11, LPD^{+}16b, MBP19, OCBJ10, PTvOR13, PDD^{+}11, YS15]. cost-constrained [LPD^{+}16b]. cost-effective [OCBJ10, PDD^{+}11]. costs [THSM^{+}18]. Could [FRF14, Bjö18, DG11, DCHN^{+}15]. Counter [CMP^{+}11, MYI12]. Counter-comment [CMP^{+}11]. Countering [HTKB19]. counterintuitive [GHPH15]. Counterpoint [Aga18b, Bas18b, Ber18a, Cla18c, Day18a, Hil18b, Obu18a, SG18a]. countries [BM16, Mat11]. Country [ZRU^{+}15]. counts [GBM^{+}19, LDT^{+}11, WLN^{+}13]. coupled [HDCH^{+}11, MBC11]. coupling [Tow14]. covariability [KBT14]. covariates [SDH^{+}18]. cover [GSP12, PH12a, SEDO19]. coverage [Ham14, SHS^{+}18]. covered [Mil10]. covered-codend [Mil10]. Coviability [May14]. Cox [TKJB19]. cpue [LWT^{+}11, MHHK13]. cpue-based [LWT^{+}11]. CPUEs [PCB^{+}14]. crab [BHS11a, BLB^{+}12, COR15, DD10, DD13, ÉSMGB15, FPRA11, FLCQ19, GM17, GMKS11, HCR13, HZZ^{+}15, Hje14, HKS12, LSF16, MR14a, MDMC11, NHL^{+}19, OCBJ10, OSM18, PWCS17, PRF^{+}17, PSGY^{+}12, PFD^{+}16, SZM^{+}10, SQ14, SLF16b, SLF17, SP13, TBC^{+}14, Urb15, VBGG^{+}11, WHNS14, Win15]. crabs [BMVPN10, KTLB16, KZS10, LVSF17, LPSF19, MQ11]. Crangon [HTS^{+}10, RGFT14, SDH^{+}18, SCD^{+}15, TH15, TCH^{+}16]. crassidens [PDS15b]. Crassostrea [BWM^{+}16]. crassus [HdBR^{+}16]. creaks [TMS^{+}15]. create [BHB^{+}19]. creates [LNH^{+}15]. Creating [PNE^{+}14]. creek [HdBR^{+}16]. crepuscular [TBD^{+}18]. Creseis [RRM^{+}12]. crisis [TSP^{+}19]. crispus [FLCQ19]. cristata [AWS^{+}13]. criterion [BP13, XS12]. Critical [LBJ15, Aks15, DBS^{+}15, Fra15, KE14, Lév15, MA13, RMJF14, RFT^{+}16, SJB15, SJ15]. Critical-Depth-Hypothesis [LBJ15]. critically [VQGDAMdL14]. critique [FWP^{+}16a]. cross [PSF12]. cross-correlations [PSF12]. crowding [TVO12]. Crozet [GTGD15]. Crustacea [DI11]. crustacean [BÁOMRV13, DCHMHQ12, Dig19, FDC13, GFMO11, HMP^{+}15, PHM13]. crustaceans [TWHB19]. crustose [DKPM16]. Cryptic [SHH^{+}17, HGF^{+}18]. CT [CGK^{+}17, EMW^{+}16]. Ctenolabrus [OHLK19a, OHLK19b, SJ16]. Cu [LLLB^{+}10]. cucumber [PLCC18]. Cues [TMS^{+}15]. cultivated [COR15]. cultural [TGH^{+}12]. cultured [KIA^{+}18]. cumulative [AHK^{+}18, KJW^{+}18, ZDF^{+}19]. Cunene [VBA10]. Curiosity [Omm18]. Current [BM16, BWR^{+}15, BLM^{+}17, BSMB18, CKD^{+}17, DCS^{+}11, ETM^{+}16, MBLS15, ØHN14, PWS^{+}11, SFG^{+}15, SDSN14, TMW19, CS15, DZ14, KAH^{+}11, KYK^{+}13, KEM^{+}17, KDF^{+}19b, MFS^{+}14, MOG^{+}14a, MOG^{+}14b, RKB^{+}12, SRS11]. currently [BFS16]. currents [NDP^{+}16, SMH^{+}19, TUUC14]. curve [Nel19, Win15]. curves [CC17, TCBD10]. Cusk [HMN^{+}12]. customary [THR16]. cut [HBC^{+}11]. Cutting [CPSCCE^{+}17]. cuttlefish [GRC^{+}14, HFGT18, LLLB^{+}10, OBJA^{+}18, SLF^{+}16a]. Cuvier [GRG^{+}10]. cycle [BS13, DPD^{+}12, FBL^{+}16, GOPG10, MGPC^{+}14, MLC16, MDPH12, MR14a, PSM^{+}10, PDCP14, WHL^{+}17, MLC16]. Cycles [SGQR^{+}12, BG18, DL12, GKR10]. cycling [TWB^{+}19]. Cyclonic [MFS^{+}14, GHD^{+}10]. Cyclopterus [EDP14, KJKÓ15, KJÓK16, PSO^{+}14]. cygnus [PCdL15, dL14]. cylindrica [BWH16]. Cynoscion [FBL^{+}16]. Cystophora [AWS^{+}13].

D [CBC^{+}11a, SJ16]. dab [vdRMC^{+}17]. dabneyi [BHCSP^{+}11]. daily [BSW^{+}11a, BSW^{+}11b, GMS^{+}18, GL11, MSMW18, PE18, RPDF14, WLN^{+}13, ZGT13, BSW^{+}11a]. daily- [WLN^{+}13]. dam [FBM^{+}18, WABZ16]. Damage [BMSC16, VMAMAG11]. dams [NSS15, TCD^{+}16]. D’Ancona [Can16]. Danish [AUEC12, EMP11, MUEO17, NMM^{+}17, SSP12, UOB^{+}15]. Daphnia [GMM^{+}12]. dark [LBJ15]. Darwin [PSKM^{+}16]. Dascyllus [BBL^{+}15]. Dasyatis [LLM12]. Data [Ben15, Mac12, RR11, WRMJ13, ANU11, AST^{+}15a, ALWH19, AH19, ASI^{+}16a, ATLC16, AHR^{+}19, AJ12, AH15, ASB^{+}19, APV18, APHC15, BHD13, BSK^{+}11, BMGB19, BCK^{+}15, BT15, BGA^{+}19, BL15, BPC10, CSS^{+}15, DMF13, DFG^{+}17, DGB^{+}15, ECG15, EDBMB14, FCG^{+}16, FWC^{+}18, FWC^{+}19a, FWC^{+}19b, FFRR12, GVA19, GB15b, GL11, GML13, GWS17, GKR14, GT19, HGS11, HMP^{+}15, HBC^{+}11, HOV^{+}15, HPCW19, HCE^{+}11, HMR15, JCSC10, JHB^{+}17, JL12b, JVPM12, KJÓK16, KOK^{+}14, KET^{+}17, LJH^{+}12, LSJ10, LCS^{+}15, LS10, Loh11, LPT10, LÓGR11, MKD^{+}18, Man17b, MONS14, MS15, MYI12, MSE12, MHM^{+}11, MRD19, MSR^{+}12, MLC^{+}19, MBP19, MHHK13, Nel19, OSC^{+}17, OCBJ10, ØS15, OLM^{+}15, OIMP18, OKDJ18, OFB^{+}17, PBE^{+}19, PCLQ^{+}16, PE18, PKR^{+}19, PHV^{+}15, PLSD16, PSS11, RML^{+}16, RTS19, RDKK15, SOK^{+}12, SOGT^{+}19, SDH^{+}18, STM19, SCY^{+}10]. data [SHS^{+}18, SSM^{+}18, SPS^{+}10, SDDA11, SHS11, SJUE11, SBR11, SI15, TRSM15, TBC15, Tho14, TFC^{+}19a, TFC^{+}19b, TCF^{+}13, VQB^{+}11, WGCL^{+}19, WJM16, WWCO15, WNW^{+}15, Wil11, WRDF10, ZKD^{+}14, ZPS^{+}18, ZDF^{+}19, dGGW^{+}11, vdKFS^{+}16]. data-based [SCY^{+}10]. data-collection [HGS11, VQB^{+}11]. data-enhancement [HGS11]. Data-limited [Mac12, ASB^{+}19, DFG^{+}17, GKR14, MRD19, PKR^{+}19, WGCL^{+}19, ZDF^{+}19]. data-poor [AH15, APV18, BT15, BPC10, GB15b, HOV^{+}15, JHB^{+}17, MHM^{+}11, PHV^{+}15, PSS11, RTS19, TCF^{+}13, WNW^{+}15]. database [GTBT18, LSD^{+}11, PVQS12]. datasets [BJMG19, KNH11, MBA^{+}19]. date [VHS^{+}10]. dates [BGM^{+}11]. Day [Aga18b, VBA10, MAA^{+}19a, MAA^{+}19b, Day18b]. daylength [SiTiM^{+}11]. Daytime [PMM^{+}14, MPL^{+}14]. Dead [TIS^{+}19, WS13, CZBFLÁ^{+}15, OKDJ18, RCH^{+}15]. dead-zone [OKDJ18]. deal [CEP^{+}18]. Dealing [MN18]. Death [Pep16, DGPR11, Rot15]. DEB [TWHB19]. Debating [PAB^{+}18]. debris [FBL^{+}16]. Decadal [MPBH11, Mäl12, ZRN18]. Decadal-scale [MPBH11]. decade [LBKTW19, RKB^{+}12]. decades [Ben13, EPKR11, OBG12, OTO^{+}15, VAP12, WVG^{+}13]. decapod [DI11]. decapods [GFSC17, PPHK19]. DeCarlo [STF^{+}17]. Decision [CAOG15, DOB^{+}17, LPD^{+}16b, MLMJ17, DMS^{+}12, LLSJ18, MTCF^{+}17, Pun17, Ric11]. decision-making [MTCF^{+}17, Pun17, Ric11]. decision-support [DMS^{+}12]. decisions [AMS^{+}12, SSMD^{+}13, THKP11]. decisive [GHM^{+}16]. deck [FGRR11]. decline [ATK^{+}10, AGMC14, AHM^{+}15, BCC^{+}14, BTA^{+}18, CD16, CCA^{+}18, FLCQ19, HGGJ15, LLST15, SPM^{+}18]. declines [MBIH11, SSF^{+}18a]. Declining [Rey12, ACM^{+}16, BFC^{+}16, KDF^{+}19b, RRM^{+}12]. decommissioning [RKCP18]. decoupled [MS19]. decoys [TMS^{+}15]. decrease [KTLB16]. Decreased [SLF17, JTJ18, NKSE14, SSM^{+}16]. Decreases [GvRKB19, KKH14, LVSF17]. Decreasing [RHV14, TH15, PH12a]. Dedication [Ano11a, Ano12e]. deduced [MLC^{+}19]. Deep [BKMdMS13, HKRM18, KRKG16, NdlPIR18, TBHK17, URE^{+}18, AL10b, BHJ14, BCL^{+}11, BHCSP^{+}11, BFH13, CNBK11, COM11, COS11, CAS^{+}16, DHS19, DFB^{+}10, FDJ18, FDJ19, HPS^{+}11, HR14, HHA^{+}11, HLD10, HBF14, JBE14, KTLB16, LAD^{+}11, MIJ^{+}17, MNWB18, MMAS11, ÖPPM19, PGN^{+}11, PHK^{+}19, RDS12, RWG^{+}18, SSM^{+}18, SO10, TBG16, VMG11, WOW^{+}17, WPvBS15, XCRP13]. deep-dwelling [XCRP13]. deep-living [HPS^{+}11]. Deep-scattering [KRKG16]. Deep-sea [BKMdMS13, BHCSP^{+}11, COM11, CAS^{+}16, DHS19, FDJ18, FDJ19, HR14, HHA^{+}11, HLD10, HBF14, JBE14, KTLB16, MIJ^{+}17, PGN^{+}11, RWG^{+}18, TBG16]. Deep-water [HKRM18, AL10b, BHJ14, CNBK11, LAD^{+}11, MMAS11, VMG11, WOW^{+}17]. deepwater [MV13]. default [JMM^{+}15b]. deficiency [KUM^{+}12]. Defining [JL12b, LFFL13, BFH13, GFSC17, ØHS10, SBB19]. Definition [AGH^{+}11, DUM^{+}12, SV10]. delayed [KVFK19]. Delineating [DPL16, vdKKS11]. deliver [Rid18]. Delphinus [FCCLA10, SCC^{+}13]. delphis [FCCLA10, SCC^{+}13]. delta [TSWS15]. delta-generalized [TSWS15]. demand [BMR10a, BDF^{+}14]. Demersal [GRF^{+}12, AIA13, BHJ14, Cad14, CNBK11, CTH^{+}19a, CTH^{+}19b, CD11, DSA13, DMT10b, EBB^{+}16a, EBB^{+}16b, FBCD^{+}17, FCD^{+}11, GS11, GFR^{+}12, HMS^{+}13, IHL19, KYK^{+}13, LLPV^{+}12, MWJ11, MG15, MST^{+}18, NS17, NPGT15, OI16, OSBG16, ÖPPM19, PRKASR12, PGN^{+}11, QOH^{+}13, RDD^{+}10, SHW^{+}18, SJUE11, SR17, TGDG15, TGG^{+}15, TDHC18, URV^{+}11, UVD^{+}17, WSGD^{+}17, WP13, WHMP15, ZH13, tHR11, vdKKS11]. Demographic [TSPL14, BMS^{+}19, GKC^{+}15, GTGD15, MHE^{+}16, MSA19, NSS15, PG16, TLPS15, VQGDAMdL14, VELT14]. demography [PHB^{+}16, WNM10]. Demonstration [GLC15, DBS^{+}15]. Denmark [DTA^{+}13]. dense [HFGT18]. densities [GVA19, KVSV14, SR16, TVO12, TPHdJ19, VBA10]. Density [ALBR14, LYW^{+}18, SO10, TCBD10, BMEBM^{+}16, CGK^{+}17, DJW^{+}15, DED13, EISJ12, FWT^{+}14, FW10, HS17, HJT12, KRKG16, KIP14, NLB^{+}12, OFVF14, PBJ^{+}14, Pow14, RKL^{+}10, RBMAOCL18, STB^{+}11, VE13, WCT^{+}16, WVG^{+}13, vGA18b, vdSSM^{+}18]. Density-dependence [ALBR14]. Density-dependent [SO10, HS17, KIP14, LYW^{+}18, Pow14, RKL^{+}10, WVG^{+}13, vGA18b, vdSSM^{+}18]. Density-independent [LYW^{+}18, TCBD10, HS17]. dentatus [KSJ14]. depend [BMR10b, PWCS17]. dependence [ALBR14, DED13, FA10]. Dependent [GGK^{+}11, AS14, BOP^{+}19, Bru10, CN16, DGB^{+}15, ECG15, EF15, GBB15, HS17, HBC^{+}11, KIP14, LCSI17, LYW^{+}18, LÓGR11, Mac12, MTP^{+}12, Mat11, MNWB18, PCLQ^{+}16, Pep16, Pow14, RKL^{+}10, SHZ^{+}16, SO10, SBKA16, SI15, TCBD10, WVG^{+}13, ZH13, vGA18b, vdSSM^{+}18]. depending [SY16]. Depleted [GNFM11, AL10a, Coo19, HKKS10, Hol10, Mur10, NNY17]. depletion [ABJ^{+}19, BHC^{+}15, CMH19, DMO10, HJT12, RU12, RFMR10, SBS^{+}19c, TJH15]. depletions [HK14]. deployed [HLA^{+}11]. deposition [LHJ15]. depredating [TGDG15, TTR^{+}19]. Depredation [PDS15b, CJP^{+}14, GTGD15, HCM^{+}15, OSL^{+}15, PMH^{+}13, PH17, RBM15, SOL^{+}15, TMS^{+}15, TGG^{+}15, WNPY15]. depressions [Cor13]. Depth [GSMA15, LBJ15, Aks15, DBS^{+}15, FWH^{+}15, Fra15, JPPY11, KJÓK16, Lév15, MNWB18, MJS^{+}15b, PMM^{+}14, SJB15]. depth-dependent [MNWB18]. depths [SJ15]. derived [AST^{+}15a, MMDSP15, MHHK13, SHS11]. Deriving [WWCO15, Cad13, WNW^{+}15]. Dermochelys [GHB^{+}10]. describe [BLS^{+}18, HOS^{+}15, LJH^{+}12]. described [AGG^{+}18]. describing [CPVS14, GFR^{+}12]. description [GHD^{+}10]. descriptor [CAC^{+}16]. descriptors [dSMGKP12]. Design [SKKM15, vH13, CH16, DOB^{+}17, EBB^{+}16a, EBB^{+}16b, EF15, Hum17, Mat19, POIM12, Par13, PH11, RBM15, RAH^{+}16, SH14a, TRH10, YS15]. designation [GGTW17]. designations [SFNdH16]. despite [GM17]. destinies [SHS15]. destratification [SZOL19]. destroys [ELBS17]. detect [VPH18]. detectable [MGS^{+}15]. detected [CTH^{+}19a, CTH^{+}19b]. Detecting [JTE19, UBvH11]. Detection [SCM19, AGH^{+}11, HPDB19, JOvdM^{+}14, OSC^{+}17, vdHBSM10]. determinants [FFPL14, KdS14]. determination [BGM^{+}11, BT15, BPF19c, Cor12, CAOG15, DWPS11, Hüs10, KMF13, PVQS12, VCF^{+}12]. determine [DPBM12, EHD^{+}15, GW16, LB13, PBW15, SLS15, TJMC11, vDKR14]. determined [APK11, COS11, CMC^{+}12, Elv15, ØHS10]. Determining [GIDP^{+}18, HHQ12, LPD^{+}16a, SCD^{+}15, SMB^{+}16a]. deterrent [OSL^{+}15]. deterrents [CJP^{+}14]. Developing [GT19, HB19b, LSJ10, SKM18, ZCW11, AGSPH14, Mat11, SHL^{+}14]. Development [BCSG13, GRR^{+}11, LWT^{+}11, LSI^{+}15, MVP^{+}14, RHOJ10, SCS^{+}11, TSP12, BA10, BHM^{+}18, CLdR^{+}15, DPOL13, DPRG^{+}18, DDF^{+}12, EMP11, EDP14, FHC^{+}12, GS12, HBC^{+}11, IKF17, KPL^{+}19, LCSI18, MPR12, MWP^{+}16, OTO^{+}15, OSM18, PSH13, PVQS12, PCdL15, QR15, SDDA11, SLF16b, WWMF17]. developmental [ES18]. developments [BV19b, BV19a, SCL^{+}19]. device [FDC13, HBHR18, RBM15, TGDG15]. Devices [MKB^{+}17, BMS^{+}18, BMO^{+}19, CAL14, FS14, FWRM12, HCM^{+}15, MV10, NDF^{+}15, WSGD^{+}17]. dFADs [MKB^{+}17]. Diacavolinia [RRM^{+}12]. Diadema [RBMAOCL18]. diadromous [NSS15]. diagnose [CR11]. Diagnosis [McB15]. diagnostic [Bak14, TJH15]. diagnostics [MRSG^{+}17]. diamond [FMK10]. diamond-mesh [FMK10]. Diatom [LBG18, JBSD11, LFH^{+}19, SY16, VKT^{+}12, WPvBS15]. Dicentrarchus [BvdMA^{+}18, PPS17, SCC^{+}13]. dichotomous [SFNdH16]. Did [MFT16]. Didemnum [KHH^{+}17a]. Diel [ISL19, DHZA14, GDS15, GJWS19, SV10]. Dieppe [DPL10]. Diet [SOGT^{+}19, BÖBL12, CFHC10, FJSJ15, HBD^{+}19a, HBD^{+}19b, JCLO17, KGW^{+}12, LHLK10, TFC^{+}19a, TFC^{+}19b, XPC11, YPN^{+}18]. Dietary [GKC^{+}15]. diets [CFT^{+}19, DPBM12, MRP^{+}16, RHP^{+}12]. differ [BHS11a, BHM^{+}18, CGAD16, MHE^{+}16, SY16, WPH^{+}17]. differences [Ada17, aFKA11, SMB^{+}16a, TNS13, TUUC14, VBA10, WMG11]. different [BFDGLÁ15, BBD^{+}13, CBD^{+}10a, CMP^{+}11, ESR^{+}10, GKV^{+}15, HNAK12, HFH10, IGGSAL^{+}15, MMWC16, MAP^{+}11, MPBH11, OAT^{+}12, PWCS17, PR16, SAB^{+}16, SDW19, SLM^{+}11, SPS^{+}10, SHS15, SSF18b, ZRU^{+}15, ZCW11, ZHV^{+}16, vPFF^{+}16]. Differential [HAF^{+}16, BMM^{+}19]. differentiating [DZ14]. differentiation [BPPB12, GSL^{+}11, VCF^{+}12, VPSV^{+}10]. differing [HMFW14, RST18]. difficult [Hüs10]. digestion [ELBS17]. digital [Ott10]. dilemma [SFNdH16]. dimension [KBH^{+}18]. dimensional [ASI^{+}16a, FOC^{+}18, LSI^{+}15]. dimensions [EBB^{+}16a, EBB^{+}16b, MBC11]. dimorphic [RvDW10]. dimorphism [HSD^{+}16, HCF^{+}12b, TLPS15]. dinoflagellate [IZP^{+}16, MMH17, VKT^{+}12]. diodes [NHL^{+}19]. dioxide [BLMW17, NDM^{+}16, SFD^{+}16]. Diplodus [SLM^{+}11]. diplomacy [HD18]. diploproa [LTGP18]. Direct [AR10, SEDO19, FLCQ19, LSF16]. directed [MKHK18]. Directional [MIM^{+}19]. Directive [GHWD13, GRF^{+}12, PKK13, SRDC^{+}14, SGP^{+}15]. Directly [GBM^{+}19]. disadvantages [CMG^{+}17]. disagreeing [STF^{+}17]. Discard [BCN16, Urb15, BPKH13, CMH^{+}16, CFM^{+}14, CEP^{+}18, CCG14, DVV^{+}11, EPVC14, FCPJ10, KFN15, MSM17, PMP^{+}16, SDT15]. Discarded [UvHS^{+}14, ACA19, Ben13, BMSC16, KVFK19, TBG16, UBvH11]. Discarding [UOB^{+}15, CR19, CFM^{+}14, DMO^{+}15a, MCAM14, PBQ^{+}10]. Discards [CJFS16, BCK^{+}15, BLB^{+}12, FCD^{+}11, GD15, HBC^{+}11, MCAM14, MUEO17, MHMP^{+}18, NDF^{+}15, RCC14a, SRGF15, TPV14, TSBV^{+}15, VGWJ11]. discernible [Ohm19]. discharge [LHJ15, WABZ16]. Disciplinary [MHH^{+}18]. disciplining [Sum13]. Discontinuous [GSP12]. discrepancies [DBA^{+}19]. discriminant [XMM12]. Discriminating [DTL^{+}11]. Discrimination [GOS^{+}13, KIA^{+}18, UBV^{+}11, GNKS18, HA15, MIM^{+}19, NPB^{+}15]. Disease [BHS11a, DJW^{+}15, MDMC11, SZM^{+}10]. diseases [RCC^{+}14b]. disentangles [BHS^{+}16]. Disentangling [BRH^{+}15, CMG^{+}17, VDK^{+}18, CJC^{+}15]. disguised [FO13]. Disko [LHJ12]. disorder [KUM^{+}12, MKC^{+}11]. Dispersal [WHNS14, AHC15, BJH^{+}16, CSB18, HPN^{+}17, LKS^{+}14b, MRC^{+}10, MJAS14, NMO^{+}17, SBL^{+}17]. dispersion [MRMK11, SAS^{+}18]. displacement [BNE^{+}15, VHS^{+}10]. displaying [TKA^{+}16]. Disrupting [PBE^{+}19]. Disruption [MQ11]. disrupts [KTLB16]. dissemination [Pin17]. dissipation [ETG^{+}15]. dissolved [LHJ15]. Dissostichus [GLP^{+}11]. distance [SMB^{+}18]. distant [AAP14a, AAP14b, GGT^{+}14]. distemper [DTA^{+}13]. Distinct [CCSG18, CCC^{+}14, SLTL15]. distinctive [MFP^{+}19]. distinctness [CNBK11]. distorted [SRW13]. distress [EHB^{+}15]. Distribute [Dek16]. Distribution [CMA^{+}16, GKV^{+}15, HFNH13, JHB^{+}12, JLS^{+}15, LNWS18, MMAS11, RHB^{+}12, RHB^{+}13, SDBB15, THR16, VHH10, Ada17, ABHT^{+}16, AFLvDH15, AGG^{+}18, AVGÓ12, BCBI13, BHD13, BSK^{+}11, BRH^{+}15, BM16, BSC14, BMC^{+}13, BFH^{+}18, BLS^{+}18, BBD^{+}13, CRHM19, CN16, DZC^{+}13, DMT10b, DHCE13, EBH^{+}17, EPKR11, EISJ12, EDP14, FPSF11, GCBI17, GL11, GIPS12, GGIC^{+}15, GTBT18, HAR19, HSK^{+}14, JBSD11, JOSO16, KJÓK16, KO12, LDD^{+}10, LVM^{+}11, LYW^{+}18, LCG^{+}18, LVPK10, LLF18, LHS^{+}19, MPL^{+}14, MGR^{+}13, MBL^{+}17, MDMC11, MJS^{+}15b, MJS^{+}16, NSO^{+}19, NUÓ^{+}16, OJK^{+}19, OSM18, PWCS17, PPHK19, QZH^{+}17, RSBC13, RBB^{+}15, RWG^{+}18, SG17, STKT16, SRHJ12, SiTiM^{+}11, STF^{+}16, SRRZ16, SS17, SO16, SO10, SSP^{+}11, TB17, TTJ^{+}18, TTJ^{+}19, TSS^{+}13, TID^{+}10, URE^{+}18, VAP12, VCRPS13, VHGTFR10, VHS^{+}10, VBO^{+}15, YLLG18, YJBL17, YCZY19, ZDD^{+}19]. Distributional [HMM12, GKC^{+}15, UHB12]. distributions [BMR^{+}11, BTTV^{+}17, FCG^{+}16, FMML19, GZS^{+}19, HCC^{+}14, KPL^{+}19, KYK^{+}12, LSJ10, PDH^{+}14, PBL11, ROBC11, SQ14, SWBJ13, TMR^{+}16, WHAA^{+}18]. disturbance [BMEBM^{+}16, DDE^{+}19, GIPS12, SKA15, SKA22, SMB^{+}16b]. disturbances [MBS^{+}15]. disturbs [PCBO^{+}18]. diurnal [EWH16, MDPH12]. divaricata [BIKP^{+}17]. Dive [FWH^{+}15, TTR^{+}19]. divergence [CBD^{+}10a, CMP^{+}11, MAP^{+}11]. divers [DFB^{+}10]. diverse [BAFRJ18, KNH11, MWS^{+}17, MPR12, WGCL^{+}19]. diversity [BKMdMS13, BBD^{+}13, BYQ^{+}19, BMEBM^{+}16, CNBK11, CBL19, DCHMHQ12, FFK^{+}11, GPG14, HCC^{+}14, KN17, MHH^{+}18, OHLK19a, OHLK19b, PSKM^{+}16, SR17, WHAA^{+}18]. diving [FBFP^{+}13, SRC11]. division [GMN^{+}17]. Divisions [FCPJ10, NPGT15]. DNA [FFN^{+}16, BPPB12, DMF13, ES18, FFN^{+}13, GSL^{+}11, GÁE16, IEL^{+}15, NNY16, TNH^{+}10, VPSV^{+}10, VSP^{+}14]. Do [BCRM^{+}18, Eri16, HHA^{+}11, NLB^{+}12, RGR^{+}11, BS14, Cad14, DWW^{+}10, DW10, DW11, HK14, PHO13, RRCT^{+}17, RFT^{+}16, SC16, XCRP13]. Documented [UOB^{+}15, KLKD11]. Does [ESAV15, GÖN^{+}12, HSHÖ14, KJG^{+}15, RDGP13, TNH^{+}10, vDKR14, BMR10b, DFG^{+}17, HSS^{+}16, HHT13, MCO^{+}17, MHE^{+}16, VOM11]. dogfish [HBHR18, VMG11, dlTQMUE10]. dolphin [CJP^{+}14, SCC^{+}13]. dolphinfish [MAA^{+}19a, MAA^{+}19b]. dolphins [FCCLA10, PWPW18, RBM15, TBC15]. Domesticated [GUN^{+}19]. dominance [VKT^{+}12]. dominated [MS19]. Dominica [PEN^{+}19b]. donacium [RHOL11]. doorstep [JWS^{+}18]. Doryteuthis [BNP19, MAS19]. Dosidicus [FBD^{+}19, KAT11, YYCC16]. double [BKTS11]. double-tagged [BKTS11]. down [ÉSMGB15, MHE^{+}16, RFF^{+}14, RC15]. downs [LÓGR11]. downstream [DHKE16, TCD^{+}16]. downwards [SWBJ13]. downwards-looking [SWBJ13]. Drake [STF^{+}16]. drawbacks [GPS^{+}17]. drawn [SDH^{+}18]. dredge [JH13, OCG10, OCG13, OCWG16]. dredges [EBB^{+}16a, EBB^{+}16b, TJR^{+}10]. dredging [BMF14, TTG^{+}15, VMAMAG11]. drift [EAOB15]. drifting [DW10, FS14, MKB^{+}17, dQCD^{+}10]. driftnet [SBR11]. Drinkwater [FPLB19]. drive [SGK^{+}11]. driven [ALWH18, BHMR14, BFH^{+}18, ET15, HGGJ15, HHM^{+}11, HF14, HP11, KTLB16, KKH14, LTGP18, RBAB17, SPE14, TSY^{+}11, YCZY19]. Drivers [LLF18, NSO^{+}19, PGG^{+}17, CFM^{+}14, EKD^{+}15, ÉSMGB15, FMML19, KRG^{+}12, LWS^{+}18, LYS^{+}10, OBG12, OSBG16, QCA^{+}18a, QCA^{+}18b, RFN^{+}19, RHOL11, SSF18b, SMB^{+}16b, TCQ^{+}15]. drives [CSJ16a, EPR^{+}14, RCG^{+}18, SFH^{+}12, SLF^{+}16a]. driving [BCC^{+}14, PRO14, RMJF14]. drop [BPH16]. drop-chain [BPH16]. drum [AS14, LEE17, LBTS^{+}19]. due [AGMC14, DDE^{+}19, RCF^{+}18, RDKK15, WAH^{+}16, vGA18a]. dummy [BDHC19]. Dungeness [SQ14, TBC^{+}14]. durable [Cri12]. duration [MST^{+}18, OSC^{+}17, SS17]. during [BMR10a, BHLS18, BDHC19, CPB19a, CPB19b, DWW^{+}10, Der18, DFS^{+}19, Eri16, ESR^{+}10, GMN^{+}17, GCBI17, HUT^{+}11, HMS^{+}13, HHM^{+}16, HNAK12, IKF17, KR12, MTP^{+}12, MYI12, OBG12, PGS^{+}19a, PGS^{+}19b, PVCB17, PT19, RHV14, SBL^{+}17, SKA^{+}12, SBLS15, SWBJ13, TPHdJ19, TSP^{+}19, VAP12, WPvBS15]. Dutch [BUDM15, Gro11, HKB16, UBvH11, dJBLH15]. dux [GRNG^{+}10]. dwelling [SMZ^{+}14, XCRP13]. dynamic [ABRL12, DGC12, IFU11, LPK12, Peñ19a, PB14, RC15, TWHB19, TMP^{+}17, WBW^{+}18]. Dynamics [BVS^{+}16, MKD^{+}18, ALMFFBP19, BMH^{+}16, CTH^{+}19a, CTH^{+}19b, CDSP11, CBC^{+}11a, DHH^{+}14, DIS^{+}12, DK19, DD13, Eer12, EKR^{+}19, FVSL^{+}10, FLP^{+}13, FPLB19, FOT^{+}17, GMS^{+}18, GW16, GCS^{+}17, HFGT18, HSB18, HSNO12, HHJB14, HMFW14, Hüs11, JCLO17, KHF14, KBC^{+}12, KCS10, KHH^{+}17b, KBH^{+}18, KRD^{+}18, LYW^{+}18, LFMJ19a, LFMJ19b, Lor11, MFP^{+}19, MWP^{+}15, MSOM10, MAS19, MJSB14, MSAS14, OLW^{+}14, OBY^{+}14, OFVF14, PNS18, Pay10, PVD14, PMOH13, PT19, RKL^{+}10, RHOL11, RTJS14, SGV18, SMH^{+}19, SCS^{+}11, SBH^{+}14, SS17, SSB^{+}10b, SSF18b, SPE14, SHZ^{+}16, SSGH18, TCQ^{+}15, TMP12, WÓG^{+}16, WMG^{+}17, dCWMH13, dL14, dMBD11, dPJGB13, DB16]. dypterygia [LDD^{+}10].

E. [HD11]. EAF [JdFBB^{+}10]. earlier [GRR^{+}12]. earliest [PKHM12]. Early [BTB^{+}17, JCS16, MLMS15, VWMS19, BS14, CYLT16, CPVS14, CCD^{+}19, COHdP19, DKC^{+}17, ÉSMGB15, GKV^{+}15, GHM^{+}16, GES^{+}12, HGGJ15, HHH12a, KBT14, LØM^{+}18, LSF16, LEE17, MA13, MWP^{+}16, NKSE14, OBY^{+}14, Pin17, PKHG14, RDCPvH11, RCF^{+}17, SiTiM^{+}11, SLF^{+}16a, SLF16b, TDN^{+}19, VRL^{+}14]. early-career [DKC^{+}17]. early-life [CYLT16, RCF^{+}17]. early-stage [RDCPvH11]. Earth [AAP14a, GGT^{+}14, AAP14b]. easily [TJH15]. East [KYK^{+}12, JMF12, MS10, KCP^{+}11, MFS^{+}14, MUW^{+}19, RKB^{+}12, STKT16, Sas19, SHG^{+}18]. eastern [AEC11, BNK10, BBH^{+}10, CDAN^{+}14, CN16, DW11, DMS^{+}12, DPL10, EPK^{+}18, FSN^{+}11, FGMC^{+}13, FS14, GMH^{+}14, HFSB^{+}14, HHM^{+}11, HPN^{+}17, HCE^{+}11, HMM12, HRP^{+}16, HGH^{+}16, HHE^{+}16, IHH^{+}11, JHB^{+}17, KKZ^{+}12, LVM^{+}11, LGBA11, LCMR^{+}18, LMG^{+}14, LWT^{+}11, MPL^{+}14, MS14, MHLW12, MSL^{+}11, MARD13, MBIH11, NKSE14, OLW^{+}14, PGV^{+}17, PMM^{+}14, RBA^{+}19, SDAM^{+}18, SRRZ16, SHZ^{+}16, SMK15, SP13, THSM^{+}18, VPSV^{+}10, WKK^{+}15, WWI^{+}16, CC17, CECL16, CTH^{+}19a, CTH^{+}19b, EHB^{+}15, KHH^{+}17b, LBTS^{+}19, PE18, RTAT17, WKL13, ZCH19]. easy [Cam18, SDH^{+}18]. EBFM [SKH^{+}12]. Echinodermata [GIPS12]. echinoderms [RPB16]. Echinoidea [GIPS12]. echo [CSD16, DBA^{+}19, MSOF18, RDKK15]. echo-integration [DBA^{+}19, MSOF18]. echoes [FFF16, RD16]. echogenic} [BAFRJ18]. echosounder [BDW18, BAFRJ18, CSS^{+}15, CBD10b, DW10, FA10, LCM10, OSP^{+}13, SCJI10]. echosounders [CD14, DBA^{+}19, HKRM18, LBLJ17, MSOF18, SWBJ13]. echotrace [FA10]. echotraces [GRKF16]. Ecological [BNB^{+}14, CLS^{+}18, DMZC17, GWS17, PH13, SWB^{+}14, Ast15b, BNB19, CMA^{+}16, CBL19, CBO15, CSY^{+}10, CLdR^{+}15, DKC^{+}17, EBSW15, FLLG15, FCB17, FSS^{+}17, GBB15, HSB16, HKK^{+}17, HHM^{+}19, iIOKW13, JW18, LFFL13, LS10, LFMJ19a, LFMJ19b, LVPK10, MWS^{+}17, MS19, MTCF^{+}17, NBS^{+}10, PNS18, PVD14, RDT^{+}17, SBS^{+}10, SS10a, SSB^{+}10a, SMK15, TRH10, TLK^{+}17, TDR^{+}16, TPL11, VHQ^{+}11]. ecologically [RFT^{+}16, SWB^{+}14]. Ecology [GCL^{+}18, VF17, ABHT^{+}16, BCKA11, CTH^{+}19a, CTH^{+}19b, COHdP19, DJW^{+}15, Dil17, DJHO10, ETM^{+}16, HD11, HKKS10, HSNO12, HTT^{+}17, JMM19, JOF14, KHF14, KVA18, KST14, Loh11, LBCOJ19, MRC^{+}18a, MRC^{+}18b, RGC^{+}19, RS10, dPLF^{+}19]. econometric [CMLVGLP14]. Economic [ACS15, HFU^{+}10, PS13, SSR^{+}16, BNE^{+}17, CLS^{+}18, CGR16, DG11, EPVC14, ETG^{+}15, GTGD15, LKHK11, LM10, OCG10, PDR19, Pre19, RML^{+}16, RDT^{+}17, SBH^{+}14, SBB^{+}17, SLT^{+}17, TPL11, TSP^{+}19, TMG16, VQB^{+}11, VHQ^{+}11]. economically [RCG^{+}18]. economics [HP12, KHF14, OLW^{+}14]. Ecopath [LHP^{+}19]. ecophysiological [SG16b]. ecophysiology [CDSP11, SY16]. EcoQO [GRR^{+}11]. ecoregions [OSBG16]. Ecosim [LHP^{+}19]. Ecosystem [DCER^{+}14, GFF^{+}19, KYK^{+}13, LLL^{+}13, LTS^{+}17, LLF18, PGG^{+}12, RSV^{+}17, ZHAG17, ZRN18, AHM^{+}15, ABF^{+}16, AHK^{+}18, AFK15, Bak14, BCRM^{+}18, BPF^{+}18a, BPF^{+}18b, BWR^{+}15, BPF19c, BSB^{+}19, BLJ^{+}17, BSR^{+}10, CMM^{+}15, CLT^{+}14, CLS^{+}18, CBR^{+}11, CCC^{+}14, CFT^{+}19, CBC^{+}11a, CKA^{+}17, CGGH^{+}17, DIS^{+}12, DGL^{+}17, DC14, DPL16, DK18, ECFL15, EAB^{+}17, EP14, FPRA11, Fle15, FWP^{+}16b, GNFM11, GFF^{+}17, GRC16b, GRP10, GGP11, GKV^{+}15, GWS17, GDJ11b, GHB^{+}14, GW16, GES^{+}12, GTBT18, HPDB19, HKS17, HDF^{+}19, HHA^{+}11, HBiI^{+}11, HNAK12, JLH12, JIB^{+}12, JIB^{+}14, JLS^{+}15, KM16, Kim12, KAH^{+}11, KWF11, KPJ^{+}15, KK19, LGL^{+}12, LPFH13, LKS^{+}14c, LYS^{+}10, LB14, LB17, LDCR^{+}19, LSB^{+}10, Llo17, LLSJ18, LBCOJ19, MLS^{+}17, MKL^{+}19, Mat11, MLB^{+}14, NBS^{+}10, OD17, ÖHSNV17, PNE^{+}14, PVD14, PBG14, PA18, RCH^{+}15, RBB^{+}15, RMF^{+}15, RD14]. ecosystem [RBB^{+}16, RRH^{+}15, SSR^{+}16, SHL^{+}14, SCY^{+}10, SRCR12, SRDC^{+}14, SRRZ16, SMOH18, SHST18, SKH^{+}12, SFA^{+}17, SWRC^{+}18, TLR^{+}17, TED16, Tow14, TMP^{+}17, TSP^{+}19, VF17, WM14, YCY^{+}13, ZGK^{+}17, vDHvKR15, vPGFT13]. Ecosystem-based [DCER^{+}14, GFF^{+}19, AHK^{+}18, CLT^{+}14, CKA^{+}17, CGGH^{+}17, DPL16, EP14, Fle15, GNFM11, GRC16b, GWS17, GTBT18, HKS17, LB14, LB17, LDCR^{+}19, LBCOJ19, MKL^{+}19, MLB^{+}14, PNE^{+}14, PVD14, RCH^{+}15, SHL^{+}14, SMOH18, SKH^{+}12, SFA^{+}17, TLR^{+}17, TED16, VF17]. ecosystem-wide [KPJ^{+}15]. Ecosystems [KLC13, AGH^{+}11, BCT^{+}10, BMNKA^{+}16, CP15, CSY^{+}10, CD17, DCS^{+}11, DPE^{+}17, DHAH12, DK19, FPSF11, FPLB19, FMM^{+}12, GG13, GBJ^{+}15, LPK12, MPJ12, MST^{+}10, OTO^{+}15, PBC^{+}16, Pla16, RGAS^{+}18, RHP^{+}12, RFT^{+}16, SSS^{+}12, SBNL12, SOB^{+}11, She15, SBS^{+}10, SS10a, SSB^{+}10a, SBS19a, SBS19b, SP13, TLR^{+}17, TUUC14, TBHK17, WRC^{+}12, vPBLR19]. EcoTroph [GG13]. Ecotypes [EEP^{+}11]. eddies [GHD^{+}10, KCP^{+}11]. eddy [KYK^{+}12]. edge [FFPL14]. edible [BHS11a]. education [CHB^{+}14]. edulis [COR15]. edwardsii [GFMO11, HGGJ15, LMG^{+}14, MLMJ17]. eel [ACM^{+}16, AFP12, ADB16, BPST^{+}16, BM16, BVS^{+}16, BMS^{+}19, BLB^{+}18, BFS16, CCA^{+}18, DB16, Dek16, JTM^{+}16, LØM^{+}18, MM12, PR16, PFH16, PMOH13, SSP12, TCD^{+}16, WBW^{+}18]. eelgrass [MQ11]. eels [CD16, CMC^{+}12, DHKE16, HdBR^{+}16, JTM^{+}16, KIA^{+}18, MMWC16, PDT^{+}18, SC16, WABZ16, WW16b]. Effect [AFLvDH15, MRC^{+}10, MGL^{+}15, MSR^{+}19, NHL^{+}19, SZM^{+}10, VHS^{+}10, WMJ13, AS14, BSB^{+}13, BLMW17, BMR10a, BL15, BMG15, FWC^{+}19b, GHF^{+}11, GM17, GDHFS^{+}18, KKMS14, KPJ^{+}15, LVSF17, McE17, MGS^{+}15, NUO^{+}10, NC11, OE16, OSM18, PG16, PGS^{+}19a, PGS^{+}19b, PH12a, PST14, QR15, RLD^{+}12, RFT^{+}16, SHS^{+}18, SFW^{+}12, SJ16, THM15, dHFF^{+}16]. Effective [Ber18b, Cla18d, Day18c, DSL14, Hil18a, HHQ11, HHQ12, OCBJ10, OD17, Obu18c, Obu18d, PDD^{+}11, SG18b, WKW^{+}10, vHCP15]. Effectiveness [EAB^{+}17, Aga18c, Bas18c, CBW15, CJP^{+}14, GKR14, HHDB14, JRG^{+}16, MHHK13, OCS14, PAB^{+}18, SRB^{+}14, WESS18]. Effects [AIA13, Alf10, APV18, BNE^{+}17, BPBD15, BUDM15, BLM^{+}17, BM15, COR15, CCE17, CPP17, DJW^{+}15, GS15, HBiI^{+}11, HBG^{+}16, HHQ11, HFM13, HLMT16, KLCC18, LK17, LKG^{+}19, LGB^{+}18, LSF16, LPSF19, MWP^{+}16, MSAS14, PWCS17, PEV^{+}16, PFD^{+}16, QHG^{+}12, Rod10, RYS11, SSH^{+}17, SGM^{+}17, SY16, SLF16b, THC16, AR10, BIKP^{+}17, BWM^{+}16, BRH^{+}15, BPKH13, BSHC17, BMM^{+}19, BMEBM^{+}16, CRAPMN12, CYLT16, CPT^{+}10, CLS^{+}18, CB12, CGK^{+}17, CIJ18, CPLH16, DIS^{+}12, DI11, DRB10, DHAH12, FPRA11, FRM^{+}16, FWH^{+}15, FHD^{+}19, FB19, FLCQ19, FWRM12, GIPS12, GBJ^{+}15, GF14, GRH^{+}14, HABV19, HP12, HGS^{+}19, HWK^{+}15, HBG^{+}19a, HBG^{+}19b, HFHS10, Hut14, IZP^{+}16, JDM^{+}16, JC15, KKC19, KOT^{+}17, KIP14, KTMV16, KCMS19, LLLB^{+}10, LCSI18, LPH^{+}15, LCG^{+}18, LHS^{+}19, MS12, MQ11, MSOM10, MMA^{+}10, MGVH^{+}10]. effects [MS15, McB15, MS14, MPB11, MDDP10, MGBvD14, MLLL11, NSS15, OSJ^{+}16, PH13, PMH^{+}13, PBG14, PGN^{+}11, QOH^{+}13, RKL^{+}10, RGAS^{+}18, RCF^{+}18, RRH^{+}15, RFHS14, SAB^{+}16, SMB^{+}18, SA14, SSM^{+}16, SHW^{+}15, SiTiM^{+}11, SFD15, SEDO19, SLT^{+}17, TNFC16, TM15, VG15, WPvBS15, WABZ16, YYCC16, vPFF^{+}16, vdSSM^{+}18, ZE17]. Efficacy [NDF^{+}15, XS12, DH13, PEV^{+}16, SI15]. efficiencies [BLB^{+}12, ZKD^{+}14]. efficiency [DMT10b, EPH15, GHV^{+}18, HJT12, HWK^{+}15, KFL^{+}15, MWJ11, OCG10, PH11, TJR^{+}10, TRPH16, WMLJ17]. efficient [GB15a, KFL^{+}15]. effort [AB16, BNE^{+}15, BUNB10, BKSB12, CSFS10, EMP11, GL11, HSHÖ14, HFU^{+}10, LSJ10, LH19, LPT10, MKD^{+}18, PSB16, PBQ^{+}10, RGFT14, SGV18, STM19, SPM^{+}17, TBD^{+}18, THH18, THC16, VE13, VKN^{+}11, WWCO15, ZCH19, dPJGB13]. effort-management [BUNB10]. efforts [BFH13, GR15]. Egg [LSY^{+}14, VHS^{+}11, ABJ^{+}19, BHDB12, BSW^{+}11a, BSW^{+}11b, BHFH14, BMG15, CMH19, CPB19a, CPB19b, DPOL13, EAOB15, ES18, HHH12a, MSMW18, NPB^{+}15, OBJA^{+}18, PBJ^{+}14, PvdKE^{+}12, RC12, SBS^{+}19c]. egg-based [NPB^{+}15]. eggs [AIA^{+}15, GHPH15, HPN^{+}17, HLMT16, LLLB^{+}10, LCSI18, MSAS14, MJAS14, NPB^{+}15, PBS14, PBJ^{+}14, SDW19, SRKV15]. eightbar [WNM^{+}13]. elasmobranch [DKC12]. elasmobranchs [BCC^{+}14, BCKA11, HSGL18]. electrical [SCD^{+}15, dHFF^{+}16]. electricity [SBB19]. electrified [SLV16]. electro [DDE^{+}19]. electro-fitted [DDE^{+}19]. Electronic [NDB^{+}15, RBC^{+}15, BGA^{+}19, Loh11, SKKM15, vHCP15, vHCP17]. electrotrawling [SBB19]. electrotrawls [SLV16]. eleginoides [GLP^{+}11]. element [FHOK14, TRH10]. elemental [BWR^{+}15, KSJ14]. elements [SBB^{+}17]. elephant [FGRR11]. Elevated [SLF^{+}16a, BIKP^{+}17, BLMW17, BSHC17, BWH16, CGK^{+}17, CPP17, DI11, DKPM16, EMW^{+}16, HWR^{+}16, HLMT16, IZP^{+}16, KPD^{+}16, MWP^{+}16, MFB17, NDM^{+}16, PWCS17, PPS17, RCG^{+}18, RCF^{+}17, SSH^{+}17, SSM^{+}16, SJ16]. Elevations [BT10]. eliminate [RKJ14]. elongatus [GKC^{+}16]. elucidation [LYS^{+}10]. embayment [RC15]. Embedding [MGPC^{+}14]. embrace [Cam18]. embryogenesis [LSF16]. embryonic [DPOL13, SLF16b]. Emergence [CKV^{+}16, PCFR13, TPHC15]. emergent [SLS18]. emerging [ACT^{+}18, LK17, RG11, SML14]. emigrating [MHE^{+}16]. emission [HGC^{+}19a, HGC^{+}19b, Tur19a, Tur19b]. emitting [NHL^{+}19]. emphasis [Kim12]. Empirical [BBD^{+}13, vPFF^{+}16, HOV^{+}15, KCF^{+}17, MDS13, PDH^{+}14, PDD^{+}11, THHH15, THHH18]. empirically [PDD^{+}10a]. encounter [DWW^{+}10, DW10, DW11]. encrasicolus [AIA^{+}15, BJH^{+}14, BCBI13, BPPB12, DBL^{+}16, LMPP10, MOVU10, RBAB17, TID^{+}10]. End [RFT^{+}16, KM16, SCL^{+}19]. end-to-end [KM16]. Endangered [HMN^{+}12, BFH^{+}18, BT10, CAMM12, CST^{+}19, HB19a, HSSB14, VQGDAMdL14]. endemic [CAMM12]. Endogenous [DRGCGJ17, DCG10]. Energy [BMR10a, BGM^{+}19, FWRM12, KOK^{+}14, RRH^{+}15, RIBB^{+}19, SMSR10, TWHB19]. enforcement [DSL14, vH10]. engagement [BA17, OD17]. engaging [SPG^{+}17]. engineering [OSC^{+}17]. England [BMVPN10, CLdR^{+}15, GLC15, KCK14, RKJ14, SA14, SG16a, WDOJ15]. English [BCSG13, BPBD15, CEP^{+}18, CCG14, DMS^{+}12, DPL10, GRC^{+}14, HFSB^{+}14, LVM^{+}11, LGBA11, MLC16, NMO^{+}17, THKP11]. Engraulis [AIA^{+}15, BJH^{+}14, BCBI13, BPPB12, CAC^{+}18, DBL^{+}16, LMPP10, MMG^{+}17, MOVU10, RBAB17, TID^{+}10]. enhance [BKMdMS13, JTE19, RLQ^{+}15]. enhancement [HGS11]. enhances [GQX^{+}19, KVSV14, OMTY14]. enhancing [CAL14]. Enlightening [KLA19]. enough [HHA^{+}11, JMHS14]. enriched [PDD^{+}10a, dBWMD^{+}14]. enrichment [BMM^{+}19, TNFC16]. Ensemble [ASI^{+}16a, RTS19, JC15, SSP^{+}13]. ensuing [TNY^{+}13]. Ensuring [VQB^{+}11]. entangle [BT10]. entanglement [Moo19]. entering [MWS^{+}17]. Enteroctopus [CAOG15]. entire [MWS^{+}17]. entitlements [Cri12]. entrainment [MFS^{+}14]. entrance [BMFSH12]. entry [BSB^{+}13, THKP11, vGA18b]. enumeration [GHB^{+}10]. environment [BCC^{+}14, BPST^{+}16, BCL^{+}11, BBB^{+}19, BPvHR10, CD11, GRF^{+}16, GF14, IHH^{+}11, MGR^{+}13, PNC15, PG10, PHH^{+}16, RAB^{+}12, SGQR^{+}12, SGK^{+}11, SCM19, WCT^{+}16]. Environmental [BSC14, CRAPMN12, DKB14, FTM^{+}17, GIPS12, GGIC^{+}15, HTM11, ME11, RHOL11, STF^{+}16, SPE14, ZD14, Ano19, AHK^{+}18, BPKH13, CAC^{+}18, CAC^{+}16, CSC^{+}13, CGGH^{+}17, DL12, DCMGB^{+}17, DHCE13, DAB^{+}12, EKR^{+}19, EKD^{+}15, FKH^{+}16, FFPL14, FDR^{+}17, Fri19, GKV^{+}15, GPG14, HFHS10, JL12a, Jen13, KKAV^{+}19, LFFL13, LWS^{+}18, LTQ14, LCG^{+}18, LYS^{+}10, LCTB^{+}15, LHS^{+}19, PCB^{+}14, PAB^{+}14, QHG^{+}12, RKL^{+}10, RFN^{+}19, Rid18, RML^{+}16, RDCPvH11, Rod10, RYS11, RMKG11, RLD^{+}12, RLP^{+}13, Sca18, SG16b, SSMD^{+}13, SSB^{+}10b, SBS19a, SBS19b, SFG^{+}15, SSP^{+}11, TCS^{+}19, TKB^{+}15, TMG16, WVG^{+}13, YCZY19, ZOQS10]. environmental-driven [YCZY19]. Environmentally [ALWH18, SBD^{+}15, BSW^{+}11b, KKH14]. environments [SMS^{+}16, WDE^{+}18]. enzymology [Pac18]. epaulette [HWR^{+}16]. epibenthic [BKMdMS13, YFG^{+}17]. epidemiological [BSMB18]. epidemiology [DTA^{+}13]. Epifauna [DD10, BMF14, RDD^{+}10, SLS18]. Epinephelidae [WNM^{+}13]. Epinephelus [LGR^{+}19]. Epipelagic [ROBC11]. equally [CZBFLÁ^{+}15]. equator [FC19]. equatorial [CSC^{+}13, COHdP19, LLL^{+}11, OBLP^{+}19, PDWH11]. equilibrium [KWF11, NCH17, Pla17]. equipped [dMBD11]. equivalence [KSJ14]. era [SLG^{+}15]. Erie [ZRN18]. Erratum [Ano11b, Ano11c, Ano14d, Ano18k, Ano18l, Ano18m, Ano18n, SKA22]. Errina [BHCSP^{+}11]. Error [FFH^{+}11, FHH^{+}13a, FHH^{+}13b, DBM^{+}15, HLCC16, MPB11]. errors [HHQ11, ZCR15]. Escape [BHLS18, BPH16, HMS^{+}13, HWK^{+}15, MWS^{+}17, SPS15, SHS15]. Escaped [DFS^{+}19, KDF^{+}16, SHAM10, SHS15]. escapees [GCS^{+}18, GUN^{+}19, MWS^{+}17, MHMP^{+}18, ŠBGT^{+}17]. Escapement [AFP12, CMC^{+}12, BFS16, DHKE16, FMK10, NMM^{+}17, PMOH13]. escapes [JOvdM^{+}14]. esculentus [KS17]. esmarkii [NWM^{+}12, NLB^{+}12]. essential [BFS16, BDF^{+}14, MQ11, OD17, PCLQ^{+}16]. establish [CAOG15, PSM^{+}10]. established [Hin15]. Establishment [FPRA11, DHCE13]. estimate [AH19, BMC^{+}13, BGA^{+}19, FWD15, FBM^{+}18, HJT12, MHHK13, PGOM11, SWvSR11, SSP12, VKN^{+}11, ZRU^{+}15]. estimated [Ben13, MUW^{+}19, OBLP^{+}19, SG17, TSWS15, HSHÖ14]. Estimates [DHKE16, KPJC14, AST^{+}15a, AAP14a, AAP14b, ACC^{+}12, BHC^{+}15, BSW^{+}11a, BSW^{+}11b, DCHN^{+}15, FMC10, FCPJ10, GÖN^{+}12, GGT^{+}14, GML13, HMHL19, HHT13, HKS12, KCF^{+}17, KRKG16, KFBG15, KIP14, LSJ10, MKR13, Mil10, MMDSP15, NPB^{+}15, OOD13, OKDJ18, PGOM11, RK16, SGM^{+}17, SGV18, SHS^{+}18, TB17, TID^{+}10, WOW^{+}17, WW16b, WMN^{+}11, WWI^{+}16, ZGT13]. Estimating [ALWH19, BMGB19, BDF^{+}14, CMH^{+}16, DMT10b, EBB^{+}16a, EBB^{+}16b, ES18, HTS^{+}10, IS15, JdFBB^{+}10, KET^{+}17, KFBG15, LSF^{+}18, LCG^{+}18, NMM^{+}17, SRGF15, SHL10, TPHdJ19, THH18, TRPH16, VPO17, WMLJ17, vdHdGL16, DPRG^{+}18, DMBD10, FWC^{+}18, FWC^{+}19a, FWC^{+}19b, HTB^{+}16, HPCW19, MTDP11, MONS14, MR10, PCRD18, PDD^{+}10a, SPS^{+}10, SJUE11, TSP12, UFLH^{+}13, WGCL^{+}19, WS13, WFS^{+}11, ZBE^{+}15]. Estimation [ANU11, Sas19, Ben15, BSW^{+}11a, BSW^{+}11b, CC17, CIJ18, DKK15, DVV^{+}11, ESD^{+}16, FFH^{+}11, FHH^{+}13a, FHH^{+}13b, HOV^{+}15, HRP^{+}16, JvdM15, PDMG11, PHK^{+}19, SOK^{+}12, Sub18a, Sub18b, TLF16, TJR^{+}10, THM15, WKP^{+}17, Win15, dlPG15]. estimations [GMS^{+}18, HCL^{+}18, PCLQ^{+}16, SMD^{+}16]. estimator [MSMW18]. estimators [SWT^{+}19, THHH15, THHH18]. Estonia [BVS^{+}16]. Estonian [VEA^{+}10]. estuaries [CSR11, MS19, WBW^{+}18]. Estuarine [SHK17, FRM^{+}16, HGS^{+}19, LEE17, RST18, TDN^{+}19, WAH^{+}16]. Estuary [VWMS19, ADB16, BHDB12, BBHC11, CCA^{+}18, DAB^{+}12, FBL^{+}16, MDPH12, MAB12, TBD^{+}18, VEA^{+}10]. Ethical [HHM^{+}19]. ethics [TGH^{+}12]. Etmopterus [HD11, MIJ^{+}17]. EU [CEP^{+}18, Pas14, PDR19, PKK13, vH13]. eubrachyuran [PRF^{+}17]. eulachon [LGB^{+}18]. Eunicea [EMW^{+}16]. Euphausia [CK18, CPP17, CWRB11, FWT^{+}14, KOC15]. euphausiid [LLF18, MDS13, OCR14]. euphausiids [BDW18, ESD^{+}16, SRRZ16, SRW13, WLL^{+}13]. Europe [Bro14, HBB^{+}16, JL12a, OVW^{+}17, TCH^{+}16, TCF^{+}13, WFM^{+}15]. European [ACM^{+}16, ADB16, BJH^{+}14, BCRM^{+}18, BvdMA^{+}18, BVS^{+}16, BMS^{+}19, BLB^{+}18, BPPB12, CMU^{+}17, CMC^{+}12, CSTJ19, CPB19a, CPB19b, CCA^{+}18, DJW^{+}15, DB16, DHKE16, DBL^{+}16, EBH^{+}17, EHD^{+}15, EKD^{+}15, FMLM^{+}14, FWS^{+}13, GHWD13, GCW^{+}18, LVM^{+}11, LØM^{+}18, LÁG^{+}19, LSF^{+}18, Llo17, LAD^{+}11, MM12, MVH^{+}11, MGVH^{+}10, MAGK18, MDdPMQ10, MSM17, ØU13, PMP^{+}16, PFH16, PDT^{+}18, PPS17, PMOH13, RSGM17, RSB^{+}15, RBAB17, RRCT^{+}17, RML^{+}16, ŠBGT^{+}17, SRDC^{+}14, SDDA11, SRB^{+}14, SSGH18, TID^{+}10, UvHS^{+}14, VCF^{+}12, VSP^{+}14, WPGW12, WKW^{+}10, WSG^{+}17, ZRU^{+}15, dPJGB13, dPLF^{+}19, vdSSM^{+}18]. Eurytemora [BHDB12]. eutrophic [DAB^{+}12, LGBA11]. eutrophication [BMC^{+}18]. euxinus [SG17]. evacuation [SZOL19]. evaluate [ACCA19, CFM^{+}14, DCG10, GBJ^{+}15, HCR13, JGCH18, KTMV16, LMPP10, Nee15, SSRT10, TSP12, TCF^{+}13]. evaluated [CLS^{+}18, PFPG12]. Evaluating [CBW15, DPRG^{+}18, Elv15, FRM^{+}16, FFK^{+}11, GGV^{+}18, HDCH^{+}11, HLA^{+}11, IHH^{+}11, JCA10, KHH^{+}17a, LH19, MV13, NC11, OCS14, OE16, OCWG16, PSGY^{+}12, RRH^{+}15, RJP^{+}17, SPRL18, SI15, THHH15, THHH18, WNW^{+}15, WWI^{+}16, WHP16, WESS18, ACT^{+}18, BBD^{+}13, CGRM^{+}13, CJP^{+}14, EP14, FDC13, GLH14, GT19, HBiI^{+}11, Jok16, KPJ^{+}15, MS10, Por11, SS10a, SSB^{+}10a, SSMD^{+}13, WOW^{+}17, WFM^{+}15, ZCW11]. Evaluation [FDC^{+}16, FS14, GKR14, KdS14, PH11, BNK10, BUNB10, BWP10, BBD^{+}10, CSS^{+}15, CCS10, CKV^{+}16, DHS19, EDBMB14, FMM^{+}12, GRR^{+}12, HFHS10, LKHK11, LBB15, LWT^{+}11, MMRG18, MSR14, OLW^{+}14, PEV^{+}16, PMBM18, RR10, SDDA11, SP13, TD19, Tho11, THKP11, WHMP15]. evaluations [HB10, SBH^{+}14]. evasion [BHM^{+}18]. Event [FMML19, MS14]. events [Hol10, LSR^{+}13, SHS15, ZHL^{+}19]. Evidence [CB12, LFGW13, MO18, SMNE14, WVG^{+}13, dCWMH13, ACT^{+}18, BJH^{+}14, CBD^{+}10a, CMP^{+}11, CCC^{+}14, FCG^{+}16, GRG^{+}10, HS17, IZP^{+}16, JQD^{+}17, JCS16, KVFK19, MAP^{+}11, MQ11, NJFH19, PSDG12, PDZ^{+}13, SLT^{+}17, VF17, WZZ^{+}16, vPFF^{+}16]. evidence-based [ACT^{+}18, NJFH19]. Evolution [Ric14, TMW19, GAB^{+}14, HBB^{+}13b, MSC14, OSP^{+}13, ÖHSNV17, QG12, RPS11, VP19, vH13]. Evolutionary [CBO15, MSA19, BSHE18, JQD^{+}17, MDDP10, PH13, SBS19a, SBS19b]. evolutionary-based [BSHE18]. evolving [BKVT16, KDF^{+}19b, SML14]. ex [MCOS17]. ex-vessel [MCOS17]. exaggerated [DGPR11]. Examining [DL11, MBLS15, KCS10]. example [BUNB10, BRH^{+}15, CP15, DGL^{+}17, GMC17, HHA^{+}11, MNGA^{+}17, RBCH10, SCS^{+}11, TD19, WRDF10]. examples [KOC15, LAD^{+}11]. excess [IGGSAL^{+}15]. exchanges [RDB^{+}18a]. exclusion [RCF^{+}18]. exemplary [PSF12]. exhaustion [CR19]. exhibit [VGWJ11]. exhibits [MFP^{+}19]. existence [Ósk18, RMJF14]. existing [EPH15, Fle15, Sca18, TPV14]. Exit [THKP11, BSB^{+}13]. exoskeleton [PWCS17]. expanded [SV10]. Expanding [PGV^{+}17]. expansion [ASV^{+}19, GLK^{+}10, PDWH11, vdKFS^{+}16]. Expected [MBIH11]. Expedition [Hub14]. experience [AGSPH14, EHHH14, LÓGR11, MV10, POIM12, Par13]. experiences [HGS11, PPD17]. experiment [FFH^{+}11, OAT^{+}12, PH13, SHK17]. Experimental [CH16, ETG^{+}15, GPP17, GLP^{+}11, MMM11, SMS^{+}17, DBS^{+}15, FAP^{+}17, Hum17, LRJ18, LSF^{+}18, RBBC11, SHJI10, SBR11, dlPG15]. experiments [BKTS11, HJT12, MV10, MMA^{+}10, Mat19, McE17, Mil10, RAH^{+}16, SSRT10]. expert [CBL19]. explain [LFMJ19a, LFMJ19b, MRD^{+}12, PHH^{+}16, WCT^{+}16]. explaining [vPFF^{+}16]. explicit [BOP^{+}19, BSW^{+}11b, DMBD10, GLC15, KRG^{+}12, MV13, MR14a, MBA^{+}19, PDMG11]. exploitation [ADB16, BHJ14, DHKE16, HFGT18, LPK12, MO18, MSA19, RFHS14, SWA^{+}16, WGCL^{+}19]. exploited [CSY^{+}10, GLK^{+}10, GMH^{+}14, HSB18, LPB^{+}14, NUO^{+}10, NPGT15, OHLK19a, OHLK19b, Sei14, SS10a, SSB^{+}10a, TB16, VELT14, WMN^{+}11]. Exploiting [LBLJ17, PCBO^{+}18, SSM^{+}18]. Exploration [Lév15, LJB16]. explorations [HOS^{+}15]. exploratory [DCMC19, Tow14]. explore [GL11, RC15]. explored [BJMG19]. Exploring [BPST^{+}16, CBL19, Pre19, SG10, WRDF10, Ano19, BLD17, BMR10a, DCMGB^{+}17, EEP^{+}11, GJWS19, HRB^{+}15b, Pac18, TTJ^{+}18, TTJ^{+}19]. exponential [BMR10a]. Exportation [IGGSAL^{+}15]. exposure [APHC15, KPJ^{+}15, LLLB^{+}10, LVSF17, MFB17, PFD^{+}16, RCF^{+}17, SJ16, SLF16b]. exposure-effect [KPJ^{+}15]. expression [FHD^{+}19, KPGH16]. extends [BTA^{+}18]. extensively [PCBO^{+}18]. extent [BMC^{+}18, DSA13, PGN^{+}11]. Externally [BHMR14]. extinction [FFPL14, MO18, SS12]. extra [LJB16]. Extracting [RBCH10]. Extraction [ELBS17, DPL10, HBB^{+}18, HKB^{+}18, HLL18, TNH^{+}10]. Extreme [SMK^{+}11]. Exxon [SHW^{+}18]. eye [VDK15].

F [PO14]. F. [Arn11]. face [GRS^{+}14, LB11, PCRD18, WDOJ15]. face-to-face [GRS^{+}14]. facilities [RLQ^{+}15]. factor [ALWH18, WPGW12]. Factors [HAH^{+}16, KHPI15, SPM^{+}18, Ast15b, BPH16, BPKH13, CAB^{+}10, Cri12, DKB14, DHCE13, aFKA11, GFSC17, HBB^{+}16, HvDH^{+}10, HSK^{+}14, KHH^{+}17b, LYW^{+}18, LEB^{+}14, MSE12, NC12, RGFT14, RIBB^{+}19, RLP^{+}13, SBFC10, SG16b, THM15, THKP11]. FAD [EGC^{+}17, LCMR^{+}18]. FADs [BMS^{+}18, CAL14, LCMR^{+}18]. fail [KH19]. fails [GDD^{+}18]. Failure [RKJ14, HvDH^{+}10, JDFN12, KGZH16, LSB^{+}10, LGR^{+}14, OTAK15]. fair [Job17]. Falkland [RU12]. false [PDS15b]. families [ÖPPM19]. Family [SDBB15]. far [ACA19, BJH^{+}16, LFMJ19a, LFMJ19b, Pep16]. far-field [BJH^{+}16]. farm [BUDM15, GCS^{+}18, JOvdM^{+}14, MRMK11, RCF^{+}18, SHS15]. farmed [DFS^{+}19, GDJ11a, GUN^{+}19, HFH10, JOvdM^{+}14, KDF^{+}16, ŠBGT^{+}17, SHAM10, SHS15, UBV^{+}11, VOS10]. Farmer [PGS^{+}19a, PGS^{+}19b]. farming [BPPP17, OS14, TKB^{+}15]. farms [FJSJ15, IGGSAL^{+}15, SO16]. Faroe [BUNB10, JHB^{+}12, LHHK12, Mag11, SO10, SMR^{+}10]. fasciatus [CC17]. fast [WRMJ13, Bjö18]. fat [GSMRM^{+}16, KUM^{+}12, MSK^{+}11]. Fate [MBBC11, LNWS18, RBBC11]. fatty [BDF^{+}14, GSMRM^{+}16, LHK18, ÖPPM19, PFPG12, YPN^{+}18]. fauna [CBK18, DMS18, UHB12]. faunal [GNB^{+}16]. FCMSY [TD19]. Fcube [URV^{+}11]. FcubEcon [HFU^{+}10]. feature [OSC^{+}17]. features [KKZ^{+}12]. Fecundity [KS18b, MMWC16, vDTF^{+}14, CPB19a, CPB19b, DDR10, DPRG^{+}18, EKD^{+}15, GMN^{+}17, GMS^{+}18, MM12, PRF^{+}17, RC12, VBGG^{+}11]. fed [PCBO^{+}18]. FEDs [CAL14]. feed [IGGSAL^{+}15, KR12]. feeders [RCG^{+}18]. Feeding [CSF^{+}17, COHdP19, FP12, GS11, HD11, HSSB14, MUW^{+}19, ML12, NDP^{+}16, PHO13, RCG^{+}18, BI13, DBDP10, Eri16, GKV^{+}15, GSMRM^{+}16, GW16, HTT^{+}17, KUM^{+}12, KVSV14, KBH^{+}18, LPSF19, LLPV^{+}12, LGR^{+}14, MTP^{+}12, Mag11, MRD^{+}12, MOVU10, MR14b, OLR^{+}18, PRB^{+}15, PvdKE^{+}12, RHV14, SKA^{+}12, SBLS15, SJRB18, VVP^{+}14, vDKR14]. Feels [TPBL17]. feet [BM12]. Female [OTAK15, KJKÓ15, KJÓK16, MHSG19, PRF^{+}17, SS10b, SLF16b, THSM^{+}18]. ferruginea [BWP10]. fertilisation [GRF^{+}16]. fertilization [BLM^{+}17, BMG15]. festivus [ZSC16]. few [GB15b, OYIN18]. Fewer [GHPH15]. fidelity [DTA^{+}13, NS17, SMK^{+}11, ZHD^{+}14]. field [BJH^{+}16, CRC^{+}16, SD17]. Fiji [PLCC18]. filter [RCG^{+}18]. fin [Elv15, GZS^{+}19, TOA^{+}16, WVG^{+}13]. Final [Aga18a, Bas18a, Cla18b, Day18b, Hil18c, Obu18b, SG18c]. Finding [KHF14]. Fine [DHZA14, ACS15, HAR19, JÓT^{+}12, MPR17, SMH^{+}19, WSG^{+}17]. Fine-scale [DHZA14, ACS15, HAR19, JÓT^{+}12, MPR17, SMH^{+}19]. finfish [FLLG15, Ham14, Jon14, NdlPIR18, PST14, WLB11]. finite [FHOK14]. finmarchicus [FHD^{+}19, GPG14, HSS^{+}16, JFJ^{+}17, RFT^{+}16]. finning [JHB^{+}17, Pas14]. First [WMN^{+}11, BTL17, LSD^{+}11, OJK^{+}19]. Firth [JMF12]. Fish [AHR^{+}19, BPH16, Bjö18, DH13, FFRR12, GS14, GRR^{+}12, GR15, LMPP10, MKB^{+}17, PNE^{+}14, SMS^{+}16, SWS13, AAP14a, AAP14b, AMS^{+}12, AIA13, ACCA19, AFLvDH15, AFH^{+}17, AML^{+}19, AL10a, AL10b, ACS15, AFK15, BI13, BGM^{+}11, Bar19, BTA^{+}18, BTL17, BHL^{+}15, BRH^{+}15, BCB^{+}18, BPKH13, BHJ14, BSHC17, BKTS11, BSP15, BÖBL12, BMS^{+}18, BMO^{+}19, BCSG13, BS14, BVW^{+}18, Bru10, BP13, BFB^{+}11, CAL14, CPT^{+}10, CNBK11, COM11, CTH^{+}19a, CTH^{+}19b, CMLVGLP14, CD11, CCC^{+}14, CFT^{+}19, CSR11, CBO15, CGRM^{+}13, CBS15, CST^{+}19, Dav11, DKK15, DDR10, DWW^{+}10, DW10, DW11, DLSJ^{+}19, DKC12, DCER^{+}14, DMT10b, DBL^{+}16, EEP^{+}11, EPR^{+}14, ETM^{+}16, EISJ12, EP14, EBSW15, FR16, FKH^{+}16, FJSJ15, FB19, FFK^{+}11, FOC^{+}18, FMML19, FS14, FA10, GNFM11, GLH14, GGT^{+}14, GPP17, GLK^{+}10]. fish [GNKS18, GMC17, GOH^{+}19, GS11, GHD^{+}10, GS15, GKC^{+}15, GRR^{+}11, GRF^{+}12, GFR^{+}12, GJWS19, HSV^{+}17, HKKS10, HHR17, HBB^{+}18, HKB^{+}18, HMN^{+}12, HUT^{+}11, HLL18, HGS^{+}19, HJT12, HHJB14, HBiI^{+}11, HBB^{+}13c, HLA^{+}11, Hut14, IGGSAL^{+}15, JTE19, JCLO17, JCS16, JOF14, JBE14, JVPM12, KT18, KUM^{+}12, KHC^{+}17, Kim12, KYK^{+}13, KS18b, KEM^{+}17, KH19, LAD^{+}19, LAG^{+}16, LPH^{+}15, LKS^{+}14b, LHV^{+}16, LBB15, LEB^{+}14, LSB^{+}10, LEE17, LOOO16, MIM^{+}19, MKL^{+}19, MCO^{+}17, MCSL16, MHLW12, MSK^{+}11, MUW^{+}19, MKHK18, MARD13, MKC^{+}11, MDV^{+}15, MVP^{+}14, MLB^{+}14, MN18, MWP^{+}16, Mur10, MFB17, MJS^{+}15b, MJS^{+}16, Nel19, NSS15, NS17, NNY17, NMM^{+}17, NPGT15, NGTT16, NBS^{+}10, OdJN^{+}12, OMTY14, OBG12, OG16, OAT^{+}12, OSBG16, ÖPPM19, PBS14, PBJ^{+}14, PSKM^{+}16, PH13, PNC15, Peñ19a, PRB^{+}15, Pep16, PRKASR12, PDMG11]. fish [PDH^{+}14, PFPG12, PRBF18, PWPW18, PGN^{+}11, PRO14, PSF12, PKK13, PHK^{+}19, Pur18, RBM15, RSGM17, ROBC11, RFF^{+}14, RFHG19, RDD^{+}10, RDGP13, RMJF14, RJP^{+}17, RTJS14, ROK^{+}18, RMKG11, RJ12, RFHS14, SDW19, Sch14, SZOL19, SHS^{+}17, SGK^{+}11, SRG11, SRDC^{+}14, SMZ^{+}14, SRGF15, SiTiM^{+}11, SSM^{+}18, SDAM^{+}18, SHS11, SO16, SR17, SJRB18, SWvSR11, SP14, SCJI10, SWBJ13, SBSE14, TNY^{+}13, TPBL17, TPHdJ19, THHH15, THHH18, TD19, TSP12, TUUC14, TK18, UBvH11, UvHS^{+}14, VAP12, VQGDAMdL14, VVP^{+}14, VDK15, VOM11, VPO17, VPH18, VEA^{+}10, VRL^{+}14, VE13, VP19, WSGD^{+}17, WGCL^{+}19, WHAA^{+}18, Wel11, WESS18, XCRP13, XMM12, YST^{+}10, dBWMD^{+}14, tHR11, vDKR14, vdKKS11, vdHBSM10]. fish-aggregating [FS14]. fish-based [RFF^{+}14]. fish-eye [VDK15]. fish-length [SWvSR11]. fish-stock [RJ12]. fishable [ZBE^{+}15]. fished [BHS11a, BHFH14, HCL^{+}18, MCOS17, ZGK^{+}17]. Fisher [BSF^{+}19, BHB^{+}19, CS15, DMO^{+}15a, EPVC14, Gro11, Hin15, JW18, LGRC14, SDT15, SPP^{+}16, STD11]. Fisheries [BSK^{+}11, HTKB19, Hub14, KBvZP16, NJFH19, PAB^{+}14, RG11, SP13, UOB^{+}15, VQB^{+}11, WFM^{+}15, ASB^{+}11, AB14, AB16, AAP14a, AAP14b, ACCA19, ABHT^{+}16, ACC^{+}12, And15, ASV^{+}19, ACS15, ASB^{+}19, APV18, BLD17, BCRM^{+}18, BSHE18, BCC^{+}14, BNE^{+}15, BUNB10, BMC^{+}18, BKSB12, Ben15, BGA^{+}19, BPC10, Bro14, BLJ^{+}17, CDC15, CMM^{+}15, CAM^{+}15, CGR16, CBR^{+}11, CDSP11, CFT^{+}19, CKD^{+}17, CAS^{+}16, Coc17, CLMP16, CLdR^{+}15, CD17, Cri12, CBD10b, DGPR11, DMO^{+}15a, dMADLP^{+}16, DMJ^{+}14, DWFD13, DCS^{+}11, DUM^{+}12, DPL16, ECG15, EP19, ECFL15, EAB^{+}17, ESAV15, EPR^{+}14, ETM^{+}16, EPPL^{+}11, EP14, FCG^{+}16, FWS^{+}13, FWH^{+}15, FCB17, FVSL^{+}10, FP12, FSS^{+}17, GRC16b, GRC16a, GGT^{+}14, GHF^{+}11, GB15a, GB15b, GWS17, GDHFS^{+}18, GIPS12, GHV^{+}18, GRH^{+}14, GTBT18, GFF^{+}19, HHMF19, HCM^{+}15, HB19a, Har14, HP12, HBB^{+}13b, HSHÖ14]. fisheries [HHA^{+}11, HO14, Hil19a, Hil19b, Hin15, HRB^{+}15b, HFU^{+}10, HMP^{+}15, HBiI^{+}11, HBB^{+}13c, HHH^{+}19, HOV^{+}15, HJS^{+}17, HHZ^{+}18, HB10, HK14, HHHE16, IOK17, IGGSAL^{+}15, JBSD11, Jen13, JMM^{+}15b, JCS16, JKS^{+}18, KHF14, KCA^{+}14, KKZ^{+}12, KKAV^{+}19, KEM^{+}17, Koe11, KST14, KTMV16, KRG^{+}12, KHMS19, KH19, KFN15, Lap11, LHS^{+}18, LGL^{+}12, LWS^{+}18, LGRC14, Lev11, LWN^{+}13, LPD^{+}16b, LH19, LCTB^{+}15, LM10, LLSJ18, LAD^{+}11, LBKTW19, MSC14, MMM11, MS12, MLT11, MV13, MKL^{+}19, MPR17, MKB^{+}17, MP15, May14, MDD^{+}14, MBC11, MSR^{+}19, MTLG14, MO13, MLB^{+}14, MBT^{+}19, MNGA^{+}17, MN18, Mor18, MST^{+}18, Mur11, MHHK13, MHMP^{+}18, NCH17, NdlPIR18, OCS14, OdS19, PGG^{+}17, PPD17, PHGK^{+}16, PVD14, PMH^{+}13, PvOP10, PEN^{+}19b, PWS^{+}11, PFH16, PE18, PBQ^{+}10, PHV^{+}15, PFD^{+}16, PHA^{+}10, RLQ^{+}15, RBM15, RR11]. fisheries [RCH^{+}15, RFHG19, Ric11, Ric14, RD14, RvDW10, RDT^{+}17, RGGFA16, RKCP18, SMR^{+}11, dÁMNGS^{+}11, STM19, SOB^{+}11, SDDA11, SBH^{+}14, SHG^{+}18, SSMD^{+}13, SKH^{+}12, SFD15, SBB^{+}17, SM15, SPS11, SPSP11, SWRC^{+}18, SH16, TDR^{+}16, TJD17, TM15, TJH15, TED16, TGG^{+}15, TMW19, Tow14, TMP^{+}17, TDHC18, THR16, TCF^{+}13, URV^{+}11, UVD^{+}17, VFR^{+}16, VOM11, VRF^{+}16, VG15, WP13, WGCL^{+}19, WHAA^{+}18, WNPY15, WNW^{+}15, WRDF10, WHP16, WCR^{+}19, ZHLK11, ZPS^{+}18, ZHV^{+}16, ZH13, dGGW^{+}11, vGA18a, vHCP15, vH13, vdHdGL16, vdRMC^{+}17, RBCH10]. fisheries-independent [HRB^{+}15b]. fisheries-induced [HBB^{+}13b, HK14, RvDW10]. fishermen [FCB17, Hil16, DMZC17]. fishers [CBL19, Cla18a, GRS^{+}14, Mat11, MHMP^{+}18, PTvOR13, YS15]. Fishery [GGK^{+}11, Hje14, LPFH13, Mat11, PCLQ^{+}16, VGWJ11, ZCH19, ACM^{+}16, AGSPH14, AMS^{+}12, APOG11, AUEC12, ABRL12, APK11, BNK10, BHS11a, BCK^{+}15, BOP^{+}19, BSC14, BNB19, BSP15, BHS^{+}16, BT10, BMSC16, BMC^{+}17, CRAPMN12, CRAM^{+}14, CAL14, CJFS16, CMH^{+}16, CMU^{+}17, CC11, CHD^{+}18, CJP^{+}14, DGC12, DSN16, DPD^{+}15, DMT^{+}10a, DCVC15, DGB^{+}15, ECG15, ECW^{+}17, EMP11, EHT17, EHD^{+}15, ETG^{+}15, EBSW15, EF15, FBCD^{+}17, Fau11, FLLG15, FCD^{+}11, FBFP^{+}13, FMK10, GFF^{+}17, Gar11, GHPH15, GPP17, GG13, GRS^{+}14, GD15, GTGD15, HSD^{+}16, HHDB14, HTM11, Har13, HMHL19, HLCC16, HSHÖ14, HMS^{+}13, HDCH^{+}11, HJS14, HBC^{+}11, HHQ11, HMR15, IS15, JRG^{+}16, JHB^{+}17, JLH12, JL12a, JRWM11, JWS^{+}18, JBE14, Kam14, KHF14, KPL^{+}19, KLKD11, KFL^{+}15, KVFK19, LLG18, LMPP10, LGvPH15, LJB16, LMG^{+}14, LWT^{+}11, LÓGR11]. fishery [Mac17a, MVK18, MKF^{+}15, MWJ11, MBLS15, MLMJ17, MKHK18, MSM17, MNGA^{+}17, ME11, NHL^{+}19, NDF^{+}15, NdlPIR18, ODL16, OCBJ10, OLW^{+}14, OE16, OSSL16, PVQS12, PSB16, PDS15a, PDS15b, PNS18, PCdL15, PH12a, PCD^{+}13, PDR19, PQR10, Pre19, PGN^{+}11, PLSD16, PDD^{+}10b, PLCC18, RH17, RGRL11, RU12, RCF^{+}18, RFMR10, RBP15, RKJ14, RBC^{+}15, SPG^{+}17, SPRL18, SMH^{+}19, SBH^{+}14, SDT15, SPE14, SMMK11, SKKM15, STSP17, SBR11, STD11, SI15, SSZH12, SH15, SWRC^{+}18, SP13, TCS^{+}19, TDR^{+}16, Tho11, TMW19, THC16, TPV14, TPS^{+}11, TSBV^{+}15, UBvH11, Urb15, VF17, VQGDAMdL14, VKN^{+}11, WWCO15, WS13, WLB11, Wil11, WHMP15, ZCW11, ZKD^{+}14, dGGW^{+}11, dLCF^{+}15, dMBD11, vGA18b, vHCP17, OdS19]. Fishery-Dependent [GGK^{+}11, Mat11, PCLQ^{+}16, BOP^{+}19, DGB^{+}15, ECG15, EF15, HBC^{+}11, LÓGR11, SI15]. fishery-driven [SPE14]. Fishery-induced [Hje14]. fishes [BDW18, CKH19, Cla18a, COHdP19, DKB14, FC19, GvRKB19, GCL^{+}18, HP17, KDF^{+}19b, RGC^{+}19, SMB^{+}18, SEDO19, TB17]. Fishing [BMNKA^{+}16, Cor12, EHG^{+}14, GHV^{+}18, HKB16, KTMV16, MMA^{+}10, RCS^{+}17, SGK^{+}11, ANU11, ATK^{+}10, BNE^{+}17, Bas18c, Bjö18, BHS^{+}11b, BVC15, BCC^{+}19, BM15, CRAM^{+}14, CR19, CPT^{+}10, CMC^{+}12, CIJ18, DRGCGJ17, dMADLP^{+}16, DL11, DMO^{+}15b, DSA13, EP19, EPVC14, FBFP^{+}13, FWP^{+}16a, FS14, GML13, GBB16, GLP^{+}11, HTM11, HCC^{+}14, HSHÖ14, HMS^{+}13, HAR19, Hol14, JLH12, JL12b, JCS16, JKS^{+}18, KHH^{+}17a, KRG^{+}12, KHMS19, KGP^{+}15, KKH14, LSD^{+}11, LJH^{+}12, LLL^{+}11, LFFL13, LSJ10, LCMR^{+}18, LYS^{+}10, Loh11, LGB^{+}18, MVH^{+}11, MWJ11, MRC^{+}18a, MRC^{+}18b, MSR^{+}19, MPB11, MDDP10, Moo19, MSR14, MGS^{+}15, MMAS11, NGTT16, OSC^{+}17, OSP^{+}13, OI16, OYI17, PSH^{+}11, PSDG12, PMP^{+}16, PRF^{+}17, PH12b, PBG14, PBQ^{+}10, PQR10, PO14, PLCC18, QHG^{+}12, QOH^{+}13, RGFT14, RRH^{+}15, RJP^{+}17, RLD^{+}12, SGV18, SBNL12, SFH^{+}12, SRGF15]. fishing [SHS^{+}18, SWA^{+}16, SSP12, SMS^{+}17, SPM^{+}17, Str10, SPV^{+}16, SMB^{+}16b, TBD^{+}18, THH18, TJR^{+}10, TKJB19, THKP11, TGG^{+}15, THC16, TSP^{+}19, TMG16, VMAMAG11, VE13, WGCL^{+}19, WWCO15, XCRP13, YFP^{+}19, ZMF12, ZKD^{+}14, ZBE^{+}15, ZDF^{+}19, dHFF^{+}16, tHR11]. fishing-effort [LSJ10]. fishing-suitable [PMP^{+}16]. Fitness [HFH10]. Fitness-related [HFH10]. fitted [DDE^{+}19]. Fitting [Cad13, ØS15]. five [FMK10, ÖPPM19, SDBB15]. fixation [FGMC^{+}13]. fixed [HLL^{+}12, ZDD^{+}19]. fixed-point [HLL^{+}12]. fjord [BJH^{+}16, CSF^{+}17, HSNO12, MVB^{+}18, MSAS14, SMNE14, SAS^{+}18, SDBB15, SHAM10, WAH^{+}16, PR16]. fjords [BLJ^{+}17, GSL^{+}11, MMH17, VHGTFR10, dlTQMUE10]. flatfish [ALBR14, DVH^{+}18, Gro11, HKB16, HLMT16, IHL19, KVFK19, MLT11, UTA^{+}16, VCH16]. flavescens [VOS10, ZRN18]. Flawed [VF17]. flaws [BBD^{+}10, RR10]. fleet [BNK10, CR19, CMM^{+}15, EMP11, KHF14, MTDP11, OCS14, OCG10, OCG13, PNS18, PH11, QG12, RGFT14, THKP11, TPL11]. fleet-based [BNK10]. fleets [MVH^{+}11, SPV^{+}16, TRPH16]. Flemish [MMAS11, PRKASR12]. Flesh [Lav15]. Flesh-footed [Lav15]. fleshy [KPD^{+}16]. flesus [HPN^{+}17]. flexibility [vPGFT13]. flexible [DOB^{+}17, DC14]. flexuosa [EMW^{+}16]. flight [SMB^{+}18]. floated [AFH^{+}17]. floating [HRB15a]. flora [DMS18]. Florida [BM15, FTM^{+}17]. flounder [AGMC14, BHMR14, BWP10, GLC15, HMS^{+}13, HPN^{+}17, JMM19, KSJ14, OFB^{+}17]. flounders [HRB15a]. flourishing [SH15]. Flow [MRSG^{+}17, DVH^{+}18, GDBM^{+}17, HZZ^{+}15, HG10, PHB^{+}16, QG12]. Flower [Lev11]. flows [RRH^{+}15]. FLR [HB10]. fluctuating [BMM^{+}19, DGPC^{+}15]. fluctuation [SS17]. fluctuations [ALWH18, BHFH14, Bro14, EWH16, JIB^{+}12, JIB^{+}14, Sch14, SMB^{+}16a, TUUC14]. fluxes [AML^{+}19, MOG^{+}14a, MOG^{+}14b, RIBB^{+}19]. flying [Che10, YYCC16, YCZY19]. focal [Tho11]. focus [HTB^{+}16, ZBE^{+}15]. focused [CKV^{+}16]. follow [RRCT^{+}17]. following [FLP^{+}13, MGVH^{+}10, RMKG11, SHW^{+}18, TNY^{+}13, TPBL17]. follows [WPH^{+}17]. Food [CST^{+}19, Mag11, RDL^{+}17, AML^{+}19, ASV^{+}19, BHB^{+}19, BSF^{+}19, BHM13, CK18, FRF14, Hil16, KLCC18, KCL18, KBvZP16, LBM19, LLL^{+}13, PCFR13, RFF^{+}14, RDB^{+}18b, RG11, SAB^{+}16, SMSR10, TLR^{+}17, WMJ13, XCRP13, vDKR14]. food-related [WMJ13]. food-web [LBM19, LLL^{+}13, TLR^{+}17]. food-webs [RDB^{+}18b]. foodweb [AHM^{+}15, BPF^{+}19a, BPF^{+}19b, BNB^{+}14, FRM^{+}16, LPB^{+}14, MNV^{+}11, PHGK^{+}16, PBG14]. foodwebs [ASB^{+}11, CLT^{+}14, HCE^{+}11, LD17]. footed [Lav15, MARD13]. footprint [EBH^{+}17]. footprints [JLH12, SHS^{+}18]. Forage [EPR^{+}14, PNE^{+}14, RTJS14, CFT^{+}19, CST^{+}19, DCER^{+}14, EP14, EBSW15, JTE19, KEM^{+}17, LAG^{+}16, LEB^{+}14, MFB17, RD14, SMSR10]. Foraging [HWR^{+}16, VRF^{+}16, BS13, HP11, LTT^{+}18, PWPW18, WCT^{+}16]. foraminifera [RKW^{+}17]. forcing [KAH^{+}11, LTQ14, SSB^{+}10b, VHS^{+}10, YCY^{+}13, ZOQS10]. Forecasted [EGC^{+}17]. Forecasting [RBAB17, CLMP16, GB15b, HHM^{+}19, HBiI^{+}11, SDSN14]. forests [COL^{+}13, FFF16, RIBB^{+}19]. forests/seaweed [COL^{+}13]. formation [LSF^{+}15]. formations [WPH^{+}17]. forming [GR15]. forms [JWS^{+}18]. Forth [JMF12]. Forty [SWA^{+}16]. forward [AH19, SP14]. fossil [GHF^{+}11]. found [STB^{+}11]. foundation [PVD14]. four [BRH^{+}15, Cri12, MSK^{+}11, ML12, OBG12, PWCS17, SG16a, SFW^{+}12, SFN^{+}19, WVG^{+}13]. Fram [GDBM^{+}17]. Framework [GHWD13, GRF^{+}12, HvDH^{+}10, PKK13, SRDC^{+}14, SGP^{+}15, APOG11, BNK10, Fle15, GAB^{+}14, LKS14a, LGRC14, LLSJ18, LVPK10, MV13, MGPC^{+}14, MBP19, PGV^{+}17, Pas14, QG12, RBB^{+}16, SSMD^{+}13, SWRC^{+}18, TM15, TMW19, Tur19a, Tur19b, URV^{+}11, WM14]. frameworks [DSL14, GRC16b, HKS17, Rid18, TD19]. France [FBM^{+}18, RLD^{+}12, CMC^{+}12, VQB^{+}11, dPLF^{+}19]. Free [SSA^{+}18, DW10, MUEO17, SHL10]. free-drifting [DW10]. Free-swimming [SSA^{+}18, SHL10]. French [GVA19, HFSB^{+}14, JTM^{+}16, MTDP11, MPR17, QG12]. frequency [ANU11, BDW18, BMGB19, Bjö18, BAFRJ18, CD14, FB11, FHK14, FWC^{+}18, FWC^{+}19a, FWC^{+}19b, HPCW19, KRKG16, Nel19, SG17, SHS^{+}18, ZE17]. frequently [MWS^{+}17]. fresh [BDHC19, CMLVGLP14]. freshwater [DPD^{+}12, GCS^{+}18, RAB^{+}12, VP19, vdHdGL16]. Froese [HPCW19]. Front [RB10]. frontal [ITT19]. fronts [BdSP11, ROBC11]. frozen [vdHBSM10]. Fucus [MMC19]. fuel [DMO^{+}15b, PGG^{+}17, PTvOR13, PLCC18]. Fuelling [ATK^{+}10]. full [DGB^{+}15, MR14a]. Fully [KLKD11, UOB^{+}15]. fulmars [FBCD^{+}17]. fumatus [MLMS15, OTM^{+}16]. function [BHB^{+}19, DDR10, HG10, HKK^{+}17, KTMV16, LCSI17, PWCS17, XMM12]. functional [FOC^{+}18, JDM^{+}16, KGZH16, LAD^{+}19, LPB^{+}14, LBK17, MBLS15, RDGP13]. functioning [GFR^{+}12, KTMV16, LLL^{+}13]. functions [CK18, PDH^{+}14]. fundamental [GTGD15]. Fundy [BT10, FFK^{+}11]. Funka [LSI^{+}15]. Furman [Man17a]. further [FFRR12]. fuscus [TSBV^{+}15]. Future [ASI^{+}16b, GQX^{+}19, JDH^{+}14, KLC13, SHST18, BJR17, Bas18c, BOP^{+}19, BLM^{+}17, BFH13, DCS^{+}11, DWPS11, DKC^{+}17, Har14, HTKB19, Hin15, Hum17, iIOKW13, Jon14, KT18, KE14, LK17, LBG18, LBCOJ19, MDD^{+}14, MBIH11, MBL^{+}17, NPB^{+}15, PSH13, Pla16, PFD^{+}16, RIBB^{+}19, SHW^{+}15, SHZ^{+}16, SCL^{+}19, SPS11, SSP^{+}13, SDSN14, TLK^{+}17, TTJ^{+}18, TTJ^{+}19, VGB^{+}17]. Future-proofing [JDH^{+}14]. fuzzy [Eid18].

Gadget [AFQ^{+}11, HB10]. Gadget/FLR [HB10]. gadids [LKR16]. Gadoid [HMFW14, KST14, AHM^{+}15, AFH^{+}17]. gadoids [JDFN12, Wri14]. Gadus [ABJ^{+}19, BGM^{+}14, BKTS11, BHLS18, BCOS11, CMH^{+}16, CMH19, CTH^{+}19a, CTH^{+}19b, CKD^{+}17, DHZA14, FWH^{+}15, FCB17, GDJ11a, GIW16, GSC10, GHV^{+}18, GF11, GCS^{+}17, HCC^{+}14, HRC^{+}13, HWK^{+}15, HHM^{+}11, HSK^{+}14, HBD^{+}19a, HBD^{+}19b, HUPBL18, HMM12, Hüs10, Hüs11, HHH12b, HGH^{+}16, IKF17, JKS^{+}18, JOvdM^{+}14, KOK^{+}14, KLCC18, KCL18, KR12, LSY^{+}14, LCSI18, LFGW13, LYW^{+}18, LWN^{+}13, Mag11, MCSL16, MSR14, NUO^{+}10, NRTK^{+}19, OAT^{+}12, OBY^{+}14, OBL12, PDAC^{+}16, SSRT10, SG10, SHG^{+}18, SSP12, SBD^{+}15, SO10, SMR^{+}10, SH15, SBS^{+}19c, TCBD10, TNH^{+}10, UBV^{+}11, WMG11, ZMKC14, ZHD^{+}14, ZLC^{+}17, dHFF^{+}16]. gahi [MAS19, RU12]. gain [MJSB14]. Gaining [IFU11]. Galicia [FBFP^{+}13, GRS^{+}14]. Galician [ALd12]. Gama [Pin17]. game [RJP^{+}17]. game-theoretic [RJP^{+}17]. gamma [HKS12]. gamma-binomial [HKS12]. gammarus [CPB19a, CPB19b, EKD^{+}15, RCF^{+}18]. gannets [GKC^{+}15]. gap [BLD17, CLT^{+}14]. Gaps [MAGK18, ALWH19, Pay10]. Garden [Lev11]. gardens [CGM^{+}19, BFH13]. garfish [BBH^{+}10]. Garibaldi [AAP14a]. gas [BM12, KRKG16, RLQ^{+}15, RKCP18]. gas-bladder [KRKG16]. Gasterosteus [OJK^{+}19]. gastric [GBM^{+}19]. gastropod [BHCSP^{+}11, CGK^{+}17, HCL^{+}18, RCF^{+}17, SSM^{+}16, ZSC16]. Gaussian [TKJB19]. gayi [FWD15, FWD15]. gear [BMC^{+}17, CMM^{+}15, DMT10b, EP19, EBB^{+}16a, EBB^{+}16b, FCG^{+}16, GML13, HKB16, HCR13, KHPI15, KHMS19, KGP^{+}15, Moo19, MUEO17, MHMP^{+}18, WFM^{+}15, ZKD^{+}14]. gear-based [WFM^{+}15]. gears [OSP^{+}13, OI16, TWB^{+}19, WMLJ17]. gelatinous [BÁOMRV13, KE14, RLW^{+}15]. gender [DHZA14]. gender-based [DHZA14]. Gene [BHL^{+}15, DVH^{+}18, FHD^{+}19, HZZ^{+}15, KPGH16, PHB^{+}16]. Gene-associated [BHL^{+}15]. General [MWP^{+}15, AH19, BMC^{+}17, KWF11, UvHS^{+}14]. generalized [BCK^{+}15, HKS12, LPT10, MPR12, RU12, TSWS15]. generated [MFS^{+}14]. generation [GUN^{+}19, OSP^{+}13, OAT^{+}12]. Generic [GB15b]. genes [OAT^{+}12]. Genetic [BHS^{+}16, CAMM12, GDJ11a, GKJ^{+}15, HGF^{+}18, JQD^{+}17, JOvdM^{+}14, KN17, MMRG18, MAS19, PSDG12, PSO^{+}14, SBG10, ŠBGT^{+}17, VPSV^{+}10, VSP^{+}14, ABJ^{+}19, BPPB12, BCSG13, CMH19, FSC^{+}12, GCW^{+}18, GSL^{+}11, GRG^{+}10, JWS^{+}18, KDF^{+}16, KJG^{+}15, KH19, LHS^{+}19, MWS^{+}17, OHD^{+}16, OVW^{+}17, PG16, PSH^{+}11, SLKS10, SG16b, SMH^{+}19, SBS^{+}19c, SFN^{+}19, WKW^{+}10, WSW^{+}19]. genetically [Elv15, ØHS10]. genetics [DVH^{+}18, GÁE16, HCC^{+}14, JMM19, MJSB14, OTM^{+}16, SHK^{+}15, SG16b]. Genome [LÁG^{+}19]. Genome-wide [LÁG^{+}19]. Genomic [MMSS18]. genomics [BSHE18]. genus [CPSCCE^{+}17]. Geographic [EKD^{+}15, AGG^{+}18, EISJ12, MMWC16, WSW^{+}19, ZHD10]. geographical [OAT^{+}12, RBAB17, YCZY19]. geographically [WRDF10]. geolocation [LFM^{+}18, LBK^{+}19]. Georges [DMZC17, FLLG15, GF11, KHH^{+}17a, LF14, TCS^{+}19]. Georgia [FWT^{+}14, SQ14, CWRB11]. geospatial [HSB16, KNH11]. Geostatistical [TSWS15, TCBD10]. Geraldton [TMG16]. German [PVQS12, RGFT14, Str10, SSZH12]. GES [ASV^{+}19]. get [GRH^{+}14]. Getting [ZBE^{+}15]. ghost [BM15]. Ghosts [HK14]. Giakoumi [Hil18b, SG18c]. Giant [BLB^{+}12, GVA19, GRNG^{+}10, KPJC14, MR14a]. giganteus [TBG16]. gigas [BWM^{+}16, FBD^{+}19, KAT11, YYCC16]. gillnet [AUEC12]. gillnets [GHV^{+}18, SKM18, SSF^{+}12]. gilthead [FSC^{+}12]. Girella [BBHC11]. Gironde [RLD^{+}12]. GIS [LVM^{+}11, RSY11, SMR^{+}11, SWRC^{+}18]. GIS-based [RSY11, SWRC^{+}18]. giving [Omm18]. glacialis [BTB^{+}17, CSF^{+}17, HSS^{+}16]. gladius [CSC^{+}13, GMD^{+}11, SSA^{+}18]. Glass [ADB16, JTM^{+}16, WBW^{+}18]. glauca [CJFS16, CMH^{+}11]. glider [TRSM15, ZDD^{+}19]. glider-mounted [ZDD^{+}19]. GLMM [BMR10b]. GLMM-based [BMR10b]. GLMs [HKS12]. global [AB16, ABRL12, Bak14, COL^{+}13, Fri19, Gar11, JPPY11, JC15, Kam14, KYK^{+}13, LSD^{+}11, MCOS17, PDWH11, PHK^{+}19, SiTiM^{+}11, SHST18, WJM16, Wel11]. Gmelin [Par13, POIM12]. go [CRAM^{+}14, GRH^{+}14, WHMP15]. goals [VF17]. goby [MNGA^{+}17]. goes [MHE^{+}16]. going [Bro14, BCC^{+}19]. golden [GSMRM^{+}16]. Goldilocks [Hil19a, Hil19b]. goldsinny [JQD^{+}17, OHLK19a, OHLK19b]. Gonad [ESR^{+}10]. gonadosomatic [FWD15]. Gonatidae [KSZ10]. Gonatopsis [KSZ10]. Gonatus [KSZ10]. gone [RB14]. Good [TGH^{+}12, Ano19, BSR^{+}10, DCMGB^{+}17, Job17, Mur10, PO14, RCS^{+}17, RDL^{+}17, TJD17]. Google [AAP14a, GGT^{+}14, AAP14b]. goose [POIM12, Par13]. gooseneck [RGGFA16]. Gordian [CPSCCE^{+}17]. gorgonian [CSBHS^{+}11]. governance [HG10, Rid18, STSP17]. governed [Cri12]. grab [GHG^{+}10]. graceful [Hol14]. Gracilechinus [GIPS12]. gradient [MSS^{+}19, Sub18a, Sub18b, VCH16]. gradient-based [Sub18a, Sub18b]. gradients [FGMC^{+}13, HTRF17, HGRAR10, LHS^{+}19, RRCT^{+}17]. graduate [CHB^{+}14]. Grand [NPGT15, BWP10, MMAS11]. grande [Pla16]. grandis [VSB^{+}16]. graphic [LKG^{+}19]. grass [LAD^{+}17]. gravity [Dav11, MRC^{+}10]. grazing [HSS^{+}16]. Great [JMHS14, PRO14, Bro14, CFT^{+}19, ODSC10, Day18c, HBB^{+}13a, LBB15, PEV^{+}16]. Greater [DSA13]. greatly [DGPR11, HBB^{+}13a, SGV18]. Greek [KMK10]. green [FLCQ19, GQX^{+}19, MQ11]. Greenland [BHS^{+}16, MMDSP15, RHB^{+}13, RKB^{+}12, SLKS10, AV15, DPOL13, FBCD^{+}17, HGRAR10, JW18, JBE14, LHJ12, MS10, MGR^{+}13, ØHN14, RHB^{+}12, SHH^{+}17, YFG^{+}17]. grenadier [KRML11, NST^{+}14]. grey [BMR^{+}11, FHH^{+}13a, FHH^{+}13b, GÖN^{+}12, Ger17, HSSB14, LDT^{+}11, LHLK10, ØFNH12, RB10, SRH^{+}14, TDHC18, VRF^{+}16]. grey-headed [RB10]. grid [GW16, SHG^{+}18]. groenlandicus [FFH^{+}11, SS10b, SBKA16]. gross [DFB^{+}10]. Ground [AAP14a, CSFS10, DHZA14, TSP^{+}19]. ground-truth [AAP14a]. Ground-truthing [AAP14a, CSFS10]. groundfish [FB11, FLP^{+}13, GMC17, LGB^{+}18, ML12, PMH^{+}13, RHSH16, RBCH10, RKJ14, SA14, SG16a, SHS11, SKKM15, TPL11, WKL13]. groundfishes [TSWS15]. groundlines [BT10]. grounds [BKMdMS13, CPVS14, DMZC17, GCG^{+}10, GIW16, HSK^{+}14, JL12b, KHH^{+}17a, NGTT16, PSH^{+}11, PSDG12, SLH^{+}17, SJRB18, SMB^{+}16b]. group [APV18, BAFRJ18, EISJ12, HMM12]. grouper [LGR^{+}19, WNM^{+}13]. groups [LPB^{+}14, RGRL11, SEDO19]. grow [WRMJ13]. Growth [HS17, LKR16, MDdPMQ10, MKAR^{+}18, AEC11, APB^{+}10, BMGB19, BMR10a, BMR10b, BWH16, CC17, CGAD16, CK18, CPP17, CPLH16, DLL12, DMBD10, EWH16, FMC10, FJSJ15, GMH^{+}14, GMC17, GM17, GMM^{+}12, GOPG10, GS14, GFML10, GOS^{+}13, GES^{+}12, HB19a, HGRAR10, HSNO12, HJS14, HCL^{+}18, HP11, HHT13, HFM13, HCF^{+}12b, JÓT^{+}12, KAT11, KLCC18, KDBP11, LCSI17, LTGP18, LSI^{+}15, LVSF17, LK15, LAD^{+}17, MPR12, MSS^{+}19, NLB^{+}12, PRB^{+}15, RPDF14, RCF^{+}17, SiTiM^{+}11, SLF^{+}16a, SJ15, TWHB19, THC16, VWMS19, WOW^{+}17, WPH^{+}17, WPvBS15, WMN^{+}11, ZSC16, ZGT13, vGA18b, vdSSM^{+}18]. grypus [FHH^{+}13a, BMR^{+}11, FHH^{+}13b, GÖN^{+}12, LDT^{+}11, LHLK10, ØFNH12, VRF^{+}16]. Guidance [LKS^{+}14c]. guide [MGVH^{+}10]. Guidelines [SBB19, EP14]. guiding [MTLG14]. guinea [KM16, HTRF17]. guitarfish [CPSCCE^{+}17]. Gujarat [BCKA11]. Gulf [BTA^{+}18, DHZA14, LBTS^{+}19, MGS^{+}15, RRH^{+}15, RHP^{+}12, ALWH18, BMR^{+}11, BDW18, Ben13, BNP19, BBB^{+}19, BCOS11, CTC14, CMH^{+}16, CCD^{+}19, CKD^{+}17, CCSG18, DEH^{+}19, Fau11, FBD^{+}19, GPP17, GCL^{+}18, GCS^{+}17, HRB15a, JFJ^{+}17, Kan11, LKR16, LFGW13, LFM^{+}18, LSWD12, Lev11, LYW^{+}18, LCG^{+}18, LGR^{+}19, LBK^{+}19, Llo17, MO18, MPL^{+}14, MDdPMQ10, MLLL11, NBS^{+}10, OSL^{+}15, OSBG16, OSM18, PMM^{+}14, RLQ^{+}15, RKL^{+}10, RGC^{+}19, RRH^{+}15, SBG10, SLT^{+}14, SG10, SRRZ16, SMK15, TCBD10, TCS^{+}19, TCQ^{+}15, WHAA^{+}18, WMJ13, ZGK^{+}17, ZHD^{+}14, ZLC^{+}17, ZCW11]. gull [MRC^{+}18a, MRC^{+}18b]. gulls [TSBV^{+}15]. gulper [DHS19]. gun [PHO13]. gut [XCRP13]. guttatus [BFDGLÁ15]. GWR [WRDF10]. Gyre [PPHK19, SRHJ12].

Habitat [AWS^{+}13, YCZY19, ASI^{+}16a, ASI^{+}16b, Ano19, BPF^{+}18a, BPF^{+}18b, BPBD15, BMF14, BVW^{+}18, BFH13, Cad14, CSS^{+}15, CSC^{+}13, CCL^{+}14, CMD17, CSD16, DCMGB^{+}17, EISJ12, Eri16, EPH15, FFK^{+}11, FBM^{+}18, GHD^{+}10, GGV^{+}18, GT19, HBB^{+}13a, HR14, HLS^{+}15, HHH12b, JLH12, JRWM11, LVM^{+}11, MMWC16, MVB^{+}18, MQ11, MSOM10, MS15, MS19, MLLL11, MBL^{+}17, NRTK^{+}19, OFVF14, PGV^{+}17, PWPW18, PMM^{+}14, PCS^{+}11, RBA^{+}19, SHJI10, SLS18, SDBB15, SBR11, SH14b, SPM^{+}17, SCL^{+}19, SBSE14, TCBD10, TB17, TMR^{+}16, VELT14, VCH16, WHAA^{+}18, WMJ13, YYCC16, ZHD10, ZOQS10, ZED11, dJBLH15, vdKKS11]. habitat-specific [VELT14]. habitats [CBD^{+}10a, CMP^{+}11, CBO15, FJSJ15, GFSC17, HGGJ15, HCDAF15, KJW^{+}18, KE14, LFM^{+}18, MAP^{+}11, dSMGKP12, PWCS17, PCLQ^{+}16, PDMG11, RDS12, RDCPvH11, SWB^{+}14, Sei14, SMS^{+}17, SSP^{+}13, TDN^{+}19, TSY^{+}11]. habits [CST^{+}19, GS11, Mag11, MUW^{+}19]. Habituation [TGDG15]. haddock [BGM^{+}14, CPVS14, LSY^{+}14, LF14, OAM^{+}10]. Hake [MTDP11, CCC^{+}14, CLV^{+}14, CKV^{+}16, DCG10, DG11, DMBD10, FCPJ10, FMLM^{+}14, FWD15, GES^{+}12, HSSB14, IEL^{+}15, JCA10, LÁG^{+}19, MDdPMQ10, SSGH18, TDN^{+}19, WSG^{+}17, dPJGB13, dlTQMUE10]. Half [Mäl12]. halibut [AV15, DPOL13, FBCD^{+}17, JBE14, KMF13, LFM^{+}18, LBK^{+}19, Loh11, MS10, MGR^{+}13, NS17, OIMP18, SMNE14, SFNdH16, TB16]. Halichoerus [FHH^{+}13a, BMR^{+}11, FHH^{+}13b, GÖN^{+}12, LDT^{+}11, LHLK10, ØFNH12, VRF^{+}16]. Haliotis [BLM^{+}17, HSB18]. hand [CJP^{+}14]. hand-jig [CJP^{+}14]. handle [dMBD11]. hands [Dek16]. Hansson [HLL18]. hapuku [WNM10]. harassment [TGDG15]. Harbour [Alf10, DTA^{+}13, STB^{+}11, TRH10, VG15, WDE^{+}18, PSKM^{+}16]. hard [BWH16, PHM13]. hard-to-age [PHM13]. Harden [Arn11]. harengus [LHK18, OTAK15, RHV14, AFLvDH15, GKR10, GOS^{+}13, HNAK12, HP11, KWF11, LGR^{+}14, MSK^{+}11, OSJ^{+}16, PHM^{+}14, PKHM12, PKHG14, RHOJ10, RTJS14, SOK^{+}12, SKA^{+}12]. Harmful [TTJ^{+}18, TTJ^{+}19]. harmonizing [LÁG^{+}19]. harp [FFH^{+}11, HS10, HS17, LLCJ10, ØHN14, SS10b, SBKA16]. harpacticoid [SSH^{+}17]. Harvest [Eid18, BCN16, BPvHR10, CCS10, DFG^{+}17, GRC16b, HMHL19, JRG^{+}16, KGZH16, KCMS19, LWT^{+}11, OFVF14, PRF^{+}17, PDD^{+}11, SDH12, SLKS10, SSH^{+}14, SSP12, ZCW11]. harvestable [HDF^{+}19]. harvested [HGF^{+}18]. harvester [POIM12, Par13]. Harvesting [HAH^{+}14, JDM^{+}16, PDCP14, BPBD15, CGR16, FWP^{+}16a, GRC16a, HSV^{+}17, HHBSM16, JCSC10, MCO^{+}17, MPBH11, PH13, Pla17, RGS^{+}16, SMBH16]. harvests [BPD^{+}18a, BPD^{+}18b]. hatch [HFM13]. hatched [WRMJ13]. hatching [PDD^{+}10a, RHHH^{+}12, SLF16b, VHS^{+}10, VKS^{+}12]. haul [HMS^{+}13, LPT10]. haul-back [HMS^{+}13]. haul-by-haul [LPT10]. haunt [HK14]. Having [Pau16]. Hawaii [STD11]. Hawaiian [BJR17]. Hazard [DCPTN14]. hazardous [EHG^{+}14]. headed [RB10]. headlights [SPS15]. Health [BMR^{+}11, BPF19c, BP13, EHG^{+}14, GSMRM^{+}16]. healthier [PBE^{+}19]. Heat [SM12, SSA^{+}18, TIS^{+}19]. heating [ET15]. heatwave [HSB18]. heavily [GMH^{+}14]. height [PGS^{+}19a, PGS^{+}19b]. Heikinheimo [HKB^{+}18]. held [WRC^{+}12]. helicina [TRHB13, WHL^{+}17]. hell [Pau16]. Heller [KdS14]. help [CCC^{+}14, KGP^{+}15, PWPW18]. Hematodinium [CR11]. Hemiscyllium [HWR^{+}16]. hemisphere [MFT16]. Henry [HBF14]. herding [HRB15a]. Heritage [MS12]. hermaphrodite [DMF13]. hermaphrodites [MSA19]. hermit [KTLB16]. herpesvirus [PGS^{+}19a, PGS^{+}19b]. Herring [SS17, TMR^{+}16, AFLvDH15, BGS^{+}14, BHL^{+}15, BDCvD^{+}10, Bow14, BPF19c, CCS10, Cor13, Der18, DCNB^{+}10, GÖN^{+}12, GNDC11, GKR10, GOS^{+}13, HG10, HNAK12, HP11, KWF11, LHK18, LFMJ19a, LFMJ19b, LGR^{+}14, LNH^{+}15, MO18, MSK^{+}11, MGBvD14, OTAK15, OLR^{+}18, OSJ^{+}16, OAM^{+}10, Ósk18, PNS18, PHM^{+}14, Pay10, PKHM12, PHO13, PKHG14, RKL^{+}10, RHV14, RHOJ10, RDCPvH11, RTJS14, SOK^{+}12, SBFC10, SSF^{+}18a, SKA^{+}12, SKA15, SKA22, SBLS15, SWS13, TVO12, TMR^{+}17, VHS^{+}10, VKS^{+}12, VBO^{+}15]. heterogeneity [SGK^{+}11, THC16]. heterogeneous [GHPH15]. heteroscedasticity [HMR15]. heterotrophic [DCMC19]. heterozygosity [GMKS11]. heuristic [MFP^{+}19]. Hide [LAD^{+}19]. hides [SWS13]. Hierarchical [CC17, DMT^{+}10a, MSOM10, BMGB19, JOSO16, MGPC^{+}14, NSO^{+}19, SBCF16, SHS11, SR16]. hierarchy [STR16]. High [GHF^{+}11, KPGH16, PRF^{+}17, TRSM15, AGH^{+}11, BAFRJ18, CSF^{+}17, DSL14, DVH^{+}18, DKPM16, FWS^{+}13, GL11, GM17, HB19a, HDF^{+}19, HHM^{+}16, HBG^{+}16, IHL19, LSJ10, LSF16, LEE17, MVK18, ME11, ROBC11, RCH^{+}19, SMOH18, SJ18, SMSR10, SCJI10, SMS^{+}17, SFD^{+}16, SY16, STB^{+}11, SLF16b, TBD^{+}18, TVO12, VSB^{+}16, WHL^{+}17, EPH15]. high- [SMSR10]. high-density [STB^{+}11]. High-latitude [GHF^{+}11, ME11]. High-resolution [TRSM15, LSJ10, ROBC11, RCH^{+}19, SCJI10]. higher [GvRKB19, WNM^{+}13]. highlights [HCC^{+}14]. Highly [MPR17, CD11, DAB^{+}12, LSCL18, MBLS15, YPN^{+}18]. hikes [BM12]. Hilborn [SG18a, Hil18c]. hindcasted [MW19]. hippoglossoides [DPOL13, DPRG^{+}18, JBE14]. Hippoglossus [KMF13, LFM^{+}18, Loh11, SMNE14]. hippos [PPM13]. hippurus [MAA^{+}19a, MAA^{+}19b]. Histioteuthis [HLD10]. histograms [PDMG11]. Historical [DMO^{+}15a, MAA^{+}19a, MAA^{+}19b, MWJ11, Pin17, PJDN12, ATLC16, BJMG19, BM16, DB16, ETM^{+}16, JHB^{+}12, MAS19, MSR^{+}12, PHB^{+}16, RHB^{+}12, RHB^{+}13, TNH^{+}10, TMG16]. histories [GCJL16, YDT14]. History [Lor11, Abl16, AL10b, BS14, CPVS14, CK18, DL12, ETM^{+}16, GSMRM^{+}16, GRNG^{+}10, Ham15, HTB^{+}16, HOS^{+}15, HP11, KDF^{+}19a, KRO^{+}16, LGR^{+}19, LSF16, MMO16, MWS^{+}17, MWP^{+}16, NUO^{+}10, OHLK19a, OHLK19b, OdS19, OBY^{+}14, PGV^{+}17, PHV^{+}15, RC12, RPDF14, RTS19, SMR12, SG10, SG16a, SLF^{+}16a, SLF16b, VFR^{+}16, VG15, WNM^{+}13, Wri14]. hitching [RLW^{+}15]. Hjort [Sch14, AB14, Bro14, Hol14, Hub14, JMHS14, KOK^{+}14]. hockey [MR10]. Hokkaido [LSR^{+}13, RSY11]. holoplanktonic [GCL^{+}18]. Holt [Pla17, PHV^{+}15]. Homarus [CPB19a, CPB19b, EKD^{+}15, RCF^{+}18, THSM^{+}18, WWMF17, ZCW11, FFN^{+}13, FFN^{+}16]. home [ALMFFBP19, TPBL17, WÓG^{+}16, WMG^{+}17]. homing [FL14, GSMA15, TPBL17]. Honeycomb [Nee15]. Hood [PSS11, TBC^{+}14]. hooded [AWS^{+}13]. hook [GPP17, HCM^{+}15, MAM^{+}15, SKKM15]. hookeri [HB19a]. hooking [CJFS16]. hooks [BBOE10, MBBC11]. Hordyk [FWC^{+}19b]. Horizontal [AMQRC10, SHAM10, SQ14, VBW^{+}18, HLS^{+}15, STSP17]. horizontally [CK12, VPH18]. horse [DDR10, GMN^{+}17, GKV^{+}15, VBA10, vDTF^{+}14]. hospite [BIKP^{+}17]. host [BHCSP^{+}11, BHM^{+}18, GMD^{+}11]. hotspot [KCP^{+}11]. hotspots [KYK^{+}13, ML12]. hub [AHC15]. Hudson [GSP12]. Human [MBC11, ALMFFBP19, BPMR17, BVW^{+}18, GBJ^{+}15, GF14, JIB^{+}12, JIB^{+}14, KJW^{+}18, KPJ^{+}15, MMO16, MVH^{+}11, SLG^{+}15]. human-altered [ALMFFBP19]. human-caused [BPMR17]. Humans [LTS^{+}17, HBB^{+}18, HKB^{+}18, HLL18, Pur18]. Humboldt [QCA^{+}18a, QCA^{+}18b, JBSD11]. humpback [PWPW18]. hunt [BMR^{+}11, HS10, LLCJ10]. hurricanes [PEN^{+}19b]. hybrids [HFH10]. hydraulic [TJR^{+}10]. hydroacoustic [GHG^{+}10]. Hydrobiological [Can16]. hydroclimatic [BR12]. hydrodynamic [WCT^{+}16]. hydrodynamics [HHH12a]. Hydrographic [HHH12b, VAP12]. Hydrography [JV12a, CDAN^{+}14, HHH12a, KGW^{+}12, VHS^{+}11]. hydrological [LGBA11]. hydrophone [GZS^{+}19]. hydrothermal [BCL^{+}11, CBC^{+}11b]. Hydrozoa [BHCSP^{+}11]. hyperdepletion [RU12]. Hyperiidae [PCFR13]. hyperstability [RU12]. hyperstability-hyperdepletion [RU12]. Hyporhamphus [BBH^{+}10]. Hyporthodus [WNM^{+}13]. hypotheses [Elv15, Har14, JOF14, LVPK10]. Hypothesis [HCE^{+}11, LBJ15, Fra15, JFJ^{+}17, Lév15, MR14b, OCR14, SJB15]. hypoxic [GRF^{+}16].

Iberian [BSW^{+}11a, BSW^{+}11b, GMN^{+}17, JCA10, SGQR^{+}12]. ice [BGM^{+}19, BA12, DMS18, GSP12, KT18, PH12a, SH14b, SBKA16, SEDO19]. ice-dependent [SBKA16]. icefish [FFF16]. Iceland [GS12, GPG14, KJKÓ15, KJÓK16, PB11, PGG^{+}12, PSDG12, PFPG12]. Icelandic [AJ12, AVGÓ12, BSP15, JV12b, OEG^{+}16, Ósk18, PSH^{+}11, PSDG12, SHS11, TJMC11, VAP12, WHMP15]. ICES [BF17, CMP^{+}11, GGK^{+}11, MAP^{+}11, Man17a, PGN^{+}11, BCRM^{+}18, BMNKA^{+}16, CNBK11, DSN16, DBM^{+}15, DGB^{+}15, ETM^{+}16, FCPJ10, GMN^{+}17, HHGtIWGoOH12, HRP^{+}16, LKS14a, Mäl12, RGRL11, WM14, WRC^{+}12]. ICES/NAFO [HHGtIWGoOH12, WRC^{+}12]. ICES40G2 [SWS13]. Ichkeul [DHKE16]. ichthyoplankton [MFS^{+}14, SDAM^{+}18]. iconic [JRG^{+}16]. icons [BFLÁ15]. ideas [Rot15]. Identification [GCS^{+}18, AHR^{+}19, BCSG13, CSS^{+}12, GMD^{+}11, GCW^{+}18, GDJ11a, GHB^{+}10, HKRM18, HBF14, Kor10, MSR^{+}19, OdJN^{+}12, OVW^{+}17, SPFM^{+}10, SMS^{+}16, SFN^{+}19, ZMKC14, ZGTL13]. identify [CPVS14, LPB^{+}14, PWPW18, PBG14]. Identifying [ALB^{+}19, BGS^{+}14, KYK^{+}13, MPB11, PMP^{+}16, SG16b, ZDD^{+}19, JMM^{+}15b, RGPG10, TLR^{+}17, XPC11]. identity [vPBLR19]. IEA [HKS17]. if [HCR13, JMM^{+}15a]. IFQ [WHP16]. IFRAME [ZHLK11]. ignore [Cad14]. Ignoring [KBH^{+}18, TLPS15]. II [LSF16, SDAM^{+}18]. Illex [CAOG15]. illuminate [KH19]. illumination [LGB^{+}18, WABZ16]. illusion [FLP^{+}13, Pla16]. illustrated [BG18, WCR^{+}19]. Image [DMS16, CSD16, FFH^{+}11, FHH^{+}13a, FHH^{+}13b, HBF14, MBQ10]. Image-based [DMS16, CSD16, FFH^{+}11, FHH^{+}13a, FHH^{+}13b, MBQ10]. imagery [SMMK11, SCJI10, WJM16]. imaging [CGK^{+}17, FOJ13, KGP^{+}15]. immature [VOM11, Yar10]. Immediate [LM10, ODSC10]. immersed [BMVPN10]. Impact [BMF14, CMU^{+}17, JBE14, KGN^{+}10, KYK^{+}12, SPdJFMS17, SKA^{+}12, SH14a, VSB^{+}16, WAH^{+}16, AB14, AIA^{+}15, ACCA19, BOP^{+}19, BG18, CJC^{+}15, CRC^{+}16, CGGH^{+}17, DID^{+}16, DPL10, DAB^{+}12, ESAV15, EHD^{+}15, GW16, HP17, JPPY11, JCA10, JIB^{+}12, JIB^{+}14, Jok16, JCS16, KKC19, KKAV^{+}19, LMPP10, MALM17, OI16, PCFR13, PvdKE^{+}12, PSGY^{+}12, RBB^{+}16, RBP15, RMKG11, RCC^{+}14b, SRHJ12, SDDA11, SBKA16, TKB^{+}15, VE13, WPeEA16, WFM^{+}15, YFG^{+}17, vDHvKR15]. impacted [GVA19, GML13, WGCL^{+}19, ZMF12]. impacting [KHH^{+}17b]. Impacts [CTC14, CGAD16, DDF^{+}12, FBD^{+}19, KCL18, PNC15, RFHG19, SLS18, ASB^{+}11, ASI^{+}16b, AFK15, BNB19, BVW^{+}18, BMNKA^{+}16, BHM13, BLB^{+}12, CAL14, CSC^{+}13, CPS13, CAS^{+}16, CCG14, CKV^{+}16, CD17, CJP^{+}14, DCS^{+}11, ECFL15, EISJ12, GG13, GIDP^{+}18, GF14, HTRF17, HBiI^{+}11, HBB^{+}13c, HF14, Hum17, HCE^{+}11, IS15, JLH12, KJW^{+}18, KGP^{+}15, LJH^{+}12, LGL^{+}12, LSR^{+}13, MSA19, MFT16, NDM^{+}16, PBC^{+}16, PDR19, PPS17, RJP^{+}17, RKB^{+}12, SMS^{+}17, SSP^{+}11, SIP18, TED16, TWB^{+}19, TTG^{+}15, TPL11, TDHC18, TMG16, VMAMAG11, ZHLK11]. impairing [RCS^{+}17]. impairment [CPT^{+}10, MSM17, TBG16, UTA^{+}16]. impeding [RMJF14]. implanted [BDHC19]. implement [JCA10]. implementation [CKA^{+}17, DSL14, LKS^{+}14c, PCdL15, SGP^{+}15, SKKM15]. Implementing [HKS17, MTLG14, MLB^{+}14, RGS^{+}16, SSH^{+}14, SFA^{+}17, CCG14, LB17, VQB^{+}11]. implication [vDHvKR15]. Implications [EHHH14, FSN^{+}11, HHRW12, LJH^{+}12, RvDW10, TED16, vGA18b, BCS^{+}17, BSMB18, CAMM12, CAL14, CSTJ19, DMJ^{+}14, Dav11, DKC12, DCHMHQ12, DTA^{+}13, FWS^{+}13, GMS^{+}18, GMD^{+}11, GRC16a, GRS^{+}14, GOPG10, HSD^{+}16, HLCC16, HRC^{+}13, HMFW14, HFNH13, Hum17, HHE^{+}16, HHHE16, JBSD11, Lev11, MWJ11, MRC^{+}10, MS19, MAS19, OFVF14, PGOM11, PMH^{+}13, PDAC^{+}16, PVCB17, PDCP14, RGS^{+}16, RRM^{+}12, REFJ12, SHW^{+}15, SG10, SJRB18, SHZ^{+}16, SCC^{+}13, SSZH12, TIS^{+}19, TPHdJ19, URE^{+}18, WNM10, ZHD^{+}14, dPLF^{+}19]. Importance [BHS^{+}11b, HMR15, BGS^{+}14, BNB^{+}14, BG18, CSF^{+}17, DEH^{+}19, EAOB15, HRB^{+}15b, HJS14, KDH^{+}10, LSF^{+}15, MMRG18, NC12, OLM^{+}15, RSB^{+}15, RFF^{+}14, RDS12, SMK15]. important [BVW^{+}18, CSTJ19, PWPW18, RCG^{+}18, RCC^{+}14b, SWB^{+}14, SRRZ16, SMB^{+}16a, WPGW12, XPC11]. imprinted [OAT^{+}12]. improve [AGSPH14, Fri19, Mor18, Sca18, TSWS15]. Improved [RK16, WLL^{+}13, RCC14a, TSP^{+}19]. Improvement [LSR^{+}13]. improvements [KMF13]. improves [AAP14a, AAP14b, BSF^{+}19, GGT^{+}14, OHD^{+}16, RHV14, dBWMD^{+}14]. Improving [AMS^{+}12, CGG^{+}11, DPD^{+}15, OSC^{+}17, SMD^{+}16, PSS11]. improvisus [EWH16, PSH13]. in-season [BHC^{+}15, RU12]. inadequate [KK19]. inadvertent [JQD^{+}17]. incentive [AGSPH14]. incentives [ACS15, EPVC14]. incentivising [KRG^{+}12]. incentivize [HSHÖ14]. incidence [CD14]. Incidental [ACA19, FBCD^{+}17, FCCLA10, TMR^{+}16]. include [TWHB19]. including [CAB^{+}10, GSL^{+}11]. Inclusion [LHP^{+}19, MKL^{+}19, RDT^{+}17]. income [PLCC18]. incomplete [AH19]. Inconsistencies [MSK^{+}11]. inconspicua [CPLH16]. incorporate [TCS^{+}19]. Incorporating [ATLC16, DED13, RGGFA16, SGV18, WÓG^{+}16, WMG^{+}17, YS15, FCPJ10, LTGP18]. increase [EWH16, HBB^{+}13a, MKB^{+}17, PCRD18, TSY^{+}11, ZBE^{+}15]. Increased [KKH14, BTB^{+}17, BWM^{+}16, DRVV^{+}17, KS17, NKSE14, SLF^{+}16a, SLF17]. increases [BMC^{+}18, BBL^{+}15, GIR^{+}19, HFGT18, HUPBL18, KTLB16, KKH14, SGV18]. Increasing [WJM16, AGMC14, BKVT16, TH15, YS15]. increment [DWPS11, HRF^{+}10, WLN^{+}13]. increments [ZGT13, vdSSM^{+}18]. independent [HS17, HRB^{+}15b, LYW^{+}18, PCLQ^{+}16, TCBD10]. indeterminate [GMN^{+}17]. Index [GIDP^{+}18, CPT^{+}10, FWD15, HMR15, OIMP18, WDG^{+}17, TCF^{+}13]. India [BCKA11]. Indian [FFN^{+}13, FFN^{+}16, HGF^{+}18, ACC^{+}12, CLMP16, DMJ^{+}14, EGC^{+}17, GRG^{+}10, KCA^{+}14, MKB^{+}17, SPM^{+}18, ZGTL13]. indicate [CST^{+}19, KSZ10]. indicative [LGR^{+}14]. Indicator [FFRR12, GRR^{+}12, PKK13, SWS13, ASV^{+}19, BMC^{+}13, Kam14, MRD19, MVP^{+}14, SRG11]. Indicator-based [PKK13]. Indicators [PH12b, BHB^{+}19, BCT^{+}10, BSB^{+}19, CCC^{+}14, Eid18, FLLG15, GRF^{+}12, JdFBB^{+}10, LFFL13, LYS^{+}10, LLSJ18, MPB11, MNWB18, OLW^{+}14, OFB^{+}17, PvOP10, PSF12, PO14, RSV^{+}17, RFF^{+}14, SCY^{+}10, SGP^{+}15, SBS^{+}10, SS10a, SSB^{+}10a, TLR^{+}17, Wri14]. indices [EDP14, MONS14, MSK^{+}11, OYIN18, OFB^{+}17, OBL12, RDB^{+}18a, RBCH10, RHOJ10, RPDF14, TSP12, TSWS15]. indigenous [DHNL12, FLCQ19, WLB11]. indirect [AR10, DMJ^{+}14, FLCQ19, SEDO19, Win15]. indiscriminate [APV18]. IndiSeas [LLSJ18, SS10a]. indispensable [DVV^{+}11]. Individual [APB^{+}10, CSTJ19, PBQ^{+}10, RPDF14, TSBV^{+}15, ALMFFBP19, BGM^{+}11, BPKH13, CJC^{+}15, DKPM16, HP11, JKS^{+}18, KHF14, KTLB16, MR14a, PVD14, ROK^{+}18, SMB^{+}18, SFW^{+}12, TLF16, TPL11]. individual- [PVD14]. individual-based [HP11, MR14a, SFW^{+}12, TLF16]. individuals [DKK15]. Indo [GFML10, PWPW18, VF17, YJBL17]. Indo-Pacific [GFML10, PWPW18, VF17]. Indo-West [YJBL17]. Indonesia [AHC15, JHB^{+}17]. induced [GQX^{+}19, HBB^{+}13b, Hje14, HK14, IZP^{+}16, LBG18, LNWS18, MVH^{+}11, RvDW10]. induction [dBWMD^{+}14]. industry [Fau11, LPT10, LÓGR11, NJFH19, SOL^{+}15]. inermis [HUPBL18]. infauna [RDD^{+}10]. Infaunal [CGM^{+}19]. infected [CZBFLÁ^{+}15, MWS^{+}17]. Infection [HdBR^{+}16, Mac17b]. infections [CR11]. infectious [RCC^{+}14b]. infectivity [BSMB18]. infer [BWR^{+}15]. inference [HKS12, SS12, URE^{+}18]. inferences [SDH^{+}18]. inferred [BCKA11, CFM^{+}14, CCD^{+}19, NNY16, SFN^{+}19]. Inferring [BTTV^{+}17, JRWM11, VQGDAMdL14, CMG^{+}17]. infestation [GKC^{+}16, SAS^{+}18]. infestations [VHF^{+}17]. inflation [NNY17]. inflow [MGBvD14]. Influence [BKSB12, EMP11, HDF^{+}19, JTJ18, MNV^{+}11, OFVF14, TSY^{+}11, BvdMA^{+}18, BFDGLÁ15, Cri12, FBM^{+}18, GCS^{+}17, HHMF19, HG10, HHH12b, KFBG15, LF14, MPL^{+}14, MOVU10, ODL16, OCG13, OCWG16, OS14, OFYS^{+}11, RC12, RAB^{+}12, SLM^{+}11, SDH^{+}18, SCL^{+}19, TFC^{+}19a, TFC^{+}19b, VCRPS13]. influenced [SBS19a, SBS19b]. influences [BWH16, GF11, HS17, HTM11, Kam14, LVM^{+}11, LWS^{+}18, LBJ15, STD11]. influencing [DKB14, EPVC14, JTM^{+}16, LEB^{+}14, RGFT14]. inform [CAM^{+}15, FCB17, HBC^{+}11, JLH12, MHM^{+}11, SHL^{+}14, SM15]. Information [RBCH10, BVC15, CECL16, DGB^{+}15, GOS^{+}13, HG10, HHA^{+}11, HRB^{+}15b, IFU11, JCA10, JRWM11, MKL^{+}19, Mat11, MHM^{+}11, RGRL11, RTS19, SSMD^{+}13, THHH15, THHH18, ZBE^{+}15, GGK^{+}11]. informing [GTBT18, PHV^{+}15]. ingested [BBOE10]. ingesting [MBBC11]. ingredients [SKKM15]. inhabits [GRG^{+}10]. inherent [KCMS19, SCL^{+}19]. inhibit [STSP17]. initial [BMR10b]. initialization [EAOB15]. initiation [CHMY15, SMB^{+}18, TRSM15]. initiative [DBM^{+}15, HBC^{+}11, JMM^{+}15a]. injuries [dHFF^{+}16]. Injury [UTA^{+}16, ODSC10]. inland [NC12, Wel11]. inner [NST^{+}14]. innovative [GM18]. inorganic [BMM^{+}19]. input [PCdL15]. insects [CTT14]. inshore [APK11, BVC15, ZHD10]. insight [MJSB14]. insights [BSK^{+}11, CCA^{+}18, HP11, JQD^{+}17, JWS^{+}18, TFC^{+}19a, TFC^{+}19b, dPJGB13, dPLF^{+}19]. Insonification [MBQ10]. inspections [ANU11]. inspiring [EP19]. instability [HTKB19, OYI17]. Installing [COL^{+}13]. institutional [EPVC14, RDT^{+}17, STSP17, SBB^{+}17]. institutions [DMS18]. instrument [Mos18]. insurance [RML^{+}16]. Integrated [MTCF^{+}17, ZMF12, Coc17, DGL^{+}17, DC14, GLC15, GM18, HPDB19, HKS17, HHQ12, HFSV^{+}15, LKS^{+}14c, MLS^{+}17, MGPC^{+}14, MLB^{+}14, MBT^{+}19, PDZ^{+}13, PA18, PHM13, SHL^{+}14, THW16, WM14]. Integrating [BSHE18, CDSP11, GL11, JGCH18, JOSO16, LB14, OIMP18, PDD^{+}10b, SBH^{+}14, SPP^{+}16, TBC15, WCR^{+}19, BJMG19, CJC^{+}15, HHZ^{+}18, LBCOJ19, Mur11, SBB^{+}17]. Integration [JL12a, LKHK11, DBA^{+}19, MSOF18, MHH^{+}18, Por11, RDKK15]. integrity [EBH^{+}17, PH12b]. intensification [LLST15]. intensity [Cor12, EBH^{+}17, MRC^{+}18a, MRC^{+}18b, PRF^{+}17]. intensively [dJBLH15]. interacting [FR16]. interaction [GF14, JWS^{+}18, KHH^{+}17a, MMAS11, PB14, SFD15]. Interactions [HLCC16, PNE^{+}14, ALB^{+}19, BNP19, CMM^{+}15, CLV^{+}14, DMJ^{+}14, EBSW15, FOT^{+}17, GLK^{+}10, GG13, GLP^{+}11, GRS^{+}14, GCL^{+}18, GJWS19, HFH10, HDG^{+}17, JIB^{+}12, JIB^{+}14, LAD^{+}19, LAG^{+}16, LLG18, LWS^{+}18, LBG18, LOOO16, MMO16, MKF^{+}15, PFPG12, PHH^{+}16, RBM15, RKCP18, STSP17, SPV^{+}16, TBC^{+}14, TPS^{+}11, WSGD^{+}17]. Interactive [WPvBS15]. Interannual [BNP19, FWT^{+}14, GDBM^{+}17, KAT11, KGW^{+}12, NS17, PBS14, PT19, STKT16, EPK^{+}18, HBD^{+}19a, HBD^{+}19b, ISL19, KRD^{+}18, SDAM^{+}18, TCQ^{+}15, TID^{+}10, VHS^{+}10, WHL^{+}17]. Intercalibration [TKJB19]. interconnected [GLC15]. interdisciplinarity [Omm18]. interdisciplinary [LKHK11, ZMKC14]. Interesting [Ful11]. interests [RSB^{+}15]. interference [CS15]. interior [THW16]. intermediate [OHLK19a, OHLK19b]. Intermittent [BHJ14]. intermixing [LBB15]. internal [AGG^{+}18]. International [DG17, LKG^{+}19, MLS^{+}17, MNG15, DUM^{+}12, Mos18, Ric14, Hub14]. Internet [Gar11]. Interpretation [RAH^{+}16]. interpreted [TJH15]. Interpreting [RFF^{+}14, RKW^{+}17, SRG11]. intersite [LLL^{+}13]. Interspecific [MMC19]. intertidal [BPBD15, MS19, NMF^{+}16, SSH^{+}17]. interval [KFBG15, RD16]. interview [GRS^{+}14, SSP12]. interview-based [SSP12]. Intra [KBB13]. Intra-patch [KBB13]. Intrastock [WMG11]. intrinsic [PCRD18]. introduced [AHC15, DD10, WHNS14]. Introduction [BMBE11, HAH^{+}14, PB14, Rod10, WRC^{+}12, BMNKA^{+}16, DHAH12, GGK^{+}11, HHRW12, Sei14, SPSP11]. introgression [KDF^{+}16]. invader [MQ11]. invariants [PHV^{+}15]. invasive [CPB19a, CPB19b, DD13, FPRA11, HdBR^{+}16, KHH^{+}17a, VF17]. inventory [GNB^{+}16]. inverse [AH19, MNV^{+}11]. invertebrate [BWM^{+}16, ECFL15, ÉSMGB15, JDM^{+}16, RGGFA16, SMH^{+}19]. invertebrates [HP17, HJT12, MGS^{+}15, NMM^{+}17, PG16, SCD^{+}15]. inverted [KBvZP16]. investigate [FB19, RGAS^{+}18]. Investigating [AHM^{+}15, CPB19a, CPB19b, MKHK18, PWPW18, RCH^{+}19, SPS^{+}10, YPN^{+}18, AUEC12, BCL^{+}11, ÉSMGB15, MKD^{+}18]. investigation [BCSG13, DUM^{+}12, DRB10, GNDC11, KCP^{+}11, LAD^{+}19, MVH^{+}11, SFD15]. investigations [EP19]. investment [THKP11]. involved [Jok16]. involvement [BLJ^{+}17]. involving [Tho11]. Ireland [BSK^{+}11, CR19, DMF13]. Irish [CRHM19, DL11, FLCQ19, GKC^{+}16, HGS11, LÓGR11, RCC^{+}14b, VGWJ11, VBO^{+}15, WKW^{+}10]. iron [SY16]. irradiance [DKPM16]. Irruptive [FLP^{+}13]. Ise [TMW19]. ish [BSR^{+}10]. ISIS [LMPP10]. Island [GTGD15, PG10, RBP15]. islandica [TJR^{+}10]. Islands [CAMM12, RWG^{+}18, ZHD10, BUNB10, DMO10, GRG^{+}10, JHB^{+}12, MNGA^{+}17, Pac18, RU12]. Isles [GNDC11, LDD^{+}10, MSE12]. isodon [PHB^{+}16]. isolated [RFN^{+}19]. isopod [TBG16]. isotope [DBDP10, GRNG^{+}10, JvdM15, KIA^{+}18, KSJ14, MUW^{+}19, PBW15, PDD^{+}10a, PFPG12, YDT14, YPN^{+}18]. isotopes [BCKA11, CCC^{+}14, DPD^{+}12, MTP^{+}12, OBLP^{+}19, RKW^{+}17, TVJ^{+}14, TMP12, dBWMD^{+}14]. Isozoanthus [CSBHS^{+}11]. Issue [Ano17, Ano14a, Ano15, Ano16a, Ano16b, BMNKA^{+}16, Rod10]. issues [Dig19, MP15, RG11]. Isurus [CJFS16, TSPL14]. Italy [DKC12, FSC^{+}12]. ITQs [MSC14, WP13, vH13]. IXa [FCPJ10, GMN^{+}17].

J [GPP17]. jack [PGV^{+}17, STKT16, VCRPS13]. Japan [KYK^{+}12, KGN^{+}10, LSR^{+}13, LSI^{+}15, MS12, RSY11, TNS13, TUUC14]. Japanese [HFHS10, IOK17, KIA^{+}18, KGN^{+}10, KKMS14, LSR^{+}13, LSI^{+}15, NNY16, RSY11, RYS11, STKT16, TIS^{+}19, TMW19]. japonica [HZZ^{+}15, LSR^{+}13, RSY11]. japonicus [STKT16, TKA^{+}16]. Jasus [GFMO11, HGGJ15, LMG^{+}14, MLMJ17]. jellyfish [CS15, EHT17, Eri16, GBB16, GHB^{+}10, KPJC14, LHP^{+}19, Pur18, QCA^{+}18a, QCA^{+}18b, RRH^{+}15]. jig [CJP^{+}14]. jigging [CRAPMN12, PG10]. Johan [Sch14, AB14, Hol14, Hub14, KOK^{+}14]. John [Ban12]. joint [WHP16]. Jones [Arn11]. Journal [BF17, CMP^{+}11, MAP^{+}11, Man17a]. journey [TDN^{+}19]. joyneri [DI11]. jumbo [FBD^{+}19, KAT11, YYCC16]. jurisdiction [Fri19, GM18, HD18, Mos18, Rid18, Sca18]. Justified [Aga18c]. Juvenile [CN16, AS14, AGG^{+}18, BCBI13, BGM^{+}14, BWR^{+}15, BGM^{+}19, BMC^{+}13, BMR10b, BMSC16, CHM15, CSTJ19, CDAN^{+}14, FJSJ15, GM17, GCL^{+}18, GBB15, HTT^{+}17, IKF17, Jon14, LCSI17, LGB^{+}18, LPSF19, LLPV^{+}12, LK15, LBK17, MHK19, MLC^{+}19, NDM^{+}16, OIMP18, PDAC^{+}16, RFN^{+}19, SDAM^{+}18, SBD^{+}15, TJMC11, VHH10]. juveniles [CMH^{+}11, PDCP14, VWMS19].

Kajikia [HLS^{+}15, KDBP11, MMRG18, SSP^{+}13, VBW^{+}18]. Katsuwanus [BMS^{+}18]. Kattegat [DTA^{+}13, HJS^{+}17, MV10, NLB^{+}12, PBS14, PBJ^{+}14]. keeper [BBHC11]. Keeping [LTS^{+}17]. kelp [COL^{+}13, HGGJ15, LSR^{+}13, RSY11, SMBH16]. kelts [MHE^{+}16]. key [CMA^{+}16, CFHC10, ECG15, ÉSMGB15, HTB^{+}16, NPB^{+}15]. Keys [BM15, AH19]. keystone [GRP10]. kHz [JLL17, Mel16, SOK^{+}12, SHL10]. Kii [YDT14]. kill [Moo14]. Killer [KKLM13, PMH^{+}13, GTGD15, PDS15b, TGDG15, TGG^{+}15, TTR^{+}19]. killing [Moo19]. king [FPRA11, GMKS11, Hje14, HKS12, KZS10, LVSF17, LPSF19, OLW^{+}14, PSGY^{+}12, SLF17, SMB^{+}16b, WHNS14, Win15]. kingfish [MWP^{+}16, RBBC11]. Kishinouye [KPJC14]. kisutch [CHD^{+}18]. knee [IHL19]. knee-high [IHL19]. knot [CPSCCE^{+}17]. Know [Man17b, GRH^{+}14]. Knowledge [DMZC17, BHB^{+}19, BSF^{+}19, BSMB18, CBL19, DCS^{+}11, FCB17, HB19b, Hin15, JW18, LGRC14, LKHK11, LAD^{+}11, Mat11, RMJF14, RCC14a, SPP^{+}16, TPV14, WFM^{+}15]. knowledge-based [WFM^{+}15]. known [FFH^{+}11, FHH^{+}13a, FHH^{+}13b, PCRD18, SV10]. known-age [FFH^{+}11, FHH^{+}13a, FHH^{+}13b]. Kola [PT19]. Kolbeinsey [JV12a]. Korean [COL^{+}13, KLN^{+}10]. Krill [BS13, MDD^{+}14, SSS^{+}12, BGM^{+}19, CK12, CK18, CPP17, CWRB11, FFF16, FWT^{+}14, HUPBL18, JLL17, KOC15, MPL^{+}14, PMM^{+}14, RT19, SRS11, STF^{+}16]. Kristiansen [FPLB19]. Krøyer [BSMB18]. kroyeri [BMSC16, KdS14]. Kuroshio [MUW^{+}19, YCY^{+}13]. Kuroshio/Oyashio [YCY^{+}13].

l [LHV^{+}16, MRD19]. L. [ACM^{+}16, AJM19, BCBI13, BvdMA^{+}18, BPPB12, GKC^{+}16, GCW^{+}18, GRC^{+}14, GOS^{+}13, JOvdM^{+}14, KMF13, LFGW13, LGR^{+}14, OAT^{+}12, OS14, OVW^{+}17, PSDG12, PHM^{+}14, PKHG14, RTJS14, SDH^{+}18, SCD^{+}15, UFJ^{+}18, WÓG^{+}16, WMG^{+}17]. labelled [SSM^{+}18]. LabHorta [CBC^{+}11b]. laboratory [CRC^{+}16]. Labrador [KR12]. labrax [BvdMA^{+}18, PPS17, SCC^{+}13]. Labridae [SDBB15]. Labrus [DMF13]. Lack [IZP^{+}16, KSJ14, MKAR^{+}18]. lacking [BPD^{+}18a, BPD^{+}18b]. lagged [PSF12]. Lagoon [MAB12, BMS^{+}19, CMC^{+}12, PKHG14]. lagoons [ACM^{+}16]. Lagrangian [GHD^{+}10, MRSG^{+}17]. Lake [DHKE16, LBB15, SR16, ZRN18]. Lakes [JMHS14, LBB15, PRO14]. lalandi [MWP^{+}16, RBBC11]. Laminaria [RSY11]. Lamna [BCS^{+}17, CRHM19, CJFS16, SRC11]. lance [FRF14]. land [BCL^{+}11, BVC15, KS14, TBC^{+}14]. land-based [BCL^{+}11, KS14]. landed [Ben13, SPFM^{+}10]. Landing [CEP^{+}18, PDR19, CLS^{+}18, GPS^{+}17, PFH16]. Landings [KMK10, BVS^{+}16, DKC12, DZ14, EMP11, FB11, JLH12, LPT10, MHMP^{+}18, SDH^{+}18, TH15, TCH^{+}16]. landscape [RGGFA16]. landscapes [BF17, Man16, Man17a]. language [RSB^{+}15]. lanternfishes [OBLP^{+}19]. lanternshark [HD11]. Lanzarote [CAMM12]. Large [FFRR12, GRR^{+}12, KLC13, KYK^{+}13, LMG^{+}14, LLF18, RB10, WSG^{+}17, AIA13, BSHC17, BPF19c, CP15, GKR10, HHR17, HSGL18, HLA^{+}11, LKS^{+}14b, MVP^{+}14, MWP^{+}16, NST^{+}14, OJK^{+}19, PBJ^{+}14, QCA^{+}18a, QCA^{+}18b, RKW^{+}17, RLW^{+}15, SBNL12, SRG11, SOB^{+}11, She15, TPBL17, TB16, WNM^{+}13, WBW^{+}18, SWS13]. large-bodied [TPBL17]. Large-scale [LMG^{+}14, AIA13, GKR10, OJK^{+}19, PBJ^{+}14, SBNL12]. largely [Cad14]. larger [GHPH15]. largest [SFN^{+}19]. Larus [TSBV^{+}15]. larva [PRB^{+}15]. larvae [BSHC17, BCSG13, DCHMHQ12, FAP^{+}17, HPN^{+}17, HLMT16, KPGH16, KRD^{+}18, LCSI18, LKS^{+}14b, LHV^{+}16, LSF16, MUW^{+}19, MDV^{+}15, MOVU10, OMTY14, PRB^{+}15, QR15, SDW19, STKT16, SHK17, SBLS15, SQ14, VWMS19, VHS^{+}10, VGB^{+}17]. Larval [PSH13, AFLvDH15, BvdMA^{+}18, BSMB18, BMG15, CMU^{+}17, CGAD16, CSB18, CKD^{+}17, DDF^{+}12, GHD^{+}10, GCL^{+}18, HP11, HFM13, HHHE16, KCF^{+}17, KLCC18, KCL18, KVSV14, LBM19, LBK17, LHH^{+}13, MRSG^{+}17, NMO^{+}17, ØU13, PHM^{+}14, Pep16, RHOJ10, RPDF14, RMJF14, Sas19, SBL^{+}17, SP14, VCRPS13, WWMF17, dBWMD^{+}14]. larval-stage [CKD^{+}17]. late [CMU^{+}17, DK18, Eri16, Pin17, vGA18b]. late-in-life [vGA18b]. late-larval [CMU^{+}17]. lateralization [SJ16]. latissima [NMF^{+}16]. latitude [GHF^{+}11, ME11, SMOH18, WPH^{+}17]. latitudes [HDF^{+}19, WNM^{+}13]. Latitudinal [TID^{+}10, HGRAR10, MSS^{+}19, SiTiM^{+}11, TNS13, WNM^{+}13, WPH^{+}15]. latus [GDS15]. Laurentian [JMHS14, PRO14]. laurettae [PBW15]. lavaretus [LHV^{+}16]. Law [GM18]. Lawrence [BCOS11, OSBG16, OSM18, VWMS19, Ben13, BBB^{+}19, CCD^{+}19, LFGW13, LFM^{+}18, LSWD12, MPL^{+}14, MGS^{+}15, PMM^{+}14, TCBD10]. layer [JPPY11, KRKG16, MMG^{+}17]. layers [Fra15, GCBI17, PHK^{+}19]. lead [PCLQ^{+}16]. leading [vGA18b]. leads [BDCvD^{+}10]. Leaper [HS10]. learn [DMS16]. learned [BKTS11, Cam18, DGL^{+}17, DCNB^{+}10, HRP^{+}16, KHC^{+}17, Mac12, MFP^{+}19, RGRL11, SHL^{+}14]. learning [KM16, NdlPIR18, Pet19, SR17]. lease [LGvPH15]. leasing [EHG^{+}14]. leatherback [GHB^{+}10]. led [HDG^{+}17, MKHK18, WLL^{+}13]. LEDs [NHL^{+}19]. Lee [KHMS19]. leeward [SPE14]. legacy [KOK^{+}14]. legal [DSL14, Pas14, TMW19]. Length [ASB^{+}19, WPWH11, AST^{+}15a, AH19, BGM^{+}14, BMGB19, DZC^{+}13, DMBD10, EISJ12, FWC^{+}18, FWC^{+}19a, FWC^{+}19b, GFMO11, GvRKB19, GIR^{+}19, HOV^{+}15, HOS^{+}15, HPCW19, HTS^{+}10, HHT13, HHQ11, KKMS14, MRD19, OSJ^{+}16, OLM^{+}15, PGOM11, PDMG11, PKR^{+}19, SHS^{+}17, SWvSR11, SMD^{+}16, THH18, ZGT13]. length-at-age [ZGT13]. Length-based [ASB^{+}19, HOV^{+}15, HHT13, MRD19, PKR^{+}19]. Length-selective [WPWH11]. length-structured [DMBD10]. Lengths [MSE12, UBvH11]. Lepeophtheirus [BSMB18, GKC^{+}16]. Lepidopsetta [CDAN^{+}14, CN16]. Lepidorhombus [MCAM14]. leptocephalus [MDV^{+}15]. less [CHM15, dMADLP^{+}16, PDD^{+}11, ZE17]. lesser [JRWM11, KO12, TSBV^{+}15]. Lessons [And15, DG11, DCNB^{+}10, KHC^{+}17, SHL^{+}14, Bow14, Cam18, CEP^{+}18, DGL^{+}17, Ger17, HKB16, HRP^{+}16, MG15, PH13, Pet19, RGRL11, SFA^{+}17, SJ18, SBR11, BKTS11]. Lethrinus [HGF^{+}18]. level [ASV^{+}19, CJC^{+}15, GFF^{+}17, GTM12, GZS^{+}19, GGP11, GRF^{+}12, Kam14, LGB^{+}18, MTP^{+}12, PVD14, SKA15, SKA22]. levels [BHS^{+}11b, BSB^{+}19, DVH^{+}18, FSC^{+}12, GKC^{+}16, HLMT16, JvdM15, LK17, MGVH^{+}10, MPBH11, MSR14, PSB16, RSV^{+}17, SFW^{+}12, SY16, TBD^{+}18, TH15, LB14]. LFI [SWS13]. libellula [PCFR13]. lice [BSMB18, BHM^{+}18, GKC^{+}16, MRMK11, SAS^{+}18, VHF^{+}17, VDK^{+}18]. lies [Har14]. Life [FFPL14, GRNG^{+}10, AL10b, BTB^{+}17, BS14, CYLT16, Can16, CPVS14, CK18, COHdP19, DL12, DPD^{+}12, ÉSMGB15, FBL^{+}16, GCJL16, GKV^{+}15, Ger17, GHM^{+}16, GOPG10, Ham15, HHH12a, HTB^{+}16, HOS^{+}15, HP11, KRO^{+}16, LGR^{+}19, LSF16, LEE17, MMO16, MGPC^{+}14, MR14a, MWP^{+}16, MFB17, NUO^{+}10, NKSE14, OHLK19a, OHLK19b, OBY^{+}14, PGV^{+}17, PKHM12, PSM^{+}10, PDCP14, PHV^{+}15, RTS19, RCF^{+}17, SMR12, SG10, SG16a, SiTiM^{+}11, SLF^{+}16a, SHS15, SLF16b, VRL^{+}14, VG15, WNM^{+}13, WM14, WHL^{+}17, Wri14, YDT14, vGA18b]. life-cycle [DPD^{+}12, PSM^{+}10]. Life-history [GRNG^{+}10, AL10b, DL12, HP11, NUO^{+}10, PHV^{+}15, RTS19, SMR12, SG10, SG16a]. light [HUPBL18, KYK^{+}13, NHL^{+}19, SPS15, WPvBS15, WLL^{+}13]. light-emitting [NHL^{+}19]. lights [MKHK18]. lightscape [VDK15]. like [TPBL17]. likelihood [HHQ11]. likely [HF14]. Limacina [BLMW17, TRHB13, WHL^{+}17]. Limanda [BWP10, vdRMC^{+}17]. limbs [BFDGLÁ15]. limit [Eri16]. limitation [LBG18, RFF^{+}14]. limitations [KKC19, PFH16, RFMR10]. Limited [BWM^{+}16, ASB^{+}19, APHC15, CR19, DFG^{+}17, FHD^{+}19, GKR14, HB19a, KET^{+}17, Mac12, MRD19, MR14b, ØU13, PKR^{+}19, SSR^{+}16, SSM^{+}18, WGCL^{+}19, ZDF^{+}19]. limiting [SBFC10]. limits [BT15, BPMR17, DMO10, GHPH15, OCR14, RDT^{+}17, SBSE14]. Limnocalanus [EKR^{+}19]. line [LGB^{+}18, SLS18, SKKM15]. Linear [BTA^{+}18, FB19, JMM^{+}15a, TSWS15]. linearity [DBA^{+}19]. ling [GKJ^{+}15, LDD^{+}10, LPT10]. link [HKK^{+}17, LD17]. linkages [PB11, PVD14, TMP12]. linked [CN16, DDR10, KGH^{+}16, MM17]. Linking [Ast15b, Bas18c, MJSB14, OLW^{+}14, WWMF17, ZGK^{+}17, HSB16, KHF14, SG16b]. links [BSC14, CLT^{+}14, LMG^{+}14, PRBF18]. Linnaeus [FFN^{+}16, FFN^{+}13, JTM^{+}16]. lion [HB19a, dlTQMUE10]. lionfish [VF17]. lions [VOS10, MDdPMQ10]. lipid [ÖPPM19, RTJS14]. lipped [HTM11]. List [Ano10a, Ano10b, Ano10c, Ano10d, Ano11d, Ano11e, Ano11f, Ano11g, Ano11h, Ano12a, Ano12b, Ano12c, Ano12d, Ano18o]. literature [MSE12]. lithodid [DD13]. litter [GHWD13]. little [BMM^{+}19]. Live [BMVPN10, SHL10]. lived [MONS14, TB16, dMBD11]. liver [KOK^{+}14]. lividus [CRC^{+}16, OFVF14]. living [CFA^{+}16, HPS^{+}11, Mer18]. lixula [VGB^{+}17]. Liza [GSMRM^{+}16]. lobster [BHC^{+}15, BT10, BM15, BMG15, CAMM12, CGG^{+}11, CPB19a, CPB19b, DJW^{+}15, EAB^{+}17, EHD^{+}15, EKD^{+}15, FFN^{+}13, FFN^{+}16, GBM^{+}19, HGGJ15, HJS^{+}17, LHJ15, LGvPH15, LCG^{+}18, LMG^{+}14, LCTB^{+}15, MLMJ17, ØU13, PSH^{+}11, PGG^{+}17, PCdL15, QR15, RCF^{+}18, STR16, SLG^{+}15, TCS^{+}19, THSM^{+}18, WDOJ15, WWMF17, XS12, YJBL17, ZCW11, dL14, dLCF^{+}15]. Lobsters [BFLÁ15, BFDGLÁ15, BBM15, CZBFLÁ^{+}15, CHM15, GDS15, GBB15, KFBG15]. Local [ABJ^{+}19, BBB^{+}19, SBS^{+}19c, Bas18c, BUDM15, CBL19, DVH^{+}18, FCB17, FB19, GRS^{+}14, GHG^{+}10, IGGSAL^{+}15, KCS10, OAT^{+}12, PNC15, RLQ^{+}15, VP19, dPLF^{+}19, CMH19]. located [DPE^{+}17]. location [BGS^{+}14, GBB15, JRWM11, NWM^{+}12]. location-dependent [GBB15]. locations [LHK18, MMWC16, SBL^{+}17]. loch [GCS^{+}18, WDG^{+}17]. Lofoten [BV19b, BV19a, MO13]. log [TKJB19]. log-Gaussian [TKJB19]. logbook [JCSC10, MONS14]. logbooks [GL11, MHHK13, ZBE^{+}15]. Loggerhead [KCP^{+}11, dQCD^{+}10]. logging [LBKTW19, RD16]. logic [Eid18, HTB^{+}16]. logistic [MS15]. Loligo [DRB10, GRP10, GOPG10, MRC^{+}10, PG10, RU12, RDS12, SHL10, TNS13]. Long [CAC^{+}18, DLSJ^{+}19, GPG14, GTGD15, LGBA11, MCSL16, PB11, PDT^{+}18, RLP^{+}13, RCF^{+}17, AFLvDH15, BCK^{+}15, CCA^{+}18, DG11, DL11, FKH^{+}16, GAB^{+}14, HHM^{+}11, HBD^{+}19a, HBD^{+}19b, HLD10, KOK^{+}14, KFBG15, MONS14, MBT^{+}19, OS14, PEN^{+}19b, RB14, Ric19, RHOJ10, SGV18, SWA^{+}16, TNFC16, TB16]. long-lived [MONS14, TB16]. Long-term [DLSJ^{+}19, GPG14, GTGD15, LGBA11, MCSL16, PB11, PDT^{+}18, RLP^{+}13, RCF^{+}17, AFLvDH15, BCK^{+}15, CAC^{+}18, DG11, DL11, FKH^{+}16, GAB^{+}14, HHM^{+}11, HBD^{+}19a, HBD^{+}19b, HLD10, KOK^{+}14, KFBG15, OS14, PEN^{+}19b, PFD^{+}16, RHOJ10, SGV18, SWA^{+}16, SLF16b, TNFC16]. Long-time [CAC^{+}18]. longevity [GOPG10]. longfin [BNP19, MAS19]. longirostris [RRM^{+}12]. longitudinal [BCK^{+}15, FGMC^{+}13]. Longline [MAM^{+}15, BSP15, CJFS16, DMT^{+}10a, FBCD^{+}17, HCM^{+}15, LJB16, PDS15a, PDS15b, PMH^{+}13, RBM15, TGG^{+}15, WNPY15]. longliners [KMK10]. longlines [GLP^{+}11, OSL^{+}15, TGDG15, dQCD^{+}10]. longstanding [ABJ^{+}19, CMH19, SBS^{+}19c]. longtail [GFML10]. Looking [BG18, HFSV^{+}15, RDCPvH11, SWBJ13]. Lophius [CSS^{+}12]. Lopholatilus [FFPL14]. losers [Ful11]. loss [BBH^{+}10, CPB19a, CPB19b, KE14, SH14b, WAH^{+}16]. lost [SS12, dCVB14]. lot [BS14]. Louisiana [KP13]. Low [FSC^{+}12, GMKS11, MDV^{+}15, Bjö18, BFS16, CYLT16, DKPM16, GQX^{+}19, HHJB14, KVSV14, NNY17, RBMAOCL18, SMSR10, SMS^{+}17, SFD^{+}16, dLCF^{+}15, dCVB14]. low-density [RBMAOCL18]. low-frequency [Bjö18]. low-quality [SMSR10]. low-salinity [GQX^{+}19]. low-tech [dCVB14]. lower [CSJ16a, MGVH^{+}10, SLF^{+}16a, XPC11]. luck [Job17]. Lucky [CBC^{+}11a]. luderick [BBHC11]. Lumpfish [EDP14, KJKÓ15, KJÓK16, PSO^{+}14]. lumpfishing [KDF^{+}19a]. lumpus [EDP14, KJKÓ15, KJÓK16, PSO^{+}14]. lunar [WABZ16]. lupus [PSDG12]. Lury [RFMR10]. Lutjanus [BTA^{+}18, SBG10].

M [FWC^{+}19b]. M. [GES^{+}12]. M74 [KUM^{+}12, MKC^{+}11]. maccoyii [WLN^{+}13]. machine [SR17]. Mackerel [SBLS15, TMR^{+}16, AVGÓ12, DDR10, FCG^{+}16, FHK14, GMN^{+}17, GMS^{+}18, GKV^{+}15, Han13, Jan14, JKvdK^{+}15, Kor10, NSO^{+}19, NDP^{+}16, NUÓ^{+}16, OSJ^{+}16, PGV^{+}17, RSBC13, RPDF14, STKT16, SGM^{+}17, SPS^{+}10, SCS^{+}11, TMR^{+}17, VCRPS13, VBA10, vDTF^{+}14, vdKFS^{+}16]. mackerels [PKR^{+}19]. macleayi [GFMO11]. macroalgae [KPD^{+}16, NMF^{+}16]. macroalgal [KKMS14]. macrobenthic [SLM^{+}11]. macrocephalus [GLP^{+}11, HMM12, NUO^{+}10]. Macroctopus [DWPS11]. macrofaunal [DHCE13]. macroinvertebrate [FMML19]. macroscope [APOG11]. Macrourus [Lap11]. macrozoobenthos [dJBLH15]. Macrozooplankton [AIA^{+}15, PFPG12]. Macruronus [KRML11]. macrurus [EKR^{+}19]. made [FB11]. maenas [FLCQ19]. Magadalena [NdlPIR18]. Magadalena-Almejas [NdlPIR18]. magellanicus [KHH^{+}17a, SPS15]. magister [SQ14]. magna [GMM^{+}12]. magnetic [FOJ13]. magnets [Qui18]. magnitude [LF14]. main [ALB^{+}19, GDBM^{+}17, HSK^{+}14, PB11]. Maine [MOG^{+}14a, MOG^{+}14b, CTC14, CMH^{+}16, CKD^{+}17, CCSG18, DHZA14, DEH^{+}19, GCS^{+}17, HRB15a, JFJ^{+}17, Kan11, LYW^{+}18, LCG^{+}18, LBK^{+}19, NSS15, NBS^{+}10, SG10, TCS^{+}19, TCQ^{+}15, ZHD^{+}14, ZLC^{+}17, ZCW11]. maintain [PSM^{+}10]. Maja [VBGG^{+}11]. major [AV15, BHS11a, FBF^{+}17, MFS^{+}14, NBS^{+}10, SS17, SEDO19, XCRP13]. Makaira [SSP^{+}11]. make [CPS13, Sum13]. Making [BPPP17, KOK^{+}14, OSSL16, MTCF^{+}17, Pun17, Ric11]. mako [CJFS16, TSPL14, TLPS15]. maladaptation [Cri12]. Male [HSD^{+}16, HLD10]. Male-biased [HSD^{+}16]. Mallotus [CMD17, GHM^{+}16, HGRAR10, OCR14, PGG^{+}12, SSF18b]. mammal [PDS15a]. mammals [BPF^{+}19a, BPF^{+}19b, BPMR17, HBB^{+}18, HKB^{+}18, HLL18, LHLK10, PMBM18, TTG^{+}15, WNPY15]. manage [HB19b]. managed [KKC19]. Management [Dek16, FLLG15, HHRW12, RD14, ACT^{+}18, AV15, AMS^{+}12, BCRM^{+}18, BSHE18, BNK10, BUNB10, Ber18b, BMS^{+}19, BFC^{+}16, BBD^{+}10, CAMM12, CLT^{+}14, CKB^{+}16, CBR^{+}11, CCC^{+}14, CHB^{+}14, CCS10, CSTJ19, CKA^{+}17, DG11, DGC12, DMJ^{+}14, Day18c, DB16, DKC12, DMO10, DFG^{+}17, DCER^{+}14, DPL16, DGB^{+}15, DHCE13, EAB^{+}17, EHHH14, EP14, FKH^{+}16, FWS^{+}13, FGRR11, Fle15, GNFM11, GFF^{+}17, Gar11, GRC16b, GRC16a, GB15a, GB15b, GB15c, GWS17, GRS^{+}14, GKJ^{+}15, GM18, GBJ^{+}15, Gro11, GTBT18, GAB^{+}14, GFF^{+}19, HHMF19, HABV19, HKKS10, HG10, HKS17, HBB^{+}13b, HSHÖ14, HRC^{+}13, HHA^{+}11, HKK^{+}17, HO14, HDCH^{+}11, HFU^{+}10, HBiI^{+}11, HBC^{+}11, HMFW14, HNAK12, HHH^{+}19, HFNH13, HB10, HFHS10, HHHE16, IHH^{+}11, JRG^{+}16, JCA10, JL12a, Jen13, JLS^{+}15, KCA^{+}14, KM16, KPL^{+}19, KOC^{+}16, KDF^{+}19a, KCK14, KLKD11]. management [KJG^{+}15, KST14, KRG^{+}12, KRO^{+}16, Lap11, LPFH13, LMPP10, LÁG^{+}19, Lev11, LKHK11, LWN^{+}13, LB14, LB17, LDCR^{+}19, LPD^{+}16b, LBKTW19, LBCOJ19, Mac12, MSC14, MHHT10, MVK18, MV13, MWJ11, MLS^{+}17, MKL^{+}19, MLMJ17, MHM^{+}11, MAS19, MLB^{+}14, MNG15, MNGA^{+}17, Mor18, MUEO17, MMSS18, ME11, Nee15, NMO^{+}17, NS17, NJFH19, NCH17, OLW^{+}14, ØHS10, POIM12, Par13, PPD17, PNE^{+}14, PCdL15, PVD14, PvOP10, PVCB17, PDCP14, PEV^{+}16, PWS^{+}11, PFH16, PDD^{+}10b, PAB^{+}14, PFD^{+}16, Pun17, RCH^{+}15, RBB^{+}15, RGGFA16, RCF^{+}18, RR10, RMKG11, SHL^{+}14, SPG^{+}17, SA14, SBH^{+}14, SSMD^{+}13, SMOH18, SKH^{+}12, SFA^{+}17, SJRB18, SRB^{+}14, SSF^{+}12, SFG^{+}15, SPP^{+}16, SBB^{+}17, SM15, Str10, SSZH12, SPS11, SWRC^{+}18, SP13, SH16, SIP18, TLR^{+}17, TH15, TDR^{+}16, TD19, TJH15, TED16, TMW19, Tow14, THW16]. management [UVD^{+}17, VF17, WNM10, WHAA^{+}18, WÓG^{+}16, WMG^{+}17, Wil11, WFM^{+}15, WCR^{+}19, ZHV^{+}16, dCWMH13, dMBD11, vGA18a, vH10, vH13]. Managing [AAP14a, AAP14b, ECG15, GGT^{+}14, HRB^{+}15b, SDH12, TLK^{+}17, CAL14, HS10, HSB16, LLCJ10, OHLK19a, OHLK19b, WWCO15]. mandates [LDCR^{+}19]. Manderson [BF17]. Mangawhai [Alf10]. mangrove [Alf10, KGP^{+}15, MS19, RFHG19]. mangroves [HGS^{+}19]. Manila [LK17]. manned [CBK18]. manual [SHS^{+}17]. many [MP15]. maorum [DWPS11]. map [DMZC17]. mapped [CBD10b]. Mapping [BVC15, CJC^{+}15, JW18, MBA^{+}19, RSGM17, BBD^{+}13, BFH13, CSS^{+}15, CSFS10, EDBMB14, EPH15, GIDP^{+}18, LGRC14, SGV18, SRB^{+}14, SCL^{+}19]. maps [GTBT18, SHJI10]. margin [AGMC14, PCS^{+}11, WKL13]. mariculture [KKAV^{+}19]. Marine [Ano18p, AHK^{+}18, BF17, CMP^{+}11, DOB^{+}17, GHWD13, GDD^{+}18, GHB^{+}14, GRF^{+}12, HAH^{+}14, KLC13, KYK^{+}13, Lev11, LLF18, MAP^{+}11, Man17a, MDDP10, Por11, PKK13, SDSA15, SRDC^{+}14, SGP^{+}15, Tur19a, Tur19b, ZHD10, ACT^{+}18, ASB^{+}11, AIA13, ABRL12, ALB^{+}19, Ast15b, AFK15, AGH^{+}11, APHC15, AHC15, BHC^{+}15, Bak14, BNE^{+}15, BPST^{+}16, BRH^{+}15, BCB^{+}18, BPKH13, Ber18b, BSHC17, BCT^{+}10, BPF^{+}19a, BPF^{+}19b, BA17, BTTV^{+}17, BPF19c, BPMR17, CSS^{+}15, Can16, CMA^{+}16, CP15, CJC^{+}15, Cha12, CKH19, CFA^{+}16, CCC^{+}14, CHB^{+}14, CBK18, CSY^{+}10, CGRM^{+}13, DAP^{+}12, DJW^{+}15, DBDP10, DCS^{+}11, DPL10, Dil17, DPD^{+}12, DPL16, DKC^{+}17, ÉSMGB15, ETM^{+}16, EPPL^{+}11, FKH^{+}16, FWS^{+}13, FHD^{+}19, FB19, FFK^{+}11, FVSL^{+}10, FMM^{+}12, FSS^{+}17, GNFM11, GGTW17, GSP12, GMC17, GB15c, GBJ^{+}15, GGV^{+}18, GF14, GRH^{+}14, GKR14, GJWS19, HD18, HMN^{+}12]. marine [HMLKR18, HSB18, HP17, HUT^{+}11, HHM^{+}19, HBB^{+}13c, HDG^{+}17, Hut14, IZP^{+}16, JDH^{+}14, JRG^{+}16, JLH12, JvdM15, JÓT^{+}12, JC15, JW18, KM16, KCF^{+}17, KHC^{+}17, KKC19, KS14, KS18b, KPBB15, LAD^{+}19, LD17, LGL^{+}12, LPK12, LPD^{+}16a, LWS^{+}18, LRJ18, LMPP10, LØM^{+}18, LNWS18, LYS^{+}10, LB14, LDCR^{+}19, LHLK10, MTP^{+}12, MPR12, MHH^{+}18, MGVH^{+}10, Mat19, MAGK18, MHM^{+}11, MG15, MTCF^{+}17, Mer18, MR14a, MGBvD14, MSR^{+}12, MMDSP15, Mos18, MST^{+}10, NDB^{+}16, OE16, OLM^{+}15, OAT^{+}12, PG16, PDS15a, PBC^{+}16, PNE^{+}14, PAB^{+}18, PCD^{+}13, Pin17, PBL11, Pla16, PRO14, PCS^{+}11, PMBM18, RSGM17, RBB^{+}15, Rid18, ROK^{+}18, RIBB^{+}19, Rot15, RAB^{+}12, SSR^{+}16, SMR^{+}11, SG18b, SHL^{+}14, SSS^{+}12, SOB^{+}11, She15, SBS^{+}10, SS10a, SSB^{+}10a, SMH^{+}19, SFA^{+}17, SBB19, SRB^{+}14, SBS19a, SBS19b, SLS15, SFG^{+}15, SSTB15, SCL^{+}19, TLR^{+}17, TLK^{+}17]. marine [TTG^{+}15, TB16, TMR^{+}16, VQGDAMdL14, VELT14, VRF^{+}16, VDK^{+}18, WPGW12, WNPY15, WRC^{+}12, XPC11, YS15, YPN^{+}18, ZSC16, ZMF12, vPFF^{+}16, vPBLR19, vdHdGL16, LLM12]. marinus [KO12, WCR^{+}19]. maritime [AHK^{+}18]. Mark [WKP^{+}17, BKTS11, SHK17]. Mark-recapture [WKP^{+}17, SHK17]. Marked [WPH^{+}15]. markers [Bak14, BHL^{+}15, GCS^{+}18, OHD^{+}16]. Market [SA14, AGSPH14, CMLVGLP14]. marketability [CGG^{+}11]. markets [LGvPH15]. marking [JOvdM^{+}14, PDD^{+}10a, dBWMD^{+}14]. marlin [HLS^{+}15, KDBP11, MMRG18, SSP^{+}11, SSP^{+}13, VBW^{+}18]. Marxan [DMS^{+}12]. mass [CFHC10, JV12b, MFS^{+}14, SRHJ12]. Massachusetts [HRC^{+}13, MOG^{+}14a, MOG^{+}14b]. masses [vdHBSM10]. Massive [MKB^{+}17, GVA19, Van11]. match [Loh11]. Mate [BBM15]. Material [BW14]. Maternal [SHW^{+}15]. mating [TLPS15]. matrix [TSPL14]. mats [MLMS15]. matters [dMADLP^{+}16, HP12]. maturation [ESR^{+}10, GvRKB19, HCF^{+}12b, MDDP10, OEK^{+}18, WMG11]. mature [OSJ^{+}16, TPBL17, Yar10]. maturity [AEC11, CAOG15, FHC^{+}12, FWD15, HPS^{+}15, HBB^{+}16, MSE12, NLB^{+}12, WMN^{+}11]. Mauritania [MG15]. mauve [BSK^{+}11]. maxima [KN17]. Maximum [FR16, FO13, BSB^{+}19, CDSP11, LB13, LM10, NCH17, PCRD18, RCH^{+}15, UVD^{+}17, vGA18b]. maximus [NMO^{+}17, SMB^{+}16b]. May [WRC^{+}12, HJS^{+}17, LFMJ19a, LFMJ19b, LGR^{+}14, MIM^{+}19, MRD^{+}12, PA18, SMB^{+}16a, VF17]. maya [BMR10a, BMR10b]. ME70 [CBD10b]. MEA [GHB^{+}14]. meadows [DTL^{+}11]. mean [ASV^{+}19, DFG^{+}17, SGM^{+}17, SMB^{+}16a, THH18, TK18, UBvH11, EWH16]. means [BCC^{+}19]. measure [KFN15]. measured [CGK^{+}17, CD14]. measurement [MSK^{+}11, SHS^{+}17]. Measurements [LCM10, ZD13, BDW18, DWFD13, DBA^{+}19, DLSJ^{+}19, DBL^{+}16, FHOK14, KLN^{+}10, KRML11, Mac11, SHS^{+}17]. measures [CFM^{+}14, CKA^{+}17, DVV^{+}11, JGCH18, LMPP10, MHHT10, SI15, TNY^{+}13, WFM^{+}15, XS12]. Measuring [DID^{+}16, Hil19a, Hil19b, KOC15, UHB12, CPT^{+}10, FCB17]. mechanical [CRC^{+}16, DDE^{+}19, TRHB13]. mechanism [TLPS15]. Mechanisms [Mer18, PRO14, CHMY15, GM18, PSM^{+}10, SD17]. mechanistic [MBL^{+}17, PVD14]. mediated [HHH12a, SBD^{+}15]. Mediterranean [ACM^{+}16, ABHT^{+}16, ALMFFBP19, BJH^{+}14, BPF^{+}18a, BPF^{+}18b, BMS^{+}19, CLS^{+}18, CMC^{+}12, CGRM^{+}13, DMO^{+}15a, DFB^{+}10, DCMC19, FMLM^{+}14, FTM^{+}17, FVSL^{+}10, FL14, GCG^{+}10, GvRKB19, GDS15, IGGSAL^{+}15, KE14, MAA^{+}19a, MAA^{+}19b, MHHT10, MVK18, MGL^{+}15, May14, MOVU10, MHMP^{+}18, SMR12, TPV14, TID^{+}10, VFR^{+}16, VPSV^{+}10, VSP^{+}14, dQCD^{+}10]. medium [APOG11]. medium-resolution [APOG11]. meet [BFS16, DKC^{+}17]. meets [ETM^{+}16]. megabenthic [JPTC16]. megabenthos [BMEBM^{+}16, PEV^{+}16]. megafauna [BKMdMS13, JLS^{+}15, LAD^{+}19, YFG^{+}17]. megafaunal [CMA^{+}16]. megalocyathus [CAOG15]. Meganyctiphanes [CK12, JLL17, KOC15, MPL^{+}14, MDS13]. megascale [TNY^{+}13]. Megrey [Ano12e]. megrim [MCAM14]. meiofauna [SPdJFMS17]. Melanogrammus [BGM^{+}14, CPVS14, LSY^{+}14]. Melicertus [OLW^{+}14]. melops [HSD^{+}16, HSV^{+}17]. melt [HMLKR18]. membras [LHK18, OTAK15, RHV14]. Menhaden [SBL^{+}17, ALWH18, BMH^{+}16, HAH^{+}16, RRH^{+}15]. Mental [BA17]. mentella [CC17, CBD^{+}10a, CMP^{+}11, MAP^{+}11, PJDN12]. mercury [CCC^{+}14]. meridional [Kam14]. meristic [OHD^{+}16]. merits [CMG^{+}17]. Merlangius [BGM^{+}14, dCWMH13]. merlangus [BGM^{+}14, dCWMH13]. Merluccius [CLV^{+}14, FMLM^{+}14, FWD15, GES^{+}12, IEL^{+}15, JCA10, MDdPMQ10, SSGH18, TDN^{+}19, WSG^{+}17, WRMJ13]. merobenthic [DWPS11]. meroplanktonic [DDF^{+}12]. mesh [FMK10, HWK^{+}15, SHG^{+}18, SWvSR11, dlPG15]. meso [PFPG12, RGC^{+}19]. meso- [PFPG12, RGC^{+}19]. Mesodesma [RHOL11]. Mesopelagic [Peñ19a, AML^{+}19, COHdP19, Dav11, DKK15, DCMC19, FC19, HB19b, KDF^{+}19b, LOOO16, MUW^{+}19, Pre19, PHK^{+}19, SV10]. mesophotic [TBHK17]. mesopredator [GGV^{+}18]. Mesoscale [DCHMHQ12, GHD^{+}10, KCP^{+}11, SSS^{+}12]. mesozooplankton [ASDSM18, DAB^{+}12, ISL19, LD17, MYI12, MDPH12]. messenger [BCRM^{+}18]. meta [Bru10, Ham15]. meta-analysis [Bru10]. meta-analytical [Ham15]. metabarcoding [BYQ^{+}19]. Metabolic [FAP^{+}17, VF17, DI11, RCG^{+}18]. metabolism [TWB^{+}19]. metal [BWM^{+}16]. Metanephrops [TPHC15]. Metapenaeus [DI11, GFMO11]. metapopulation [GMKS11]. metapopulations [PG16]. Metazoan [GMD^{+}11, UPF^{+}10]. Method [HHR17, ABRL12, BMC^{+}17, CPB19a, CPB19b, DZ14, DSA13, DPRG^{+}18, DBL^{+}16, Fle15, FBM^{+}18, Ham15, HOV^{+}15, ØHS10, OBL12, RFMR10, SKM18, UBvH11, VPH18, ZPS^{+}18, ZDF^{+}19, dCVB14, dlPG15]. methodological [ESD^{+}16, PA18]. methodology [Eid18, KNH11]. methods [AH15, BSW^{+}11a, BSW^{+}11b, CDC15, CR11, DUM^{+}12, DBM^{+}15, HPDB19, HHQ11, LJH^{+}12, LSJ10, LS10, MRP^{+}16, Nee15, Nel19, PKR^{+}19, SCJI10, SFN^{+}19, TSP12, TKJB19, VQB^{+}11, WOW^{+}17, WLL^{+}13, Win15, ZRU^{+}15]. métier [AUEC12, DUM^{+}12, TCF^{+}13]. métiers [KMK10]. metric [GW16]. metrics [BKSB12, Cor12, CSD16, GFR^{+}12, RGGFA16]. Mexican [CPSCCE^{+}17]. Mexico [ALWH18, BTA^{+}18, EPPL^{+}11, GPP17, GCL^{+}18, Lev11, LBTS^{+}19, MLLL11, RGC^{+}19, RRH^{+}15, SBG10, SLT^{+}14, RR11, dÁMNGS^{+}11]. MHz [KLN^{+}10]. mice [Hil16]. Micro [EMW^{+}16, WSW^{+}19, CGK^{+}17, LD17]. micro- [LD17]. Micro-CT [EMW^{+}16, CGK^{+}17]. Micro-geographic [WSW^{+}19]. Microbial [CAC^{+}16, SAB^{+}16, WZZ^{+}16]. microchemistry [GNDC11, HGH^{+}16, LPD^{+}16a, RDB^{+}18a]. Micromesistius [OOD13, PGOM11]. micronekton [LCS^{+}15]. microplankton [WDG^{+}17]. Microplastic [LOOO16, ELBS17]. Microprocessor [HBHR18]. Microprocessor-based [HBHR18]. Microsatellite [GSL^{+}11, IEL^{+}15, BPPB12, GCW^{+}18, OVW^{+}17]. microsatellites [SHH^{+}17]. microstructure [BWR^{+}15, LCM10, RHHH^{+}12]. microzooplankton [HBG^{+}16, LD17]. Mid [SZOL19, HPS^{+}11, JCS16, Pin17, SLS18, BFH^{+}18]. mid-20th [Pin17]. Mid-Atlantic [SZOL19, SLS18, BFH^{+}18]. mid-North [HPS^{+}11]. mid-west [JCS16]. midnight [GCBI17]. midwater [BSC14, KDF^{+}19b, SWBJ13, WPWH11, WMJ13]. might [Cla18d]. migrant [CTT14, CCSG18, WABZ16]. Migrating [HHM^{+}16, DHKE16, KJÓK16, KBH^{+}18, MHE^{+}16, VHF^{+}17]. Migration [FL14, íHHJ^{+}13, LFM^{+}18, BCS^{+}17, CCD^{+}19, DMBD10, HUT^{+}11, HAF^{+}16, KGN^{+}10, LWS^{+}18, LSWD12, LØM^{+}18, Mac12, MFP^{+}19, MHSG19, MGBvD14, MSR^{+}12, MGH^{+}12, OEK^{+}18, SRC11, SV10, STC^{+}17, SFN^{+}19, TDN^{+}19, TOA^{+}16, UFJ^{+}18, VWMS19, WKP^{+}17]. migration-timing [HAF^{+}16]. Migrations [Lap11, AJ12, BPST^{+}16, DLL12, MALM17, ODL16]. Migratory [TVJ^{+}14, AL10b, BLS^{+}18, BG18, GSMA15, JWS^{+}18, LFGW13, LSCL18, NS17, SSRT10, SJRB18, dL14]. Mikawa [TMW19]. mill [GBM^{+}19]. millennium [Pas14]. Mind [Pay10]. mineralogy [PWCS17]. Minho [CCA^{+}18]. Minimum [PFH16]. Mining [Van11]. minispinosa [AEC11]. Minor [HTRF17]. Mira [LTQ14]. Miramichi [CB12, RC12, STC^{+}17]. miranda [HLD10]. mirror [HFSV^{+}15]. misconception [DGPC^{+}15]. misinterpretation [Bar19, RR10]. misinterpretations [BBD^{+}10]. mislabelled [vdHBSM10]. mislabelling [BCSG13]. Misleading [vdHBSM10]. mismatch [KCK14, Loh11]. mismatches [KHC^{+}17, KRO^{+}16, MMSS18]. misreporting [HSHÖ14]. Missing [SS12, BFDGLÁ15, ECG15, Mac17b]. Misspent [VOM11]. misuse [DCPTN14]. Mitigating [TGG^{+}15, WNPY15, CMH^{+}16, HCM^{+}15, WLL^{+}13]. mitigation [CFM^{+}14, COL^{+}13, RBM15, RCC14a, TNY^{+}13, WSGD^{+}17]. Mitochondrial [BPPB12, FFN^{+}13, FFN^{+}16, DMF13, VPSV^{+}10, VSP^{+}14].

Mixed [TM15, AUEC12, BUNB10, BMF14, BHS^{+}16, CMM^{+}15, DGC12, FCD^{+}11, FB19, HAF^{+}16, HKS12, JPPY11, KFN15, LH19, MLT11, MV13, MS15, MPR17, MDPH12, OHD^{+}16, PDR19, PBQ^{+}10, STM19, SLT^{+}17, TJD17, TSWS15, URV^{+}11, UVD^{+}17, WGCL^{+}19, vHCP15, Fra15]. mixed-effects [MS15]. mixed-fisheries [BUNB10, MV13, URV^{+}11]. mixed-fishery [DGC12]. mixed-species [KFN15]. mixed-stock [BHS^{+}16, OHD^{+}16]. mixed-substrate [BMF14]. mixing [BNE^{+}17, BDCvD^{+}10, BL15, CTC14, ET15, GNDC11, HHE^{+}16, JRWM11, SWA^{+}16, SJ15]. mixture [BMGB19]. mobile [GML13, HSGL18, MBLS15, YPN^{+}18]. modal [KAT11]. mode [RDB^{+}18b]. Model [MJS^{+}15a, PMOH13, TCQ^{+}15, ASI^{+}16a, AHM^{+}15, AML^{+}19, AFQ^{+}11, AFK15, AHC15, BHC^{+}15, BGS^{+}14, BMC^{+}18, BPST^{+}16, BCK^{+}15, BSF^{+}19, BN16, BMS^{+}19, BNAH13, Cad13, CC17, CPVS14, CIJ18, DRGCGJ17, DDR10, DCPTN14, DWPS11, DMBD10, Eid18, EAOB15, EP14, EBSW15, FFK^{+}11, GG13, GLC15, GRC^{+}14, HJT12, HFU^{+}10, HB10, HKS12, IFU11, JMM^{+}15a, JC15, KDH^{+}10, KWF11, KRML11, KRD^{+}18, LCS^{+}15, Lév15, LSR^{+}13, LSI^{+}15, LVPK10, MFP^{+}19, MWP^{+}15, MDS13, MR10, MR14a, Nee15, PBJ^{+}14, Pay10, Pla17, RDB^{+}18a, RC15, RFMR10, RCH^{+}19, SLKS10, SBCF16, SCS^{+}11, SFW^{+}12, SHST18, SRW13, SWT^{+}19, TCS^{+}19, TBC15, TJR^{+}10, TSP12, Tow14, TMP^{+}17, TSPL14, TLPS15, WFM^{+}15, ZCR15, vPGFT13]. model-based [CC17, Eid18, SWT^{+}19, TCS^{+}19]. modeling [LJB16, SR16]. Modelled [PDAC^{+}16, SAS^{+}18, FOJ13, KRD^{+}18]. Modelling [BLB^{+}18, CAL14, CSC^{+}13, Cla18a, CK18, DMO^{+}15b, EAOB15, HMS^{+}13, HF14, iIOKW13, JSOO12, KPBB15, LPH^{+}15, LHJ15, MPL^{+}14, MGH^{+}12, MJAS14, MHMP^{+}18, NMO^{+}17, ØFNH12, OFB^{+}17, PG16, PSB16, PPD17, PDH^{+}14, PLSD16, RU12, RPS11, SDT15, SRH^{+}14, SHZ^{+}16, SP14, SSP^{+}11, SDSN14, TPL11, YYCC16, ZKD^{+}14, ACCA19, ABHT^{+}16, ALBR14, BJH^{+}16, BNE^{+}15, BGA^{+}19, BMR10b, BSMB18, BBD^{+}13, CCL^{+}14, GGP11, GJWS19, HCC^{+}14, HSB16, HP11, HMR15, JVPM12, KCS10, LHS^{+}18, LVM^{+}11, LHS^{+}19, MSOM10, MHHT10, MLT11, MV13, MGPC^{+}14, MS15, MTCF^{+}17, NSO^{+}19, RBB^{+}15, RFMR10, Sub18a, Sub18b, TM15, TJH15, TBC^{+}14, TDHC18, VHQ^{+}11, YCZY19, ZMF12]. models [ANT18, ABF^{+}16, BSHE18, BMGB19, BFC^{+}16, BA17, BHS^{+}11b, BFH^{+}18, BPC10, CAM^{+}15, CTT14, CIJ18, DGC12, DHH^{+}14, DBM^{+}15, EP14, FB19, FWP^{+}16b, GDJ11b, GMM^{+}12, GÁE16, GTBT18, GT19, HLCC16, HDCH^{+}11, HPN^{+}17, HHQ12, HFSV^{+}15, JSOO12, JOSO16, JMM^{+}15b, KM16, LHP^{+}19, LLL^{+}13, LPB^{+}14, LTGP18, LPT10, MPB11, MBA^{+}19, MBL^{+}17, NdlPIR18, ØS15, OYIN18, OLM^{+}15, PB14, PHM13, RSY11, RMF^{+}15, RU12, RWG^{+}18, RTS19, SHW^{+}18, SSM^{+}18, SPS^{+}10, SBH^{+}14, Tho11, TSWS15, TB17, TFC^{+}19a, TFC^{+}19b, THC16, TMR^{+}16, ZCH19, dlPG15]. Moderate [OMTY14]. modern [SLG^{+}15]. modernization [Hub14]. modes [CBD10b]. modification [HLA^{+}11]. modifications [HCR13]. modified [MBBC11, SCJI10, VKT^{+}12]. modifies [GOH^{+}19, RCF^{+}17]. modify [SMZ^{+}14]. MODIS [LSI^{+}15]. modulates [SMB^{+}18]. Molecular [NPB^{+}15, BCOS11, CPSCCE^{+}17, MJSB14, RGPG10]. mollusc [PHM13]. Mollusca [SLF^{+}16a]. molluscs [RPB16]. Molva [GKJ^{+}15, LDD^{+}10]. MoMAR [CBC^{+}11a]. MoMAR-D [CBC^{+}11a]. monitor [CBC^{+}11a]. monitored [TOA^{+}16]. Monitoring [BTL17, FBFP^{+}13, GL11, GML13, LJH^{+}12, LSJ10, MKD^{+}18, MHHK13, NDB^{+}15, OSP^{+}13, ACT^{+}18, CBC^{+}11b, ESD^{+}16, EF15, GVA19, GUN^{+}19, HRC^{+}13, JL12b, JWS^{+}18, KOC^{+}16, LPB^{+}14, LGBA11, MGVH^{+}10, PGS^{+}19a, PGS^{+}19b, RBC^{+}15, SKH^{+}12, SMMK11, SKKM15, TMS^{+}15, vHCP15, vHCP17]. Monte [LH19]. Monterey [UHB12]. monthly [CR19]. Montipora [BJR17]. moored [OdJN^{+}12]. moratoria [EGC^{+}17]. Moreton [KS17, WWCO15]. morhua [ABJ^{+}19, BGM^{+}14, BKTS11, BHLS18, BCOS11, CMH^{+}16, CMH19, CTH^{+}19a, CTH^{+}19b, CKD^{+}17, DHZA14, FWH^{+}15, FCB17, GDJ11a, GIW16, GHV^{+}18, GF11, GCS^{+}17, HCC^{+}14, HRC^{+}13, HWK^{+}15, HHM^{+}11, HSK^{+}14, HBD^{+}19a, HBD^{+}19b, HUPBL18, Hüs10, Hüs11, HHH12b, HGH^{+}16, IKF17, JKS^{+}18, JOvdM^{+}14, KOK^{+}14, KR12, LSY^{+}14, LFGW13, LYW^{+}18, LWN^{+}13, Mag11, MCSL16, MSR14, NRTK^{+}19, OAT^{+}12, OBY^{+}14, OBL12, SSRT10, SG10, SHG^{+}18, SSP12, SBD^{+}15, SO10, SMR^{+}10, SH15, SBS^{+}19c, TCBD10, TNH^{+}10, UBV^{+}11, WMG11, ZMKC14, ZHD^{+}14, ZLC^{+}17, dHFF^{+}16]. Moroccan [RFMR10]. Morone [VWMS19]. Morphological [GRG^{+}10, VGB^{+}17, CPSCCE^{+}17, KN17, SSA^{+}18]. morphology [BHDB12, KSZ10, LVSF17, MIM^{+}19, UBV^{+}11]. Morphometric [GNKS18, SPFM^{+}10]. morphometrics [WNW^{+}15, ZGTL13]. mortalities [DRGCGJ17, GVA19, OYI17]. mortality [ACA19, AST^{+}15a, AIA^{+}15, AS14, BPKH13, BCK^{+}15, BSW^{+}11a, BPMR17, BMSC16, BSMB18, BBH^{+}10, BBOE10, CJFS16, CMH^{+}16, CECL16, Cha12, CFHC10, CIJ18, CHD^{+}18, FHD^{+}19, GOPG10, Ham15, HB19a, HCR13, HSNO12, HTB^{+}16, HBC^{+}11, Hol10, HTS^{+}10, HHT13, JTE19, JSOO12, JMM^{+}15b, JTM^{+}16, KVFK19, KHMS19, LSY^{+}14, LKR16, LSF^{+}18, MLMS15, MSM17, MSR14, ODSC10, Pep16, PGS^{+}19a, PGS^{+}19b, PH17, PvdKE^{+}12, Pow14, PO14, RLD^{+}12, SRGF15, SSF^{+}18a, SHZ^{+}16, TBG16, THHH15, THH18, THHH18, Urb15, VWMS19, VHS^{+}11, WSGD^{+}17, WGCL^{+}19, WS13, WKP^{+}17, WNW^{+}15, Win15, ZBE^{+}15, ZD13, NLB^{+}12]. moss [FLCQ19]. most [SG18b]. mostly [Mur10]. motility [GRF^{+}16]. motion [TKA^{+}16]. moulting [CPP17, TWHB19]. mountains [TBC^{+}14]. mounted [ZDD^{+}19]. mouth [BPH16]. mouths [MAM^{+}15]. move [AGH^{+}11, MSS^{+}19]. move-on [AGH^{+}11]. Movement [MLC^{+}19, KCP^{+}11, PDD^{+}10a]. Movements [AGG^{+}18, DTA^{+}13, KJKÓ15, LBK^{+}19, TTR^{+}19, AMQRC10, BGA^{+}19, GJWS19, KJÓK16, ØU13, SKM18, SHAM10, VBW^{+}18, ZLC^{+}17]. movescape [LBKTW19]. Moving [CKA^{+}17]. mp [WDG^{+}17]. MPA [Aga18c, Bas18c, EPH15, JWS^{+}18]. MPAs [Cla18d, DSL14, Hil18a, SJ18]. MRI [FOJ13]. MSC [AGSPH14]. MSFD [PKK13]. MSI25 [TCF^{+}13]. MSVPA [GLK^{+}10]. MSY [Cad13, DOB^{+}17, HNAK12, May14, MR10, RCS^{+}17, TD19]. MSY-based [DOB^{+}17]. MSY-orientated [HNAK12]. mtDNA [VCF^{+}12]. much [GML13]. mud [BLB^{+}12, MR14a]. mullet [GSMRM^{+}16]. Multi [JC15, BMC^{+}17, CMM^{+}15, CCD^{+}19, EAB^{+}17, MLS^{+}17, NCH17, RBBC11, SG17, SDSA15, SLS18, TK18, VPO17]. multi-beam [TK18, VPO17]. multi-experimental [RBBC11]. multi-frequency [SG17]. multi-gear [BMC^{+}17, CMM^{+}15]. Multi-model [JC15]. multi-sectoral [MLS^{+}17]. multi-species [EAB^{+}17, NCH17]. multi-stock [CCD^{+}19]. multi-trap [SLS18]. multi-use [SDSA15]. multi-year [CCD^{+}19]. multiannual [Str10]. multibeam [CSS^{+}15, CBD10b, DTL^{+}11, MBQ10, Mel16, OSP^{+}13, PPM13, SHJI10, TMR^{+}17]. Multidecadal [EKR^{+}19, Kam14, MLC16]. Multidisciplinary [MMO16, DGL^{+}17, LBTS^{+}19]. multifleet [PKR^{+}19]. Multifrequency [CK12, KBB13, MDS13]. Multiple [RSB^{+}15, ALWH19, Ben13, DSN16, GF14, GT19, Ham15, Har14, HGC^{+}19a, HGC^{+}19b, KRG^{+}12, MBA^{+}19, MMSS18, MFB17, NDM^{+}16, OIMP18, PPD17, SDH^{+}18, ZDD^{+}19, ZKD^{+}14]. multiplexed [MSOF18]. multiscale [dSMGKP12]. Multispecies [OKDJ18, BSB^{+}19, DCVC15, GFF^{+}17, HFU^{+}10, HDG^{+}17, JVPM12, KDH^{+}10, KTMV16, PQR10, RDL^{+}17, RDT^{+}17, SA14, SDSA15, TDR^{+}16, TJD17, TB17, TPL11, TFC^{+}19a, TFC^{+}19b, WGCL^{+}19, WWCO15, WHMP15, WHP16, ZCR15]. multistock [GPS^{+}17]. Multivariate [BPF19c, WOW^{+}17]. Multiyear [Kan11]. mums [AJM19]. Munidopsis [CAMM12]. Murman [MMC19]. murphyi [VCRPS13]. mussel [BCL^{+}11, COR15, CBC^{+}11b, KPGH16, PCBO^{+}18]. mussels [NDM^{+}16]. Mustelus [FMC10]. Mycale [VSB^{+}16]. myctophid [YST^{+}10]. Myctophidae [OBLP^{+}19]. myctophids [Sas19]. mysid [KBB13]. Mysis [SR16]. mysterious [Ger17]. myth [CSJ16a, EP19]. myths [CSJ16b, Ove19, WHH16]. Mytilus [COR15, KPGH16].

N [RGGFA16, DKB14, FGMC^{+}13]. NAFO [NPGT15, FDC^{+}16, HHGtIWGoOH12, Mäl12, WRC^{+}12]. Namibian [WRMJ13]. narrow [NST^{+}14, Qui18]. nascent [WOW^{+}17]. Nash [NCH17]. Nassarius [ZSC16]. nasus [BCS^{+}17, CRHM19, CJFS16, SRC11]. natal [LPD^{+}16a, PDZ^{+}13, VCF^{+}12]. National [Lev11, Dek16, Fri19, GM18, HD18, HGS11, Mos18, Rid18, Sca18]. Natural [Abl16, BBL^{+}15, MS12, SMB^{+}16b, WKK^{+}15, BFS16, CECL16, DFB^{+}10, FJSJ15, GS14, GS15, GOS^{+}13, Ham15, HTRF17, HTB^{+}16, JSOO12, JMM^{+}15b, Pow14, RGAS^{+}18, Sch14, SMB^{+}16a, THHH15, THH18, THHH18, WÓG^{+}16, WMG^{+}17, Win15, ZBE^{+}15, ZD13]. Nature [JCSC10, Cur19, DC14, GGTW17]. nautilus [RS10]. naval [KKLM13, SKA^{+}12]. navigation [SP14]. Near [BJH^{+}16, PSH13, Pep16, RB10, SMD^{+}16, VGB^{+}17]. Near- [BJH^{+}16]. near-future [PSH13, VGB^{+}17]. Nearshore [BdSP11, BSP15, SJRB18]. nebulosus [HGF^{+}18]. necessity [PvOP10]. need [BKVT16, DVV^{+}11, GNFM11, GRH^{+}14, LÁG^{+}19, SC16]. needed [HB19b]. needs [Abl16]. negative [RFT^{+}16]. neighbouring [DLL12, EHHH14]. nekton [FW10, HPS^{+}11, LEB^{+}14]. Nemopilema [KPJC14]. Neogoniolithon [FTM^{+}17]. neon [Che10, YCZY19]. Nephrops [BMC^{+}17, HJS^{+}17, MKHK18, NDF^{+}15, PSH^{+}11]. Nephrops-directed [MKHK18]. nest [HSD^{+}16]. nestedness [DCHMHQ12]. net [HHDB14, TMR^{+}17, TRSM15, Tur19a, Tur19b, WLL^{+}13]. net-based [WLL^{+}13]. net-zero [Tur19a, Tur19b]. Netherlands [vdHdGL16]. nets [BBHC11, RGR^{+}11, SSF^{+}12]. netting [dlPG15]. network [AHR^{+}19, AHC15, DHS19, EPH15, HG10, MAGK18, RMF^{+}15, SSM^{+}18, TMP^{+}17, WZZ^{+}16]. Networks [MRSG^{+}17, AFQ^{+}11, ALB^{+}19, JDH^{+}14, LKHK11, Hub14]. neural [AHR^{+}19, SSM^{+}18]. Neutral [ÖPPM19]. never [SH15]. Newfoundland [CMD17, Ham14, KR12, MDMC11, MMAS11, OCR14, PNS18]. newly [HdBR^{+}16, OFB^{+}17, WMN^{+}11]. next [BMC^{+}13, HD18]. niche [MSS^{+}19]. Night [KKMS14, VBA10]. Night-time [KKMS14]. nighttime [MMG^{+}17]. nigricans [SSP^{+}11]. Nilsson [NWM^{+}12]. Nine [EPKR11, DLL12]. Niño [KKAV^{+}19]. nitrogen [JvdM15, LHJ15, LBG18]. No [CPLH16, GM17, MGS^{+}15, SG18b, CGRM^{+}13, FHD^{+}19, FFRR12, PCBO^{+}18, SJ16]. no-fed [PCBO^{+}18]. No-take [SG18b, CGRM^{+}13]. NOAA [OdS19]. noctiluca [BSK^{+}11]. nocturnal [GFSC17, TBD^{+}18]. noise [DW11, DH13, DWFD13, DLSJ^{+}19, HP17, PPS17, RDKK15]. noise-reduced [DW11, DH13, DWFD13, DLSJ^{+}19]. nomurai [KPJC14]. non [BHS11a, Cad13, CIJ18, CPB19a, CPB19b, DHNL12, FRM^{+}16, FLCQ19, HSHÖ14, JSOO12, JBE14, LCG^{+}18, NMF^{+}16, SRGF15, SWS13, SH16, TWHB19, TH15, WGCL^{+}19, WLB11, WRDF10, XMM12, Yar10]. non-calcifying [NMF^{+}16]. non-commercial [JBE14]. non-compliance [HSHÖ14]. non-continuous [TWHB19]. non-fish [XMM12]. non-fished [BHS11a]. non-indigenous [DHNL12, FLCQ19, WLB11]. non-invasive [CPB19a, CPB19b]. non-management [TH15]. non-parametric [Cad13, CIJ18]. non-reproductive [Yar10]. non-stationarity [WRDF10]. non-stationary [JSOO12, LCG^{+}18, SH16]. non-target [SRGF15, WGCL^{+}19]. non-trawled [SWS13]. non-trophic [FRM^{+}16]. Nord [SMBH16]. Nordic [NSO^{+}19, NUÓ^{+}16, RKB^{+}12]. norms [GRC16b]. North [CCG14, DSA13, DK19, HFSB^{+}14, LGBA11, ME11, SRCR12, GTM12, GPG14, KGP^{+}15, RRCT^{+}17, ASI^{+}16b, AFLvDH15, BGM^{+}14, BR12, BJMG19, BDCvD^{+}10, BNB19, BGA^{+}19, BT10, COM11, Cha12, Che10, CFT^{+}19, CCL^{+}14, CHMY15, CLV^{+}14, CKV^{+}16, Cor13, DBDP10, DUM^{+}12, DCNB^{+}10, DCER^{+}14, DK18, Elv15, EPKR11, FGMC^{+}13, FPLB19, FFRR12, GIW16, GSC10, GMKS11, GRR^{+}11, GFR^{+}12, GRR^{+}12, GKR10, HCC^{+}14, HPS^{+}11, HLS^{+}15, HLMT16, HHHE16, JPPY11, Jan14, JV12b, KDH^{+}10, KDF^{+}19a, KJW^{+}18, KBT14, KKAV^{+}19, KJG^{+}15, KRD^{+}18, LLSJ18, LVPK10, LOOO16, LGR^{+}14, LHH^{+}13, MCAM14, MSOM10, MALM17, MDP^{+}15, NWM^{+}12, NLB^{+}12, NRTK^{+}19, PPHK19, PSO^{+}14, Pay10, PNC15, PDWH11, PSF12, PKK13, RDD^{+}10, RvDW10, RDCPvH11, RKCP18, RFT^{+}16, SHK^{+}15, SRHJ12, SBH^{+}14, SDT15, SHG^{+}18, SSGH18, STSP17, SSP^{+}13, SWS13, SPV^{+}16]. North [TH15, TD19, THKP11, URV^{+}11, UVD^{+}17, VVP^{+}14, WHL^{+}17, WKK^{+}15, WDG^{+}17, WRC^{+}12, WVG^{+}13, vdRMC^{+}17]. North-East [SHG^{+}18]. north-west [RRCT^{+}17]. Northeast [MMM11, PDCP14, PGN^{+}11, GLH14, HSK^{+}14, KOK^{+}14, LSY^{+}14, MDMC11, RSBC13, RHOJ10, Yar10, ASB^{+}11, AV15, AGG^{+}18, BR12, BW14, BRH^{+}15, BCS^{+}17, Bru10, BFH13, CNBK11, CSS^{+}12, COS11, CPVS14, CDSP11, CCC^{+}14, CAB^{+}10, DMO10, EPH15, FMC10, FHC^{+}12, FMML19, GMC17, GSL^{+}11, GKJ^{+}15, Han13, HLL^{+}12, íHHJ^{+}13, HCF^{+}12a, HCF^{+}12b, JKvdK^{+}15, JÓT^{+}12, LLL^{+}13, LK15, LLF18, MGPC^{+}14, MPJ12, MJSB14, MVP^{+}14, MGH^{+}12, NSO^{+}19, NUÓ^{+}16, OSJ^{+}16, RFHS14, SRC11, SCS^{+}11, SHST18, VSP^{+}14, WSG^{+}17, WMN^{+}11, vdKFS^{+}16]. northeastern [BFB^{+}11, Dav11, MUW^{+}19, SCC^{+}13, PCB^{+}14]. Northern [HRB15a, JLL17, RGAS^{+}18, ALWH18, Alf10, BCC^{+}14, BGM^{+}14, BBB^{+}19, Bro14, CK12, CAC^{+}18, CDAN^{+}14, CN16, DG11, DLL12, DPD^{+}15, DHKE16, DK18, EMP11, FBCD^{+}17, GPP17, GKV^{+}15, GCL^{+}18, JTJ18, KJG^{+}15, KKLM13, LFGW13, MCAM14, MFT16, MKAR^{+}18, MR14b, OEK^{+}18, OFYS^{+}11, PDD^{+}10b, RHV14, RHOL11, RRH^{+}15, RLP^{+}13, SBG10, TCBD10, UFJ^{+}18, VHGTFR10, WKP^{+}17, YLLG18, BTA^{+}18, CS15, HBB^{+}16, HKS12, MGS^{+}15, QCA^{+}18a, QCA^{+}18b, VFR^{+}16]. northernmost [OVW^{+}17]. northward [VHS^{+}10]. Northwest [FFK^{+}11, BCOS11, CCD^{+}19, GTM12, LDD^{+}10, ML12, YCZY19, AWS^{+}13, BKMdMS13, BSC14, BMH^{+}16, Der18, EPPL^{+}11, FLP^{+}13, FMM^{+}12, KPJC14, KLC13, MOVU10, MMAS11, NGTT16, OFYS^{+}11, ÖPPM19, PRKASR12, RJ12, SBNL12, SFNdH16, SRCR12, SS10b, SBKA16, TDN^{+}19, TSPL14, WRDF10, YST^{+}10]. northwestern [FCCLA10, FMLM^{+}14, Lev11, SMR12, SLH^{+}17, SWA^{+}16, TSS^{+}13, TCH^{+}16]. norvegica [CK12, JLL17, KOC15, MPL^{+}14, MDS13]. norvegicus [HJS^{+}17, PSH^{+}11, SHH^{+}17]. Norway [AASD12, BNB19, BV19b, BV19a, CKV^{+}16, DMF13, FBCD^{+}17, FBF^{+}17, GDJ11a, HJS^{+}17, HKS12, KMF13, KKLM13, NWM^{+}12, NLB^{+}12, PSH^{+}11, SHS15, SMBH16, UFJ^{+}18, VKN^{+}11]. Norwegian [Aks15, AJM19, ACS15, DFS^{+}19, GZS^{+}19, GSL^{+}11, GUN^{+}19, GHV^{+}18, GAB^{+}14, HSNO12, Hje14, JDFN12, KGH^{+}16, MJAS14, NDP^{+}16, ØFNH12, Ott10, PH11, Rey12, SMNE14, SJUE11, SM12, SBLS15, SDBB15, SHAM10, TKB^{+}15, VHS^{+}10, VKS^{+}12, WHNS14, Win15, ZH13]. Notice [Ano10e]. notional [RML^{+}16]. nouryi [RS10]. Nova [EAB^{+}17]. novaezealandiae [THR16]. novaezelandiae [KRML11]. novel [EBSW15, HOV^{+}15, LKG^{+}19, MBS^{+}15, UFLH^{+}13, WOW^{+}17]. NPZ [Lév15]. nucleotide [GCS^{+}18, SHK^{+}15]. number [DHKE16, KFBG15, LCMR^{+}18]. numbers [GUN^{+}19, JMF12, PDT^{+}18, SHS15]. Numerical [ET15]. Nurseries [GCG^{+}10, Jon14, WDOJ15]. Nursery [CDAN^{+}14, JKvdK^{+}15, NST^{+}14, SBSE14, AV15, BGM^{+}14, HCDAF15, dSMGKP12, PGV^{+}17, PHM^{+}14, PDAC^{+}16]. nutrient [KYK^{+}12, MOG^{+}14a, MOG^{+}14b, TCQ^{+}15, TWB^{+}19]. nutrients [DRVV^{+}17]. Nutritional [PHM^{+}14, GES^{+}12, SY16]. NW [GRS^{+}14, BYQ^{+}19, MGL^{+}15]. nylon [GHV^{+}18].

obesus [BMO^{+}19, WLN^{+}13]. OBIS [RBCH10]. Objective [BGM^{+}11, CSFS10]. objectives [CKA^{+}17, DCER^{+}14, GFF^{+}17, PPD17, SDSA15, SWRC^{+}18, WLB11]. obligation [CLS^{+}18, GPS^{+}17, CEP^{+}18, PDR19]. observation [DHAH12, HHQ11, PMOH13, ZCR15]. Observations [FCD^{+}11, KJÓK16, Mel16, BN16, CRC^{+}16, DHH^{+}14, DW10, DLSJ^{+}19, Eid18, GCBI17, HHGtIWGoOH12, THH18, VPO17, VHQ^{+}11]. observatory [HLL^{+}12]. observe [FA10, HHR17]. Observed [MW19, AJ12, BAFRJ18, GZS^{+}19, RKB^{+}12, SCM19, TMR^{+}17]. observer [APK11, Fau11, FB11, GWS17, PLSD16]. observers [HMP^{+}15]. Observing [HSB16, THW16]. Obstinate [Cur19]. obtained [FA10, MRSG^{+}17]. Obura [Bas18b, Obu18b]. occupancy [WDE^{+}18]. occur [GHB^{+}14]. occurrence [AVGÓ12, BFH^{+}18, MDV^{+}15, PLSD16, TIS^{+}19, TBC15]. occurring [NPB^{+}15]. Ocean [BFC^{+}16, FFN^{+}16, HSS^{+}16, HHGtIWGoOH12, LLL^{+}11, NDM^{+}16, NF16, OFYS^{+}11, PGN^{+}11, SZOL19, SMH^{+}19, STC^{+}17, ASI^{+}16a, ASI^{+}16b, BFLÁ15, BMM^{+}19, BL15, Bro16, Bro17b, BDF^{+}14, BHM13, CGAD16, CRC^{+}16, CH16, CD17, CPLH16, CSJ16a, DG17, DCS^{+}11, DDF^{+}12, DHAH12, FAP^{+}17, GQX^{+}19, GDJ11b, HKK^{+}17, Hum17, HFM13, JPPY11, Jok16, KLA19, Kam14, KOC^{+}16, KBT14, KVA18, KGH^{+}16, KRD^{+}18, LBG18, LGB^{+}18, LVSF17, LPSF19, LAD^{+}17, LBKTW19, LNH^{+}15, MMO16, Mac17b, MGL^{+}15, McE17, OSSL16, Pac18, PSH13, PBE^{+}19, PFD^{+}16, RGAS^{+}18, RAH^{+}16, RRCT^{+}17, RB10, RPB16, RDKK15, RCH^{+}19, SAB^{+}16, SG18b, SFCS^{+}17, SD17, SS12, SFD^{+}16, TIS^{+}19, TJR^{+}10, THW16, TMP12, VGB^{+}17, WZZ^{+}16, Wil11, ZSC16, ACC^{+}12, AWS^{+}13, BJH^{+}14, BMFSH12, BA12, BMH^{+}16, COM11, CAM^{+}15, CSC^{+}13, CLMP16, DMJ^{+}14, Dav11, FFF16]. Ocean [FFN^{+}13, FMC10, FWT^{+}14, FS14, GMD^{+}11, GDBM^{+}17, GRG^{+}10, HGF^{+}18, HLS^{+}15, JPPY11, KT18, KCA^{+}14, KKZ^{+}12, KDBP11, KN17, LLL^{+}11, LCMR^{+}18, LLST15, LSCL18, MAA^{+}19a, MAA^{+}19b, MDP^{+}15, MMAS11, PDS15a, PDS15b, PE18, PKR^{+}19, PHB^{+}16, RBCH10, SLTL15, SLH^{+}17, SSP^{+}11, SSP^{+}13, TRSM15, TNS13, TSS^{+}13, TPS^{+}11, VCRPS13, VSP^{+}14, YYCC16, YCZY19]. ocean-colour [Wil11]. oceanic [LCM10, MFT16, PGV^{+}17, SDW19, TDN^{+}19]. oceanica [DTL^{+}11]. oceanographic [BvdMA^{+}18, BGA^{+}19, KKZ^{+}12, NC12, VCRPS13, ZHD10]. oceanography [ABHT^{+}16, Har14, HSB16]. Oceans [RBCH10, BKVT16, DPE^{+}17, HBG^{+}19a, HBG^{+}19b, LBM19, MKB^{+}17, SPM^{+}18, EGC^{+}17, LKG^{+}19, ZGTL13]. ocellatum [HWR^{+}16]. ocellatus [AS14, LEE17]. octocoral [EMW^{+}16]. octofasciatus [WNM^{+}13]. octopods [BA10]. octopus [ALMFFBP19, HRF^{+}10, PHH^{+}16, RFMR10, SPE14, BMR10a, BMR10b, ESR^{+}10, GCG^{+}10, GRG^{+}10, HMHL19, HRF^{+}10, SPE14, EHT17]. octopus-catshark [PHH^{+}16]. octopuses [DWPS11]. odontocete [HCM^{+}15]. odontocetes [MAM^{+}15]. odour [SJ16]. off [ACA19, BM12, BS13, BCKA11, CC17, CAC^{+}18, CLdR^{+}15, CST^{+}19, DCHMHQ12, FCCLA10, FOT^{+}17, GMH^{+}14, GLC15, GS11, GHB^{+}10, HD11, HKS12, JBE14, KCK14, KCP^{+}11, KR12, MG15, MS10, MBP19, NST^{+}14, RHP^{+}12, dÁMNGS^{+}11, TDN^{+}19, VBW^{+}18, VBA10, VHH10, YYCC16, ZOQS10, dPLF^{+}19]. officinalis [GRC^{+}14, LLLB^{+}10, OBJA^{+}18, SLF^{+}16a]. offs [ECFL15, EBSW15, FR16, HDG^{+}17, JLH12, KEM^{+}17, LPD^{+}16b, MS10, Pun17]. offshore [AIA13, BTL17, LSB^{+}10, MSL^{+}11, SG10, SJRB18, TMW19, ZHD10]. offspring [EAOB15, LFMJ19a, LFMJ19b, RBA^{+}19, ROK^{+}18]. OGCM [LSI^{+}15]. ogive [FWD15, JTJ18]. oil [BM12, BOP^{+}19, MGVH^{+}10, MO13, RLQ^{+}15, RKCP18, SLT^{+}14, SDW19, SHW^{+}18, VRL^{+}14]. oil-spill [BOP^{+}19]. Okhotsk [Kim12]. okutanii [KSZ10]. old [GÁE16, HJS14, SH15]. old-growth [HJS14]. olfactory [KTLB16]. oligotrophic [LBM19]. omega [BDF^{+}14, CSJ16a]. omega-3 [BDF^{+}14]. omission [WNM^{+}13]. Ommastrephes [Che10, YCZY19]. Ommastrephidae [RYS11]. omnidirectional [SWBJ13, TK18]. on-board [HMP^{+}15]. Once [PRB^{+}15]. Oncorhynchus [CHD^{+}18, KBT14]. ongrowing [ESR^{+}10]. online [vdHdGL16]. only [ZPS^{+}18]. Onset [LLST15]. Ontario [SR16]. ontogenesis [PKHG14]. Ontogenetic [IKF17, LWS^{+}18, HBD^{+}19a, HBD^{+}19b]. Ontogenic [SV10]. ontogeny [MOVU10]. Oocyte [FHC^{+}12, DPRG^{+}18]. oocytes [GMN^{+}17]. oogenesis [LSF16]. open [CD17, KHH^{+}17a, RDKK15]. open-ocean [CD17, RDKK15]. Opening [BNB19, dlPG15]. operated [CBD10b]. operating [PDS15a]. operation [HBC^{+}11]. operational [AGSPH14, ABHT^{+}16, CKA^{+}17, GFF^{+}17, SMR^{+}11, TLR^{+}17]. Operationalizing [DGL^{+}17, LB17, HKS17, PB14]. Operations [GIDP^{+}18, QHG^{+}12, TKJB19, Wil11]. opilio [HCR13, MDMC11, OSM18, Urb15]. opportunist [Bak14]. opportunistic [BHD13]. Opportunistically [vdKFS^{+}16]. Opportunities [SMOH18, BPF^{+}18a, BPF^{+}18b, CMA^{+}16, MRD^{+}12]. optical [KRML11, KRKG16, MYI12, MSL^{+}11, RK16]. optically [OOD13, Ohm19]. Optimal [LHV^{+}16, WWCO15]. Optimization [LCS^{+}15, DCG10, DGC12]. optimize [SHG^{+}18, ZED11]. optimized [ZPS^{+}18]. Optimizing [RD16, STM19, TRH10]. optimum [RKJ14]. options [SRB^{+}14, TDR^{+}16]. oral [OdS19]. orange [LGR^{+}19, MKR13, RK16]. orange-spotted [LGR^{+}19]. orbits [SSB^{+}10b]. orca [PDS15b, PMH^{+}13]. orchestrate [Dek16]. Orcinus [PDS15b, PMH^{+}13]. Öresund [LWN^{+}13, PBS14, SWS13]. organic [BJH^{+}16]. organism [Man17b]. organisms [COM11, Ohm19]. organization [MGVH^{+}10, SRS11]. organizations [HHMF19]. organized [Bro16]. Oria [ADB16]. orientalis [SLH^{+}17]. orientated [HNAK12]. orientation [KOC15, MBQ10, dL14]. Origin [FAR15, OEG^{+}16, BHL^{+}15, GCS^{+}18, JHB^{+}12, JKS^{+}18, PDZ^{+}13, SLKS10, VCF^{+}12]. ornatus [YJBL17]. orthogonal [PDH^{+}14]. Oscillating [HCE^{+}11]. oscillations [VGWJ11]. Osmotic [dBWMD^{+}14]. OSPAR [ASV^{+}19]. ossicles [GBM^{+}19]. Otaria [VOS10]. other [COM11, FCG^{+}16, LFMJ19a, LFMJ19b, MB14, Obu18c, Obu18d, RGS^{+}16, SFCS^{+}17, SD17, dQCD^{+}10]. Otolith [DGPC^{+}15, MMDSP15, MNG15, RHHH^{+}12, vdSSM^{+}18, BJH^{+}14, BTA^{+}18, BWR^{+}15, Cam18, CSS^{+}12, FJSJ15, GNDC11, HA15, HGH^{+}16, IKF17, KIA^{+}18, KSJ14, LPD^{+}16a, MIM^{+}19, RDB^{+}18a, SBD^{+}15, TJMC11, WOW^{+}17, WNW^{+}15]. Otolith-derived [MMDSP15]. otoliths [GHF^{+}11, PBW15, TNH^{+}10, WPH^{+}17, WLN^{+}13, ZGT13, dBWMD^{+}14]. otter [CCG14, DUM^{+}12, MBS^{+}15, MSM17, PHA^{+}10]. our [Dek16, GBB16, KCK14, RMJF14]. outbreak [PGS^{+}19a, PGS^{+}19b]. outcomes [GHPH15]. outer [BNP19, WKL13]. output [OBJA^{+}18, PCdL15, ROK^{+}18]. output-based [PCdL15]. ovarian [OTAK15]. over-quota [PBQ^{+}10]. Overcoming [MBT^{+}19, GPS^{+}17]. Overconfidence [BNAH13]. overexploitation [Lor11]. overexploited [MBLS15, NGTT16]. overfished [MHSG19, WPeEA16]. Overfishing [MG15, Cor12, GAB^{+}14, MTLG14, RKJ14, SML14, Sum13, VEA^{+}10]. overflowing [DPL10]. overlap [GW16, HF14, KDH^{+}10, PHH^{+}16, TMR^{+}16, VKS^{+}12, VRL^{+}14, VE13]. overlaps [VRF^{+}16]. overturning [Kam14]. Overview [Cha12, COL^{+}13, AB16, GGP11, Wel11]. Overwinter [HTT^{+}17, AS14]. oxygen [BFC^{+}16, CYLT16, HHM^{+}11, KDF^{+}19b, RAH^{+}16]. oxygeneios [WNM10]. oxyrinchus [CJFS16, TSPL14]. Oyashio [YCY^{+}13]. oyster [BWM^{+}16, GOH^{+}19, HTM11, RC15, RST18, TBC^{+}14]. oysters [PGS^{+}19a, PGS^{+}19b].

P. [LPSF19]. Pacific [JPPY11, KKZ^{+}12, KDBP11, PPHK19, PE18, SSP^{+}13, YYCC16, AMQRC10, ASB^{+}11, ASI^{+}16b, BWR^{+}15, BW14, Bow14, CPSCCE^{+}17, Che10, CCS10, CPP17, Dav11, DZC^{+}13, DZ14, DMO10, FRF14, FS14, GMC17, GSC10, GMKS11, GFML10, Hol10, HMM12, HCDAF15, HLMT16, IaF11, iIOKW13, IEL^{+}15, KPJC14, LCMR^{+}18, LFMJ19a, LFMJ19b, Loh11, MG12, MARD13, MW19, MTO^{+}18, NUO^{+}10, NS17, NC12, OIMP18, PGV^{+}17, PCB^{+}14, PWPW18, PDWH11, RTAT17, SBFC10, SLH^{+}17, SSF^{+}18a, SSP^{+}11, TSP12, TNS13, TRPH16, TBC^{+}14, TSPL14, TSY^{+}11, TSS^{+}13, VF17, VCRPS13, WHL^{+}17, WKK^{+}15, YST^{+}10, YJBL17, YCZY19, ZGTL13, ZED11, ZD13, ZD14, dGGW^{+}11]. pacifica [CPP17]. pacificus [KGN^{+}10, RYS11]. paddle [HZZ^{+}15]. Pagellus [Lor11, PDCP14]. Pagophilus [FFH^{+}11, SS10b, SBKA16]. pagurus [BMVPN10, BHS11a, CR11, HBB^{+}16]. pair [IZP^{+}16, KSJ14]. paired [McB15, Mil10, TKJB19]. paired-trawl [Mil10]. pairtrawlers [FCCLA10]. Palamós [DCVC15]. Palmaria [MMC19, NMF^{+}16]. palmata [MMC19, NMF^{+}16]. Pandalus [EMP11, JTJ18, KJG^{+}15, OFYS^{+}11]. panel [HWK^{+}15, SHG^{+}18, vdHdGL16]. panels [SLV16, dlPG15]. panmixia [PSDG12]. Panulirus [BFDGLÁ15, BMG15, CHM15, FFN^{+}13, FFN^{+}16, GBM^{+}19, GBB15, PCdL15, YJBL17, dL14]. PAP [HLL^{+}12]. PAP-SO [HLL^{+}12]. paper [RS10]. Papua [HTRF17]. parabolic [WPH^{+}17]. Paracentrotus [CRC^{+}16, OFVF14]. paradigm [TH15]. paradigms [SM15]. paradox [CN14, RPM14a, RPM14b]. paradoxus [GES^{+}12]. paralarva [SV10]. paralarvae [GOPG10, MRC^{+}10, RGPG10, VHH10]. Paralichthys [KSJ14]. Paralithodes [DD13, FPRA11, GMKS11, Hje14, HKS12, LPSF19, SLF17, WHNS14, Win15]. Parameter [Sub18a, Sub18b, BMVPN10, WOW^{+}17, dlPG15]. Parameterizing [PB14]. parameters [BMGB19, GMH^{+}14, HTB^{+}16, KPS^{+}16, Sas19, SG16a, SS10b, Yar10]. parametric [Cad13, CIJ18]. paramount [DVV^{+}11]. parasite [HdBR^{+}16, MLC^{+}19]. parasites [GMD^{+}11, UPF^{+}10, ZGTL13]. Parasitic [Mac17b]. Parasitism [FMLM^{+}14, RHOL11]. Parent [LFMJ19a, LFMJ19b]. Parent-offspring [LFMJ19a, LFMJ19b]. parentage [CMG^{+}17]. parental [BMG15, PVCB17, ZD14]. park [LPD^{+}16a]. parr [SRKV15]. Part [EPK^{+}18, LSF16, SDAM^{+}18, SLF16b, JV12a, URE^{+}18, VVP^{+}14]. Partial [MHSG19, TDN^{+}19, LWS^{+}18, VWMS19]. Participatory [MVK18, SGV18]. particle [BGS^{+}14, BPST^{+}16]. particles [BAFRJ18]. particulate [vDKR14]. partitioning [LHJ12, MB14, dSMGKP12]. Pass [MMAS11]. passability [TCD^{+}16]. passage [WABZ16, STF^{+}16]. passed [Abl16]. passing [RB14]. Passive [LAG^{+}16, OSL^{+}15, TMS^{+}15, TBC15, TOA^{+}16]. past [Hin15, MO18, NRTK^{+}19, OTO^{+}15, RKB^{+}12, VQGDAMdL14]. Patagonia [NST^{+}14, TDN^{+}19, VHGTFR10, dlTQMUE10]. Patagonian [DRVV^{+}17, GLP^{+}11, GTGD15, HLCC16, NST^{+}14]. Patch [HJT12, KBB13]. Patchiness [STR16, GCBI17]. Patchwork [RLQ^{+}15]. pathogens [KPBB15, UPF^{+}10]. pathway [APHC15]. pathways [GDBM^{+}17, HG10, LBM19, MHSG19]. Patos [MAB12]. patronus [ALWH18, RRH^{+}15]. pattern [APB^{+}10, UBV^{+}11, YLLG18]. Patterns [CKD^{+}17, VELT14, WKL13, BGM^{+}14, BCS^{+}17, BMC^{+}13, BFB^{+}11, CR19, CPSCCE^{+}17, CCC^{+}14, CN16, DHZA14, DCHN^{+}15, EPK^{+}18, FFH^{+}11, FHH^{+}13a, FHH^{+}13b, GGTW17, GDS15, GKJ^{+}15, HMS^{+}13, HFNH13, HBD^{+}19a, HBD^{+}19b, HMM12, HFSV^{+}15, ISL19, KOK^{+}14, LFM^{+}18, LKS^{+}14b, LMG^{+}14, LHH^{+}13, MPR17, MGBvD14, PB11, QCA^{+}18a, QCA^{+}18b, RHHH^{+}12, RDD^{+}10, RGFT14, RRCT^{+}17, SV10, SDAM^{+}18, SAS^{+}18, SFN^{+}19, SIP18, THSM^{+}18, TBC15, TCQ^{+}15, TPHC15, UBvH11, UvHS^{+}14, WDE^{+}18, WKW^{+}10, YPN^{+}18]. Paul [GRG^{+}10]. Pauly [GGT^{+}14]. PaV1 [CZBFLÁ^{+}15, KPBB15]. PaV1-infected [CZBFLÁ^{+}15]. PBSAT [GHV^{+}18]. pCO [BTB^{+}17, BIKP^{+}17, BSHC17, CCE17, CPP17, DKPM16, EMW^{+}16, GM17, KPGH16, KPD^{+}16, LSF16, LEE17, PWCS17, RCF^{+}17, SLF^{+}16a, SMB^{+}16a, SJ16, SLF16b, VSB^{+}16, WWMF17]. PDF [Ano15, Ano16b, Ano17, Ano14a, Ano16a]. peaks [KS18b]. pealeii [BNP19]. pearl [HTM11]. Pecten [MLMS15, NMO^{+}17, OTM^{+}16, SMB^{+}16b, THR16]. Pedicularia [BHCSP^{+}11]. Pediculariidae [BHCSP^{+}11]. Pelagia [BSK^{+}11]. Pelagic [ABHT^{+}16, Ano19, DCMGB^{+}17, RBA^{+}19, RDB^{+}18b, BI13, BTL17, CJFS16, CTH^{+}19a, CTH^{+}19b, DL12, DBL^{+}16, Eri16, EAOB15, GNKS18, GCBI17, HCM^{+}15, HPS^{+}11, HSHÖ14, HPN^{+}17, HLA^{+}11, JCLO17, JVPM12, Kam14, KOC^{+}16, LAD^{+}19, LJB16, LEB^{+}14, LLPV^{+}12, LK15, MNV^{+}11, MAM^{+}15, MHLW12, MWP^{+}16, OSM18, ÖPPM19, PPHK19, PDS15a, PDS15b, PCFR13, RBM15, RB10, RHP^{+}12, RS10, REFJ12, SRCR12, SRDC^{+}14, SWBJ13, URE^{+}18, UHB12, VVP^{+}14, ZH13, dMBD11]. pelagics [VFR^{+}16]. pelamis [BMS^{+}18]. pen [GGIC^{+}15]. Penaeidae [DI11]. Penaeus [KS17]. penguin [BS13, RBP15]. penguins [LTT^{+}18]. Peninsula [BS13, BFB^{+}11, VBW^{+}18]. Penobscot [NSS15]. pens [BPF^{+}18a, BPF^{+}18b]. people [DMS18]. Peprilus [Ada17, JLL17]. Perca [ZRN18]. perception [DMO^{+}15a, KCK14]. Perceptions [vPBLR19, BV19b, BV19a, CS15, Coo19, HMFW14, KP13, MLS^{+}17]. perch [HLL18, ZRN18]. perfect [MFT16]. Performance [BPvHR10, CKB^{+}16, CCS10, PKR^{+}19, TFC^{+}19a, TFC^{+}19b, WSGD^{+}17, DHS19, FLLG15, GPP17, HHMF19, HABV19, Hil19a, Hil19b, HOV^{+}15, HHT13, LBG18, MSOF18, OYIN18, OCWG16, OHD^{+}16, PAB^{+}14, QHG^{+}12, SIP18, TWHB19, THHH15, THHH18, WNW^{+}15]. performed [BF17, Man16, Man17a, SHS11]. period [AFP12, DK18, Eri16, Ósk18, PRKASR12, PVCB17, PT19, RMJF14]. periodic [MGBvD14]. periods [HHM^{+}16, Loh11, tHR11]. permanent [BA10, BSP15, LGvPH15]. Persistence [HAR19, JFJ^{+}17]. Persistent [KGH^{+}16, BdSP11, FMML19, Hol10]. persistently [PCBO^{+}18]. personal [Ban12, Koe11]. perspective [Bro17b, DB16, DPL16, LKS^{+}14c, Llo17, RDB^{+}18b, SBS19a, SBS19b, SPS11, ZHV^{+}16]. perspectives [EAB^{+}17, FOC^{+}18, HHZ^{+}18, KCA^{+}14, MMO16, Pep16, RG11, SFG^{+}15, SDSN14, SH15]. Peru [KKAV^{+}19, YYCC16]. Peruvian [ZHD10]. petrification [Cam18]. petroleum [BV19b, BV19a]. pH [BMM^{+}19, CYLT16, CGK^{+}17, EWH16, HTRF17, KTLB16, RKW^{+}17, SSM^{+}16, SLF17]. phase [CMC^{+}12, CTT14, MMWC16, WABZ16]. phases [DPD^{+}12]. Phenological [TFM^{+}12]. phenology [LF14, MM17, ME11, RTAT17, SLTL15, ZPT^{+}11]. phenomena [MDP^{+}15]. phenomenon [KHMS19]. phenotypes [CCSG18, KH19]. phenotypic [TFM^{+}12]. philippinarum [LK17]. Philippines [NC12]. Phoca [DTA^{+}13]. Phocarctos [HB19a]. phocine [DTA^{+}13]. phosphatase [LLLB^{+}10]. photogrammetry [KOC15]. photoperiod [KT18]. photophysiology [BIKP^{+}17]. photosynthesis [BWH16, CCE17, KPS^{+}16]. Phyllosomata [RLW^{+}15]. phylogeny [BPKH13]. Phylogeographic [DMF13, VPSV^{+}10]. Physeter [GLP^{+}11]. physical [BPST^{+}16, DID^{+}16, GGTW17, OI16, TNH^{+}10, VHS^{+}10]. Physiological [HP11, LFH^{+}19, LBJ15, RFN^{+}19, ZSC16, BCL^{+}11, CBC^{+}11b, JGCH18, MSM17, ODSC10, RS10, SMB^{+}16a, SSA^{+}18]. physiology [BBHC11, WWMF17]. phyto [DRVV^{+}17]. phyto- [DRVV^{+}17]. Phytoplankton [FPSF11, MDP^{+}15, ZPT^{+}11, BL15, ET15, HFSB^{+}14, Kan11, KPS^{+}16, LGBA11, LSF^{+}15, LLST15, RRCT^{+}17, RPS11, SJ15, TCQ^{+}15, UFLH^{+}13, VKS^{+}12, WAH^{+}16]. PI [WDG^{+}17]. picture [BG18]. picturing [TLK^{+}17]. piece [Mac17b]. pig [KM16]. pikeperch [HLL18]. pilchardus [CMU^{+}17, MMA^{+}10, MOVU10, SGQR^{+}12, SSB^{+}10b, TID^{+}10]. pile [FWC^{+}19b]. pile-up [FWC^{+}19b]. pilot [PSH^{+}11]. pink [aFKA11]. pinniped [HB19a]. pipelines [RKCP18]. piscatorius [CSS^{+}12]. Pisces [GMD^{+}11, OBLP^{+}19, WNM^{+}13]. piscivore [EBSW15, OE16]. piscivorous [BLS^{+}18]. Pitfalls [EP14, HA15, SPRL18]. place [CBK18]. placing [ETM^{+}16]. Placopecten [KHH^{+}17a, SPS15]. plaice [EPKR11, HMS^{+}13, LVM^{+}11, LVPK10, MSM17, PvdKE^{+}12, RvDW10, WKW^{+}10, vdRMC^{+}17, vdSSM^{+}18]. Plain [HLL^{+}12]. plan [BNK10, DG11, DL11, JCA10, KZS10, Str10]. Planktivorous [KT18, KGW^{+}12, SEDO19]. Plankton [RPM14b, ZHL^{+}19, BR12, BJMG19, BÁOMRV13, CDSP11, CN14, GFF^{+}19, JDFN12, MYI12, MGL^{+}15, RPM14a, TMP12]. planktonic [BSMB18, FRF14, MNV^{+}11, Ohm19, RFT^{+}16]. planning [ATLC16, AHK^{+}18, DMS^{+}12, EPH15, JLH12, KS14, Por11, YS15]. plans [AR10, DCG10, LKHK11, POIM12, Par13, SDDA11]. Plastic [FBL^{+}16, NDB^{+}16, ELBS17]. plasticity [RCG^{+}18]. Plateau [SO10, SMR^{+}10]. platessa [LVM^{+}11, MSM17, WKW^{+}10, vdRMC^{+}17, vdSSM^{+}18]. platforms [SLT^{+}14]. Platichthys [HPN^{+}17]. platypus [LPSF19]. play [Mat11]. plebejus [OLW^{+}14]. plei [GRP10, PG10]. Pleuronectes [LVM^{+}11, MSM17, WKW^{+}10, vdRMC^{+}17, vdSSM^{+}18]. plocamia [QCA^{+}18a, QCA^{+}18b]. plot [Hil19a, Hil19b]. plots [BKSB12]. plume [LEB^{+}14, OSP^{+}13]. Poey [HLS^{+}15]. point [DPL16, GML13, HLL^{+}12, LM10, RCH^{+}15]. points [ABRL12, BPC10, Cad13, CAOG15, DGC12, GCJL16, GPS^{+}17, GFF^{+}19, Har13, HBB^{+}13b, LB13, Mac12, MR10, MRD19, MSR14, RHSH16, WÓG^{+}16, WMG^{+}17, ZCW11]. Polar [RB10, DHAH12, FRF14, ÖPPM19, SSS^{+}12, TRHB13]. poleward [CCL^{+}14, YFP^{+}19]. policies [HJS^{+}17]. Policy [HTKB19, VQB^{+}11, WFM^{+}15, CKA^{+}17, DVV^{+}11, ETM^{+}16, FR16, FO13, Gro11, Mer18, RSGM17, SBB^{+}17, TMW19]. Polish [FCB17]. political [KS14, TGH^{+}12]. Pollachius [AJ12, CLV^{+}14, GHV^{+}18, íHHJ^{+}13, OS14, SHK^{+}15, SO16]. pollack [HSNO12]. Pollicipes [JCSC10, POIM12, Par13]. pollock [BHD13, DWW^{+}10, DW11, GSC10, HFM13, IHH^{+}11, IS15, KLCC18, KCL18, KHPI15, LCSI18, MBIH11, PH12a, SHZ^{+}16, WPWH11, WMJ13, WWI^{+}16, PDAC^{+}16]. polykrikoides [KLN^{+}10]. polymers [ELBS17]. polymorpha [CAMM12]. polymorphic [GCS^{+}18]. polymorphisms [SHK^{+}15]. Polynesia [GVA19]. Polyprion [WNM10]. polyxystra [CDAN^{+}14, CN16].

Pomatomus [BLS^{+}18, SMR12]. pooling [SDH12]. Poor [RMJF14, AH15, APV18, BT15, BPC10, CST^{+}19, GB15b, HOV^{+}15, JHB^{+}17, LGR^{+}14, MHM^{+}11, MR14b, PHV^{+}15, PSS11, RTS19, STM19, TCF^{+}13, WNW^{+}15, ZPS^{+}18]. pop [CAM^{+}15, HLA^{+}11, LFM^{+}18, STC^{+}17]. pop-up [CAM^{+}15, HLA^{+}11, LFM^{+}18, STC^{+}17]. Population [BJH^{+}14, CBD^{+}10a, CMP^{+}11, CRHM19, CHD^{+}18, DD13, HB19a, Lav15, MAP^{+}11, PHB^{+}16, VMG11, YJBL17, ABJ^{+}19, ALMFFBP19, BSHE18, BHJ14, BPF19c, CC17, CMH19, CTH^{+}19a, CTH^{+}19b, CTT14, CMG^{+}17, DJW^{+}15, DCHMHQ12, Eer12, FFN^{+}13, FFN^{+}16, FSC^{+}12, GSL^{+}11, GMKS11, GÁE16, GHG^{+}10, GT19, HCC^{+}14, HSNO12, HRB^{+}15b, HMFW14, IEL^{+}15, JTE19, JQD^{+}17, JOSO16, JKS^{+}18, KBC^{+}12, KCS10, KHC^{+}17, KJG^{+}15, LFGW13, LPH^{+}15, LLM12, LÁG^{+}19, LVPK10, MMRG18, MWP^{+}15, MSOM10, MAS19, MSR14, NUO^{+}10, OYI17, Ósk18, OFVF14, OTM^{+}16, PGG^{+}12, PDZ^{+}13, PGV^{+}17, PCRD18, PDD^{+}10a, PVD14, PB14, PSF12, RHOL11, RC15, SHW^{+}15, SBH^{+}14, SKA15, SKA22, SSF18b, SR16, SBS^{+}19c, SH16, TRH10, VCRPS13, VSP^{+}14, WKW^{+}10, WKK^{+}15, WSG^{+}17, WÓG^{+}16, WMG^{+}17, WSW^{+}19, WCR^{+}19, ZHD^{+}14, ZMF12, dPJGB13, dPLF^{+}19]. population- [PVD14]. population-level [SKA15, SKA22]. population-of-origin [JKS^{+}18]. Population-specific [CHD^{+}18]. populations [BHS11a, BHFH14, CBS15, DKK15, DEH^{+}19, FWRM12, GNDC11, GMKS11, GR15, HSB18, HHJB14, HJS14, Hol10, KDF^{+}16, KKC19, Kim12, LTQ14, LBB15, LNH^{+}15, MTP^{+}12, MAA^{+}19a, MAA^{+}19b, ME11, NNY17, ØS15, OAT^{+}12, OFYS^{+}11, PBL11, Rod10, RCC^{+}14b, SLG^{+}15, TFM^{+}12, TVJ^{+}14, TMP12, VE13, ZKD^{+}14]. porbeagle [BCS^{+}17, CJFS16, SRC11]. porbeagles [CRHM19]. porcelain [PWCS17]. Porcupine [HLL^{+}12, PGN^{+}11]. Porites [BIKP^{+}17, BWH16]. porpoise [WDE^{+}18]. porpoises [STB^{+}11, VG15]. ports [DHNL12]. Portugal [CCA^{+}18, LTQ14, Pin17, BMVPN10, JCSC10, SPE14, ZOQS10]. Portuguese [CMM^{+}15, MKF^{+}15, OCG10, OCG13, OCWG16, SPFM^{+}10, SWRC^{+}18, VMG11]. Posidonia [DTL^{+}11]. position [CFHC10, DPD^{+}12, HWK^{+}15, LD17, OBLP^{+}19, SHS^{+}18]. positioning [Peñ19a]. positive [GM17]. possibilities [Fri19, Sca18]. Possible [OS14, SiTiM^{+}11, Lev11]. Post [BBHC11, FDR^{+}17, CJFS16, CCD^{+}19, FWH^{+}15, HUT^{+}11, HLA^{+}11, JÓT^{+}12, KKMS14, KVFK19, LSWD12, LSF^{+}18, MRD^{+}12, MA13, MHE^{+}16, MLMS15, MGH^{+}12, OEK^{+}18, PDD^{+}10a, RPM14b, TBG16, VHF^{+}17, WS13]. post-hatching [PDD^{+}10a]. post-paradox [RPM14b]. Post-release [BBHC11, CJFS16, FWH^{+}15, HLA^{+}11, KVFK19, LSF^{+}18, TBG16, WS13]. post-settlement [KKMS14, MA13, MLMS15]. Post-smolt [FDR^{+}17, CCD^{+}19, HUT^{+}11, LSWD12, MGH^{+}12]. post-smolts [JÓT^{+}12, MRD^{+}12, OEK^{+}18, VHF^{+}17]. post-spawn [MHE^{+}16]. pot [NHL^{+}19, SPM^{+}17]. pot-fishing [SPM^{+}17]. Potential [ASB^{+}11, BHM13, DCS^{+}11, Gar11, LK15, OHLK19a, OHLK19b, QR15, SKA15, ASI^{+}16b, AFQ^{+}11, Bak14, Ben15, BNP19, BT10, CAC^{+}16, CDSP11, CSF^{+}17, CBK18, CKV^{+}16, DSL14, DRB10, ECW^{+}17, FKH^{+}16, FBM^{+}18, GRC16b, Ham14, HG10, HHM^{+}11, Hje14, HOV^{+}15, HOS^{+}15, HHHE16, JQD^{+}17, KCA^{+}14, KMK10, KRO^{+}16, LK17, LB13, LKHK11, MKD^{+}18, MKHK18, NUO^{+}10, NS17, OSBG16, PGV^{+}17, PCFR13, PBG14, RLQ^{+}15, RvDW10, RCF^{+}18, RRM^{+}12, RTAT17, RCC^{+}14b, SPRL18, SSRT10, SSF^{+}12, TLPS15, TSY^{+}11, VCF^{+}12, ZOQS10, SKA22]. potentially [LGBA11, SDH^{+}18]. pots [AFH^{+}17, COR15, HUPBL18, KFL^{+}15, MSR^{+}19]. potting [SMS^{+}17]. pout [BNB19, CKV^{+}16, NWM^{+}12, NLB^{+}12]. poutassou [PGOM11]. poverty [Ben15]. power [HTM11, MMRG18, TPBL17, ZBE^{+}15]. power-station [TPBL17]. Powys [ZSC16]. Practical [SBB^{+}17, GGP11, JdFBB^{+}10, KHC^{+}17, RGS^{+}16]. practice [CLT^{+}14, Gar11, HR14, ZGK^{+}17]. practices [DL11, EPVC14, TGG^{+}15]. pragmatic [vH13]. prawn [DPD^{+}15, GD15, KS17, OLW^{+}14, PDD^{+}10b, ZBE^{+}15]. prawns [DPD^{+}15]. Pre [JTM^{+}16, SSM^{+}18, URE^{+}18]. pre-capture [URE^{+}18]. Pre-release [JTM^{+}16]. pre-trained [SSM^{+}18]. precaution [Har13]. precautionary [GNFM11, HS10, LKS14a, LLCJ10]. Precision [ACC^{+}12, McB15, TSWS15, WOW^{+}17]. precocious [OEK^{+}18]. Predation [DHH^{+}14, GHM^{+}16, AIA^{+}15, CPT^{+}10, GÖN^{+}12, KKMS14, MRD^{+}12, MLMS15, NKSE14, PvdKE^{+}12, SFW^{+}12, SBLS15, SHZ^{+}16, TH15, TDHC18, VOS10, VHS^{+}11, dlTQMUE10]. Predator [CTT14, ALB^{+}19, BLS^{+}18, CFHC10, CKV^{+}16, EBSW15, FOT^{+}17, GLK^{+}10, GW16, HUT^{+}11, IZP^{+}16, KDH^{+}10, LAG^{+}16, LWS^{+}18, OHLK19a, OHLK19b, SJ16, TB16, XPC11]. predator-prey [FOT^{+}17]. predators [Der18, EPR^{+}14, FDR^{+}17, KEM^{+}17, LGL^{+}12, LLPV^{+}12, YPN^{+}18]. predatory [LSB^{+}10, MDV^{+}15, RGC^{+}19, SBSE14]. predict [AFQ^{+}11, FS14, Jon14, LVM^{+}11, SMH^{+}19, SSP^{+}13, TMR^{+}16]. Predictable [HFGT18, WCT^{+}16]. predicted [SPdJFMS17]. Predicting [BOP^{+}19, BPD^{+}18a, BPD^{+}18b, GCJL16, Jok16, MLLL11, TMP^{+}17, ZED11, AMS^{+}12, ABHT^{+}16, FBM^{+}18, KDH^{+}10, MMA^{+}10, RH17]. Prediction [KVFK19, DPD^{+}15, DHAH12, LSI^{+}15, MTO^{+}18, NdlPIR18, PMOH13, SR17]. predictions [BSF^{+}19, CMH^{+}11, FFK^{+}11, PMOH13, TWHB19]. predictive [BBD^{+}13, DMO^{+}15b, THHH15, THHH18]. predictors [HCR13]. predicts [MARD13, VHF^{+}17]. preference [RMJF14]. preferences [SDBB15, SCC^{+}13]. pregnancy [WVG^{+}13]. prehistory [SLG^{+}15]. preindustrial [JCS16]. preliminary [GRC16a]. preparation [WOW^{+}17]. preparations [BA10]. prepared [HMLKR18]. Preparing [LBCOJ19]. Presence [BWH16, KKLM13, MS15]. present [BJR17, Hin15, KJG^{+}15, MAA^{+}19a, MAA^{+}19b, MDV^{+}15]. present-day [MAA^{+}19a, MAA^{+}19b]. Prespawning [VBO^{+}15]. press [HKB^{+}18]. pressure [CJC^{+}15, EBB^{+}16a, EBB^{+}16b, GBJ^{+}15, MPBH11, PH12b, PSF12, SMS^{+}17, tHR11]. pressure-based [GBJ^{+}15]. pressures [KJW^{+}18, LFFL13]. presumed [SS12]. pretty [RCS^{+}17, RDL^{+}17, TJD17]. prevalence [MDMC11]. prevalent [ZH13]. prevent [FDC13, GAB^{+}14]. Preventing [SML14, MTLG14]. prevention [SDT15]. Prey [SCC^{+}13, ALB^{+}19, CTT14, CFHC10, CSF^{+}17, EBSW15, FLP^{+}13, FOT^{+}17, GLK^{+}10, GSP12, GGV^{+}18, GHB^{+}10, GW16, IZP^{+}16, KUM^{+}12, KDH^{+}10, KVSV14, KKLM13, LAG^{+}16, LLPV^{+}12, LBK17, MHLW12, MOVU10, NDP^{+}16, PB11, RMJF14, RGPG10, SRH^{+}14, VRF^{+}16]. prey-related [MHLW12]. price [BM12]. prices [ACS15, MCOS17]. pricing [ZH13]. primarily [PA18]. primary [AML^{+}19, BA12, FPSF11, HSFiS12, JPPY11, MOG^{+}14a, MOG^{+}14b, TMP^{+}17, VHH10]. primer [GB15c]. primnoidus [CSBHS^{+}11]. Principal [PA18]. principle [DBS^{+}15]. principles [Jen13]. Prionace [CJFS16, CMH^{+}11]. prior [Ham15, KS18b]. priorities [MRP^{+}16, YS15]. priors [FWC^{+}19b]. Probabilistic [SLKS10]. Probabilistic-based [SLKS10]. Probability [TJH15, VKN^{+}11, CMH^{+}11]. Probability-based [VKN^{+}11]. probe [KRKG16]. problem [GBB16, MCAM14, SFCS^{+}17, SD17]. problems [AGH^{+}11, BBD^{+}10, CH16, ETG^{+}15, RR10]. procedure [GFF^{+}17, dMBD11]. procedures [CKB^{+}16, GB15a, GB15b]. proceedings [GGK^{+}11, WRC^{+}12]. process [ALBR14, FBFP^{+}13, HMS^{+}13, HHQ11, MJS^{+}15a, VSB^{+}16, ZCR15]. processed [BCSG13]. processes [ASV^{+}19, GHB^{+}14, HSB16, JTM^{+}16, KH19, LRJ18, RFF^{+}14, SRRZ16, SH16, TKJB19, VELT14, VCRPS13, VGWJ11, WCR^{+}19, ZGK^{+}17]. processing [KVFK19, LBLJ17]. produced [PDT^{+}18]. Producing [GTBT18]. product [ACS15, ZHV^{+}16]. Production [BSMB18, DG17, AML^{+}19, BSW^{+}11a, BSW^{+}11b, BHFH14, BA12, BMG15, CN16, DED13, FKH^{+}16, FPSF11, GMM^{+}12, GS14, HLD10, JPPY11, KPS^{+}16, MNV^{+}11, MSMW18, MOG^{+}14a, MOG^{+}14b, MAB12, ØHN14, RLQ^{+}15, Sas19, SHK17, TRSM15, TCH^{+}16, VHH10, WHP16]. productivity [BNE^{+}17, GW16, GFF^{+}19, HABV19, HSFiS12, KPL^{+}19, MSR14, NF16, OCG13, RCG^{+}18, SOB^{+}11, STD11, TRPH16, TMP^{+}17]. products [BCSG13, SCL^{+}19, vdHBSM10]. productus [IEL^{+}15, SQ14]. profile [MWS^{+}17]. profiles [AHK^{+}18, BHS11a, KMK10, KPS^{+}16, LSCL18, WFS^{+}11]. profit [MUEO17]. program [POIM12, Par13]. programme [CAB^{+}10, GUN^{+}19, JTM^{+}16, UBvH11, VP19]. programmes [GWS17, HGS11, KFBG15, MBP19, RCC14a]. programs [HKS17]. progress [CPS13, DHAH12, MKL^{+}19, Mur11, OSSL16, Sum13]. progression [KAT11]. Project [COL^{+}13, SOL^{+}15, HGS11, SS10a]. Projected [HGC^{+}19a, HGC^{+}19b, HBB^{+}13c, LNH^{+}15, PDWH11, ASI^{+}16b, CDSP11, NRTK^{+}19, RBP15, SSTB15]. Projecting [FKH^{+}16, Pla16, CPS13, PBC^{+}16, PBL11]. Projections [MBL^{+}17, Bar19, BNAH13, CFA^{+}16, FCPJ10, JC15]. projects [BA17]. promise [SPRL18]. promising [PDD^{+}10a]. promotes [BLJ^{+}17]. pronounced [BDCvD^{+}10]. proof [FGRR11, Tho11, WFS^{+}11]. proof-of-concept [Tho11, WFS^{+}11]. proofing [JDH^{+}14]. propagating [HHR17]. Proper [JCLO17]. Properties [AST^{+}15a, BW14, CRC^{+}16, FHK14, JH13, WPGW12, WCK^{+}10]. proportional [HFH10]. proportions [BSP15, GUN^{+}19]. proposal [Moo19]. proposed [CKD^{+}17, DSA13]. prospects [HHJB14]. Protected [Ano18p, DOB^{+}17, LLM12, AIA13, BHC^{+}15, Ber18b, DHS19, FVSL^{+}10, GGTW17, GRH^{+}14, HSB18, JDH^{+}14, KKC19, LMPP10, LNWS18, MAGK18, MHM^{+}11, MR14a, NS17, OE16, PAB^{+}18, SG18b, VRF^{+}16, YS15]. protecting [Mos18, SSRT10]. protection [Dek16, Fri19, FMM^{+}12, FDC^{+}16, GR15, HDG^{+}17, SMB^{+}18]. protections [Sca18]. Protistan [LD17]. protists [DCMC19]. protocols [GVA19, HRP^{+}16]. protogynous [DMF13, KS18b]. prototype [HBHR18]. proven [EP19]. provide [BSK^{+}11, CCC^{+}14, FCG^{+}16, KCMS19, LSJ10, MBP19, TWHB19]. provides [AV15]. providing [EAOB15]. provision [RGRL11]. provisions [WHMP15, WHP16]. proxies [PSS^{+}14]. Proximate [SMSR10]. proxy [LB13]. Pseudocollapse [Jan14]. Pseudorca [PDS15b]. Pseudosagitta [KN17]. Pteropod [MG12, BLMW17, TRHB13, WHL^{+}17]. pterotum [Sas19]. Public [KP13]. published [SOGT^{+}19]. Puerto [RBMAOCL18]. puerulus [EF15, LMG^{+}14, dLCF^{+}15]. puffer [TMW19]. Puffinus [Lav15]. Puget [BHM13, SSF^{+}18a]. Pulse [dHFF^{+}16, BHJ14, HKB16, vdRMC^{+}17]. pulse-trawl [vdRMC^{+}17]. pulsed [ZHL^{+}19]. pulses [RU12, SCD^{+}15]. PulseWing [DDE^{+}19]. pup [DED13, HB19a, ØHN14]. Purported [BBD^{+}10, RR10]. purpose [BTL17]. purpuratus [RFN^{+}19]. purse [ACC^{+}12, CHD^{+}18, CLMP16, LCMR^{+}18, MMA^{+}10, MKF^{+}15, MKB^{+}17, PH11, RBC^{+}15, TVO12, TMR^{+}17, TPHdJ19, TRPH16]. purse-seine [ACC^{+}12, LCMR^{+}18, MMA^{+}10, MKF^{+}15, PH11, RBC^{+}15, TMR^{+}17, TRPH16]. purse-seines [TVO12]. pursuit [Har14]. putative [LFM^{+}18]. puzzle [Mac17b]. Pygoscelis [LTT^{+}18]. pyramid [KBvZP16].

Qatari [WHAA^{+}18]. quad [BMC^{+}17]. quad-rig [BMC^{+}17]. quahogs [TJR^{+}10]. Qualitative [RMF^{+}15, CLT^{+}14, Fle15, LPB^{+}14, LAD^{+}11, ZHAG17]. Quality [HBF14, BMG15, DAB^{+}12, OBJA^{+}18, RC15, RTJS14, ROK^{+}18, SMSR10, TRH10, Tow14]. quantifiable [CAOG15]. Quantification [MKF^{+}15]. quantify [SKM18]. Quantifying [ACM^{+}16, AML^{+}19, DVV^{+}11, Gro11, Har13, NUÓ^{+}16, PCRD18, RDB^{+}18a, RH17, RBP15, SR16, VE13, dlPG15, GLK^{+}10, GHB^{+}14, LD17, MJAS14, Pun17, WBW^{+}18, ZHV^{+}16]. Quantitative [SHJI10, SFG^{+}15, CLS^{+}18, CGGH^{+}17, LPB^{+}14, LAD^{+}11, RBB^{+}16, RHP^{+}12, TRHB13, WW16b]. quantities [FS14]. quantity [FB11]. quest [Ger17]. Quirky [DCHN^{+}15]. Quo [Bro12, Bro17a]. Quota [CR19, MLT11, EHG^{+}14, KHF14, KLKD11, LGvPH15, MCAM14, MLMJ17, OCBJ10, PBQ^{+}10, WHMP15]. quotas [CCG14, OCS14, ØHN14, OCG13, PBQ^{+}10, TPL11]. quotas/caps [OCS14].

R [Arn11, Cad13]. radar [KGP^{+}15]. radiated [DWFD13]. radically [EWH16]. radiocarbon [BTA^{+}18]. radiometric [COS11]. radionuclide [BWM^{+}16]. Radionuclides [COM11]. ragworm [SCD^{+}15]. Rajidae [MSE12]. rake [BPBD15]. RAMP [HCR13]. Ramster [Ban12]. ranching [Bjö18]. Random [THM15, FFF16]. range [AGMC14, ALMFFBP19, BCS^{+}17, BLB^{+}18, FHK14, Lap11, MRP^{+}16, RBAB17, RD16, RHOL11, SCD^{+}15, WNM^{+}13, WPH^{+}15, vPFF^{+}16]. ranges [PHK^{+}19, RCS^{+}17]. Ranina [OCBJ10]. Ranking [CSY^{+}10]. Rapid [ZGT13, BCSG13, DCVC15]. rapidly [SMOH18]. rare [RS10, ZMF12]. raschii [MPL^{+}14, MDS13]. rate [BHLS18, CMH^{+}11, DHKE16, DFB^{+}10, GIR^{+}19, Ham15, HTB^{+}16, HCL^{+}18, Hum17, KKMS14, OCBJ10, PCRD18, RCG^{+}18, SSA^{+}18, THHH15, THHH18, WVG^{+}13, Wri14]. rates [BSB^{+}13, BJR17, BTTV^{+}17, COR15, COS11, CB12, CSJ16a, DMBD10, FWS^{+}13, LCSI18, MONS14, MOG^{+}14a, MOG^{+}14b, MDV^{+}15, NHL^{+}19, PMH^{+}13, Pow14, RFT^{+}16, SDH^{+}18, THH18, Urb15, WGCL^{+}19, WKP^{+}17, WNW^{+}15]. ratio [ESR^{+}10, HOV^{+}15, HOS^{+}15, HHQ11, LB13, OKDJ18, SLH^{+}17]. rational [Hol14]. ratios [HOS^{+}15, KIA^{+}18, SPFM^{+}10, WMLJ17]. rattail [Lap11]. rays [BNB^{+}14, DOB^{+}17]. re [GRR^{+}12]. re-evaluation [GRR^{+}12]. react [PHO13]. reaction [TVO12]. reactions [REFJ12, Str10]. readability [FHH^{+}13a, FHH^{+}13b]. readers [WOW^{+}17]. reading [FFH^{+}11, FHH^{+}13a, FHH^{+}13b]. Real [Mac11, NC11, WM14, WESS18, JWS^{+}18]. Real-time [Mac11, NC11, WESS18, JWS^{+}18]. realistic [ABF^{+}16, EAOB15, FWP^{+}16b, LM10]. realities [KS14]. reality [PMOH13]. Realized [RDCPvH11, DDR10]. really [CSJ16a, STF^{+}17]. realm [MAGK18, RS10]. rear [HFSV^{+}15]. rear-view [HFSV^{+}15]. reared [OEK^{+}18]. Reassessment [WHL^{+}17, GMN^{+}17, GRR^{+}12]. rebuild [WPeEA16]. Rebuilding [HKKS10, Mur10, AJM19, AR10, HHJB14, Jan14, KST14, KZS10, MS10, OCR14, SPM^{+}18, SZM^{+}10, WPeEA16, ZHD^{+}14]. recall [SSP12]. recapture [BKTS11, SHK17, WKP^{+}17, WHNS14]. receding [WDOJ15]. reclamation [TNY^{+}13]. Recommendations [HR14, DCS^{+}11, LWN^{+}13]. recommended [ELBS17]. reconcile [KHC^{+}17]. Reconciling [SM15, URV^{+}11, WW16b]. Reconstructing [Eer12]. Reconstruction [MCOS17, KOK^{+}14]. Record [CN14, vHCP17]. recorded [DW11, GRNG^{+}10, vdKFS^{+}16]. recording [SHS^{+}18]. records [APK11, BLS^{+}18]. recover [Bow14, Hol10]. recoveries [JHB^{+}12, RHB^{+}12, RHB^{+}13]. recovering [PMBM18]. Recovery [AL10a, KPS^{+}16, KZS10, NKSE14, AFK15, BMVPN10, BNK10, BPF^{+}18a, BPF^{+}18b, BWP10, Coo19, CAB^{+}10, DCG10, DCNB^{+}10, EHHH14, FFRR12, GNFM11, HvDH^{+}10, HHJB14, JCA10, KDH^{+}10, KKH14, MV10, MKAR^{+}18, MR14b, OE16, PSM^{+}10, RvDW10, SBFC10, SWS13, TB16, Wri14, vGA18a]. recreational [BCK^{+}15, BCC^{+}19, CRAPMN12, CRAM^{+}14, CMH^{+}16, CBL19, DFB^{+}10, ESAV15, FWS^{+}13, FWH^{+}15, GPP17, JRG^{+}16, JKS^{+}18, PE18, SSP12, SSZH12, TBD^{+}18, WS13, ZRU^{+}15, vdHdGL16]. recreationally [SLS15]. recruit [HOS^{+}15, MR10, PR16, SCS^{+}11, THC16]. recruited [HdBR^{+}16]. recruiting [GMN^{+}17]. Recruitment [Cor13, LSB^{+}10, RKL^{+}10, AGMC14, ADB16, ALBR14, BHC^{+}15, BHJ14, BLB^{+}18, BHFH14, BMC^{+}13, BFS16, BBB^{+}19, Bru10, BMH^{+}16, Cad13, CCA^{+}18, DCHN^{+}15, ÉSMGB15, GCG^{+}10, GF11, HHT13, HMM12, Hüs11, HHE^{+}16, JKvdK^{+}15, JDFN12, JOF14, KS17, KHH^{+}17b, KKH14, LKR16, LF14, LB13, LMG^{+}14, LGR^{+}14, MA13, MFT16, MBIH11, MTO^{+}18, NKSE14, NNY17, OFYS^{+}11, OFVF14, OTM^{+}16, PNC15, PRBF18, PRO14, RB14, RCS^{+}17, RU12, RYS11, SGQR^{+}12, SHW^{+}15, SiTiM^{+}11, STR16, SFW^{+}12, SMR^{+}10, SDSN14, ZRN18, ZD14]. recruits [AV15]. recycling [EP14]. red [AS14, BTA^{+}18, BLM^{+}17, FPRA11, GMKS11, Hje14, KZS10, LPSF19, LEE17, LBTS^{+}19, MS15, PSGY^{+}12, SBG10, SQ14, SMK15, SWS13, SLF17, WHNS14, Win15, BdSP11]. Redefining [Rid18]. redfish [CC17, CBD^{+}10a, CMP^{+}11, MAP^{+}11, PJDN12, SHG^{+}18]. redox [Ham14]. reduce [BSP15, ETG^{+}15, GLP^{+}11, HCR13, HRB15a, HBC^{+}11, KFN15, OSL^{+}15, OCS14, OYI17, PDR19]. Reduced [HRB15a, CGK^{+}17, DW11, DH13, DWFD13, DLSJ^{+}19, HTRF17, RCG^{+}18, SPS15, TMR^{+}16, ZE17]. reduces [HBHR18, MOG^{+}14a, MOG^{+}14b, PRF^{+}17, SJ16, SLF16b]. Reducing [BMC^{+}18, GD15, MUEO17, RMKG11, RDKK15, SLV16, SIP18, NDF^{+}15, OBJA^{+}18, WESS18, YS15]. reduction [FDC13, HBHR18, MKHK18, WSGD^{+}17]. reductions [Hol10, MFT16]. Redundancy [GFR^{+}12, RDGP13]. Redux [Bro17a]. Reef [HBB^{+}13a, ASB^{+}19, BJR17, BTL17, BIKP^{+}17, BFB^{+}11, CGRM^{+}13, GPP17, GOH^{+}19, GFSC17, GS15, KNH11, MCO^{+}17, SPdJFMS17, SMS^{+}17, WLB11, Day18c, PEV^{+}16]. reefs [BdSP11, FTM^{+}17, GS14, GS15, JDH^{+}14, Jok16, KK18, LGR^{+}19, PS13, RBMAOCL18, SLT^{+}14, SLM^{+}11, SLT^{+}17]. reel [CMH^{+}16]. Referees [Ano10b, Ano11d, Ano11f, Ano12b, Ano12c, Ano12d, Ano10a, Ano10c, Ano10d, Ano11e, Ano11g, Ano11h, Ano12a]. Reference [BSB^{+}19, DGC12, MRD19, ABRL12, BPC10, Cad13, CAOG15, DPL16, DMT10b, GCJL16, GPS^{+}17, GFF^{+}19, Har13, HLL18, HBB^{+}13b, LB13, LM10, Mac12, MR10, MSR14, PH11, RCH^{+}15, RHSH16, TNS13, WÓG^{+}16, WMG^{+}17, ZCW11, dGGW^{+}11]. refine [CCC^{+}14]. refinement [DZ14, Fle15]. reflect [CFT^{+}19, PA18]. reflected [RPDF14]. reflecting [VGWJ11]. Reflections [HO14, Mac17a, KS14]. reflects [GKC^{+}15]. Reflex [MSM17, HCR13, TBG16, UTA^{+}16]. reflexes [JGCH18, KVFK19]. refuge [HHM^{+}16, MLMS15]. refugia [KK18, OFVF14]. regard [FDC^{+}16]. regeneration [GKR10]. regime [FBD^{+}19, KGN^{+}10, LSB^{+}10, Mos18, RBMAOCL18, SCM19, SP13, TNS13, vPBLR19]. regime-based [SP13]. regimes [CDAN^{+}14, HNAK12, SLTL15, SJ18]. region [BS13, BÁOMRV13, HHGtIWGoOH12, JVRT10, NdlPIR18, PGV^{+}17, SOGT^{+}19, VPSV^{+}10, VSP^{+}14, WMJ13]. Regional [DUM^{+}12, aFKA11, JBSD11, MTO^{+}18, SKH^{+}12, VCF^{+}12, WDE^{+}18, ASV^{+}19, EDBMB14, GPG14, HHMF19, KCS10, Llo17, MPJ12, PCB^{+}14, RTAT17, SHST18]. Regional-scale [JBSD11, MTO^{+}18, SKH^{+}12, WDE^{+}18]. regions [LGR^{+}19, OAT^{+}12, RST18, dCWMH13]. regression [SR17, WRDF10, dlPG15]. regression-based [SR17]. Regrowth [SMBH16]. regulate [SMR^{+}10]. regulated [HSD^{+}16]. Regulation [HSB18, LH19, PMOH13, vDTF^{+}14]. regulations [KJG^{+}15, MSAS14]. regulatory [FDC^{+}16]. Reinhardtius [DPOL13, DPRG^{+}18, JBE14]. related [AVGÓ12, CCC^{+}14, Dig19, HFH10, KUM^{+}12, MM12, MHLW12, MUW^{+}19, SPM^{+}18, VAP12, WMJ13]. Relating [LYS^{+}10, SBCF16, MS15]. relation [BS13, BÖBL12, BT10, CKD^{+}17, CDAN^{+}14, EKR^{+}19, EPVC14, FMM^{+}12, GPG14, HSK^{+}14, KKLM13, KGW^{+}12, LK15, MKAR^{+}18, MRMK11, NLB^{+}12, PBJ^{+}14, QCA^{+}18a, QCA^{+}18b, ROBC11, SMR12, SOB^{+}11, VHH10]. relations [BI13]. relationship [BHDB12, CR19, KN17, LTT^{+}18, MB14, Mos18, PRB^{+}15, PG10, PSF12, SGQR^{+}12, SRKV15, TKA^{+}16, WPH^{+}15, WPH^{+}17, YLLG18]. Relationships [BR12, MKC^{+}11, dJBLH15, BNP19, CBBM14, DHCE13, MMC19, dSMGKP12, PRBF18, ZRN18]. Relative [MBS^{+}15, BNP19, BLB^{+}12, CMG^{+}17, CSY^{+}10, DEH^{+}19, GS14, HTKB19, IOK17, NC12, NDP^{+}16, PSS^{+}14, RDS12, TSP12, VBA10, VE13]. relatives [Elv15, ØHS10]. release [BBHC11, CJFS16, CPT^{+}10, FWS^{+}13, FWH^{+}15, HWK^{+}15, HLA^{+}11, JTM^{+}16, KVFK19, LSF^{+}18, ODSC10, PVCB17, SLV16, TBG16, WS13]. released [BBH^{+}10, BBOE10, CHD^{+}18, PST14, RBBC11, TPHdJ19]. releases [BCK^{+}15]. relevance [DGPC^{+}15, GGIC^{+}15, KYK^{+}13, MBQ10, RBB^{+}15]. relevant [BP13, CAOG15, FCG^{+}16, GKJ^{+}15]. Reliability [Mil10]. reliable [LSJ10, SBS^{+}10]. relicta [SR16]. relies [NJFH19]. remain [MP15]. Remembrance [Ban12]. Remote [CSD16, NDB^{+}15, QFM^{+}10, SPSP11, APOG11, CBR^{+}11, CCSG18, DMT^{+}10a, DMS16, HSFiS12, JBSD11, RSY11, ROBC11, RPS11, SMR^{+}11, SOB^{+}11, SPS11, WFS^{+}11]. remote-sensing [ROBC11, SPS11]. removal [Alf10, FBM^{+}18]. Renaissance [Hut14]. renewal [EMP11]. rent [ETG^{+}15]. repair [CPLH16]. repeatable [LSJ10]. Reply [DK19, FWP^{+}16b, AAP14a, RPM14b, STF^{+}17]. Report [HHGtIWGoOH12]. reporting [ACT^{+}18]. reports [DGPR11, GUN^{+}19]. represent [SBS^{+}10]. representations [HFH10]. representative [VQB^{+}11]. representativeness [EPH15]. reproducing [OBJA^{+}18]. Reproduction [MARD13, CK18, FMLM^{+}14, GS14, LSF16, RST18, SBKA16, SLF16b, WNM10]. Reproductive [DPOL13, HBB^{+}16, NNY16, BHDB12, BBM15, Hje14, HLD10, KUM^{+}12, LPH^{+}15, LFMJ19a, LFMJ19b, MFP^{+}19, MKC^{+}11, NUO^{+}10, OTAK15, OG16, QZH^{+}17, RvDW10, ROK^{+}18, Sas19, SS10b, THSM^{+}18, VKT^{+}12, WPH^{+}15, XS12, Yar10]. reptile [WSGD^{+}17]. require [Ber18b]. requirements [GGIC^{+}15, RSGM17]. requires [DC14, EHD^{+}15, HSGL18]. research [AB16, BKVT16, BTTV^{+}17, Bro16, Bro17b, BCC^{+}19, Cam18, CTH^{+}19a, CTH^{+}19b, CH16, DW11, DH13, DWFD13, DLSJ^{+}19, DCS^{+}11, DKC^{+}17, GD15, Hin15, Mac17a, MHH^{+}18, NDB^{+}16, Pin17, RB14, STF^{+}17, SPP^{+}16, VP19, Wil11]. researchers [DKC^{+}17]. Reserve [JCSC10, DHS19, HDG^{+}17, PRF^{+}17]. reserves [CGRM^{+}13, GKR14, MSK^{+}11, MDDP10, SG18b]. residence [URE^{+}18]. residency [FL14]. resident [CCSG18, OS14]. residual [SGQR^{+}12]. resilience [Cri12, GQX^{+}19, KCMS19, KKH14, LPH^{+}15, RDB^{+}18b, SSM^{+}16, WZZ^{+}16]. resiliency [HHH^{+}19]. resilient [MMH17]. resistance [dlPG15]. Resolution [MST^{+}18, APOG11, GL11, KRD^{+}18, LSJ10, MSR^{+}19, ROBC11, RCH^{+}19, SCJI10, TRSM15, WHL^{+}17]. resolve [NCH17, RHHH^{+}12]. resolved [GBJ^{+}15, NRTK^{+}19, Ohm19]. Reson [PPM13]. resonance [BHFH14, FOJ13, SCJI10]. resource [CHB^{+}14, CK18, LLPV^{+}12, Mer18, Tow14]. Resources [MB14, CFA^{+}16, CGRM^{+}13, GHF^{+}11, GDHFS^{+}18, HDF^{+}19, HB19b, HHM^{+}19, HHT13, MG15, OND^{+}18, PNE^{+}14, RMKG11]. respiration [CCE17, LK17, LBJ15, LPSF19]. respond [GOH^{+}19]. responds [PSH13]. Response [ABJ^{+}19, CSJ16b, HKB^{+}18, JPPY11, Man17a, WW16a, BDW18, BM12, BHS^{+}11b, CAC^{+}18, DL11, FWC^{+}19b, GDD^{+}18, HSB18, HWK^{+}15, HBF14, JGCH18, Kim12, KDF^{+}19b, LGvPH15, PvOP10, PMM^{+}14, PTvOR13, RCG^{+}18, dÁMNGS^{+}11, SDT15, TVO12, TNS13, TFM^{+}12, VGB^{+}17]. Responses [BMM^{+}19, JMF12, AFQ^{+}11, Bak14, CFA^{+}16, CCSG18, CGM^{+}19, DK18, DKPM16, EEP^{+}11, GM17, HBiI^{+}11, iIOKW13, KPD^{+}16, LFFL13, LFH^{+}19, LYS^{+}10, MKHK18, NMF^{+}16, RFN^{+}19, RH17, RPB16, SBNL12, SFD^{+}16, TMP^{+}17, ZSC16, ZHL^{+}19]. responsible [KRG^{+}12]. Responsive [KCMS19]. restoration [KP13]. restored [GOH^{+}19]. restoring [dCVB14]. restrictions [BNE^{+}15]. result [MUEO17, PCFR13]. result-based [MUEO17]. resulting [ZCR15]. results [DBM^{+}15, FGRR11, FFH^{+}11, FHH^{+}13a, FHH^{+}13b, GRS^{+}14, HGS11, LPK12]. results-based [FGRR11]. resurrected [Abl16]. retained [ANU11, MPR17]. retention [BDHC19, BHM^{+}18, MSAS14, ØU13, SHK17, WPWH11]. reticulatus [BBL^{+}15]. Retraction [Ano10e]. retrospective [AH15, HFSV^{+}15, LHK18, SKKM15, SIP18, TVJ^{+}14]. retroversa [BLMW17]. Return [BCS^{+}17, UFJ^{+}18, CB12, GIR^{+}19]. returning [WÓG^{+}16, WMG^{+}17]. reveal [GKJ^{+}15, HGH^{+}16, MTP^{+}12, SHW^{+}18, TBD^{+}18, TMP12, WKK^{+}15, WRMJ13, vdSSM^{+}18]. revealed [ABJ^{+}19, ALBR14, CMH19, GHF^{+}11, GRNG^{+}10, JBSD11, KBB13, LFM^{+}18, SHH^{+}17, SBS^{+}19c, VPSV^{+}10, YDT14]. revealing [YCZY19]. reveals [DVH^{+}18, EBSW15, GSL^{+}11, HGF^{+}18, JQD^{+}17, JWS^{+}18, KOK^{+}14, LÁG^{+}19, MAS19, MMSS18, PSH^{+}11, PGS^{+}19a, PGS^{+}19b, WNM^{+}13]. revenue [EHG^{+}14]. Review [DSL14, Dig19, Fle15, Hüs11, PHM13, AL10a, CKB^{+}16, CAS^{+}16, CD17, DHH^{+}14, DH13, DL12, DCNB^{+}10, ECW^{+}17, FPRA11, Hin15, KM16, Kim12, MV10, MRP^{+}16, NDB^{+}16, OAM^{+}10, OBY^{+}14, PCdL15, PGN^{+}11, RCC^{+}14b, SBFC10, SLG^{+}15, SMMK11, SCL^{+}19, TTG^{+}15, TPV14, TBHK17].

Reviewers [Ano18o]. reviews [HHMF19]. revised [ABRL12, PGOM11]. revision [BSW^{+}11a, BSW^{+}11b]. revisited [DW11, KHH^{+}17b, MR14b]. Revisiting [CSB18, MSR^{+}12, PHV^{+}15, SJB15, PRB^{+}15]. reynaudii [DRB10, MRC^{+}10, RDS12, SHL10]. Rhizostoma [EHT17]. rhythm [GDS15, GKR10, HGH^{+}16]. Ria [LTQ14]. ribbonfish [TKA^{+}16]. Rican [RBMAOCL18]. rich [DFG^{+}17]. Richness [CBBM14, CNBK11, JVRT10]. rides [RLW^{+}15]. Ridge [JV12a, OYI17, Lap11]. ridge-scaled [Lap11]. riding [DCER^{+}14]. Riemann [BF17, Man17a, Man16]. rig [BMC^{+}17]. Riga [MO18, RKL^{+}10]. right [BT10]. ring [BNP19]. ringens [CAC^{+}18]. ripening [Yar10]. Rising [KS17, MM17, PTvOR13, WWMF17]. Risk [CBS15, GB15c, SDW19, TKB^{+}15, AHK^{+}18, Ast15b, ASB^{+}19, BV19b, BV19a, CPT^{+}10, CLdR^{+}15, Fle15, GLH14, GCJL16, PGS^{+}19a, PGS^{+}19b, SDH12, SFG^{+}15, Van11, ZMF12, ZDF^{+}19]. risk-based [GLH14]. Risks [TJD17, APHC15, KPJ^{+}15, LTGP18, WHP16]. Risso [CJP^{+}14]. River [CB12, FBM^{+}18, LEB^{+}14, ME11, NSS15, RC12, STC^{+}17, Bow14, BFS16, LKS^{+}14b, LNH^{+}15, MWS^{+}17, MSAS14, PDT^{+}18, RDB^{+}18a, WABZ16]. rivers [DLL12, DFS^{+}19, GUN^{+}19, SRKV15]. road [GNFM11, Wil11]. roadmap [CGGH^{+}17]. roam [MVB^{+}18]. Robben [RBP15]. Roberts [HBF14]. Robin [PSS11]. robustly [PSH13]. Robustness [SHS11, AH15, DBM^{+}15, MRSG^{+}17]. rock [CDAN^{+}14, CN16, HGGJ15, KFBG15, LGvPH15, LMG^{+}14, PGG^{+}17, SQ14, dL14, dLCF^{+}15]. Rockall [SRHJ12]. Rockfish [LTGP18, Fau11, FDC13, KKMS14, OE16, TSP12]. rocky [CGRM^{+}13, Wil11]. rocky-reef [CGRM^{+}13]. rod [CMH^{+}16]. rod-and-reel [CMH^{+}16]. rodgersii [DDF^{+}12]. roei [HSB18]. role [Aks15, BCC^{+}14, BIKP^{+}17, BPST^{+}16, BPF^{+}19a, BPF^{+}19b, FLCQ19, FP12, GRP10, MA13, Mat11, MBLS15, MLMS15, MDPH12, MST^{+}18, PHGK^{+}16, RKW^{+}17, RLP^{+}13, SDDA11, VDK^{+}18, vPGFT13]. roles [DSN16]. rollercoaster [DCER^{+}14]. Rosa [KHMS19]. Roskilde [PR16]. Ross [PBG14, SJ15]. rostrata [BM16, HdBR^{+}16]. roughness [RT19]. roughy [MKR13, RK16]. round [SG16a]. routes [LSWD12]. routine [dGGW^{+}11]. Ruditapes [BPBD15]. rufescens [BLM^{+}17]. rule [AGH^{+}11, LWT^{+}11, MHM^{+}11]. Rules [Eid18, BF17, BPvHR10, CCS10, KCMS19, LLL^{+}13, Man16, Man17a, MSA19, MLMJ17, PAB^{+}14, ZCW11]. run [GUN^{+}19, MBT^{+}19]. rupestris [OHLK19a, OHLK19b, SJ16]. rural [Bas18c]. Russell [MLC16].

S [HLL18, LHV^{+}16, ACA19, PG10]. S. [SCD^{+}15]. Sablefish [PH17, HABV19]. sac [HPN^{+}17, LCSI18]. Saccharina [LSR^{+}13, NMF^{+}16]. Sackville [FMM^{+}12]. safety [LØM^{+}18, SCD^{+}15]. sagax [DZC^{+}13, DZ14, ZED11, ZD13, ZD14]. sagittal [KSJ14, ZGT13]. Sagmariasus [GFMO11]. saida [KLCC18, KCL18, LCSI17, LCSI18, WSW^{+}19]. Saildrone [DLSJ^{+}19]. sailfish [PE18]. Saint [GRG^{+}10]. saira [iIOKW13, TSY^{+}11, TSS^{+}13]. saithe [AJ12, CLV^{+}14, CKV^{+}16, GHV^{+}18, HSNO12, íHHJ^{+}13, OAM^{+}10, OS14, SHK^{+}15, SBH^{+}14, SDT15, SO16]. Sala [Hil18b]. salar [BDHC19, CB12, Cha12, CCD^{+}19, DPD^{+}12, FBM^{+}18, GCW^{+}18, GSMA15, JHB^{+}12, KUM^{+}12, MKC^{+}11, MMDSP15, MGH^{+}12, OEG^{+}16, OVW^{+}17, RHB^{+}12, RHB^{+}13, RC12, SHAM10, TMP12, UFJ^{+}18, VCF^{+}12, WÓG^{+}16, WMG^{+}17]. sale [LGvPH15]. salinity [GQX^{+}19, PMM^{+}14, RHV14, RCG^{+}18, ZSC16]. Salish [IEL^{+}15]. Salmo [BDHC19, CB12, Cha12, CCD^{+}19, DLL12, DPD^{+}12, FBM^{+}18, GKC^{+}16, GCW^{+}18, GSMA15, JHB^{+}12, KUM^{+}12, MKC^{+}11, MMDSP15, MGH^{+}12, OEG^{+}16, OVW^{+}17, RHB^{+}12, RHB^{+}13, RC12, SHAM10, TMP12, UFJ^{+}18, UPF^{+}10, VCF^{+}12, WÓG^{+}16, WMG^{+}17, WKP^{+}17]. Salmon [BHS^{+}16, BHM^{+}18, HHRW12, VHF^{+}17, BGM^{+}11, BJH^{+}16, BR12, Bow14, BDHC19, CB12, Cha12, CCD^{+}19, CHD^{+}18, DBDP10, DFS^{+}19, DPD^{+}12, FSN^{+}11, FBM^{+}18, FBF^{+}17, FDR^{+}17, aFKA11, GKC^{+}16, GCS^{+}18, GCW^{+}18, GUN^{+}19, GSMA15, GIR^{+}19, HHDB14, HUT^{+}11, HTT^{+}17, Hol10, HFH10, IS15, IaF11, JHB^{+}12, JÓT^{+}12, KDF^{+}16, KUM^{+}12, KBT14, LSWD12, LKHK11, MTP^{+}12, MWS^{+}17, MVB^{+}18, MRD^{+}12, MGPC^{+}14, MKC^{+}11, MSR^{+}12, MMDSP15, MGH^{+}12, MRMK11, ME11, NSS15, OEG^{+}16, OEK^{+}18, OS14, OSSL16, OVW^{+}17, Qui18, RHB^{+}12, RHB^{+}13, RC12, RH17, RAB^{+}12, SLKS10, SRCR12, SAS^{+}18, SHAM10, SHS15, SI15, STC^{+}17, SFN^{+}19, TKB^{+}15, TFM^{+}12, TVJ^{+}14, TMP12, UFJ^{+}18, VCF^{+}12, VOS10, VDK^{+}18, WÓG^{+}16, WMG^{+}17]. salmonid [BHM^{+}18, HSV^{+}17, VP19]. salmonis [BSMB18, GKC^{+}16]. Salpa [WCK^{+}10]. SALSEA [SRCR12]. salt [BDHC19]. saltatrix [BLS^{+}18, SMR12]. Same [ZHV^{+}16]. sample [HHQ11, HHQ12, LHV^{+}16]. sampled [GSL^{+}11]. samples [UBvH11]. Sampling [DZC^{+}13, CSFS10, GHG^{+}10, HPS^{+}11, HGS11, MBP19, POIM12, Par13, PH11, TID^{+}10, UBvH11, VKN^{+}11, ZED11]. sampling-effort [CSFS10]. Samsonfish [PPM13]. Sanctuary [Lev11]. sand [DTL^{+}11, FRF14]. sandbar [HBHR18]. sandeel [FHOK14, GHG^{+}10, JRWM11, KO12, WCR^{+}19]. sandeels [JH13, LHH^{+}13]. sands [DPL10]. Santander [WRC^{+}12]. São [PG10]. sapidus [GM17]. SAR [KGP^{+}15]. Sardina [CMU^{+}17, MMA^{+}10, MOVU10, SGQR^{+}12, SSB^{+}10b, TID^{+}10]. Sardine [ZOQS10, BSW^{+}11a, BSW^{+}11b, CMU^{+}17, DZC^{+}13, DZ14, HFHS10, MMA^{+}10, MKF^{+}15, dSMGKP12, MOVU10, NNY16, RBP15, SGQR^{+}12, SWRC^{+}18, TID^{+}10, ZED11, ZD13, ZD14, dMBD11]. Sardinella [dSMGKP12]. Sardinops [DZC^{+}13, DZ14, ZED11, ZD13, ZD14]. Sargasso [ISL19]. Satellite [BFH^{+}18, CBR^{+}11, HSFiS12, AGG^{+}18, CAM^{+}15, HLA^{+}11, LFM^{+}18, LDT^{+}11, PE18, RSY11, RPS11, SMR^{+}11, SOB^{+}11, STC^{+}17, STB^{+}11, WFS^{+}11, Wil11]. Satellites [Koe11]. Satisficing [MS10]. saturation [BSB^{+}13, CSJ16b, RRM^{+}12, WHH16]. Saudi [RLQ^{+}15]. saury [iIOKW13, TSY^{+}11, TSS^{+}13]. saxatilis [VWMS19]. scabbardfish [MMM11]. Scale [BBH^{+}10, AIA13, APOG11, ACS15, BNP19, CAL14, dMADLP^{+}16, DHZA14, DCVC15, FMML19, GNB^{+}16, GW16, GKR10, GF11, HHR17, HAR19, HOV^{+}15, JBSD11, JRG^{+}16, JRWM11, JÓT^{+}12, JH13, KRO^{+}16, LGRC14, LMG^{+}14, MVK18, MPR17, MHM^{+}11, MPJ12, MSR^{+}19, MPBH11, dSMGKP12, MTO^{+}18, MJS^{+}15b, NdlPIR18, OJK^{+}19, PBJ^{+}14, PVQS12, PDD^{+}11, PLCC18, RR11, RFHG19, SBNL12, SMH^{+}19, SKH^{+}12, UBV^{+}11, VGWJ11, WDE^{+}18, WSG^{+}17, WCR^{+}19, dGGW^{+}11]. scale-circuli [UBV^{+}11]. scale-less [PDD^{+}11]. scaled [Lap11]. scalefish [LWT^{+}11, MWJ11]. scales [GHB^{+}14, LTQ14, STR16, SSF18b, TVJ^{+}14, dPLF^{+}19]. Scaling [RT19]. scallop [BMF14, KHH^{+}17a, LSI^{+}15, MLMS15, MHHK13, OTM^{+}16, SPS15, SG16b, SMB^{+}16b, THC16, THR16]. scalloped [FFN^{+}13, FFN^{+}16]. Scallops [SPS15, RFN^{+}19, SH14a]. Scampi [TPHC15]. Scandinavian [TRH10]. scanning [VPH18]. scans [FOJ13]. scarce [ØS15]. scarer [HHDB14]. scats [WRMJ13]. Scattering [FHK14, GCBI17, KRKG16, LCM10, MMG^{+}17, PHK^{+}19]. Scavenging [SMZ^{+}14, ZSC16]. scenarios [BJR17, FKH^{+}16, HGC^{+}19a, HGC^{+}19b, May14, MDD^{+}14, SPdJFMS17, SHZ^{+}16]. scene [SSB^{+}10a]. scepticism [Bro16]. schedules [WMG11]. scheduling [HP11]. scheme [RML^{+}16]. schemes [SCL^{+}19]. school [TK18]. schooling [OMTY14, SS17]. schools [HHR17, PHO13, VPO17]. sciaenids [MPBH11]. Sciaenops [AS14, LEE17]. Science [BF17, CMP^{+}11, HG10, MAP^{+}11, Man17a, AB14, AB16, And15, BLD17, BLS^{+}18, CHB^{+}14, FWS^{+}13, FO13, HKKS10, HD18, Hin15, KH19, Mac17a, MSC14, Mer18, MTLG14, Mur11, NDB^{+}15, OdS19, Ove19, Pau16, Pin17, Ric11, Rot15, SPP^{+}16, SPS11, TGH^{+}12, Tur19a, Tur19b]. science-based [MTLG14]. Science-to-management [HG10]. sciences [MHH^{+}18]. Scientific [HHRW12, Hub14, BHB^{+}19, Cam18, Dig19, JdFBB^{+}10, KPL^{+}19, KCMS19, Paf18, PSGY^{+}12, WM14]. scientists [DSN16, GB15c, Pet19, SOL^{+}15]. scleractinian [CCE17]. Scomber [AVGÓ12, Jan14, JKvdK^{+}15, NSO^{+}19, NUÓ^{+}16, OSJ^{+}16, PDZ^{+}13, RSBC13, SGM^{+}17, SCS^{+}11]. scombrus [AVGÓ12, Jan14, JKvdK^{+}15, NSO^{+}19, NUÓ^{+}16, OSJ^{+}16, PDZ^{+}13, RSBC13, SGM^{+}17, SCS^{+}11]. scope [BBD^{+}13, DI11, ZSC16]. scores [KVFK19]. Scotia [EAB^{+}17, STF^{+}16]. Scotian [FFK^{+}11, SSTB15, ZPT^{+}11]. Scotland [AHM^{+}15, DMF13, HHA^{+}11, JMF12, JCS16, TDHC18, WKW^{+}10]. Scottish [FCD^{+}11, GGTW17, GCS^{+}18, GSMA15, LCTB^{+}15, NDB^{+}15, UPF^{+}10]. screening [CLdR^{+}15]. scriba [APB^{+}10]. scuba [DFB^{+}10]. Scylla [BLB^{+}12, MR14a]. Scyllarides [GDS15]. scyphomedusae [GCL^{+}18]. scyphozoan [Eri16, GHB^{+}10]. SDWBA [MDS13]. Sea [BPF^{+}18a, BPF^{+}18b, BNB19, GTM12, GKC^{+}16, HHRW12, HKB^{+}18, HLL18, JIB^{+}14, KYK^{+}12, MRMK11, PBG14, SFN^{+}19, WSW^{+}19, ANU11, BSB^{+}13, BGM^{+}11, BKMdMS13, BvdMA^{+}18, Ber18b, BGM^{+}19, BSC14, BCL^{+}11, BdSP11, BHCSP^{+}11, BSMB18, BA12, BBH^{+}10, COM11, CEP^{+}18, CAS^{+}16, CRC^{+}16, DHS19, DLL12, DDF^{+}12, FBFP^{+}13, FDJ18, FDJ19, FSC^{+}12, GHB^{+}14, GvRKB19, GHB^{+}10, GGIC^{+}15, HB19a, HR14, HHA^{+}11, HLD10, HBF14, JHB^{+}12, JBE14, Kam14, KHH^{+}17a, KLC13, KTLB16, KS18b, LHJ15, LCMR^{+}18, LHV^{+}16, LSF^{+}18, Llo17, Lor11, MIJ^{+}17, MM17, MNWB18, NST^{+}14, Ohm19, OEG^{+}16, OEK^{+}18, OFVF14, PH12a, PDCP14, PPS17, PGN^{+}11, PLCC18, RWG^{+}18, SPRL18, Sch14, ŠBGT^{+}17, SBCF16, SPS15, SH14a, SCC^{+}13, SEDO19, TBG16, TBC^{+}14, THC16, TSY^{+}11, UPF^{+}10, VOS10, VHF^{+}17, VDK^{+}18, WKP^{+}17, WDG^{+}17, ZRU^{+}15, dQCD^{+}10, dlTQMUE10]. Sea [ASDSM18, AEC11, AMS^{+}12, AFLvDH15, AL10a, BJH^{+}14, BCC^{+}14, BNE^{+}15, BPF^{+}18a, BPF^{+}18b, BGM^{+}14, BJMG19, BDCvD^{+}10, BPF^{+}19a, BPF^{+}19b, BdSP11, BKT12, BA12, BBD^{+}13, BMEBM^{+}16, CMU^{+}17, CTH^{+}19a, CTH^{+}19b, CP15, CFT^{+}19, CCG14, CDAN^{+}14, CN16, CLV^{+}14, CKV^{+}16, Cor13, DIS^{+}12, DW11, DLL12, DUM^{+}12, DMO10, DCNB^{+}10, DCER^{+}14, DSA13, DCMC19, DJHO10, DD10, DD13, Eer12, EKR^{+}19, EPK^{+}18, EPKR11, EISJ12, EDP14, ESD^{+}16, Eri16, FRF14, FPRA11, FSN^{+}11, FMLM^{+}14, FTM^{+}17, FOC^{+}18, FL14, FFRR12, GÖN^{+}12, GZS^{+}19, GKC^{+}16, GCG^{+}10, GRKF16, GS12, GHM^{+}16, GDBM^{+}17, GNB^{+}16, GIW16, GMKS11, GRR^{+}11, GFR^{+}12, GRR^{+}12, GKR10, HZZ^{+}15, HBB^{+}18, HCC^{+}14, HSHÖ14, HGC^{+}19a, HGC^{+}19b, HFSB^{+}14, HMS^{+}13, HHM^{+}11, HGS11, HNAK12, HBD^{+}19a, HBD^{+}19b, HBG^{+}16, HB10, HF14, HHBSM16, HCE^{+}11, HMM12]. Sea [HCDAF15, HHHE16, IHH^{+}11, IKF17, ISL19, ITT19, IEL^{+}15, Jan14, JIB^{+}12, JMM19, JLS^{+}15, JPTC16, KUM^{+}12, KDH^{+}10, KJW^{+}18, KGN^{+}10, Kim12, KOT^{+}17, KJG^{+}15, KCP^{+}11, KGH^{+}16, KRD^{+}18, LGBA11, LHK18, LSB^{+}10, LLSJ18, LVPK10, LHLK10, LGR^{+}14, LHH^{+}13, MO18, MCAM14, MAA^{+}19a, MAA^{+}19b, MKD^{+}18, MHHT10, MVK18, MALM17, MDP^{+}15, MPR17, MYI12, MGL^{+}15, MS14, MHK19, MUW^{+}19, MSL^{+}11, MGBvD14, MKC^{+}11, MLB^{+}14, MBIH11, MJS^{+}15b, NWM^{+}12, NHL^{+}19, NLB^{+}12, NDP^{+}16, NRTK^{+}19, ØHN14, OTAK15, OLR^{+}18, OND^{+}18, OAM^{+}10, OBG12, OTO^{+}15, OJK^{+}19, OEK^{+}18, OFB^{+}17, Ott10, OBY^{+}14, PGG^{+}12, PHM^{+}14, Pay10, PNC15, PFPG12, PH12a, PCFR13, PJDN12, PKHG14, PSF12, PKK13, PT19, PHA^{+}10, RHV14, RDD^{+}10, RvDW10, RDCPvH11, RTJS14, RKCP18, RLP^{+}13, RCC^{+}14b, RCH^{+}19, SG17, STKT16, Sas19, SOGT^{+}19, SGK^{+}11, SRG11]. Sea [SFH^{+}12, SDAM^{+}18, STF^{+}16, SBH^{+}14, SDT15, SRRZ16, SJUE11, SM12, SBLS15, SRW13, SJ15, SR17, SSF18b, SHZ^{+}16, SSGH18, STSP17, SEDO19, Str10, SSZH12, SWS13, SPV^{+}16, SP13, TH15, TD19, TNS13, THKP11, TVJ^{+}14, TOA^{+}16, URV^{+}11, UVD^{+}17, Urb15, VVP^{+}14, VGWJ11, VSP^{+}14, VG15, VBO^{+}15, WKW^{+}10, WKP^{+}17, WKL13, WWI^{+}16, YLLG18, ZGT13, vdRMC^{+}17]. sea-ice [PH12a]. Sea-level [GTM12]. sea-spawning [LHV^{+}16]. seabass [CSTJ19, ŠBGT^{+}17, dPLF^{+}19]. Seabat [PPM13]. Seabed [CD14, CSD16, DSA13, DMS16, EBB^{+}16a, EBB^{+}16b, EBH^{+}17, GML13, GHG^{+}10, RD16, RBB^{+}16, SKM18, SMD^{+}16, TED16, vdKKS11]. seabeds [SMZ^{+}14]. Seabight [PGN^{+}11]. seabird [ACA19, BG18, CST^{+}19, DPBM12, FWRM12, LLG18, LJB16, TPS^{+}11]. Seabirds [SSF^{+}12, CFT^{+}19, DMJ^{+}14, GLP^{+}11, KGW^{+}12, MB14, SRS11, WCT^{+}16, XCRP13]. seaboard [Bow14]. seabob [BMSC16]. seafloor [PH12b, Van11]. seafood [HHZ^{+}18, KK19]. seagrass [Jon14, TNFC16]. seahorses [QZH^{+}17]. seal [HS10, HHDB14, LLCJ10, LDT^{+}11, ØS15, SBKA16, TDHC18, WRMJ13]. seals [AWS^{+}13, BMR^{+}11, DTA^{+}13, FFH^{+}11, FHH^{+}13a, FHH^{+}13b, GÖN^{+}12, GRKF16, HSSB14, HS17, LHLK10, ØFNH12, ØHN14, SOGT^{+}19, SS10b, SRH^{+}14, SH14b, TRH10, VRF^{+}16]. seamount [CCSG18, OdJN^{+}12, PGV^{+}17]. Searching [BCOS11, LLL^{+}13]. Seas [EPH15, AGH^{+}11, DSL14, NUÓ^{+}16, RRCT^{+}17, SJ18, HSFiS12, NSO^{+}19, Rey12, RKB^{+}12]. Seascape [DVH^{+}18, GOH^{+}19, SHK^{+}15, ABHT^{+}16, GFSC17, HSGL18, RIBB^{+}19, SG16b]. Seascapes [BF17, KOC^{+}16, Man16, Man17a, HGS^{+}19, HSB16]. season [BHC^{+}15, GMN^{+}17, LEB^{+}14, MWS^{+}17, RU12, SiTiM^{+}11]. Seasonal [BJR17, BHDB12, EPK^{+}18, HSNO12, JCLO17, MJS^{+}15b, MJS^{+}16, QZH^{+}17, SDAM^{+}18, ZLC^{+}17, Ada17, ALd12, BTTV^{+}17, CTC14, CBW15, CSF^{+}17, HBD^{+}19a, HBD^{+}19b, HHT13, HGH^{+}16, ISL19, LLM12, MDPH12, ODL16, PH12a, RHHH^{+}12, RGFT14, RTJS14, SLS15, TCQ^{+}15, VHGTFR10]. Seasonality [VG15]. seasonally [BdSP11]. SEASWAP [SOL^{+}15]. seawater [BTB^{+}17, BWH16, LK17, LBG18, LFH^{+}19, SSM^{+}16, SMB^{+}16a]. seaweed [COL^{+}13]. seaweeds [BMM^{+}19]. Sebastes [CC17, CBD^{+}10a, CMP^{+}11, KKMS14, LTGP18, MAP^{+}11, PJDN12, SHH^{+}17, SHG^{+}18, TSP12]. Sebastião [PG10]. second [CB12, GUN^{+}19]. second-generation [GUN^{+}19]. secondary [AHC15, GMM^{+}12]. Section [PT19]. sector [PVQS12, PEN^{+}19b, RG11]. sectoral [MLS^{+}17]. secure [GAB^{+}14]. security [RG11]. sedentary [GDHFS^{+}18, PG16]. Sediment [COS11, Ham14, Alf10, BMEBM^{+}16, EDBMB14, OSP^{+}13]. sedimentation [DFB^{+}10]. sediments [DDE^{+}19, JW18, LNWS18, OI16]. seed [PCBO^{+}18]. Seeing [Bra18]. seek [LAD^{+}19]. seeps [HTRF17, SFD^{+}16]. segregation [HFNH13, MTP^{+}12, PQR10]. seine [ACC^{+}12, CHD^{+}18, CLMP16, IHL19, LCMR^{+}18, MMA^{+}10, MKF^{+}15, MKB^{+}17, MNGA^{+}17, NMM^{+}17, PH11, RBC^{+}15, TMR^{+}17, TPHdJ19, TRPH16]. seines [EBB^{+}16a, EBB^{+}16b, TVO12]. seismic [FHD^{+}19, PHO13]. SELECT [TJR^{+}10]. Selecting [OLR^{+}18, PSS^{+}14]. selection [AUEC12, AWS^{+}13, CIJ18, CMD17, DVH^{+}18, LSR^{+}13, LVPK10, MVH^{+}11, MHMP^{+}18, RSY11, SHG^{+}18, TCBD10]. Selective [IHL19, VVP^{+}14, HSV^{+}17, MV10, MSA19, NDF^{+}15, OE16, PvdKE^{+}12, SFH^{+}12, SWA^{+}16, WPWH11]. Selectivity [FMK10, CAM^{+}15, DZC^{+}13, GPP17, HLCC16, KHMS19, LHS^{+}18, LB11, Mil10, ODL16, SWvSR11, STD11, TJR^{+}10]. self [HGS11, PH11, STSP17, UBvH11, VKN^{+}11]. self-governance [STSP17]. self-sampling [HGS11, PH11, UBvH11, VKN^{+}11]. Sélune [FBM^{+}18]. semelparous [PFH16]. semi [MDPH12, SHS^{+}17]. semi-automatic [SHS^{+}17]. semi-diurnal [MDPH12]. Sendai [Mur11]. seniors [DKC^{+}17]. Senja [BV19b, BV19a]. sensing [APOG11, CBR^{+}11, CSD16, DMS16, HSFiS12, JBSD11, QFM^{+}10, RSY11, ROBC11, RPS11, SMR^{+}11, SOB^{+}11, SPS11, SPSP11, WFS^{+}11]. sensitivities [LBM19]. Sensitivity [KRD^{+}18, MRSG^{+}17, DRVV^{+}17, EBSW15, FTM^{+}17, FWRM12, GIDP^{+}18, HCC^{+}14, Hum17, JLH12, LMPP10, McE17, SMB^{+}16a, TBC^{+}14]. sentiments [Str10]. separable [DDR10]. separation [MAA^{+}19a, MAA^{+}19b, MFS^{+}14]. Sepia [GRC^{+}14, LLLB^{+}10, OBJA^{+}18, SLF^{+}16a]. sequence [DMF13, VSP^{+}14]. sequences [NNY16]. sequential [MSA19, PSB16]. series [BYQ^{+}19, CLMP16, CCA^{+}18, DCHN^{+}15, HLL^{+}12, KOK^{+}14, MG12, NBS^{+}10, SGQR^{+}12, SPE14, SCM19, THH18, WHL^{+}17]. Seriola [MWP^{+}16, PPM13, RBBC11]. Serranus [APB^{+}10]. serrata [BLB^{+}12, MR14a]. services [DPE^{+}17, SSR^{+}16]. sessile [BWM^{+}16, HJT12, PEV^{+}16]. set [AFH^{+}17, GLP^{+}11, KMK10, SKM18, Sei14, SSF^{+}12]. set-nets [SSF^{+}12]. Setting [SSB^{+}10a, MLMJ17, OCBJ10, PSGY^{+}12, RDT^{+}17]. settled [ALMFFBP19]. Settlement [BGM^{+}14, HGGJ15, BvdMA^{+}18, EF15, HTRF17, IKF17, KKMS14, LMG^{+}14, MA13, MLMS15, RHHH^{+}12, dLCF^{+}15]. settling [LKR16]. seven [dLCF^{+}15]. severity [AS14, NF16]. sewage [DAB^{+}12]. Sex [HSV^{+}17, BMR^{+}11, BFDGLÁ15, ESR^{+}10, JTJ18, KS18b, MSA19, PRF^{+}17, RvDW10, RJP^{+}17, SLH^{+}17, TSPL14, TLPS15]. Sex- [HSV^{+}17, PRF^{+}17]. sex-changing [RJP^{+}17]. sex-dimorphic [RvDW10]. Sexual [HCF^{+}12b, HSD^{+}16, HFNH13, NLB^{+}12, TLPS15]. shad [RLD^{+}12]. shads [RDB^{+}18a]. shaken [OTM^{+}16]. shallow [DI11, MDPH12, MJS^{+}16, RDS12]. shallow-water [DI11]. Shandong [SLT^{+}17]. shape [BJH^{+}14, CSS^{+}12, HA15, IKF17, MIM^{+}19, Mer18, RT19, SPG^{+}17]. shaped [SPG^{+}17]. Shaping [DKC^{+}17, HHZ^{+}18, RIBB^{+}19]. share [Job17, SA14]. Sharing [Han13, PBE^{+}19]. Shark [JHB^{+}17, BCS^{+}17, BPD^{+}18a, BPD^{+}18b, CJFS16, CMH^{+}11, CAM^{+}15, DHS19, GMH^{+}14, HWR^{+}16, HBHR18, JGCH18, LWT^{+}11, MIJ^{+}17, MHSG19, Pas14, PHB^{+}16, SRC11, TSPL14, TLPS15, VMG11]. sharks [BNB^{+}14, KFN15, RHP^{+}12]. Shearwaters [Lav15]. Shelf [FMML19, LLF18, RRCT^{+}17, SSTB15, AIA^{+}15, AHM^{+}15, BRH^{+}15, BNP19, BYQ^{+}19, EPK^{+}18, GHD^{+}10, HCDAF15, JV12b, KdS14, LPB^{+}14, LK15, Mos18, NPB^{+}15, NST^{+}14, PCFR13, RHOJ10, RDS12, RHP^{+}12, SPFM^{+}10, SDAM^{+}18, YFG^{+}17, FFK^{+}11, KGW^{+}12, ZPT^{+}11]. shelf-spawned [NPB^{+}15]. Shell [VMAMAG11, CGK^{+}17, CPLH16, PWCS17, RRM^{+}12, RCF^{+}17, TRHB13]. shelled [BLMW17]. shellfish [BPPP17, FRM^{+}16, RCC^{+}14b]. shelves [LLL^{+}13]. shift [DPD^{+}12, FBD^{+}19, HBB^{+}13b, KGN^{+}10, KGH^{+}16, LSB^{+}10, MCSL16, MHLW12, TNS13, YFP^{+}19, YCZY19, vPBLR19]. Shifting [DMS18, GSC10, CCL^{+}14, ITT19, KPL^{+}19]. Shifting-balance [GSC10]. Shifts [BÁOMRV13, MM17, Ful11, GKC^{+}15, HTT^{+}17, MJS^{+}16, SiTiM^{+}11, SCM19, TB17, vPFF^{+}16]. shipment [BMVPN10]. ships [DWW^{+}10, DW10, DW11]. Shiretoko [MS12]. shoaling [REFJ12, TSP12]. shoot [BCRM^{+}18]. shore [TBD^{+}18]. shore-based [TBD^{+}18]. Shoreline [LKS^{+}14b]. shorelines [HGS^{+}19]. Short [AUEC12, MS14, OEK^{+}18, BCK^{+}15, CCG14, FCCLA10, JTM^{+}16, LLM12, NMF^{+}16, OSC^{+}17, dMBD11]. short- [BCK^{+}15]. short-beaked [FCCLA10]. short-duration [OSC^{+}17]. short-lived [dMBD11]. short-tailed [LLM12]. Short-term [AUEC12, MS14, CCG14, JTM^{+}16, NMF^{+}16]. shortfin [CJFS16, TSPL14, TLPS15]. Should [BKTS11, GBB16, GRH^{+}14, GR15]. show [McE17, MNWB18, NMF^{+}16]. shows [GMN^{+}17]. shrimp [AMS^{+}12, BMSC16, BBB^{+}19, EMP11, HRB15a, JTJ18, KJG^{+}15, LHS^{+}18, LGB^{+}18, LAD^{+}17, OFYS^{+}11, RGFT14, dÁMNGS^{+}11, SDH^{+}18, SCD^{+}15, STSP17, TH15, TCH^{+}16, ZHV^{+}16]. siblings [BMR10b]. Sicily [GCG^{+}10]. sicula [BHCSP^{+}11]. sidescan [SHJI10]. sighting [BVC15]. sigmoid [Hol14]. signal [LBLJ17]. signals [HGH^{+}16, LKR16, RDCPvH11, SKA^{+}12]. signatures [DBDP10, GRNG^{+}10]. significant [BHFH14]. significantly [DSA13]. signs [ASV^{+}19]. Silbiger [STF^{+}17]. Silent [DWW^{+}10, DW10, DW11]. silica [VSB^{+}16]. silicate [Rey12]. silos [Coc17]. silver [BPST^{+}16, BFS16, CMC^{+}12, HTM11, MMWC16, PDT^{+}18, PMOH13]. silver-lipped [HTM11]. silver-phase [CMC^{+}12, MMWC16]. Similar [RST18]. similarities [YDT14]. similarity [GS15]. simple [AML^{+}19, CKB^{+}16, GG13, GB15a, JMM^{+}15a, KFN15, Lév15, OCBJ10, PDD^{+}11, SBS^{+}10]. simplistic [PFH16]. Simrad [CBD10b]. Simulated [EWH16, CAB^{+}10, EGC^{+}17, SHAM10, SHS15]. simulating [GF14, MPR12, SCS^{+}11, VMAMAG11]. Simulation [DHS19, DBM^{+}15, KCS10, Elv15, HPDB19, MMA^{+}10, McB15, Nee15, PAB^{+}14, SSMD^{+}13, TK18]. Simulation-based [DHS19]. simulations [ET15, FHOK14, GFF^{+}17, HLCC16]. simultaneous [TJR^{+}10]. since [Bro14, FFPL14, KGH^{+}16]. Single [PDZ^{+}13, ANT18, EDBMB14, GCS^{+}18, MLB^{+}14, QFM^{+}10, SHK^{+}15, SHJI10, TB17, TLPS15, URV^{+}11]. single- [TB17]. single-beam [EDBMB14, QFM^{+}10, SHJI10]. single-nucleotide [GCS^{+}18]. single-sex [TLPS15]. single-species [MLB^{+}14, URV^{+}11]. single-stock [ANT18]. singular [WABZ16]. sinicus [YLLG18]. sink [SMH^{+}19]. sinking [BLMW17, LBJ15]. siphonophores [PHK^{+}19]. site [CRAM^{+}14, DPL10, DTA^{+}13, LSR^{+}13, NS17, RSY11, SMK^{+}11, ZHD^{+}14]. site-selection [LSR^{+}13]. sites [BGS^{+}14, BCOS11, Ham14, MAGK18, SHS15]. situ [BMO^{+}19, BL15, FA10, KRML11, KPS^{+}16, KOC15, Mac11, MKR13, Mel16, OOD13, Ohm19, PMOH13, TNFC16]. situation [PHM^{+}14]. situations [GKR14, SCS^{+}11]. Six [Cla18d, PMH^{+}13, SPFM^{+}10]. Sixteen [Ger17]. Sixth [DG17]. Size [dMADLP^{+}16, FC19, MM12, MNWB18, SFH^{+}12, ABF^{+}16, AS14, ACS15, AFK15, BRH^{+}15, BCB^{+}18, BMR10b, BFDGLÁ15, BMG15, CKH19, CCC^{+}14, DJW^{+}15, FPSF11, FWP^{+}16b, GHPH15, HPS^{+}15, HBB^{+}16, HSD^{+}16, HSV^{+}17, HGRAR10, HHQ11, HHQ12, HFM13, HCF^{+}12b, KPGH16, KAT11, KRKG16, LHS^{+}18, LCSI17, LPK12, MSOM10, MYI12, MSA19, MDDP10, Mil10, MHMP^{+}18, OSJ^{+}16, OFB^{+}17, PRF^{+}17, Pep16, Pla17, PFH16, Pow14, PDD^{+}10b, PHM13, RC12, RPS11, RT19, SMB^{+}18, SGK^{+}11, SMZ^{+}14, SLH^{+}17, SWA^{+}16, SHG^{+}18, SMR^{+}10, TJR^{+}10, TDN^{+}19, WKW^{+}10, ZCR15, ZH13, vDKR14, vGA18b]. size-at-age [OSJ^{+}16]. size-at-entry [vGA18b]. Size-based [MNWB18, ABF^{+}16, FWP^{+}16b, OFB^{+}17, Pow14]. size-biased [PRF^{+}17]. size-dependent [AS14, Pep16, ZH13]. Size-related [MM12, CCC^{+}14]. Size-selective [SFH^{+}12, HSV^{+}17, MSA19, SWA^{+}16]. size-selectivity [Mil10]. size-spectrum [ZCR15]. size-structured [Pla17, PDD^{+}10b, PHM13]. sized [KOT^{+}17]. sizes [BGM^{+}11, FS14, PB11, PR16, SBSE14]. sizing [MBAGVG^{+}18]. Skagerrak [ABJ^{+}19, CMH19, AASD12, HJS^{+}17, JDFN12, KJG^{+}15, NLB^{+}12, SBS^{+}19c, ZHV^{+}16]. skate [AEC11, Ben13, SPFM^{+}10]. skates [DOB^{+}17, KFN15, MSE12]. skeletal [CRC^{+}16]. skew [NNY16]. skills [OCWG16]. skipjack [BMS^{+}18]. skipper [OCWG16]. skua [CFT^{+}19]. Slave [GKR10, HGH^{+}16]. slicing [AST^{+}15a]. slipper [GDS15]. slippery [DB16]. slipping [Dek16]. slope [CNBK11, DB16, VCH16, RCH^{+}19]. slope-spawning [VCH16]. slow [ÖHSNV17]. Small [JH13, KRO^{+}16, PLCC18, APOG11, APV18, BI13, CAL14, DCVC15, FBCD^{+}17, FCG^{+}16, HOV^{+}15, JRG^{+}16, Kam14, KOT^{+}17, LAD^{+}19, LGRC14, MB14, MVK18, MCO^{+}17, MHM^{+}11, MHLW12, MSR^{+}19, MJS^{+}15b, NdlPIR18, PVQS12, PKR^{+}19, PHB^{+}16, RBM15, RR11, RFHG19, SBR11, VFR^{+}16, VHH10, WESS18, dGGW^{+}11]. Small-scale [JH13, KRO^{+}16, PLCC18, APOG11, CAL14, DCVC15, HOV^{+}15, JRG^{+}16, LGRC14, MVK18, MHM^{+}11, MSR^{+}19, MJS^{+}15b, NdlPIR18, PVQS12, RR11, RFHG19, dGGW^{+}11]. small-sized [KOT^{+}17]. small-vessel [FBCD^{+}17]. smaller [dMADLP^{+}16]. smaller-scale [dMADLP^{+}16]. smolt [CCD^{+}19, FDR^{+}17, GCS^{+}18, GIR^{+}19, HUT^{+}11, LSWD12, MGH^{+}12]. smolts [BDHC19, JÓT^{+}12, MRD^{+}12, OEK^{+}18, RAB^{+}12, VHF^{+}17]. smooth [VMAMAG11]. smoothhound [FMC10]. snapper [BTA^{+}18, SBG10, WPH^{+}15, WPH^{+}17]. snappers [MONS14, WOW^{+}17]. snow [ÉSMGB15, HCR13, MDMC11, NHL^{+}19, OSM18, SP13, Urb15]. SNP [LÁG^{+}19, VCF^{+}12]. social [BLD17, BFDGLÁ15, CHM15, FSS^{+}17, HKKS10, KS14, MTCF^{+}17, PNS18, RDT^{+}17, SBB^{+}17, TMG16]. socially [MFP^{+}19]. societies [Bas18c]. society [BLD17]. Socio [EPVC14, DKC^{+}17, TLK^{+}17]. socio-ecological [DKC^{+}17, TLK^{+}17]. Socio-economic [EPVC14]. sociological [LKHK11]. sockeye [FSN^{+}11]. soft [DDE^{+}19, MS14, OI16]. soft-bottom [MS14]. software [GÁE16]. solandri [ZGT13, ZGTL13]. Solar [DRVV^{+}17]. sole [BMM^{+}19, CDAN^{+}14, CN16, EPKR11, FB19, vHCP17, vdRMC^{+}17]. Solea [FB19, vHCP17, vdRMC^{+}17]. solutions [ACT^{+}18, Ben15, CH16]. solve [GBB16]. Some [HOS^{+}15, SMR^{+}11, DBM^{+}15, Dig19, FCPJ10, GSP12, GS11, KS14, MG15, NF16, SFG^{+}15, VAP12]. sometimes [DWW^{+}10, DW10]. sonar [DTL^{+}11, HHR17, KKLM13, Mel16, PPM13, SKA^{+}12, SKA15, SKA22, TMR^{+}17, TK18, VPH18]. sonars [SWBJ13, VPO17]. sorting [SHG^{+}18]. sound [Bjö18, FW10, HP17, TBC^{+}14, BHM13, SSF^{+}18a]. sound-speed [FW10]. source [BMM^{+}19, GZS^{+}19, SMH^{+}19, TJMC11, XCRP13]. source-sink [SMH^{+}19]. Sources [Der18, MDMC11, dQCD^{+}10, Ben13, HRB^{+}15b, LHHK12, LPD^{+}16a, SPS^{+}10]. Sousa [PWPW18]. South [MRC^{+}10, GPG14, LMG^{+}14, WNM10, ADB16, APK11, CAM^{+}15, CWRB11, CST^{+}19, EHT17, FWT^{+}14, GRP10, GD15, GKC^{+}15, Lav15, PGV^{+}17, RDS12, RBP15, VOS10, dGGW^{+}11, dMBD11, dlTQMUE10, vdHBSM10]. south-eastern [LMG^{+}14]. southeast [GHD^{+}10, JCS16, AMQRC10, SOL^{+}15, YYCC16, dGGW^{+}11]. southeastern [GS11, HD11, HCDAF15, NPB^{+}15, PCFR13, PG10, VCRPS13]. Southern [FWT^{+}14, TRSM15, AGMC14, ACA19, Ben13, BNB^{+}14, CGG^{+}11, DCG10, DCHMHQ12, GES^{+}12, GRG^{+}10, HD11, HGGJ15, JV12a, KKZ^{+}12, KDF^{+}19b, LMG^{+}14, LWT^{+}11, LSR^{+}13, MPBH11, NDM^{+}16, OOD13, PWS^{+}11, PFD^{+}16, RSY11, RHP^{+}12, SPRL18, SMNE14, STF^{+}16, SPE14, TDN^{+}19, TOA^{+}16, TCF^{+}13, VHH10, VOS10, WDOJ15, WLN^{+}13, ZHL^{+}19, dlTQMUE10, FFF16, HFSB^{+}14, LGBA11, LLST15, NPGT15, SLTL15]. southernmost [Llo17]. southwest [JCLO17, RLD^{+}12, AIA13, CMH^{+}11, GLP^{+}11, HGF^{+}18, KdS14, KDBP11, Lap11]. southwestern [KGH^{+}16, CMA^{+}16, PDS15a, PDS15b]. Sowerby [CPLH16]. sp [CR11]. space [AAP14a, AAP14b, BNE^{+}15, BN16, BPPP17, DRGCGJ17, GGT^{+}14, MWP^{+}15, MSS^{+}19, ØS15, Pay10, PSM^{+}10, PST14, RFMR10, SSS^{+}12, SBCF16, SPS^{+}10]. Spain [DCVC15, GRS^{+}14, RGGFA16, WRC^{+}12, FCCLA10, GTM12]. Spanish [CMLVGLP14, PHA^{+}10]. spanner [OCBJ10]. Sparidae [WPH^{+}15]. Sparus [FSC^{+}12]. Spatial [BCBI13, BMH^{+}16, CMH^{+}11, CLV^{+}14, CWRB11, EPPL^{+}11, FOT^{+}17, GMC17, HSGL18, KCA^{+}14, KBT14, LDD^{+}10, LTQ14, LHH^{+}13, PQR10, QCA^{+}18a, QCA^{+}18b, RSBC13, RDD^{+}10, RHSH16, RRCT^{+}17, SG17, SRS11, SR17, SSGH18, SPM^{+}17, TSS^{+}13, VHGTFR10, Ada17, ACCA19, AHK^{+}18, BNE^{+}15, BMC^{+}18, BTTV^{+}17, BMC^{+}13, BHS^{+}16, CRHM19, CN16, DCHMHQ12, DMT10b, EPK^{+}18, EDP14, FOC^{+}18, GZS^{+}19, GL11, GR15, GCS^{+}17, GJWS19, HPS^{+}15, Ham14, HCC^{+}14, HAR19, JLH12, JOSO16, JH13, JVPM12, KCS10, KS14, KRO^{+}16, LJB16, LCG^{+}18, LVPK10, LBK17, LBCOJ19, MFP^{+}19, MTP^{+}12, MPR17, MRC^{+}18a, MRC^{+}18b, MLMJ17, MSR^{+}19, MSL^{+}11, MJSB14, dSMGKP12, NMO^{+}17, NUÓ^{+}16, OLW^{+}14, OJK^{+}19, OFVF14, PSH^{+}11, PVQS12, PGV^{+}17, PDH^{+}14, PBL11, Por11, PGN^{+}11, PCS^{+}11, PHH^{+}16, RGFT14, SBG10, SGV18, SGK^{+}11, SDAM^{+}18, STR16, SRB^{+}14, SFG^{+}15]. spatial [SWT^{+}19, TBC15, THC16, URE^{+}18, VRF^{+}16, WHAA^{+}18, WKW^{+}10, WRDF10, YS15]. spatial-scale [dSMGKP12]. Spatially [NRTK^{+}19, PDMG11, BOP^{+}19, CN16, DMBD10, GHPH15, GLC15, GBJ^{+}15, MV13, MONS14, MR14a, MBA^{+}19, SHZ^{+}16, BSW^{+}11b]. Spatio [CTH^{+}19a, CTH^{+}19b, HHH12a, HHE^{+}16, PCB^{+}14, SHW^{+}18, SRRZ16, THSM^{+}18, VKS^{+}12, VRL^{+}14, YLLG18, CAC^{+}18, FVSL^{+}10, GHB^{+}14, GTBT18, GT19, HSGL18, KRG^{+}12, LYW^{+}18, Nee15, OIMP18, VE13, ZCH19]. Spatio-temporal [CTH^{+}19a, CTH^{+}19b, HHH12a, HHE^{+}16, PCB^{+}14, SHW^{+}18, SRRZ16, THSM^{+}18, VKS^{+}12, VRL^{+}14, YLLG18, CAC^{+}18, FVSL^{+}10, GHB^{+}14, GTBT18, GT19, HSGL18, LYW^{+}18, Nee15, OIMP18, VE13, ZCH19]. spatio-temporally [KRG^{+}12]. spatiotemporal [HFGT18, PWPW18]. spawn [MHE^{+}16]. spawned [NPB^{+}15]. spawners [GMN^{+}17, GR15]. Spawning [GIW16, NWM^{+}12, RC12, SBL^{+}17, ZHD^{+}14, AIA^{+}15, ABHT^{+}16, BGS^{+}14, BSW^{+}11a, BDCvD^{+}10, BCOS11, CPVS14, CB12, CBO15, CBW15, CMD17, DPBM12, DMZC17, DHZA14, DRB10, FB19, GMN^{+}17, GMS^{+}18, GCG^{+}10, GOS^{+}13, GRH^{+}14, GKR14, GR15, HFGT18, HSK^{+}14, HOV^{+}15, HOS^{+}15, HHH12b, KS18b, KR12, LDD^{+}10, LFM^{+}18, LB13, LØM^{+}18, LHV^{+}16, LFMJ19a, LFMJ19b, Loh11, LVPK10, LBTS^{+}19, MO18, MM17, MGBvD14, dSMGKP12, MLLL11, NLB^{+}12, OTAK15, Ósk18, OAT^{+}12, PGV^{+}17, Pay10, PDAC^{+}16, PKHG14, PO14, RDS12, RTAT17, ROK^{+}18, RYS11, RCH^{+}19, Sas19, SMNE14, SLH^{+}17, SiTiM^{+}11, SS17, SMK^{+}11, SJRB18, SRKV15, TJMC11, VCH16, VHS^{+}10, VKS^{+}12, WNM^{+}13, ZLC^{+}17, ZDD^{+}19]. spawning-per-recruit [HOS^{+}15]. spawning-site [SMK^{+}11]. spawning-stock [PO14]. spawning/nursery [PGV^{+}17]. spear [TNS13]. Spearfishing [SMB^{+}18]. speargun [SMB^{+}18]. Special [Ano12e, BMNKA^{+}16, HLL18, Kim12, Rod10, TNS13, dGGW^{+}11]. specialization [CSTJ19, GMKS11, TSBV^{+}15]. specializations [SSA^{+}18]. specie [RSV^{+}17]. Species [AFH^{+}17, CNBK11, CCC^{+}14, OdJN^{+}12, URE^{+}18, ASDSM18, AHR^{+}19, ACS15, APV18, APHC15, AHC15, Bak14, BPKH13, Ben13, BTTV^{+}17, BVW^{+}18, BBD^{+}13, BHM^{+}18, BG18, CMA^{+}16, CSBHS^{+}11, CGK^{+}17, CSTJ19, CPP17, DMF13, DPRG^{+}18, EAB^{+}17, Eri16, FMK10, FOT^{+}17, FS14, GGTW17, GRP10, GMH^{+}14, GFMO11, GIDP^{+}18, GGIC^{+}15, GRF^{+}12, GES^{+}12, GJWS19, HMN^{+}12, HKRM18, HGF^{+}18, IGGSAL^{+}15, JCLO17, JMM19, JBE14, JPTC16, JVRT10, KN17, KFN15, Lap11, LCG^{+}18, LHS^{+}19, LSCL18, MMC19, MPR12, MPL^{+}14, MRC^{+}18a, MRC^{+}18b, MBLS15, MDS13, MCOS17, ML12, MLB^{+}14, NCH17, NBS^{+}10, OLM^{+}15, PWCS17, PMH^{+}13, PMM^{+}14, PFH16, PGN^{+}11, PKK13, PLSD16, PHM13, PMBM18, RD14, RvDW10, RFT^{+}16, SHH^{+}17, SSM^{+}16, SWB^{+}14, Sei14, SPFM^{+}10, SV10, SFH^{+}12, SSM^{+}18, SDBB15, SMSR10, SBKA16, SSTB15, THHH15, THHH18, URV^{+}11, VAP12]. species [VELT14, WGCL^{+}19, WMN^{+}11, WHMP15, WHP16, YDT14, YFP^{+}19, ZGK^{+}17, ZMF12, dMBD11]. Species- [CCC^{+}14]. species-based [APHC15]. species-level [GRF^{+}12]. Species-specific [URE^{+}18]. species-transformation [WHP16]. specific [Ada17, AFH^{+}17, CHD^{+}18, Dav11, HAF^{+}16, LWN^{+}13, MRC^{+}10, Pow14, URE^{+}18, VELT14]. spectra [FC19, JLL17, LPK12, LS10, MYI12, RPS11]. spectral [GZS^{+}19]. spectrum [ZCR15]. speed [BBL^{+}15, FW10, LØM^{+}18]. Sperm [GRF^{+}16, SOL^{+}15, GLP^{+}11, HLD10, OSL^{+}15, PRF^{+}17, TMS^{+}15, TTR^{+}19]. spermatophore [BMG15]. spermatophores [HLD10]. Sphyraena [ODSC10]. spider [VBGG^{+}11]. spill [BOP^{+}19, SHW^{+}18]. spills [MGVH^{+}10, VRL^{+}14]. spinax [MIJ^{+}17]. spined [OJK^{+}19]. spiny [BHC^{+}15, BFDGLÁ15, BMG15, BBM15, CZBFLÁ^{+}15, CHM15, GBM^{+}19, GBB15, HBHR18, LHJ15, YJBL17, dlTQMUE10]. Spitsbergen [KGW^{+}12]. split [CD14, FA10]. split-beam [CD14, FA10]. Splitnose [LTGP18]. spoke [AHC15]. Spondyliosoma [PVCB17]. sponge [BKMdMS13, RWG^{+}18, VSB^{+}16]. spots [SSR^{+}16]. spotted [BFDGLÁ15, LGR^{+}19]. spp [CCL^{+}14, FGMC^{+}13, Ham14, KGH^{+}16, NPB^{+}15, SLM^{+}11, SHG^{+}18, TSP12]. SPR [HOV^{+}15]. Sprat [PvdKE^{+}12, Eer12, GSL^{+}11, JMF12, KBH^{+}18, OLR^{+}18, RTJS14, VHQ^{+}11]. sprattus [Eer12, Eer12, GSL^{+}11, KBH^{+}18, RTJS14]. spread [KPBB15]. Spring [LHJ12, BL15, CHMY15, ET15, GPG14, GOS^{+}13, KOT^{+}17, LSF^{+}15, LK15, MGBvD14, Ósk18, RHV14, TRSM15, VKT^{+}12, VHS^{+}10, VKS^{+}12, WAH^{+}16, YCZY19]. spring-bloom [VKT^{+}12]. spring-spawning [GOS^{+}13, MGBvD14, Ósk18, VHS^{+}10, VKS^{+}12]. Spur [FMM^{+}12]. spurdog [AJM19, DED13]. squaloid [VMG11]. Squalus [AJM19, DED13]. square [FMK10, HWK^{+}15, SHG^{+}18]. square- [FMK10]. squat [CAMM12]. squid [ASI^{+}16a, ASI^{+}16b, AL10b, BPH16, BNP19, CRAPMN12, CRAM^{+}14, Che10, CJP^{+}14, DRB10, FBD^{+}19, GRP10, GRNG^{+}10, HLD10, KAT11, KGN^{+}10, MRC^{+}10, MAS19, PG10, RDS12, RB10, RYS11, SV10, SFW^{+}12, SHL10, TNS13, YYCC16, YCZY19]. Sr [PBW15, PBW15]. Sr/ [PBW15]. SSB [PO14]. SST [PCB^{+}14]. St [Ben13, BBB^{+}19, BCOS11, CCD^{+}19, LFGW13, LFM^{+}18, LSWD12, MPL^{+}14, MGS^{+}15, OSBG16, OSM18, PMM^{+}14, TCBD10]. St. [VWMS19]. stability [HTKB19, MPJ12, RBMAOCL18, SSB^{+}10b]. stabilize [RDGP13]. Stable [DBDP10, MTP^{+}12, TMP12, BCKA11, CCC^{+}14, DPD^{+}12, GRNG^{+}10, JvdM15, KIA^{+}18, KSJ14, OBLP^{+}19, PFPG12, RKW^{+}17, TVJ^{+}14, YDT14, YPN^{+}18, dBWMD^{+}14]. stage [CKD^{+}17, DDF^{+}12, ES18, GRC^{+}14, HHH12a, IFU11, MSMW18, RDCPvH11, TSPL14]. stage-based [MSMW18, TSPL14]. stages [BTB^{+}17, CYLT16, COHdP19, CAOG15, ÉSMGB15, FRF14, GKV^{+}15, GHM^{+}16, GES^{+}12, LEE17, MFB17, NKSE14, OSM18, PKHM12, SiTiM^{+}11, SHS15, VRL^{+}14]. stainless [MBBC11]. stainless- [MBBC11]. stakeholder [LAD^{+}11, OD17]. stakeholders [BLJ^{+}17, HHZ^{+}18, SPG^{+}17]. Stalked [JCSC10]. standard [AGSPH14, FMK10]. standardization [ZCH19]. standardizations [BKSB12]. standardized [MONS14, OFB^{+}17]. Standardizing [LPT10, Tho14]. standing [SLT^{+}14]. starry [FMC10]. STARS [SCM19]. start [GFF^{+}17]. starvation [BSHC17]. state [BN16, CSJ16b, DRGCGJ17, JIB^{+}12, JIB^{+}14, MWP^{+}15, ØS15, PVQS12, Pay10, Pla16, PSF12, RFMR10, SRDC^{+}14, SBCF16, SPS^{+}10, WHH16]. state-space [BN16, DRGCGJ17, MWP^{+}15, Pay10, RFMR10, SBCF16, SPS^{+}10]. states [GFF^{+}19, SCY^{+}10, SBS^{+}10, BT15, GS11, GHD^{+}10, MTLG14, RSBC13]. Station [Can16, TPBL17]. stationarity [WRDF10]. stationary [JSOO12, KGP^{+}15, LCG^{+}18, SH16]. stations [Can16]. Statistical [GJWS19, DUM^{+}12, HLCC16, JOSO16, MMRG18, OLM^{+}15, TM15]. statistics [Wel11]. status [ABRL12, Ano19, BCT^{+}10, CAC^{+}16, Cha12, CSY^{+}10, DCMGB^{+}17, FWC^{+}18, FWC^{+}19a, FWC^{+}19b, GCJL16, GSMRM^{+}16, HPCW19, IOK17, KMF13, LLL^{+}13, Lav15, ODSC10, ØHN14, PKK13, SFNdH16, SS10a, SSB^{+}10a, SY16, Tho11, WGCL^{+}19, dGGW^{+}11]. statutory [Tur19a, Tur19b]. stealing [RLW^{+}15]. steel [MBBC11]. steelhead [HAF^{+}16, HHM^{+}16, MHE^{+}16]. Steindachner [dSMGKP12]. stenolepis [Loh11]. step [NPB^{+}15]. steps [Cam18, SBB^{+}17]. stereo [KOC15, SHS^{+}17]. stereoscopic [MBAGVG^{+}18]. stern [SJUE11]. stewardship [Rid18]. stick [MR10]. stickleback [OJK^{+}19]. still [BS14, HK14, MP15]. stimulation [HWK^{+}15, dHFF^{+}16]. stinger [BSK^{+}11]. stingray [LLM12]. stingrays [LLM12]. Stirred [OTM^{+}16]. Stochastic [MHHT10, SBH^{+}14, TJH15]. Stock [CDC15, Coo19, FCPJ10, PSM^{+}10, THM15, WW16b, ZMKC14, ZGTL13, AJM19, ANT18, ACCA19, ABRL12, AFQ^{+}11, ALBR14, AFK15, BNE^{+}17, BN16, BCS^{+}17, BHS^{+}16, BPD^{+}18a, BPD^{+}18b, BWP10, Cad13, CSS^{+}12, CTC14, CMU^{+}17, CAM^{+}15, CECL16, CCD^{+}19, CIJ18, DED13, DEH^{+}19, DB16, DBM^{+}15, DMO10, DCNB^{+}10, EHHH14, EHB^{+}15, Elv15, FWC^{+}18, FWC^{+}19a, FWC^{+}19b, GÖN^{+}12, GFF^{+}17, GMD^{+}11, GCW^{+}18, GHM^{+}16, GLC15, GKJ^{+}15, GSC10, GRC^{+}14, HvDH^{+}10, HA15, HTM11, HAF^{+}16, HMFW14, HPCW19, HFSV^{+}15, HRP^{+}16, HHE^{+}16, HMR15, JMM^{+}15a, JOSO16, JMM^{+}15b, JKS^{+}18, KHC^{+}17, KET^{+}17, KZS10, LTGP18, LB13, LGvPH15, LWN^{+}13, MV10, MIM^{+}19, MWP^{+}15, MKL^{+}19, MGPC^{+}14, MHE^{+}16, MP15, MIJ^{+}17, MR10, MKC^{+}11, MJS^{+}15a, MBT^{+}19, MBP19, OLW^{+}14, OSJ^{+}16, OdS19, OHD^{+}16, OVW^{+}17, PBW15, PB11, PNS18, PRBF18, Pow14, PMOH13, PO14]. stock [PSS11, PHM13, RKL^{+}10, RBA^{+}19, RvDW10, RJ12, SPS^{+}10, SCS^{+}11, SWA^{+}16, SH14a, SHZ^{+}16, SMR^{+}10, SM15, SSZH12, SDSN14, Sub18a, Sub18b, SBSE14, SC16, SFN^{+}19, SIP18, Tho11, Tho14, TJH15, VE13, Wri14, ZRN18, ZHV^{+}16, ZE17, vGA18a]. stock-assessment [ABRL12]. stock-recovery [HvDH^{+}10]. stock-specific [HAF^{+}16]. stock-status [Tho11]. Stocking [BFS16, BMVPN10, EHD^{+}15, JTM^{+}16, PR16, PST14, SRKV15]. stocks [AV15, AHM^{+}15, AL10a, BRH^{+}15, BVS^{+}16, BBB^{+}19, Bru10, BP13, Cad14, Che10, DZ14, FR16, GNFM11, GNKS18, GLC15, GKC^{+}15, HKKS10, HSSB14, JCS16, JVPM12, Mac12, MHM^{+}11, MG15, MJSB14, MRD19, MLB^{+}14, MSR14, Mur10, NMO^{+}17, OFB^{+}17, PDZ^{+}13, POIM12, Par13, PNC15, PSS11, RFHS14, Sch14, SBFC10, SPM^{+}18, SZM^{+}10, SFN^{+}19, VEA^{+}10, WPeEA16, WÓG^{+}16, WMG^{+}17]. stomach [BCBI13, Der18, MDV^{+}15, YPN^{+}18]. stones [GLP^{+}11]. stop [Moo19]. storage [AJ12, HLD10, KJÓK16]. stories [CSJ16b, WHH16]. storm [MS14, MFT16, SZOL19]. strain [FLCQ19]. Strait [OTM^{+}16, GCG^{+}10, GDBM^{+}17, SQ14, ZHL^{+}19]. strandings [VG15]. strange [Ric19]. Strategic [Pun17, DBM^{+}15, YS15]. strategies [DFG^{+}17, GFF^{+}19, HABV19, HvDH^{+}10, HKKS10, HMHL19, HBiI^{+}11, IHH^{+}11, JCA10, LCMR^{+}18, MV13, MS10, Nee15, PEV^{+}16, PDD^{+}11, SDT15, SSH^{+}14, Tow14, WPeEA16, YFP^{+}19, dlPG15]. Strategy [GHWD13, GRF^{+}12, PKK13, SRDC^{+}14, SGP^{+}15, BNK10, BHDB12, Bow14, BBD^{+}10, DPOL13, DPRG^{+}18, HLD10, HB10, HFHS10, KGZH16, NDP^{+}16, QZH^{+}17, RR10, SP13, TD19]. stratification [CSFS10]. stratified [CD11]. stream [FWRM12, WCT^{+}16, WDE^{+}18, BNP19]. streamline [MJS^{+}15a]. strength [BMS^{+}18, BMO^{+}19, CK12, CD14, DBL^{+}16, FOJ13, FA10, GHM^{+}16, KPJC14, KRML11, KO12, Mac11, MKR13, MMG^{+}17, OOD13, PGOM11, PJDN12, RMJF14, SGM^{+}17, SRW13, SHL10, TRHB13, TLF16, TKA^{+}16, TK18, VMAMAG11, YST^{+}10]. Strengths [KKC19, GRKF16, RFMR10]. stress [AS14, BMVPN10, BSHC17, CPT^{+}10, GQX^{+}19, JGCH18, MSM17, NLB^{+}12, RHOL11, TVO12]. Stressful [OBJA^{+}18]. stressor [NDM^{+}16]. striata [BSB^{+}13, KS18b, SBCF16]. Strike [CBC^{+}11a]. striped [KDBP11, SHK17, SSP^{+}13, VWMS19]. strobe [WLL^{+}13]. stronger [LSF16]. Strontium [PBW15]. structural [BNB^{+}14]. structuration [PG16]. Structure [FVSL^{+}10, BJH^{+}14, BCB^{+}18, BHB^{+}19, BS13, BCS^{+}17, BHS^{+}11b, Bru10, BP13, BFB^{+}11, CBD^{+}10a, CMP^{+}11, CRHM19, CKH19, CD11, CKD^{+}17, CWRB11, DMF13, Elv15, EPPL^{+}11, FFN^{+}13, FFN^{+}16, FC19, FPSF11, FSC^{+}12, FMM^{+}12, FDC^{+}16, GSC10, GMKS11, GFR^{+}12, GCS^{+}17, HG10, HSNO12, HRB^{+}15b, HJS14, IEL^{+}15, JDM^{+}16, KCS10, KHC^{+}17, KAT11, KJG^{+}15, Lap11, LLM12, LÁG^{+}19, MAA^{+}19a, MAA^{+}19b, MAP^{+}11, MMRG18, MNV^{+}11, MIJ^{+}17, MST^{+}18, PGG^{+}12, PSH^{+}11, PSDG12, PSO^{+}14, PGV^{+}17, PNS18, PBG14, PHB^{+}16, PCS^{+}11, PLSD16, RvDW10, RRM^{+}12, RTAT17, SGK^{+}11, SMZ^{+}14, SG16b, SMH^{+}19, SWA^{+}16, STSP17, VCRPS13, VMG11, VSP^{+}14, WKW^{+}10, WKK^{+}15, WSG^{+}17, WSW^{+}19, ZHD^{+}14]. structured [BHL^{+}15, BPC10, DGC12, DMBD10, HHQ12, HFSV^{+}15, OYIN18, OLM^{+}15, Pla17, PDD^{+}10b, PHM13, RC15, SBH^{+}14]. structures [GOH^{+}19, TUUC14]. structuring [JH13, SSS^{+}12]. studies [CCD^{+}19, DHAH12, GTGD15, HTB^{+}16, SD17, SCL^{+}19]. study [AFLvDH15, AFK15, BMC^{+}18, BFC^{+}16, BBB^{+}19, Can16, CAM^{+}15, CLS^{+}18, CGG^{+}11, CSR11, CLMP16, DFB^{+}10, DCMC19, DD10, EHT17, Elv15, FB19, FBM^{+}18, FHH^{+}13a, FHH^{+}13b, GB15a, HPDB19, HR14, HMS^{+}13, HHBSM16, KRML11, LGL^{+}12, LAG^{+}16, LLG18, LHV^{+}16, LBB15, LCG^{+}18, LHK18, MCAM14, MR14a, MHMP^{+}18, NSS15, PBJ^{+}14, PSH^{+}11, PNS18, PNC15, PCdL15, Pin17, PDR19, RGGFA16, Sas19, SBH^{+}14, SOL^{+}15, SWRC^{+}18, TCS^{+}19, TCQ^{+}15, TMW19, TLPS15, TMG16, WPvBS15, WHNS14]. studying [Qui18, TMS^{+}15]. sturgeon [BFH^{+}18]. Stylasteridae [BHCSP^{+}11]. stylet [DWPS11, HRF^{+}10]. stylets [BA10]. sub [ASV^{+}19, DEH^{+}19, GW16, LWN^{+}13, SAS^{+}18]. sub-Arctic [SAS^{+}18]. sub-grid [GW16]. sub-populations [DEH^{+}19]. sub-regional [ASV^{+}19]. sub-stock-specific [LWN^{+}13]. Subantarctic [TRSM15]. Subarctic [DMS18, LHJ12, GS12, KKZ^{+}12]. Subarea [PGN^{+}11]. subcontinent [SLKS10]. subdecadal [MW19]. Subdivision [CNBK11, YJBL17]. subject [PRF^{+}17]. subjected [EMW^{+}16]. Sublethal [CPT^{+}10, FHD^{+}19]. sublittoral [GDD^{+}18]. Submerged [HBB^{+}13a]. submesoscale [SS17]. subpolar [DHAH12]. subpopulations [FB19, MJAS14, ØU13]. subsequent [RAB^{+}12]. subsidies [Sum13]. Subsistence [MCO^{+}17, CBL19]. substantially [HUPBL18, MCO^{+}17]. substituted [vdHBSM10]. Substituting [Eid18]. substock [dCWMH13]. substrata [VKS^{+}12]. substrate [BMF14, CCSG18]. substratum [DTL^{+}11]. subsurface [WSGD^{+}17]. subtidal [GFSC17]. subtle [SHW^{+}18]. Subtropical [PPHK19, FGMC^{+}13, KdS14, LBM19, PDWH11, RC15, WPH^{+}17]. success [BvdMA^{+}18, BLM^{+}17, BMG15, GRF^{+}16, HvDH^{+}10, HO14, LGR^{+}14, PRB^{+}15, SKKM15, SLF16b, VKT^{+}12]. successful [Ove19, SOL^{+}15]. such [TSP12]. sufficient [McE17]. suggest [FFRR12, HSNO12, SV10]. suggestions [SBS19a, SBS19b, SCL^{+}19]. suggests [KRO^{+}16, MKL^{+}19, RDB^{+}18b, RCF^{+}18]. suitability [BPF^{+}18a, BPF^{+}18b, CSC^{+}13, EP14, HDCH^{+}11, HHH12b, RCH^{+}19, YYCC16, YCZY19]. suitable [EF15, GIPS12, NRTK^{+}19, PMP^{+}16]. sulphides [Ham14, Van11]. summer [BHMR14, HHM^{+}16, KSJ14, KYK^{+}12, NSO^{+}19, SKA^{+}12, SBLS15]. summer-distribution [NSO^{+}19]. summers [MYI12]. Summing [Mur11]. SumWing [DDE^{+}19]. sun [GCBI17]. superba [CK18, CWRB11, FWT^{+}14, KOC15]. supervised [CSS^{+}15]. supplementary [GJWS19]. Supply [CC11, BDF^{+}14, HHZ^{+}18, JFJ^{+}17]. Support [DOB^{+}17, AHK^{+}18, BFH13, DMS^{+}12, Gro11, HKS17, Jen13, KLKD11, KJG^{+}15, LLSJ18, MV10, MNV^{+}11, MLMJ17, MTCF^{+}17, MNG15, RMF^{+}15, SDDA11, THW16, vdKFS^{+}16]. supporting [BMC^{+}18, LBM19, PKHM12, VF17]. surface [BdSP11, CD14, DLSJ^{+}19, ET15, KLC13, MTO^{+}18, OFYS^{+}11, PMM^{+}14, RB10, RT19, TSY^{+}11]. surfclam [MBC11]. surgically [BDHC19]. surrogate [PCS^{+}11]. surrounding [THR16]. Surveillance [SGP^{+}15]. Survey [HKS12, CTH^{+}19a, CTH^{+}19b, DMT10b, GRS^{+}14, GHG^{+}10, GTBT18, HHQ11, HMR15, JVPM12, KFBG15, KHPI15, NBS^{+}10, OYIN18, OFB^{+}17, RBCH10, SRCR12, SHS11, SH14a, SR16, TRH10, TKJB19, URE^{+}18, UPF^{+}10, VQB^{+}11, WMLJ17, WW16b, WRDF10, ZRU^{+}15, ZKD^{+}14, vdHdGL16]. surveyed [SRGF15]. surveying [VKN^{+}11, WLL^{+}13]. surveys [ABJ^{+}19, BMC^{+}13, CMH19, CSR11, CSFS10, CBK18, DWW^{+}10, DZC^{+}13, EAOB15, FCG^{+}16, FHD^{+}19, HKRM18, JdFBB^{+}10, KJÓK16, KIP14, Loh11, MST^{+}18, Nel19, OIMP18, PHO13, SSP12, SMD^{+}16, SBS^{+}19c, STB^{+}11, WWI^{+}16, ZE17]. Survival [HPN^{+}17, LVSF17, PKHG14, vdRMC^{+}17, ANT18, BDHC19, BBHC11, COR15, CGAD16, CCD^{+}19, CPP17, FWH^{+}15, FDR^{+}17, HGGJ15, HHH12a, Hol10, KBT14, KLCC18, KFBG15, LCSI18, MRD^{+}12, MA13, MRMK11, OSM18, PWCS17, PKHM12, PDAC^{+}16, RBA^{+}19, RC12, RAB^{+}12, TVO12, TPHdJ19, TMP12, UTA^{+}16, VDK^{+}18]. survive [FWH^{+}15]. survivors [GES^{+}12]. susceptibilities [BPKH13]. susceptibility [BHM^{+}18, WESS18]. suspension [LNWS18]. Sustainability [TCF^{+}13, AGSPH14, ASB^{+}19, GAB^{+}14, HHZ^{+}18, JHB^{+}17, NJFH19, Ric14, VOM11, ZMF12, ZHV^{+}16]. Sustainable [KK19, OND^{+}18, BNE^{+}17, BPD^{+}18a, BPD^{+}18b, BSB^{+}19, CDC15, FR16, LB13, LWN^{+}13, MTLG14, NCH17, RCH^{+}15, SMR^{+}11, UVD^{+}17, vGA18b]. sustainably [HB19b]. sustained [FO13, HLL^{+}12, dPJGB13]. Sustaining [She15]. Svalbard [AV15]. Svedäng [CMH19, WW16a]. Sverdrup [Aks15, Fra15, LBJ15, SJB15]. SW [BJH^{+}14, CLdR^{+}15]. swarms [KBB13, RT19]. Sweden [AFP12]. Swedish [ABJ^{+}19, CMH19, BLJ^{+}17, HSHÖ14, HJS^{+}17, ÖHSNV17, SBS^{+}19c, WP13]. Swimbladder [YST^{+}10, FOJ13, HdBR^{+}16]. swimbladders [SCJI10]. swimming [BLMW17, BBL^{+}15, CK12, CGAD16, SPS15, SHL10, SSA^{+}18, SJ16, TKA^{+}16]. switching [SRH^{+}14]. swordfish [AMQRC10, BGA^{+}19, CSC^{+}13, GMD^{+}11, SSA^{+}18, VPSV^{+}10]. sympatric [CCSG18, MPL^{+}14, MDS13, PMM^{+}14, RHP^{+}12, YDT14]. sympatry [MRC^{+}18a, MRC^{+}18b]. Symphodus [HSD^{+}16, HSV^{+}17]. symposia [Mäl12]. Symposium [GGK^{+}11, LKG^{+}19, BMNKA^{+}16, DGB^{+}15, PNE^{+}14, WRC^{+}12, DG17, MNG15, Rod10]. synchronicity [BSW^{+}11a]. synchronies [Kam14, TUUC14]. Synchronous [QOH^{+}13]. Synchrony [TNS13, JOSO16]. synchrony/asynchrony [JOSO16]. syndrome [KUM^{+}12, MKC^{+}11]. Synergistic [DI11, WABZ16]. synopsis [PWS^{+}11]. synoptic [HSB16]. Synthesis [THM15, DGB^{+}15, DK18, GGK^{+}11, HCE^{+}11, SOGT^{+}19]. synthetic [AHR^{+}19, KGP^{+}15, SCM19]. synthetic-aperture [KGP^{+}15]. System [GML13, LJH^{+}12, LSJ10, MKD^{+}18, RBCH10, BUNB10, BFS16, BPMR17, CR19, CBC^{+}11b, DAP^{+}12, Ful11, GLP^{+}11, JBSD11, JL12b, KLKD11, MNV^{+}11, MBC11, MTLG14, MBAGVG^{+}18, Peñ19a, PCS^{+}11, ROBC11, RLD^{+}12, RK16, SHST18, URE^{+}18, WLL^{+}13, WHMP15, vPBLR19, QCA^{+}18a, QCA^{+}18b]. system-identity [vPBLR19]. Systematic [DMS^{+}12, EPH15, TBHK17]. Systems [GL11, MHHK13, BJH^{+}16, BBM15, CJC^{+}15, DKC^{+}17, FSS^{+}17, HHZ^{+}18, LGBA11, MSC14, MTCF^{+}17, NST^{+}14, PCdL15, PDT^{+}18, PRO14, RAH^{+}16, SCY^{+}10, SMB^{+}16a, TLK^{+}17].

T. [WLN^{+}13]. TAC [ESAV15]. TACs [PSGY^{+}12, URV^{+}11]. Tag [BDHC19, CAB^{+}10, GLC15, GOS^{+}13, JHB^{+}12, MSR^{+}12, RHB^{+}12, RHB^{+}13, WHNS14]. tag-integrated [GLC15]. tag-recapture [WHNS14]. tagged [BKTS11, STC^{+}17, VBW^{+}18]. tagging [BKTS11, BGA^{+}19, CAB^{+}10, KFBG15, MDdPMQ10, MJSB14, PE18, SSRT10, dPJGB13]. tags [AJ12, AGG^{+}18, CAM^{+}15, HLA^{+}11, JRG^{+}16, KJÓK16, LFM^{+}18, Loh11, SLS15]. tailed [LLM12]. Taiwan [KCP^{+}11, ZHL^{+}19]. take [BM12, CGRM^{+}13, SG18b]. taken [SSZH12]. Taking [OdS19]. tale [JMM19, KM16]. tanker [ÖHSNV17]. tanks [BMVPN10]. Tanner [HCR13, LSF16, PFD^{+}16, SZM^{+}10, SLF16b]. Target [BMS^{+}18, GRKF16, KO12, ACS15, BMO^{+}19, BFS16, CK12, DBL^{+}16, FOJ13, FA10, KPJC14, KRML11, Mac11, MKR13, MMG^{+}17, MVP^{+}14, OOD13, PGOM11, SGM^{+}17, SRGF15, SRW13, SHL10, TKA^{+}16, WGCL^{+}19, YST^{+}10]. targeted [IGGSAL^{+}15, PST14]. targeting [GDHFS^{+}18, NC11, SRCR12]. targets [GRF^{+}12, GR15, RDT^{+}17, SDSA15, SMD^{+}16]. tariff [KRG^{+}12]. tariff-based [KRG^{+}12]. Tasmania [MHLW12]. taxa [HBF14]. taxonomic [CNBK11]. taxonomical [RMJF14]. team [DGL^{+}17]. teamwork [Pet19]. tech [dCVB14]. Technical [OCG10, CMM^{+}15, KFN15, MTDP11, TRPH16]. technique [BA10, Bjö18, BCSG13]. techniques [BCOS11, CSS^{+}15, COS11, GHG^{+}10, ROBC11, SR17]. technological [CBC^{+}11a]. technologies [ZDD^{+}19, vdKKS11]. technology [Mac17a, MVH^{+}11, MJSB14]. technology-induced [MVH^{+}11]. teeth [FFH^{+}11]. telemetry [CCD^{+}19, DRB10, LDT^{+}11]. teleost [SLS15]. teleosts [MBBC11]. tell [GDJ11b]. Telomere [GFMO11]. temperate [BHDB12, BBM15, CBBM14, DJW^{+}15, GMH^{+}14, GD15, GFML10, HDG^{+}17, KPD^{+}16, PDWH11, RCF^{+}17, SSS^{+}12, SJ16, WPH^{+}17, WHL^{+}17, YST^{+}10]. Temperature [CMD17, LCSI17, VHQ^{+}11, BWM^{+}16, BMFSH12, BBL^{+}15, BSHC17, BdSP11, BKT12, CGK^{+}17, CCSG18, CK18, CN16, DPBM12, DI11, DRB10, GDD^{+}18, KLCC18, KCL18, LCSI18, LBG18, LSCL18, MSOM10, MMDSP15, MTO^{+}18, NDP^{+}16, Ott10, OSM18, QR15, RAH^{+}16, RFHS14, SSH^{+}17, SSM^{+}16, TMP^{+}17, TSY^{+}11, VHS^{+}10, WPH^{+}15, WWMF17, WPvBS15, ZSC16, vdSSM^{+}18]. Temperature-based [CMD17]. Temperature-dependent [LCSI17]. temperatures [AGMC14, GvRKB19, HHM^{+}16, KLC13, KS17, MM17, SLF17, VSB^{+}16]. template [FFK^{+}11]. Temporal [AGMC14, GZS^{+}19, HPS^{+}15, HFSB^{+}14, OTO^{+}15, ROK^{+}18, SMB^{+}16a, WKW^{+}10, vDHvKR15, BGM^{+}14, BHJ14, BÖBL12, BHS^{+}16, BMH^{+}16, CAC^{+}18, CTH^{+}19a, CTH^{+}19b, CWRB11, FVSL^{+}10, FOC^{+}18, FOT^{+}17, GHB^{+}14, GTBT18, GT19, HSGL18, HHH12a, HHE^{+}16, JCSC10, KBT14, LDD^{+}10, LTQ14, LYW^{+}18, LBK17, MSR^{+}19, NUO^{+}10, Nee15, OIMP18, PHGK^{+}16, PCB^{+}14, RSBC13, SHW^{+}18, SRRZ16, SSGH18, SPM^{+}17, THSM^{+}18, TSS^{+}13, VGWJ11, VKS^{+}12, VRL^{+}14, VE13, WCT^{+}16, YLLG18, ZCH19]. temporally [KRG^{+}12, MONS14]. temporary [DLL12, RCF^{+}18]. tentacles [Ohm19]. term [AFLvDH15, AUEC12, BCK^{+}15, CAC^{+}18, CCG14, DG11, DL11, DLSJ^{+}19, FKH^{+}16, GPG14, GTGD15, GAB^{+}14, HHM^{+}11, HBD^{+}19a, HBD^{+}19b, HLD10, JTM^{+}16, KOK^{+}14, KFBG15, LGBA11, MCSL16, MS14, NMF^{+}16, OS14, PB11, PEN^{+}19b, PDT^{+}18, PFD^{+}16, RHOJ10, RLP^{+}13, RCF^{+}17, SGV18, SWA^{+}16, SLF16b, TNFC16]. terminal [OYI17]. tern [JMF12]. terrestrial [DMS16, SBS19a, SBS19b]. territories [BM16]. test [RD14]. test-bed [RD14]. tested [Fra15]. Testing [AH15, GMM^{+}12, OSL^{+}15, OG16, SSM^{+}16, DBM^{+}15, KKC19, MAM^{+}15, PMOH13, SKM18, SCM19]. tests [CMG^{+}17, HOV^{+}15]. tetricus [HMHL19]. Teuthida [KSZ10]. Thailand [APV18]. Thalassiosira [LFH^{+}19, SY16, WPvBS15].

Thecosomata [RRM^{+}12]. their [AGMC14, BSB^{+}13, BT10, BMM^{+}19, CAL14, CJC^{+}15, DGPR11, DPE^{+}17, EPR^{+}14, EISJ12, GMN^{+}17, GG13, GSP12, GCL^{+}18, GTGD15, HHRW12, HSB18, Jon14, KEM^{+}17, KFL^{+}15, KGW^{+}12, Lap11, MB14, Mat11, MRP^{+}16, OSBG16, PSM^{+}10, RB10, RAB^{+}12, SGQR^{+}12, SPFM^{+}10, SGP^{+}15, WPH^{+}15, YDT14, dMBD11]. theme [Sei14]. Themisto [PCFR13]. theoretic [RJP^{+}17]. theoretical [GGP11, JdFBB^{+}10]. Theory [VF17, AH19, CLT^{+}14, Hol14, vdKFS^{+}16]. Theragra [BHD13, HFM13, IHH^{+}11, MBIH11, WPWH11, WMJ13]. there [ALWH19, BT15, BS14, GRH^{+}14, Wri14]. Thermal [EISJ12, PKHM12, AS14, CDAN^{+}14, HHM^{+}16, HMM12, ITT19, NF16, NRTK^{+}19, PMM^{+}14, ROBC11]. Thermographic [TBD^{+}18]. thermohaline [dSMGKP12]. thermoregulation [SSA^{+}18]. thiamine [KUM^{+}12]. thicket [dCVB14]. thickets [SMK15]. thickness [MSL^{+}11]. things [BCC^{+}19]. thinking [TBHK17]. thompsoni [WCK^{+}10]. Thoreau [Pac18]. thought [RDL^{+}17]. threat [HMN^{+}12]. threats [FBF^{+}17, Lav15]. three [ASI^{+}16a, Ben13, DKB14, FOC^{+}18, GMH^{+}14, GLC15, GGIC^{+}15, LGBA11, LHK18, LSI^{+}15, MMC19, MFS^{+}14, MBBC11, MSR14, OLR^{+}18, OJK^{+}19, OSBG16, SV10, SHS11, TUUC14, YDT14, ZHV^{+}16]. three-dimensional [ASI^{+}16a, FOC^{+}18, LSI^{+}15]. three-spined [OJK^{+}19]. thresholds [LFFL13]. throughout [BLB^{+}18, GSL^{+}11, LEB^{+}14, MWS^{+}17, YDT14, YJBL17]. thumb [LLL^{+}13]. Thunnus [BMO^{+}19, CAB^{+}10, FAR15, GFML10, LLL^{+}11, Mel16, MLLL11, MMSS18, SLH^{+}17, SWA^{+}16, WKK^{+}15, WLN^{+}13]. thynnus [FAR15, Mel16, MLLL11, SWA^{+}16]. Thysanoessa [HUPBL18, MPL^{+}14, MDS13]. tickler [DDE^{+}19]. tickler-chain [DDE^{+}19]. tidal [BS13, FWRM12, HdBR^{+}16, MDPH12, PGS^{+}19a, PGS^{+}19b, WCT^{+}16, WDE^{+}18]. tide [GQX^{+}19]. tier [BPMR17]. tiger [DPD^{+}15, KS17, TMW19]. tilefish [FFPL14]. tiles [HTRF17]. tilt [FCG^{+}16, FA10, KO12, KOC15]. tilt-angle [FA10, KO12]. Time [BYQ^{+}19, JMM^{+}15b, Abl16, Ben15, CRAM^{+}14, CAC^{+}18, CLMP16, CCA^{+}18, DCHN^{+}15, HLL^{+}12, KKMS14, KOK^{+}14, LHV^{+}16, LB11, LVSF17, Mac11, MG12, MNWB18, MN18, NC11, NBS^{+}10, OCS14, OAT^{+}12, PSM^{+}10, PDH^{+}14, PSF12, QR15, RB14, RPS11, SSF18b, SPE14, SH14b, SCM19, Tho11, THM15, WHL^{+}17, WESS18, JWS^{+}18]. time-evolution [RPS11]. time-lagged [PSF12]. Time-series [BYQ^{+}19, CCA^{+}18, DCHN^{+}15, HLL^{+}12, KOK^{+}14, MG12, NBS^{+}10, SPE14]. Time-varying [JMM^{+}15b, LB11, Tho11, THM15]. time/area [OCS14]. times [BFLÁ15, FB19, Ful11, LFMJ19a, LFMJ19b, MBT^{+}19, URE^{+}18]. Timing [WPH^{+}17, DPBM12, HAF^{+}16, LFMJ19a, LFMJ19b, WAH^{+}16]. Tinkering [ÖHSNV17]. tinro [KSZ10]. tissue [GS14]. tissues [MNG15]. Todarodes [KGN^{+}10, RYS11]. tolerance [SG16b]. tolerant [LEE17, SPRL18]. tonggol [GFML10]. tonsa [MAB12]. Tool [DOB^{+}17, CGRM^{+}13, GPS^{+}17, GJWS19, KCS10, KLKD11, NMO^{+}17, PDD^{+}10a, TMS^{+}15]. toolkit [LLG18]. tools [BKSB12, BFC^{+}16, DMS^{+}12, HHZ^{+}18, MNG15, PNE^{+}14]. tooth [FHH^{+}13a, FHH^{+}13b]. toothfish [GLP^{+}11, GTGD15, HLCC16, TTR^{+}19]. toothfish-depredating [TTR^{+}19]. Top [ÉSMGB15, CFHC10, LGL^{+}12, RFF^{+}14, RC15]. Top-down [ÉSMGB15, RFF^{+}14, RC15]. toppled [SLT^{+}14]. Total [PCD^{+}13, SMR^{+}10, Win15, ANU11, HTS^{+}10, SRGF15, WWI^{+}16]. tourist [VKN^{+}11]. tow [MST^{+}18, RGR^{+}11]. towed [OSP^{+}13, OI16]. towing [BHLS18, HMS^{+}13]. towns [TMG16]. Toxic [MMH17]. toxicity [MGVH^{+}10]. trace [BWR^{+}15]. tracers [WKK^{+}15]. Trachurus [DDR10, GMN^{+}17, GKV^{+}15, NPB^{+}15, STKT16, VCRPS13, VBA10]. track [GDD^{+}18, HG10]. Tracking [TNY^{+}13, WDG^{+}17, Bak14, BPST^{+}16, STB^{+}11]. tracks [OSC^{+}17, XMM12]. Trade [ECFL15, HDG^{+}17, KEM^{+}17, EBSW15, FR16, JLH12, LPD^{+}16b, MS10, MBP19, Pun17]. trade-off [MBP19]. Trade-offs [ECFL15, HDG^{+}17, KEM^{+}17, EBSW15, FR16, JLH12, LPD^{+}16b, MS10, Pun17]. tradeoff [Har13]. traditional [HO14, SMB^{+}16b]. tragedy [STSP17]. trained [AHR^{+}19, SSM^{+}18]. training [OSC^{+}17]. trait [KVA18, LSF^{+}15, OHD^{+}16, RDB^{+}18b]. trait-based [KVA18, RDB^{+}18b]. traits [AL10b, BPKH13, CAC^{+}18, GRNG^{+}10, HBB^{+}16, KJW^{+}18, KTMV16, LPH^{+}15, LGR^{+}19, NUO^{+}10, Ohm19, PG16, SMR12]. trajectories [SHW^{+}15]. transatlantic [FFH^{+}11, FHH^{+}13a, FHH^{+}13b]. transboundary [SFNdH16]. Transdisciplinary [CHB^{+}14, OSSL16]. transect [FC19, MS15, SMD^{+}16]. transfer [SSA^{+}18]. transferable [TPL11]. transform [ACT^{+}18]. transformation [WHMP15, WHP16]. Transgenerational [PDD^{+}10a, RPB16]. transhipment [ANU11]. transient [GKR14, GR15]. transition [BDHC19, GDD^{+}18, HHHE16, KS18b, TPL11]. transitions [HKB16]. translocation [CGG^{+}11, JQD^{+}17, TPBL17]. transmit [RD16]. transmitters [BDHC19]. transparency [YS15]. transparent [MNGA^{+}17]. transport [BMFSH12, DHNL12, JFJ^{+}17, JV12b, KRD^{+}18, MPL^{+}14, PBS14, PBJ^{+}14, PPHK19]. trap [BSB^{+}13, FDC13, GBB15, SLS18, SBCF16]. traps [BLB^{+}12, BM15, FDC13, MSR^{+}19]. travelling [OLW^{+}14]. trawl [APK11, BPH16, BSC14, BMSC16, CTH^{+}19a, CTH^{+}19b, DZC^{+}13, DDE^{+}19, DUM^{+}12, DMT10b, EMP11, FCD^{+}11, GD15, HKB16, HCR13, JdFBB^{+}10, JVPM12, KHF14, KJÓK16, KIP14, KHPI15, KVFK19, KFN15, LHS^{+}18, LGB^{+}18, MSM17, Mil10, MHMP^{+}18, NDF^{+}15, OKDJ18, OFB^{+}17, PDR19, PTvOR13, PHA^{+}10, QHG^{+}12, REFJ12, SHG^{+}18, SWvSR11, TWB^{+}19, THKP11, TPL11, UBvH11, URE^{+}18, WMLJ17, WWCO15, WWI^{+}16, dHFF^{+}16, vHCP15, vHCP17, vdRMC^{+}17]. trawl-survey [JVPM12]. trawled [SMZ^{+}14, SJUE11, SWS13, UTA^{+}16]. trawler [MTDP11]. trawlermen [BM12]. trawlers [CCG14, SJUE11]. Trawling [BMEBM^{+}16, ACA19, BUDM15, BMC^{+}17, DID^{+}16, EBH^{+}17, GIPS12, GIDP^{+}18, JPTC16, LNWS18, MKD^{+}18, MS14, PEV^{+}16, RBB^{+}16, REFJ12, SWBJ13, YFG^{+}17, vDHvKR15]. trawling-induced [LNWS18]. trawlnets [RK16]. trawls [EBB^{+}16a, EBB^{+}16b, HRB15a, HPS^{+}11, HWK^{+}15, MV10, MKHK18, SLV16, WSGD^{+}17, WPWH11]. treatise [Bro14]. treatment [Ast15b]. trecae [VBA10]. tree [CAOG15, LLSJ18, MS15, SMK15]. Trend [BCT^{+}10, HPDB19, TRH10, TNS13, ZMF12]. Trends [CKH19, DKC12, ALd12, Bak14, BA12, BSR^{+}10, Cha12, Che10, DCVC15, GRF^{+}12, GRR^{+}12, GT19, HLL^{+}12, HHM^{+}11, HMFW14, HHE^{+}16, IaF11, KOK^{+}14, LFFL13, LGR^{+}19, LHH^{+}13, LSCL18, MLC16, NGTT16, NBS^{+}10, OSBG16, PRKASR12, RKB^{+}12, SGQR^{+}12, SCY^{+}10, SBD^{+}15, TB17, VGWJ11]. trends-based [GRF^{+}12]. triacanthus [Ada17, JLL17]. trial [CEP^{+}18, MUEO17]. trials [BMC^{+}17, RBC^{+}15, UOB^{+}15]. tribute [Ban12]. Trichiurus [TKA^{+}16]. Trichodesmium [FGMC^{+}13]. tricuspidata [BBHC11]. trigger [GKR10]. triggerfish [Ger17]. triggering [ET15]. trip [Ric19]. tripartite [BSR^{+}10]. trips [ANU11]. Trisopterus [NWM^{+}12, NLB^{+}12]. Trøndelag [SMBH16]. Trophic [BCKA11, DJHO10, JvdM15, MALM17, OBLP^{+}19, PFPG12, RGC^{+}19, ASV^{+}19, BHS^{+}11b, CB12, CFHC10, CSTJ19, CKV^{+}16, DPD^{+}12, FRM^{+}16, FPSF11, GRP10, GGP11, JIB^{+}12, JIB^{+}14, LD17, LS10, MTP^{+}12, MRC^{+}18a, MRC^{+}18b, MSS^{+}19, PB11, PB14, RSV^{+}17, SAB^{+}16, SMZ^{+}14, SFW^{+}12, TUUC14, VRF^{+}16]. trophic-level-based [GGP11]. Trophodynamic [YDT14]. Trophodynamics [MPJ12]. Tropical [DKPM16, RTAT17, ASB^{+}19, BSHC17, BdSP11, BMSC16, BBM15, CCE17, COHdP19, FBL^{+}16, GFML10, HCM^{+}15, HGS^{+}19, JDM^{+}16, KCA^{+}14, LHJ15, MB14, MONS14, MKB^{+}17, MARD13, MW19, NF16, OBLP^{+}19, RRM^{+}12, RBC^{+}15, SZOL19, SFD^{+}16, TNFC16, WNW^{+}15, YJBL17]. troubled [Dil17]. Trough [SRHJ12]. trout [DLL12, GKC^{+}16, SRKV15, UPF^{+}10, VHF^{+}17, WKP^{+}17]. true [MLC16]. truth [AAP14a]. truthing [AAP14a, CSFS10]. truths [SH15]. trutta [DLL12, GKC^{+}16, UPF^{+}10, WKP^{+}17]. TS [MMG^{+}17]. tshawytscha [KBT14]. Tuamotu [GVA19]. Tuna [MBAGVG^{+}18, ZCH19, ALWH19, ABHT^{+}16, ACC^{+}12, AGG^{+}18, BWR^{+}15, BMS^{+}18, BMO^{+}19, FAR15, FL14, GFML10, KCA^{+}14, LLL^{+}11, LD17, LBM19, MALM17, MKB^{+}17, Mel16, MLC^{+}19, MLLL11, MBL^{+}17, MTO^{+}18, MMSS18, RBA^{+}19, RBC^{+}15, RCH^{+}19, SLH^{+}17, SWA^{+}16, WLN^{+}13, XCRP13]. tunas [FS14, PKR^{+}19]. tunicate [KHH^{+}17a]. tunicates [DL12]. Tunisia [DHKE16]. tunnelling [Coc17]. turbidity [OMTY14]. turbines [FWRM12]. Turbulence [KVSV14]. turbulent [Fra15]. turtle [KCP^{+}11]. turtles [GHB^{+}10, NDB^{+}16, dQCD^{+}10]. Twelve [Cam18]. Twentieth [Kam14]. twilight [KLA19]. twist [FDC13]. Two [Ben13, HCM^{+}15, NMF^{+}16, BHC^{+}15, BIKP^{+}17, CGK^{+}17, Che10, DBA^{+}19, DZ14, FR16, GRC^{+}14, GKC^{+}15, HTRF17, HGF^{+}18, IFU11, JMM19, KRKG16, LGR^{+}19, MSOF18, MPL^{+}14, MRC^{+}18a, MRC^{+}18b, MDS13, MJSB14, NPB^{+}15, NGTT16, OTO^{+}15, OAT^{+}12, PHM^{+}14, PR16, PMM^{+}14, PDT^{+}18, RRM^{+}12, SSM^{+}16, SJ18, SCD^{+}15, Str10, SPV^{+}16, TVJ^{+}14, TMP12, TSPL14]. two-frequency [KRKG16]. two-generation [OAT^{+}12]. two-sex [TSPL14]. two-stage [GRC^{+}14, IFU11]. type [GMN^{+}17]. types [ACCA19, CJC^{+}15, GT19, MMWC16, MFS^{+}14]. typical [CHMY15, DD13]. tyrannus [BMH^{+}16, HAH^{+}16].

U.S [LLF18]. UAVs [CBK18]. ubiquitously [MDV^{+}15]. ugly [BSR^{+}10, Mur10]. UK [ATK^{+}10, SRH^{+}14, SMS^{+}17]. Ulleung [KYK^{+}12]. Umberto [Can16]. umbrella [GLP^{+}11]. umbrella-and-stones [GLP^{+}11]. unaccounted [WSGD^{+}17]. unaffected [BTB^{+}17]. unanticipated [HHM^{+}19]. uncertain [FR16, HHBSM16, SCS^{+}11]. Uncertainties [PBC^{+}16, PBL11, LBM19]. Uncertainty [CIJ18, KCF^{+}17, AH15, Ben13, Cad13, DAP^{+}12, GCJL16, iIOKW13, JvdM15, KET^{+}17, KCMS19, KKH14, LMPP10, MPB11, PCRD18, PHK^{+}19, PSGY^{+}12, PAB^{+}14, RMKG11, RTS19, SR16, TID^{+}10, WWI^{+}16, ZCR15]. unchanged [KT18]. uncontrolled [SMS^{+}16]. undatum [HPS^{+}15, HCL^{+}18]. underlying [BHJ14, FR16, KH19, VGWJ11]. undermine [VF17]. undermining [Fri19, Sca18]. underpin [GWS17]. undersized [BSP15, vdRMC^{+}17]. understand [Ger17, SBR11]. Understanding [CD16, GVA19, HWK^{+}15, LHLK10, Mor18, VP19, BKSB12, GOPG10, GTGD15, HP12, HSGL18, HBG^{+}19a, HBG^{+}19b, NJFH19, Pun17, SRRZ16, CRAM^{+}14]. Underwater [DWFD13, HP17, OdJN^{+}12, QFM^{+}10, SMS^{+}16, SHS^{+}17, SSM^{+}18, SKH^{+}12, SMD^{+}16]. Unexpectedly [FWS^{+}13]. unicohort [LHV^{+}16]. unicolor [HD11]. unifying [TM15]. Union [BCRM^{+}18]. unique [FLCQ19, JWS^{+}18]. unit [BKSB12, EHHH14, EMP11, LPT10, ZCH19]. unite [JOF14]. United [BT15, GS11, GHD^{+}10, MTLG14, RSBC13]. units [CCC^{+}14, KCK14, KHC^{+}17, KRO^{+}16, LÁG^{+}19, MMSS18, ØHS10]. unknown [KOK^{+}14]. unknowns [PCRD18, SC16]. unmanned [CBK18, DLSJ^{+}19]. unobserved [HCR13, SWT^{+}19]. unravelling [CPSCCE^{+}17]. unsupervised [CSS^{+}15]. unusual [DPRG^{+}18]. unwanted [TPHdJ19]. upon [PRB^{+}15]. upper [BHS^{+}11b, BL15, RRCT^{+}17, XPC11]. upper-ocean [BL15, RRCT^{+}17]. ups [LÓGR11]. upstream [MHE^{+}16, TCD^{+}16, WABZ16]. uptake [EP19]. Upwelling [QCA^{+}18a, QCA^{+}18b, BLM^{+}17, BÁOMRV13, GKV^{+}15, JBSD11, LEB^{+}14, PDWH11, RFN^{+}19, ROBC11, RAH^{+}16, SCY^{+}10, VHH10, ZHL^{+}19]. urchin [CRC^{+}16, DDF^{+}12, FBFP^{+}13, OFVF14, SPRL18]. urchins [CGAD16]. urgent [MHH^{+}18]. Uruguayan [PDS15a, PDS15b]. USA [LEB^{+}14, SHL^{+}14, SG16a]. usage [GT19]. Use [CSS^{+}12, DMO10, MV10, PDR19, AFQ^{+}11, BMM^{+}19, Cor12, DMO^{+}15b, DLL12, DPRG^{+}18, DHCE13, ETG^{+}15, FB19, GSMA15, HLS^{+}15, KOK^{+}14, KTMV16, MVB^{+}18, MKB^{+}17, MRD19, MMDSP15, MNG15, MBL^{+}17, OND^{+}18, PVQS12, PGG^{+}17, RGRL11, SDSA15, SGP^{+}15, SBB19, SLS15, SBR11, SPM^{+}17, SWBJ13, TJMC11, ZHD10]. used [Bjö18, CR11, DBA^{+}19, FHD^{+}19, GGTW17, GHV^{+}18, HSV^{+}17, HKRM18, HJT12, MBLS15, MHM^{+}11, MRP^{+}16, TJR^{+}10, dJBLH15]. useful [Cad13, SBS^{+}10, Wri14]. uses [GGP11, SMR^{+}11]. Using [BBOE10, CAM^{+}15, CECL16, CFM^{+}14, CGRM^{+}13, DMZC17, FCB17, FWD15, HCR13, HBC^{+}11, LAD^{+}11, OCBJ10, PSF12, RGAS^{+}18, RC15, RBBC11, SSRT10, SS10a, SSB^{+}10a, SFW^{+}12, SHS15, SWvSR11, SSP12, SBR11, TMR^{+}16, vHCP17, ASI^{+}16a, AHM^{+}15, AHR^{+}19, AML^{+}19, ASV^{+}19, AFK15, BMGB19, BGS^{+}14, BPST^{+}16, BF17, BVC15, BBD^{+}13, BCOS11, CSS^{+}15, CPT^{+}10, CTH^{+}19a, CTH^{+}19b, CLS^{+}18, CGK^{+}17, CPB19a, CPB19b, CD14, CSD16, DMF13, DUM^{+}12, DTL^{+}11, DPD^{+}12, DWPS11, DRB10, Eid18, Elv15, FFF16, FFH^{+}11, FHH^{+}13a, FHH^{+}13b, GCS^{+}18, GTBT18, GT19, Ham15, HHR17, HG10, HKS17, HRF^{+}10, Hil19a, Hil19b, HSFiS12, HHQ11, HKS12, KJW^{+}18, KRKG16, KVFK19, LJH^{+}12, LLL^{+}13, LCM10, LAG^{+}16, LPD^{+}16a, LGRC14, LSI^{+}15, LPT10, Man16, Man17a, MV13, MYI12, MSR^{+}19, MPB11, MSR^{+}12, MBA^{+}19, MBAGVG^{+}18, MHHK13, OdJN^{+}12]. using [OSP^{+}13, OBLP^{+}19, OFB^{+}17, OSSL16, PE18, PLSD16, Pun17, QFM^{+}10, RSY11, ROBC11, RCS^{+}17, RK16, RCH^{+}19, SHK^{+}15, SG17, ŠBGT^{+}17, SPS^{+}10, SSMD^{+}13, SKH^{+}12, SFD15, SCJI10, SCM19, SI15, SR16, SFN^{+}19, TBG16, TMR^{+}17, THHH15, THHH18, TMS^{+}15, TRSM15, TD19, TB17, TKJB19, TVJ^{+}14, TSPL14, TLPS15, URV^{+}11, URE^{+}18, VPH18, VSP^{+}14, VKN^{+}11, WOW^{+}17, WDG^{+}17, WLL^{+}13, WRDF10, XMM12, YPN^{+}18, ZDD^{+}19, ZKD^{+}14, ZCH19, ZGT13, ZGTL13, vdHdGL16, LVM^{+}11, SBS^{+}10]. Utility [PCS^{+}11, SOK^{+}12, BTA^{+}18, NS17, SPFM^{+}10, SWT^{+}19, VCF^{+}12]. utilization [HSGL18, LVM^{+}11, MRC^{+}10, PST14, SLG^{+}15, VCH16]. utilize [HHM^{+}16]. utilizing [BVW^{+}18]. utopia [KGZH16]. UVR [DRVV^{+}17].

Vadimus [Bro12, Bro17a]. Valdez [SHW^{+}18]. validated [GBM^{+}19]. Validation [RWG^{+}18, BTA^{+}18, HRF^{+}10, JCSC10, MHK19, MDS13]. valorization [ECW^{+}17]. valuation [SSR^{+}16]. Value [Sei14, JLH12, MVK18, OAM^{+}10, PS13, RML^{+}16, SWB^{+}14, XS12, ZHAG17]. Variability [BMFSH12, FW10, HSK^{+}14, HHGtIWGoOH12, RGFT14, AASD12, APB^{+}10, ALd12, BMGB19, BHJ14, BDCvD^{+}10, BNP19, CWRB11, DVV^{+}11, EKR^{+}19, EISJ12, FWT^{+}14, GHF^{+}11, GMC17, GDBM^{+}17, GDJ11b, HHH12a, HPN^{+}17, HHT13, ITT19, JV12b, Kam14, Kan11, KBC^{+}12, KAT11, KKH14, KRO^{+}16, LHHK12, LSF^{+}15, LCTB^{+}15, LFMJ19a, LFMJ19b, LBK17, MDPH12, MSL^{+}11, MW19, MTO^{+}18, MDMC11, OFYS^{+}11, PBS14, PDH^{+}14, PMM^{+}14, PRO14, RSV^{+}17, Rod10, ROK^{+}18, RYS11, dÁMNGS^{+}11, TCS^{+}19, TCQ^{+}15, TSS^{+}13, UHB12, VAP12, VELT14, VHGTFR10, WÓG^{+}16, WMG^{+}17, WRC^{+}12, YLLG18, YYCC16, YCZY19]. Variable [KPD^{+}16, AS14, BPvHR10, HHBSM16, HMM12, NMF^{+}16, SGM^{+}17]. variables [CAOG15, PCB^{+}14]. variably [PSH13]. variance [ALBR14, HHQ11, EWH16]. Variation [HGRAR10, AGMC14, BHDB12, BDCvD^{+}10, BHS^{+}16, CBL19, CSC^{+}13, Der18, GNKS18, HPS^{+}15, JCSC10, JTE19, JKvdK^{+}15, KTLB16, KRD^{+}18, LPH^{+}15, LTGP18, LJB16, MM12, MVH^{+}11, MSS^{+}19, MJS^{+}15b, PSKM^{+}16, PWPW18, PDT^{+}18, QR15, RFF^{+}14, RAH^{+}16, RJ12, SBG10, SGQR^{+}12, SG10, SBS19a, SBS19b, SSP^{+}11, VHS^{+}10, ZRN18]. variations [AVGÓ12, CAC^{+}18, GTM12, HLL^{+}12, LLL^{+}11, MUW^{+}19, PJDN12, STKT16, WCT^{+}16, WPH^{+}15]. varies [WSGD^{+}17]. various [SDSA15]. varying [GFF^{+}19, JMM^{+}15b, LB11, MGR^{+}13, MSR14, Tho11, THM15, WVG^{+}13]. Vasco [Pin17]. vegetation [QFM^{+}10]. vehicle [SKH^{+}12]. vehicles [CBK18, DLSJ^{+}19]. Venerupis [LK17]. vent [BCL^{+}11, CBC^{+}11b, CBC^{+}11a]. vents [TNFC16]. verification [HCF^{+}12a, KRML11]. verified [MKR13, OOD13]. vermetid [FTM^{+}17]. vernacular [KOC^{+}16]. verreauxi [GFMO11]. versatile [DGC12]. versus [CBBM14, CD14]. vertebral [MIJ^{+}17]. Vertical [AJ12, HLS^{+}15, OSM18, PPHK19, SO16, SJ15, AMQRC10, CD14, GJWS19, KJÓK16, KBH^{+}18, KRD^{+}18, MJS^{+}16, SV10, TKA^{+}16, UHB12, VBW^{+}18, VHS^{+}10, WFS^{+}11]. vertically [KBH^{+}18]. verticillata [CSBHS^{+}11]. Vessel [DWW^{+}10, MHHK13, BTTV^{+}17, BVC15, DW11, DWFD13, DLSJ^{+}19, FBCD^{+}17, FB11, JL12b, MCOS17, OSC^{+}17, Peñ19a, SHS^{+}18, GL11, GML13, LJH^{+}12, LSJ10, MKD^{+}18]. vessel-based [BVC15]. vessels [DMO^{+}15b, DW10, DH13, OCWG16, PQR10, THKP11, XCRP13]. Vesterålen [BV19b, BV19a, MO13]. vexillum [KHH^{+}17a]. via [JQD^{+}17, CNBK11]. viability [BNE^{+}17, Pre19]. Video [HMP^{+}15, HBF14, OdJN^{+}12, SHS^{+}17, SMMK11]. videos [SMS^{+}16, SSM^{+}18]. view [Coc17, FOC^{+}18, HFSV^{+}15, Koe11, KBH^{+}18, MNG15, SKKM15, TRSM15, VDK15]. Viewpoints [STF^{+}17]. VII [PGN^{+}11]. VIIIc [FCPJ10]. villosus [CMD17, GHM^{+}16, HGRAR10, OCR14, PGG^{+}12, SSF18b]. virens [AJ12, CLV^{+}14, GHV^{+}18, íHHJ^{+}13, OS14, SHK^{+}15, SO16, SCD^{+}15]. virtual [FFPL14, OYI17]. virus [DTA^{+}13, MWS^{+}17]. virus-infected [MWS^{+}17]. Vision [URE^{+}18, MBAGVG^{+}18, SKM18]. visual [SMD^{+}16, TBC15]. Visualizing [PPM13]. visually [MKR13]. vital [RFT^{+}16]. vitality [KVFK19]. vitulina [DTA^{+}13]. VME [FDC^{+}16]. VMS [GL11, GML13, LJH^{+}12, LSJ10, MKD^{+}18, ANU11, SHS^{+}18, SJUE11]. vocalizations [GZS^{+}19]. voices [OdS19]. volcanic [SFD^{+}16, TNFC16]. volume [JLL17, TPHdJ19, TK18]. volumes [VPO17]. voluntary [EP19]. Voyage [RS10]. VPA [DDR10]. vs [DDE^{+}19, ÉSMGB15, Fra15, GPP17, GOS^{+}13, HMFW14, MVH^{+}11, MBL^{+}17, Pow14, PMOH13, SLT^{+}14, SMB^{+}16b]. vulgaris [ESR^{+}10, GCG^{+}10, GOPG10, GRG^{+}10, HRF^{+}10, SPE14]. Vulnerability [JPTC16, AHK^{+}18, CJC^{+}15, GLH14, HFGT18, PEN^{+}19b, PGN^{+}11, SSF^{+}12, SSTB15]. vulnerable [AGH^{+}11, FMM^{+}12, PMBM18].

WA [TBC^{+}14]. wahoo [ZGT13, ZGTL13]. Waiting [SH15]. Waldbusser [CSJ16b]. Wales [EHT17]. Walleye [PDAC^{+}16, BHD13, DW11, GSC10, HFM13, IHH^{+}11, KLCC18, KCL18, KHPI15, LCSI18, MBIH11, SHZ^{+}16, WPWH11, WMJ13, WWI^{+}16]. warm [BNP19, FBD^{+}19, PDAC^{+}16]. warm-core [BNP19]. warm-water [FBD^{+}19]. warmer [VSB^{+}16]. warming [ASI^{+}16b, COL^{+}13, CRC^{+}16, FRF14, FSN^{+}11, GM17, HBG^{+}16, JPPY11, KPD^{+}16, LAD^{+}17, LNH^{+}15, MGL^{+}15, NDM^{+}16, PDWH11, PT19, RCF^{+}17, SiTiM^{+}11, SSTB15, TIS^{+}19, VGB^{+}17]. warning [DCPTN14]. waste [BJH^{+}16]. Water [DPBM12, JV12b, Aks15, AL10b, BHJ14, BMFSH12, BBL^{+}15, BKT12, BDHC19, CNBK11, CBBM14, DHNL12, Dil17, DI11, DMS18, FBD^{+}19, GDBM^{+}17, HKRM18, HHM^{+}16, JDH^{+}14, LHHK12, LRJ18, LAD^{+}11, LSCL18, MFS^{+}14, MHSG19, MMAS11, QR15, RC15, SRHJ12, Tow14, VMG11, WOW^{+}17, WPGW12]. water-column [LRJ18]. waterborne [KPBB15]. waters [AJM19, AASD12, ALMFFBP19, AJ12, AVGÓ12, BVS^{+}16, CRHM19, DRVV^{+}17, EBH^{+}17, GMN^{+}17, GSMA15, GFML10, KLN^{+}10, MGVH^{+}10, MUW^{+}19, MSL^{+}11, MJSB14, NC12, OMTY14, OEG^{+}16, OAM^{+}10, PSH^{+}11, PHB^{+}16, RLQ^{+}15, RSB^{+}15, SHH^{+}17, SRB^{+}14, SO10, SOL^{+}15, TJMC11, UvHS^{+}14, UPF^{+}10, VAP12, WPH^{+}17, WHAA^{+}18, WÓG^{+}16, WMG^{+}17, WHNS14, Win15, ZMKC14]. wave [ACT^{+}18, FWRM12, HD18, SS17, SRW13]. waves [HHR17]. way [BT15, SP14]. weakening [ITT19]. weakfish [FBL^{+}16]. weakly [BHL^{+}15]. wearing [DSN16]. weather [LTT^{+}18, SDW19, VG15]. web [AML^{+}19, ASV^{+}19, BHB^{+}19, BSF^{+}19, BHM13, LBM19, LLL^{+}13, PCFR13, RFF^{+}14, SAB^{+}16, TLR^{+}17]. webs [RDB^{+}18b]. weight [BMR10a, DDR10, MUW^{+}19, OSJ^{+}16, RC15]. weight-at-length [OSJ^{+}16]. weight-related [MUW^{+}19]. weight-structured [RC15]. weighted [WRDF10]. weights [BCBI13]. weissflogii [LFH^{+}19, SY16, WPvBS15]. Welfare [BCC^{+}19, Dig19]. well [KVFK19]. West [FFN^{+}16, HABV19, JW18, KHF14, LHJ12, RH17, AHM^{+}15, BLJ^{+}17, GNDC11, JCS16, KBT14, LDD^{+}10, MWJ11, RRCT^{+}17, TPL11, WKW^{+}10, YFG^{+}17, BHS^{+}16, FFN^{+}13, HGRAR10, HHA^{+}11, JBE14, KGW^{+}12, MMDSP15, SLKS10, STD11, TSWS15, TDHC18, YJBL17, dGGW^{+}11, dPLF^{+}19]. Westerberg [SC16]. Western [Lav15, TMG16, ALWH19, ASI^{+}16b, CGRM^{+}13, CST^{+}19, DMJ^{+}14, DDR10, FL14, FOT^{+}17, GOS^{+}13, HHH12a, HLS^{+}15, Hüs11, HHH12b, HHE^{+}16, IS15, KKZ^{+}12, MLC16, PHM^{+}14, PKHG14, RHHH^{+}12, SHAM10, SSZH12, TID^{+}10, VF17, VPSV^{+}10, ZLC^{+}17, ZOQS10, dL14, dLCF^{+}15, dQCD^{+}10, ABHT^{+}16, BGS^{+}14, BS13, BCSG13, DHZA14, HMHL19, JRG^{+}16, KVFK19, MWJ11, May14, MOG^{+}14a, MOG^{+}14b, MHK19, MGBvD14, NST^{+}14, PCdL15, PHB^{+}16, TRPH16, WNM10, dL14, dLCF^{+}15]. wetland [KP13]. WGFTFB [RGRL11]. Whale [SOL^{+}15, GZS^{+}19, KKLM13, Moo19, OSL^{+}15, PDS15b, PMH^{+}13, PH17, TMS^{+}15, TGG^{+}15, TTR^{+}19]. whales [BT10, Elv15, GRKF16, GLP^{+}11, GTGD15, LHJ12, Moo14, Moo19, TGDG15, TOA^{+}16, WVG^{+}13]. whelk [HPS^{+}15]. Where [CRAM^{+}14, MVB^{+}18, RB14, Bro14, BCC^{+}19, GRH^{+}14]. which [Cla18d, GHB^{+}14, LB13, MBP19]. whiffiagonis [MCAM14]. while [RD16]. white [HSSB14, HLS^{+}15, MMRG18, VBW^{+}18]. whitebrow [AEC11]. whitefish [LHV^{+}16, LBB15, OHD^{+}16]. whiting [BGM^{+}14, DJHO10, OOD13, PGOM11, dCWMH13]. Who [JKS^{+}18, EPR^{+}14]. whole [SSA^{+}18, vPGFT13]. whom [EPR^{+}14]. whose [Abl16]. Wickström [SC16]. wide [CHMY15, FHK14, KPJ^{+}15, LÁG^{+}19, WPH^{+}15]. Wideband [JLL17]. widely [DBA^{+}19]. Wider [BM16, JVRT10]. Widespread [KDF^{+}16]. width [SMD^{+}16]. wild [FJSJ15, GDJ11a, GOPG10, HHDB14, HFH10, KIA^{+}18, KDF^{+}16, OS14, ŠBGT^{+}17, UBV^{+}11, UPF^{+}10]. Will [Hol10, BNB19, CRAM^{+}14, GBB16]. William [Ban12]. wind [BUDM15, ET15, RCF^{+}18, WAH^{+}16]. wind-driven [ET15]. windows [PKHM12]. winner [MMH17]. winners [Ful11]. Winter [KOT^{+}17, SRC11, AGMC14, AS14, BHD13, BHMR14, BGM^{+}19, DWW^{+}10, Loh11, LK15, PH12a, YCZY19]. within [CKD^{+}17, DGL^{+}17, Fau11, Fle15, GFF^{+}17, GHWD13, GFSC17, HDG^{+}17, HKS12, KSJ14, MGPC^{+}14, OE16, QOH^{+}13, Sca18, SLS18, SRDC^{+}14, Tur19a, Tur19b, THR16, WSW^{+}19, dCWMH13]. without [BHC^{+}15, FB11, GKC^{+}16, MUEO17, RCS^{+}17]. wolffish [BFC^{+}16, PSDG12]. woods [Pac18]. Word [Aga18a, Bas18a, Cla18b, Day18b, Hil18c, Obu18b, SG18c]. work [BS14]. worked [DGL^{+}17]. workflow [DUM^{+}12]. working [LÓGR11, RGRL11]. World [LKG^{+}19, MS12, BCT^{+}10, DGPR11, FR16, FP12, HBG^{+}19a, HBG^{+}19b, HHBSM16, IOK17, KS18a, Pun17, SOB^{+}11, She15]. worldwide [SLG^{+}15]. wrasse [DMF13, HSD^{+}16, HSV^{+}17, JQD^{+}17, OHLK19a, OHLK19b, SDBB15, SJ16]. wrong [WHMP15].

Xiphias [CSC^{+}13, GMD^{+}11, SSA^{+}18]. Xiphiidae [GMD^{+}11]. Xiphopenaeus [BMSC16, KdS14].

year [AFP12, BUDM15, BDCvD^{+}10, BMC^{+}13, CCD^{+}19, CCA^{+}18, DD10, Ger17, GHM^{+}16, KOK^{+}14, OYI17, Ósk18, PJDN12, RMJF14, SBG10, SG16a, SLF17, TLF16]. year-class [GHM^{+}16, PJDN12, RMJF14, TLF16]. year-round [SG16a]. years [ACM^{+}16, ALWH19, BGM^{+}19, Bro14, CFHC10, EPK^{+}18, FCPJ10, LHP^{+}19, PDAC^{+}16, SDAM^{+}18, SWA^{+}16, SSH^{+}14, dLCF^{+}15]. yellow [WABZ16, ZRN18, HZZ^{+}15, YLLG18]. yellow-phase [WABZ16]. yellowfin [BBOE10, LLL^{+}11, MMSS18, WLN^{+}13]. yellowtail [BWP10, GLC15, MWP^{+}16, RBBC11]. Yield [PR16, BSB^{+}19, FR16, FO13, Har13, KKZ^{+}12, LB13, LM10, RCS^{+}17, RDL^{+}17, RKJ14, TJD17, THC16, UVD^{+}17, vGA18b]. yield-per-recruit [THC16]. yielding [dMADLP^{+}16]. yields [ACM^{+}16, NCH17, SOB^{+}11]. yolk [HPN^{+}17, LCSI18, MRC^{+}10]. yolk-sac [HPN^{+}17, LCSI18]. Young [AJM19, Pet19, SBG10, SLF17]. young-of-the-year [SBG10, SLF17]. youth [VOM11]. Yucatán [VBW^{+}18]. Yukon [ME11].

Zapteryx [CPSCCE^{+}17]. Zealand [Alf10, CPLH16, ECFL15, EAB^{+}17, HB19a, MSC14, SG16b, THR16]. zeehaani [DHS19]. zero [Tur19a, Tur19b]. Zoantharia [CSBHS^{+}11]. zoanthid [CSBHS^{+}11]. Zonation [DMS^{+}12, GFSC17]. zone [BUDM15, BVW^{+}18, DJW^{+}15, DCHMHQ12, GDD^{+}18, GFSC17, KLA19, OKDJ18, WPH^{+}17, YST^{+}10, dJBLH15]. zones [ITT19, MJS^{+}15b]. zoobenthos [RLP^{+}13]. Zooming [APOG11]. Zooplankton [DG17, KBC^{+}12, AFLvDH15, BW14, BYQ^{+}19, CBBM14, DHH^{+}14, DHNL12, EPK^{+}18, FW10, GS12, GDBM^{+}17, GMM^{+}12, GHD^{+}10, KGW^{+}12, LCM10, LK15, MOG^{+}14a, MOG^{+}14b, MLC16, PB14, PT19, RLW^{+}15, SEDO19]. zooxanthellae [BIKP^{+}17].

References

Al-Abdulrazzak:2014:GTG

[AAP14a] Dalal Al-Abdulrazzak and Daniel Pauly. Ground-truthing the ground-truth: reply to Garibaldi et al.’s comment on “Managing fisheries from space: Google Earth improves estimates of distant fish catches”. ICES Journal of Marine Science, 71(7):1927–1931, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/7/1927/2804455. See [AAP14b, GGT^{+}14].

Al-Abdulrazzak:2014:MFS

[AAP14b] Dalal Al-Abdulrazzak and Daniel Pauly. Managing fisheries from space: Google Earth improves estimates of distant fish catches. ICES Journal of Marine Science, 71(3): 450–454, March 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/71/3/450/638656. See com-

ment [GGT^{+}14] and reply [AAP14a].

Albretsen:2012:CVS

[AASD12] Jon Albretsen, Jan Aure, Roald Sætre, and Didrik S. Danielssen. Climatic variability in the Skagerrak and coastal waters of Norway. ICES Journal of Marine Science, 69(5):758–763, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/69/5/758/646533.

Aksnes:2014:JHI

[AB14] Dag W. Aksnes and Howard I. Browman. Johan Hjort’s impact on fisheries science: a bibliometric analysis. ICES Journal of Marine Science, 71(8):2012–2016, October 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/8/2012/2804446.

Aksnes:2016:OGR

[AB16] Dag W. Aksnes and Howard I. Browman. An overview of global research effort in fisheries science. ICES Journal of Marine Science, 73(4):1004–1011, March 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/4/1004/2458729.

Andersen:2016:ABS

[ABF^{+}16] Ken H. Andersen, Julia L. Blanchard, Elizabeth A. Fulton, Henrik Gislason, Nis Sand Jacobsen, and Tobias van Kooten. Assumptions behind size-based ecosystem models are realistic. ICES Journal of Marine Science, 73(6): 1651–1655, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/73/6/1651/2458714. See re-

ply [FWP^{+}16b].

Alvarez-Berastegui:2016:PSE

[ABHT^{+}16] Diego Alvarez-Berastegui, Manuel Hidalgo, María Pilar Tugores, Patricia Reglero, Alberto Aparicio-González, Lorenzo Ciannelli, Mélanie Juza, Baptiste Mourre, Ananda Pascual, José Luís López-Jurado, Alberto García, José María Rodríguez, Joaquín Tintoré, and Francisco Alemany. Pelagic seascape ecology for operational fisheries oceanography: modelling and predicting spawning distribution of Atlantic bluefin tuna in Western Mediterranean. ICES Journal of Marine Science, 73(7):1851–1862, July 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/7/1851/2458726.

Andre:2019:RCC

[ABJ^{+}19] Carl André, Julia M. I. Barth, Patrik Jonsson, Sissel Jentoft, Halvor Knutsen, and Henrik Svedäng. Response to comments by Cardinale et al. on “Local cod (Gadus morhua) revealed by egg surveys and population genetic analysis after longstanding depletion on the Swedish Skagerrak coast” by Svedäng et al. (2019). ICES Journal of Marine Science, 76(4):1212–1213, July 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/4/1212/5528145.

Able:2016:NHA

[Abl16] Kenneth W. Able. Natural history: an approach whose time has come, passed, and needs to be resurrected. ICES Journal of Marine Science, 73(9):2150–2155, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/9/2150/2198256.

Anderson:2012:AGM

[ABRL12] Sean C. Anderson, Trevor A. Branch, Daniel Ricard, and Heike K. Lotze. Assessing global marine fishery status with a revised dynamic catch-based method and stock-assessment reference points. ICES Journal of Marine Science, 69(8):1491–1500, September 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/8/1491/702072.

Adasme:2019:ISM

[ACA19] Luis M. Adasme, Cristian M. Canales, and Nicolás A. Adasme. Incidental seabird mortality and discarded catches from trawling off far southern Chile (39–57^{o}S). ICES Journal of Marine Science, 76(4):848–858, July 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/4/848/5307404.

Amande:2012:PBE

[ACC^{+}12] Monin J. Amandè, Emmanuel Chassot, Pierre Chavance, Hilario Murua, Alicia Delgado de Molina, and Nicolas Bez. Precision in bycatch estimates: the case of tuna purse-seine fisheries in the Indian Ocean. ICES Journal of Marine Science, 69(8):1501–1510, September 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/8/1501/702198.

Alos:2019:MAE

[ACCA19] Josep Alós, Andrea Campos-Candela, and Robert Arlinghaus. A modelling approach to evaluate the impact of fish spatial behavioural types on fisheries stock assessment. ICES Journal of Marine Science, 76(2):489–500, March 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/2/489/5211192.

Aalto:2016:QYD

[ACM^{+}16] Emilius Aalto, Fabrizio Capoccioni, Juan Terradez Mas, Marcello Schiavina, Chiara Leone, Giulio De Leo, and Eleonora Ciccotti. Quantifying 60 years of declining European eel (Anguilla anguilla L., 1758) fishery yields in Mediterranean coastal lagoons. ICES Journal of Marine Science, 73(1):101–110, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
1/101/2458643.

Asche:2015:EIT

[ACS15] Frank Asche, Yanyou Chen, and Martin D. Smith. Economic incentives to target species and fish size: prices and fine-scale product attributes in Norwegian fisheries. ICES Journal of Marine Science, 72(3):733–740, March 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/3/733/700626.

Addison:2018:NWM

[ACT^{+}18] P. F. E. Addison, D. J. Collins, R. Trebilco, S. Howe, N. Bax, P. Hedge, G. Jones, P. Miloslavich, C. Roelfsema, M. Sams, R. D. Stuart-Smith, P. Scanes, P. von Baumgarten, and A. McQuatters-Gollop. A new wave of marine evidence-based management: emerging challenges and solutions to transform monitoring, evaluating, and reporting. ICES Journal of Marine Science, 75(3):941–952, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/3/941/4739744.

Adams:2017:ASD

[Ada17] Charles F. Adams. Age-specific differences in the seasonal spatial distribution of butterfish (Peprilus triacanthus). ICES Journal of Marine Science, 74(1):170–179, January 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/1/170/2669550.

Aranburu:2016:GER

[ADB16] Aizkorri Aranburu, Estibaliz Díaz, and Cédric Briand. Glass eel recruitment and exploitation in a South European estuary (Oria, Bay of Biscay). ICES Journal of Marine Science, 73(1):111–121, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/1/111/2458722.

Ainsley:2011:AGM

[AEC11] Shaara M. Ainsley, David A. Ebert, and Gregor M. Cailliet. Age, growth, and maturity of the whitebrow skate, Bathyraja minispinosa, from the eastern Bering Sea. ICES Journal of Marine Science, 68(7):1426–1434, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/7/1426/657341.

Anders:2017:SSB

[AFH^{+}17] Neil Anders, Anders Fernö, Odd-Børre Humborstad, Svein Løkkeborg, and Anne Christine Utne-Palm. Species specific behaviour and catchability of gadoid fish to floated and bottom set pots. ICES Journal of Marine Science, 74(3): 769–779, March 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/74/3/769/2656161.

Audzijonyte:2015:IFB

[AFK15] Asta Audzijonyte, Elizabeth A. Fulton, and Anna Kuparinen. The impacts of fish body size changes on stock recovery: a case study using an Australian marine ecosystem model. ICES Journal of Marine Science, 72(3):782–792, March 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/3/782/693470.

Fukuwaka:2011:RDC

[aFKA11] Masa aki Fukuwaka, Toshiki Kaga, and Tomonori Azumaya. Regional differences in climate factors controlling chum and pink salmon abundance. ICES Journal of Marine Science, 68(6):1131–1137, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/68/
6/1131/702939.

Alvarez-Fernandez:2015:EZF

[AFLvDH15] S. Alvarez-Fernandez, P. Licandro, C. J. G. van Damme, and M. Hufnagl. Effect of zooplankton on fish larval abundance and distribution: a long-term study on North Sea herring (Clupea harengus). ICES Journal of Marine Science, 72(9):2569–2577, November 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/9/2569/2458363.

Andersson:2012:EEA

[AFP12] Jan Andersson, Ann-Britt Florin, and Erik Petersson. Escapement of eel (Anguilla anguilla) in coastal areas in Sweden over a 50-year period. ICES Journal of Marine Science, 69(6):991–999, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/69/6/991/622665.

Andonegi:2011:PUG

[AFQ^{+}11] Eider Andonegi, Jose Antonio Fernandes, Iñaki Quincoces, Xabier Irigoien, Andrés Uriarte, Aritz Pérez, Daniel Howell, and Gunnar Stefánsson. The potential use of a gadget model to predict stock responses to climate change in combination with Bayesian networks: the case of Bay of Biscay anchovy. ICES Journal of Marine Science, 68(6): 1257–1269, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/6/1257/715402.

Agardy:2018:AFW

[Aga18a] Tundi Agardy. Agardy’s final word. ICES Journal of Marine Science, 75(3):1187, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/75/
3/1187/4098830.

Agardy:2018:CD

[Aga18b] Tundi Agardy. Counterpoint to day. ICES Journal of Marine Science, 75(3):1191, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/75/
3/1191/4098833.

Agardy:2018:JAA

[Aga18c] Tundi Agardy. Justified ambivalence about MPA effectiveness. ICES Journal of Marine Science, 75(3):1183–1185, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/3/1183/4098825.

Arregui:2018:MGD

[AGG^{+}18] Igor Arregui, Benjamin Galuardi, Nicolas Goñi, Chi Hin Lam, Igaratza Fraile, Josu Santiago, Molly Lutcavage, and Haritz Arrizabalaga. Movements and geographic distribution of juvenile bluefin tuna in the Northeast Atlantic, described through internal and satellite archival tags. ICES Journal of Marine Science, 75(5):1560–1572, September 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/5/1560/5004410.

Auster:2011:DDV

[AGH^{+}11] Peter J. Auster, Kristina Gjerde, Eric Heupel, Les Watling, Anthony Grehan, and Alex David Rogers. Definition and detection of vulnerable marine ecosystems on the high seas: problems with the “move-on” rule. ICES Journal of Marine Science, 68(2):254–264, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/2/254/615543.

Able:2014:TVW

[AGMC14] K. W. Able, T. M. Grothues, J. M. Morson, and K. E. Coleman. Temporal variation in winter flounder recruitment at the southern margin of their range: is the decline due to increasing temperatures? ICES Journal of Marine Science, 71(8):2186–2197, October 2014.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/8/2186/2804399.

Agnew:2014:MED

[AGSPH14] D. J. Agnew, N. L. Gutiérrez, A. Stern-Pirlot, and D. D. Hoggarth. The MSC experience: developing an operational certification standard and a market incentive to improve fishery sustainability. ICES Journal of Marine Science, 71(2):216–225, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/71/2/216/778266.

Arnold:2015:TRD

[AH15] Linsey M. Arnold and Selina S. Heppell. Testing the robustness of data-poor assessment methods to uncertainty in catch and biology: a retrospective approach. ICES Journal of Marine Science, 72(1):243–250, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/1/243/824522.

Ailloud:2019:GTA

[AH19] Lisa E. Ailloud and John M. Hoenig. A general theory of age–length keys: combining the forward and inverse keys to estimate age composition from incomplete data. ICES Journal of Marine Science, 76(6):1515–1523, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/6/1515/5480389.

Azmi:2015:HSN

[AHC15] Fauziah Azmi, Chad L. Hewitt, and Marnie L. Campbell. A hub and spoke network model to analyse the secondary dispersal of introduced marine species in Indonesia. ICES Journal of Marine Science, 72(3):1069–1077, March 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/3/1069/688875.

Aps:2018:MEV

[AHK^{+}18] Robert Aps, Kristjan Herkül, Jonne Kotta, Roland Cormier, Kirsi Kostamo, Leena Laamanen, Juho Lappalainen, Külli Lokko, Anneliis Peterson, and Riku Varjopuro. Marine environmental vulnerability and cumulative risk profiles to support ecosystem-based adaptive maritime spatial planning. ICES Journal of Marine Science, 75(7):2488–2500, December 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/75/
7/2488/5070430.

Alexander:2015:IRD

[AHM^{+}15] Karen A. Alexander, Johanna J. Heymans, Shona Magill, Maciej T. Tomczak, Steven J. Holmes, and Thomas A. Wilding. Investigating the recent decline in gadoid stocks in the west of Scotland shelf ecosystem using a foodweb model. ICES Journal of Marine Science, 72(2):436–449, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/2/436/2801443.

Allken:2019:FSI

[AHR^{+}19] Vaneeda Allken, Nils Olav Handegard, Shale Rosen, Tiffanie Schreyeck, Thomas Mahiout, and Ketil Malde. Fish species identification using a convolutional neural network trained on synthetic data. ICES Journal of Marine Science, 76(1):342–349, January 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/1/342/5137114.

Alemany:2013:ELS

[AIA13] Daniela Alemany, Oscar O. Iribarne, and Eduardo M. Acha. Effects of a large-scale and offshore marine protected area on the demersal fish assemblage in the Southwest Atlantic. ICES Journal of Marine Science, 70(1):123–134, January 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/1/123/661945.

Albaina:2015:MPI

[AIA^{+}15] Aitor Albaina, Xabier Irigoien, Unai Aldalur, Guillermo Boyra, María Santos, and Andone Estonba. Macrozooplankton predation impact on anchovy (Engraulis encrasicolus) eggs mortality at the Bay of Biscay shelf break spawning centre. ICES Journal of Marine Science, 72(5): 1370–1379, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/72/5/1370/758123.

Armannsson:2012:VMS

[AJ12] Hlynur Armannsson and Sigurdur Tór Jónsson. Vertical migrations of saithe (Pollachius virens) in Icelandic waters as observed with data storage tags. ICES Journal of Marine Science, 69(8):1372–1381, September 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/8/1372/703486.

Albert:2019:YMR

[AJM19] Ole Thomas Albert, Claudia Junge, and Marlén Knutsen Myrlund. Young mums are rebuilding the spurdog stock (Squalus acanthias L.) in Norwegian waters. ICES Journal of Marine Science, 76(7):2193–2204, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/7/2193/5543885.

Aksnes:2015:SCD

[Aks15] D. L. Aksnes. Sverdrup critical depth and the role of water clarity in Norwegian coastal water. ICES Journal of Marine Science, 72(6):2041–2050, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/6/2041/918847.

Aps:2010:RDB

[AL10a] Robert Aps and Hans Lassen. Recovery of depleted Baltic Sea fish stocks: a review. ICES Journal of Marine Science, 67(9):1856–1860, December 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/67/
9/1856/620692.

Arkhipkin:2010:CLH

[AL10b] Alexander I. Arkhipkin and Vladimir V. Laptikhovsky. Convergence in life-history traits in migratory deep-water squid and fish. ICES Journal of Marine Science, 67 (7):1444–1451, October 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/67/7/1444/665269.

Astarloa:2019:IMI

[ALB^{+}19] Amaia Astarloa, Maite Louzao, Guillermo Boyra, Udane Martinez, Anna Rubio, Xabier Irigoien, Francis K. C. Hui, and Guillem Chust. Identifying main interactions in marine predator–prey networks of the Bay of Biscay. ICES Journal of Marine Science, 76(7):2247–2259, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/7/2247/5541029.

Archambault:2014:DDC

[ALBR14] B. Archambault, O. Le Pape, N. Bousquet, and E. Rivot. Density-dependence can be revealed by modelling the variance in the stock–recruitment process: an application to flatfish. ICES Journal of Marine Science, 71(8):2127–2140, October 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/71/8/2127/749918.

Alvarez:2012:ACC

[ALd12] I. Alvarez, M. N. Lorenzo, and M. deCastro. Analysis of chlorophyll a concentration along the Galician coast: seasonal variability and trends. ICES Journal of Marine Science, 69(5):728–738, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/69/5/728/650677.

Alfaro:2010:EMR

[Alf10] Andrea C. Alfaro. Effects of mangrove removal on benthic communities and sediment characteristics at Mangawhai Harbour, northern New Zealand. ICES Journal of Marine Science, 67(6):1087–1104, September 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/6/1087/738149.

Arechavala-Lopez:2019:COS

[ALMFFBP19] P. Arechavala-Lopez, M. Minguito-Frutos, G. Follana-Berná, and M. Palmer. Common octopus settled in human-altered Mediterranean coastal waters: from individual home range to population dynamics. ICES Journal of Marine Science, 76(2):585–597, March 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/2/585/4917675.

Adams:2018:EDF

[ALWH18] Grant D. Adams, Robert T. Leaf, Wei Wu, and Frank J. Hernandez. Environmentally driven fluctuations in condition factor of adult Gulf menhaden (Brevoortia patronus) in the northern Gulf of Mexico. ICES Journal of Marine Science, 75(4):1269–1279, July 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/75/
4/1269/4831067.

Ailloud:2019:EAC

[ALWH19] Lisa E. Ailloud, Matthew V. Lauretta, John F. Walter III, and John M. Hoenig. Estimating age composition for multiple years when there are gaps in the ageing data: the case of western Atlantic bluefin tuna. ICES Journal of Marine Science, 76(6):1690–1701, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/6/1690/5480393.

Anderson:2019:QCF

[AML^{+}19] Thomas R. Anderson, Adrian P. Martin, Richard S. Lampitt, Clive N. Trueman, Stephanie A. Henson, and Daniel J. Mayor. Quantifying carbon fluxes from primary production to mesopelagic fish using a simple food web model. ICES Journal of Marine Science, 76(3):690–701, May 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/3/690/4791420.

Abascal:2010:HVM

[AMQRC10] Francisco J. Abascal, Jaime Mejuto, Manuel Quintans, and Ana Ramos-Cartelle. Horizontal and vertical movements of swordfish in the Southeast Pacific. ICES Journal of Marine Science, 67(3):466–474, April 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/3/466/731722.

Aldrin:2012:IMD

[AMS^{+}12] Magne Aldrin, Bjørnar Mortensen, Geir Storvik, Kjell Nedreaas, Asgeir Aglen, and Sondre Aanes. Improving management decisions by predicting fish bycatch in the Barents Sea shrimp fishery. ICES Journal of Marine Science, 69(1): 64–74, January 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/69/1/64/671190.

Anderson:2015:LCF

[And15] Emory D. Anderson. Lessons from a career in fisheries science. ICES Journal of Marine Science, 72(8):2169– 2179, September 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/72/8/2169/2458760.

Anonymous:2010:LRa

[Ano10a] Anonymous. List of referees. ICES Journal of Marine Science, 67(1):186–189, January 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/67/
1/186/596690.

Anonymous:2010:LRb

[Ano10b] Anonymous. List of referees. ICES Journal of Marine Science, 67(7):1524, October 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/67/7/1524/664457.

Anonymous:2010:LRc

[Ano10c] Anonymous. List of referees. ICES Journal of Marine Science, 67(8):1822–1823, November 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/67/
8/1822/608554.

Anonymous:2010:LRd

[Ano10d] Anonymous. List of referees. ICES Journal of Marine Science, 67(9):2051, December 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/67/
9/2051/623648.

Anonymous:2010:RN

[Ano10e] Anonymous. Retraction notice. ICES Journal of Marine Science, 67(9):2052, December 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/67/
9/2052/620412.

Anonymous:2011:D

[Ano11a] Anonymous. Dedication. ICES Journal of Marine Science, 68(6):983, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/6/983/715094.

Anonymous:2011:Ea

[Ano11b] Anonymous. Erratum. ICES Journal of Marine Science, 68(4):801, March 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/4/801/652068.

Anonymous:2011:Eb

[Ano11c] Anonymous. Erratum. ICES Journal of Marine Science, 68(7):1592, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/7/1592/658267.

Anonymous:2011:LRa

[Ano11d] Anonymous. List of referees. ICES Journal of Marine Science, 68(2):425, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/68/2/425/617415.

Anonymous:2011:LRb

[Ano11e] Anonymous. List of referees. ICES Journal of Marine Science, 68(4):800, March 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/68/4/800/652065.

Anonymous:2011:LRc

[Ano11f] Anonymous. List of referees. ICES Journal of Marine Science, 68(6):1373, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/68/6/1373/716742.

Anonymous:2011:LRd

[Ano11g] Anonymous. List of referees. ICES Journal of Marine Science, 68(8):1825, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/68/
8/1825/757661.

Anonymous:2011:LRe

[Ano11h] Anonymous. List of referees. ICES Journal of Marine Science, 68(10):2277–2278, November 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/68/
10/2277/617351.

Anonymous:2012:LRa

[Ano12a] Anonymous. List of referees. ICES Journal of Marine Science, 69(3):492, May 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/69/3/492/601240.

Anonymous:2012:LRb

[Ano12b] Anonymous. List of referees. ICES Journal of Marine Science, 69(5):923, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/69/5/923/654158.

Anonymous:2012:LRc

[Ano12c] Anonymous. List of referees. ICES Journal of Marine Science, 69(7):1328, September 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/69/
7/1328/752079.

Anonymous:2012:LRd

[Ano12d] Anonymous. List of referees. ICES Journal of Marine Science, 69(9):1699, November 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/69/
9/1699/640747.

Anonymous:2012:SDB

[Ano12e] Anonymous. Special dedication to Bernard Megrey. ICES Journal of Marine Science, 69(7):1119, September 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/7/1119/752976.

Anonymous:2014:CIP

[Ano14a] Anonymous. Complete issue PDF 71-8. ICES Journal of Marine Science, 71(8):1989–2369, October 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/8/1989/763846.

Anonymous:2014:Ca

[Ano14b] Anonymous. Corrigendum. ICES Journal of Marine Science, 71(2):427, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/71/2/427/791257.

Anonymous:2014:Cb

[Ano14c] Anonymous. Corrigendum. ICES Journal of Marine Science, 71(3):739, March 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/71/3/739/639939.

Anonymous:2014:E

[Ano14d] Anonymous. Erratum. ICES Journal of Marine Science, 71(9):2643, November 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/71/9/2643/595608.

Anonymous:2015:CIP

[Ano15] Anonymous. Complete issue pdf 72 suppl 1. ICES Journal of Marine Science, 72(S1):S1–S260, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/72/
suppl_1/i1/622867.

Anonymous:2016:CIPa

[Ano16a] Anonymous. Complete issue PDF 73-3. ICES Journal of Marine Science, 73(3):529–990, February 2016.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/3/529/2459153.

Anonymous:2016:CIPb

[Ano16b] Anonymous. Complete issue pdf 73 suppl 1. ICES Journal of Marine Science, 73(suppl 1):S1–S138, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/73/suppl_1/i1/2573998.

Anonymous:2017:CIP

[Ano17] Anonymous. Complete issue pdf 74-4. ICES Journal of Marine Science, 74(4):889–1229, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
4/889/3852374.

Anonymous:2018:Ca

[Ano18a] Anonymous. Corrigendum. ICES Journal of Marine Science, 75(3):1143, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/75/3/1143/3192368.

Anonymous:2018:Cb

[Ano18b] Anonymous. Corrigendum. ICES Journal of Marine Science, 75(3):1144, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/75/3/1144/3094706.

Anonymous:2018:Cc

[Ano18c] Anonymous. Corrigendum. ICES Journal of Marine Science, 75(3):1145, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/75/3/1145/3737534.

Anonymous:2018:Cd

[Ano18d] Anonymous. Corrigendum. ICES Journal of Marine Science, 75(3):1146, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/75/3/1146/3820952.

Anonymous:2018:Ce

[Ano18e] Anonymous. Corrigendum. ICES Journal of Marine Science, 75(3):1147, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/75/3/1147/3820953.

Anonymous:2018:Cf

[Ano18f] Anonymous. Corrigendum. ICES Journal of Marine Science, 75(3):1148, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/75/3/1148/3926534.

Anonymous:2018:Cg

[Ano18g] Anonymous. Corrigendum. ICES Journal of Marine Science, 75(3):1149, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/75/3/1149/4774612.

Anonymous:2018:Ch

[Ano18h] Anonymous. Corrigendum. ICES Journal of Marine Science, 75(3):1150, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/75/3/1150/4939312.

Anonymous:2018:Ci

[Ano18i] Anonymous. Corrigendum. ICES Journal of Marine Science, 75(3):1151, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/75/3/1151/5003289.

Anonymous:2018:Cj

[Ano18j] Anonymous. Corrigendum. ICES Journal of Marine Science, 75(3):1152, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/75/3/1152/3737838.

Anonymous:2018:Ea

[Ano18k] Anonymous. Erratum. ICES Journal of Marine Science, 75(3):1153, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/3/1153/4596636.

Anonymous:2018:Eb

[Ano18l] Anonymous. Erratum. ICES Journal of Marine Science, 75(3):1154, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/3/1154/4823421.

Anonymous:2018:Ec

[Ano18m] Anonymous. Erratum. ICES Journal of Marine Science, 75(3):1155, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/3/1155/4915975.

Anonymous:2018:Ed

[Ano18n] Anonymous. Erratum. ICES Journal of Marine Science, 75(6):2286–2287, November 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/75/
6/2286/5116195.

Anonymous:2018:LR

[Ano18o] Anonymous. List of reviewers 2017. ICES Journal of Marine Science, 75(1):451–454, January 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/1/451/4828025.

Anonymous:2018:MPA

[Ano18p] Anonymous. Marine protected areas: all articles. ICES Journal of Marine Science, 75(3):903–1201, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/3/903/4098845.

Anonymous:2019:PHE

[Ano19] Anonymous. Pelagic habitat: exploring the concept of good environmental status. ICES Journal of Marine Science, 76(6):1939, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
6/1939/5525265.

Albertsen:2018:CSS

[ANT18] Christoffer Moesgaard Albertsen, Anders Nielsen, and Uffe Høgsbro Thygesen. Connecting single-stock assessment models through correlated survival. ICES Journal of Marine Science, 75(1):235–244, January 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/1/235/4057589.

Aanes:2011:ETR

[ANU11] Sondre Aanes, Kjell Nedreaas, and Sigbjørn Ulvatn. Estimation of total retained catch based on frequency of fishing trips, inspections at sea, transhipment, and VMS data. ICES Journal of Marine Science, 68(8):1598–1605, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/68/8/1598/759811.

Alos:2010:IGP

[APB^{+}10] Josep Alós, Miquel Palmer, Salvador Balle, Antoni Maria Grau, and Beatriz Morales-Nin. Individual growth pattern and variability in Serranus scriba: a Bayesian analysis. ICES Journal of Marine Science, 67(3):502–512, April 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/3/502/733216.

Azmi:2015:AMB

[APHC15] Fauziah Azmi, Carmen Primo, Chad L. Hewitt, and Marnie L. Campbell. Assessing marine biosecurity risks when data are limited: bioregion pathway and species-based exposure analyses. ICES Journal of Marine Science, 72(3):1078–1091, March 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/72/
3/1078/702518.

Attwood:2011:BSA

[APK11] Colin G. Attwood, Samantha L. Petersen, and Sven E. Kerwath. Bycatch in South Africa’s inshore trawl fishery as determined from observer records. ICES Journal of Marine Science, 68(10):2163–2174, November 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/10/2163/616873.

Amoroso:2011:ZMM

[APOG11] Ricardo O. Amoroso, Ana M. Parma, J. M. (Lobo) Orensanz, and Domingo A. Gagliardini. Zooming the macroscope: medium-resolution remote sensing as a framework for the assessment of a small-scale fishery. ICES Journal of Marine Science, 68(4):696–706, March 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/4/696/644808.

Aylesworth:2018:EIF

[APV18] Lindsay Aylesworth, Ratanavaree Phoonsawat, and Amanda C. J. Vincent. Effects of indiscriminate fisheries on a group of small data-poor species in Thailand. ICES Journal of Marine Science, 75(2):642–652, March 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/75/
2/642/4557572.

Andersen:2010:DIC

[AR10] Ken H. Andersen and Jake C. Rice. Direct and indirect community effects of rebuilding plans. ICES Journal of Marine Science, 67(9):1980–1988, December 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/9/1980/618840.

Arnold:2011:FRH

[Arn11] Geoff Arnold. F. R. Harden Jones. ICES Journal of Marine Science, 68(10):2005–2006, November 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/10/2005/616461.

Anderson:2014:EVW

[AS14] Deena A. Anderson and Frederick S. Scharf. The effect of variable winter severity on size-dependent overwinter mortality caused by acute thermal stress in juvenile red drum (Sciaenops ocellatus). ICES Journal of Marine Science, 71(4):1010–1021, May 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/71/
4/1010/664098.

Ainsworth:2011:PIC

[ASB^{+}11] C. H. Ainsworth, J. F. Samhouri, D. S. Busch, W. W. L. Cheung, J. Dunne, and T. A. Okey. Potential impacts of climate change on Northeast Pacific marine foodwebs and fisheries. ICES Journal of Marine Science, 68(6):1217– 1229, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/68/6/1217/707044.

Ault:2019:LBR

[ASB^{+}19] Jerald S. Ault, Steven G. Smith, James A. Bohnsack, Jiangang Luo, Molly H. Stevens, Gerard T. DiNardo, Matthew W. Johnson, and David R. Bryan. Lengthbased risk analysis for assessing sustainability of datalimited tropical reef fisheries. ICES Journal of Marine Science, 76(1):165–180, January 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
1/165/5107807.

Aarflot:2018:CCS

[ASDSM18] Johanna Myrseth Aarflot, Hein Rune Skjoldal, Padmini Dalpadado, and Mette Skern-Mauritzen. Contribution of Calanus species to the mesozooplankton biomass in the Barents Sea. ICES Journal of Marine Science, 75(7):2342– 2354, December 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/7/2342/4748804.

Alabia:2016:ESH

[ASI^{+}16a] Irene D. Alabia, Sei-Ichi Saitoh, Hiromichi Igarashi, Yoichi Ishikawa, Norihisa Usui, Masafumi Kamachi, Toshiyuki Awaji, and Masaki Seito. Ensemble squid habitat model using three-dimensional ocean data. ICES Journal of Marine Science, 73(7):1863–1874, July 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
7/1863/2458778.

Alabia:2016:FPI

[ASI^{+}16b] Irene D. Alabia, Sei-Ichi Saitoh, Hiromichi Igarashi, Yoichi Ishikawa, Norihisa Usui, Masafumi Kamachi, Toshiyuki Awaji, and Masaki Seito. Future projected impacts of ocean warming to potential squid habitat in western and central North Pacific. ICES Journal of Marine Science, 73(5): 1343–1356, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/73/5/1343/2240659.

Ailloud:2015:PAC

[AST^{+}15a] Lisa E. Ailloud, Matthew W. Smith, Amy Y. Then, Kristen L. Omori, Gina M. Ralph, and John M. Hoenig. Properties of age compositions and mortality estimates derived from cohort slicing of length data. ICES Journal of Marine Science, 72(1):44–53, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/72/
1/44/825624.

Astles:2015:LRF

[Ast15b] K. L. Astles. Linking risk factors to risk treatment in ecological risk assessment of marine biodiversity. ICES Journal of Marine Science, 72(3):1116–1132, March 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/3/1116/699498.

Arroyo:2019:TCG

[ASV^{+}19] Nina-Larissa Arroyo, Georges Safi, Pauline Vouriot, Lucía López-López, Nathalie Niquil, Francois Le Loc’h, Tarek Hattab, and Izaskun Preciado. Towards coherent GES assessments at sub-regional level: signs of fisheries expansion processes in the Bay of Biscay using an OSPAR food web indicator, the mean trophic level. ICES Journal of Marine Science, 76(6):1543–1553, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/6/1543/5369193.

Abernethy:2010:FDU

[ATK^{+}10] Kirsten E. Abernethy, Paul Trebilcock, Bereket Kebede, Edward H. Allison, and Nicholas K. Dulvy. Fuelling the decline in UK fishing communities? ICES Journal of Marine Science, 67(5):1076–1085, July 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/67/
5/1076/608158.

Alleway:2016:IHD

[ATLC16] Heidi K. Alleway, Ruth H. Thurstan, Peter R. Lauer, and Sean D. Connell. Incorporating historical data into aquaculture planning. ICES Journal of Marine Science, 73(5): 1427–1436, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/73/5/1427/2296358.

Andersen:2012:STC

[AUEC12] Bo Sølgaard Andersen, Clara Ulrich, Ole Ritzau Eigaard, and Anne-Sofie Christensen. Short-term choice behaviour in a mixed fishery: investigating métier selection in the Danish gillnet fishery. ICES Journal of Marine Science, 69(1):131–143, January 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/69/1/131/673193.

Albert:2015:MNA

[AV15] Ole Thomas Albert and Tone Vollen. A major nursery area around the Svalbard archipelago provides recruits for the stocks in both Greenland halibut management areas in the Northeast Atlantic. ICES Journal of Marine Science, 72(3): 872–879, March 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/72/3/872/694942.

Astthorsson:2012:CRV

[AVGÓ12] Olafur S. Astthorsson, Hédinn Valdimarsson, Asta Gudmundsdottir, and Gudmundur J. Óskarsson. Climaterelated variations in the occurrence and distribution of mackerel (Scomber scombrus) in Icelandic waters. ICES Journal of Marine Science, 69(7):1289–1297, September 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/7/1289/745283.

Andersen:2013:HSH

[AWS^{+}13] Julie M. Andersen, Yolanda F. Wiersma, Garry B. Stenson, Mike O. Hammill, Aqqalu Rosing-Asvid, and Mette Skern-Maurizen. Habitat selection by hooded seals (Cystophora cristata) in the Northwest Atlantic Ocean. ICES Journal of Marine Science, 70(1):173–185, January 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/1/173/660807.

Barratt:2010:AOS

[BA10] Iain M. Barratt and A. Louise Allcock. Ageing octopods from stylets: development of a technique for permanent preparations. ICES Journal of Marine Science, 67(7):1452– 1457, October 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/67/7/1452/662949.

Brown:2012:CTS

[BA12] Zachary W. Brown and Kevin R. Arrigo. Contrasting trends in sea ice and primary production in the Bering Sea and Arctic Ocean. ICES Journal of Marine Science, 69(7):1180–1193, September 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/69/
7/1180/751549.

Boschetti:2017:MMC

[BA17] Fabio Boschetti and Matthew Andreotta. Mental models, communication, and engagement in marine projects. ICES Journal of Marine Science, 74(7):2034–2039, September 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/7/2034/2907793.

Briseno-Avena:2018:DGE

[BAFRJ18] Christian Briseño-Avena, Peter J. S. Franks, Paul L. D. Roberts, and Jules S. Jaffe. A diverse group of echogenic particles observed with a broadband, high frequency echosounder. ICES Journal of Marine Science, 75(2):471– 482, March 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/2/471/4161436.

Bakun:2014:AOS

[Bak14] Andrew Bakun. Active opportunist species as potential diagnostic markers for comparative tracking of complex marine ecosystem responses to global trends. ICES Journal of Marine Science, 71(8):2281–2292, October 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/8/2281/752158.

Bannister:2012:RJW

[Ban12] Colin Bannister. In remembrance: John William Ramster, 1937–2011. A personal tribute. ICES Journal of Marine Science, 69(4):694–696, May 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/69/4/694/635811.

Bode:2013:SBG

[BÁOMRV13] Antonio Bode, Maria Teresa Álvarez-Ossorio, Ana Miranda, and Manuel Ruiz-Villarreal. Shifts between gelatinous and crustacean plankton in a coastal upwelling region. ICES Journal of Marine Science, 70(5):934–942, September 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/5/934/643530.

Barange:2019:AMC

[Bar19] Manuel Barange. Avoiding misinterpretation of climate change projections of fish catches. ICES Journal of Marine Science, 76(6):1390–1392, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/6/1390/5475453.

Basurto:2018:BFW

[Bas18a] Xavier Basurto. Basurto’s final word. ICES Journal of Marine Science, 75(3):1197, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/75/
3/1197/4098832.

Basurto:2018:CO

[Bas18b] Xavier Basurto. Counterpoint to obura. ICES Journal of Marine Science, 75(3):1200, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/75/
3/1200/4098839.

Basurto:2018:LME

[Bas18c] Xavier Basurto. Linking MPA effectiveness to the future of local rural fishing societies. ICES Journal of Marine Science, 75(3):1193–1194, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/75/
3/1193/4098824.

Brosset:2019:LEA

[BBB^{+}19] Pablo Brosset, Hugo Bourdages, Marjolaine Blais, Michael Scarratt, and Stéphane Plourde. Local environment affecting northern shrimp recruitment: a comparative study of Gulf of St. Lawrence stocks. ICES Journal of Marine Science, 76(4):974–986, July 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
4/974/5237773.

Butterworth:2010:PFM

[BBD^{+}10] Douglas S. Butterworth, Nokome Bentley, José A. A. De Oliveira, Gregory P. Donovan, Laurence T. Kell, Ana M. Parma, André E. Punt, Keith J. Sainsbury, Anthony D. M. Smith, and T. Kevin Stokes. Purported flaws in management strategy evaluation: basic problems or misinterpretations? ICES Journal of Marine Science, 67(3):567–574, April 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/67/3/567/734358.

Bucas:2013:EMB

[BBD^{+}13] M. Bučas, U. Bergström, A-L. Downie, G. Sundblad, M. Gullström, M. von Numers, A. Šiaulys, and M. Lindegarth. Empirical modelling of benthic species distribution, abundance, and diversity in the Baltic Sea: evaluating the scope for predictive mapping using different modelling approaches. ICES Journal of Marine Science, 70(6):1233– 1243, September 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/70/6/1233/634285.

Butcher:2010:SLM

[BBH^{+}10] Paul A. Butcher, Matt K. Broadhurst, Karina C. Hall, Brian R. Cullis, and Robert G. Nicoll. Scale loss and mortality in angled-and-released eastern sea garfish (Hyporhamphus australis). ICES Journal of Marine Science, 67 (3):522–529, April 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/67/3/522/732685.

Butcher:2011:PRS

[BBHC11] Paul A. Butcher, Matt K. Broadhurst, Karina C. Hall, and Steven J. Cooke. Post-release survival and physiology of angled luderick (Girella tricuspidata) after confinement in keeper nets in an Australian estuary. ICES Journal of Marine Science, 68(3):572–579, March 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/3/572/657876.

Beyan:2015:NSS

[BBL^{+}15] Cigdem Beyan, Bastian J. Boom, Jolanda M. P. Liefhebber, Kwang-Tsao Shao, and Robert B. Fisher. Natural swimming speed of Dascyllus reticulatus increases with water temperature. ICES Journal of Marine Science, 72(8):2506– 2511, September 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/72/8/2506/2458916.

Butler:2015:MCT

[BBM15] Mark Butler IV, Rodney Bertelsen, and Alison Macdiarmid. Mate choice in temperate and tropical spiny lobsters with contrasting reproductive systems. ICES Journal of Marine Science, 72(S1):S101–S114, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/72/
suppl_1/i101/615695.

Butcher:2010:UBA

[BBOE10] Paul A. Butcher, Matt K. Broadhurst, Beverley A. Orchard, and Megan T. Ellis. Using biotelemetry to assess the mortality and behaviour of yellowfin bream (Acanthopagrus australis) released with ingested hooks. ICES Journal of Marine Science, 67(6):1175–1184, September 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/6/1175/736983.

Bell:2018:CSS

[BCB^{+}18] Richard J. Bell, Jeremy S. Collie, Trevor A. Branch, Michael J. Fogarty, Coilin Minto, and Daniel Ricard. Changes in the size structure of marine fish communities. ICES Journal of Marine Science, 75(1):102–112, January 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/1/102/4055905.

Bachiller:2013:SDS

[BCBI13] Eneko Bachiller, Unai Cotano, Guillermo Boyra, and Xabier Irigoien. Spatial distribution of the stomach weights of juvenile anchovy (Engraulis encrasicolus L.) in the Bay of Biscay. ICES Journal of Marine Science, 70(2):362–378, March 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/70/2/362/795222.

Barausse:2014:RFE

[BCC^{+}14] Alberto Barausse, Vittoria Correale, Aleksia Curkovic, Licia Finotto, Emilio Riginella, Eleonora Visentin, and Carlotta Mazzoldi. The role of fisheries and the environment in driving the decline of elasmobranchs in the northern Adriatic Sea. ICES Journal of Marine Science, 71(7):1593– 1603, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/71/7/1593/663758.

Browman:2019:WAA

[BCC^{+}19] Howard I. Browman, Steven J. Cooke, Ian G. Cowx, Stuart W. G. Derbyshire, Alexander Kasumyan, Brian Key, James D. Rose, Alexander Schwab, Anne Berit Skiftesvik, E. Don Stevens, Craig A. Watson, and Robert Arlinghaus. Welfare of aquatic animals: where things are, where they are going, and what it means for research, aquaculture, recreational angling, and commercial fishing. ICES Journal of Marine Science, 76(1):82–92, January 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/1/82/5037898.

Benoit:2015:GML

[BCK^{+}15] Hugues P. Benoît, Connor W. Capizzano, Ryan J. Knotek, David B. Rudders, James A. Sulikowski, Micah J. Dean, William Hoffman, Douglas R. Zemeckis, and John W. Mandelman. A generalized model for longitudinal shortand long-term mortality data for commercial fishery discards and recreational fishery catch-and-releases. ICES Journal of Marine Science, 72(6):1834–1847, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/6/1834/921232.

Borrell:2011:TEE

[BCKA11] Asunción Borrell, Luis Cardona, Ramanathan P. Kumarran, and Alejandro Aguilar. Trophic ecology of elasmobranchs caught off Gujarat, India, as inferred from stable isotopes. ICES Journal of Marine Science, 68(3):547–554, March 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/68/3/547/656056.

Bettencourt:2011:DSL

[BCL^{+}11] Raul Bettencourt, Valentina Costa, Mário Laranjo, Domitília Rosa, Luís Pires, Ana Colaço, Humberto Lopes, and Ricardo Serrão Santos. Out of the deep sea into a land-based aquarium environment: investigating physiological adaptations in the hydrothermal vent mussel Bathymodiolus azoricus. ICES Journal of Marine Science, 68(2): 357–364, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/2/357/617195.

Borges:2016:DBB

[BCN16] Lisa Borges, Luis Cocas, and Kåre Nolde Nielsen. Discard ban and balanced harvest: a contradiction? ICES Journal of Marine Science, 73(6):1632–1639, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/6/1632/2459122.

Bui:2011:SAC

[BCOS11] Alice O. V. Bui, Martin Castonguay, Patrick Ouellet, and Jean-Marie Sévigny. Searching for Atlantic cod (Gadus morhua) spawning sites in the northwest Gulf of St Lawrence (Canada) using molecular techniques. ICES Journal of Marine Science, 68(5):911–918, May 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/5/911/648117.

Ballesteros:2018:DSM

[BCRM^{+}18] Marta Ballesteros, Rosa Chapela, Paulina Ramírez-Monsalve, Jesper Raakjaer, Troels J. Hegland, Kåre N. Nielsen, Unn Laksá, and Poul Degnbol. Do not shoot the messenger: ICES advice for an ecosystem approach to fisheries management in the European union. ICES Journal of Marine Science, 75(2):519–530, March 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/75/
2/519/4344803.

Biais:2017:RMP

[BCS^{+}17] Gérard Biais, Yann Coupeau, Bernard Séret, Beatriz Calmettes, Rémy Lopez, Stuart Hetherington, and David Righton. Return migration patterns of porbeagle shark (Lamna nasus) in the Northeast Atlantic: implications for stock range and structure. ICES Journal of Marine Science, 74(5):1268–1276, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
5/1268/2883378.

Brechon:2013:DRG

[BCSG13] Amanda L. Bréchon, Stephen H. Coombs, David W. Sims, and Andrew M. Griffiths. Development of a rapid genetic technique for the identification of clupeid larvae in the Western English Channel and investigation of mislabelling in processed fish products. ICES Journal of Marine Science, 70(2):399–407, March 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/70/2/399/795543.

Blanchard:2010:TAI

[BCT^{+}10] Julia L. Blanchard, Marta Coll, Verena M. Trenkel, Rémi Vergnon, Dawit Yemane, Didier Jouffre, Jason S. Link, and Yunne-Jai Shin. Trend analysis of indicators: a comparison of recent changes in the status of marine ecosystems around the world. ICES Journal of Marine Science, 67(4):732– 744, May 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/67/4/732/681988.

Bierman:2010:BYV

[BDCvD^{+}10] Stijn M. Bierman, Mark Dickey-Collas, Cindy J. G. van Damme, Harriët M. J. van Overzee, M. G. Pennock-Vos, Silja V. Tribuhl, and Lotte A. W. Clausen. Between-year variability in the mixing of North Sea herring spawning components leads to pronounced variation in the composition of the catch. ICES Journal of Marine Science, 67(5): 885–896, July 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/67/5/885/609315.

Budge:2014:ECE

[BDF^{+}14] Suzanne M. Budge, Emmanuel Devred, Marie-Hélène Forget, Venetia Stuart, M. Kurtis Trzcinski, Shubha Sathyendranath, and Trevor Platt. Estimating concentrations of essential omega-3 fatty acids in the ocean: supply and demand. ICES Journal of Marine Science, 71(7):1885– 1893, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/71/7/1885/668208.

Brunsdon:2019:TRS

[BDHC19] Eric Blake Brunsdon, Jason Daniels, Alex Hanke, and Jonathan Carr. Tag retention and survival of Atlantic salmon (Salmo salar) smolts surgically implanted with dummy acoustic transmitters during the transition from fresh to salt water. ICES Journal of Marine Science, 76(7):2471–2480, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
7/2471/5532215.

Blythe:2011:NSP

[BdSP11] Jonathan N. Blythe, José C. B. da Silva, and Jesús Pineda. Nearshore, seasonally persistent fronts in sea surface temperature on Red Sea tropical reefs. ICES Journal of Marine Science, 68(9):1827–1832, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/9/1827/666439.

Bassett:2018:BEM

[BDW18] Christopher Bassett, Alex De Robertis, and Christopher D. Wilson. Broadband echosounder measurements of the frequency response of fishes and euphausiids in the Gulf of Alaska. ICES Journal of Marine Science, 75(3):1131–1142, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/3/1131/4626990.

Benoit:2013:TDA

[Ben13] Hugues P. Benoît. Two decades of annual landed and discarded catches of three southern Gulf of St Lawrence skate species estimated under multiple sources of uncertainty. ICES Journal of Marine Science, 70(3):554–563, April 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/70/3/554/914859.

Bentley:2015:DTP

[Ben15] Nokome Bentley. Data and time poverty in fisheries estimation: potential approaches and solutions. ICES Journal of Marine Science, 72(1):186–193, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/1/186/822177.

Bergseth:2018:CC

[Ber18a] Brock Bergseth. Counterpoint to claudet. ICES Journal of Marine Science, 75(3):1175–1176, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/75/
3/1175/4098853.

Bergseth:2018:EMP

[Ber18b] Brock J. Bergseth. Effective marine protected areas require a sea change in compliance management. ICES Journal of Marine Science, 75(3):1178–1180, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/3/1178/4098828.

Bell:2017:SLA

[BF17] Susan S. Bell and Bradley T. Furman. Seascapes are landscapes after all; comment on Manderson (2016): Seascapes are not landscapes: an analysis performed using Bernhard Riemann’s rules. ICES Journal of Marine Science, 73:1831–1838. ICES Journal of Marine Science, 74(8):2276–2279, September 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/

8/2276/3872164. See [Man16, Man17a].

Burt:2011:BPR

[BFB^{+}11] John A. Burt, David A. Feary, Andrew G. Bauman, Paolo Usseglio, Georgenes H. Cavalcante, and Peter F. Sale. Biogeographic patterns of reef fish community structure in the northeastern Arabian Peninsula. ICES Journal of Marine Science, 68(9):1875–1883, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/9/1875/671133.

Bianucci:2016:OBM

[BFC^{+}16] Laura Bianucci, Katja Fennel, Denis Chabot, Nancy Shackell, and Diane Lavoie. Ocean biogeochemical models as management tools: a case study for Atlantic wolffish and declining oxygen. ICES Journal of Marine Science, 73(2): 263–274, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/73/2/263/2614459.

Briones-Fourzan:2015:ABS

[BFDGLÁ15] Patricia Briones-Fourzán, Roberto Domínguez-Gallegos, and Enrique Lozano-Álvarez. Aggressive behaviour of spotted spiny lobsters (Panulirus guttatus) in different social contexts: the influence of sex, size, and missing limbs. ICES Journal of Marine Science, 72(S1):S155–S163, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/72/suppl_1/i155/615217.

Bullimore:2013:CCB

[BFH13] Ross D. Bullimore, Nicola L. Foster, and Kerry L. Howell. Coral-characterized benthic assemblages of the deep Northeast Atlantic: defining “Coral Gardens” to support future habitat mapping efforts. ICES Journal of Marine Science, 70(3):511–522, April 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/70/3/511/914657.

Breece:2018:SDD

[BFH^{+}18] Matthew W. Breece, Dewayne A. Fox, Danielle E. Haulsee, Isaac I. Wirgin, and Matthew J. Oliver. Satellite driven distribution models of endangered Atlantic sturgeon occurrence in the mid-Atlantic Bight. ICES Journal of Marine Science, 75(2):562–571, March 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/75/
2/562/4222666.

Briones-Fourzan:2015:LOI

[BFLÁ15] Patricia Briones-Fourzán and Enrique Lozano-Álvarez. Lobsters: ocean icons in changing times. ICES Journal of Marine Science, 72(S1):S1–S6, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/72/
suppl_1/i1/622676.

Bramick:2016:SEM

[BFS16] Uwe Brämick, Erik Fladung, and Janek Simon. Stocking is essential to meet the silver eel escapement target in a river system with currently low natural recruitment. ICES Journal of Marine Science, 73(1):91–100, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/1/91/2458715.

Busch:2018:LBP

[BG18] Malte Busch and Stefan Garthe. Looking at the bigger picture: the importance of considering annual cycles in impact assessments illustrated in a migratory seabird species. ICES Journal of Marine Science, 75(2):690–700, March 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/2/690/4104491.

Braun:2019:AET

[BGA^{+}19] Camrin D. Braun, Peter Gaube, Pedro Afonso, Jorge Fontes, Gregory B. Skomal, and Simon R. Thorrold. Assimilating electronic tagging, oceanographic modelling, and fisheries data to estimate movements and connectivity of swordfish in the North Atlantic. ICES Journal of Marine Science, 76(7):2305–2317, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/7/2305/5523039.

Bacon:2011:ODS

[BGM^{+}11] Philip J. Bacon, William S. C. Gurney, Eddie McKenzie, Bryce Whyte, Ronald Campbell, Robert Laughton, Gordon Smith, and Julian MacLean. Objective determination of the sea age of Atlantic salmon from the sizes and dates of capture of individual fish. ICES Journal of Marine Science, 68(1):130–143, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/68/1/130/628967.

Bastrikin:2014:SLT

[BGM^{+}14] Dorota K. Bastrikin, Alejandro Gallego, Colin P. Millar, Imants G. Priede, and Emma G. Jones. Settlement length and temporal settlement patterns of juvenile cod (Gadus morhua), haddock (Melanogrammus aeglefinus), and whiting (Merlangius merlangus) in a northern North Sea coastal nursery area. ICES Journal of Marine Science, 71(8):2101– 2113, October 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/71/8/2101/753583.

Bernard:2019:CIA

[BGM^{+}19] Kim S. Bernard, Lacey A. Gunther, Sean H. Mahaffey, Katelyn M. Qualls, Monisha Sugla, Benjamin T. Saenz, Anthony M. Cossio, Jennifer Walsh, and Christian S. Reiss. The contribution of ice algae to the winter energy budget of juvenile Antarctic krill in years with contrasting sea ice conditions. ICES Journal of Marine Science, 76(1):206–216, January 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/1/206/5138278.

Bauer:2014:ILI

[BGS^{+}14] R. K. Bauer, U. Gräwe, D. Stepputtis, C. Zimmermann, and C. Hammer. Identifying the location and importance of spawning sites of Western Baltic herring using a particle backtracking model. ICES Journal of Marine Science, 71(3):499–509, March 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/71/3/499/636471.

Bentley:2019:CSF

[BHB^{+}19] Jacob W. Bentley, David E. Hines, Stuart R. Borrett, Natalia Serpetti, Gema Hernandez-Milian, Clive Fox, Johanna J. Heymans, and David G. Reid. Combining scientific and fishers’ knowledge to co-create indicators of food web structure and function. ICES Journal of Marine Science, 76(7):2218–2234, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/7/2218/5535676.

Babcock:2015:BDM

[BHC^{+}15] Elizabeth A. Babcock, William J. Harford, Robin Coleman, Janet Gibson, Julio Maaz, James R. Foley, and Mauro Gongora. Bayesian depletion model estimates of spiny lobster abundance at two marine protected areas in Belize with or without in-season recruitment. ICES Journal of Marine Science, 72(S1):S232–S243, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/72/
suppl_1/i232/615437.

Braga-Henriques:2011:ABD

[BHCSP^{+}11] Andreia Braga-Henriques, Marina Carreiro-Silva, Filipe M. Porteiro, Valentina de Matos, Íris Sampaio, Oscar Ocaña, and Sérgio P. Ávila. The association between a deep-sea gastropod Pedicularia sicula (Caenogastropoda: Pediculariidae) and its coral host Errina dabneyi (Hydrozoa: Stylasteridae) in the Azores. ICES Journal of Marine Science, 68(2):399–407, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/68/2/399/614681.

Barbeaux:2013:CWP

[BHD13] Steven J. Barbeaux, John K. Horne, and Martin W. Dorn. Characterizing walleye pollock (Theragra chalcogramma) winter distribution from opportunistic acoustic data. ICES Journal of Marine Science, 70(6):1162–1173, September 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/6/1162/635729.

Berasategui:2012:SVE

[BHDB12] Anabela A. Berasategui, Mónica S. Hoffmeyer, M. Sofía Dutto, and Florencia Biancalana. Seasonal variation in the egg morphology of the copepod Eurytemora americana and its relationship with reproductive strategy in a temperate estuary in Argentina. ICES Journal of Marine Science, 69 (3):380–388, May 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/69/3/380/599866.

Botsford:2014:CRS

[BHFH14] Louis W. Botsford, Matthew D. Holland, John C. Field, and Alan Hastings. Cohort resonance: a significant component of fluctuations in recruitment, egg production, and catch of fished populations. ICES Journal of Marine Science, 71(8):2158–2170, October 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/8/2158/759719.

Bergstad:2014:IRE

[BHJ14] Odd Aksel Bergstad, Hege Øverbø Hansen, and Terje Jørgensen. Intermittent recruitment and exploitation pulse underlying temporal variability in a demersal deep-water fish population. ICES Journal of Marine Science, 71(8):2088–2100, October 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/71/
8/2088/749525.

Bekkevold:2015:GAM

[BHL^{+}15] Dorte Bekkevold, Sarah J. Helyar, Morten T. Limborg, Einar E. Nielsen, Jakob Hemmer-Hansen, Lotte A. W. Clausen, and Gary R. Carvalho. Gene-associated markers can assign origin in a weakly structured fish, Atlantic herring. ICES Journal of Marine Science, 72(6):1790–1801, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/6/1790/917973.

Brinkhof:2018:ERC

[BHLS18] Jesse Brinkhof, Bent Herrmann, Roger B. Larsen, and Manu Sistiaga. Escape rate for cod (Gadus morhua) from the codend during buffer towing. ICES Journal of Marine Science, 75(2):805–813, March 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/75/
2/805/4582356.

Busch:2013:PIO

[BHM13] D. Shallin Busch, Chris J. Harvey, and Paul McElhany. Potential impacts of ocean acidification on the Puget Sound food web. ICES Journal of Marine Science, 70(4):823– 833, July 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/70/4/823/727377.

Bui:2018:SLE

[BHM^{+}18] Samantha Bui, Elina Halttunen, Agnes M. Mohn, Tone Vågseth, and Frode Oppedal. Salmon lice evasion, susceptibility, retention, and development differ amongst host salmonid species. ICES Journal of Marine Science, 75(3): 1071–1079, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/3/1071/4735125.

Bell:2014:EDC

[BHMR14] Richard J. Bell, Jonathan A. Hare, John P. Manderson, and David E. Richardson. Externally driven changes in the abundance of summer and winter flounder. ICES Journal of Marine Science, 71(9):2416–2428, November 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/9/2416/594657.

Bateman:2011:DPD

[BHS11a] Kelly S. Bateman, Ruth J. Hicks, and Grant D. Stentiford. Disease profiles differ between non-fished and fished populations of edible crab (Cancer pagurus) from a major commercial fishery. ICES Journal of Marine Science, 68(10):2044– 2052, November 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/10/2044/612790.

Botsford:2011:IAS

[BHS^{+}11b] Louis W. Botsford, Matthew D. Holland, Jameal F. Samhouri, J. Wilson White, and Alan Hastings. Importance of age structure in models of the response of upper trophic levels to fishing and climate change. ICES Journal of Marine Science, 68(6):1270–1283, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/6/1270/706727.

Bradbury:2016:GMS

[BHS^{+}16] Ian R. Bradbury, Lorraine C. Hamilton, Timothy F. Sheehan, Gerald Chaput, Martha J. Robertson, J. Brian Dempson, David Reddin, Vicki Morris, Timothy King, and Louis Bernatchez. Genetic mixed-stock analysis disentangles spatial and temporal variation in composition of the West Greenland Atlantic salmon fishery. ICES Journal of Marine Science, 73(9):2311–2321, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/9/2311/2198471.

Bachiller:2013:ARC

[BI13] Eneko Bachiller and Xabier Irigoien. Allometric relations and consequences for feeding in small pelagic fish in the Bay of Biscay. ICES Journal of Marine Science, 70(1):232–243, January 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/70/1/232/662552.

Bedwell-Ivers:2017:RHZ

[BIKP^{+}17] Hayley E. Bedwell-Ivers, Marguerite S. Koch, Katherine E. Peach, Luke Joles, Elizabeth Dutra, and Carrie Manfrino. The role of in hospite zooxanthellae photophysiology and reef chemistry on elevated pCO_{2} effects in two branching Caribbean corals: Acropora cervicornis and Porites divaricata. ICES Journal of Marine Science, 74(4):1103–1112, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/4/1103/2907918.

Bacha:2014:PSE

[BJH^{+}14] Mahmoud Bacha, Sherif Jemaa, Azzedine Hamitouche, Khalef Rabhi, and Rachid Amara. Population structure of the European anchovy, Engraulis encrasicolus, in the SW Mediterranean Sea, and the Atlantic Ocean: evidence from otolith shape analysis. ICES Journal of Marine Science, 71(9):2429–2435, November 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/9/2429/2798174.

Bannister:2016:NFF

[BJH^{+}16] Raymond J. Bannister, Ingrid A. Johnsen, Pia K. Hansen, Tina Kutti, and Lars Asplin. Near- and far-field dispersal modelling of organic waste from Atlantic salmon aquaculture in fjord systems. ICES Journal of Marine Science, 73(9):2408–2419, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
9/2408/2198194.

Bedford:2019:CCN

[BJMG19] Jacob Bedford, David Johns, and Abigail McQuatters-Gollop. A century of change in North Sea plankton communities explored through integrating historical datasets. ICES Journal of Marine Science, 76(1):104–112, January 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/1/104/5127715.

Bjornsson:2018:FAS

[Bjö18] Björn Björnsson. Fish aggregating sound technique (FAST): how low-frequency sound could be used in fishing and ranching of cod. ICES Journal of Marine Science, 75(4):1258–1268, July 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/75/
4/1258/4817094.

Bahr:2017:SAC

[BJR17] Keisha D. Bahr, Paul L. Jokiel, and Ku‘ulei S. Rodgers. Seasonal and annual calcification rates of the Hawaiian reef coral, Montipora capitata, under present and future climate change scenarios. ICES Journal of Marine Science, 74(4): 1083–1091, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/74/4/1083/2907932.

Beazley:2013:DSS

[BKMdMS13] Lindsay I. Beazley, Ellen L. Kenchington, Francisco Javier Murillo, and María del Mar Sacau. Deep-sea sponge grounds enhance diversity and abundance of epibenthic megafauna in the Northwest Atlantic. ICES Journal of Marine Science, 70(7):1471–1490, November 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/7/1471/610631.

Bentley:2012:IPM

[BKSB12] Nokome Bentley, Terese H. Kendrick, Paul J. Starr, and Paul A. Breen. Influence plots and metrics: tools for better understanding fisheries catch-per-unit-effort standardizations. ICES Journal of Marine Science, 69(1):84–88, January 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/69/1/84/671405.

Boitsov:2012:AWT

[BKT12] Vladimir D. Boitsov, Alexey L. Karsakov, and Alexander G. Trofimov. Atlantic water temperature and climate in the Barents Sea, 2000–2009. ICES Journal of Marine Science, 69(5):833–840, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/69/5/833/653020.

Bjornsson:2011:SAF

[BKTS11] Björn Björnsson, Hjalti Karlsson, Vilhjálmur Thorsteinsson, and Jón Solmundsson. Should all fish in mark– recapture experiments be double-tagged? Lessons learned from tagging coastal cod (Gadus morhua). ICES Journal of Marine Science, 68(3):603–610, March 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/68/
3/603/658482.

Barange:2016:EIN

[BKVT16] Manuel Barange, Jacquelynne King, Luis Valdés, and Alexander Turra. The evolving and increasing need for climate change research on the oceans. ICES Journal of Marine Science, 73(5):1267–1271, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
5/1267/2240705.

Brody:2015:CUO

[BL15] Sarah R. Brody and M. Susan Lozier. Characterizing upper-ocean mixing and its effect on the spring phytoplankton bloom with in situ data. ICES Journal of Marine Science, 72(6):1961–1970, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/72/
6/1961/918062.

Butcher:2012:GMC

[BLB^{+}12] Paul A. Butcher, Jesse C. Leland, Matt K. Broadhurst, Brian D. Paterson, and David G. Mayer. Giant mud crab (Scylla serrata): relative efficiencies of common baited traps and impacts on discards. ICES Journal of Marine Science, 69(8):1511–1522, September 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/8/1511/702327.

Bornarel:2018:MRE

[BLB^{+}18] Virginie Bornarel, Patrick Lambert, Cédric Briand, Carlos Antunes, Claude Belpaire, Eleonora Ciccotti, Estibaliz Diaz, Ola Diserud, Denis Doherty, Isabel Domingos, Derek Evans, Martin de Graaf, Ciara O’Leary, Michael Pedersen, Russell Poole, Alan Walker, Håkan Wickström, Laurent Beaulaton, and Hilaire Drouineau. Modelling the recruitment of European eel (Anguilla anguilla) throughout its European range. ICES Journal of Marine Science, 75(2): 541–552, March 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/2/541/4259273.

Bailey:2017:BGB

[BLD17] Jennifer L. Bailey, Yajie Liu, and Jan Grimsrud Davidsen. Bridging the gap between fisheries science and society: exploring fisheries science as a social activity. ICES Journal of Marine Science, 74(2):598–611, March 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/2/598/2734643.

Bryhn:2017:CIS

[BLJ^{+}17] Andreas C. Bryhn, Karl Lundström, Amelie Johansson, Henrik Ragnarsson Stabo, and Henrik Svedäng. A continuous involvement of stakeholders promotes the ecosystem approach to fisheries in the 8-fjords area on the Swedish west coast. ICES Journal of Marine Science, 74(1):431– 442, January 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/74/1/431/2734730.

Boch:2017:ECF

[BLM^{+}17] Charles A. Boch, Steven Y. Litvin, Fiorenza Micheli, Giulio De Leo, Emil A. Aalto, Christopher Lovera, C. Brock Woodson, Stephen Monismith, and James P. Barry. Effects of current and future coastal upwelling conditions on the fertilization success of the red abalone (Haliotis rufescens). ICES Journal of Marine Science, 74(4):1125–1134, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/4/1125/3094691.

Bergan:2017:EEC

[BLMW17] Alexander J. Bergan, Gareth L. Lawson, Amy E. Maas, and Zhaohui Aleck Wang. The effect of elevated carbon dioxide on the sinking and swimming of the shelled pteropod Limacina retroversa. ICES Journal of Marine Science, 74(7):1893–1905, September 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
7/1893/3096456.

Brodie:2018:CSR

[BLS^{+}18] S. Brodie, L. Litherland, J. Stewart, H. T. Schilling, J. G. Pepperell, and I. M. Suthers. Citizen science records describe the distribution and migratory behaviour of a piscivorous predator , Pomatomus saltatrix. ICES Journal of Marine Science, 75(5):1573–1582, September 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/5/1573/4996701.

Beare:2012:BTT

[BM12] Doug Beare and Marcel Machiels. Beam trawlermen take feet off gas in response to oil price hikes. ICES Journal of Marine Science, 69(6):1064–1068, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/6/1064/618777.

Butler:2015:EGF

[BM15] Casey B. Butler and Thomas R. Matthews. Effects of ghost fishing lobster traps in the Florida keys. ICES Journal of Marine Science, 72(S1):S185–S198, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/72/
suppl_1/i185/616418.

Benchetrit:2016:CHD

[BM16] José Benchetrit and James D. McCleave. Current and historical distribution of the American eel Anguilla rostrata in the countries and territories of the wider Caribbean. ICES Journal of Marine Science, 73(1):122–134, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/1/122/2457886.

Brock:2011:I

[BMBE11] Robert J. Brock, Gui Manuel Machado Menezes, Odd Aksel Bergstad, and Elizabethann English. Introduction. ICES Journal of Marine Science, 68(2):253, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/2/253/617417.

Boyra:2013:ASJ

[BMC^{+}13] G. Boyra, U. Martínez, U. Cotano, M. Santos, X. Irigoien, and A. Uriarte. Acoustic surveys for juvenile anchovy in the Bay of Biscay: abundance estimate as an indicator of the next year’s recruitment and spatial distribution patterns. ICES Journal of Marine Science, 70(7):1354–1368, November 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/70/7/1354/608431.

Browne:2017:GCC

[BMC^{+}17] D. Browne, C. Minto, R. Cosgrove, B. Burke, D. Mc-Donald, R. Officer, and M. Keatinge. A general catch comparison method for multi-gear trials: application to a quad-rig trawling fishery for Nephrops. ICES Journal of Marine Science, 74(5):1458–1468, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/5/1458/2936868.

Bauer:2018:REI

[BMC^{+}18] Barbara Bauer, H. E. Markus Meier, Michele Casini, Ayoe Hoff, Piotr Margoński, Alessandro Orio, Sofia Saraiva, Jeroen Steenbeek, and Maciej T. Tomczak. Reducing eutrophication increases spatial extent of communities supporting commercial fisheries: a model case study. ICES Journal of Marine Science, 75(4):1306–1317, July 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/4/1306/4828189.

Buhl-Mortensen:2016:TDM

[BMEBM^{+}16] Lene Buhl-Mortensen, Kari E. Ellingsen, Pål Buhl-Mortensen, Kristian L. Skaar, and Genoveva Gonzalez-Mirelis. Trawling disturbance on megabenthos and sediment in the Barents Sea: chronic effects on density, diversity, and composition. ICES Journal of Marine Science, 73 (suppl 1):S98–S114, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/73/suppl_1/i98/
2573995.

Boulcott:2014:ISD

[BMF14] Philip Boulcott, Colin P. Millar, and Rob J. Fryer. Impact of scallop dredging on benthic epifauna in a mixedsubstrate habitat. ICES Journal of Marine Science, 71(4): 834–844, May 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/71/4/834/669416.

Beszczynska-Moller:2012:VAW

[BMFSH12] Agnieszka Beszczynska-Möller, Eberhard Fahrbach, Ursula Schauer, and Edmond Hansen. Variability in Atlantic water temperature and transport at the entrance to the Arctic Ocean, 1997–2010. ICES Journal of Marine Science, 69 (5):852–863, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/69/5/852/651648.

Butler:2015:EPS

[BMG15] Mark J. Butler IV, Alison Macdiarmid, and Gaya Gnanalingam. The effect of parental size on spermatophore production, egg quality, fertilization success, and larval characteristics in the Caribbean spiny lobster, Panulirus argus. ICES Journal of Marine Science, 72(S1):S115–S123, July 2015. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/72/suppl_1/i115/618433.

Batts:2019:EGP

[BMGB19] Luke Batts, Cóilín Minto, Hans Gerritsen, and Deirdre Brophy. Estimating growth parameters and growth variability from length frequency data using hierarchical mixture models. ICES Journal of Marine Science, 76(7):2150–2163, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/7/2150/5528148.

Buchheister:2016:STD

[BMH^{+}16] André Buchheister, Thomas J. Miller, Edward D. Houde, David H. Secor, and Robert J. Latour. Spatial and temporal dynamics of Atlantic menhaden (Brevoortia tyrannus) recruitment in the Northwest Atlantic Ocean. ICES Journal of Marine Science, 73(4):1147–1159, March 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/4/1147/2458747.

Britton:2019:RSU

[BMM^{+}19] Damon Britton, Craig N. Mundy, Christina M. McGraw, Andrew T. Revill, and Catriona L. Hurd. Responses of seaweeds that use CO_{2} as their sole inorganic carbon source to ocean acidification: differential effects of fluctuating pH but little benefit of CO_{2} enrichment. ICES Journal of Marine Science, 76(6):1860–1870, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/6/1860/5479978.

Buhl-Mortensen:2016:FIB

[BMNKA^{+}16] Lene Buhl-Mortensen, Francis Neat, Mariano Koen-Alonso, Carsten Hvingel, and Børge Holte. Fishing impacts on benthic ecosystems: an introduction to the 2014 ICES symposium special issue. ICES Journal of Marine Science, 73(suppl 1):S1–S4, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/73/
suppl_1/i1/2573997.

Boyra:2019:STS

[BMO^{+}19] G. Boyra, G. Moreno, B. Orue, B. Sobradillo, and I. Sancristobal. In situ target strength of bigeye tuna (Thunnus obesus) associated with fish aggregating devices. ICES Journal of Marine Science, 76(7):2446–2458, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/7/2446/5537349.

Briceno:2010:EDD

[BMR10a] Felipe Briceño, Maite Mascaró, and Carlos Rosas. Energy demand during exponential growth of Octopus maya: exploring the effect of age and weight. ICES Journal of Marine Science, 67(7):1501–1508, October 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/7/1501/663564.

Briceno:2010:GBM

[BMR10b] Felipe Briceño, Maite Mascaró, and Carlos Rosas. GLMMbased modelling of growth in juvenile Octopus maya siblings: does growth depend on initial size? ICES Journal of Marine Science, 67(7):1509–1516, October 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/7/1509/661957.

Backlin:2011:HAS

[BMR^{+}11] Britt-Marie Bäcklin, Charlotta Moraeus, Anna Roos, Eva Eklöf, and Ylva Lind. Health and age and sex distributions of Baltic grey seals (Halichoerus grypus) collected from bycatch and hunt in the Gulf of Bothnia. ICES Journal of Marine Science, 68(1):183–188, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/1/183/628139.

Boyra:2018:TSS

[BMS^{+}18] Guillermo Boyra, Gala Moreno, Bea Sobradillo, Isabel Pérez-Arjona, Igor Sancristobal, and David A. Demer. Target strength of skipjack tuna (Katsuwanus pelamis) associated with fish aggregating devices (FADs). ICES Journal of Marine Science, 75(5):1790–1802, September 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/5/1790/4973732.

Bevacqua:2019:DMC

[BMS^{+}19] Daniele Bevacqua, Paco Melià, Marcello Schiavina, Alain J. Crivelli, Giulio A. De Leo, and Marino Gatto. A demographic model for the conservation and management of the European eel: an application to a Mediterranean coastal lagoon. ICES Journal of Marine Science, 76(7):2164–2178, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/7/2164/5532120.

Broadhurst:2016:DMJ

[BMSC16] Matt K. Broadhurst, Russell B. Millar, Henry L. Spach, and Nathalia Colombo. Damage and mortality of juvenile seabob shrimp (Xiphopenaeus kroyeri) discarded in a tropical artisanal trawl fishery. ICES Journal of Marine Science, 73(9):2364–2369, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/9/2364/2198433.

Barrento:2010:LSI

[BMVPN10] Sara Barrento, António Marques, Paulo Vaz-Pires, and Maria Leonor Nunes. Live shipment of immersed crabs Cancer pagurus from England to Portugal and recovery in stocking tanks: stress parameter characterization. ICES Journal of Marine Science, 67(3):435–443, April 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/3/435/733388.

Berg:2016:ACO

[BN16] Casper W. Berg and Anders Nielsen. Accounting for correlated observations in an age-based state-space stock assessment model. ICES Journal of Marine Science, 73(7): 1788–1797, July 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/73/7/1788/2458744.

Brander:2013:OMP

[BNAH13] Keith Brander, Anna Neuheimer, Ken Haste Andersen, and Martin Hartvig. Overconfidence in model projections. ICES Journal of Marine Science, 70(6):1065–1068, September 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/6/1065/636646.

Bornatowski:2014:EIS

[BNB^{+}14] Hugo Bornatowski, Andrés Felipe Navia, Raul Rennó Braga, Vinícius Abilhoa, and Marco Fábio Maia Corrêa. Ecological importance of sharks and rays in a structural foodweb analysis in southern Brazil. ICES Journal of Marine Science, 71(7):1586–1592, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/7/1586/671404.

Bigne:2019:ONP

[BNB19] Matthieu Bigné, J. Rasmus Nielsen, and François Bastardie. Opening of the Norway pout box: will it change the ecological impacts of the North Sea Norway pout fishery? ICES Journal of Marine Science, 76(1):136–152, January 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/1/136/5115364.

Bastardie:2015:CMS

[BNE^{+}15] Francois Bastardie, J. Rasmus Nielsen, O. R. Eigaard, H. O. Fock, P. Jonsson, and V. Bartolino. Competition for marine space: modelling the Baltic Sea fisheries and effort displacement under spatial restrictions. ICES Journal of Marine Science, 72(3):824–840, March 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/3/824/701817.

Bastardie:2017:ECS

[BNE^{+}17] François Bastardie, J. Rasmus Nielsen, Margit Eero, Federico Fuga, and Anna Rindorf. Effects of changes in stock productivity and mixing on sustainable fishing and economic viability. ICES Journal of Marine Science, 74(2): 535–551, March 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/74/2/535/2669542.

Bastardie:2010:EBC

[BNK10] Francois Bastardie, J. Rasmus Nielsen, and Gerd Kraus. The eastern Baltic cod fishery: a fleet-based management strategy evaluation framework to assess the cod recovery plan of 2008. ICES Journal of Marine Science, 67(1):71–86, January 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/67/1/71/596160.

Bisagni:2019:IVG

[BNP19] James J. Bisagni, Owen C. Nichols, and Roger Pettipas. Interannual variability of Gulf Stream warm-core ring interactions with the outer continental shelf and potential broad scale relationships with longfin squid (Doryteuthis pealeii) relative abundance, 1981–2004. ICES Journal of Marine Science, 76(5):1257–1270, 09- 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/5/1257/5573649.

Bostrom:2012:CDR

[BÖBL12] Maria K. Boström, Örjan Östman, Mikaela A. J. Bergenius, and Sven-Gunnar Lunneryd. Cormorant diet in relation to temporal changes in fish communities. ICES Journal of Marine Science, 69(2):175–183, March 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/2/175/704394.

Berenshtein:2019:PIF

[BOP^{+}19] Igal Berenshtein, Shay O’Farrell, Natalie Perlin, James N. Sanchirico, Steven A. Murawski, Larry Perruso, and Claire B. Paris. Predicting the impact of future oilspill closures on fishery-dependent communities — a spatially explicit approach. ICES Journal of Marine Science, 76(7):2276–2285, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
7/2276/5536328.

Bowden:2014:TCS

[Bow14] Alison A. Bowden. Towards a comprehensive strategy to recover river herring on the Atlantic seaboard: lessons from Pacific salmon. ICES Journal of Marine Science, 71(3): 666–671, March 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/71/3/666/632784.

Brunel:2013:ASR

[BP13] Thomas Brunel and GerJan J. Piet. Is age structure a relevant criterion for the health of fish stocks? ICES Journal of Marine Science, 70(2):270–283, March 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/2/270/796652.

Beck:2015:ECR

[BPBD15] Florence Beck, Jean-Philippe Pezy, Alexandrine Baffreau, and Jean-Claude Dauvin. Effects of clam rake harvesting on the intertidal Ruditapes habitat of the English Channel. ICES Journal of Marine Science, 72(9):2663–2673, November 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/9/2663/2457907.

Brooks:2010:ARP

[BPC10] Elizabeth N. Brooks, Joseph E. Powers, and Enric Cortés. Analytical reference points for age-structured models: application to data-poor fisheries. ICES Journal of Marine Science, 67(1):165–175, January 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/67/
1/165/595670.

Bradshaw:2018:PSSa

[BPD^{+}18a] Corey J. A. Bradshaw, Thomas A. A. Prowse, Michael Drew, Bronwyn M. Gillanders, Steven C. Donnellan, and Charlie Huveneers. Predicting sustainable shark harvests when stock assessments are lacking. ICES Journal of Marine Science, 75(5):1591–1601, September 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/5/1591/4951540.

Bradshaw:2018:PSSb

[BPD^{+}18b] Corey J. A. Bradshaw, Thomas A. A. Prowse, Michael Drew, Bronwyn M. Gillanders, Steven C. Donnellan, and Charlie Huveneers. Predicting sustainable shark harvests when stock assessments are lacking. ICES Journal of Marine Science, 75(5):1840, September 2018.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/5/1840/5053963.

Bastari:2018:SPMa

[BPF^{+}18a] Azzurra Bastari, Daniela Pica, Francesco Ferretti, Fiorenza Micheli, and Carlo Cerrano. Sea pens in the Mediterranean Sea: habitat suitability and opportunities for ecosystem recovery. ICES Journal of Marine Science, 75(5):1722– 1732, September 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/5/1722/4904207.

Bastari:2018:SPMb

[BPF^{+}18b] Azzurra Bastari, Daniela Pica, Francesco Ferretti, Fiorenza Micheli, and Carlo Cerrano. Sea pens in the Mediterranean Sea: habitat suitability and opportunities for ecosystem recovery. ICES Journal of Marine Science, 75(6):2289–2291, November 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/6/2289/5182255.

Blanchet:2019:RMMa

[BPF^{+}19a] Marie-Anne Blanchet, Raul Primicerio, André Frainer, Susanne Kortsch, Mette Skern-Mauritzen, Andrey V. Dolgov, and Michaela Aschan. The role of marine mammals in the Barents Sea foodweb. ICES Journal of Marine Science, 76(S1):S37–S53, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/76/
Supplement_1/i37/5554618.

Blanchet:2019:RMMb

[BPF^{+}19b] Marie-Anne Blanchet, Raul Primicerio, André Frainer, Susanne Kortsch, Mette Skern-Mauritzen, Andrey V. Dolgov, and Michaela Aschan. The role of marine mammals in the Barents Sea foodweb. ICES Journal of Marine Science, 76(S1):S54, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/76/
Supplement_1/i54/5606754.

Boyce:2019:MDA

[BPF19c] Daniel G. Boyce, Brian Petrie, and Kenneth T. Frank. Multivariate determination of Atlantic herring population health in a large marine ecosystem. ICES Journal of Marine Science, 76(4):859–869, July 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
4/859/5301644.

Bayse:2016:FSB

[BPH16] Shannon M. Bayse, Michael V. Pol, and Pingguo He. Fish and squid behaviour at the mouth of a drop-chain trawl: factors contributing to capture or escape. ICES Journal of Marine Science, 73(6):1545–1556, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/6/1545/2458756.

Benoit:2013:CAM

[BPKH13] Hugues P. Benoît, Sébastien Plante, Molly Kroiz, and Thomas Hurlbut. A comparative analysis of marine fish species susceptibilities to discard mortality: effects of environmental factors, individual traits, and phylogeny. ICES Journal of Marine Science, 70(1):99–113, January 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/1/99/660595.

Brandon:2017:TTS

[BPMR17] John R. Brandon, André E. Punt, Paula Moreno, and Randall R. Reeves. Toward a tier system approach for calculating limits on human-caused mortality of marine mammals. ICES Journal of Marine Science, 74(3):877–887, March 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/3/877/2736330.

Borrell:2012:MDM

[BPPB12] Yaisel J. Borrell, Jorge A. Piñera, José A. Sánchez Prado, and Gloria Blanco. Mitochondrial DNA and microsatellite genetic differentiation in the European anchovy Engraulis encrasicolus L. ICES Journal of Marine Science, 69(8):1357–1371, September 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/69/
8/1357/704144.

Brigolin:2017:MSS

[BPPP17] Daniele Brigolin, Erika Maria Diletta Porporato, Giuseppe Prioli, and Roberto Pastres. Making space for shellfish farming along the Adriatic coast. ICES Journal of Marine Science, 74(6):1540–1551, July 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
6/1540/3071823.

Beguer-Pon:2016:ERP

[BPST^{+}16] Mélanie Béguer-Pon, Shiliang Shan, Keith R. Thompson, Martin Castonguay, Jinyu Sheng, and Julian J. Dodson. Exploring the role of the physical marine environment in silver eel migrations using a biophysical particle tracking model. ICES Journal of Marine Science, 73(1):57–74, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/1/57/2458881.

Brunel:2010:PHC

[BPvHR10] Thomas Brunel, Gerjan J. Piet, Ralf van Hal, and Christine Röckmann. Performance of harvest control rules in a variable environment. ICES Journal of Marine Science, 67(5): 1051–1062, July 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/67/5/1051/608950.

Beaugrand:2012:RBN

[BR12] Grégory Beaugrand and Philip C. Reid. Relationships between North Atlantic salmon, plankton, and hydroclimatic change in the Northeast Atlantic. ICES Journal of Marine Science, 69(9):1549–1562, November 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/9/1549/640973.

Brander:2018:ST

[Bra18] Keith M. Brander. Seeing through. ICES Journal of Marine Science, 75(5):1536–1545, September 2018.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/5/1536/4980378.

Bell:2015:DEC

[BRH^{+}15] Richard J. Bell, David E. Richardson, Jonathan A. Hare, Patrick D. Lynch, and Paula S. Fratantoni. Disentangling the effects of climate, abundance, and size on the distribution of marine fish: an example based on four stocks from the Northeast US shelf. ICES Journal of Marine Science, 72(5):1311–1322, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/5/1311/760258.

Browman:2012:QV

[Bro12] Howard I. Browman. Quo vadimus. ICES Journal of Marine Science, 69(1):1–2, January 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/69/1/1/675458.

Browman:2014:CYS

[Bro14] Howard I. Browman. Commemorating 100 years since Hjort’s 1914 treatise on fluctuations in the great fisheries of northern Europe: where we have been, where we are, and where we are going. ICES Journal of Marine Science, 71(8):1989–1992, October 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/71/
8/1989/763523.

Browman:2016:AOS

[Bro16] Howard I. Browman. Applying organized scepticism to ocean acidification research. ICES Journal of Marine Science, 73(3):529–536, February 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
3/529/2459146.

Browman:2017:QVR

[Bro17a] Howard I. Browman. Quo vadimus — redux. ICES Journal of Marine Science, 74(1):1–2, January 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/1/1/2967557.

Browman:2017:TBP

[Bro17b] Howard I. Browman. Towards a broader perspective on ocean acidification research. ICES Journal of Marine Science, 74(4):889–894, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
4/889/3852373.

Brunel:2010:ASD

[Bru10] Thomas Brunel. Age-structure-dependent recruitment: a meta-analysis applied to Northeast Atlantic fish stocks. ICES Journal of Marine Science, 67(9):1921–1930, December 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/67/9/1921/617906.

Bernard:2013:KBA

[BS13] Kim S. Bernard and Deborah K. Steinberg. Krill biomass and aggregation structure in relation to tidal cycle in a penguin foraging region off the Western Antarctic Peninsula. ICES Journal of Marine Science, 70(4):834–849, July 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/4/834/729071.

Browman:2014:ELH

[BS14] Howard I. Browman and Anne Berit Skiftesvik. The early life history of fish — there is still a lot of work to do! ICES Journal of Marine Science, 71(4):907–908, May 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/4/907/669653.

Bacheler:2013:WTT

[BSB^{+}13] Nathan M. Bacheler, Zeb H. Schobernd, David J. Berrane, Christina M. Schobernd, Warren A. Mitchell, and Nathan R. Geraldi. When a trap is not a trap: converging entry and exit rates and their effect on trap saturation of black sea bass (Centropristis striata). ICES Journal of Marine Science, 70(4):873–882, July 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/70/
4/873/727488.

Briton:2019:RLE

[BSB^{+}19] Florence Briton, Lynne Shannon, Nicolas Barrier, Philippe Verley, and Yunne-Jai Shin. Reference levels of ecosystem indicators at multispecies maximum sustainable yield. ICES Journal of Marine Science, 76(7):2070–2081, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/7/2070/5522151.

Bethoney:2014:ELA

[BSC14] N. David Bethoney, Kevin D. E. Stokesbury, and Steven X. Cadrin. Environmental links to alosine at-sea distribution and bycatch in the Northwest Atlantic midwater trawl fishery. ICES Journal of Marine Science, 71(5):1246–1255, July 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/71/5/1246/637251.

Bentley:2019:FKI

[BSF^{+}19] Jacob W. Bentley, Natalia Serpetti, Clive Fox, Johanna J. Heymans, and David G. Reid. Fishers’ knowledge improves the accuracy of food web model predictions. ICES Journal of Marine Science, 76(4):897–912, July 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/4/897/5304545.

Bignami:2017:CEE

[BSHC17] Sean Bignami, Su Sponaugle, Martha Hauff, and Robert K. Cowen. Combined effects of elevated pCO_{2}, temperature, and starvation stress on larvae of a large tropical marine fish. ICES Journal of Marine Science, 74(4):1220–1229, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/4/1220/2742041.

Baltazar-Soares:2018:IPG

[BSHE18] Miguel Baltazar-Soares, Hans-Harald Hinrichsen, and Christophe Eizaguirre. Integrating population genomics and biophysical models towards evolutionary-based fisheries management. ICES Journal of Marine Science, 75(4): 1245–1257, July 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/4/1245/4791960.

Bastian:2011:FBD

[BSK^{+}11] Thomas Bastian, David Stokes, Jane E. Kelleher, Graeme C. Hays, John Davenport, and Thomas K. Doyle. Fisheries bycatch data provide insights into the distribution of the mauve stinger (Pelagia noctiluca) around Ireland. ICES Journal of Marine Science, 68(3):436–443, March 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/3/436/657228.

Brooker:2018:PMI

[BSMB18] A. J. Brooker, R. Skern-Mauritzen, and J. E. Bron. Production, mortality, and infectivity of planktonic larval sea lice, Lepeophtheirus salmonis (Krøyer, 1837): current knowledge and implications for epidemiological modelling. ICES Journal of Marine Science, 75(4):1214–1234, July 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/4/1214/4877008.

Bjornsson:2015:CPC

[BSP15] Björn Björnsson, Jón Sólmundsson, and Ólafur K. Pálsson. Can permanent closures of nearshore areas reduce the proportions of undersized fish in the Icelandic longline fishery? ICES Journal of Marine Science, 72(3):841–850, March 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/3/841/2835877.

Bundy:2010:GIB

[BSR^{+}10] Alida Bundy, Lynne J. Shannon, Marie-Joëlle Rochet, Sergio Neira, Yunne-Jai Shin, Louize Hill, and Kerim Aydin. The good(ish), the bad, and the ugly: a tripartite classification of ecosystem trends. ICES Journal of Marine Science, 67(4):745–768, May 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/67/4/745/682056.

Bernal:2011:RDEa

[BSW^{+}11a] Miguel Bernal, Yorgos Stratoudakis, Simon Wood, Leire Ibaibarriaga, Andres Uriarte, Luis Valdés, and David Borchers. A revision of daily egg production estimation methods, with application to Atlanto–Iberian sardine. 1. Daily spawning synchronicity and estimates of egg mortality. ICES Journal of Marine Science, 68(3):519–527, March 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/68/3/519/660797.

Bernal:2011:RDEb

[BSW^{+}11b] Miguel Bernal, Yorgos Stratoudakis, Simon Wood, Leire Ibaibarriaga, Luis Valdés, and David Borchers. A revision of daily egg production estimation methods, with application to Atlanto–Iberian sardine. 2. Spatially and environmentally explicit estimates of egg production. ICES Journal of Marine Science, 68(3):528–536, March 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/3/528/661120.

Brillant:2010:ELF

[BT10] Sean W. Brillant and Edward A. Trippel. Elevations of lobster fishery groundlines in relation to their potential to entangle endangered North Atlantic right whales in the Bay of Fundy, Canada. ICES Journal of Marine Science, 67(2): 355–364, March 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/67/2/355/691705.

Berkson:2015:DDP

[BT15] Jim Berkson and James T. Thorson. The determination of data-poor catch limits in the United States: is there a better way? ICES Journal of Marine Science, 72(1):237– 242, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/72/1/237/825263.

Barnett:2018:LDR

[BTA^{+}18] Beverly K. Barnett, Laura Thornton, Robert Allman, Jeffrey P. Chanton, and William F. Patterson III. Linear decline in red snapper (Lutjanus campechanus) otolith Δ 14C extends the utility of the bomb radiocarbon chronometer for fish age validation in the Northern Gulf of Mexico. ICES Journal of Marine Science, 75(5):1664–1671, September 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/5/1664/4992258.

Bailey:2017:ELS

[BTB^{+}17] Allison Bailey, Peter Thor, Howard I. Browman, David M. Fields, Jeffrey Runge, Alexander Vermont, Reidun Bjelland, Cameron Thompson, Steven Shema, Caroline M. F. Durif, and Haakon Hop. Early life stages of the Arctic copepod Calanus glacialis are unaffected by increased seawater pCO_{2}. ICES Journal of Marine Science, 74(4):996– 1004, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/4/996/2907927.

Becker:2017:MRA

[BTL17] Alistair Becker, Matthew D. Taylor, and Michael B. Lowry. Monitoring of reef associated and pelagic fish communities on Australia’s first purpose built offshore artificial reef. ICES Journal of Marine Science, 74(1):277–285, January 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/1/277/2669552.

Bourdaud:2017:IAS

[BTTV^{+}17] Pierre Bourdaud, Morgane Travers-Trolet, Youen Vermard, Xochitl Cormon, and Paul Marchal. Inferring the annual, seasonal, and spatial distributions of marine species from complementary research and commercial vessels’ catch rates. ICES Journal of Marine Science, 74(9):2415–2426, November 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/9/2415/3858384.

Bergman:2015:EYT

[BUDM15] Magda J. N. Bergman, Selma M. Ubels, Gerard C. A. Duineveld, and Erik W. G. Meesters. Effects of a 5-year trawling ban on the local benthic community in a wind farm in the Dutch coastal zone. ICES Journal of Marine Science, 72(3):962–972, March 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/72/3/962/696364.

Baudron:2010:CEM

[BUNB10] Alan Baudron, Clara Ulrich, J. Rasmus Nielsen, and Jesper Boje. Comparative evaluation of a mixed-fisheries effortmanagement system based on the Faroe Islands example. ICES Journal of Marine Science, 67(5):1036–1050, July 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/5/1036/607659.

Bjorkan:2019:BCPb

[BV19a] Maiken Bjørkan and Siri Veland. Beyond concensus: perceptions of risk from petroleum developments in Lofoten, Vesterålen, and Senja, Norway. ICES Journal of Marine Science, 76(6):1935, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
6/1935/5513065.

Bjorkan:2019:BCPa

[BV19b] Maiken Bjørkan and Siri Veland. Beyond consensus: perceptions of risk from petroleum developments in Lofoten, Vesterålen, and Senja, Norway. ICES Journal of Marine Science, 76(6):1393–1403, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/6/1393/5435725.

Breen:2015:MIF

[BVC15] Patricia Breen, Koen Vanstaen, and Robert W. E. Clark. Mapping inshore fishing activity using aerial, land, and vessel-based sighting information. ICES Journal of Marine Science, 72(2):467–479, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/72/
2/467/2801298.

Beraud:2018:IOC

[BvdMA^{+}18] Claire Beraud, Johan van der Molen, Mike Armstrong, Ewan Hunter, Leila Fonseca, and Kieran Hyder. The influence of oceanographic conditions and larval behaviour on settlement success — the European sea bass Dicentrarchus labrax (L.). ICES Journal of Marine Science, 75(2): 455–470, March 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/2/455/4608921.

Bernotas:2016:DEE

[BVS^{+}16] P. Bernotas, M. Vetemaa, L. Saks, R. Eschbaum, A. Verliin, and A. Järvalt. Dynamics of European eel landings and stocks in the coastal waters of Estonia. ICES Journal of Marine Science, 73(1):84–90, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/1/84/2458922.

Brown:2018:CCZ

[BVW^{+}18] Elliot J. Brown, Rita P. Vasconcelos, Håkan Wennhage, Ulf Bergström, Josianne G. Støttrup, Karen van de Wolfshaar, Giacomo Millisenda, Francesco Colloca, and Olivier Le Pape. Conflicts in the coastal zone: human impacts on commercially important fish species utilizing coastal habitat. ICES Journal of Marine Science, 75(4):1203–1213, July 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/4/1203/4788357.

Becker:2014:MPN

[BW14] Kaylyn N. Becker and Joseph D. Warren. Material properties of Northeast Pacific zooplankton. ICES Journal of Marine Science, 71(9):2550–2563, November 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/9/2550/2798180.

Brien:2016:PCI

[BWH16] H. V. Brien, S.-A. Watson, and M. O. Hoogenboom. Presence of competitors influences photosynthesis, but not growth, of the hard coral Porites cylindrica at elevated seawater CO_{2}. ICES Journal of Marine Science, 73(3):659– 669, February 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/73/3/659/2458741.

Belivermis:2016:LEI

[BWM^{+}16] Murat Belivermiş, Michel Warnau, Marc Metian, François Oberhänsli, Jean-Louis Teyssié, and Thomas Lacoue-Labarthe. Limited effects of increased CO_{2} and temperature on metal and radionuclide bioaccumulation in a sessile invertebrate, the oyster Crassostrea gigas. ICES Journal of Marine Science, 73(3):753–763, February 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/3/753/2459096.

Brodie:2010:ECR

[BWP10] William B. Brodie, Stephen J. Walsh, and Dawn Maddock Parsons. An evaluation of the collapse and recovery of the yellowtail flounder (Limanda ferruginea) stock on the Grand Bank. ICES Journal of Marine Science, 67(9):1887– 1895, December 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/67/9/1887/620852.

Baumann:2015:COM

[BWR^{+}15] Hannes Baumann, R. J. D. Wells, Jay R. Rooker, Saijin Zhang, Zofia Baumann, Daniel J. Madigan, Heidi Dewar, Owyn E. Snodgrass, and Nicholas S. Fisher. Combining otolith microstructure and trace elemental analyses to infer the arrival of juvenile Pacific bluefin tuna in the California current ecosystem. ICES Journal of Marine Science, 72(7):2128–2138, October 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/7/2128/2457866.

Bucklin:2019:TSM

[BYQ^{+}19] Ann Bucklin, Heidi D. Yeh, Jennifer M. Questel, David E. Richardson, Bo Reese, Nancy J. Copley, and Peter H. Wiebe. Time-series metabarcoding analysis of zooplankton diversity of the NW Atlantic continental shelf. ICES Journal of Marine Science, 76(4):1162–1176, July 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/4/1162/5333157.

Cosgrove:2010:SAT

[CAB^{+}10] Ronan Cosgrove, Igor Arregi, Deirdre Brophy, Haritz Arrizabalaga, Victoria Ortiz de Zarate, and Nigel Griffin. A simulated archival tagging programme for albacore (Thunnus alalunga) in the Northeast Atlantic, including an analysis of factors affecting tag recovery. ICES Journal of Marine Science, 67(6):1216–1221, September 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/6/1216/737567.

Caruso:2016:MCP

[CAC^{+}16] Gabriella Caruso, Maurizio Azzaro, Carmela Caroppo, Franco Decembrini, Luis Salvador Monticelli, Marcella Leonardi, Giovanna Maimone, Renata Zaccone, and Rosabruna La Ferla. Microbial community and its potential as descriptor of environmental status. ICES Journal of Marine Science, 73(9):2174–2177, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
9/2174/2199434.

Canales:2018:LTS

[CAC^{+}18] Cristian M. Canales, Nicolás A. Adasme, Luis A. Cubillos, Maria Jose Cuevas, and Nazareth Sánchez. Longtime spatio-temporal variations in anchovy (Engraulis ringens) biological traits off northern Chile: an adaptive response to long-term environmental change? ICES Journal of Marine Science, 75(6):1908–1923, November 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/6/1908/5052273.

Cadigan:2013:FNP

[Cad13] Noel G. Cadigan. Fitting a non-parametric stock– recruitment model in R that is useful for deriving MSY reference points and accounting for model uncertainty. ICES Journal of Marine Science, 70(1):56–67, January 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/1/56/663017.

Caddy:2014:WDA

[Cad14] John F. Caddy. Why do assessments of demersal stocks largely ignore habitat? ICES Journal of Marine Science, 71(8):2114–2126, October 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/71/
8/2114/743879.

Cabral:2014:MIF

[CAL14] Reniel B. Cabral, Porfirio M. Aliño, and May T. Lim. Modelling the impacts of fish aggregating devices (FADs) and fish enhancing devices (FEDs) and their implications for managing small-scale fishery. ICES Journal of Marine Science, 71(7):1750–1759, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/7/1750/664488.

Carvalho:2015:UPS

[CAM^{+}15] Felipe Carvalho, Robert Ahrens, Debra Murie, Keith Bigelow, Alexandre Aires-Da-Silva, Mark N. Maunder, and Fábio Hazin. Using pop-up satellite archival tags to inform selectivity in fisheries stock assessment models: a case study for the blue shark in the South Atlantic Ocean. ICES Journal of Marine Science, 72(6):1715–1730, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/6/1715/918447.

Campana:2018:TES

[Cam18] Steven E. Campana. Twelve easy steps to embrace or avoid scientific petrification: lessons learned from a career in otolith research. ICES Journal of Marine Science, 75(1): 22–29, January 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/1/22/4100496.

Cabezas:2012:GCE

[CAMM12] Patricia Cabezas, Fernando Alda, Enrique Macpherson, and Annie Machordom. Genetic characterization of the endangered and endemic anchialine squat lobster Munidopsis polymorpha from Lanzarote (Canary Islands): management implications. ICES Journal of Marine Science, 69(6):1030– 1037, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/69/6/1030/619441.

Canadelli:2016:BSS

[Can16] Elena Canadelli. Biological stations and the study of marine life: Umberto D’Ancona and the Hydrobiological Station of Chioggia (1940–1964). ICES Journal of Marine Science, 73(5):1447–1457, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
5/1447/2296472.

Crespi-Abril:2015:DTA

[CAOG15] Augusto César Crespi-Abril, Nicolás Ortiz, and David Edgardo Galván. Decision tree analysis for the determination of relevant variables and quantifiable reference points to establish maturity stages in Enteroctopus megalocyathus and Illex argentinus. ICES Journal of Marine Science, 72(5): 1449–1461, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/72/5/1449/757624.

Clark:2016:IDS

[CAS^{+}16] Malcolm R. Clark, Franziska Althaus, Thomas A. Schlacher, Alan Williams, David A. Bowden, and Ashley A. Rowden. The impacts of deep-sea fisheries on benthic communities: a review. ICES Journal of Marine Science, 73(suppl 1):S51– S69, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/73/suppl_1/i51/2573992.

Chaput:2012:EBT

[CB12] Gérald Chaput and Hugues P. Benoît. Evidence for bottom–up trophic effects on return rates to a second spawning for Atlantic salmon (Salmo salar) from the Miramichi River, Canada. ICES Journal of Marine Science, 69(9):1656–1667, November 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/69/
9/1656/636674.

Chan:2014:RAR

[CBBM14] Farrah T. Chan, Elizabeta Briski, Sarah A. Bailey, and Hugh J. MacIsaac. Richness–abundance relationships for zooplankton in ballast water: temperate versus Arctic comparisons. ICES Journal of Marine Science, 71(7):1876– 1884, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/71/7/1876/671307.

Colaco:2011:MDT

[CBC^{+}11a] A. Colaço, J. Blandin, M. Cannat, T. Carval, V. Chavagnac, D. Connelly, M. Fabian, S. Ghiron, J. Goslin, J. M. Miranda, G. Reverdin, J. Sarrazin, C. Waldmann, and P. M. Sarradin. MoMAR-d: a technological challenge to monitor the dynamics of the lucky strike vent ecosystem. ICES Journal of Marine Science, 68(2):416–424, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/2/416/615925.

Colaco:2011:LCA

[CBC^{+}11b] Ana Colaço, Raul Bettencourt, Valentina Costa, Silvia Lino, Humberto Lopes, Inês Martins, Luis Pires, Catarina Prieto, and Ricardo Serrão Santos. LabHorta: a controlled aquarium system for monitoring physiological characteristics of the hydrothermal vent mussel Bathymodiolus azoricus. ICES Journal of Marine Science, 68(2):349–356, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/68/2/349/617247.

Cadrin:2010:PSB

[CBD^{+}10a] Steven X. Cadrin, Matthias Bernreuther, Anna Kristín Daníelsdóttir, Einar Hjörleifsson, Torild Johansen, Lisa Kerr, Kristjan Kristinsson, Stefano Mariani, Kjell Nedreaas, Christophe Pampoulie, Benjamin Planque, Jákup Reinert, Fran Saborido-Rey, Thorsteinn Sigurdsson, and Christoph Stransky. Population structure of beaked redfish, Sebastes mentella: evidence of divergence associated with different habitats. ICES Journal of Marine Science, 67(8):1617–1630, November 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/67/

8/1617/603289. See comment [MAP^{+}11] and reply [CMP^{+}11].

Cutter:2010:CBM

[CBD10b] George R. Cutter, Jr., Laurent Berger, and David A. Demer. A comparison of bathymetry mapped with the Simrad ME70 multibeam echosounder operated in bathymetric and fisheries modes. ICES Journal of Marine Science, 67(6):1301–1309, September 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/67/
6/1301/733705.

Colefax:2018:PUA

[CBK18] Andrew P. Colefax, Paul A. Butcher, and Brendan P. Kelaher. The potential for unmanned aerial vehicles (UAVs) to conduct marine fauna surveys in place of manned aircraft. ICES Journal of Marine Science, 75(1):1–8, January 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/1/1/3862123.

Chan:2019:EDE

[CBL19] Maggie N. Chan, Anne H. Beaudreau, and Philip A. Loring. Exploring diversity in expert knowledge: variation in local ecological knowledge of Alaskan recreational and subsistence fishers. ICES Journal of Marine Science, 76(4): 913–924, July 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/76/4/913/5259367.

Ciannelli:2015:EEC

[CBO15] Lorenzo Ciannelli, Kevin Bailey, and Esben Moland Olsen. Evolutionary and ecological constraints of fish spawning habitats. ICES Journal of Marine Science, 72(2):285–296, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/2/285/2801424.

Chassot:2011:SRS

[CBR^{+}11] Emmanuel Chassot, Sylvain Bonhommeau, Gabriel Reygondeau, Karen Nieto, Jeffrey J. Polovina, Martin Huret, Nicholas K. Dulvy, and Herve Demarcq. Satellite remote sensing for an ecosystem approach to fisheries management. ICES Journal of Marine Science, 68(4):651–666, March 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/68/4/651/650740.

Cortes:2015:RAC

[CBS15] Enric Cortés, Elizabeth N. Brooks, and Kyle W. Shertzer. Risk assessment of cartilaginous fish populations. ICES Journal of Marine Science, 72(3):1057–1068, March 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/3/1057/690224.

Clarke:2015:EES

[CBW15] J. Clarke, D. M. Bailey, and P. J. Wright. Evaluating the effectiveness of a seasonal spawning area closure. ICES Journal of Marine Science, 72(9):2627–2637, November 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/9/2627/2458713.

Connolly:2011:SCF

[CC11] P. L. Connolly and L. Caffrey. Supply chaining fishery advice. ICES Journal of Marine Science, 68(8):1706– 1711, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/8/1706/761228.

Cadigan:2017:HMB

[CC17] Noel G. Cadigan and Steven E. Campana. Hierarchical model-based estimation of population growth curves for redfish (Sebastes mentella and Sebastes fasciatus) off the Eastern coast of Canada. ICES Journal of Marine Science, 74(3):687–697, March 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
3/687/2742060.

Correia:2018:DRE

[CCA^{+}18] Maria João Correia, José Lino Costa, Carlos Antunes, Giulio De Leo, and Isabel Domingos. The decline in recruitment of the European eel: new insights from a 40-year-long time-series in the Minho estuary (Portugal). ICES Journal of Marine Science, 75(6):1975–1983, November 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/6/1975/5046594.

Chouvelon:2014:SSR

[CCC^{+}14] T. Chouvelon, F. Caurant, Y. Cherel, B. Simon-Bouhet, J. Spitz, and P. Bustamante. Species- and size-related patterns in stable isotopes and mercury concentrations in fish help refine marine ecosystem indicators and provide evidence for distinct management units for hake in the Northeast Atlantic. ICES Journal of Marine Science, 71(5):1073– 1087, July 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/71/5/1073/644162.

Chaput:2019:ASS

[CCD^{+}19] Gérald Chaput, Jonathan Carr, Jason Daniels, Steve Tinker, Ian Jonsen, and Frederick Whoriskey. Atlantic salmon (Salmo salar) smolt and early post-smolt migration and survival inferred from multi-year and multi-stock acoustic telemetry studies in the Gulf of St. Lawrence, northwest Atlantic. ICES Journal of Marine Science, 76(4):1107– 1121, July 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/4/1107/5230889.

Comeau:2017:EPP

[CCE17] S. Comeau, R. C. Carpenter, and P. J. Edmunds. Effects of pCO_{2} on photosynthesis and respiration of tropical scleractinian corals and calcified algae. ICES Journal of Marine Science, 74(4):1092–1102, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
4/1092/2907909.

Condie:2014:STI

[CCG14] Harriet M. Condie, Thomas L. Catchpole, and Alastair Grant. The short-term impacts of implementing catch quotas and a discard ban on English North Sea otter trawlers. ICES Journal of Marine Science, 71(5):1266–1276, July 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/5/1266/643693.

Chust:2014:CSS

[CCL^{+}14] Guillem Chust, Claudia Castellani, Priscilla Licandro, Leire Ibaibarriaga, Yolanda Sagarminaga, and Xabier Irigoien. Are Calanus spp. shifting poleward in the North Atlantic? A habitat modelling approach. ICES Journal of Marine Science, 71(2):241–253, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/2/241/787087.

Cleary:2010:PEH

[CCS10] J. S. Cleary, S. P. Cox, and J. F. Schweigert. Performance evaluation of harvest control rules for Pacific herring management in British Columbia, Canada. ICES Journal of Marine Science, 67(9):2005–2011, December 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/9/2005/623391.

Conroy:2018:DRS

[CCSG18] Christian W. Conroy, Jay Calvert, Graham D. Sherwood, and Jonathan H. Grabowski. Distinct responses of sympatric migrant and resident Atlantic cod phenotypes to substrate and temperature at a remote Gulf of Maine seamount. ICES Journal of Marine Science, 75(1):122–134, January 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/1/122/3866450.

Chouinard:2011:SDF

[CD11] P-M. Chouinard and J-D. Dutil. The structure of demersal fish assemblages in a cold, highly stratified environment. ICES Journal of Marine Science, 68(9):1896–1908, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/68/9/1896/669629.

Cutter:2014:SCU

[CD14] George R. Cutter, Jr. and David A. Demer. Seabed classification using surface backscattering strength versus acoustic frequency and incidence angle measured with vertical, splitbeam echosounders. ICES Journal of Marine Science, 71 (4):882–894, May 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/71/4/882/666881.

Castonguay:2016:UDA

[CD16] Martin Castonguay and Caroline M. F. Durif. Understanding the decline in anguillid eels. ICES Journal of Marine Science, 73(1):1–4, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/73/1/1/2458925.

Crespo:2017:RIF

[CD17] Guillermo Ortuño Crespo and Daniel C. Dunn. A review of the impacts of fisheries on open-ocean ecosystems. ICES Journal of Marine Science, 74(9):2283–2297, November 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/9/2283/3855115.

Cooper:2014:NAJ

[CDAN^{+}14] Daniel W. Cooper, Janet T. Duffy-Anderson, Brenda L. Norcross, Brenda A. Holladay, and Phyllis J. Stabeno. Nursery areas of juvenile northern rock sole (Lepidopsetta polyxystra) in the eastern Bering Sea in relation to hydrography and thermal regimes. ICES Journal of Marine Science, 71(7):1683–1695, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/71/
7/1683/663174.

Cadrin:2015:SAM

[CDC15] Steven X. Cadrin and Mark Dickey-Collas. Stock assessment methods for sustainable fisheries. ICES Journal of Marine Science, 72(1):1–6, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/72/
1/1/2804348.

Cheung:2011:IEP

[CDSP11] William W. L. Cheung, John Dunne, Jorge L. Sarmiento, and Daniel Pauly. Integrating ecophysiology and plankton dynamics into projected maximum fisheries catch potential under climate change in the Northeast Atlantic. ICES Journal of Marine Science, 68(6):1008–1018, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/6/1008/699298.

Casini:2016:UAB

[CECL16] Michele Casini, Margit Eero, Sofia Carlshamre, and Johan Lövgren. Using alternative biological information in stock assessment: condition-corrected natural mortality of Eastern Baltic cod. ICES Journal of Marine Science, 73(10):2625–2631, November 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
10/2625/2647113.

Catchpole:2018:HDE

[CEP^{+}18] Thomas L. Catchpole, Sam Elliott, Dave Peach, Stephen C. Mangi, and Tim S. Gray. How to deal with the EU Landing Obligation: lessons from an English discard ban sea trial. ICES Journal of Marine Science, 75(1):270–278, January 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/1/270/3953901.

Cheung:2016:BCP

[CFA^{+}16] William W. L. Cheung, Thomas L. Frölicher, Rebecca G. Asch, Miranda C. Jones, Malin L. Pinsky, Gabriel Reygondeau, Keith B. Rodgers, Ryan R. Rykaczewski, Jorge L. Sarmiento, Charles Stock, and James R. Watson. Building confidence in projections of the responses of living marine resources to climate change. ICES Journal of Marine Science, 73(5):1283–1296, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
5/1283/2240697.

Chiaradia:2010:CDT

[CFHC10] André Chiaradia, Manuela G. Forero, Keith A. Hobson, and J. Mike Cullen. Changes in diet and trophic position of a top predator 10 years after a mass mortality of a key prey. ICES Journal of Marine Science, 67(8):1710–1720, November 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/67/8/1710/604480.

Catchpole:2014:UID

[CFM^{+}14] T. L. Catchpole, J. P. Feekings, N. Madsen, A. Palialexis, V. Vassilopoulou, J. Valeiras, T. Garcia, N. Nikolic, and M.-J. Rochet. Using inferred drivers of discarding behaviour to evaluate discard mitigation measures. ICES Journal of Marine Science, 71(5):1277–1285, July 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/5/1277/643538.

Church:2019:CNS

[CFT^{+}19] Gabriella E. Church, Robert W. Furness, Glen Tyler, Lucy Gilbert, and Stephen C. Votier. Change in the North Sea ecosystem from the 1970s to the 2010s: great skua diets reflect changing forage fish, seabirds, and fisheries. ICES Journal of Marine Science, 76(4):925–937, July 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/4/925/5237030.

Chan:2016:IOA

[CGAD16] Kit Yu Karen Chan, Daniel Grünbaum, Maj Arnberg, and Sam Dupont. Impacts of ocean acidification on survival, growth, and swimming behaviours differ between larval urchins and brittlestars. ICES Journal of Marine Science, 73(3):951–961, February 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
3/951/2457914.

Chandrapavan:2011:IMT

[CGG^{+}11] Arani Chandrapavan, Caleb Gardner, Bridget S. Green, Adrian Linnane, and David Hobday. Improving marketability through translocation: a lobster case study from southern Australia. ICES Journal of Marine Science, 68(9):1842–1851, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/68/
9/1842/670993.

Coston-Guarini:2017:RQE

[CGGH^{+}17] J. Coston-Guarini, J.-M. Guarini, Shawn Hinz, Jeff Wilson, and L. Chauvaud. A roadmap for a quantitative ecosystem-based environmental impact assessment. ICES Journal of Marine Science, 74(7):2012–2023, September 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/7/2012/3037985.

Chatzinikolaou:2017:CER

[CGK^{+}17] Eva Chatzinikolaou, Panagiotis Grigoriou, Kleoniki Keklikoglou, Sarah Faulwetter, and Nafsika Papageorgiou. The combined effects of reduced pH and elevated temperature on the shell density of two gastropod species measured using micro-CT imaging. ICES Journal of Marine Science, 74(4):1135–1149, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
4/1135/2739036.

Cox:2019:ICR

[CGM^{+}19] Kieran D. Cox, Travis G. Gerwing, Tara Macdonald, Margot Hessing-Lewis, Ben Millard-Martin, Rylan J. Command, Francis Juanes, and Sarah E. Dudas. Infaunal community responses to ancient clam gardens. ICES Journal of Marine Science, 76(7):2362–2373, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/7/2362/5542618.

Charles:2016:BHF

[CGR16] A. Charles, S. M. Garcia, and J. Rice. Balanced harvesting in fisheries: economic considerations. ICES Journal of Marine Science, 73(6):1679–1689, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/6/1679/2458695.

Coll:2013:UNT

[CGRM^{+}13] Josep Coll, Antoni Garcia-Rubies, Gabriel Morey, Olga Reñones, Diego Álvarez-Berastegui, Oliver Navarro, and Antoni M. Grau. Using no-take marine reserves as a tool for evaluating rocky-reef fish resources in the western Mediterranean. ICES Journal of Marine Science, 70(3):578–590, April 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/70/3/578/919323.

Cornwall:2016:EDO

[CH16] Christopher E. Cornwall and Catriona L. Hurd. Experimental design in ocean acidification research: problems and solutions. ICES Journal of Marine Science, 73(3):572–581, February 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/3/572/2458712.

Chaput:2012:OSA

[Cha12] Gérald Chaput. Overview of the status of Atlantic salmon (Salmo salar) in the North Atlantic and trends in marine mortality. ICES Journal of Marine Science, 69(9):1538– 1548, November 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/69/9/1538/635167.

Ciannelli:2014:TGE

[CHB^{+}14] Lorenzo Ciannelli, Mary Hunsicker, Anne Beaudreau, Kevin Bailey, Larry B. Crowder, Carmel Finley, Colleen Webb, John Reynolds, Kay Sagmiller, John M. Anderies, David Hawthorne, Julia Parrish, Selina Heppell, Flaxen Conway, and Paulinus Chigbu. Transdisciplinary graduate education in marine resource science and management. ICES Journal of Marine Science, 71(5):1047–1051, July 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/5/1047/648379.

Cook:2018:PSM

[CHD^{+}18] Katrina V. Cook, Scott G. Hinch, S. Matthew Drenner, Edmund A. Halfyard, Graham D. Raby, and Steven J. Cooke. Population-specific mortality in Coho salmon (Oncorhynchus kisutch) released from a purse seine fishery. ICES Journal of Marine Science, 75(1):309–318, January 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/1/309/4061561.

Chen:2010:ATT

[Che10] Chih-Shin Chen. Abundance trends of two neon flying squid (Ommastrephes bartramii) stocks in the North Pacific. ICES Journal of Marine Science, 67(7):1336–1345, October 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/67/7/1336/663807.

Childress:2015:JCS

[CHM15] Michael J. Childress, Katherine A. Heldt, and Scott D. Miller. Are juvenile Caribbean spiny lobsters (Panulirus argus) becoming less social? ICES Journal of Marine Science, 72(S1):S170–S176, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/72/
suppl_1/i170/619684.

Cole:2015:BWM

[CHMY15] Harriet S. Cole, Stephanie Henson, Adrian P. Martin, and Andrew Yool. Basin-wide mechanisms for spring bloom initiation: how typical is the North Atlantic? ICES Journal of Marine Science, 72(6):2029–2040, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/6/2029/917853.

Citores:2018:UEM

[CIJ18] Leire Citores, Leire Ibaibarriaga, and Ernesto Jardim. Uncertainty estimation and model selection in stock assessment models with non-parametric effects on fishing mortality. ICES Journal of Marine Science, 75(2):585–595, March 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/2/585/4430992.

Certain:2015:MVA

[CJC^{+}15] Grégoire Certain, Lis Lindahl Jørgensen, Isadora Christel, Benjamin Planque, and Vincent Bretagnolle. Mapping the vulnerability of animal community to pressure in marine systems: disentangling pressure types and integrating their impact from the individual to the community level. ICES Journal of Marine Science, 72(5):1470–1482, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/5/1470/772310.

Campana:2016:DHP

[CJFS16] Steven E. Campana, Warren Joyce, Mark Fowler, and Mark Showell. Discards, hooking, and post-release mortality of porbeagle (Lamna nasus), shortfin mako (Isurus oxyrinchus), and blue shark (Prionace glauca) in the Canadian pelagic longline fishery. ICES Journal of Marine Science, 73(2):520–528, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
2/520/2614471.

Cruz:2014:RDD

[CJP^{+}14] Maria João Cruz, Vera Leal Jordão, João Gil Pereira, Ricardo Serrão Santos, and Mónica A. Silva. Risso’s dolphin depredation in the Azorean hand-jig squid fishery: assessing the impacts and evaluating effectiveness of acoustic deterrents. ICES Journal of Marine Science, 71(9):2608– 2620, November 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/71/9/2608/2798168.

Calise:2012:MTS

[CK12] Lucio Calise and Tor Knutsen. Multifrequency target strength of northern krill (Meganyctiphanes norvegica) swimming horizontally. ICES Journal of Marine Science, 69(1):119–130, January 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/69/1/119/670723.

Constable:2018:MGR

[CK18] Andrew John Constable and So Kawaguchi. Modelling growth and reproduction of Antarctic krill, Euphausia superba, based on temperature, food and resource allocation amongst life history functions. ICES Journal of Marine Science, 75(2):738–750, March 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/2/738/4558664.

Cormier:2017:MEB

[CKA^{+}17] Roland Cormier, Christopher R. Kelble, M. Robin Anderson, J. Icarus Allen, Anthony Grehan, and Ólavur Gregersen. Moving from ecosystem-based policy objectives to operational implementation of ecosystem-based management measures. ICES Journal of Marine Science, 74(1): 406–413, January 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/74/1/406/2444580.

Carruthers:2016:PRS

[CKB^{+}16] Thomas R. Carruthers, Laurence T. Kell, Doug D. S. Butterworth, Mark N. Maunder, Helena F. Geromont, Carl Walters, Murdoch K. McAllister, Richard Hillary, Polina Levontin, Toshihide Kitakado, and Campbell R. Davies. Performance review of simple management procedures. ICES Journal of Marine Science, 73(2):464–482, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/2/464/2614445.

Churchill:2017:PLS

[CKD^{+}17] J. H. Churchill, J. P. Kritzer, M. J. Dean, J. H. Grabowski, and G. D. Sherwood. Patterns of larval-stage connectivity of Atlantic cod (Gadus morhua) within the Gulf of Maine in relation to current structure and a proposed fisheries closure. ICES Journal of Marine Science, 74(1):20–30, January 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/1/20/2669556.

Charbonneau:2019:TSA

[CKH19] Julie A. Charbonneau, David M. Keith, and Jeffrey A. Hutchings. Trends in the size and age structure of marine fishes. ICES Journal of Marine Science, 76(4):938– 945, July 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/4/938/5248192.

Cormon:2016:ENP

[CKV^{+}16] Xochitl Cormon, Alexander Kempf, Youen Vermard, Morten Vinther, and Paul Marchal. Emergence of a new predator in the North Sea: evaluation of potential trophic impacts focused on hake, saithe, and Norway pout. ICES Journal of Marine Science, 73(5):1370–1381, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/5/1370/2296623.

Clark:2018:MBF

[Cla18a] Colin W. Clark. Modelling the behaviour of fishers and fishes. ICES Journal of Marine Science, 75(3):932–940, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/3/932/4718118.

Claudet:2018:CFW

[Cla18b] Joachim Claudet. Claudet’s final word. ICES Journal of Marine Science, 75(3):1177, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/75/
3/1177/4098837.

Claudet:2018:CB

[Cla18c] Joachim Claudet. Counterpoint to bergeseth. ICES Journal of Marine Science, 75(3):1181–1182, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/3/1181/4098841.

Claudet:2018:SCU

[Cla18d] Joachim Claudet. Six conditions under which MPAs might not appear effective (when they are). ICES Journal of Marine Science, 75(3):1172–1174, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/3/1172/4098823.

Cotter:2015:DER

[CLdR^{+}15] John Cotter, William Lart, Nathan de Rozarieux, Al Kingston, Richard Caslake, Will Le Quesne, Simon Jennings, Alex Caveen, and Mary Brown. A development of ecological risk screening with an application to fisheries off SW England. ICES Journal of Marine Science, 72(3):1092–1104, March 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/3/1092/2835879.

Coro:2016:AFF

[CLMP16] Gianpaolo Coro, Scott Large, Chiara Magliozzi, and Pasquale Pagano. Analysing and forecasting fisheries time series: purse seine in Indian Ocean as a case study. ICES Journal of Marine Science, 73(10):2552–2571, November 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/10/2552/2647120.

Celic:2018:EEE

[CLS^{+}18] Igor Celić, Simone Libralato, Giuseppe Scarcella, Saša Raicevich, Bojan Marčeta, and Cosimo Solidoro. Ecological and economic effects of the landing obligation evaluated using a quantitative ecosystem approach: a Mediterranean case study. ICES Journal of Marine Science, 75(6):1992– 2003, November 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/6/1992/5045241.

Carey:2014:CCF

[CLT^{+}14] Michael P. Carey, Phillip S. Levin, Howard Townsend, Thomas J. Minello, Glen R. Sutton, Tessa B. Francis, Chris J. Harvey, Jodie E. Toft, Katie K. Arkema, Jennifer L. Burke, Choong-Ki Kim, Anne D. Guerry, Mark Plummer, Georgi Spiridonov, and Mary Ruckelshaus. Characterizing coastal foodwebs with qualitative links to bridge the gap between the theory and the practice of ecosystem-based management. ICES Journal of Marine Science, 71(3):713–724, March 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/3/713/630828.

Cormon:2014:SIB

[CLV^{+}14] Xochitl Cormon, Christophe Loots, Sandrine Vaz, Youen Vermard, and Paul Marchal. Spatial interactions between saithe (Pollachius virens) and hake (Merluccius merluccius) in the North Sea. ICES Journal of Marine Science, 71(6):1342–1355, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/71/
6/1342/2835591.

Carman:2016:DMS

[CMA^{+}16] V. González Carman, A. Mandiola, D. Alemany, M. Dassis, J. P. Seco Pon, L. Prosdocimi, A. Ponce de León, H. Mianzan, E. M. Acha, D. Rodríguez, M. Favero, and S. Copello. Distribution of megafaunal species in the Southwestern Atlantic: key ecological areas and opportunities for marine conservation. ICES Journal of Marine Science, 73(6):1579– 1588, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/6/1579/2458920.

Charrier:2012:ESP

[CMC^{+}12] Fabien Charrier, Virgile Mazel, Jean-Marie Caraguel, Yann Abdallah, L. Laëtitia Le Gurun, Antoine Legault, and Pascal Laffaille. Escapement of silver-phase European eels, Anguilla anguilla, determined from fishing activities in a Mediterranean lagoon (Or, France). ICES Journal of Marine Science, 69(1):30–33, January 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/1/30/670528.

Crook:2017:TBS

[CMD17] Kevin A. Crook, Emily Maxner, and Gail K. Davoren. Temperature-based spawning habitat selection by capelin (Mallotus villosus) in Newfoundland. ICES Journal of Marine Science, 74(6):1622–1629, July 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
6/1622/3066285.

Christie:2017:DRM

[CMG^{+}17] Mark R. Christie, Patrick G. Meirmans, Oscar E. Gaggiotti, Robert J. Toonen, and Crow White. Disentangling the relative merits and disadvantages of parentage analysis and assignment tests for inferring population connectivity. ICES Journal of Marine Science, 74(6):1749–1762, July 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/6/1749/3192370.

Carvalho:2011:SPB

[CMH^{+}11] Felipe C. Carvalho, Debra J. Murie, Fábio H. V. Hazin, Humberto G. Hazin, Bruno Leite-Mourato, and George H. Burgess. Spatial predictions of blue shark (Prionace glauca) catch rate and catch probability of juveniles in the Southwest Atlantic. ICES Journal of Marine Science, 68(5):890– 900, May 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/68/5/890/653584.

Capizzano:2016:EMD

[CMH^{+}16] Connor W. Capizzano, John W. Mandelman, William S. Hoffman, Micah J. Dean, Douglas R. Zemeckis, Hugues P. Benoît, Jeff Kneebone, Emily Jones, Marc J. Stettner, Nicholas J. Buchan, Joseph A. Langan, and James A. Sulikowski. Estimating and mitigating the discard mortality of Atlantic cod (Gadus morhua) in the Gulf of Maine recreational rod-and-reel fishery. ICES Journal of Marine Science, 73(9):2342–2355, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/9/2342/2198306.

Cardinale:2019:CLC

[CMH19] Massimiliano Cardinale, Stefano Mariani, and Joakim Hjelm. Comments on Local cod (Gadus morhua) revealed by egg surveys and population genetic analysis after longstanding depletion on the Swedish Skagerrak Coast by Svedäng et al. ICES Journal of Marine Science, 76(4): 1209–1211, July 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/76/4/1209/5509965.  See

[SBS^{+}19c].

Castillo-Manzano:2014:EAS

[CMLVGLP14] José I. Castillo-Manzano, Lourdes López-Valpuesta, Fernando Gonzalez-Laxe, and Diego J. Pedregal. An econometric analysis of the Spanish fresh fish market. ICES Journal of Marine Science, 71(3):628–635, March 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/3/628/639619.

Cardoso:2015:EAM

[CMM^{+}15] Inês Cardoso, Teresa Moura, Hugo Mendes, Cristina Silva, and Manuela Azevedo. An ecosystem approach to mixed fisheries: technical and biological interactions in the Portuguese multi-gear fleet. ICES Journal of Marine Science, 72(9):2618–2626, November 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/72/
9/2618/2457915.

Cadrin:2011:CCC

[CMP^{+}11] Steven X. Cadrin, Stefano Mariani, Christophe Pampoulie, Matthias Bernreuther, Anna Kristín Daníelsdóttir, Torild Johanssen, Lisa Kerr, Kjell Nedreaas, Jákup Reinert, Thorsteinn Sigurdsson, and Christoph Stransky. Countercomment on: Cadrin et al. (2010) “Population structure of beaked redfish, Sebastes mentella: evidence of divergence associated with different habitats. ICES Journal of Marine Science, 67: 1617–1630.”. ICES Journal of Marine Science, 68(10):2016–2018, November 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/10/2016/609741. See [CBD^{+}10a, MAP^{+}11].

Carpi:2017:IFL

[CMU^{+}17] Piera Carpi, Elisabetta B. Morello, Andres Uriarte, Monica Panfili, Beatriz Roel, Alberto Santojanni, Fortunata Donato, and Enrico Arneri. Impact of the fishery for late-larval European sardine (Sardina pilchardus) on the adult stock in the Adriatic Sea. ICES Journal of Marine Science, 74(3): 728–740, March 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/74/3/728/2742039.

Cropp:2014:CPP

[CN14] Roger Cropp and John Norbury. Comment on “The paradox of the ‘paradox of the plankton”’ by Record et al. ICES Journal of Marine Science, 71(2):293–295, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/2/293/790573.

Cooper:2016:JNR

[CN16] Daniel W. Cooper and Daniel G. Nichol. Juvenile northern rock sole (Lepidopsetta polyxystra) spatial distribution and abundance patterns in the eastern Bering Sea: spatially dependent production linked to temperature. ICES Journal of Marine Science, 73(4):1138–1146, March 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/4/1138/2459086.

Campbell:2011:SRT

[CNBK11] Neil Campbell, Francis Neat, Finlay Burns, and Phil Kunzlik. Species richness, taxonomic diversity, and taxonomic distinctness of the deep-water demersal fish community on the Northeast Atlantic continental slope (ICES Subdivision VIa). ICES Journal of Marine Science, 68(2):365–376, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/68/2/365/614751.

Cochrane:2017:IVF

[Coc17] Kevern L. Cochrane. An integrated view of fisheries: tunnelling between silos. ICES Journal of Marine Science, 74(3):625–634, March 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
3/625/2674179.

Contreras:2019:FEE

[COHdP19] Tabit Contreras, M. Pilar Olivar, P. Alexander Hulley, and M. Luz Fernández de Puelles. Feeding ecology of early life stages of mesopelagic fishes in the equatorial and tropical Atlantic. ICES Journal of Marine Science, 76(3):673– 689, May 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/3/673/5045242.

Chung:2013:IKF

[COL^{+}13] Ik Kyo Chung, Jung Hyun Oak, Jin Ae Lee, Jong Ahm Shin, Jong Gyu Kim, and Kwang-Seok Park. Installing kelp forests/seaweed beds for mitigation and adaptation against global warming: Korean project overview. ICES Journal of Marine Science, 70(5):1038–1044, September 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/5/1038/644026.

Carvalho:2011:RDS

[COM11] Fernando P. Carvalho, João M. Oliveira, and Margarida Malta. Radionuclides in deep-sea fish and other organisms from the North Atlantic Ocean. ICES Journal of Marine Science, 68(2):333–340, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/2/333/616694.

Cook:2019:SCS

[Coo19] R. M. Cook. Stock collapse or stock recovery? Contrasting perceptions of a depleted cod stock. ICES Journal of Marine Science, 76(4):787–793, July 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/4/787/5299883.

Cordue:2012:FIM

[Cor12] P. L. Cordue. Fishing intensity metrics for use in overfishing determination. ICES Journal of Marine Science, 69(4): 615–623, May 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/69/4/615/635007.

Corten:2013:RDN

[Cor13] Ad Corten. Recruitment depressions in North Sea herring. ICES Journal of Marine Science, 70(1):1–15, January 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/70/1/1/663444.

Calderwood:2015:EBC

[COR15] Julia Calderwood, Nessa E. O’Connor, and Dai Roberts. Effects of baited crab pots on cultivated mussel (Mytilus edulis) survival rates. ICES Journal of Marine Science, 72(6):1802–1810, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/72/6/1802/922927.

Carvalho:2011:SAB

[COS11] Fernando P. Carvalho, João M. Oliveira, and António M. M. Soares. Sediment accumulation and bioturbation rates in the deep Northeast Atlantic determined by radiometric techniques. ICES Journal of Marine Science, 68(3): 427–435, March 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/3/427/661886.

Certain:2015:BBL

[CP15] Grégoire Certain and Benjamin Planque. Biodiversity baseline for large marine ecosystems: an example from the Barents Sea. ICES Journal of Marine Science, 72(6):1756– 1768, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/6/1756/922746.

Coleman:2019:IFEa

[CPB19a] Matthew T. Coleman, Joanne S. Porter, and Michael C. Bell. Investigating fecundity and egg loss using a non-invasive method during brooding in European lobster (Homarus gammarus). ICES Journal of Marine Science, 76(6):1871–1881, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
6/1871/5430873.

Coleman:2019:IFEb

[CPB19b] Matthew T. Coleman, Joanne S. Porter, and Michael C. Bell. Investigating fecundity and egg loss using a non-invasive method during brooding in European lobster (Homarus gammarus). ICES Journal of Marine Science, 76(6):1934, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
6/1934/5498597.

Cross:2016:NOA

[CPLH16] Emma L. Cross, Lloyd S. Peck, Miles D. Lamare, and Elizabeth M. Harper. No ocean acidification effects on shell growth and repair in the New Zealand brachiopod Calloria inconspicua (Sowerby, 1846). ICES Journal of Marine Science, 73(3):920–926, February 2016.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/3/920/2457901.

Cooper:2017:EEP

[CPP17] Helen L. Cooper, Donald C. Potts, and Adina Paytan. Effects of elevated pCO_{2} on the survival, growth, and moulting of the Pacific krill species, Euphausia pacifica. ICES Journal of Marine Science, 74(4):1005–1012, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/4/1005/2907916.

Cheung:2013:HMP

[CPS13] William W. L. Cheung, Daniel Pauly, and Jorge L. Sarmiento. How to make progress in projecting climate change impacts. ICES Journal of Marine Science, 70(6): 1069–1074, September 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/70/6/1069/639516.

Castillo-Paez:2017:CTG

[CPSCCE^{+}17] Ana Castillo-Páez, Jonathan Sandoval-Castillo, David Corro-Espinosa, Javier Tovar-Ávila, María-Del-Pilar Blanco-Parra, Nancy C. Saavedra-Sotelo, Oscar Sosa-Nishizaki, Felipe Galván-Magaña, and Axayácatl Rocha-Olivares. Cutting through the Gordian knot: unravelling morphological, molecular, and biogeographical patterns in the genus Zapteryx (guitarfish) from the Mexican Pacific. ICES Journal of Marine Science, 74(6):1630–1638, July 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/6/1630/3061538.

Campbell:2010:SEC

[CPT^{+}10] Matthew D. Campbell, Reynaldo Patino, James Tolan, Richard Strauss, and Sandra L. Diamond. Sublethal effects of catch-and-release fishing: measuring capture stress, fish impairment, and predation risk using a condition index. ICES Journal of Marine Science, 67(3):513–521, April 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/3/513/732324.

Castano-Primo:2014:MAI

[CPVS14] Rocío Castaño-Primo, Frode Bendiksen Vikebø, and Svein Sundby. A model approach to identify the spawning grounds and describing the early life history of Northeast Arctic haddock (Melanogrammus aeglefinus). ICES Journal of Marine Science, 71(9):2505–2514, November 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/9/2505/2798169.

Chualain:2011:CAM

[CR11] Ciara Ní Chualáin and Martin Robinson. Comparison of assessment methods used to diagnose Hematodinium sp. infections in Cancer pagurus. ICES Journal of Marine Science, 68(3):454–462, March 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/68/3/454/660429.

Calderwood:2019:QED

[CR19] Julia Calderwood and David G. Reid. Quota exhaustion and discarding: how Ireland’s monthly quota system has a limited relationship with discarding patterns in the commercial fishing fleet. ICES Journal of Marine Science, 76(1):244–254, January 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
1/244/5168505.

Cabanellas-Reboredo:2014:WWW

[CRAM^{+}14] Miguel Cabanellas-Reboredo, Josep Alós, David March, Margarita Palmer, Gabriel Jordà, and Miquel Palmer. Where and when will they go fishing? Understanding fishing site and time choice in a recreational squid fishery. ICES Journal of Marine Science, 71(7):1760–1773, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/7/1760/662927.

Cabanellas-Reboredo:2012:EER

[CRAPMN12] Miguel Cabanellas-Reboredo, Josep Alós, Miquel Palmer, and Beatriz Morales-Nin. Environmental effects on recreational squid jigging fishery catches. ICES Journal of Marine Science, 69(10):1823–1830, December 2012.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/10/1823/624040.

Collard:2016:IOA

[CRC^{+}16] Marie Collard, Samuel P. S. Rastrick, Piero Calosi, Yoann Demolder, Jean Dille, Helen S. Findlay, Jason Michael Hall-Spencer, Marco Milazzo, Laure Moulin, Steve Widdicombe, Frank Dehairs, and Philippe Dubois. The impact of ocean acidification and warming on the skeletal mechanical properties of the sea urchin Paracentrotus lividus from laboratory and field observations. ICES Journal of Marine Science, 73(3):727–738, February 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/3/727/2457884.

Cameron:2019:PSS

[CRHM19] Luke W. J. Cameron, William K. Roche, Jonathan D. R. Houghton, and Paul J. Mensink. Population structure and spatial distribution of porbeagles (Lamna nasus) in Irish waters. ICES Journal of Marine Science, 76(6):1581–1590, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/6/1581/5475854.

Criddle:2012:AMF

[Cri12] Keith R. Criddle. Adaptation and maladaptation: factors that influence the resilience of four Alaskan fisheries governed by durable entitlements. ICES Journal of Marine Science, 69(7):1168–1179, September 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/7/1168/746054.

Conley:2015:CFP

[CS15] Keats R. Conley and Kelly R. Sutherland. Commercial fishers’ perceptions of jellyfish interference in the Northern California Current. ICES Journal of Marine Science, 72(5): 1565–1575, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/72/5/1565/772713.

Chan:2018:RLD

[CSB18] Kit Yu Karen Chan, Mary A. Sewell, and Maria Byrne. Revisiting the larval dispersal black box in the Anthropocene. ICES Journal of Marine Science, 75(6):1841–1848, November 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/6/1841/5070415.

Carreiro-Silva:2011:IPN

[CSBHS^{+}11] M. Carreiro-Silva, A. Braga-Henriques, I. Sampaio, V. de Matos, F. M. Porteiro, and O. Ocaña. Isozoanthus primnoidus, a new species of zoanthid (Cnidaria: Zoantharia) associated with the gorgonian Callogorgia verticillata (Cnidaria: Alcyonacea). ICES Journal of Marine Science, 68(2):408–415, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/68/
2/408/615426.

Chang:2013:MIE

[CSC^{+}13] Yi-Jay Chang, Chi-Lu Sun, Yong Chen, Su-Zan Yeh, Gerard DiNardo, and Nan-Jay Su. Modelling the impacts of environmental variation on the habitat suitability of swordfish, Xiphias gladius, in the equatorial Atlantic Ocean. ICES Journal of Marine Science, 70(5):1000–1012, September 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/70/5/1000/642761.

Cutter:2016:RSH

[CSD16] George R. Cutter, Jr., Kevin L. Stierhoff, and David A. Demer. Remote sensing of habitat characteristics using echo metrics and image-based seabed classes. ICES Journal of Marine Science, 73(8):1965–1974, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/8/1965/2198330.

Cleary:2017:FCG

[CSF^{+}17] Alison C. Cleary, Janne E. Søreide, Daniela Freese, Barbara Niehoff, and Tove M. Gabrielsen. Feeding by Calanus glacialis in a high Arctic fjord: potential seasonal importance of alternative prey. ICES Journal of Marine Science, 74(7):1937–1946, September 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/7/1937/4036290.

Clements:2010:OSS

[CSFS10] Annika J. Clements, James A. Strong, Clare Flanagan, and Matthew Service. Objective stratification and samplingeffort allocation of ground-truthing in benthic-mapping surveys. ICES Journal of Marine Science, 67(4):628–637, May 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/4/628/681787.

Cyronak:2016:OMW

[CSJ16a] Tyler Cyronak, Kai G. Schulz, and Paul L. Jokiel. The Omega myth: what really drives lower calcification rates in an acidifying ocean. ICES Journal of Marine Science, 73(3):558–562, February 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
3/558/2457920.

Cyronak:2016:RWA

[CSJ16b] Tyler Cyronak, Kai G. Schulz, and Paul L. Jokiel. Response to Waldbusser et al. (2016): “Calcium carbonate saturation state: on myths and this or that stories”. ICES Journal of Marine Science, 73(3):569–571, February 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/3/569/2458969.

Churnside:2011:ASF

[CSR11] James H. Churnside, Alexei F. Sharov, and Ronald A. Richter. Aerial surveys of fish in estuaries: a case study in Chesapeake Bay. ICES Journal of Marine Science, 68(1): 239–244, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/1/239/628435.

Canas:2012:UOS

[CSS^{+}12] Lucía Cañás, Christoph Stransky, Jürgen Schlickeisen, M. Paz Sampedro, and A. Celso Fariña. Use of the otolith shape analysis in stock identification of anglerfish (Lophius piscatorius) in the Northeast Atlantic. ICES Journal of Marine Science, 69(2):250–256, March 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/2/250/705841.

Calvert:2015:ESU

[CSS^{+}15] Jay Calvert, James Asa Strong, Matthew Service, Chris McGonigle, and Rory Quinn. An evaluation of supervised and unsupervised classification techniques for marine benthic habitat mapping using multibeam echosounder data. ICES Journal of Marine Science, 72(5):1498–1513, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/5/1498/761886.

Crawford:2019:FHE

[CST^{+}19] Robert J. M. Crawford, William J. Sydeman, Sarah Ann Thompson, Richard B. Sherley, and Azwianewi B. Makhado. Food habits of an endangered seabird indicate recent poor forage fish availability off western South Africa. ICES Journal of Marine Science, 76(5):1344–1352, 09- 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/5/1344/5491592.

Cobain:2019:ITS

[CSTJ19] Matthew R. D. Cobain, Will Steward, Clive N. Trueman, and Antony Jensen. Individual trophic specialization in juvenile European seabass: implications for the management of a commercially important species. ICES Journal of Marine Science, 76(6):1784–1793, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/6/1784/5445396.

Coll:2010:RER

[CSY^{+}10] Marta Coll, Lynne J. Shannon, Dawit Yemane, Jason S. Link, Henn Ojaveer, Sergio Neira, Didier Jouffre, Pierre Labrosse, Johanna J. Heymans, Elizabeth A. Fulton, and Yunne-Jai Shin. Ranking the ecological relative status of exploited marine ecosystems. ICES Journal of Marine Science, 67(4):769–786, May 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/67/4/769/678259.

Cao:2014:ISS

[CTC14] Jie Cao, Samuel B. Truesdell, and Yong Chen. Impacts of seasonal stock mixing on the assessment of Atlantic cod in the Gulf of Maine. ICES Journal of Marine Science, 71(6):1443–1457, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/6/1443/2835579.

Casini:2019:STDa

[CTH^{+}19a] Michele Casini, Huidong Tian, Martin Hansson, Włodzimierz Grygiel, Guntars Strods, Romas Statkus, Elor Sepp, Tomas Gröhsler, Alessandro Orio, and Niklas Larson. Spatiotemporal dynamics and behavioural ecology of a “demersal” fish population as detected using research survey pelagic trawl catches: the Eastern Baltic Sea cod (Gadus morhua). ICES Journal of Marine Science, 76(6):1591– 1600, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/76/6/1591/5364299.

Casini:2019:STDb

[CTH^{+}19b] Michele Casini, Huidong Tian, Martin Hansson, Włodzimierz Grygiel, Guntars Strods, Romas Statkus, Elor Sepp, Tomas Gröhsler, Alessandro Orio, and Niklas Larson. Spatiotemporal dynamics and behavioural ecology of a “demersal” fish population as detected using research survey pelagic trawl catches: the Eastern Baltic Sea cod (Gadus morhua). ICES Journal of Marine Science, 76(6):1931, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/6/1931/5453297.

Cheke:2014:PPP

[CTT14] Robert A. Cheke, Sanyi Tang, and Jamie A. Tratalos. Predator–prey population models of migrant insects with phase change. ICES Journal of Marine Science, 71(8):2221–2230, October 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/71/
8/2221/745321.

Cury:2019:ON

[Cur19] Philippe Cury. Obstinate nature. ICES Journal of Marine Science, 76(2):384–391, March 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
2/384/5258065.

Cox:2011:STV

[CWRB11] Martin J. Cox, Jonathan L. Watkins, Keith Reid, and Andrew S. Brierley. Spatial and temporal variability in the structure of aggregations of Antarctic krill (Euphausia superba) around South Georgia, 1997–1999. ICES Journal of Marine Science, 68(3):489–498, March 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/3/489/660620.

Campanati:2016:CEL

[CYLT16] Camilla Campanati, Stella Yip, Ackley Lane, and Vengatesen Thiyagarajan. Combined effects of low pH and low oxygen on the early-life stages of the barnacle Balanus amphitrite. ICES Journal of Marine Science, 73(3):791–802, February 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/3/791/2458960.

Candia-Zulbaran:2015:CSL

[CZBFLÁ^{+}15] Rebeca I. Candia-Zulbarán, Patricia Briones-Fourzán, Enrique Lozano-Álvarez, Cecilia Barradas-Ortiz, and Fernando Negrete-Soto. Caribbean spiny lobsters equally avoid dead and clinically PaV1-infected conspecifics. ICES Journal of Marine Science, 72(S1):S164–S169, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/72/suppl_1/i164/617804.

Dutto:2012:ISE

[DAB^{+}12] M. Sofía Dutto, M. Celeste López Abbate, Florencia Biancalana, Anabela A. Berasategui, and Mónica S. Hoffmeyer. The impact of sewage on environmental quality and the mesozooplankton community in a highly eutrophic estuary in Argentina. ICES Journal of Marine Science, 69(3):399– 409, May 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/69/3/399/600660.

Santamaria-del-Angel:2011:RSF

[dÁMNGS^{+}11] Eduardo Santamaría del Ángel, Roberto Millán-Núñez, Adriana González-Silvera, Mariana Callejas-Jiménez, Ramón Cajal-Medrano, and Manuel S. Galindo-Bect. The response of shrimp fisheries to climate variability off Baja California, México. ICES Journal of Marine Science, 68(4):766–772, March 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/68/4/766/648701.

Dankel:2012:AUU

[DAP^{+}12] Dorothy J. Dankel, Robert Aps, Gurpreet Padda, Christine Röckmann, Jeroen P. van der Sluijs, Douglas C. Wilson, and Poul Degnbol. Advice under uncertainty in the marine system. ICES Journal of Marine Science, 69(1):3–7, January 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/69/1/3/672262.

Davison:2011:SGM

[Dav11] Peter Davison. The specific gravity of mesopelagic fish from the northeastern Pacific Ocean and its implications for acoustic backscatter. ICES Journal of Marine Science, 68(10):2064–2074, November 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/68/
10/2064/610449.

Day:2018:CA

[Day18a] Jon C. Day. Counterpoint to Agardy. ICES Journal of Marine Science, 75(3):1186, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/75/
3/1186/4098829.

Day:2018:DFW

[Day18b] Jon C. Day. Day’s final word. ICES Journal of Marine Science, 75(3):1192, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/75/3/1192/4098834.

Day:2018:HEM

[Day18c] Jon C. Day. How effective is the management of the Great Barrier Reef? ICES Journal of Marine Science, 75(3): 1188–1190, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/3/1188/4098826.

Dekker:2016:CBW

[DB16] Willem Dekker and Laurent Beaulaton. Climbing back up what slippery slope? Dynamics of the European eel stock and its management in historical perspective. ICES Journal of Marine Science, 73(1):5–13, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/1/5/2458735.

DeRobertis:2019:ALA

[DBA^{+}19] Alex De Robertis, Christopher Bassett, Lars Nonboe Andersen, Ivar Wangen, Scott Furnish, and Michael Levine. Amplifier linearity accounts for discrepancies in echo-integration measurements from two widely used echosounders. ICES Journal of Marine Science, 76(6): 1882–1892, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/76/6/1882/5430872.

Dempson:2010:SIA

[DBDP10] J. Brian Dempson, Victoria A. Braithwaite, Denis Doherty, and Michael Power. Stable isotope analysis of marine feeding signatures of Atlantic salmon in the North Atlantic. ICES Journal of Marine Science, 67(1):52–61, January 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/67/1/52/595843.

Doray:2016:MCT

[DBL^{+}16] Mathieu Doray, Laurent Berger, Naig Le Bouffant, Jean Yves Coail, Jean Philippe Vacherot, Xavier de La Bernardie, Pierre Morinière, Elisabeth Lys, Romain Schwab, and Pierre Petitgas. A method for controlled target strength measurements of pelagic fish, with application to European anchovy (Engraulis encrasicolus). ICES Journal of Marine Science, 73(8):1987–1997, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/8/1987/2198469.

Deroba:2015:STR

[DBM^{+}15] J. J. Deroba, D. S. Butterworth, R. D. Methot, Jr., J. A. A. De Oliveira, C. Fernandez, A. Nielsen, S. X. Cadrin, M. Dickey-Collas, C. M. Legault, J. Ianelli, J. L. Valero, C. L. Needle, J. M. O’Malley, Y-J. Chang, G. G. Thompson, C. Canales, D. P. Swain, D. C. M. Miller, N. T. Hintzen, M. Bertignac, L. Ibaibarriaga, A. Silva, A. Murta, L. T. Kell, C. L. de Moor, A. M. Parma, C. M. Dichmont, V. R. Restrepo, Y. Ye, E. Jardim, P. D. Spencer, D. H. Hanselman, J. Blaylock, M. Mood, and P.-J. F. Hulson. Simulation testing the robustness of stock assessment models to error: some results from the ICES strategic initiative on stock assessment methods. ICES Journal of Marine Science, 72(1):19–30, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/72/1/19/815567.

Diehl:2015:EDC

[DBS^{+}15] Sebastian Diehl, Stella A. Berger, Quentin Soissons, Darren P. Giling, and Herwig Stibor. An experimental demonstration of the critical depth principle. ICES Journal of Marine Science, 72(6):2051–2060, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/6/2051/919890.

deBraux:2014:OII

[dBWMD^{+}14] Emmanuel de Braux, Fletcher Warren-Myers, Tim Dempster, Per Gunnar Fjelldal, Tom Hansen, and Stephen E. Swearer. Osmotic induction improves batch marking of larval fish otoliths with enriched stable isotopes. ICES Journal of Marine Science, 71(9):2530–2538, November 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/9/2530/2798171.

Dickey-Collas:2014:WCN

[DC14] Mark Dickey-Collas. Why the complex nature of integrated ecosystem assessments requires a flexible and adaptive approach. ICES Journal of Marine Science, 71(5):1174–1182, July 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/71/5/1174/647715.

Dickey-Collas:2014:EBM

[DCER^{+}14] M. Dickey-Collas, G. H. Engelhard, A. Rindorf, K. Raab, S. Smout, G. Aarts, M. van Deurs, T. Brunel, A. Hoff, R. A. M. Lauerburg, S. Garthe, K. Haste Andersen, F. Scott, T. van Kooten, D. Beare, and M. A. Peck. Ecosystem-based management objectives for the North Sea: riding the forage fish rollercoaster. ICES Journal of Marine Science, 71(1):128–142, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/1/128/642315.

DaRocha:2010:EBO

[DCG10] José-María Da Rocha, Santiago Cerviño, and María-José Gutiérrez. An endogenous bioeconomic optimization algorithm to evaluate recovery plans: an application to southern hake. ICES Journal of Marine Science, 67(9):1957– 1962, December 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/67/9/1957/620414.

Diaz-Cabrera:2012:MDS

[DCHMHQ12] Ernesto Díaz-Cabrera, Eduardo Hernández-Miranda, Cristián E. Hernández, and Renato A. Quiñones. Mesoscale β diversity and spatial nestedness of crustacean larvae in the coastal zone off central southern Chile: population and community implications. ICES Journal of Marine Science, 69(3):429– 438, May 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/69/3/429/599379.

Dickey-Collas:2015:QPT

[DCHN^{+}15] M. Dickey-Collas, N. T. Hintzen, R. D. M. Nash, P-J. Schön, and M. R. Payne. Quirky patterns in timeseries of estimates of recruitment could be artefacts. ICES Journal of Marine Science, 72(1):111–116, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/1/111/821856.

Dolan:2019:ESH

[DCMC19] John R. Dolan, Maria Ciobanu, Sophie Marro, and Laurent Coppola. An exploratory study of heterotrophic protists of the mesopelagic Mediterranean Sea. ICES Journal of Marine Science, 76(3):616–625, May 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/3/616/4717462.

Dickey-Collas:2017:PHE

[DCMGB^{+}17] Mark Dickey-Collas, Abigail McQuatters-Gollop, Eileen Bresnan, Alexandra C. Kraberg, John P. Manderson, Richard D. M. Nash, Saskia A. Otto, Anne F. Sell, Jacqueline F. Tweddle, and Verena M. Trenkel. Pelagic habitat: exploring the concept of good environmental status. ICES Journal of Marine Science, 74(9):2333–2341, November 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/9/2333/4085765.

Dickey-Collas:2010:LLS

[DCNB^{+}10] Mark Dickey-Collas, Richard D. M. Nash, Thomas Brunel, Cindy J. G. van Damme, C. Tara Marshall, Mark R. Payne, Ad Corten, Audrey J. Geffen, Myron A. Peck, Emma M. C. Hatfield, Niels T. Hintzen, Katja Enberg, Laurence T. Kell, and E. John Simmonds. Lessons learned from stock collapse and recovery of North Sea herring: a review. ICES Journal of Marine Science, 67(9):1875–1886, December 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/9/1875/618208.

Dickey-Collas:2014:HWM

[DCPTN14] Mark Dickey-Collas, Mark R. Payne, Verena M. Trenkel, and Richard D. M. Nash. Hazard warning: model misuse ahead. ICES Journal of Marine Science, 71(8):2300–2306, October 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/71/8/2300/750508.

Denman:2011:PIF

[DCS^{+}11] Kenneth Denman, James R. Christian, Nadja Steiner, Hans-Otto Pörtner, and Yukihiro Nojiri. Potential impacts of future ocean acidification on marine ecosystems and fisheries: current knowledge and recommendations for future research. ICES Journal of Marine Science, 68(6): 1019–1029, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/6/1019/713847.

delaCruz:2014:CBL

[dCVB14] Dexter W. dela Cruz, Ronald D. Villanueva, and Maria Vanessa B. Baria. Community-based, low-tech method of restoring a lost thicket of Acropora corals. ICES Journal of Marine Science, 71(7):1866–1875, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/7/1866/664250.

Dimitriadis:2015:RAT

[DCVC15] Caterina Dimitriadis, Alvar Carranza, Raúl Vilela, and Margarida Casadevall. A rapid assessment of trends in the multispecies small-scale fishery of Palamós (Catalonia, Spain). ICES Journal of Marine Science, 72(9):2638–2649, November 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/9/2638/2458749.

deCastro:2013:ESD

[dCWMH13] Carlos de Castro, Peter J. Wright, Colin P. Millar, and Steven J. Holmes. Evidence for substock dynamics within whiting (Merlangius merlangus) management regions. ICES Journal of Marine Science, 70(6):1118– 1127, September 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/70/6/1118/633552.

Dvoretsky:2010:EAI

[DD10] Alexander G. Dvoretsky and Vladimir G. Dvoretsky. Epifauna associated with an introduced crab in the Barents Sea: a 5-year study. ICES Journal of Marine Science, 67(2):204–214, March 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/67/2/204/694306.

Dvoretsky:2013:PDI

[DD13] Alexander G. Dvoretsky and Vladimir G. Dvoretsky. Population dynamics of the invasive lithodid crab, Paralithodes camtschaticus, in a typical bay of the Barents Sea. ICES Journal of Marine Science, 70(6):1255–1262, September 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/6/1255/634857.

Depestele:2019:CMD

[DDE^{+}19] Jochen Depestele, Koen Degrendele, Moosa Esmaeili, Ana Ivanović, Silke Kröger, Finbarr G. O’Neill, Ruth Parker, Hans Polet, Marc Roche, Lorna R. Teal, Bart Vanelslander, and Adriaan D. Rijnsdorp. Comparison of mechanical disturbance in soft sediments due to tickler-chain SumWing trawl vs. electro-fitted PulseWing trawl. ICES Journal of Marine Science, 76(1):312–329, January 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/1/312/5104431.

Doo:2012:IOA

[DDF^{+}12] Steve S. Doo, Symon A. Dworjanyn, Shawna A. Foo, Natalie A. Soars, and Maria Byrne. Impacts of ocean acidification on development of the meroplanktonic larval stage of the sea urchin Centrostephanus rodgersii. ICES Journal of Marine Science, 69(3):460–464, May 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/3/460/597243.

DeOliveira:2010:LSA

[DDR10] José A. A. De Oliveira, Chris D. Darby, and Beatriz A. Roel. A linked separable–ADAPT VPA assessment model for western horse mackerel (Trachurus trachurus), accounting for realized fecundity as a function of fish weight. ICES Journal of Marine Science, 67(5):916–930, July 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/5/916/608291.

DeOliveira:2013:IDD

[DED13] José A. A. De Oliveira, James R. Ellis, and Helen Dobby. Incorporating density dependence in pup production in a stock assessment of NE Atlantic spurdog Squalus acanthias. ICES Journal of Marine Science, 70(7):1341–1353, November 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/70/7/1341/607111.

Dean:2019:RIS

[DEH^{+}19] Micah J. Dean, Scott P. Elzey, William S. Hoffman, Nicholas C. Buchan, and Jonathan H. Grabowski. The relative importance of sub-populations to the Gulf of Maine stock of Atlantic cod. ICES Journal of Marine Science, 76(6):1626–1640, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
6/1626/5490626.

Dekker:2016:MES

[Dek16] Willem Dekker. Management of the eel is slipping through our hands! Distribute control and orchestrate national protection. ICES Journal of Marine Science, 73(10):2442– 2452, November 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/73/10/2442/2647096.

Deroba:2018:SVS

[Der18] Jonathan J. Deroba. Sources of variation in stomach contents of predators of Atlantic herring in the Northwest Atlantic during 1973–2014. ICES Journal of Marine Science, 75(4):1439–1450, July 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/75/
4/1439/4866347.

DiFranco:2010:CRS

[DFB^{+}10] Antonio Di Franco, Gaetano Ferruzza, Pasquale Baiata, Renato Chemello, and Marco Milazzo. Can recreational scuba divers alter natural gross sedimentation rate? A case study from a Mediterranean deep cave. ICES Journal of Marine Science, 67(5):871–874, July 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/5/871/610287.

Dichmont:2017:DRD

[DFG^{+}17] Catherine M. Dichmont, Elizabeth A. Fulton, Rebecca Gorton, Miriana Sporcic, L. Richard Little, André E. Punt, Natalie Dowling, Malcolm Haddon, Neil Klaer, and David C. Smith. From data rich to data-limited harvest strategies — does more data mean better management? ICES Journal of Marine Science, 74(3):670–686, March 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/3/670/2741995.

Diserud:2019:EFA

[DFS^{+}19] O. H. Diserud, P. Fiske, H. Sægrov, K. Urdal, T. Aronsen, H. Lo, B. T. Barlaup, E. Niemelä, P. Orell, J. Erkinaro, R. A. Lund, F. Økland, G. M. Østborg, L. P. Hansen, and K. Hindar. Escaped farmed Atlantic salmon in Norwegian rivers during 1989–2013. ICES Journal of Marine Science, 76(4):1140–1150, July 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
4/1140/5289588.

DaRocha:2011:LLT

[DG11] José-María Da Rocha and María-José Gutiérrez. Lessons from the long-term management plan for northern hake: could the economic assessment have accepted it? ICES Journal of Marine Science, 68(9):1937–1941, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/9/1937/665637.

Dalpadado:2017:SIZ

[DG17] Padmini Dalpadado and Astthor Gislason. Sixth International Zooplankton Production Symposium: new challenges in a changing ocean. ICES Journal of Marine Science, 74(7):1839–1845, September 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
7/1839/4210358.

Dorner:2015:CDC

[DGB^{+}15] Hendrik Dörner, Norman Graham, Gabriella Bianchi, Åsmund Bjordal, Marco Frederiksen, William A. Karp, Steven J. Kennelly, Jann Thorsten Martinsohn, Kimberly Murray, Martin Pastoors, and Njård Håkon Gudbrandsen. From cooperative data collection to full collaboration and co-management: a synthesis of the 2014 ICES symposium on fishery-dependent information. ICES Journal of Marine Science, 72(4):1133–1139, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/4/1133/801305.

DaRocha:2012:RPB

[DGC12] José-María Da Rocha, María-José Gutiérrez, and Santiago Cerviño. Reference points based on dynamic optimization: a versatile algorithm for mixed-fishery management with bioeconomic age-structured models. ICES Journal of Marine Science, 69(4):660–669, May 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/4/660/631933.

deGraaf:2011:SRF

[dGGW^{+}11] Gertjan J. de Graaf, Richard J. R. Grainger, Lena Westlund, Rolf Willmann, David Mills, Kieran Kelleher, and Kwame Koranteng. The status of routine fishery data collection in Southeast Asia, central America, the South Pacific, and West Africa, with special reference to small-scale fisheries. ICES Journal of Marine Science, 68(8):1743– 1750, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/8/1743/748339.

DePiper:2017:OIE

[DGL^{+}17] Geret S. DePiper, Sarah K. Gaichas, Sean M. Lucey, Patricia Pinto da Silva, M. Robin Anderson, Heather Breeze, Alida Bundy, Patricia M. Clay, Gavin Fay, Robert J. Gamble, Robert S. Gregory, Paula S. Fratantoni, Catherine L. Johnson, Mariano Koen-Alonso, Kristin M. Kleisner, Julia Olson, Charles T. Perretti, Pierre Pepin, Fred Phelan, Vincent S. Saba, Laurel A. Smith, Jamie C. Tam, Nadine D. Templeman, and Robert P. Wildermuth. Operationalizing integrated ecosystem assessments within a multidisciplinary team: lessons learned from a worked example. ICES Journal of Marine Science, 74(8):2076–2086, September 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/8/2076/3094701.

Diaz-Gil:2015:OFA

[DGPC^{+}15] Carlos Díaz-Gil, Miquel Palmer, Ignacio A. Catalán, Josep Alós, Lee A. Fuiman, Elena García, María del Mar Gil, Amalia Grau, Andrew Kang, Rommel H. Maneja, John A. Mohan, Bernat Morro, Jason J. Schaffler, Lucie Buttay, Inmaculada Riera-Batle, Borja Tolosa, and Beatriz Morales-Nin. Otolith fluctuating asymmetry: a misconception of its biological relevance? ICES Journal of Marine Science, 72(7):2079–2089, October 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/7/2079/2457871.

Daan:2011:AWF

[DGPR11] Niels Daan, Henrik Gislason, John G. Pope, and Jake C. Rice. Apocalypse in world fisheries? The reports of their death are greatly exaggerated. ICES Journal of Marine Science, 68(7):1375–1378, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/68/
7/1375/656650.

DeRobertis:2013:FAR

[DH13] Alex De Robertis and Nils Olav Handegard. Fish avoidance of research vessels and the efficacy of noise-reduced vessels: a review. ICES Journal of Marine Science, 70(1):34–45, January 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/70/1/34/661416.

Drinkwater:2012:CSC

[DHAH12] Kenneth F. Drinkwater, George L. Hunt, Jr., Olafur S. Astthorsson, and Erica J. H. Head. Comparative studies of climate effects on polar and subpolar ocean ecosystems, progress in observation and prediction: an introduction. ICES Journal of Marine Science, 69(7):1120–1122, September 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/69/7/1120/752964.

Dutertre:2013:URB

[DHCE13] Mickaël Dutertre, Dominique Hamon, Claire Chevalier, and Axel Ehrhold. The use of the relationships between environmental factors and benthic macrofaunal distribution in the establishment of a baseline for coastal management. ICES Journal of Marine Science, 70(2):294–308, March 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/2/294/794591.

deHaan:2016:PTF

[dHFF^{+}16] D. de Haan, J. E. Fosseidengen, P. G. Fjelldal, D. Burggraaf, and A. D. Rijnsdorp. Pulse trawl fishing: characteristics of the electrical stimulation and the effect on behaviour and injuries of Atlantic cod (Gadus morhua). ICES Journal of Marine Science, 73(6):1557–1569, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
6/1557/2458917.

Daewel:2014:PCZ

[DHH^{+}14] Ute Daewel, Solfrid Sætre Hjøllo, Martin Huret, Rubao Ji, Marie Maar, Susa Niiranen, Morgane Travers-Trolet, Myron A. Peck, and Karen E. van de Wolfshaar. Predation control of zooplankton dynamics: a review of observations and models. ICES Journal of Marine Science, 71(2):254– 271, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/71/2/254/781831.

Derouiche:2016:EEE

[DHKE16] Emna Derouiche, Besma Hizem Habbechi, Med. Mejdeddine Kraïem, and Pierre Elie. Estimates of escapement, exploitation rate, and number of downstream migrating European eels Anguilla anguilla in Ichkeul Lake (northern Tunisia). ICES Journal of Marine Science, 73(1):142–149, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/1/142/2458902.

DiBacco:2012:BWT

[DHNL12] Claudio DiBacco, Donald B. Humphrey, Leslie E. Nasmith, and Colin D. Levings. Ballast water transport of non-indigenous zooplankton to Canadian ports. ICES Journal of Marine Science, 69(3):483–491, May 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/3/483/597336.

Daley:2019:SBE

[DHS19] Ross K. Daley, Alistair J. Hobday, and Jayson M. Semmens. Simulation-based evaluation of reserve network performance for Centrophorus zeehaani (Centrophoridae): a protected deep-sea gulper shark. ICES Journal of Marine Science, 76(7):2318–2328, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/7/2318/5506076.

Dean:2014:FSD

[DHZA14] Micah J. Dean, William S. Hoffman, Douglas R. Zemeckis, and Michael P. Armstrong. Fine-scale diel and gender-based patterns in behaviour of Atlantic cod (Gadus morhua) on a spawning ground in the Western Gulf of Maine. ICES Journal of Marine Science, 71(6):1474– 1489, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/71/6/1474/624423.

Dissanayake:2011:SEE

[DI11] Awantha Dissanayake and Atsushi Ishimatsu. Synergistic effects of elevated CO_{2} and temperature on the metabolic scope and activity in a shallow-water coastal decapod (Metapenaeus joyneri; Crustacea: Penaeidae). ICES Journal of Marine Science, 68(6):1147–1154, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/6/1147/696833.

Depestele:2016:MAP

[DID^{+}16] Jochen Depestele, Ana Ivanović, Koen Degrendele, Moosa Esmaeili, Hans Polet, Marc Roche, Keith Summerbell, Lorna R. Teal, Bart Vanelslander, and Finbarr G. O’Neill. Measuring and assessing the physical impact of beam trawling. ICES Journal of Marine Science, 73(suppl 1):S15–S26, January 2016. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/73/suppl_1/i15/2573918.

Diggles:2019:RSS

[Dig19] B. K. Diggles. Review of some scientific issues related to crustacean welfare. ICES Journal of Marine Science, 76(1): 66–81, January 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/76/1/66/5037899.

Dill:2017:BEM

[Dil17] Lawrence M. Dill. Behavioural ecology and marine conservation: a bridge over troubled water? ICES Journal of Marine Science, 74(6):1514–1521, July 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/6/1514/3094697.

Dalpadado:2012:CEB

[DIS^{+}12] Padmini Dalpadado, Randi B. Ingvaldsen, Leif Christian Stige, Bjarte Bogstad, Tor Knutsen, Geir Ottersen, and Bjørnar Ellertsen. Climate effects on Barents Sea ecosystem dynamics. ICES Journal of Marine Science, 69(7):1303– 1316, September 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/69/7/1303/742743.

deJong:2015:RBM

[dJBLH15] Maarten F. de Jong, Martin J. Baptist, Han J. Lindeboom, and Piet Hoekstra. Relationships between macrozoobenthos and habitat characteristics in an intensively used area of the Dutch coastal zone. ICES Journal of Marine Science, 72(8):2409–2422, September 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/8/2409/2457859.

Dolgov:2010:TEB

[DJHO10] Andrey V. Dolgov, Edda Johannesen, Mikko Heino, and Erik Olsen. Trophic ecology of blue whiting in the Barents Sea. ICES Journal of Marine Science, 67(3):483–493, April 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/67/3/483/731986.

Davies:2015:EPD

[DJW^{+}15] Charlotte E. Davies, Andrew F. Johnson, Emma C. Wootton, Spencer J. Greenwood, K. Fraser Clark, Claire L. Vogan, and Andrew F. Rowley. Effects of population density and body size on disease ecology of the European lobster in a temperate marine conservation zone. ICES Journal of Marine Science, 72(S1):S128–S138, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/72/
suppl_1/i128/616050.

Drinkwater:2018:SER

[DK18] Kenneth F. Drinkwater and Trond Kristiansen. A synthesis of the ecosystem responses to the late 20th century cold period in the northern North Atlantic. ICES Journal of Marine Science, 75(7):2325–2341, December 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/7/2325/5049598.

Drinkwater:2019:RTM

[DK19] Kenneth F. Drinkwater and Trond Kristiansen. Reply to “Towards a more balanced assessment of the dynamics of North Atlantic ecosystems — a comment on Drinkwater and Kristiansen (2018)”. ICES Journal of Marine Science, 76(7):2495–2499, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/7/2495/5554625.

Davias:2014:EFI

[DKB14] Lori A. Davias, Matthew S. Kornis, and Denise L. Breitburg. Environmental factors influencing δ ^{13}C and δ ^{15}N in three Chesapeake Bay fishes. ICES Journal of Marine Science, 71(3):689–702, March 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/71/3/689/634123.

DellApa:2012:TFE

[DKC12] Andrea Dell’Apa, David G. Kimmel, and Simona Clò. Trends of fish and elasmobranch landings in Italy: associated management implications. ICES Journal of Marine Science, 69(6):1045–1052, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/69/
6/1045/619719.

Drakou:2017:SFM

[DKC^{+}17] Evangelia G. Drakou, Charlène Kermagoret, Adrien Comte, Brita Trapman, and Jake C. Rice. Shaping the future of marine socio-ecological systems research: when early-career researchers meet the seniors. ICES Journal of Marine Science, 74(7):1957–1964, September 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/7/1957/3003307.

Davison:2015:ABE

[DKK15] Peter C. Davison, J. Anthony Koslow, and Rudy J. Kloser. Acoustic biomass estimation of mesopelagic fish: backscattering from individuals, populations, and communities. ICES Journal of Marine Science, 72(5):1413–1424, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/5/1413/777115.

Dutra:2016:TCC

[DKPM16] Elizabeth Dutra, Marguerite Koch, Katherine Peach, and Carrie Manfrino. Tropical crustose coralline algal individual and community responses to elevated pCO_{2} under high and low irradiance. ICES Journal of Marine Science, 73(3):803– 813, February 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/73/3/803/2458928.

Davie:2011:ECI

[DL11] Sarah Davie and Colm Lordan. Examining changes in Irish fishing practices in response to the cod long-term plan. ICES Journal of Marine Science, 68(8):1638–1646, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/68/8/1638/748024.

Deibel:2012:RLC

[DL12] Don Deibel and Ben Lowen. A review of the life cycles and life-history adaptations of pelagic tunicates to environmental conditions. ICES Journal of Marine Science, 69(3): 358–369, May 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/69/3/358/598151.

deLestang:2014:OMD

[dL14] Simon de Lestang. The orientation and migratory dynamics of the western rock lobster, Panulirus cygnus, in Western Australia. ICES Journal of Marine Science, 71(5):1052– 1063, July 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/71/5/1052/645017.

deLestang:2015:WCS

[dLCF^{+}15] Simon de Lestang, Nick Caputi, Ming Feng, Ainslie Denham, James Penn, Dirk Slawinski, Alan Pearce, and Jason How. What caused seven consecutive years of low puerulus settlement in the western rock lobster fishery of Western Australia? ICES Journal of Marine Science, 72(S1):S49– S58, July 2015. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/72/suppl_1/i49/613514.

Degerman:2012:CMT

[DLL12] Erik Degerman, Kjell Leonardsson, and Hans Lundqvist. Coastal migrations, temporary use of neighbouring rivers, and growth of sea trout (Salmo trutta) from nine northern Baltic Sea rivers. ICES Journal of Marine Science, 69(6): 971–980, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/69/6/971/620855.

delaPrada:2015:QMR

[dlPG15] Amelia de la Prada and Manuel González. Quantifying mesh resistance to opening of netting panels: experimental method, regression models, and parameter estimation strategies. ICES Journal of Marine Science, 72(2):697–707, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/2/697/2801333.

DeRobertis:2019:LTM

[DLSJ^{+}19] Alex De Robertis, Noah Lawrence-Slavas, Richard Jenkins, Ivar Wangen, Calvin W. Mordy, Christian Meinig, Mike Levine, Dave Peacock, and Heather Tabisola. Longterm measurements of fish backscatter from saildrone unmanned surface vehicles and comparison with observations from a noise-reduced research vessel. ICES Journal of Marine Science, 76(7):2459–2470, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/7/2459/5529308.

delaTorriente:2010:SAS

[dlTQMUE10] Ana de la Torriente, Renato A. Quiñones, Diego A. Miranda-Urbina, and Fidel Echevarría. South American sea lion and spiny dogfish predation on artisanal catches of southern hake in fjords of Chilean Patagonia. ICES Journal of Marine Science, 67(2):294–303, March 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/2/294/692438.

Damasio:2016:SMF

[dMADLP^{+}16] Ludmila de Melo Alves Damasio, Priscila Fabiana Macedo Lopes, Maria Grazia Pennino, Adriana Rosa Carvalho, and Ussif Rashid Sumaila. Size matters: fishing less and yielding more in smaller-scale fisheries. ICES Journal of Marine Science, 73(6):1494–1502, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
6/1494/2458899.

Drouineau:2010:LSS

[DMBD10] Hilaire Drouineau, Stéphanie Mahévas, Michel Bertignac, and Daniel Duplisea. A length-structured spatially explicit model for estimating hake growth and migration rates. ICES Journal of Marine Science, 67(8):1697–1709, November 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/67/8/1697/603040.

deMoor:2011:MPA

[dMBD11] Carryn L. de Moor, Douglas S. Butterworth, and José A. A. De Oliveira. Is the management procedure approach equipped to handle short-lived pelagic species with their boom and bust dynamics? The case of the South African fishery for sardine and anchovy. ICES Journal of Marine Science, 68(10):2075–2085, November 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/10/2075/617354.

DArcy:2013:PSP

[DMF13] J. D’Arcy, L. Mirimin, and R. FitzGerald. Phylogeographic structure of a protogynous hermaphrodite species, the ballan wrasse Labrus bergylta, in Ireland, Scotland, and Norway, using mitochondrial DNA sequence data. ICES Journal of Marine Science, 70(3):685–693, April 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/3/685/918519.

Danckwerts:2014:BCB

[DMJ^{+}14] D. K. Danckwerts, C. D. McQuaid, A. Jaeger, G. K. McGregor, R. Dwight, M. Le Corre, and S. Jaquemet. Biomass consumption by breeding seabirds in the western Indian Ocean: indirect interactions with fisheries and implications for management. ICES Journal of Marine Science, 71(9):2589–2598, November 2014.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/9/2589/2798172.

DiCosimo:2010:UAC

[DMO10] Jane DiCosimo, Richard D. Methot, Jr., and Olav A. Ormseth. Use of annual catch limits to avoid stock depletion in the Bering Sea and Aleutian Islands management area (Northeast Pacific). ICES Journal of Marine Science, 67(9):1861–1865, December 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/67/
9/1861/619696.

Damalas:2015:HDM

[DMO^{+}15a] Dimitrios Damalas, Christos D. Maravelias, Giacomo C. Osio, Francesc Maynou, Mario Sbrana, Paolo Sartor, and John Casey. Historical discarding in Mediterranean fisheries: a fishers’ perception. ICES Journal of Marine Science, 72(9):2600–2608, November 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/72/
9/2600/2458665.

Davie:2015:MFC

[DMO^{+}15b] Sarah Davie, Cóilín Minto, Rick Officer, Colm Lordan, and Emmet Jackson. Modelling fuel consumption of fishing vessels for predictive use. ICES Journal of Marine Science, 72(2):708–719, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/72/2/708/670365.

Delavenne:2012:SCP

[DMS^{+}12] Juliette Delavenne, Kristian Metcalfe, Robert J. Smith, Sandrine Vaz, Corinne S. Martin, Ludovic Dupuis, Franck Coppin, and André Carpentier. Systematic conservation planning in the eastern English Channel: comparing the marxan and zonation decision-support tools. ICES Journal of Marine Science, 69(1):75–83, January 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/1/75/672823.

Diesing:2016:IBS

[DMS16] Markus Diesing, Peter Mitchell, and David Stephens. Image-based seabed classification: what can we learn from terrestrial remote sensing? ICES Journal of Marine Science, 73(10):2425–2441, November 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
10/2425/2647114.

Drinkwater:2018:SBW

[DMS18] Kenneth F. Drinkwater, Franz J. Mueter, and Sei-Ichi Saitoh. Shifting boundaries of water, ice, flora, fauna, people, and institutions in the Arctic and subarctic. ICES Journal of Marine Science, 75(7):2293–2298, December 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/7/2293/5256682.

Dennard:2010:HAR

[DMT^{+}10a] Susan T. Dennard, M. Aaron MacNeil, Margaret A. Treble, Steven Campana, and Aaron T. Fisk. Hierarchical analysis of a remote, Arctic, artisanal longline fishery. ICES Journal of Marine Science, 67(1):41–51, January 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/67/1/41/595009.

Doray:2010:EGE

[DMT10b] Mathieu Doray, Stéphanie Mahévas, and Verena M. Trenkel. Estimating gear efficiency in a combined acoustic and trawl survey, with reference to the spatial distribution of demersal fish. ICES Journal of Marine Science, 67(4): 668–676, May 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/67/4/668/680831.

DeCelles:2017:UFE

[DMZC17] Gregory R. DeCelles, David Martins, Douglas R. Zemeckis, and Steven X. Cadrin. Using Fishermen’s ecological knowledge to map Atlantic cod spawning grounds on Georges Bank. ICES Journal of Marine Science, 74(6):1587–1601, July 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/6/1587/3748235.

Dedman:2017:TFD

[DOB^{+}17] Simon Dedman, Rick Officer, Deirdre Brophy, Maurice Clarke, and David G. Reid. Towards a flexible decision support tool for MSY-based marine protected area design for skates and rays. ICES Journal of Marine Science, 74(2): 576–587, March 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/74/2/576/2669563.

Davoren:2012:WTT

[DPBM12] Gail K. Davoren, Paulette Penton, Chantelle Burke, and William A. Montevecchi. Water temperature and timing of capelin spawning determine seabird diets. ICES Journal of Marine Science, 69(7):1234–1241, September 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/7/1234/741220.

Dixon:2012:CTP

[DPD^{+}12] Heather J. Dixon, Michael Power, J. Brian Dempson, Timothy F. Sheehan, and Gérald Chaput. Characterizing the trophic position and shift in Atlantic salmon (Salmo salar) from freshwater to marine life-cycle phases using stable isotopes. ICES Journal of Marine Science, 69(9):1646–1655, November 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/69/9/1646/640360.

Deng:2015:ICP

[DPD^{+}15] Roy A. Deng, André E. Punt, Catherine M. Dichmont, Rik C. Buckworth, and Charis Y. Burridge. Improving catch prediction for tiger prawns in the Australian northern prawn fishery. ICES Journal of Marine Science, 72(1): 117–129, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/72/1/117/822868.

Drakou:2017:WET

[DPE^{+}17] Evangelia G. Drakou, Linwood Pendleton, Micah Effron, Jane Carter Ingram, and Lida Teneva. When ecosystems and their services are not co-located: oceans and coasts. ICES Journal of Marine Science, 74(6):1531–1539, July 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/6/1531/3072608.

dePontual:2013:NIE

[dPJGB13] Hélène de Pontual, Aurélie Jolivet, François Garren, and Michel Bertignac. New insights on European hake biology and population dynamics from a sustained tagging effort in the Bay of Biscay. ICES Journal of Marine Science, 70(7):1416–1428, November 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/70/
7/1416/609286.

Desprez:2010:BIO

[DPL10] M. Desprez, B. Pearce, and S. Le Bot. The biological impact of overflowing sands around a marine aggregate extraction site: Dieppe (eastern English Channel). ICES Journal of Marine Science, 67(2):270–277, March 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/2/270/695074.

Dolan:2016:DCM

[DPL16] Tara E. Dolan, Wesley S. Patrick, and Jason S. Link. Delineating the continuum of marine ecosystem-based management: a US fisheries reference point perspective. ICES Journal of Marine Science, 73(4):1042–1050, March 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/4/1042/2458509.

dePontual:2019:NIB

[dPLF^{+}19] Hélène de Pontual, Maxime Lalire, Ronan Fablet, Claire Laspougeas, François Garren, Stéphane Martin, Mickael Drogou, and Mathieu Woillez. New insights into behavioural ecology of European seabass off the West Coast of France: implications at local and population scales. ICES Journal of Marine Science, 76(2):501–515, March 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/2/501/5057665.

Dominguez-Petit:2013:RSE

[DPOL13] Rosario Domínguez-Petit, Patrick Ouellet, and Yvan Lambert. Reproductive strategy, egg characteristics and embryonic development of Greenland halibut (Reinhardtius hippoglossoides). ICES Journal of Marine Science, 70(2): 342–351, March 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/70/2/342/796040.

Dominguez-Petit:2018:EUA

[DPRG^{+}18] R. Dominguez-Petit, R. M. Rideout, D. Garabana, Y. Lambert, M. Hermida, and M. J. Morgan. Evaluating the use of the autodiametric method for estimating fecundity of Reinhardtius hippoglossoides, a species with an unusual oocyte development strategy. ICES Journal of Marine Science, 75(2):831–839, March 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/75/
2/831/4096826.

deQuevedo:2010:SBL

[dQCD^{+}10] Irene Álvarez de Quevedo, Luis Cardona, Andrea De Haro, Eva Pubill, and Alex Aguilar. Sources of bycatch of loggerhead sea turtles in the western Mediterranean other than drifting longlines. ICES Journal of Marine Science, 67(4): 677–685, May 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/67/4/677/680975.

Downey:2010:ISB

[DRB10] Nicola J. Downey, Michael J. Roberts, and Dan Baird. An investigation of the spawning behaviour of the chokka squid Loligo reynaudii and the potential effects of temperature using acoustic telemetry. ICES Journal of Marine Science, 67(2):231–243, March 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/67/2/231/693114.

Da-Rocha:2017:EFM

[DRGCGJ17] José-María Da-Rocha, Javier García-Cutrín, María-José Gutiérrez, and Ernesto Jardim. Endogenous fishing mortalities: a state-space bioeconomic model. ICES Journal of Marine Science, 74(9):2437–2447, November 2017.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/9/2437/3823334.

Duran-Romero:2017:SUS

[DRVV^{+}17] Cristina Durán-Romero, Virginia E. Villafañe, Macarena S. Valiñas, Rodrigo J. Gonçalves, and E. Walter Helbling. Solar UVR sensitivity of phyto- and bacterioplankton communities from Patagonian coastal waters under increased nutrients and acidification. ICES Journal of Marine Science, 74(4):1062–1073, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
4/1062/2936869.

Diesing:2013:PMA

[DSA13] Markus Diesing, David Stephens, and John Aldridge. A proposed method for assessing the extent of the seabed significantly affected by demersal fishing in the Greater North Sea. ICES Journal of Marine Science, 70(6):1085– 1096, September 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/70/6/1085/638182.

Delfour-Samama:2014:RPL

[DSL14] Odile Delfour-Samama and Cédric Leboeuf. Review of potential legal frameworks for effective implementation and enforcement of MPAs in the high seas. ICES Journal of Marine Science, 71(5):1031–1039, July 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/5/1031/647372.

Moraes:2012:BSS

[dSMGKP12] Luiz Eduardo de Souza Moraes, Douglas Francisco Marcolino Gherardi, Mario Katsuragawa, and Eduardo Tavares Paes. Brazilian sardine (Sardinella brasiliensis Steindachner, 1879) spawning and nursery habitats: spatialscale partitioning and multiscale relationships with thermohaline descriptors. ICES Journal of Marine Science, 69 (6):939–952, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/69/6/939/619178.

Dankel:2016:WYW

[DSN16] Dorothy J. Dankel, Kari Stange, and Kåre Nolde Nielsen. What hat are you wearing? On the multiple roles of fishery scientists in the ICES community. ICES Journal of Marine Science, 73(2):209–216, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/2/209/2614413.

Dietz:2013:MSF

[DTA^{+}13] Rune Dietz, Jonas Teilmann, Signe M. Andersen, Frank Rigét, and Morten T. Olsen. Movements and site fidelity of harbour seals (Phoca vitulina) in Kattegat, Denmark, with implications for the epidemiology of the phocine distemper virus. ICES Journal of Marine Science, 70(1):186–195, January 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/70/1/186/661282.

DiMaida:2011:DBP

[DTL^{+}11] Germana Di Maida, Agostino Tomasello, Filippo Luzzu, Antonino Scannavino, Maria Pirrotta, Carla Orestano, and Sebastiano Calvo. Discriminating between Posidonia oceanica meadows and sand substratum using multibeam sonar. ICES Journal of Marine Science, 68(1):12–19, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/68/1/12/627859.

Deporte:2012:RMD

[DUM^{+}12] Nicolas Deporte, Clara Ulrich, Stéphanie Mahévas, Sébastien Demanèche, and Francois Bastardie. Regional métier definition: a comparative investigation of statistical methods using a workflow applied to international otter trawl fisheries in the North Sea. ICES Journal of Marine Science, 69(2):331–342, March 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/69/2/331/698671.

Diopere:2018:SGF

[DVH^{+}18] Eveline Diopere, Sara G. Vandamme, Pascal I. Hablützel, Alessia Cariani, Jeroen Van Houdt, Adriaan Rijnsdorp, Fausto Tinti, Filip A. M. Volckaert, and Gregory E. Maes. Seascape genetics of a flatfish reveals local selection under high levels of gene flow. ICES Journal of Marine Science, 75(2):675–689, March 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/75/
2/675/4094880.

Depestele:2011:QCD

[DVV^{+}11] Jochen Depestele, Sofie Vandemaele, Willy Vanhee, Hans Polet, Els Torreele, Herwig Leirs, and Magda Vincx. Quantifying causes of discard variability: an indispensable assistance to discard estimation and a paramount need for policy measures. ICES Journal of Marine Science, 68(8):1719– 1725, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/8/1719/746897.

DeRobertis:2010:SSSb

[DW10] Alex De Robertis and Christopher D. Wilson. Silent ships sometimes do encounter more fish. 2. concurrent echosounder observations from a free-drifting buoy and vessels. ICES Journal of Marine Science, 67(5):996–1003, July 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/5/996/609471.

DeRobertis:2011:SSD

[DW11] Alex De Robertis and Christopher D. Wilson. Silent ships do not always encounter more fish (revisited): comparison of acoustic backscatter from walleye pollock recorded by a noise-reduced and a conventional research vessel in the eastern Bering Sea. ICES Journal of Marine Science, 68(10):2229–2239, November 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/10/2229/612073.

DeRobertis:2013:URN

[DWFD13] Alex De Robertis, Christopher D. Wilson, Scott R. Furnish, and Peter H. Dahl. Underwater radiated noise measurements of a noise-reduced fisheries research vessel. ICES Journal of Marine Science, 70(2):480–484, March 2013.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/2/480/794881.

Doubleday:2011:ADM

[DWPS11] Zoë A. Doubleday, Jimmy White, Gretta T. Pecl, and Jayson M. Semmens. Age determination in merobenthic octopuses using stylet increment analysis: assessing future challenges using Macroctopus maorum as a model. ICES Journal of Marine Science, 68(10):2059–2063, November 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/10/2059/609885.

DeRobertis:2010:SSSa

[DWW^{+}10] Alex De Robertis, Christopher D. Wilson, Neal J. Williamson, Michael A. Guttormsen, and Sarah Stienessen. Silent ships sometimes do encounter more fish. 1. vessel comparisons during winter pollock surveys. ICES Journal of Marine Science, 67(5):985–995, July 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/67/
5/985/609223.

Demer:2014:CRM

[DZ14] David A. Demer and Juan P. Zwolinski. Corroboration and refinement of a method for differentiating landings from two stocks of Pacific sardine (Sardinops sagax) in the California Current. ICES Journal of Marine Science, 71(2):328–335, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/71/2/328/784419.

Demer:2013:SSA

[DZC^{+}13] David A. Demer, Juan P. Zwolinski, George R. Cutter, Kyle A. Byers, Beverly J. Macewicz, and Kevin T. Hill. Sampling selectivity in acoustic-trawl surveys of Pacific sardine (Sardinops sagax) biomass and length distribution. ICES Journal of Marine Science, 70(7):1369–1377, November 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/70/7/1369/610568.

Eddy:2017:ELF

[EAB^{+}17] Tyler D. Eddy, Julio N. Araújo, Alida Bundy, Elizabeth A. Fulton, and Heike K. Lotze. Effectiveness of lobster fisheries management in New Zealand and Nova Scotia from multi-species and ecosystem perspectives. ICES Journal of Marine Science, 74(1):146–157, January 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/1/146/2669549.

Espeland:2015:MDP

[EAOB15] Sigurd Heiberg Espeland, Jon Albretsen, Esben Moland Olsen, and Torjan Bodvin. Modelling drift of pelagic offspring: the importance of egg surveys in providing a realistic model initialization. ICES Journal of Marine Science, 72(9):2578–2589, November 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/72/
9/2578/2457872.

Eigaard:2016:CES

[EBB^{+}16a] Ole R. Eigaard, Francois Bastardie, Mike Breen, Grete E. Dinesen, Niels T. Hintzen, Pascal Laffargue, Lars O. Mortensen, J. Rasmus Nielsen, Hans Nilsson, Finbarr G. O’Neill, Hans Polet, David G. Reid, Antonello Sala, Mattias Sköld, Chris Smith, Thomas K. Sørensen, Oliver Tully, Mustafa Zengin, and Adriaan D. Rijnsdorp. A correction to “Estimating seabed pressure from demersal trawls, seines and dredges based on gear design and dimensions”. ICES Journal of Marine Science, 73(9):2420–2423, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/9/2420/2199719.

Eigaard:2016:ESP

[EBB^{+}16b] Ole R. Eigaard, Francois Bastardie, Mike Breen, Grete E. Dinesen, Niels T. Hintzen, Pascal Laffargue, Lars O. Mortensen, J. Rasmus Nielsen, Hans C. Nilsson, Finbarr G. O’Neill, Hans Polet, David G. Reid, Antonello Sala, Mattias Sköld, Chris Smith, Thomas K. Sørensen, Oliver Tully, Mustafa Zengin, and Adriaan D. Rijnsdorp. Estimating seabed pressure from demersal trawls, seines, and dredges based on gear design and dimensions. ICES Journal of Marine Science, 73(suppl 1):S27–S43, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/73/
suppl_1/i27/2573989.

Eigaard:2017:FBT

[EBH^{+}17] Ole R. Eigaard, Francois Bastardie, Niels T. Hintzen, Lene Buhl-Mortensen, Pål Buhl-Mortensen, Rui Catarino, Grete E. Dinesen, Josefine Egekvist, Heino O. Fock, Kerstin Geitner, Hans D. Gerritsen, Manuel Marín González, Patrik Jonsson, Stefanos Kavadas, Pascal Laffargue, Mathieu Lundy, Genoveva Gonzalez-Mirelis, J. Rasmus Nielsen, Nadia Papadopoulou, Paulette E. Posen, Jacopo Pulcinella, Tommaso Russo, Antonello Sala, Cristina Silva, Christopher J. Smith, Bart Vanelslander, and Adriaan D. Rijnsdorp. The footprint of bottom trawling in European waters: distribution, intensity, and seabed integrity. ICES Journal of Marine Science, 74(3):847–865, March 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/3/847/2631171.

Essington:2015:NMP

[EBSW15] Timothy E. Essington, Marissa L. Baskett, James N. Sanchirico, and Carl Walters. A novel model of predator– prey interactions reveals the sensitivity of forage fish: piscivore fishery trade-offs to ecological conditions. ICES Journal of Marine Science, 72(5):1349–1358, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/5/1349/766979.

Eddy:2015:TOB

[ECFL15] Tyler D. Eddy, Marta Coll, Elizabeth A. Fulton, and Heike K. Lotze. Trade-offs between invertebrate fisheries catches and ecosystem impacts in coastal New Zealand. ICES Journal of Marine Science, 72(5):1380–1388, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/5/1380/772875.

Eayrs:2015:MCF

[ECG15] Stephen Eayrs, Steven X. Cadrin, and Christopher W. Glass. Managing change in fisheries: a missing key to fishery-dependent data collection? ICES Journal of Marine Science, 72(4):1152–1158, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/72/
4/1152/800733.

Egerton:2017:BCA

[ECW^{+}17] Sian Egerton, Sarah Culloty, Jason Whooley, Catherine Stanton, and R. Paul Ross. Boarfish (Capros aper): review of a new capture fishery and its valorization potential. ICES Journal of Marine Science, 74(8):2059–2068, September 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/8/2059/3738531.

Elvenes:2014:ECS

[EDBMB14] Sigrid Elvenes, Margaret F. J. Dolan, Pål Buhl-Mortensen, and Valérie K. Bellec. An evaluation of compiled singlebeam bathymetry data as a basis for regional sediment and biotope mapping. ICES Journal of Marine Science, 71(4): 867–881, May 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/71/4/867/666460.

Eriksen:2014:LCL

[EDP14] Elena Eriksen, Caroline M. F. Durif, and Dmitry Prozorkevich. Lumpfish (Cyclopterus lumpus) in the Barents Sea: development of biomass and abundance indices, and spatial distribution. ICES Journal of Marine Science, 71(9):2398– 2402, November 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/71/9/2398/594340.

Engelhard:2011:ECE

[EEP^{+}11] Georg H. Engelhard, Jim R. Ellis, Mark R. Payne, Remment ter Hofstede, and John K. Pinnegar. Ecotypes as a concept for exploring responses to climate change in fish assemblages. ICES Journal of Marine Science, 68(3):580–591, March 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/68/3/580/658004.

Eero:2012:RPD

[Eer12] Margit Eero. Reconstructing the population dynamics of sprat (Sprattus sprattus balticus) in the Baltic Sea in the 20th century. ICES Journal of Marine Science, 69(6):1010– 1018, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/69/6/1010/618064.

Ewing:2015:NPC

[EF15] Graeme Ewing and Stewart Frusher. New puerulus collector design suitable for fishery-dependent settlement monitoring. ICES Journal of Marine Science, 72(S1):S225–S231, July 2015. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/72/suppl_1/i225/618656.

Escalle:2017:FCS

[EGC^{+}17] Lauriane Escalle, Daniel Gaertner, Pierre Chavance, Alicia Delgado de Molina, Javier Ariz, and Bastien Mérigot. Forecasted consequences of simulated FAD moratoria in the Atlantic and Indian Oceans on catches and bycatches. ICES Journal of Marine Science, 74(3):780–792, March 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/3/780/2674178.

Eero:2015:EBC

[EHB^{+}15] Margit Eero, Joakim Hjelm, Jane Behrens, Kurt Buchmann, Massimiliano Cardinale, Michele Casini, Pavel Gasyukov, Noél Holmgren, Jan Horbowy, Karin Hüssy, Eskild Kirkegaard, Georgs Kornilovs, Uwe Krumme, Friedrich W. Köster, Rainer Oeberst, Maris Plikshs, Krzysztof Radtke, Tiit Raid, Jörn Schmidt, Maciej T. Tomczak, Morten Vinther, Christopher Zimmermann, and Marie Storr-Paulsen. Eastern Baltic cod in distress: biological changes and challenges for stock assessment. ICES Journal of Marine Science, 72(8):2180–2186, September 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/8/2180/2458921.

Ellis:2015:ELS

[EHD^{+}15] Charlie D. Ellis, David J. Hodgson, Carly L. Daniels, Dominic P. Boothroyd, R. Colin A. Bannister, and Amber G. F. Griffiths. European lobster stocking requires comprehensive impact assessment to determine fishery benefits. ICES Journal of Marine Science, 72(S1):S35–S48, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/72/suppl_1/i35/614205.

Emery:2014:FRH

[EHG^{+}14] Timothy J. Emery, Klaas Hartmann, Bridget S. Green, Caleb Gardner, and John Tisdell. Fishing for revenue: how leasing quota can be hazardous to your health. ICES Journal of Marine Science, 71(7):1854–1865, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/7/1854/671173.

Eero:2014:ISR

[EHHH14] Margit Eero, Jakob Hemmer-Hansen, and Karin Hüssy. Implications of stock recovery for a neighbouring management unit: experience from the Baltic cod. ICES Journal of Marine Science, 71(6):1458–1466, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/6/1458/2835576.

Elliott:2017:BFC

[EHT17] Anna Elliott, Victoria Hobson, and Kam W. Tang. Balancing fishery and conservation: a case study of the barrel jellyfish Rhizostoma octopus in South Wales. ICES Journal of Marine Science, 74(1):234–241, January 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/1/234/2669573.

Eide:2018:SMB

[Eid18] Arne Eide. Substituting model-based indicators in harvest control rules by observations using fuzzy logic methodology. ICES Journal of Marine Science, 75(3):977–987, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/3/977/4775004.

Eriksen:2012:THG

[EISJ12] Elena Eriksen, Randi Ingvaldsen, Jan Erik Stiansen, and Geir Odd Johansen. Thermal habitat for 0-group fish in the Barents Sea; how climate variability impacts their density, length, and geographic distribution. ICES Journal of Marine Science, 69(5):870–879, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/5/870/646878.

Ellis:2015:GED

[EKD^{+}15] Charlie D. Ellis, Hannah Knott, Carly L. Daniels, Matthew J. Witt, and David J. Hodgson. Geographic and environmental drivers of fecundity in the European lobster (Homarus gammarus). ICES Journal of Marine Science, 72(S1):S91–S100, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/72/suppl_1/i91/
619150.

Einberg:2019:MDA

[EKR^{+}19] Heli Einberg, Riina Klais, Gunta Rubene, Georgs Kornilovs, Ivars Putnis, and Henn Ojaveer. Multidecadal dynamics of the Arctic copepod Limnocalanus macrurus in relation to environmental variability in the Baltic Sea. ICES Journal of Marine Science, 76(7):2427–2436, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/7/2427/5511709.

Enders:2017:EMB

[ELBS17] Kristina Enders, Robin Lenz, Sabrina Beer, and Colin A. Stedmon. Extraction of microplastic from biota: recommended acidic digestion destroys common plastic polymers. ICES Journal of Marine Science, 74(1):326–331, January 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/1/326/2418179.

Elvarsson:2015:ESS

[Elv15] Bjarki Tór Elvarsson. Evaluating stock structure hypotheses using genetically determined close relatives: a simulation study on North Atlantic fin whales. ICES Journal of Marine Science, 72(2):661–669, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/2/661/2801352.

Eigaard:2011:IFR

[EMP11] Ole Ritzau Eigaard and Sten Munch-Petersen. Influence of fleet renewal and trawl development on landings per unit effort of the Danish northern shrimp (Pandalus borealis) fishery. ICES Journal of Marine Science, 68(1):26–31, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/68/1/26/628876.

Enochs:2016:MCA

[EMW^{+}16] I. C. Enochs, D. P. Manzello, H. H. Wirshing, R. Carlton, and J. Serafy. Micro-CT analysis of the Caribbean octocoral Eunicea flexuosa subjected to elevated pCO_{2}. ICES Journal of Marine Science, 73(3):910–919, February 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/3/910/2458732.

Essington:2014:PGR

[EP14] Timothy E. Essington and Éva E. Plagányi. Pitfalls and guidelines for “recycling” models for ecosystem-based fisheries management: evaluating model suitability for forage fish fisheries. ICES Journal of Marine Science, 71(1):118– 127, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/71/1/118/641481.

Eayrs:2019:MVU

[EP19] Stephen Eayrs and Michael Pol. The myth of voluntary uptake of proven fishing gear: investigations into the challenges inspiring change in fisheries. ICES Journal of Marine Science, 76(2):392–401, March 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/2/392/5224774.

Evans:2015:CAB

[EPH15] Jon L. Evans, Frances Peckett, and Kerry L. Howell. Combined application of biophysical habitat mapping and systematic conservation planning to assess efficiency and representativeness of the existing High Seas MPA network in the Northeast Atlantic. ICES Journal of Marine Science, 72(5): 1483–1497, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/72/5/1483/774114.

Eisner:2018:SIS

[EPK^{+}18] Lisa B. Eisner, Alexei I. Pinchuk, David G. Kimmel, Kathryn L. Mier, Colleen E. Harpold, and Elizabeth C. Siddon. Seasonal, interannual, and spatial patterns of community composition over the eastern Bering Sea shelf in cold years. part i: zooplankton. ICES Journal of Marine Science, 75(1):72–86, January 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/75/1/72/4093372.

Engelhard:2011:NDN

[EPKR11] Georg H. Engelhard, John K. Pinnegar, Laurence T. Kell, and Adriaan D. Rijnsdorp. Nine decades of North Sea sole and plaice distribution. ICES Journal of Marine Science, 68(6):1090–1104, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/68/
6/1090/701797.

Erisman:2011:SSC

[EPPL^{+}11] Brad E. Erisman, Gustavo A. Paredes, Tomas Plomozo-Lugo, Juan J. Cota-Nieto, Philip A. Hastings, and Octavio Aburto-Oropeza. Spatial structure of commercial marine fisheries in Northwest Mexico. ICES Journal of Marine Science, 68(3):564–571, March 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/68/3/564/657590.

Engelhard:2014:FFT

[EPR^{+}14] Georg H. Engelhard, Myron A. Peck, Anna Rindorf, Sophie C. Smout, Mikael van Deurs, Kristina Raab, Ken H. Andersen, Stefan Garthe, Rebecca A. M. Lauerburg, Finlay Scott, Thomas Brunel, Geert Aarts, Tobias van Kooten, and Mark Dickey-Collas. Forage fish, their fisheries, and their predators: who drives whom? ICES Journal of Marine Science, 71(1):90–104, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/1/90/643678.

Eliasen:2014:SEI

[EPVC14] Søren Q. Eliasen, K.-Nadia Papadopoulou, Vassiliki Vassilopoulou, and Tom L. Catchpole. Socio-economic and institutional incentives influencing fishers’ behaviour in relation to fishing practices and discard. ICES Journal of Marine Science, 71(5):1298–1307, July 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/71/
5/1298/640460.

Eriksen:2016:DSJ

[Eri16] Elena Eriksen. Do scyphozoan jellyfish limit the habitat of pelagic species in the Barents Sea during the late feeding period? ICES Journal of Marine Science, 73(2):217–226, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/2/217/2614326.

Espeland:2018:ECE

[ES18] Sigurd Heiberg Espeland and Hanne Sannæs. Estimating cod egg developmental stage based on DNA concentration. ICES Journal of Marine Science, 75(2):825–830, March 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/2/825/4157379.

Eero:2015:DRC

[ESAV15] Margit Eero, Harry V. Strehlow, Charles M. Adams, and Morten Vinther. Does recreational catch impact the TAC for commercial fisheries? ICES Journal of Marine Science, 72(2):450–457, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/72/
2/450/2801306.

Eriksen:2016:BSE

[ESD^{+}16] Elena Eriksen, Hein Rune Skjoldal, Andrey V. Dolgov, Padmini Dalpadado, Emma L. Orlova, and Dmitry V. Prozorkevich. The Barents Sea euphausiids: methodological aspects of monitoring and estimation of abundance and biomass. ICES Journal of Marine Science, 73(6):1533– 1544, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/6/1533/2458930.

Emond:2015:TVB

[ÉSMGB15] Kim Émond, Bernard Sainte-Marie, Peter S. Galbraith, and Joël Bêty. Top-down vs. bottom-up drivers of recruitment in a key marine invertebrate: investigating early life stages of snow crab. ICES Journal of Marine Science, 72(5):1336– 1348, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/5/1336/765923.

Estefanell:2010:GMO

[ESR^{+}10] Juan Estefanell, Juan Socorro, Francisco J. Roo, Rafael Guirao, Hipólito Fernández-Palacios, and Marisol Izquierdo. Gonad maturation in Octopus vulgaris during ongrowing, under different conditions of sex ratio. ICES Journal of Marine Science, 67(7):1487–1493, October 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/67/
7/1487/666767.

Enriquez:2015:NSC

[ET15] Rica Mae Enriquez and John R. Taylor. Numerical simulations of the competition between wind-driven mixing and surface heating in triggering spring phytoplankton blooms. ICES Journal of Marine Science, 72(6):1926–1941, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/6/1926/926475.

Emery:2015:EAU

[ETG^{+}15] Timothy J. Emery, John Tisdell, Bridget S. Green, Klaas Hartmann, Caleb Gardner, and Rafael León. Experimental analysis of the use of fishery closures and cooperatives to reduce economic rent dissipation caused by assignment problems. ICES Journal of Marine Science, 72(9):2650– 2662, November 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/72/9/2650/2458742.

Engelhard:2016:IMM

[ETM^{+}16] Georg H. Engelhard, Ruth H. Thurstan, Brian R. MacKenzie, Heidi K. Alleway, R. Colin A. Bannister, Massimiliano Cardinale, Maurice W. Clarke, Jock C. Currie, Tomaso Fortibuoni, Poul Holm, Sidney J. Holt, Carlotta Mazzoldi, John K. Pinnegar, Saša Raicevich, Filip A. M. Volckaert, Emily S. Klein, and Ann-Katrien Lescrauwaet. ICES meets marine historical ecology: placing the history of fish and fisheries in current policy context. ICES Journal of Marine Science, 73(5):1386–1403, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/5/1386/2240665.

Eriander:2016:SDP

[EWH16] L. Eriander, A.-L. Wrange, and J. N. Havenhand. Simulated diurnal pH fluctuations radically increase variance in — but not the mean of — growth in the barnacle Balanus improvisus. ICES Journal of Marine Science, 73(3):596– 603, February 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/73/3/596/2458933.

Furusawa:2010:AEO

[FA10] Masahiko Furusawa and Kazuo Amakasu. The analysis of echotrace obtained by a split-beam echosounder to observe the tilt-angle dependence of fish target strength in situ. ICES Journal of Marine Science, 67(2):215–230, March 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/2/215/695440.

Frieder:2017:MCC

[FAP^{+}17] Christina A. Frieder, Scott L. Applebaum, T.-C. Francis Pan, Dennis Hedgecock, and Donal T. Manahan. Metabolic cost of calcification in bivalve larvae under experimental ocean acidification. ICES Journal of Marine Science, 74 (4):941–954, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/74/4/941/2660821.

Fraile:2015:OAB

[FAR15] Igaratza Fraile, Haritz Arrizabalaga, and Jay R. Rooker. Origin of Atlantic bluefin tuna (Thunnus thynnus) in the Bay of Biscay. ICES Journal of Marine Science, 72(2):625– 634, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/72/2/625/2801458.

Faunce:2011:CBI

[Fau11] Craig H. Faunce. A comparison between industry and observer catch compositions within the Gulf of Alaska rockfish fishery. ICES Journal of Marine Science, 68(8):1769– 1777, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/8/1769/754921.

Faunce:2011:FQA

[FB11] Craig H. Faunce and Steven J. Barbeaux. The frequency and quantity of Alaskan groundfish catcher-vessel landings made with and without an observer. ICES Journal of Marine Science, 68(8):1757–1763, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/8/1757/753689.

Fincham:2019:ULM

[FB19] Jennifer I. Fincham and Jon Barry. The use of linear mixed effects models to investigate local adaptation in marine fish subpopulations: case study, changing spawning times of common sole (Solea solea). ICES Journal of Marine Science, 76(7):2297–2304, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/7/2297/5580658.

Fangel:2017:IBN

[FBCD^{+}17] Kirstin Fangel, Kim Magnus Bærum, Signe Christensen-Dalsgaard, Øystein Aas, and Tycho Anker-Nilssen. Incidental bycatch of northern fulmars in the small-vessel demersal longline fishery for Greenland halibut in coastal Norway 2012–2014. ICES Journal of Marine Science, 74(1): 332–342, January 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/74/1/332/2669565.

Frawley:2019:ISW

[FBD^{+}19] Timothy H. Frawley, Dana K. Briscoe, Patrick C. Daniel, Gregory L. Britten, Larry B. Crowder, Carlos J. Robinson, and William F. Gilly. Impacts of a shift to a warm-water regime in the Gulf of California on jumbo squid (Dosidicus gigas). ICES Journal of Marine Science, 76(7):2413–2426, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/7/2413/5533276.

Forseth:2017:MTA

[FBF^{+}17] Torbjørn Forseth, Bjørn T. Barlaup, Bengt Finstad, Peder Fiske, Harald Gjøsæter, Morten Falkegård, Atle Hindar, Tor Atle Mo, Audun H. Rikardsen, Eva B. Thorstad, Leif Asbjørn Vøllestad, and Vidar Wennevik. The major threats to Atlantic salmon in Norway. ICES Journal of Marine Science, 74(6):1496–1513, July 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/6/1496/3061737.

Fernandez-Boan:2013:MFP

[FBFP^{+}13] María Fernández-Boán, Juan Freire, Ana M. Parma, Luis Fernández, and José M. Orensanz. Monitoring the fishing process in the sea urchin diving fishery of Galicia. ICES Journal of Marine Science, 70(3):604–617, April 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/3/604/915153.

Ferreira:2016:PDC

[FBL^{+}16] Guilherme V. B. Ferreira, Mario Barletta, Andre R. A. Lima, David V. Dantas, Anne K. S. Justino, and Monica F. Costa. Plastic debris contamination in the life cycle of acoupa weakfish (Cynoscion acoupa) in a tropical estuary. ICES Journal of Marine Science, 73(10):2695–2707, November 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/10/2695/2647104.

Forget:2018:NME

[FBM^{+}18] Guillaume Forget, Jean-Luc Baglinière, Frédéric Marchand, Arnaud Richard, and Marie Nevoux. A new method to estimate habitat potential for Atlantic salmon (Salmo salar): predicting the influence of dam removal on the Sélune River (France) as a case study. ICES Journal of Marine Science, 75(6):2172–2181, November 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/6/2172/5056129.

Fock:2019:SSC

[FC19] Heino O. Fock and Stephanie Czudaj. Size structure changes of mesopelagic fishes and community biomass size spectra along a transect from the equator to the Bay of Biscay collected in 1966–1979 and 2014–2015. ICES Journal of Marine Science, 76(3):755–770, May 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/3/755/5049105.

Figus:2017:ULE

[FCB17] Elizabeth Figus, Courtney Carothers, and Anne H. Beaudreau. Using local ecological knowledge to inform fisheries assessment: measuring agreement among Polish fishermen about the abundance and condition of Baltic cod (Gadus morhua). ICES Journal of Marine Science, 74(8):2213– 2222, September 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/74/8/2213/3746140.

Fernandez-Contreras:2010:IBS

[FCCLA10] M. M. Fernández-Contreras, L. Cardona, C. H. Lockyer, and A. Aguilar. Incidental bycatch of short-beaked common dolphins (Delphinus delphis) by pairtrawlers off northwestern Spain. ICES Journal of Marine Science, 67(8):1732– 1738, November 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/67/8/1732/605256.

Fernandes:2011:ODS

[FCD^{+}11] Paul G. Fernandes, Kenny Coull, Craig Davis, Peter Clark, Rui Catarino, Nick Bailey, Rob Fryer, and Alastair Pout. Observations of discards in the Scottish mixed demersal trawl fishery. ICES Journal of Marine Science, 68(8):1734– 1742, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/8/1734/762277.

Fernandes:2016:AEF

[FCG^{+}16] Paul G. Fernandes, Phillip Copland, Rafael Garcia, Tudor Nicosevici, and Ben Scoulding. Additional evidence for fisheries acoustics: small cameras and angling gear provide tilt angle distributions and other relevant data for mackerel surveys. ICES Journal of Marine Science, 73(8):2009– 2019, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/73/8/2009/2198495.

Fernandez:2010:SAP

[FCPJ10] C. Fernández, S. Cerviño, N. Pérez, and E. Jardim. Stock assessment and projections incorporating discard estimates in some years: an application to the hake stock in ICES divisions VIIIc and IXa. ICES Journal of Marine Science, 67(6):1185–1197, September 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/6/1185/737306.

Favaro:2013:TTE

[FDC13] Brett Favaro, Stefanie D. Duff, and Isabelle M. Côté. A trap with a twist: evaluating a bycatch reduction device to prevent rockfish capture in crustacean traps. ICES Journal of Marine Science, 70(1):114–122, January 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/1/114/660896.

Frojan:2016:EBA

[FDC^{+}16] Christopher Barrio Froján, Anna-Leena Downie, Mar Sacau Cuadrado, Ellen Kenchington, and Andrew Kenny. Evaluation of benthic assemblage structure in the NAFO regulatory area with regard to the protection of VME. ICES Journal of Marine Science, 73(2):405–419, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/2/405/2614339.

Finucci:2018:AAD

[FDJ18] Brittany Finucci, Matt R. Dunn, and Emma G. Jones. Aggregations and associations in deep-sea chondrichthyans. ICES Journal of Marine Science, 75(5):1613–1626, September 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/5/1613/4993370.

Finucci:2019:AAD

[FDJ19] Brittany Finucci, Matt R. Dunn, and Emma G. Jones. Aggregations and associations in deep-sea chondrichthyans. ICES Journal of Marine Science, 76(2):466, March 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/2/466/5359481.

Friedland:2017:PSS

[FDR^{+}17] Kevin D. Friedland, Johan Dannewitz, Atso Romakkaniemi, Stefan Palm, Henni Pulkkinen, Tapani Pakarinen, and Rainer Oeberst. Post-smolt survival of Baltic salmon in context to changing environmental conditions and predators. ICES Journal of Marine Science, 74(5):1344–1355, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/5/1344/3059926.

Fallon:2016:CSO

[FFF16] Niall G. Fallon, Sophie Fielding, and Paul G. Fernandes. Classification of Southern Ocean krill and icefish echoes using random forests. ICES Journal of Marine Science, 73(8):1998–2008, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
8/1998/2198397.

Frie:2011:EPA

[FFH^{+}11] Anne K. Frie, Kjell-Arne Fagerheim, Mike O. Hammill, Finn O. Kapel, Christina Lockyer, Garry B. Stenson, Aqqalu Rosing-Asvid, and Vladislav Svetochev. Error patterns in age estimation of harp seals (Pagophilus groenlandicus): results from a transatlantic, image-based, blindreading experiment using known-age teeth. ICES Journal of Marine Science, 68(9):1942–1953, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/9/1942/667345.

Fisher:2011:EHT

[FFK^{+}11] Jonathan A. D. Fisher, Kenneth T. Frank, Vladimir E. Kostylev, Nancy L. Shackell, Tracy Horsman, and Charles G. Hannah. Evaluating a habitat template model’s predictions of marine fish diversity on the Scotian Shelf and Bay of Fundy, Northwest Atlantic. ICES Journal of Marine Science, 68(10):2096–2105, November 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/68/
10/2096/612334.

Farhadi:2013:MDP

[FFN^{+}13] Ahmad Farhadi, Hamid Farhamand, Mohammad Ali Nematollahi, Andrew Jeffs, and Shane D. Lavery. Mitochondrial DNA population structure of the scalloped lobster Panulirus homarus (Linnaeus 1758) from the West Indian Ocean. ICES Journal of Marine Science, 70(7):1491–1498, November 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/70/7/1491/608518.

Farhadi:2016:CMD

[FFN^{+}16] Ahmad Farhadi, Hamid Farahmand, Mohammad Ali Nematollahi, Andrew Jeffs, and Shane D. Lavery. A correction to “Mitochondrial DNA population structure of the scalloped lobster Panulirus homarus (Linnaeus 1758) from the West Indian Ocean”. ICES Journal of Marine Science, 73(10):2747, November 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
10/2747/2647089.

Fisher:2014:LEE

[FFPL14] Jonathan A. D. Fisher, Kenneth T. Frank, Brian Petrie, and William C. Leggett. Life on the edge: environmental determinants of tilefish (Lopholatilus chamaeleonticeps) abundance since its virtual extinction in 1882. ICES Journal of Marine Science, 71(9):2371–2378, November 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/9/2371/593672.

Fung:2012:RDS

[FFRR12] Tak Fung, Keith D. Farnsworth, David G. Reid, and Axel G. Rossberg. Recent data suggest no further recovery in North Sea large fish indicator. ICES Journal of Marine Science, 69(2):235–239, March 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/69/
2/235/702015.

Fernandez:2013:CFT

[FGMC^{+}13] Ana Fernández, Rocío Graña, Beatriz Mouriño-Carballido, Antonio Bode, Manuel Varela, J. Francisco Domínguez-Yanes, José Escánez, Demetrio de Armas, and Emilio Marañón. Community N_{2} fixation and Trichodesmium spp. abundance along longitudinal gradients in the eastern subtropical North Atlantic. ICES Journal of Marine Science, 70(1):223–231, January 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/70/1/223/661237.

Fitzpatrick:2011:BPC

[FGRR11] Mike Fitzpatrick, Norman Graham, Dominic J. Rihan, and Dave G. Reid. The burden of proof in co-management and results-based management: the elephant on the deck! ICES Journal of Marine Science, 68(8):1656–1662, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/8/1656/755559.

Farrell:2012:ODM

[FHC^{+}12] Edward D. Farrell, Karin Hüssy, Julie O. Coad, Lotte W. Clausen, and Maurice W. Clarke. Oocyte development and maturity classification of boarfish (Capros aper) in the Northeast Atlantic. ICES Journal of Marine Science, 69(4): 498–507, May 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/69/4/498/635181.

Fields:2019:ABU

[FHD^{+}19] David M. Fields, Nils Olav Handegard, John Dalen, Christiane Eichner, Ketil Malde, Ørjan Karlsen, Anne Berit Skiftesvik, Caroline M. F. Durif, and Howard I. Browman. Airgun blasts used in marine seismic surveys have limited effects on mortality, and no sublethal effects on behaviour or gene expression, in the copepod Calanus finmarchicus. ICES Journal of Marine Science, 76(7):2033– 2044, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/76/7/2033/5543877.

Frie:2013:CEP

[FHH^{+}13a] Anne K. Frie, Mike O. Hammill, Erlingur Hauksson, Ylva Lind, Christina Lockyer, Olavi Stenman, and Olga Svetocheva. A correction to “Error patterns in age estimation and tooth readability assignment of grey seals (Halichoerus grypus) — results from a transatlantic, image-based blind-reading study using known-age animals”. ICES Journal of Marine Science, 70(4):913, July 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/4/913/728374.

Frie:2013:EPA

[FHH^{+}13b] Anne K. Frie, Mike O. Hammill, Erlingur Hauksson, Ylva Lind, Christina Lockyer, Olavi Stenman, and Olga Svetocheva. Error patterns in age estimation and tooth readability assignment of grey seals (Halichoerus grypus): results from a transatlantic, image-based, blind-reading study using known-age animals. ICES Journal of Marine Science, 70(2):418–430, March 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/70/2/418/794413.

Forland:2014:SPA

[FHK14] Tonje Nesse Forland, Halvor Hobæk, and Rolf J. Korneliussen. Scattering properties of Atlantic mackerel over a wide frequency range. ICES Journal of Marine Science, 71(7):1904–1912, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/71/
7/1904/2804376.

Forland:2014:BBA

[FHOK14] Tonje Nesse Forland, Halvor Hobæk, Egil Ona, and Rolf J. Korneliussen. Broad bandwidth acoustic backscattering from sandeel — measurements and finite element simulations. ICES Journal of Marine Science, 71(7):1894– 1903, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/71/7/1894/670593.

Fernandez-Jover:2015:CDO

[FJSJ15] Damian Fernandez-Jover and Pablo Sanchez-Jerez. Comparison of diet and otolith growth of juvenile wild fish communities at fish farms and natural habitats. ICES Journal of Marine Science, 72(3):916–929, March 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/3/916/689798.

Fernandes:2016:PMF

[FKH^{+}16] Jose A. Fernandes, Susan Kay, Mostafa A. R. Hossain, Munir Ahmed, William W. L. Cheung, Attila N. Lazar, and Manuel Barange. Projecting marine fish production and catch potential in Bangladesh in the 21st century under long-term environmental change and management scenarios. ICES Journal of Marine Science, 73(5):1357–1369, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/5/1357/2296500.

Fromentin:2014:MRH

[FL14] Jean-Marc Fromentin and Daniel Lopuszanski. Migration, residency, and homing of bluefin tuna in the western Mediterranean Sea. ICES Journal of Marine Science, 71(3):

510–518, March 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/71/3/510/635898.

Flynn:2019:RNI

[FLCQ19] Paula Tummon Flynn, Kassandra Devon Lynn, David K. Cairns, and Pedro A. Quijón. The role of the nonindigenous green crab (Carcinus maenas) in the decline of a unique strain of Irish moss (Chondrus crispus): direct and indirect effects. ICES Journal of Marine Science, 76(7):2338–2348, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
7/2338/5529195.

Fletcher:2015:RRE

[Fle15] Warrick (Rick) J. Fletcher. Review and refinement of an existing qualitative risk assessment method for application within an ecosystem-based management framework. ICES Journal of Marine Science, 72(3):1043–1056, March 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/3/1043/687833.

Fay:2015:MPE

[FLLG15] Gavin Fay, Jason S. Link, Scott I. Large, and Robert J. Gamble. Management performance of ecological indicators in the Georges Bank finfish fishery. ICES Journal of Marine Science, 72(5):1285–1296, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/5/1285/759512.

Frank:2013:IPD

[FLP^{+}13] Kenneth T. Frank, William C. Leggett, Brian D. Petrie, Jonathan A. D. Fisher, Nancy L. Shackell, and Christopher T. Taggart. Irruptive prey dynamics following the groundfish collapse in the Northwest Atlantic: an illusion? ICES Journal of Marine Science, 70(7):1299–1307, November 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/70/7/1299/610445.

Farrell:2010:AGE

[FMC10] Edward D. Farrell, Stefano Mariani, and Maurice W. Clarke. Age and growth estimates for the starry smoothhound (Mustelus asterias) in the Northeast Atlantic Ocean. ICES Journal of Marine Science, 67(5):931–939, July 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/5/931/608786.

Frandsen:2010:SEB

[FMK10] Rikke P. Frandsen, Niels Madsen, and Ludvig A. Krag. Selectivity and escapement behaviour of five commercial fishery species in standard square- and diamond-mesh codends. ICES Journal of Marine Science, 67(8):1721–1731, November 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/67/8/1721/603854.

Ferrer-Maza:2014:PCR

[FMLM^{+}14] Dolors Ferrer-Maza, Josep Lloret, Marta Muñoz, Elisabeth Faliex, Sílvia Vila, and Pierre Sasal. Parasitism, condition and reproduction of the European hake (Merluccius merluccius) in the northwestern Mediterranean Sea. ICES Journal of Marine Science, 71(5):1088–1099, July 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/5/1088/645845.

Frojan:2012:EBC

[FMM^{+}12] Christopher R. S. Barrio Froján, Kevin G. MacIsaac, Andrew K. McMillan, María del Mar Sacau Cuadrado, Philip A. Large, Andrew J. Kenny, Ellen Kenchington, and Enrique de Cárdenas González. An evaluation of benthic community structure in and around the Sackville Spur closed area (Northwest Atlantic) in relation to the protection of vulnerable marine ecosystems. ICES Journal of Marine Science, 69(2):213–222, March 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/2/213/705009.

Friedland:2019:ESP

[FMML19] Kevin D. Friedland, M. Conor McManus, Ryan E. Morse, and Jason S. Link. Event scale and persistent drivers of fish and macroinvertebrate distributions on the Northeast US shelf. ICES Journal of Marine Science, 76(5):1316– 1334, 09- 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/5/1316/5197071.

Finley:2013:MSY

[FO13] Carmel Finley and Naomi Oreskes. Maximum sustained yield: a policy disguised as science. ICES Journal of Marine Science, 70(2):245–250, March 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/2/245/797356.

Frelat:2018:TDV

[FOC^{+}18] Romain Frelat, Alessandro Orio, Michele Casini, Andreas Lehmann, Bastien Mérigot, Saskia A. Otto, Camilla Sguotti, and Christian Möllmann. A three-dimensional view on biodiversity changes: spatial, temporal, and functional perspectives on fish communities in the Baltic Sea. ICES Journal of Marine Science, 75(7):2463–2475, December 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/7/2463/4955416.

Fassler:2013:BCA

[FOJ13] Sascha M. M. Fässler, Ciaran O’Donnell, and J. M. Jech. Boarfish (Capros aper) target strength modelled from magnetic resonance imaging (MRI) scans of its swimbladder. ICES Journal of Marine Science, 70(7):1451–1459, November 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/70/7/1451/608068.

Fu:2017:STD

[FOT^{+}17] Caihong Fu, Norm Olsen, Nathan Taylor, Arnaud Grüss, Sonia Batten, Huizhu Liu, Philippe Verley, and Yunne-Jai Shin. Spatial and temporal dynamics of predatorprey species interactions off western Canada. ICES Journal of Marine Science, 74(8):2107–2119, September 2017.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/8/2107/3820951.

Frid:2012:FWW

[FP12] Chris L. J. Frid and Odette A. L. Paramor. Feeding the world: what role for fisheries? ICES Journal of Marine Science, 69(2):145–150, March 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/69/2/145/702270.

Frank:2019:TMB

[FPLB19] Kenneth T. Frank, Brian Petrie, William C. Leggett, and Daniel G. Boyce. Towards a more balanced assessment of the dynamics of North Atlantic ecosystems — a comment on drinkwater and kristiansen (2018). ICES Journal of Marine Science, 76(7):2489–2494, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/7/2489/5554626.

Falk-Petersen:2011:EEE

[FPRA11] Jannike Falk-Petersen, Paul Renaud, and Natalia Anisimova. Establishment and ecosystem effects of the alien invasive red king crab (Paralithodes camtschaticus) in the Barents Sea — a review. ICES Journal of Marine Science, 68(3):479–488, March 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/68/3/479/658855.

Forget:2011:PSS

[FPSF11] Marie-Hélène Forget, Trevor Platt, Shubha Sathyendranath, and Paul Fanning. Phytoplankton size structure, distribution, and primary production as the basis for trophic analysis of Caribbean ecosystems. ICES Journal of Marine Science, 68(4):751–765, March 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/68/
4/751/647809.

Farcas:2016:MSY

[FR16] Adrian Farcas and Axel G. Rossberg. Maximum sustainable yield from interacting fish stocks in an uncertain world: two policy choices and underlying trade-offs. ICES Journal of Marine Science, 73(10):2499–2508, November 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/10/2499/2647108.

Franks:2015:SCD

[Fra15] Peter J. S. Franks. Has Sverdrup’s critical depth hypothesis been tested? Mixed layers vs. turbulent layers. ICES Journal of Marine Science, 72(6):1897–1907, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/6/1897/917319.

Falardeau:2014:CPS

[FRF14] Marianne Falardeau, Dominique Robert, and Louis Fortier. Could the planktonic stages of polar cod and Pacific sand lance compete for food in the warming Beaufort Sea? ICES Journal of Marine Science, 71(7):1956–1965, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/7/1956/663450.

Friedman:2019:BUP

[Fri19] Andrew Friedman. Beyond “not undermining”: possibilities for global cooperation to improve environmental protection in areas beyond national jurisdiction. ICES Journal of Marine Science, 76(2):452–456, March 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/2/452/5304549.

Ferriss:2016:ETN

[FRM^{+}16] Bridget E. Ferriss, Jonathan C. P. Reum, P. Sean Mc-Donald, Dara M. Farrell, and Chris J. Harvey. Evaluating trophic and non-trophic effects of shellfish aquaculture in a coastal estuarine foodweb. ICES Journal of Marine Science, 73(2):429–440, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/2/429/2614240.

Fuller:2014:EFC

[FS14] Daniel W. Fuller and Kurt M. Schaefer. Evaluation of a fishing captain’s ability to predict species composition, sizes, and quantities of tunas associated with drifting fishaggregating devices in the eastern Pacific Ocean. ICES Journal of Marine Science, 71(7):1774–1780, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/7/1774/670709.

Franchini:2012:LLP

[FSC^{+}12] Paolo Franchini, Luciana Sola, Donatella Crosetti, Valentina Milana, and Anna Rita Rossi. Low levels of population genetic structure in the gilthead sea bream, Sparus aurata, along the coast of Italy. ICES Journal of Marine Science, 69(1):41–50, January 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/69/1/41/671474.

Farley:2011:IWE

[FSN^{+}11] Edward V. Farley, Alexander Starovoytov, Svetlana Naydenko, Ron Heintz, Marc Trudel, Charles Guthrie, Lisa Eisner, and Jeffrey R. Guyon. Implications of a warming eastern Bering Sea for Bristol Bay sockeye salmon. ICES Journal of Marine Science, 68(6):1138–1146, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/6/1138/699820.

Fuller:2017:CFC

[FSS^{+}17] Emma C. Fuller, Jameal F. Samhouri, Joshua S. Stoll, Simon A. Levin, and James R. Watson. Characterizing fisheries connectivity in marine social–ecological systems. ICES Journal of Marine Science, 74(8):2087–2096, September 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/8/2087/4082761.

Fine:2017:ESN

[FTM^{+}17] Maoz Fine, Rami Tsadok, Dalit Meron, Stephanie Cohen, and Marco Milazzo. Environmental sensitivity of Neogoniolithon brassica-florida associated with vermetid reefs in the Mediterranean Sea. ICES Journal of Marine Science, 74(4):1074–1082, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
4/1074/2670312.

Fulton:2011:ITW

[Ful11] E. A. Fulton. Interesting times: winners, losers, and system shifts under climate change around Australia. ICES Journal of Marine Science, 68(6):1329–1342, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/6/1329/702275.

Forcada:2010:SST

[FVSL^{+}10] Aitor Forcada, Carlos Valle, José L. Sánchez-Lizaso, Just T. Bayle-Sempere, and Fabio Corsi. Structure and spatiotemporal dynamics of artisanal fisheries around a Mediterranean marine protected area. ICES Journal of Marine Science, 67(2):191–203, March 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/67/2/191/692333.

Forman:2010:VDS

[FW10] Krissy A. Forman and Joseph D. Warren. Variability in the density and sound-speed of coastal zooplankton and nekton. ICES Journal of Marine Science, 67(1):10–18, January 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/67/1/10/594734.

Froese:2018:NAE

[FWC^{+}18] Rainer Froese, Henning Winker, Gianpaolo Coro, Nazli Demirel, Athanassios C. Tsikliras, Donna Dimarchopoulou, Giuseppe Scarcella, Wolfgang Nikolaus Probst, Manuel Dureuil, and Daniel Pauly. A new approach for estimating stock status from length frequency data. ICES Journal of Marine Science, 75(6):2004–2015, November 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/6/2004/5051296.

Froese:2019:NAE

[FWC^{+}19a] Rainer Froese, Henning Winker, Gianpaolo Coro, Nazli Demirel, Athanassios C. Tsikliras, Donna Dimarchopoulou, Giuseppe Scarcella, Wolfgang Nikolaus Probst, Manuel Dureuil, and Daniel Pauly. A new approach for estimating stock status from length frequency data. ICES Journal of Marine Science, 76(1):350–351, January 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/1/350/5156206.

Froese:2019:PEP

[FWC^{+}19b] Rainer Froese, Henning Winker, Gianpaolo Coro, Nazli Demirel, Athanassios C. Tsikliras, Donna Dimarchopoulou, Giuseppe Scarcella, Wolfgang Nikolaus Probst, Manuel Dureuil, and Daniel Pauly. On the pile-up effect and priors for L_{inf} and M /K: response to a comment by Hordyk et al. on “A new approach for estimating stock status from length frequency data”. ICES Journal of Marine Science, 76(2): 461–465, March 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/76/2/461/5298549.

Flores:2015:UGI

[FWD15] Andrés Flores, Rodrigo Wiff, and Eduardo Díaz. Using the gonadosomatic index to estimate the maturity ogive: application to Chilean hake (Merluccius gayi gayi). ICES Journal of Marine Science, 72(2):508–514, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/2/508/2801454.

Ferter:2015:DSE

[FWH^{+}15] Keno Ferter, Marc Simon Weltersbach, Odd-Børre Humborstad, Per Gunnar Fjelldal, Florian Sambraus, Harry Vincent Strehlow, and Jon Helge Vølstad. Dive to survive: effects of capture depth on barotrauma and post-release survival of Atlantic cod (Gadus morhua) in recreational fisheries. ICES Journal of Marine Science, 72(8):2467– 2481, September 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/72/8/2467/2458793.

Froese:2016:CBH

[FWP^{+}16a] Rainer Froese, Carl Walters, Daniel Pauly, Henning Winker, Olaf L. F. Weyl, Nazli Demirel, Athanassios C. Tsikliras, and Sidney J. Holt. A critique of the balanced harvesting approach to fishing. ICES Journal of Marine Science, 73(6):1640–1650, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
6/1640/2457873.

Froese:2016:RAA

[FWP^{+}16b] Rainer Froese, Carl Walters, Daniel Pauly, Henning Winker, Olaf L. F. Weyl, Nazli Demirel, Athanassios C. Tsikliras, and Sidney J. Holt. Reply to Andersen et al. (2016) “Assumptions behind size-based ecosystem models are realistic”. ICES Journal of Marine Science, 73(6):1656– 1658, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/6/1656/2458754. See [ABF^{+}16].

Furness:2012:ASS

[FWRM12] Robert W. Furness, Helen M. Wade, Alexandra M. C. Robbins, and Elizabeth A. Masden. Assessing the sensitivity of seabird populations to adverse effects from tidal stream turbines and wave energy devices. ICES Journal of Marine Science, 69(8):1466–1479, September 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/8/1466/704765.

Ferter:2013:UHC

[FWS^{+}13] Keno Ferter, Marc Simon Weltersbach, Harry Vincent Strehlow, Jon Helge Vølstad, Josep Alós, Robert Arlinghaus, Mike Armstrong, Malte Dorow, Martin de Graaf, Tessa van der Hammen, Kieran Hyder, Harold Levrel, Anton Paulrud, Krzysztof Radtke, Delphine Rocklin, Claus Reedtz Sparrevohn, and Pedro Veiga. Unexpectedly high catch-and-release rates in European marine recreational fisheries: implications for science and management. ICES Journal of Marine Science, 70(7):1319–1329, November 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/70/7/1319/609574.

Fielding:2014:IVA

[FWT^{+}14] Sophie Fielding, Jonathan L. Watkins, Philip N. Trathan, Peter Enderlein, Claire M. Waluda, Gabriele Stowasser, Geraint A. Tarling, and Eugene J. Murphy. Interannual variability in Antarctic krill (Euphausia superba) density at South Georgia, Southern Ocean: 1997–2013. ICES Journal of Marine Science, 71(9):2578–2588, November 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/9/2578/2798178.

Gullestad:2014:CAE

[GAB^{+}14] Peter Gullestad, Asgeir Aglen, Åsmund Bjordal, Geir Blom, Sverre Johansen, Jørn Krog, Ole Arve Misund, and Ingolf Røttingen. Changing attitudes 1970–2012: evolution of the Norwegian management framework to prevent overfishing and to secure long-term sustainability. ICES Journal of Marine Science, 71(2):173–182, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/2/173/778852.

Grant:2016:NDC

[GÁE16] W. Stewart Grant, Einar Árnason, and Bjarki Eldon. New DNA coalescent models and old population genetics software. ICES Journal of Marine Science, 73(9):2178–2180, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/73/9/2178/2198890.

Garcia:2011:PCI

[Gar11] Serge M. Garcia. Potential contribution of the Internet to a global community of practice for fishery management. ICES Journal of Marine Science, 68(8):1800–1804, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/68/8/1800/746089.

Geromont:2015:CAS

[GB15a] H. F. Geromont and D. S. Butterworth. Complex assessments or simple management procedures for efficient fisheries management: a comparative study. ICES Journal of Marine Science, 72(1):262–274, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/1/262/821583.

Geromont:2015:GMP

[GB15b] H. F. Geromont and D. S. Butterworth. Generic management procedures for data-poor fisheries: forecasting with few data. ICES Journal of Marine Science, 72(1):251–261, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/1/251/815189.

Gibbs:2015:RAR

[GB15c] Mark T. Gibbs and Howard I. Browman. Risk assessment and risk management: a primer for marine scientists. ICES Journal of Marine Science, 72(3):992–996, March 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/3/992/702340.

Gutzler:2015:CLD

[GBB15] Benjamin C. Gutzler, Mark J. Butler IV, and Donald C. Behringer. Casitas: a location-dependent ecological trap for juvenile Caribbean spiny lobsters, Panulirus argus. ICES Journal of Marine Science, 72(S1):S177–S184, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/72/suppl_1/i177/618966.

Gibbons:2016:WSA

[GBB16] M. J. Gibbons, F. Boero, and L. Brotz. We should not assume that fishing jellyfish will solve our jellyfish problem. ICES Journal of Marine Science, 73(4):1012–1018, March 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/4/1012/2458740.

Goodsir:2015:SRP

[GBJ^{+}15] Freya Goodsir, Helen J. Bloomfield, Adrian D. Judd, Filip Kral, Leonie A. Robinson, and Antony M. Knights. A spatially resolved pressure-based approach to evaluate combined effects of human activities and management in marine ecosystems. ICES Journal of Marine Science, 72(8):2245– 2256, September 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/72/8/2245/2458694.

Gnanalingam:2019:DAC

[GBM^{+}19] Gaya Gnanalingam, Mark J. Butler IV, Thomas R. Matthews, Emily Hutchinson, and Raouf Kilada. Directly ageing the Caribbean spiny lobster, Panulirus argus with validated band counts from gastric mill ossicles. ICES Journal of Marine Science, 76(2):442–451, March 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/2/442/5224784.

Geoffroy:2017:ABA

[GCBI17] Maxime Geoffroy, Finlo R. Cottier, Jørgen Berge, and Mark E. Inall. AUV-based acoustic observations of the distribution and patchiness of pelagic scattering layers during midnight sun. ICES Journal of Marine Science, 74(9):2342–2353, November 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
9/2342/2669574.

Garofalo:2010:NSG

[GCG^{+}10] Germana Garofalo, Luca Ceriola, Michele Gristina, Fabio Fiorentino, and Roberta Pace. Nurseries, spawning grounds and recruitment of Octopus vulgaris in the Strait of Sicily, central Mediterranean Sea. ICES Journal of Marine Science, 67(7):1363–1371, October 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/7/1363/664985.

Garcia-Carreras:2016:PRP

[GCJL16] Bernardo García-Carreras, Simon Jennings, and Will J. F. Le Quesne. Predicting reference points and associated uncertainty from life histories for risk and status assessment. ICES Journal of Marine Science, 73(2):483–493, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/2/483/2614399.

Greer:2018:EBH

[GCL^{+}18] Adam T. Greer, Luciano M. Chiaverano, Jessica Y. Luo, Robert K. Cowen, and William M. Graham. Ecology and behaviour of holoplanktonic scyphomedusae and their interactions with larval and juvenile fishes in the northern Gulf of Mexico. ICES Journal of Marine Science, 75(2): 751–763, March 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/2/751/4558893.

Guan:2017:ICS

[GCS^{+}17] Lisha Guan, Yong Chen, Kevin W. Staples, Jie Cao, and Bai Li. The influence of complex structure on the spatial dynamics of Atlantic cod (Gadus morhua) in the Gulf of Maine. ICES Journal of Marine Science, 74(9):2379–2388, November 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/9/2379/3769381.

Gilbey:2018:IFO

[GCS^{+}18] John Gilbey, Eef Cauwelier, James Sampayo, Iveta Matejusova, Charles Allan, Jennifer Graham, Lee Stradmeyer, and Stuart Middlemas. Identification of the farm of origin of Atlantic salmon smolt escapees in a freshwater Scottish loch using single-nucleotide polymorphic markers. ICES Journal of Marine Science, 75(6):2182–2192, November 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/6/2182/5046593.

Gilbey:2018:MBG

[GCW^{+}18] John Gilbey, Jamie Coughlan, Vidar Wennevik, Paulo Prodöhl, Jamie R. Stevens, Carlos Garcia de Leaniz, Dennis Ensing, Eef Cauwelier, Corrine Cherbonnel, Sofia Consuegra, Mark W. Coulson, Tom F. Cross, Walter Crozier, Eileen Dillane, Jonathan S. Ellis, Eva García-Vázquez, Andrew M. Griffiths, Sigurdur Gudjonsson, Kjetil Hindar, Sten Karlsson, David Knox, Gonzalo Machado-Schiaffino, Dorte Meldrup, Einar Eg Nielsen, Kristinn Ólafsson, Craig R. Primmer, Sergey Prusov, Lee Stradmeyer, Juha-Pekka Vähä, Alexey Je Veselov, Lucy M. I. Webster, Philip McGinnity, and Eric Verspoor. A microsatellite baseline for genetic stock identification of European Atlantic salmon (Salmo salar L.). ICES Journal of Marine Science, 75(2): 662–674, March 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/2/662/4191362.

Gorman:2015:RDT

[GD15] Daniel Gorman and Cameron Dixon. Reducing discards in a temperate prawn trawl fishery: a collaborative approach to bycatch research in South Australia. ICES Journal of Marine Science, 72(9):2609–2617, November 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/9/2609/2458734.

Gluchowska:2017:IZV

[GDBM^{+}17] Marta Gluchowska, Padmini Dalpadado, Agnieszka Beszczynska-Möller, Anna Olszewska, Randi B. Ingvaldsen, and Slawomir Kwasniewski. Interannual zooplankton variability in the main pathways of the Atlantic water flow into the Arctic Ocean (Fram Strait and Barents Sea branches). ICES Journal of Marine Science, 74(7):1921–1936, September 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/7/1921/3738489.

Gaudin:2018:MSB

[GDD^{+}18] François Gaudin, Nicolas Desroy, Stanislas F. Dubois, Caroline Broudin, Louis Cabioch, Jérôme Fournier, Franck Gentil, Jacques Grall, Céline Houbin, Patrick Le Mao, and Éric Thiébaut. Marine sublittoral benthos fails to track temperature in response to climate change in a biogeographical transition zone. ICES Journal of Marine Science, 75(6):1894–1907, November 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/75/
6/1894/5078582.

Gonzalez-Duran:2018:BAE

[GDHFS^{+}18] Enrique González-Durán, Alvaro Hernández-Flores, Juan Carlos Seijo, Alfonso Cuevas-Jiménez, and Angélica Moreno-Enriquez. Bioeconomics of the Allee effect in fisheries targeting sedentary resources. ICES Journal of Marine Science, 75(4):1362–1373, July 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/75/
4/1362/4920968.

Glover:2011:GIF

[GDJ11a] Kevin A. Glover, Geir Dahle, and Knut E. Jørstad. Genetic identification of farmed and wild Atlantic cod, Gadus morhua, in coastal Norway. ICES Journal of Marine Science, 68(5):901–910, May 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/68/5/901/654380.

Gnanadesikan:2011:WOB

[GDJ11b] A. Gnanadesikan, J. P. Dunne, and J. John. What ocean biogeochemical models can tell us about bottomup control of ecosystem variability. ICES Journal of Marine Science, 68(6):1030–1044, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/68/
6/1030/713106.

Goldstein:2015:RDA

[GDS15] Jason S. Goldstein, Elizabeth A. Dubofsky, and Ehud Spanier. Into a rhythm: diel activity patterns and behaviour in Mediterranean slipper lobsters, Scyllarides latus. ICES Journal of Marine Science, 72(S1):S147–S154, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/72/suppl_1/i147/617985.

Gerlotto:2017:SLY

[Ger17] François Gerlotto. Sixteen lessons from a 40-year quest to understand the mysterious life of the grey triggerfish. ICES Journal of Marine Science, 74(9):2321–2332, November 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/9/2321/3852699.

Grote:2012:CSG

[GES^{+}12] Britta Grote, Werner Ekau, Erling K. Stenevik, Catriona Clemmesen, Hans M. Verheye, Marek R. Lipinski, and Wilhelm Hagen. Characteristics of survivors: growth and nutritional condition of early stages of the hake species Merluccius paradoxus and M. capensis in the southern Benguela ecosystem. ICES Journal of Marine Science, 69(4):553– 562, May 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/69/4/553/633543.

Groger:2011:BSC

[GF11] Joachim P. Gröger and Michael J. Fogarty. Broad-scale climate influences on cod (Gadus morhua) recruitment on Georges Bank. ICES Journal of Marine Science, 68(3): 592–602, March 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/3/592/659944.

Griffith:2014:NAS

[GF14] Gary P. Griffith and Elizabeth A. Fulton. New approaches to simulating the complex interaction effects of multiple human impacts on the marine environment. ICES Journal of Marine Science, 71(4):764–774, May 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/4/764/669093.

Gaichas:2017:CSM

[GFF^{+}17] Sarah K. Gaichas, Michael Fogarty, Gavin Fay, Robert Gamble, Sean Lucey, and Laurel Smith. Combining stock, multispecies, and ecosystem level fishery objectives within an operational management procedure: simulations to start the conversation. ICES Journal of Marine Science, 74(2): 552–565, March 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/74/2/552/2669545.

Guo:2019:EBR

[GFF^{+}19] Chuanbo Guo, Caihong Fu, Robyn E. Forrest, Norm Olsen, Huizhu Liu, Philippe Verley, and Yunne-Jai Shin. Ecosystem-based reference points under varying plankton productivity states and fisheries management strategies. ICES Journal of Marine Science, 76(7):2045–2059, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/7/2045/5525266.

Griffiths:2010:AGL

[GFML10] Shane P. Griffiths, Gary C. Fry, Fiona J. Manson, and Dong C. Lou. Age and growth of longtail tuna (Thunnus tonggol) in tropical and temperate waters of the central Indo-Pacific. ICES Journal of Marine Science, 67(1):125– 134, January 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/67/1/125/595352.

Godwin:2011:TLA

[GFMO11] Rosamond M. Godwin, Stewart Frusher, Steven S. Montgomery, and Jennifer Ovenden. Telomere length analysis in crustacean species: Metapenaeus macleayi, Sagmariasus verreauxi, and Jasus edwardsii. ICES Journal of Marine Science, 68(10):2053–2058, November 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/10/2053/611301.

Greenstreet:2012:RMD

[GFR^{+}12] Simon P. R. Greenstreet, Helen M. Fraser, Stuart I. Rogers, Verena M. Trenkel, Stephen D. Simpson, and John K. Pinnegar. Redundancy in metrics describing the composition, structure, and functioning of the North Sea demersal fish community. ICES Journal of Marine Science, 69(1):8–22, January 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/69/1/8/674899.

Giraldes:2017:NZD

[GFSC17] Bruno Welter Giraldes, Petrônio Alves Coelho Filho, David M. Smyth, and Petrônio Alves Coelho. The nocturnal zonation of decapods in the subtidal zone within the reef seascape — abiotic factors defining habitats. ICES Journal of Marine Science, 74(8):2180–2190, September 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/8/2180/3749599.

Gasche:2013:ESM

[GG13] Loïc Gasche and Didier Gascuel. EcoTroph: a simple model to assess fishery interactions and their impacts on ecosystems. ICES Journal of Marine Science, 70(3):498–510, April 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/70/3/498/917462.

Greathead:2015:ERT

[GGIC^{+}15] C. Greathead, J. M. González-Irusta, J. Clarke, P. Boulcott, L. Blackadder, A. Weetman, and P. J. Wright. Environmental requirements for three sea pen species: relevance to distribution and conservation. ICES Journal of Marine Science, 72(2):576–586, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/2/576/675444.

Graham:2011:IPS

[GGK^{+}11] Norman Graham, Richard Grainger, William A. Karp, David N. MacLennan, Philip MacMullen, and Kjell Nedreaas. An introduction to the proceedings and a synthesis of the 2010 ICES Symposium on Fishery-Dependent Information. ICES Journal of Marine Science, 68(8):1593– 1597, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/8/1593/763960.

Gascuel:2011:TLB

[GGP11] Didier Gascuel, Sylvie Guénette, and Daniel Pauly. The trophic-level-based ecosystem modelling approach: theoretical overview and practical uses. ICES Journal of Marine Science, 68(7):1403–1416, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/68/
7/1403/656418.

Garibaldi:2014:CMF

[GGT^{+}14] Luca Garibaldi, Jennifer Gee, Sachiko Tsuji, Piero Mannini, and David Currie. Comment on: “Managing fisheries from space: Google Earth improves estimates of distant fish catches” by Al-Abdulrazzak and Pauly. ICES Journal of Marine Science, 71(7):1921–1926, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/7/1921/2804454. See [AAP14b, AAP14a].

Gallego:2017:BPC

[GGTW17] Alejandro Gallego, Fiona M. Gibb, David Tullet, and Peter J. Wright. Bio-physical connectivity patterns of benthic marine species used in the designation of Scottish nature conservation marine protected areas. ICES Journal of Marine Science, 74(6):1797–1811, July 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/6/1797/2726868.

Goyert:2018:EHP

[GGV^{+}18] Holly F. Goyert, Beth Gardner, Richard R. Veit, Andrew T. Gilbert, Emily Connelly, Melissa Duron, Sarah Johnson, and Kathryn Williams. Evaluating habitat, prey, and mesopredator associations in a community of marine birds. ICES Journal of Marine Science, 75(5):1602–1612, September 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/5/1602/4973731.

Graham:2010:AIE

[GHB^{+}10] Tanya R. Graham, James T. Harvey, Scott R. Benson, Josiah S. Renfree, and David A. Demer. The acoustic identification and enumeration of scyphozoan jellyfish, prey for leatherback sea turtles (Dermochelys coriacea), off central California. ICES Journal of Marine Science, 67(8):1739– 1748, November 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/67/8/1739/608506.

Godo:2014:MEA

[GHB^{+}14] Olav Rune Godø, Nils Olav Handegard, Howard I. Browman, Gavin J. Macaulay, Stein Kaartvedt, Jarl Giske, Egil Ona, Geir Huse, and Espen Johnsen. Marine ecosystem acoustics (MEA): quantifying processes in the sea at the spatio-temporal scales on which they occur. ICES Journal of Marine Science, 71(8):2357–2369, October 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/8/2357/761870.

Govoni:2010:MCE

[GHD^{+}10] J. J. Govoni, J. A. Hare, E. D. Davenport, M. H. Chen, and K. E. Marancik. Mesoscale, cyclonic eddies as larval fish habitat along the southeast United States shelf: a Lagrangian description of the zooplankton community. ICES Journal of Marine Science, 67(3):403–411, April 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/3/403/733500.

Geffen:2011:HLC

[GHF^{+}11] Audrey J. Geffen, Hans Høie, Arild Folkvord, Anne Karin Hufthammer, Carin Andersson, Ulysses Ninnemann, Rolf B. Pedersen, and Kjell Nedreaas. High-latitude climate variability and its effect on fisheries resources as revealed by fossil cod otoliths. ICES Journal of Marine Science, 68(6): 1081–1089, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/6/1081/699699.

Greenstreet:2010:CHS

[GHG^{+}10] Simon P. R. Greenstreet, Gayle J. Holland, Emma J. Guirey, Eric Armstrong, Helen M. Fraser, and Iain M. Gibb. Combining hydroacoustic seabed survey and grab sampling techniques to assess “local” sandeel population abundance. ICES Journal of Marine Science, 67(5):971– 984, July 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/67/5/971/608540.

Gjosaeter:2016:PEL

[GHM^{+}16] Harald Gjøsæter, Elvar H. Hallfredsson, Nina Mikkelsen, Bjarte Bogstad, and Torstein Pedersen. Predation on early life stages is decisive for year-class strength in the Barents Sea capelin (Mallotus villosus) stock. ICES Journal of Marine Science, 73(2):182–195, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/2/182/2614260.

Gardner:2015:FEL

[GHPH15] Caleb Gardner, Klaas Hartmann, André E. Punt, and Eriko Hoshino. Fewer eggs from larger size limits: counterintuitive outcomes in a spatially heterogeneous fishery. ICES Journal of Marine Science, 72(S1):S252–S259, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/72/suppl_1/i252/613301.

Grimaldo:2018:FEB

[GHV^{+}18] Eduardo Grimaldo, Bent Herrmann, Jørgen Vollstad, Biao Su, Heidi Moe Føre, Roger B. Larsen, and Ivan Tatone. Fishing efficiency of biodegradable PBSAT gillnets and conventional nylon gillnets used in Norwegian cod (Gadus morhua) and saithe (Pollachius virens) fisheries. ICES Journal of Marine Science, 75(6):2245–2256, November 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/6/2245/5094487.

Galgani:2013:MLW

[GHWD13] F. Galgani, G. Hanke, S. Werner, and L. De Vrees. Marine litter within the European marine strategy framework directive. ICES Journal of Marine Science, 70(6):1055– 1064, September 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/70/6/1055/639375.

Gonzalez-Irusta:2018:DMS

[GIDP^{+}18] Jose M. González-Irusta, Ana De la Torriente, Antonio Punzón, Marian Blanco, and Alberto Serrano. Determining and mapping species sensitivity to trawling impacts: the BEnthos Sensitivity Index to Trawling Operations (BE-SITO). ICES Journal of Marine Science, 75(5):1710– 1721, September 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/5/1710/4955966.

Gonzalez-Irusta:2012:EFE

[GIPS12] J. M. González-Irusta, A. Punzón, and A. Serrano. Environmental and fisheries effects on Gracilechinus acutus (Echinodermata: Echinoidea) distribution: is it a suitable bioindicator of trawling disturbance? ICES Journal of Marine Science, 69(8):1457–1465, September 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/8/1457/700830.

Gregory:2019:ASR

[GIR^{+}19] Stephen D. Gregory, Anton T. Ibbotson, William D. Riley, Marie Nevoux, Rasmus B. Lauridsen, Ian C. Russell, J. Robert Britton, Phillipa K. Gillingham, Olivia M. Simmons, and Etienne Rivot. Atlantic salmon return rate increases with smolt length. ICES Journal of Marine Science, 76(6):1702–1712, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
6/1702/5481490.

Gonzalez-Irusta:2016:SGA

[GIW16] José M. González-Irusta and Peter J. Wright. Spawning grounds of Atlantic cod (Gadus morhua) in the North Sea. ICES Journal of Marine Science, 73(2):304–315, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/2/304/2614292.

Guan:2019:SME

[GJWS19] Lisha Guan, Xianshi Jin, Qiang Wu, and Xiujuan Shan. Statistical modelling for exploring diel vertical movements and spatial correlations of marine fish species: a supplementary tool to assess species interactions. ICES Journal of Marine Science, 76(6):1776–1783, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/6/1776/5430864.

Green:2015:DCC

[GKC^{+}15] David B. Green, Norbert T. W. Klages, Robert J. M. Crawford, Janet C. Coetzee, Bruce M. Dyer, Gavin M. Rishworth, and Pierre A. Pistorius. Dietary change in Cape gannets reflects distributional and demographic shifts in two South African commercial fish stocks. ICES Journal of Marine Science, 72(3):771–781, March 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/3/771/698294.

Gargan:2016:SLL

[GKC^{+}16] Patrick Gargan, Egil Karlsbakk, John Coyne, Carys Davies, and William Roche. Sea lice (Lepeophtheirus salmonis and Caligus elongatus) infestation levels on sea trout (Salmo trutta L.) around the Irish Sea, an area without salmon aquaculture. ICES Journal of Marine Science, 73(9):2395– 2407, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/73/9/2395/2198216.

Gonzalez:2015:GAL

[GKJ^{+}15] Enrique Blanco Gonzalez, Halvor Knutsen, Per Erik Jorde, Kevin Alan Glover, and Odd Aksel Bergstad. Genetic analyses of ling (Molva molva) in the Northeast Atlantic reveal patterns relevant to stock assessments and management advice. ICES Journal of Marine Science, 72(2):635–641, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/2/635/2801344.

Groger:2010:SRH

[GKR10] Joachim P. Gröger, Gordon H. Kruse, and Norbert Rohlf. Slave to the rhythm: how large-scale climate cycles trigger herring (Clupea harengus) regeneration in the North Sea. ICES Journal of Marine Science, 67(3):454–465, April 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/3/454/732742.

Gruss:2014:EEM

[GKR14] Arnaud Grüss, David M. Kaplan, and Jan Robinson. Evaluation of the effectiveness of marine reserves for transient spawning aggregations in data-limited situations. ICES Journal of Marine Science, 71(3):435–449, March 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/3/435/631071.

Geist:2015:DFB

[GKV^{+}15] S. J. Geist, A. Kunzmann, H. M. Verheye, A. Eggert, A. Schukat, and W. Ekau. Distribution, feeding behaviour, and condition of Cape horse mackerel early life stages, Trachurus capensis, under different environmental conditions in the northern Benguela upwelling ecosystem. ICES Journal of Marine Science, 72(2):543–557, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/2/543/670986.

Gerritsen:2011:IVM

[GL11] Hans Gerritsen and Colm Lordan. Integrating Vessel Monitoring Systems (VMS) data with daily catch data from logbooks to explore the spatial distribution of catch and effort at high resolution. ICES Journal of Marine Science, 68(1): 245–252, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/1/245/628374.

Goethel:2015:DSE

[GLC15] Daniel R. Goethel, Christopher M. Legault, and Steven X. Cadrin. Demonstration of a spatially explicit, tagintegrated stock assessment model with application to three interconnected stocks of yellowtail flounder off of New England. ICES Journal of Marine Science, 72(1):164–177, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/1/164/817226.

Gaichas:2014:RBA

[GLH14] S. K. Gaichas, J. S. Link, and J. A. Hare. A riskbased approach to evaluating northeast US fish community vulnerability to climate change. ICES Journal of Marine Science, 71(8):2323–2342, October 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/8/2323/758638.

Garrison:2010:EMA

[GLK^{+}10] Lance P. Garrison, Jason S. Link, D. Patrick Kilduff, Matthew D. Cieri, Brandon Muffley, Douglas S. Vaughan, Alexei Sharov, Behzad Mahmoudi, and Robert J. Latour. An expansion of the MSVPA approach for quantifying predator–prey interactions in exploited fish communities. ICES Journal of Marine Science, 67(5):856–870, July 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/67/5/856/610180.

Goetz:2011:EFU

[GLP^{+}11] Sabine Goetz, Martín Laporta, Julio Martínez Portela, M. Begoña Santos, and Graham J. Pierce. Experimental fishing with an “umbrella-and-stones” system to reduce interactions of sperm whales (Physeter macrocephalus) and seabirds with bottom-set longlines for Patagonian toothfish (Dissostichus eleginoides) in the Southwest Atlantic. ICES Journal of Marine Science, 68(1):228–238, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/1/228/632600.

Glandon:2017:NEH

[GM17] Hillary L. Glandon and Thomas J. Miller. No effect of high pCO_{2} on juvenile blue crab, Callinectes sapidus, growth and consumption despite positive responses to concurrent warming. ICES Journal of Marine Science, 74(4):1201– 1209, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/4/1201/2399215.

Goodman:2018:LBB

[GM18] Camille Goodman and Holly Matley. Law beyond boundaries: innovative mechanisms for the integrated management of biodiversity beyond national jurisdiction. ICES Journal of Marine Science, 75(1):402–404, January 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/1/402/4796915.

Gertseva:2017:SGV

[GMC17] Vladlena Gertseva, Sean E. Matson, and Jason Cope. Spatial growth variability in marine fish: example from Northeast Pacific groundfish. ICES Journal of Marine Science, 74(6):1602–1613, July 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
6/1602/3003308.

Garcia:2011:MPS

[GMD^{+}11] Alexandra Garcia, Simonetta Mattiucci, Simone Damiano, Miguel N. Santos, and Giuseppe Nascetti. Metazoan parasites of swordfish, Xiphias gladius (Pisces: Xiphiidae) from the Atlantic Ocean: implications for host stock identification. ICES Journal of Marine Science, 68(1):175–182, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/68/1/175/629590.

Geraghty:2014:AGP

[GMH^{+}14] Pascal T. Geraghty, William G. Macbeth, Alastair V. Harry, Jacqueline E. Bell, Michelle N. Yerman, and Jane E. Williamson. Age and growth parameters for three heavily exploited shark species off temperate eastern Australia. ICES Journal of Marine Science, 71(3):559–573, March 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/3/559/636691.

Grant:2011:LAH

[GMKS11] W. Stewart Grant, Susan E. Merkouris, Gordon H. Kruse, and Lisa W. Seeb. Low allozyme heterozygosity in North Pacific and Bering Sea populations of red king crab (Paralithodes camtschaticus): adaptive specialization, population bottleneck, or metapopulation structure? ICES Journal of Marine Science, 68(3):499–506, March 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/3/499/658194.

Gerritsen:2013:HMS

[GML13] Hans D. Gerritsen, Cóilín Minto, and Colm Lordan. How much of the seabed is impacted by mobile fishing gear? Absolute estimates from Vessel Monitoring System (VMS) point data. ICES Journal of Marine Science, 70(3):523– 531, April 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/70/3/523/918433.

Gomez:2012:TZS

[GMM^{+}12] May Gómez, Ico Martínez, Ismael Mayo, José Miguel Morales, Angelo Santana, and Ted T. Packard. Testing zooplankton secondary production models against Daphnia magna growth. ICES Journal of Marine Science, 69(3): 421–428, May 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/69/3/421/600074.

Ganias:2017:RFT

[GMN^{+}17] Konstantinos Ganias, Foivos-Alexandros Mouchlianitis, Cristina Nunes, Ana-Maria Costa, and Maria-Manuel Angélico. A reassessment of the fecundity type of Atlantic horse mackerel (Trachurus trachurus) in Atlantic Iberian waters (ICES division IXa) shows that indeterminate spawners can cease recruiting oocytes during their spawning season. ICES Journal of Marine Science, 74(1): 31–40, January 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/74/1/31/2669562.

Ganias:2018:AMD

[GMS^{+}18] K. Ganias, D. Marmara, A. Solla, D. Garabana, and R. Dominguez-Petit. Atlantic mackerel daily spawning dynamics and implications for batch fecundity estimations. ICES Journal of Marine Science, 75(5):1647–1654, September 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/5/1647/4961371.

Gogina:2016:BSS

[GNB^{+}16] Mayya Gogina, Henrik Nygård, Mats Blomqvist, Darius Daunys, Alf B. Josefson, Jonne Kotta, Alexey Maximov, Jan Warzocha, Vadim Yermakov, Ulf Gräwe, and Michael L. Zettler. The Baltic Sea scale inventory of benthic faunal communities. ICES Journal of Marine Science, 73(4):1196–1213, March 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
4/1196/2458890.

Geffen:2011:CHP

[GNDC11] Audrey J. Geffen, Richard D. M. Nash, and Mark Dickey-Collas. Characterization of herring populations west of the British Isles: an investigation of mixing based on otolith microchemistry. ICES Journal of Marine Science, 68(7):

1447–1458, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/7/1447/653737.

Gaardmark:2011:DMF

[GNFM11] Anna Gårdmark, Anders Nielsen, Jens Floeter, and Christian Möllmann. Depleted marine fish stocks and ecosystembased management: on the road to recovery, we need to be precautionary. ICES Journal of Marine Science, 68(1): 212–220, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/1/212/631725.

Geladakis:2018:MDP

[GNKS18] George Geladakis, Nikolaos Nikolioudakis, George Koumoundouros, and Stylianos Somarakis. Morphometric discrimination of pelagic fish stocks challenged by variation in body condition. ICES Journal of Marine Science, 75(2):711–718, March 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/2/711/4157898.

Gilby:2019:SCM

[GOH^{+}19] Ben L. Gilby, Andrew D. Olds, Christopher J. Henderson, Nicholas L. Ortodossi, Rod M. Connolly, and Thomas A. Schlacher. Seascape context modifies how fish respond to restored oyster reef structures. ICES Journal of Marine Science, 76(4):1131–1139, July 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/4/1131/5333156.

Gaardmark:2012:DPG

[GÖN^{+}12] Anna Gårdmark, Örjan Östman, Anders Nielsen, Karl Lundström, Olle Karlsson, Jukka Pönni, and Teija Aho. Does predation by grey seals (Halichoerus grypus) affect Bothnian Sea herring stock estimates? ICES Journal of Marine Science, 69(8):1448–1456, September 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/8/1448/700008.

Gonzalez:2010:AGM

[GOPG10] Ángel F. González, Jaime Otero, Graham J. Pierce, and Ángel Guerra. Age, growth, and mortality of Loligo vulgaris wild paralarvae: implications for understanding of the life cycle and longevity. ICES Journal of Marine Science, 67(6):1119–1127, September 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/6/1119/733836.

Grohsler:2013:DWB

[GOS^{+}13] Tomas Gröhsler, Rainer Oeberst, Matthias Schaber, Niklas Larson, and Georgs Kornilovs. Discrimination of western Baltic spring-spawning and central Baltic herring (Clupea harengus L.) based on growth vs. natural tag information. ICES Journal of Marine Science, 70(6):1108–1117, September 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/70/6/1108/637556.

Gislason:2014:LTC

[GPG14] Astthor Gislason, Hildur Petursdottir, and Kristinn Gudmundsson. Long-term changes in abundance of Calanus finmarchicus south and north of Iceland in relation to environmental conditions and regional diversity in spring 1990– 2013. ICES Journal of Marine Science, 71(9):2539–2549, November 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/71/9/2539/2798175.

Garner:2017:EAC

[GPP17] Steven B. Garner, William F. Patterson III, and Clay E. Porch. Experimental assessment of circle vs. J hook performance and selectivity in the northern Gulf of Mexico recreational reef fish fishery. ICES Journal of Marine Science, 74(5):1437–1447, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
5/1437/2977677.

Garcia:2017:BMR

[GPS^{+}17] Dorleta García, Raúl Prellezo, Paz Sampedro, José María Da-Rocha, José Castro, Santiago Cerviño, Javier García-Cutrín, and María-José Gutiérrez. Bioeconomic multistock reference points as a tool for overcoming the drawbacks of the landing obligation. ICES Journal of Marine Science, 74(2):511–524, March 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
2/511/2669541.

Gao:2019:FCI

[GQX^{+}19] Guang Gao, Liming Qu, Tianpeng Xu, J. Grant Burgess, Xinshu Li, and Juntian Xu. Future CO_{2}-induced ocean acidification enhances resilience of a green tide alga to low-salinity stress. ICES Journal of Marine Science, 76(7):2437–2445, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
7/2437/5532214.

Gruss:2015:FPF

[GR15] Arnaud Grüss and Jan Robinson. Fish populations forming transient spawning aggregations: should spawners always be the targets of spatial protection efforts? ICES Journal of Marine Science, 72(2):480–497, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/2/480/2801346.

Gras:2014:TSB

[GRC^{+}14] Michaël Gras, Beatriz A. Roel, Franck Coppin, Eric Foucher, and Jean-Paul Robin. A two-stage biomass model to assess the English Channel cuttlefish (Sepia officinalis L.) stock. ICES Journal of Marine Science, 71(9):2457– 2468, November 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/71/9/2457/594946.

Garcia:2016:BHF

[GRC16a] S. M. Garcia, J. Rice, and A. Charles. Balanced harvesting in fisheries: a preliminary analysis of management implications. ICES Journal of Marine Science, 73(6):1668–1678, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/6/1668/2457918.

Garcia:2016:BFM

[GRC16b] S. M. Garcia, J. Rice, and A. Charles. Bridging fisheries management and biodiversity conservation norms: potential and challenges of balancing harvest in ecosystem-based frameworks. ICES Journal of Marine Science, 73(6):1659– 1667, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/6/1659/2458727.

Greenstreet:2012:DFB

[GRF^{+}12] Simon P. R. Greenstreet, Axel G. Rossberg, Clive J. Fox, William J. F. Le Quesne, Tom Blasdale, Philip Boulcott, Ian Mitchell, Colin Millar, and Colin F. Moffat. Demersal fish biodiversity: species-level indicators and trendsbased targets for the marine strategy framework directive. ICES Journal of Marine Science, 69(10):1789–1801, December 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/69/10/1789/622820.

Graham:2016:SMF

[GRF^{+}16] Helen Graham, Samuel P. S. Rastrick, Helen S. Findlay, Matthew G. Bentley, Stephen Widdicombe, Anthony S. Clare, and Gary S. Caldwell. Sperm motility and fertilisation success in an acidified and hypoxic environment. ICES Journal of Marine Science, 73(3):783–790, February 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/3/783/2458750.

Guerra:2010:MGE

[GRG^{+}10] Ángel Guerra, Álvaro Roura, Ángel F. González, Santiago Pascual, Yves Cherel, and Marcos Pérez-Losada. Morphological and genetic evidence that Octopus vulgaris Cuvier, 1797 inhabits Amsterdam and Saint Paul Islands (southern Indian Ocean). ICES Journal of Marine Science, 67(7):1401–1407, October 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/67/
7/1401/662355.

Gruss:2014:CFE

[GRH^{+}14] Arnaud Grüss, Jan Robinson, Selina S. Heppell, Scott A. Heppell, and Brice X. Semmens. Conservation and fisheries effects of spawning aggregation marine protected areas: What we know, where we should go, and what we need to get there. ICES Journal of Marine Science, 71(7):1515– 1534, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/71/7/1515/672201.

Geoffroy:2016:TSE

[GRKF16] Maxime Geoffroy, Shani Rousseau, Frank Reier Knudsen, and Louis Fortier. Target strengths and echotraces of whales and seals in the Canadian Beaufort Sea. ICES Journal of Marine Science, 73(2):451–463, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/2/451/2614313.

Guerra:2010:LHT

[GRNG^{+}10] Ángel Guerra, Alejandro B. Rodríguez-Navarro, Ángel F. González, Chris S. Romanek, Pedro Álvarez-Lloret, and Graham J. Pierce. Life-history traits of the giant squid Architeuthis dux revealed from stable isotope signatures recorded in beaks. ICES Journal of Marine Science, 67(7):1425–1431, October 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/67/
7/1425/664168.

Groeneveld:2011:QFC

[Gro11] Rolf A. Groeneveld. Quantifying fishers’ and citizens’ support for Dutch flatfish management policy. ICES Journal of Marine Science, 68(5):919–928, May 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/5/919/646167.

Gasalla:2010:TRS

[GRP10] Maria A. Gasalla, Amanda R. Rodrigues, and Felippe A. Postuma. The trophic role of the squid Loligo plei as a keystone species in the South Brazil Bight ecosystem.

ICES Journal of Marine Science, 67(7):1413–1424, October 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/7/1413/665990.

Greenstreet:2011:DEN

[GRR^{+}11] Simon P. R. Greenstreet, Stuart I. Rogers, Jake C. Rice, Gerjan J. Piet, Emma J. Guirey, Helen M. Fraser, and Rob J. Fryer. Development of the EcoQO for the North Sea fish community. ICES Journal of Marine Science, 68 (1):1–11, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/1/1/631462.

Greenstreet:2012:RTN

[GRR^{+}12] Simon P. R. Greenstreet, Stuart I. Rogers, Jake C. Rice, Gerjan J. Piet, Emma J. Guirey, Helen M. Fraser, and Rob J. Fryer. A reassessment of trends in the North Sea large fish indicator and a re-evaluation of earlier conclusions. ICES Journal of Marine Science, 69(2):343–345, March 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/69/2/343/701317.

Goetz:2014:CFI

[GRS^{+}14] Sabine Goetz, Fiona L. Read, M. Begoña Santos, C. Pita, and Graham J. Pierce. Cetacean–fishery interactions in Galicia (NW Spain): results and management implications of a face-to-face interview survey of local fishers. ICES Journal of Marine Science, 71(3):604–617, March 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/3/604/634499.

Goldman:2011:FHS

[GS11] Sarah F. Goldman and George R. Sedberry. Feeding habits of some demersal fish on the Charleston bump off the southeastern United States. ICES Journal of Marine Science, 68(2):390–398, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/68/2/390/614008.

Gislason:2012:ACD

[GS12] Astthor Gislason and Teresa Silva. Abundance, composition, and development of zooplankton in the Subarctic Iceland Sea in 2006, 2007, and 2008. ICES Journal of Marine Science, 69(7):1263–1276, September 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/7/1263/743627.

Granneman:2014:FGR

[GS14] Jennifer E. Granneman and Mark A. Steele. Fish growth, reproduction, and tissue production on artificial reefs relative to natural reefs. ICES Journal of Marine Science, 71(9):2494–2504, November 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/9/2494/595448.

Granneman:2015:ERA

[GS15] Jennifer E. Granneman and Mark A. Steele. Effects of reef attributes on fish assemblage similarity between artificial and natural reefs. ICES Journal of Marine Science, 72(8):2385–2397, September 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/72/
8/2385/2458748.

Grant:2010:SBS

[GSC10] W. Stewart Grant, Ingrid Spies, and Michael F. Canino. Shifting-balance stock structure in North Pacific walleye pollock (Gadus chalcogrammus). ICES Journal of Marine Science, 67(8):1687–1696, November 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/67/
8/1687/605882.

Glover:2011:MDR

[GSL^{+}11] Kevin A. Glover, Øystein Skaala, Morten Limborg, Cecilie Kvamme, and Else Torstensen. Microsatellite DNA reveals population genetic differentiation among sprat (Sprattus sprattus) sampled throughout the Northeast Atlantic, including Norwegian fjords. ICES Journal of Marine Science, 68(10):2145–2151, November 2011.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/10/2145/614425.

Godfrey:2015:DUM

[GSMA15] Jason D. Godfrey, David C. Stewart, Stuart J. Middlemas, and John D. Armstrong. Depth use and migratory behaviour of homing Atlantic salmon (Salmo salar) in Scottish coastal waters. ICES Journal of Marine Science, 72(2): 568–575, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/72/2/568/2801299.

Gonzalez-Silvera:2016:AFH

[GSMRM^{+}16] D. Gonzalez-Silvera, L. Martinez-Rubio, M. E. Abad Mateo, R. Rabadan-Ros, J. A. López Jiménez, and F. J. Martínez López. Assessing feeding history and health status through analysis of fatty acids and fat content in golden mullet Liza aurata. ICES Journal of Marine Science, 73(10):2632–2643, November 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/10/2632/2647094.

Gaston:2012:DCI

[GSP12] Anthony J. Gaston, Paul A. Smith, and Jennifer F. Provencher. Discontinuous change in ice cover in Hudson Bay in the 1990s and some consequences for marine birds and their prey. ICES Journal of Marine Science, 69(7):1218–1225, September 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/69/
7/1218/741796.

Gruss:2019:DST

[GT19] Arnaud Grüss and James T. Thorson. Developing spatiotemporal models using multiple data types for evaluating population trends and habitat usage. ICES Journal of Marine Science, 76(6):1748–1761, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/6/1748/5479986.

Gruss:2018:PDM

[GTBT18] Arnaud Grüss, James T. Thorson, Elizabeth A. Babcock, and Joseph H. Tarnecki. Producing distribution maps for informing ecosystem-based fisheries management using a comprehensive survey database and spatio-temporal models. ICES Journal of Marine Science, 75(1):158–177, January 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/1/158/3958179.

Guinet:2015:LTS

[GTGD15] Christophe Guinet, Paul Tixier, Nicolas Gasco, and Guy Duhamel. Long-term studies of crozet Island killer whales are fundamental to understanding the economic and demographic consequences of their depredation behaviour on the Patagonian toothfish fishery. ICES Journal of Marine Science, 72(5):1587–1597, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/5/1587/761039.

Garcia:2012:SLV

[GTM12] María Jesús García, Elena Tel, and Joaquín Molinero. Sea-level variations on the north and northwest coasts of Spain. ICES Journal of Marine Science, 69(5):720–727, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/69/5/720/652001.

Glover:2019:DER

[GUN^{+}19] K. A. Glover, K. Urdal, T. Næsje, H. Skoglund, B. Florø-Larsen, H. Otterå, P. Fiske, M. Heino, T. Aronsen, H. Sægrov, O. Diserud, B. T. Barlaup, K. Hindar, G. Bakke, I. Solberg, H. Lo, M. F. Solberg, S. Karlsson, Ø. Skaala, A. Lamberg, Ø. Kanstad-Hanssen, R. Muladal, O. T. Skilbrei, and V. Wennevik. Domesticated escapees on the run: the second-generation monitoring programme reports the numbers and proportions of farmed Atlantic salmon in >200 Norwegian rivers annually. ICES Journal of Marine Science, 76(4):1151–1161, July 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/4/1151/5303246.

Georget:2019:UCA

[GVA19] Stéphane Georget, Simon Van Wynsberge, and Serge Andréfouët. Understanding consequences of adaptive monitoring protocols on data consistency: application to the monitoring of giant clam densities impacted by massive mortalities in tuamotu atolls, French Polynesia. ICES Journal of Marine Science, 76(4):1062–1071, July 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/4/1062/5299885.

Goldberg:2019:DLM

[GvRKB19] Daphna Shapiro Goldberg, Itai van Rijn, Moshe Kiflawi, and Jonathan Belmaker. Decreases in length at maturation of Mediterranean fishes associated with higher sea temperatures. ICES Journal of Marine Science, 76(4):946–959, July 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/4/946/5310316.

Greer:2016:APP

[GW16] Adam T. Greer and C. Brock Woodson. Application of a predator–prey overlap metric to determine the impact of sub-grid scale feeding dynamics on ecosystem productivity. ICES Journal of Marine Science, 73(4):1051–1061, March 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/4/1051/2458966.

Gilman:2017:EDO

[GWS17] Eric Gilman, Mariska Weijerman, and Petri Suuronen. Ecological data from observer programmes underpin ecosystem-based fisheries management. ICES Journal of Marine Science, 74(6):1481–1495, July 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
6/1481/3098076.

Garcia:2019:TSS

[GZS^{+}19] Heriberto A. Garcia, Chenyang Zhu, Matthew E. Schinault, Anna I. Kaplan, Nils Olav Handegard, Olav Rune Godø, Heidi Ahonen, Nicholas C. Makris, Delin Wang, Wei Huang, and Purnima Ratilal. Temporal–spatial, spectral, and source level distributions of fin whale vocalizations in the Norwegian Sea observed with a coherent hydrophone array. ICES Journal of Marine Science, 76(1):268–283, January 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/1/268/5115402.

Harbitz:2015:PSD

[HA15] Alf Harbitz and Ole Thomas Albert. Pitfalls in stock discrimination by shape analysis of otolith contours. ICES Journal of Marine Science, 72(7):2090–2097, October 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/7/2090/2457856.

Haltuch:2019:AEC

[HABV19] Melissa A. Haltuch, Z. Teresa A’mar, Nicholas A. Bond, and Juan L. Valero. Assessing the effects of climate change on US West Coast sablefish productivity and on the performance of alternative management strategies. ICES Journal of Marine Science, 76(6):1524–1542, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/6/1524/5382129.

Hess:2016:DAM

[HAF^{+}16] Jon E. Hess, Michael W. Ackerman, Jeffrey K. Fryer, Daniel J. Hasselman, Craig A. Steele, Jeff J. Stephenson, John M. Whiteaker, and Shawn R. Narum. Differential adult migration-timing and stock-specific abundance of steelhead in mixed stock assemblages. ICES Journal of Marine Science, 73(10):2606–2615, November 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/10/2606/2647122.

Haug:2014:IMH

[HAH^{+}14] Tore Haug, Michaela Aschan, Alf Håkon Hoel, Torild Johansen, and Jan H. Sundet. Introduction: Marine harvesting in the Arctic. ICES Journal of Marine Science, 71(7):1932–1933, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/71/
7/1932/673373.

Houde:2016:FAA

[HAH^{+}16] Edward D. Houde, Eric R. Annis, Lawrence W. Harding, Jr., Michael E. Mallonee, and Michael J. Wilberg. Factors affecting the abundance of age-0 Atlantic menhaden (Brevoortia tyrannus) in Chesapeake Bay. ICES Journal of Marine Science, 73(9):2238–2251, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/9/2238/2198389.

Hamoutene:2014:SSR

[Ham14] Dounia Hamoutene. Sediment sulphides and redox potential associated with spatial coverage of Beggiatoa spp. at finfish aquaculture sites in Newfoundland, Canada. ICES Journal of Marine Science, 71(5):1153–1157, July 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/5/1153/646119.

Hamel:2015:MCM

[Ham15] Owen S. Hamel. A method for calculating a meta-analytical prior for the natural mortality rate using multiple life history correlates. ICES Journal of Marine Science, 72(1): 62–69, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/72/1/62/2804304.

Hannesson:2013:SNA

[Han13] Rögnvaldur Hannesson. Sharing the Northeast Atlantic mackerel. ICES Journal of Marine Science, 70(2):259–269, March 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/70/2/259/793656.

Hart:2013:QTB

[Har13] Deborah R. Hart. Quantifying the tradeoff between precaution and yield in fishery reference points. ICES Journal of Marine Science, 70(3):591–603, April 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/3/591/914936.

Hare:2014:FFO

[Har14] Jonathan A. Hare. The future of fisheries oceanography lies in the pursuit of multiple hypotheses. ICES Journal of Marine Science, 71(8):2343–2356, October 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/8/2343/752776.

Hintzen:2019:PFS

[HAR19] Niels T. Hintzen, Geert Aarts, and Adriaan D. Rijnsdorp. Persistence in the fine-scale distribution and spatial aggregation of fishing. ICES Journal of Marine Science, 76(4): 1072–1082, July 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/76/4/1072/5232289.

Howell:2010:CGF

[HB10] Daniel Howell and Bjarte Bogstad. A combined Gadget/FLR model for management strategy evaluations of the Barents Sea fisheries. ICES Journal of Marine Science, 67(9):1998–2004, December 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/67/
9/1998/623743.

Hamilton:2019:PGE

[HB19a] Sheryl Hamilton and G. Barry Baker. Population growth of an endangered pinniped — the New Zealand sea lion (Phocarctos hookeri) — is limited more by high pup mortality than fisheries bycatch. ICES Journal of Marine Science, 76(6):1794–1806, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/6/1794/5421315.

Hidalgo:2019:DKB

[HB19b] Manuel Hidalgo and Howard I. Browman. Developing the knowledge base needed to sustainably manage mesopelagic resources. ICES Journal of Marine Science, 76(3):609–615, May 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/3/609/5486347.

Harris:2013:SBG

[HBB^{+}13a] Peter T. Harris, Thomas C. L. Bridge, Robin J. Beaman, Jody M. Webster, Scott L. Nichol, and Brendan P. Brooke. Submerged banks in the Great Barrier Reef, Australia, greatly increase available coral reef habitat. ICES Journal of Marine Science, 70(2):284–293, March 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/2/284/794080.

Heino:2013:CFI

[HBB^{+}13b] Mikko Heino, Loïc Baulier, David S. Boukal, Bruno Ernande, Fiona D. Johnston, Fabian M. Mollet, Heidi Pardoe, Nina O. Therkildsen, Silva Uusi-Heikkilä, Anssi Vainikka, Robert Arlinghaus, Dorothy J. Dankel, Erin S. Dunlop, Anne Maria Eikeset, Katja Enberg, Georg H. Engelhard, Christian Jørgensen, Ane T. Laugen, Shuichi Matsumura, Sébastien Nusslé, Davnah Urbach, Rebecca Whitlock, Adriaan D. Rijnsdorp, and Ulf Dieckmann. Can fisheries-induced evolution shift reference points for fisheries management? ICES Journal of Marine Science, 70 (4):707–721, July 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/70/4/707/728577.

Hollowed:2013:PIC

[HBB^{+}13c] Anne B. Hollowed, Manuel Barange, Richard J. Beamish, Keith Brander, Kevern Cochrane, Kenneth Drinkwater, Michael G. G. Foreman, Jonathan A. Hare, Jason Holt, Shin ichi Ito, Suam Kim, Jacquelynne R. King, Harald Loeng, Brian R. MacKenzie, Franz J. Mueter, Thomas A. Okey, Myron A. Peck, Vladimir I. Radchenko, Jake C. Rice, Michael J. Schirripa, Akihiko Yatsu, and Yasuhiro Yamanaka. Projected impacts of climate change on marine fish and fisheries. ICES Journal of Marine Science, 70(5):1023– 1037, September 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/70/5/1023/645653.

Haig:2016:RTF

[HBB^{+}16] Jodie A. Haig, Snorre Bakke, Michael C. Bell, Isobel S. M. Bloor, Mike Cohen, Matthew Coleman, Samuel Dignan, Michel J. Kaiser, Julia R. Pantin, Michael Roach, Harriet Salomonsen, and Oliver Tully. Reproductive traits and factors affecting the size at maturity of Cancer pagurus across Northern Europe. ICES Journal of Marine Science, 73(10):2572–2585, November 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/10/2572/2647087.

Hansson:2018:CFF

[HBB^{+}18] Sture Hansson, Ulf Bergström, Erik Bonsdorff, Tero Härkönen, Niels Jepsen, Lena Kautsky, Karl Lundström, Sven-Gunnar Lunneryd, Maria Ovegård, Juhani Salmi, Dmitry Sendek, and Markus Vetemaa. Competition for the fish–fish extraction from the Baltic Sea by humans, aquatic mammals, and birds. ICES Journal of Marine Science, 75(3):999–1008, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/75/
3/999/4616536.

Holmes:2011:UFD

[HBC^{+}11] S. J. Holmes, N. Bailey, N. Campbell, R. Catarino, K. Barratt, A. Gibb, and P. G. Fernandes. Using fisherydependent data to inform the development and operation of a co-management initiative to reduce cod mortality and cut discards. ICES Journal of Marine Science, 68(8):1679– 1688, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/8/1679/757531.

Holt:2019:BSCa

[HBD^{+}19a] Rebecca E. Holt, Bjarte Bogstad, Joël M. Durant, Andrey V. Dolgov, and Geir Ottersen. Barents Sea cod (Gadus morhua) diet composition: long-term interannual, seasonal, and ontogenetic patterns. ICES Journal of Marine Science, 76(6):1641–1652, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/6/1641/5491593.

Holt:2019:BSCb

[HBD^{+}19b] Rebecca E. Holt, Bjarte Bogstad, Joël M. Durant, Andrey V. Dolgov, and Geir Ottersen. Barents Sea cod (Gadus morhua) diet composition: long-term interannual, seasonal, and ontogenetic patterns. ICES Journal of Marine Science, 76(6):1936, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
6/1936/5522544.

Howell:2014:QAI

[HBF14] Kerry L. Howell, Ross D. Bullimore, and Nicola L. Foster. Quality assurance in the identification of deep-sea taxa from video and image analysis: response to Henry and Roberts. ICES Journal of Marine Science, 71(4):899– 906, May 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/71/4/899/2843160.

Horn:2016:EHC

[HBG^{+}16] Henriette G. Horn, Maarten Boersma, Jessica Garzke, Martin G. J. Löder, Ulrich Sommer, and Nicole Aberle. Effects of high CO_{2} and warming on a Baltic Sea microzooplankton community. ICES Journal of Marine Science, 73(3):772– 782, February 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/73/3/772/2458875.

Hollowed:2019:RAUa

[HBG^{+}19a] Anne B. Hollowed, Manuel Barange, Véronique Garçon, Shin ichi Ito, Jason S. Link, Salvatore Aricò, Harold Batchelder, Robin Brown, Roger Griffis, and Wojciech Wawrzynski. Recent advances in understanding the effects of climate change on the world’s oceans. ICES Journal of Marine Science, 76(5):1215–1220, 09- 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/5/1215/5543601.

Hollowed:2019:RAUb

[HBG^{+}19b] Anne B. Hollowed, Manuel Barange, Véronique Garçon, Shin ichi Ito, Jason S. Link, Salvatore Aricò, Harold Batchelder, Robin Brown, Roger Griffis, and Wojciech Wawrzynski. Recent advances in understanding the effects of climate change on the world’s oceans. ICES Journal of Marine Science, 76(6):1940, November 2019.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/6/1940/5552796.

Howard:2018:MBP

[HBHR18] Sunkita Howard, Richard Brill, Chris Hepburn, and Jenny Rock. Microprocessor-based prototype bycatch reduction device reduces bait consumption by spiny dogfish and sandbar shark. ICES Journal of Marine Science, 75(6):2235– 2244, November 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/6/2235/5059521.

Hollowed:2011:ECC

[HBiI^{+}11] Anne Babcock Hollowed, Manuel Barange, Shin ichi Ito, Suam Kim, Harald Loeng, and Myron A. Peck. Effects of climate change on fish and fisheries: forecasting impacts, assessing ecosystem responses, and evaluating management strategies. ICES Journal of Marine Science, 68(6):984– 985, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/68/6/984/715097.

Heath:2014:CGS

[HCC^{+}14] Michael R. Heath, Mark A. Culling, Walter W. Crozier, Clive J. Fox, William S. C. Gurney, William F. Hutchinson, Einar E. Nielsen, Martha O’Sullivan, Katharine F. Preedy, David A. Righton, Douglas C. Speirs, Martin I. Taylor, Peter J. Wright, and Gary R. Carvalho. Combination of genetics and spatial modelling highlights the sensitivity of cod (Gadus morhua) population diversity in the North Sea to distributions of fishing. ICES Journal of Marine Science, 71(4):794–807, May 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/71/4/794/667717.

Hurst:2015:CCS

[HCDAF15] Thomas P. Hurst, Daniel W. Cooper, Janet T. Duffy-Anderson, and Edward V. Farley. Contrasting coastal and shelf nursery habitats of Pacific cod in the southeastern Bering Sea. ICES Journal of Marine Science, 72(2):515– 527, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/72/2/515/2801363.

Hunt:2011:CIE

[HCE^{+}11] George L. Hunt, Kenneth O. Coyle, Lisa B. Eisner, Edward V. Farley, Ron A. Heintz, Franz Mueter, Jeffrey M. Napp, James E. Overland, Patrick H. Ressler, Sigrid Salo, and Phyllis J. Stabeno. Climate impacts on eastern Bering Sea foodwebs: a synthesis of new data and an assessment of the oscillating control hypothesis. ICES Journal of Marine Science, 68(6):1230–1243, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/6/1230/703602.

Hussy:2012:AVB

[HCF^{+}12a] Karin Hüssy, Julie O. Coad, Edward D. Farrell, Lotte A. W. Clausen, and Maurice W. Clarke. Age verification of boarfish (Capros aper) in the Northeast Atlantic. ICES Journal of Marine Science, 69(1):34–40, January 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/1/34/670178.

Hussy:2012:SDS

[HCF^{+}12b] Karin Hüssy, Julie Olivia Coad, Edward D. Farrell, Lotte W. Clausen, and Maurice W. Clarke. Sexual dimorphism in size, age, maturation, and growth characteristics of boarfish (Capros aper) in the Northeast Atlantic. ICES Journal of Marine Science, 69(10):1729–1735, December 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/69/10/1729/623663.

Hollyman:2018:AGR

[HCL^{+}18] Philip R. Hollyman, Simon R. N. Chenery, Melanie J. Leng, Vladimir V. Laptikhovsky, Charlotte N. Colvin, and Christopher A. Richardson. Age and growth rate estimations of the commercially fished gastropod Buccinum undatum. ICES Journal of Marine Science, 75(6):2129–2144, November 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/6/2129/5061530.

Hamer:2015:TDM

[HCM^{+}15] Derek J. Hamer, Simon J. Childerhouse, John P. McKinlay, Mike C. Double, and Nick J. Gales. Two devices for mitigating odontocete bycatch and depredation at the hook in tropical pelagic longline fisheries. ICES Journal of Marine Science, 72(5):1691–1705, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/5/1691/775094.

Hammond:2013:URA

[HCR13] Carwyn F. Hammond, Loveday L. Conquest, and Craig S. Rose. Using reflex action mortality predictors (RAMP) to evaluate if trawl gear modifications reduce the unobserved mortality of Tanner crab (Chionoecetes bairdi) and snow crab (C. opilio). ICES Journal of Marine Science, 70(7):1308–1318, November 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/70/
7/1308/607342.

Hallett:2011:FES

[HD11] Chris S. Hallett and Ross K. Daley. Feeding ecology of the southern lanternshark (Etmopterus baxteri) and the brown lanternshark (E. unicolor) off southeastern Australia. ICES Journal of Marine Science, 68(1):157–165, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/68/1/157/629079.

Harden-Davies:2018:NWS

[HD18] Harriet Harden-Davies. The next wave of science diplomacy: marine biodiversity beyond national jurisdiction. ICES Journal of Marine Science, 75(1):426–434, January 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/1/426/4554490.

Hein:2016:INR

[HdBR^{+}16] Jennifer L. Hein, Isaure de Buron, William A. Roumillat, William C. Post, Allan P. Hazel, and Stephen A. Arnott. Infection of newly recruited American eels (Anguilla rostrata) by the invasive swimbladder parasite Anguillicoloides crassus in a US Atlantic tidal creek. ICES Journal of Marine Science, 73(1):14–21, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/1/14/2458698.

Hinrichsen:2011:ESC

[HDCH^{+}11] Hans-Harald Hinrichsen, Mark Dickey-Collas, Martin Huret, Myron A. Peck, and Frode B. Vikebø. Evaluating the suitability of coupled biophysical models for fishery management. ICES Journal of Marine Science, 68(7): 1478–1487, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/7/1478/654419.

Haug:2019:IEC

[HDF^{+}19] Tore Haug, Andrey Dolgov, Anatoly Filin, Maria Fossheim, Geir Huse, Evgeny Shamray, Jan Erik Stiansen, and Alexander Trofimov. Influence of ecosystem changes on harvestable resources at high latitudes. ICES Journal of Marine Science, 76(S1):S1–S2, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/76/
Supplement_1/i1/5701493.

Howarth:2017:TOM

[HDG^{+}17] Leigh M. Howarth, Pascal Dubois, Paul Gratton, Matthew Judge, Brian Christie, James J. Waggitt, Julie P. Hawkins, Callum M. Roberts, and Bryce D. Stewart. Trade-offs in marine protection: multispecies interactions within a community-led temperate marine reserve. ICES Journal of Marine Science, 74(1):263–276, January 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/1/263/2669577.

Howell:2014:MLI

[HF14] Daniel Howell and Anatoly A. Filin. Modelling the likely impacts of climate-driven changes in cod-capelin overlap in the Barents Sea. ICES Journal of Marine Science, 71(1): 72–80, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/71/1/72/646463.

Hall:2018:PSD

[HFGT18] Karina C. Hall, Anthony J. Fowler, Michael C. Geddes, and Julian D. Taylor. Predictable spatiotemporal dynamics of a dense cuttlefish spawning aggregation increases its vulnerability to exploitation. ICES Journal of Marine Science, 75(1):221–234, January 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/1/221/3867701.

Houde:2010:FRC

[HFH10] Aimee Lee S. Houde, Dylan J. Fraser, and Jeffrey A. Hutchings. Fitness-related consequences of competitive interactions between farmed and wild Atlantic salmon at different proportional representations of wild–farmed hybrids. ICES Journal of Marine Science, 67(4):657–667, May 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/67/4/657/679528.

Hurtado-Ferro:2010:AEE

[HFHS10] Felipe Hurtado-Ferro, Kazuhiko Hiramatsu, and Kunio Shirakihara. Allowing for environmental effects in a management strategy evaluation for Japanese sardine. ICES Journal of Marine Science, 67(9):2012–2017, December 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/9/2012/622949.

Hurst:2013:EOA

[HFM13] Thomas P. Hurst, Elena R. Fernandez, and Jeremy T. Mathis. Effects of ocean acidification on hatch size and larval growth of walleye pollock (Theragra chalcogramma). ICES Journal of Marine Science, 70(4):812–822, July 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/4/812/727203.

Holt:2013:DPS

[HFNH13] Rebecca E. Holt, Andrew Foggo, Francis C. Neat, and Kerry L. Howell. Distribution patterns and sexual segregation in chimaeras: implications for conservation and management. ICES Journal of Marine Science, 70(6):1198– 1205, September 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/70/6/1198/636894.

Hernandez-Farinas:2014:TCP

[HFSB^{+}14] Tania Hernández-Fariñas, Dominique Soudant, Laurent Barillé, Catherine Belin, Alain Lefebvre, and Cédric Bacher. Temporal changes in the phytoplankton community along the French coast of the eastern English Channel and the Southern Bight of the North Sea. ICES Journal of Marine Science, 71(4):821–833, May 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/4/821/668488.

Hurtado-Ferro:2015:LRV

[HFSV^{+}15] Felipe Hurtado-Ferro, Cody S. Szuwalski, Juan L. Valero, Sean C. Anderson, Curry J. Cunningham, Kelli F. Johnson, Roberto Licandeo, Carey R. McGilliard, Cole C. Monnahan, Melissa L. Muradian, Kotaro Ono, Katyana A. Vert-Pre, Athol R. Whitten, and André E. Punt. Looking in the rear-view mirror: bias and retrospective patterns in integrated, age-structured stock assessment models. ICES Journal of Marine Science, 72(1):99–110, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/1/99/2804338.

Hoff:2010:EEM

[HFU^{+}10] Ayoe Hoff, Hans Frost, Clara Ulrich, Dimitrios Damalas, Christos D. Maravelias, Leyre Goti, and Marina Santurtún. Economic effort management in multispecies fisheries: the FcubEcon model. ICES Journal of Marine Science, 67(8):1802–1810, November 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/67/
8/1802/605088.

Hartley:2010:SMP

[HG10] Troy W. Hartley and Christopher Glass. Science-to-management pathways in US Atlantic herring management: using governance network structure and function to track information flow and potential influence. ICES Journal of Marine Science, 67(6):1154–1163, September 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/6/1154/734964.

Hermann:2019:PBCa

[HGC^{+}19a] Albert J. Hermann, Georgina A. Gibson, Wei Cheng, Ivonne Ortiz, Kerim Aydin, Muyin Wang, Anne B. Hollowed, and Kirstin K. Holsman. Projected biophysical conditions of the Bering Sea to 2100 under multiple emission scenarios. ICES Journal of Marine Science, 76(5):1280– 1304, 09- 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/5/1280/5477847.

Hermann:2019:PBCb

[HGC^{+}19b] Albert J. Hermann, Georgina A. Gibson, Wei Cheng, Ivonne Ortiz, Kerim Aydin, Muyin Wang, Anne B. Hollowed, and Kirstin K. Holsman. Projected biophysical conditions of the Bering Sea to 2100 under multiple emission scenarios. ICES Journal of Marine Science, 76(6):1937, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/6/1937/5513066.

Healey:2018:GAR

[HGF^{+}18] Amy J. E. Healey, Gavin Gouws, Sean T. Fennessy, Baraka Kuguru, Warwick H. H. Sauer, Paul W. Shaw, and Niall J. McKeown. Genetic analysis reveals harvested Lethrinus nebulosus in the Southwest Indian Ocean comprise two cryptic species. ICES Journal of Marine Science, 75(4): 1465–1472, July 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/4/1465/4791425.

Hinojosa:2015:SES

[HGGJ15] Iván A. Hinojosa, Bridget S. Green, Caleb Gardner, and Andrew Jeffs. Settlement and early survival of southern rock lobster, Jasus edwardsii, under climate-driven decline of kelp habitats. ICES Journal of Marine Science, 72(S1): S59–S68, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/72/suppl_1/i59/614738.

Hussy:2016:SRS

[HGH^{+}16] K. Hüssy, J. Gröger, F. Heidemann, H.-H. Hinrichsen, and L. Marohn. Slave to the rhythm: seasonal signals in otolith microchemistry reveal age of eastern Baltic cod (Gadus morhua). ICES Journal of Marine Science, 73(4):1019– 1032, March 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/73/4/1019/2458723.

Hedeholm:2010:VSG

[HGRAR10] R. Hedeholm, P. Grønkjær, A. Rosing-Asvid, and S. Rysgaard. Variation in size and growth of West Greenland capelin (Mallotus villosus) along latitudinal gradients. ICES Journal of Marine Science, 67(6):1128–1137, September 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/67/6/1128/736600.

Hoare:2011:ISD

[HGS11] Deirdre Hoare, Norman Graham, and Pieter-Jan Schön. The Irish Sea data-enhancement project: comparison of self-sampling and national data-collection programmes — results and experiences. ICES Journal of Marine Science, 68(8):1778–1784, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/68/
8/1778/757089.

Henderson:2019:CEM

[HGS^{+}19] Christopher J. Henderson, Ben L. Gilby, Thomas A. Schlacher, Rod M. Connolly, Marcus Sheaves, Nicole Flint, Hayden P. Borland, and Andrew D. Olds. Contrasting effects of mangroves and armoured shorelines on fish assemblages in tropical estuarine seascapes. ICES Journal of Marine Science, 76(4):1052–1061, July 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/4/1052/5306605.

Heymans:2011:DWE

[HHA^{+}11] Johanna J. Heymans, Kerry L. Howell, Morag Ayers, Michael T. Burrows, John D. M. Gordon, Emma G. Jones, and Francis Neat. Do we have enough information to apply the ecosystem approach to management of deepsea fisheries? An example from the West of Scotland. ICES Journal of Marine Science, 68(2):265–280, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/2/265/614494.

Howell:2016:BHV

[HHBSM16] Daniel Howell, Cecilie Hansen, Bjarte Bogstad, and Mette Skern-Mauritzen. Balanced harvesting in a variable and uncertain world: a case study from the Barents Sea. ICES Journal of Marine Science, 73(6):1623–1631, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/6/1623/2459101.

Harris:2014:ESS

[HHDB14] R. N. Harris, C. M. Harris, C. D. Duck, and I. L. Boyd. The effectiveness of a seal scarer at a wild salmon net fishery. ICES Journal of Marine Science, 71(7):1913–1920, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/71/7/1913/663289.

Hussy:2016:STT

[HHE^{+}16] K. Hüssy, H.-H. Hinrichsen, M. Eero, H. Mosegaard, J. Hemmer-Hansen, A. Lehmann, and L. S. Lundgaard. Spatio-temporal trends in stock mixing of eastern and western Baltic cod in the Arkona Basin and the implications for recruitment. ICES Journal of Marine Science, 73(2):293– 303, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/73/2/293/2614469.

Hughes:2012:VIN

[HHGtIWGoOH12] Sarah L. Hughes, N. Penny Holliday, Fabienne Gaillard, and the Ices Working Group on Oceanic Hydrography. Variability in the ICES/NAFO region between 1950 and 2009: observations from the ICES Report on Ocean Climate. ICES Journal of Marine Science, 69(5):706–719, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/69/5/706/649782.

Hinrichsen:2012:STV

[HHH12a] H-H. Hinrichsen, K. Hüssy, and B. Huwer. Spatio-temporal variability in western Baltic cod early life stage survival mediated by egg buoyancy, hydrography and hydrodynamics. ICES Journal of Marine Science, 69(10):1744–1752, December 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/69/10/1744/622270.

Hussy:2012:HIS

[HHH12b] K. Hüssy, H.-H. Hinrichsen, and B. Huwer. Hydrographic influence on the spawning habitat suitability of western Baltic cod (Gadus morhua). ICES Journal of Marine Science, 69(10):1736–1743, December 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/10/1736/622218.

Holsman:2019:TCR

[HHH^{+}19] Kirstin K. Holsman, Elliott Lee Hazen, Alan Haynie, Sophie Gourguet, Anne Hollowed, Steven J. Bograd, Jameal F. Samhouri, and Kerim Aydin. Towards climate resiliency in fisheries management. ICES Journal of Marine Science, 76(5):1368–1378, 09- 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
5/1368/5420302.

Huwer:2016:CLC

[HHHE16] B. Huwer, H.-H. Hinrichsen, K. Hüssy, and M. Eero. Connectivity of larval cod in the transition area between North Sea and Baltic Sea and potential implications for fisheries management. ICES Journal of Marine Science, 73(7):1815– 1824, July 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/7/1815/2458738.

Hilborn:2014:DFP

[HHJB14] Ray Hilborn, Daniel J. Hively, Olaf P. Jensen, and Trevor A. Branch. The dynamics of fish populations at low abundance and prospects for rebuilding and recovery. ICES Journal of Marine Science, 71(8):2141–2151, October 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/8/2141/754978.

Hinrichsen:2011:CDL

[HHM^{+}11] Hans-Harald Hinrichsen, Bastian Huwer, Andrejs Makarchouk, Christoph Petereit, Matthias Schaber, and Rudi Voss. Climate-driven long-term trends in Baltic Sea oxygen concentrations and the potential consequences for eastern Baltic cod (Gadus morhua). ICES Journal of Marine Science, 68(10):2019–2028, November 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/10/2019/611652.

Hess:2016:MAS

[HHM^{+}16] Maureen A. Hess, Jon E. Hess, Andrew P. Matala, Rod A. French, Craig A. Steele, Jens C. Lovtang, and Shawn R. Narum. Migrating adult steelhead utilize a thermal refuge during summer periods with high water temperatures. ICES Journal of Marine Science, 73(10):2616–2624, November 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/10/2616/2647115.

Hobday:2019:ECU

[HHM^{+}19] Alistair J. Hobday, Jason R. Hartog, John P. Manderson, Katherine E. Mills, Matthew J. Oliver, Andrew J. Pershing, and Samantha Siedlecki. Ethical considerations and unanticipated consequences associated with ecological forecasting for marine resources. ICES Journal of Marine Science, 76(5):1244–1256, 09- 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/5/1244/5303214.

Haas:2019:IPR

[HHMF19] Bianca Haas, Marcus Haward, Jeffrey McGee, and Aysha Fleming. The influence of performance reviews on regional fisheries management organizations. ICES Journal of Marine Science, 76(7):2082–2089, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/7/2082/5497987.

Hulson:2011:EPO

[HHQ11] Peter-John F. Hulson, Dana H. Hanselman, and Terrance J. Quinn II. Effects of process and observation errors on effective sample size of fishery and survey age and length composition using variance ratio and likelihood methods. ICES Journal of Marine Science, 68(7):1548–1557, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/7/1548/661038.

Hulson:2012:DES

[HHQ12] Peter-John F. Hulson, Dana H. Hanselman, and Terrance J. Quinn II. Determining effective sample size in integrated age-structured assessment models. ICES Journal of Marine Science, 69(2):281–292, March 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/2/281/697956.

Handegard:2017:MOL

[HHR17] Nils Olav Handegard, Arne Johannes Holmin, and Guillaume Rieucau. Method to observe large scale behavioural waves propagating through fish schools using 4D sonar. ICES Journal of Marine Science, 74(3):804–812, March 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/3/804/2736331.

Hansen:2012:SSS

[HHRW12] Lars P. Hansen, Peter Hutchinson, David G. Reddin, and Malcolm L. Windsor. Salmon at sea: Scientific advances and their implications for management: an introduction. ICES Journal of Marine Science, 69(9):1533–1537, November 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/69/9/1533/640722.

Hufnagl:2013:HDS

[HHT13] Marc Hufnagl, Klaus B. Huebert, and Axel Temming. How does seasonal variability in growth, recruitment, and mortality affect the performance of length-based mortality and asymptotic length estimates in aquatic resources? ICES Journal of Marine Science, 70(2):329–341, March 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/2/329/793786.

Hornborg:2018:SSS

[HHZ^{+}18] Sara Hornborg, Alistair J. Hobday, Friederike Ziegler, Anthony D. M. Smith, and Bridget S. Green. Shaping sustainability of seafood from capture fisheries integrating the perspectives of supply chain stakeholders through combining systems analysis tools. ICES Journal of Marine Science, 75(6):1965–1974, November 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/6/1965/5056868.

Hilborn:2016:MFF

[Hil16] Ray Hilborn. Of mice, fishermen, and food. ICES Journal of Marine Science, 73(9):2167–2173, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/9/2167/2199405.

Hilborn:2018:ME

[Hil18a] Ray Hilborn. Are MPAs effective? ICES Journal of Marine Science, 75(3):1160–1162, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/75/
3/1160/4098822.

Hilborn:2018:CSG

[Hil18b] Ray Hilborn. Counterpoint to sala and giakoumi. ICES Journal of Marine Science, 75(3):1169–1170, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/3/1169/4098838.

Hilborn:2018:HFW

[Hil18c] Ray Hilborn. Hilborn’s final word. ICES Journal of Marine Science, 75(3):1165, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/75/
3/1165/4098836.

Hilborn:2019:MFPa

[Hil19a] Ray Hilborn. Measuring fisheries performance using the “Goldilocks plot”. ICES Journal of Marine Science, 76(1): 45–49, January 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/76/1/45/5133277.

Hilborn:2019:MFPb

[Hil19b] Ray Hilborn. Measuring fisheries performance using the “Goldilocks plot”. ICES Journal of Marine Science, 76(1): 356, January 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/76/1/356/5250413.

Hind:2015:RPP

[Hin15] Edward J. Hind. A review of the past, the present, and the future of fishers’ knowledge research: a challenge to established fisheries science. ICES Journal of Marine Science, 72(2):341–358, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/72/
2/341/2801477.

Hjelset:2014:FIC

[Hje14] Ann Merete Hjelset. Fishery-induced changes in Norwegian red king crab (Paralithodes camtschaticus) reproductive potential. ICES Journal of Marine Science, 71(2):365–373, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/71/2/365/783160.

Hixon:2014:BIC

[HJS14] Mark A. Hixon, Darren W. Johnson, and Susan M. Sogard. BOFFFFs: on the importance of conserving oldgrowth age structure in fishery populations. ICES Journal of Marine Science, 71(8):2171–2185, October 2014.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/8/2171/748104.

Hornborg:2017:NPM

[HJS^{+}17] Sara Hornborg, Patrik Jonsson, Mattias Sköld, Mats Ulmestrand, Daniel Valentinsson, Ole Ritzau Eigaard, Jordan Feekings, J. Rasmus Nielsen, Francois Bastardie, and Johan Lövgren. New policies may call for new approaches: the case of the Swedish Norway lobster (Nephrops norvegicus) fisheries in the Kattegat and Skagerrak. ICES Journal of Marine Science, 74(1):134–145, January 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/1/134/2669569.

Hennen:2012:APM

[HJT12] Daniel R. Hennen, Larry D. Jacobson, and Jiashen Tang. Accuracy of the patch model used to estimate density and capture efficiency in depletion experiments for sessile invertebrates and fish. ICES Journal of Marine Science, 69(2): 240–249, March 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/69/2/240/702914.

Hutchings:2014:GFI

[HK14] Jeffrey A. Hutchings and Anna Kuparinen. Ghosts of fisheries-induced depletions: do they haunt us still? ICES Journal of Marine Science, 71(6):1467–1473, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/6/1467/625046.

Haasnoot:2016:FGT

[HKB16] Tim Haasnoot, Marloes Kraan, and Simon R. Bush. Fishing gear transitions: lessons from the Dutch flatfish pulse trawl. ICES Journal of Marine Science, 73(4):1235–1243, March 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/4/1235/2458968.

Hansson:2018:RCH

[HKB^{+}18] Sture Hansson, Lena Kautsky, Ulf Bergström, Erik Bonsdorff, Niels Jepsen, Karl Lundström, Sven-Gunnar Lunneryd, Maria Ovegård, Juhani Salmi, Dmitry Sendek, and Markus Vetemaa. Response to comments by Heikinheimo et al. (in press) on Hansson et al. (2018): competition for the fish — fish extraction from the Baltic Sea by humans, aquatic mammals, and birds. ICES Journal of Marine Science, 75(5):1837–1839, September 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/5/1837/5051295.

Hidalgo:2017:ALB

[HKK^{+}17] Manuel Hidalgo, David M. Kaplan, Lisa A. Kerr, James R. Watson, Claire B. Paris, and Howard I. Browman. Advancing the link between ocean connectivity, ecological function and management challenges. ICES Journal of Marine Science, 74(6):1702–1707, July 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
6/1702/4004697.

Hammer:2010:RDF

[HKKS10] Cornelius Hammer, Olav Sigurd Kjesbu, Gordon H. Kruse, and Peter A. Shelton. Rebuilding depleted fish stocks: biology, ecology, social science, and management strategies. ICES Journal of Marine Science, 67(9):1825–1829, December 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/67/9/1825/619160.

Haris:2018:DWC

[HKRM18] Kunnath Haris, Rudy J. Kloser, Tim E. Ryan, and Jacques Malan. Deep-water calibration of echosounders used for biomass surveys and species identification. ICES Journal of Marine Science, 75(3):1117–1130, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/3/1117/4683721.

Hvingel:2012:SEK

[HKS12] C. Hvingel, M. C. S. Kingsley, and J. H. Sundet. Survey estimates of king crab (Paralithodes camtschaticus) abundance off Northern Norway using GLMs within a mixed generalized gamma-binomial model and Bayesian inference.

ICES Journal of Marine Science, 69(8):1416–1426, September 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/69/8/1416/702937.

Harvey:2017:IIU

[HKS17] Chris J. Harvey, Christopher R. Kelble, and Franklin B. Schwing. Implementing “the IEA”: using integrated ecosystem assessment frameworks, programs, and applications in support of operationalizing ecosystem-based management. ICES Journal of Marine Science, 74(1):398–405, January 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/1/398/2674180.

Hoolihan:2011:EPR

[HLA^{+}11] J. P. Hoolihan, J. Luo, F. J. Abascal, S. E. Campana, G. De Metrio, H. Dewar, M. L. Domeier, L. A. Howey, M. E. Lutcavage, M. K. Musyl, J. D. Neilson, E. S. Orbesen, E. D. Prince, and J. R. Rooker. Evaluating post-release behaviour modification in large pelagic fish deployed with pop-up satellite archival tags. ICES Journal of Marine Science, 68(5):880–889, May 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/68/5/880/649692.

Henriquez:2016:IBA

[HLCC16] Vania Henríquez, Roberto Licandeo, Luis A. Cubillos, and Sean P. Cox. Interactions between ageing error and selectivity in statistical catch-at-age models: simulations and implications for assessment of the Chilean Patagonian toothfish fishery. ICES Journal of Marine Science, 73(4):1074– 1090, March 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/73/4/1074/2458932.

Hoving:2010:MRS

[HLD10] Hendrik Jan T. Hoving, Marek R. Lipiński, and Lammertjan Dam. The male reproductive strategy of a deep-sea squid: sperm allocation, continuous production, and longterm storage of spermatophores in Histioteuthis miranda. ICES Journal of Marine Science, 67(7):1478–1486, October 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/7/1478/662477.

Hartman:2012:PAP

[HLL^{+}12] Susan E. Hartman, Richard S. Lampitt, Kate E. Larkin, Maureen Pagnani, Jon Campbell, Thanos Gkritzalis, Zong-Pei Jiang, Corinne A. Pebody, Henry A. Ruhl, Andrew J. Gooday, Brian J. Bett, David S. M. Billett, Paul Provost, Rob McLachlan, Jon D. Turton, and Steven Lankester. The porcupine abyssal plain fixed-point sustained observatory (PAP-SO): variations and trends from the Northeast Atlantic fixed-point time-series. ICES Journal of Marine Science, 69(5):776–783, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/69/5/776/653510.

Heikinheimo:2018:CHA

[HLL18] Outi Heikinheimo, Hannu Lehtonen, and Aleksi Lehikoinen. Comment to Hansson, S. et al. (2017): “Competition for the fish–fish extraction from the Baltic Sea by humans, aquatic mammals, and birds”, with special reference to cormorants, perch, and pikeperch. ICES Journal of Marine Science, 75(5):1832–1836, September 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/75/
5/1832/4992260.

Hurst:2016:EEC

[HLMT16] Thomas P. Hurst, Benjamin J. Laurel, Jeremy T. Mathis, and Lauren R. Tobosa. Effects of elevated CO_{2} levels on eggs and larvae of a North Pacific flatfish. ICES Journal of Marine Science, 73(3):981–990, February 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/3/981/2457911.

Hoolihan:2015:VHH

[HLS^{+}15] John P. Hoolihan, Jiangang Luo, Derke Snodgrass, Eric S. Orbesen, Ann M. Barse, and Eric D. Prince. Vertical and horizontal habitat use by white marlin Kajikia albida (Poey, 1860) in the western North Atlantic Ocean. ICES Journal of Marine Science, 72(8):2364–2373, September 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/8/2364/2458701.

Holmes:2014:GDD

[HMFW14] Steven J. Holmes, Colin P. Millar, Robert J. Fryer, and Peter J. Wright. Gadoid dynamics: differing perceptions when contrasting stock vs. population trends and its implications to management. ICES Journal of Marine Science, 71(6):1433–1442, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/6/1433/2835580.

Hart:2019:BEH

[HMHL19] Anthony M. Hart, Dave Murphy, S. Alex Hesp, and Stephen Leporati. Biomass estimates and harvest strategies for the Western Australian Octopus aff. tetricus fishery. ICES Journal of Marine Science, 76(7):2205–2217, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/7/2205/5552830.

Harris:2018:AMC

[HMLKR18] Peter T. Harris, Miles Macmillan-Lawler, Lars Kullerud, and Jake C. Rice. Arctic marine conservation is not prepared for the coming melt. ICES Journal of Marine Science, 75(1):61–71, January 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/75/1/61/4080407.

Hurst:2012:DPG

[HMM12] Thomas P. Hurst, Jamal H. Moss, and Jessica A. Miller. Distributional patterns of 0-group Pacific cod (Gadus macrocephalus) in the eastern Bering Sea under variable recruitment and thermal conditions. ICES Journal of Marine Science, 69(2):163–174, March 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/69/
2/163/707756.

Hare:2012:CBB

[HMN^{+}12] Jonathan A. Hare, John P. Manderson, Janet A. Nye, Michael A. Alexander, Peter J. Auster, Diane L. Borggaard, Antonietta M. Capotondi, Kimberly B. Damon-Randall, Eric Heupel, Ivan Mateo, Loretta O’Brien, David E. Richardson, Charles A. Stock, and Sarah T. Biegel. Cusk (Brosme brosme) and climate change: assessing the threat to a candidate marine fish species under the US Endangered Species Act. ICES Journal of Marine Science, 69(10):1753– 1768, December 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/69/10/1753/624341.

Hold:2015:VCC

[HMP^{+}15] Natalie Hold, Lee G. Murray, Julia R. Pantin, Jodie A. Haig, Hilmar Hinz, and Michel J. Kaiser. Video capture of crustacean fisheries data as an alternative to on-board observers. ICES Journal of Marine Science, 72(6):1811– 1821, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/6/1811/919590.

Hyun:2015:IMH

[HMR15] Saang-Yoon Hyun, Mark N. Maunder, and Brian J. Rothschild. Importance of modelling heteroscedasticity of survey index data in fishery stock assessments. ICES Journal of Marine Science, 72(1):130–136, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/1/130/823326.

Herrmann:2013:MTH

[HMS^{+}13] Bent Herrmann, Bernd Mieske, Daniel Stepputtis, Ludvig Ahm Krag, Niels Madsen, and Thomas Noack. Modelling towing and haul-back escape patterns during the fishing process: a case study for cod, plaice, and flounder in the demersal Baltic Sea cod fishery. ICES Journal of Marine Science, 70(4):850–863, July 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/70/4/850/725120.

Holmgren:2012:MOM

[HNAK12] Noel M. A. Holmgren, Niclas Norrström, Robert Aps, and Sakari Kuikka. MSY-orientated management of Baltic Sea herring (Clupea harengus) during different ecosystem regimes. ICES Journal of Marine Science, 69(2):257–266, March 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/69/2/257/701350.

Hilborn:2014:RST

[HO14] Ray Hilborn and Daniel Ovando. Reflections on the success of traditional fisheries management. ICES Journal of Marine Science, 71(5):1040–1046, July 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/71/
5/1040/648075.

Holt:2010:WDP

[Hol10] Carrie A. Holt. Will depleted populations of Pacific salmon recover under persistent reductions in survival and catastrophic mortality events? ICES Journal of Marine Science, 67(9):2018–2026, December 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/67/
9/2018/620513.

Holt:2014:GSJ

[Hol14] Sidney J. Holt. The graceful sigmoid: Johan Hjort’s contribution to the theory of rational fishing. ICES Journal of Marine Science, 71(8):2008–2011, October 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/8/2008/2804457.

Hordyk:2015:SEL

[HOS^{+}15] Adrian Hordyk, Kotaro Ono, Keith Sainsbury, Neil Loneragan, and Jeremy Prince. Some explorations of the life history ratios to describe length composition, spawningper-recruit, and the spawning potential ratio. ICES Journal of Marine Science, 72(1):204–216, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/1/204/815363.

Hordyk:2015:NLB

[HOV^{+}15] Adrian Hordyk, Kotaro Ono, Sarah Valencia, Neil Loneragan, and Jeremy Prince. A novel length-based empirical estimation method of spawning potential ratio (SPR), and tests of its performance, for small-scale, data-poor fisheries. ICES Journal of Marine Science, 72(1):217–231, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/1/217/815772.

Hufnagl:2011:PIB

[HP11] Marc Hufnagl and Myron A. Peck. Physiological individualbased modelling of larval Atlantic herring (Clupea harengus) foraging and growth: insights on climate-driven lifehistory scheduling. ICES Journal of Marine Science, 68(6): 1170–1188, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/6/1170/714089.

Haynie:2012:WEM

[HP12] Alan C. Haynie and Lisa Pfeiffer. Why economics matters for understanding the effects of climate change on fisheries. ICES Journal of Marine Science, 69(7):1160–1167, September 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/69/7/1160/740809.

Hawkins:2017:SAA

[HP17] Anthony D. Hawkins and Arthur N. Popper. A sound approach to assessing the impact of underwater noise on marine fishes and invertebrates. ICES Journal of Marine Science, 74(3):635–651, March 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
3/635/2739034.

Hordyk:2019:CNA

[HPCW19] Adrian R. Hordyk, Jeremy D. Prince, Thomas R. Carruthers, and Carl J. Walters. Comment on “A new approach for estimating stock status from length frequency data” by Froese et al. (2018). ICES Journal of Marine Science, 76(2):457–460, March 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
2/457/5298547.

Hardison:2019:SST

[HPDB19] Sean Hardison, Charles T. Perretti, Geret S. DePiper, and Andrew Beet. A simulation study of trend detection methods for integrated ecosystem assessment. ICES Journal of Marine Science, 76(7):2060–2069, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/7/2060/5512306.

Hinrichsen:2017:SDV

[HPN^{+}17] Hans-Harald Hinrichsen, Christoph Petereit, Anders Nissling, Isa Wallin, Didzis Ustups, and Ann-Britt Florin. Survival and dispersal variability of pelagic eggs and yolksac larvae of central and eastern Baltic flounder (Platichthys flesus): application of biophysical models. ICES Journal of Marine Science, 74(1):41–55, January 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
1/41/2444572.

Heino:2011:CPT

[HPS^{+}11] M. Heino, F. M. Porteiro, T. T. Sutton, T. Falkenhaug, O. R. Godø, and U. Piatkowski. Catchability of pelagic trawls for sampling deep-living nekton in the mid-North Atlantic. ICES Journal of Marine Science, 68(2):377–389, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/68/2/377/616796.

Haig:2015:TSV

[HPS^{+}15] Jodie A. Haig, Julia R. Pantin, Harriet Salomonsen, Lee G. Murray, and Michel J. Kaiser. Temporal and spatial variation in size at maturity of the common whelk (Buccinum undatum). ICES Journal of Marine Science, 72(9):2707– 2719, November 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/72/9/2707/2457857.

Henry:2014:RBP

[HR14] Lea-Anne Henry and J. Murray Roberts. Recommendations for best practice in deep-sea habitat classification: Bullimore et al. as a case study. ICES Journal of Marine Science, 71(4):895–898, May 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/71/4/895/666861.

He:2015:RHF

[HRB15a] Pingguo He, Christopher Rillahan, and Vincent Balzano. Reduced herding of flounders by floating bridles: application in Gulf of Maine northern shrimp trawls to reduce bycatch. ICES Journal of Marine Science, 72(5):1514–1524, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/5/1514/765689.

Hintzen:2015:MCP

[HRB^{+}15b] N. T. Hintzen, B. Roel, D. Benden, M. Clarke, A. Egan, R. D. M. Nash, N. Rohlf, and E. M. C. Hatfield. Managing a complex population structure: exploring the importance of information from fisheries-independent sources. ICES Journal of Marine Science, 72(2):528–542, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/2/528/672291.

Hernandez:2013:AMA

[HRC^{+}13] Keith M. Hernandez, Denise Risch, Danielle M. Cholewiak, Micah J. Dean, Leila T. Hatch, William S. Hoffman, Aaron N. Rice, Douglas Zemeckis, and Sofie M. Van Parijs. Acoustic monitoring of Atlantic cod (Gadus morhua) in Massachusetts Bay: implications for management and conservation. ICES Journal of Marine Science, 70(3):628–635, April 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/70/3/628/915557.

Hermosilla:2010:AVC

[HRF^{+}10] Consuelo A. Hermosilla, Francisco Rocha, Graziano Fiorito, Ángel F. González, and Ángel Guerra. Age validation in common octopus Octopus vulgaris using stylet increment analysis. ICES Journal of Marine Science, 67(7):1458– 1463, October 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/67/7/1458/663287.

Hussy:2016:CIA

[HRP^{+}16] Karin Hüssy, Krzysztof Radtke, Maris Plikshs, Rainer Oeberst, Tatjana Baranova, Uwe Krumme, Rajlie Sjöberg, Yvonne Walther, and Henrik Mosegaard. Challenging ICES age estimation protocols: lessons learned from the eastern Baltic cod stock. ICES Journal of Marine Science, 73(9):2138–2149, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
9/2138/2199530.

Hammill:2010:CTP

[HS10] Michael O. Hammill and Garry B. Stenson. Comment on “Towards a precautionary approach to managing Canada’s commercial harp seal hunt” by Leaper et al. ICES Journal of Marine Science, 67(2):321–322, March 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/2/321/696491.

Hammill:2017:GCH

[HS17] M. O. Hammill and C. Sauvé. Growth and condition in harp seals: evidence of density-dependent and density-independent influences. ICES Journal of Marine Science, 74(5):1395–1407, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
5/1395/2806280.

Hidalgo:2016:OMS

[HSB16] Manuel Hidalgo, David H. Secor, and Howard I. Browman. Observing and managing seascapes: linking synoptic oceanography, ecological processes, and geospatial modelling. ICES Journal of Marine Science, 73(7):1825–1830, July 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/7/1825/2458794.

Hart:2018:RDE

[HSB18] Anthony M. Hart, Lachlan W. S. Strain, and Jamin Brown. Regulation dynamics of exploited and protected populations of Haliotis roei, and their response to a marine heatwave. ICES Journal of Marine Science, 75(6):1924–1939, November 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/6/1924/5056871.

Halvorsen:2016:MBS

[HSD^{+}16] Kim Tallaksen Halvorsen, Tonje Knutsen Sørdalen, Caroline Durif, Halvor Knutsen, Esben Moland Olsen, Anne Berit Skiftesvik, Torborg Emmerhoff Rustand, Reidun Marie Bjelland, and Leif Asbjørn Vøllestad. Male-biased sexual size dimorphism in the nest building corkwing wrasse (Symphodus melops): implications for a size regulated fishery. ICES Journal of Marine Science, 73(10):2586–2594, November 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
10/2586/2647121.

Hirawake:2012:SRS

[HSFiS12] Toru Hirawake, Katsuhito Shinmyo, Amane Fujiwara, and Sei ichi Saitoh. Satellite remote sensing of primary productivity in the Bering and Chukchi Seas using an absorption-based approach. ICES Journal of Marine Science, 69(7):1194–1204, September 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/69/
7/1194/751103.

Henderson:2018:SCL

[HSGL18] Christopher J. Henderson, Tim Stevens, Ben L. Gilby, and Shing Y. Lee. Spatial conservation of large mobile elasmobranchs requires an understanding of spatiotemporal seascape utilization. ICES Journal of Marine Science, 75(2):553–561, March 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/75/
2/553/4554491.

Hentati-Sundberg:2014:DFM

[HSHÖ14] J. Hentati-Sundberg, J. Hjelm, and H. Österblom. Does fisheries management incentivize non-compliance? Estimated misreporting in the Swedish Baltic Sea pelagic fishery based on commercial fishing effort. ICES Journal of Marine Science, 71(7):1846–1853, September 2014.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/7/1846/672045.

Hoffle:2014:VNA

[HSK^{+}14] H. Höffle, P. Solemdal, K. Korsbrekke, M. Johannessen, K. Bakkeplass, and O. S. Kjesbu. Variability of northeast Arctic cod (Gadus morhua) distribution on the main spawning grounds in relation to biophysical factors. ICES Journal of Marine Science, 71(6):1317–1331, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/6/1317/2835599.

Heino:2012:SDG

[HSNO12] Mikko Heino, Terje Svåsand, Jarle Tryti Nordeide, and Håkon Otterå. Seasonal dynamics of growth and mortality suggest contrasting population structure and ecology for cod, pollack, and saithe in a Norwegian fjord. ICES Journal of Marine Science, 69(4):537–546, May 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/4/537/635815.

Hildebrandt:2016:OAD

[HSS^{+}16] Nicole Hildebrandt, Franz J. Sartoris, Kai G. Schulz, Ulf Riebesell, and Barbara Niehoff. Ocean acidification does not alter grazing in the calanoid copepods Calanus finmarchicus and Calanus glacialis. ICES Journal of Marine Science, 73(3):927–936, February 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/3/927/2458986.

Hammill:2014:FGS

[HSSB14] M. O. Hammill, G. B. Stenson, D. P. Swain, and H. P. Benoît. Feeding by grey seals on endangered stocks of Atlantic cod and white hake. ICES Journal of Marine Science, 71(6):1332–1341, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/6/1332/2835596.

Halvorsen:2017:SSS

[HSV^{+}17] Kim Tallaksen Halvorsen, Tonje Knutsen Sørdalen, Leif Asbjørn Vøllestad, Anne Berit Skiftesvik, Sigurd Heiberg Espeland, and Esben Moland Olsen. Sexand size-selective harvesting of corkwing wrasse (Symphodus melops) — a cleaner fish used in salmonid aquaculture. ICES Journal of Marine Science, 74(3):660–669, March 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/3/660/2705983.

Hoenig:2016:LCL

[HTB^{+}16] John M. Hoenig, Amy Y.-H. Then, Elizabeth A. Babcock, Norman G. Hall, David A. Hewitt, and Sybrand A. Hesp. The logic of comparative life history studies for estimating key parameters, with a focus on natural mortality rate. ICES Journal of Marine Science, 73(10):2453–2467, November 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/10/2453/2647091.

Harte:2019:CCI

[HTKB19] Michael Harte, Rachel Tiller, George Kailis, and Merrick Burden. Countering a climate of instability: the future of relative stability under the common fisheries policy. ICES Journal of Marine Science, 76(7):1951–1958, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/7/1951/5522962.

Hart:2011:EIS

[HTM11] Anthony M. Hart, Adrian W. Thomson, and David Murphy. Environmental influences on stock abundance and fishing power in the silver-lipped pearl oyster fishery. ICES Journal of Marine Science, 68(3):444–453, March 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/3/444/655528.

Hassenruck:2017:MIR

[HTRF17] Christiane Hassenrück, Halina E. Tegetmeyer, Alban Ramette, and Katharina E. Fabricius. Minor impacts of reduced pH on bacterial biofilms on settlement tiles along natural pH gradients at two CO_{2} seeps in Papua New Guinea. ICES Journal of Marine Science, 74(4):978–987, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/4/978/2900575.

Hufnagl:2010:ETM

[HTS^{+}10] Marc Hufnagl, Axel Temming, Volker Siegel, Ingrid Tulp, and Loes Bolle. Estimating total mortality and asymptotic length of Crangon crangon between 1955 and 2006. ICES Journal of Marine Science, 67(5):875–884, July 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/5/875/609782.

Hertz:2017:OSF

[HTT^{+}17] Eric Hertz, Marc Trudel, Strahan Tucker, Terry D. Beacham, and Asit Mazumder. Overwinter shifts in the feeding ecology of juvenile Chinook salmon. ICES Journal of Marine Science, 74(1):226–233, January 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/1/226/2669557.

Hubbard:2014:JHC

[Hub14] Jennifer Hubbard. Johan Hjort: The Canadian Fisheries Expedition, International Scientific Networks, and the challenge of modernization. ICES Journal of Marine Science, 71(8):2000–2007, October 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/71/
8/2000/2804386.

Humphreys:2017:CSR

[Hum17] Matthew P. Humphreys. Climate sensitivity and the rate of ocean acidification: future impacts, and implications for experimental design. ICES Journal of Marine Science, 74 (4):934–940, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/74/4/934/2667504.

Humborstad:2018:ALB

[HUPBL18] Odd-Børre Humborstad, Anne Christine Utne-Palm, Michael Breen, and Svein Løkkeborg. Artificial light in baited pots substantially increases the catch of cod (Gadus morhua) by attracting active bait, krill (Thysanoessa inermis). ICES Journal of Marine Science, 75(6):2257–2264, November 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/6/2257/5066377.

Hussy:2010:WAD

[Hüs10] Karin Hüssy. Why is age determination of Baltic cod (Gadus morhua) so difficult? ICES Journal of Marine Science, 67(6):1198–1205, September 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/67/
6/1198/736321.

Hussy:2011:RWB

[Hüs11] Karin Hüssy. Review of western Baltic cod (Gadus morhua) recruitment dynamics. ICES Journal of Marine Science, 68(7):1459–1471, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/68/
7/1459/658556.

Hedger:2011:BAC

[HUT^{+}11] Richard D. Hedger, Ingebrigt Uglem, Eva B. Thorstad, Bengt Finstad, Cedar M. Chittenden, Pablo Arechavala-Lopez, Arne J. Jensen, Rune Nilsen, and Finn Økland. Behaviour of Atlantic cod, a marine fish predator, during Atlantic salmon post-smolt migration. ICES Journal of Marine Science, 68(10):2152–2162, November 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/10/2152/611089.

Hutchings:2014:RCA

[Hut14] Jeffrey A. Hutchings. Renaissance of a caveat: Allee effects in marine fish. ICES Journal of Marine Science, 71 (8):2152–2157, October 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/71/8/2152/746541.

Hammer:2010:FSR

[HvDH^{+}10] Cornelius Hammer, Christian von Dorrien, Christopher C. E. Hopkins, Fritz W. Köster, Einar M. Nilssen, Michael St John, and Douglas C. Wilson. Framework of stock-recovery strategies: analyses of factors affecting success and failure. ICES Journal of Marine Science, 67(9):1849–1855, December 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/67/
9/1849/621142.

Herrmann:2015:URE

[HWK^{+}15] Bent Herrmann, Harald Wienbeck, Junita Diana Karlsen, Daniel Stepputtis, Erdmann Dahm, and Waldemar Moderhak. Understanding the release efficiency of Atlantic cod (Gadus morhua) from trawls with a square mesh panel: effects of panel area, panel position, and stimulation of escape response. ICES Journal of Marine Science, 72(2):686–696, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/2/686/2801328.

Heinrich:2016:FBE

[HWR^{+}16] Dennis D. U. Heinrich, Sue-Ann Watson, Jodie L. Rummer, Simon J. Brandl, Colin A. Simpfendorfer, Michelle R. Heupel, and Philip L. Munday. Foraging behaviour of the epaulette shark Hemiscyllium ocellatum is not affected by elevated CO_{2}. ICES Journal of Marine Science, 73(3):633– 640, February 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/73/3/633/2458696.

Han:2015:BGF

[HZZ^{+}15] Zhiqiang Han, Wei Zheng, Wenbin Zhu, Cungen Yu, Bonian Shui, and Tianxiang Gao. A barrier to gene flow in the Asian paddle crab, Charybdis japonica, in the Yellow Sea. ICES Journal of Marine Science, 72(5):1440–1448, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/5/1440/776914.

Irvine:2011:PSA

[IaF11] James R. Irvine and Masa aki Fukuwaka. Pacific salmon abundance trends and climate change. ICES Journal of Marine Science, 68(6):1122–1130, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/68/
6/1122/697357.

Iwamoto:2015:MDA

[IEL^{+}15] Eric M. Iwamoto, Anna E. Elz, Francisco J. García-De León, Claudia A. Silva-Segundo, Michael J. Ford, Wayne A. Palsson, and Richard G. Gustafson. Microsatellite DNA analysis of Pacific hake Merluccius productus population structure in the Salish Sea. ICES Journal of Marine Science, 72(9):2720–2731, November 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/9/2720/2458724.

Ibaibarriaga:2011:GIC

[IFU11] Leire Ibaibarriaga, Carmen Fernández, and Andrés Uriarte. Gaining information from commercial catch for a Bayesian two-stage biomass dynamic model: application to Bay of Biscay anchovy. ICES Journal of Marine Science, 68(7): 1435–1446, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/7/1435/659387.

Izquierdo-Gomez:2015:EEF

[IGGSAL^{+}15] David Izquierdo-Gómez, Daniel González-Silvera, Pablo Arechavala-López, José Ángel López-Jiménez, Just Tomás Bayle-Sempere, and Pablo Sánchez-Jerez. Exportation of excess feed from Mediterranean fish farms to local fisheries through different targeted fish species. ICES Journal of Marine Science, 72(3):930–938, March 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/3/930/2835886.

Ianelli:2011:EMS

[IHH^{+}11] James N. Ianelli, Anne B. Hollowed, Alan C. Haynie, Franz J. Mueter, and Nicholas A. Bond. Evaluating management strategies for eastern Bering Sea walleye pollock (Theragra chalcogramma) in a changing environment. ICES Journal of Marine Science, 68(6):1297–1304, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/6/1297/698941.

Homrum:2013:MSP

[íHHJ^{+}13] Eydna í Homrum, Bogi Hansen, Sigurdur Thór Jónsson, Kathrine Michalsen, Julian Burgos, David Righton, Petur Steingrund, Tore Jakobsen, Rógvi Mouritsen, Hjálmar Hátún, Hlynur Armannsson, and Jákup Sverri Joensen. Migration of saithe (Pollachius virens) in the Northeast Atlantic. ICES Journal of Marine Science, 70(4):782–792, July 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/70/4/782/726798.

Ingolfsson:2019:SFS

[IHL19] Ólafur Arnar Ingólfsson, Odd-Børre Humborstad, and Svein Løkkeborg. Selective flatfish seine: a knee-high demersal seine barely catches cod. ICES Journal of Marine Science, 76(4):1200–1208, July 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
4/1200/5299880.

Ito:2013:MER

[iIOKW13] Shin ichi Ito, Takeshi Okunishi, Michio J. Kishi, and Muyin Wang. Modelling ecological responses of Pacific saury (Cololabis saira) to future climate change and its uncertainty. ICES Journal of Marine Science, 70(5):980– 990, September 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/70/5/980/646796.

Irgens:2017:ODO

[IKF17] Christian Irgens, Olav S. Kjesbu, and Arild Folkvord. Ontogenetic development of otolith shape during settlement of juvenile Barents Sea cod (Gadus morhua). ICES Journal of Marine Science, 74(9):2389–2397, November 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/9/2389/3858383.

Ichinokawa:2017:SJF

[IOK17] Momoko Ichinokawa, Hiroshi Okamura, and Hiroyuki Kurota. The status of Japanese fisheries relative to fisheries around the world. ICES Journal of Marine Science, 74(5):

1277–1287, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/74/5/1277/3003306.

Ianelli:2015:EIP

[IS15] James N. Ianelli and Diana L. Stram. Estimating impacts of the pollock fishery bycatch on western Alaska Chinook salmon. ICES Journal of Marine Science, 72(4):1159–1172, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/4/1159/800295.

Ivory:2019:DSI

[ISL19] Jami A. Ivory, Deborah K. Steinberg, and Robert J. Latour. Diel, seasonal, and interannual patterns in mesozooplankton abundance in the Sargasso Sea. ICES Journal of Marine Science, 76(1):217–231, January 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/1/217/5145715.

Ivshin:2019:BST

[ITT19] Viktor A. Ivshin, Alexander G. Trofimov, and Oleg V. Titov. Barents Sea thermal frontal zones in 1960–2017: variability, weakening, shifting. ICES Journal of Marine Science, 76(S1):S3–S9, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/76/
Supplement_1/i3/5554545.

Isari:2016:LEE

[IZP^{+}16] Stamatina Isari, Soultana Zervoudaki, Janna Peters, Georgia Papantoniou, Carles Pelejero, and Enric Saiz. Lack of evidence for elevated CO_{2}-induced bottom-up effects on marine copepods: a dinoflagellate–calanoid prey–predator pair. ICES Journal of Marine Science, 73(3):650–658, February 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/3/650/2457980.

Jansen:2014:PRN

[Jan14] Teunis Jansen. Pseudocollapse and rebuilding of North Sea mackerel (Scomber scombrus). ICES Journal of Marine Science, 71(2):299–307, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/71/
2/299/787847.

Jorgensen:2014:IDS

[JBE14] Ole A. Jørgensen, François Bastardie, and Ole R. Eigaard. Impact of deep-sea fishery for Greenland halibut (Reinhardtius hippoglossoides) on non-commercial fish species off West Greenland. ICES Journal of Marine Science, 71(4): 845–852, May 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/71/4/845/668420.

Jackson:2011:RSC

[JBSD11] Thomas Jackson, Heather A. Bouman, Shubha Sathyendranath, and Emmanuel Devred. Regional-scale changes in diatom distribution in the Humboldt upwelling system as revealed by remote sensing: implications for fisheries. ICES Journal of Marine Science, 68(4):729–736, March 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/4/729/647744.

Jones:2015:MME

[JC15] Miranda C. Jones and William W. L. Cheung. Multi-model ensemble projections of climate change effects on global marine biodiversity. ICES Journal of Marine Science, 72(3): 741–752, March 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/72/3/741/2835882.

Jardim:2010:EMS

[JCA10] Ernesto Jardim, Santiago Cerviño, and Manuela Azevedo. Evaluating management strategies to implement the recovery plan for Iberian hake (Merluccius merluccius); the impact of censored catch information. ICES Journal of Marine Science, 67(2):258–269, March 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/67/
2/258/692183.

Jakubaviciute:2017:SDD

[JCLO17] Eglė Jakubavičiūtė, Michele Casini, Linas Ložys, and Jens Olsson. Seasonal dynamics in the diet of pelagic fish species in the southwest Baltic proper. ICES Journal of Marine Science, 74(3):750–758, March 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
3/750/2739037.

Jones:2016:EEI

[JCS16] Peter Jones, Alison Cathcart, and Douglas C. Speirs. Early evidence of the impact of preindustrial fishing on fish stocks from the mid-west and southeast coastal fisheries of Scotland in the 19th century. ICES Journal of Marine Science, 73(5):1404–1414, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
5/1404/2240616.

Jacinto:2010:SBP

[JCSC10] David Jacinto, Teresa Cruz, Teresa Silva, and João J. Castro. Stalked barnacle (Pollicipes pollicipes) harvesting in the Berlengas Nature Reserve, Portugal: temporal variation and validation of logbook data. ICES Journal of Marine Science, 67(1):19–25, January 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/1/19/595759.

Jouffre:2010:EEI

[JdFBB^{+}10] Didier Jouffre, Maria de Fatima Borges, Alida Bundy, Marta Coll, Ibrahima Diallo, Elizabeth A. Fulton, Jérôme Guitton, Pierre Labrosse, Khairdine ould Mohamed Abdellahi, Bora Masumbuko, and Djiga Thiao. Estimating EAF indicators from scientific trawl surveys: theoretical and practical concerns. ICES Journal of Marine Science, 67 (4):796–806, May 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/67/4/796/682787.

Johannessen:2012:CRF

[JDFN12] Tore Johannessen, Einar Dahl, Tone Falkenhaug, and Lars J. Naustvoll. Concurrent recruitment failure in gadoids and changes in the plankton community along the Norwegian Skagerrak coast after 2002. ICES Journal of Marine Science, 69(5):795–801, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/69/
5/795/646643.

Jackson:2014:FPM

[JDH^{+}14] E. L. Jackson, A. J. Davies, K. L. Howell, P. J. Kershaw, and J. M. Hall-Spencer. Future-proofing marine protected area networks for cold water coral reefs. ICES Journal of Marine Science, 71(9):2621–2629, November 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/9/2621/2798176.

Jimenez:2016:HEF

[JDM^{+}16] Haizea Jimenez, Pascal Dumas, David Mouillot, Lionel Bigot, and Jocelyne Ferraris. Harvesting effects on functional structure and composition of tropical invertebrate assemblages. ICES Journal of Marine Science, 73(2):420– 428, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/73/2/420/2614283.

Jennings:2013:WCP

[Jen13] S. Jennings. When can “principles” support advice on fisheries and environmental management? ICES Journal of Marine Science, 70(4):726–733, July 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/4/726/726221.

Ji:2017:CAS

[JFJ^{+}17] Rubao Ji, Zhixuan Feng, Benjamin T. Jones, Cameron Thompson, Changsheng Chen, Nicholas R. Record, and Jeffrey A. Runge. Coastal amplification of supply and transport (CAST): a new hypothesis about the persistence of Calanus finmarchicus in the Gulf of Maine. ICES Journal of Marine Science, 74(7):1865–1874, September 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/7/1865/2926067.

Jerome:2018:IRP

[JGCH18] J. M. Jerome, A. J. Gallagher, S. J. Cooke, and N. Hammerschlag. Integrating reflexes with physiological measures to evaluate coastal shark stress response to capture. ICES Journal of Marine Science, 75(2):796–804, March 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/2/796/4587596.

Johnsen:2013:SSS

[JH13] Espen Johnsen and Alf Harbitz. Small-scale spatial structuring of burrowed sandeels and the catching properties of the dredge. ICES Journal of Marine Science, 70(2): 379–386, March 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/70/2/379/798452.

Jacobsen:2012:DOS

[JHB^{+}12] Jan Arge Jacobsen, Lars P. Hansen, Vegar Bakkestuen, Rune Halvorsen, David G. Reddin, Jonathan White, Niall Ó Maoiléidigh, Ian C. Russell, E. C. E. (Ted) Potter, Mark Fowler, Gordon W. Smith, Kjell A. Mork, Arni Isaksson, Sumarlidi Oskarsson, Lars Karlsson, and Stig Pedersen. Distribution by origin and sea age of Atlantic salmon (Salmo salar) in the sea around the Faroe Islands based on analysis of historical tag recoveries. ICES Journal of Marine Science, 69(9):1598–1608, November 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/9/1598/639760.

Jaiteh:2017:SFE

[JHB^{+}17] Vanessa F. Jaiteh, Adrian R. Hordyk, Matías Braccini, Carol Warren, and Neil R. Loneragan. Shark finning in eastern Indonesia: assessing the sustainability of a datapoor fishery. ICES Journal of Marine Science, 74(1):242– 253, January 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/74/1/242/2528220.

Johannesen:2012:CBS

[JIB^{+}12] Edda Johannesen, Randi B. Ingvaldsen, Bjarte Bogstad, Padmini Dalpadado, Elena Eriksen, Harald Gjøsæter, Tor Knutsen, Mette Skern-Mauritzen, and Jan Erik Stiansen. Changes in Barents Sea ecosystem state, 1970–2009: climate fluctuations, human impact, and trophic interactions. ICES Journal of Marine Science, 69(5):880–889, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/69/5/880/650923.

Johannesen:2014:CCB

[JIB^{+}14] Edda Johannesen, Randi B. Ingvaldsen, Bjarte Bogstad, Padmini Dalpadado, Elena Eriksen, Harald Gjøsæter, Tor Knutsen, Mette Skern-Mauritzen, and Jan Erik Stiansen. A correction to “Changes in Barents Sea ecosystem state, 1970–2009: climate fluctuations, human impact, and trophic interactions”. ICES Journal of Marine Science, 71(5):1308–1309, July 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/71/
5/1308/648376.

Jorde:2018:WFW

[JKS^{+}18] Per Erik Jorde, Alf Ring Kleiven, Marte Sodeland, Esben Moland Olsen, Keno Ferter, Sissel Jentoft, and Halvor Knutsen. Who is fishing on what stock: populationof-origin of individual cod (Gadus morhua) in commercial and recreational fisheries. ICES Journal of Marine Science, 75(6):2153–2162, November 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/75/
6/2153/5057978.

Jansen:2015:NAR

[JKvdK^{+}15] Teunis Jansen, Kasper Kristensen, Jeroen van der Kooij, Søren Post, Andrew Campbell, Kjell Rong Utne, Pablo Carrera, Jan Arge Jacobsen, Asta Gudmundssdottir, Beatriz A. Roel, and Emma M. C. Hatfield. Nursery areas and recruitment variation of Northeast Atlantic mackerel (Scomber scombrus). ICES Journal of Marine Science, 72(6):1779–1789, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/72/
6/1779/917521.

Jennings:2012:IEF

[JL12a] Simon Jennings and Will J. F. Le Quesne. Integration of environmental and fishery management in Europe. ICES Journal of Marine Science, 69(8):1329–1332, September 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/8/1329/701991.

Jennings:2012:DFG

[JL12b] Simon Jennings and Janette Lee. Defining fishing grounds with vessel monitoring system data. ICES Journal of Marine Science, 69(1):51–63, January 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/69/
1/51/671252.

Jennings:2012:AFF

[JLH12] S. Jennings, J. Lee, and J. G. Hiddink. Assessing fishery footprints and the trade-offs between landings value, habitat sensitivity, and fishing impacts to inform marine spatial planning and an ecosystem approach. ICES Journal of Marine Science, 69(6):1053–1063, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/6/1053/617910.

Jech:2017:WKA

[JLL17] J. Michael Jech, Gareth L. Lawson, and Andone C. Lavery. Wideband (15–260 kHz) acoustic volume backscattering spectra of northern krill (Meganyctiphanes norvegica) and butterfish (Peprilus triacanthus). ICES Journal of Marine Science, 74(8):2249–2261, September 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/8/2249/4096423.

Jorgensen:2015:DBM

[JLS^{+}15] Lis Lindal Jørgensen, Pavel Ljubin, Hein Rune Skjoldal, Randi B. Ingvaldsen, Natalia Anisimova, and Igor Manushin. Distribution of benthic megafauna in the Barents Sea: baseline for an ecosystem approach to management. ICES Journal of Marine Science, 72(2):595–613, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/2/595/672811.

Jennings:2012:RCS

[JMF12] Gemma Jennings, Derek J. McGlashan, and Robert W. Furness. Responses to changes in sprat abundance of common tern breeding numbers at 12 colonies in the Firth of Forth, east Scotland. ICES Journal of Marine Science, 69 (4):572–577, May 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/69/4/572/633907.

Janssen:2014:LGL

[JMHS14] John Janssen, J. Ellen Marsden, Thomas R. Hrabik, and Jason D. Stockwell. Are the Laurentian Great Lakes great enough for Hjort? ICES Journal of Marine Science, 71(8):2242–2251, October 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/71/
8/2242/750919.

Jardim:2015:WIS

[JMM^{+}15a] Ernesto Jardim, Colin P. Millar, Iago Mosqueira, Finlay Scott, Giacomo Chato Osio, Marco Ferretti, Nekane Alzorriz, and Alessandro Orio. What if stock assessment is as simple as a linear model? The a4a initiative. ICES Journal of Marine Science, 72(1):232–236, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/1/232/2804293.

Johnson:2015:TVN

[JMM^{+}15b] Kelli F. Johnson, Cole C. Monnahan, Carey R. McGilliard, Katyana A. Vert-pre, Sean C. Anderson, Curry J. Cunningham, Felipe Hurtado-Ferro, Roberto R. Licandeo, Melissa L. Muradian, Kotaro Ono, Cody S. Szuwalski, Juan L. Valero, Athol R. Whitten, and A. E. Punt. Time-varying natural mortality in fisheries stock assessment models: identifying a default approach. ICES Journal of Marine Science, 72(1):137–150, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/1/137/823490.

Jokinen:2019:EGB

[JMM19] Henri Jokinen, Paolo Momigliano, and Juha Merilä. From ecology to genetics and back: the tale of two flounder species in the Baltic Sea. ICES Journal of Marine Science, 76(7):2267–2275, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
7/2267/5552826.

Jobling:2017:MTF

[Job17] Malcolm Jobling. More than a fair share of good luck. ICES Journal of Marine Science, 74(5):1249–1255, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/5/1249/2999344.

Jorgensen:2014:CBE

[JOF14] Christian Jørgensen, Anders Frugård Opdal, and Øyvind Fiksen. Can behavioural ecology unite hypotheses for fish recruitment? ICES Journal of Marine Science, 71(4):909– 917, May 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/71/4/909/664709.

Jokiel:2016:PIO

[Jok16] Paul L. Jokiel. Predicting the impact of ocean acidification on coral reefs: evaluating the assumptions involved. ICES Journal of Marine Science, 73(3):550–557, February 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/3/550/2458699.

Jones:2014:CWP

[Jon14] Cynthia M. Jones. Can we predict the future: juvenile finfish and their seagrass nurseries in the Chesapeake Bay. ICES Journal of Marine Science, 71(3):681–688, March 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/3/681/633484.

Jiao:2016:ISS

[JOSO16] Yan Jiao, Rob O’Reilly, Eric Smith, and Don Orth. Integrating spatial synchrony/asynchrony of population distribution into stock assessment models: a spatial hierarchical Bayesian statistical catch-at-age approach. ICES Journal of Marine Science, 73(7):1725–1738, July 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/7/1725/2458716.

Jensen:2012:AFS

[JÓT^{+}12] Arne J. Jensen, Niall Ó Maoiléidigh, Katie Thomas, Sigurdur M. Einarsson, Monika Haugland, Jaakko Erkinaro, Peder Fiske, Kevin D. Friedland, Asta K. Gudmundsdottir, Jari Haantie, Marianne Holm, Jens Christian Holst, Jan Arge Jacobsen, Jan G. Jensås, Jorma Kuusela, Webjørn Melle, Kjell Arne Mork, Vidar Wennevik, and Gunnel M. Østborg. Age and fine-scale marine growth of Atlantic salmon post-smolts in the Northeast Atlantic. ICES Journal of Marine Science, 69(9):1668–1677, November 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/69/9/1668/637962.

Jorstad:2014:GMF

[JOvdM^{+}14] Knut E. Jørstad, Håkon Otterå, Terje van der Meeren, Geir Dahle, Ole I. Paulsen, Gunnar Bakke, and Terje Svåsand. Genetic marking of farmed Atlantic cod (Gadus morhua L.) and detection of escapes from a commercial cod farm. ICES Journal of Marine Science, 71(3):574–584, March 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/3/574/638263.

Jang:2011:ROM

[JPPY11] Chan Joo Jang, Jisoo Park, Taewook Park, and Sinjae Yoo. Response of the ocean mixed layer depth to global warming and its impact on primary production: a case for the North Pacific Ocean. ICES Journal of Marine Science, 68(6):996– 1007, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/68/6/996/712683.

Jorgensen:2016:VMS

[JPTC16] Lis Lindal Jørgensen, Benjamin Planque, Trude Hauge Thangstad, and Grégoire Certain. Vulnerability of megabenthic species to trawling in the Barents Sea. ICES Journal of Marine Science, 73(suppl 1):S84–S97, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/73/suppl_1/i84/2573990.

Jansson:2017:GAG

[JQD^{+}17] Eeva Jansson, María Quintela, Geir Dahle, Jon Albretsen, Halvor Knutsen, Carl André, Åsa Strand, Stein Mortensen, John B. Taggart, Egil Karlsbakk, Bjørn Olav Kvamme, and Kevin A. Glover. Genetic analysis of goldsinny wrasse reveals evolutionary insights into population connectivity and potential evidence of inadvertent translocation via aquaculture. ICES Journal of Marine Science, 74(8):2135– 2147, September 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/74/8/2135/3738529.

Jackson:2016:AEH

[JRG^{+}16] Gary Jackson, Karina L. Ryan, Timothy J. Green, Kenneth H. Pollock, and Jeremy M. Lyle. Assessing the effectiveness of harvest tags in the management of a smallscale, iconic marine recreational fishery in Western Australia. ICES Journal of Marine Science, 73(10):2666–2676, November 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/10/2666/2647095.

Jensen:2011:ILS

[JRWM11] Henrik Jensen, Anna Rindorf, Peter J. Wright, and Henrik Mosegaard. Inferring the location and scale of mixing between habitat areas of lesser sandeel through information from the fishery. ICES Journal of Marine Science, 68(1): 43–51, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/1/43/631084.

Jiao:2012:MNS

[JSOO12] Yan Jiao, Eric P. Smith, Rob O’Reilly, and Donald J. Orth. Modelling non-stationary natural mortality in catch-at-age models. ICES Journal of Marine Science, 69(1):105–118, January 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/69/1/105/674280.

Jacobsen:2019:DMV

[JTE19] Nis S. Jacobsen, James T. Thorson, and Timothy E. Essington. Detecting mortality variation to enhance forage fish population assessments. ICES Journal of Marine Science, 76(1):124–135, January 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
1/124/5181372.

Jonsdottir:2018:IDB

[JTJ18] Ingibjörg G. Jónsdóttir, Gudrún G. Thórarinsdóttir, and Jónas P. Jonasson. Influence of decreased biomass on the ogive of sex change of northern shrimp (Pandalus borealis). ICES Journal of Marine Science, 75(3):1054–1062, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/3/1054/4633711.

Josset:2016:PRP

[JTM^{+}16] Q. Josset, T. Trancart, V. Mazel, F. Charrier, L. Frotté, A. Acou, and E. Feunteun. Pre-release processes influencing short-term mortality of glass eels in the French eel (Anguilla anguilla, Linnaeus 1758) stocking programme. ICES Journal of Marine Science, 73(1):150–157, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/1/150/2458188.

Jonsson:2012:HCS

[JV12a] Steingrímur Jónsson and Hédinn Valdimarsson. Hydrography and circulation over the southern part of the kolbeinsey ridge. ICES Journal of Marine Science, 69(7):1255– 1262, September 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/69/7/1255/749843.

Jonsson:2012:WMT

[JV12b] Steingrímur Jónsson and Hédinn Valdimarsson. Water mass transport variability to the North Icelandic shelf, 1994–2010. ICES Journal of Marine Science, 69(5):809– 815, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/69/5/809/647908.

Jennings:2015:TLM

[JvdM15] Simon Jennings and Johan van der Molen. Trophic levels of marine consumers from nitrogen stable isotope analysis: estimation and uncertainty. ICES Journal of Marine Science, 72(8):2289–2300, September 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/8/2289/2459088.

Juntunen:2012:BSM

[JVPM12] Teppo Juntunen, Jarno Vanhatalo, Heikki Peltonen, and Samu Mäntyniemi. Bayesian spatial multispecies modelling to assess pelagic fish stocks from acoustic- and trawl-survey data. ICES Journal of Marine Science, 69(1):95–104, January 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/69/1/95/673665.

Judkins:2010:CSR

[JVRT10] Heather L. Judkins, Michael Vecchione, Clyde F. E. Roper, and Joseph Torres. Cephalopod species richness in the wider Caribbean region. ICES Journal of Marine Science, 67(7):1392–1400, October 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/67/
7/1392/664222.

Jorgensbye:2018:MMS

[JW18] Helle Jørgensbye and Susse Wegeberg. Mapping of marine sediments on the Greenland West Coast: contributions of fishers’ ecological knowledge. ICES Journal of Marine Science, 75(5):1768–1778, September 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/5/1768/4958791.

Johansen:2018:RTG

[JWS^{+}18] Torild Johansen, Jon-Ivar Westgaard, Bjørghild B. Seliussen, Kjell Nedreaas, Geir Dahle, Kevin A. Glover, Roger Kvalsund, and Asgeir Aglen. “Real-time” genetic monitoring of a commercial fishery on the doorstep of an MPA reveals unique insights into the interaction between coastal and migratory forms of the Atlantic cod. ICES Journal of Marine Science, 75(3):1093–1104, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/75/
3/1093/4748806.

King:2011:CFC

[KAH^{+}11] Jacquelynne R. King, Vera N. Agostini, Christopher J. Harvey, Gordon A. McFarlane, Michael G. G. Foreman, James E. Overland, Emanuele Di Lorenzo, Nicholas A. Bond, and Kerim Y. Aydin. Climate forcing and the California Current ecosystem. ICES Journal of Marine Science, 68(6):1199–1216, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/68/
6/1199/698284.

Kamykowski:2014:TCA

[Kam14] Daniel Kamykowski. Twentieth century Atlantic meridional overturning circulation as an indicator of global ocean multidecadal variability: influences on sea level anomalies and small pelagic fishery synchronies. ICES Journal of Marine Science, 71(3):455–468, March 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/3/455/636969.

Kane:2011:MVP

[Kan11] Joseph Kane. Multiyear variability of phytoplankton abundance in the Gulf of Maine. ICES Journal of Marine Science, 68(9):1833–1841, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/68/
9/1833/668901.

Keyl:2011:IVS

[KAT11] Friedemann Keyl, Juan Argüelles, and Ricardo Tafur. Interannual variability in size structure, age, and growth of jumbo squid (Dosidicus gigas) assessed by modal progression analysis. ICES Journal of Marine Science, 68(3):

507–518, March 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/3/507/655651.

Kaltenberg:2013:IPC

[KBB13] Amanda M. Kaltenberg and Kelly J. Benoit-Bird. Intrapatch clustering in mysid swarms revealed through multifrequency acoustics. ICES Journal of Marine Science, 70 (4):883–891, July 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/70/4/883/725234.

Keister:2012:ZPC

[KBC^{+}12] Julie E. Keister, Delphine Bonnet, Sanae Chiba, Catherine L. Johnson, David L. Mackas, and Ruben Escribano. Zooplankton population connections, community dynamics, and climate variability. ICES Journal of Marine Science, 69(3):347–350, May 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/69/3/347/601237.

Kulke:2018:IVD

[KBH^{+}18] Rini Kulke, Viola Bödewadt, Kristin Hänselmann, Jens-Peter Herrmann, and Axel Temming. Ignoring the vertical dimension: biased view on feeding dynamics of vertically migrating sprat (Sprattus sprattus). ICES Journal of Marine Science, 75(7):2450–2462, December 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/7/2450/5123541.

Kilduff:2014:STC

[KBT14] D. Patrick Kilduff, Louis W. Botsford, and Steven L. H. Teo. Spatial and temporal covariability in early ocean survival of Chinook salmon (Oncorhynchus tshawytscha) along the west coast of North America. ICES Journal of Marine Science, 71(7):1671–1682, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/7/1671/671710.

Kolding:2016:FIF

[KBvZP16] Jeppe Kolding, Alida Bundy, Paul A. M. van Zwieten, and Michael J. Plank. Fisheries, the inverted food pyramid. ICES Journal of Marine Science, 73(6):1697–1713, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/6/1697/2458720.

Kaplan:2014:SMI

[KCA^{+}14] David M. Kaplan, Emmanuel Chassot, Justin M. Amandé, Sibylle Dueri, Hervé Demarcq, Laurent Dagorn, and Alain Fonteneau. Spatial management of Indian Ocean tropical tuna fisheries: potential and perspectives. ICES Journal of Marine Science, 71(7):1728–1749, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/7/1728/665606.

Kaplan:2017:UEE

[KCF^{+}17] David M. Kaplan, Marion Cuif, Cécile Fauvelot, Laurent Vigliola, Tri Nguyen-Huu, Josina Tiavouane, and Christophe Lett. Uncertainty in empirical estimates of marine larval connectivity. ICES Journal of Marine Science, 74(6):1723–1734, July 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
6/1723/2741993.

Kerr:2014:CMB

[KCK14] Lisa A. Kerr, Steven X. Cadrin, and Adrienne I. Kovach. Consequences of a mismatch between biological and management units on our perception of Atlantic cod off New England. ICES Journal of Marine Science, 71(6):1366– 1381, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/71/6/1366/2835585.

Koenker:2018:ITF

[KCL18] Brittany L. Koenker, Louise A. Copeman, and Benjamin J. Laurel. Impacts of temperature and food availability on the condition of larval Arctic cod (Boreogadus saida) and walleye pollock (Gadus chalcogrammus). ICES Journal of Marine Science, 75(7):2370–2385, December 2018.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/7/2370/4992259.

Kritzer:2019:RHC

[KCMS19] J. P. Kritzer, C. Costello, T. Mangin, and S. L. Smith. Responsive harvest control rules provide inherent resilience to adverse effects of climate change and scientific uncertainty. ICES Journal of Marine Science, 76(6):1424–1435, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/6/1424/5425355.

Kobayashi:2011:LTC

[KCP^{+}11] Donald R. Kobayashi, I-Jiunn Cheng, Denise M. Parker, Jeffrey J. Polovina, Naoki Kamezaki, and George H. Balazs. Loggerhead turtle (Caretta caretta) movement off the coast of Taiwan: characterization of a hotspot in the East China Sea and investigation of mesoscale eddies. ICES Journal of Marine Science, 68(4):707–718, March 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/4/707/648332.

Kerr:2010:SMT

[KCS10] Lisa A. Kerr, Steven X. Cadrin, and Dave H. Secor. Simulation modelling as a tool for examining the consequences of spatial structure and connectivity on local and regional population dynamics. ICES Journal of Marine Science, 67(8):1631–1639, November 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/8/1631/604016.

Kopf:2011:AGS

[KDBP11] R. Keller Kopf, Peter S. Davie, Donald Bromhead, and Julian G. Pepperell. Age and growth of striped marlin (Kajikia audax) in the Southwest Pacific Ocean. ICES Journal of Marine Science, 68(9):1884–1895, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/9/1884/666536.

Karlsson:2016:WGI

[KDF^{+}16] Sten Karlsson, Ola H. Diserud, Peder Fiske, Kjetil Hindar, and W. Stewart Grant. Widespread genetic introgression of escaped farmed Atlantic salmon in wild salmon populations. ICES Journal of Marine Science, 73(10):2488–2498, November 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/10/2488/2647116.

Kennedy:2019:BHL

[KDF^{+}19a] James Kennedy, Caroline M. F. Durif, Ann-Britt Florin, Alain Fréchet, Johanne Gauthier, Karin Hüssy, Sigurdur Thór Jónsson, Halldór Gunnar Ólafsson, Søren Post, and Rasmus B. Hedeholm. A brief history of lumpfishing, assessment, and management across the North Atlantic. ICES Journal of Marine Science, 76(1):181–191, January 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/1/181/5133607.

Koslow:2019:ERM

[KDF^{+}19b] J. Anthony Koslow, Pete Davison, Erica Ferrer, S. Patricia A. Jiménez Rosenberg, Gerardo Aceves-Medina, and William Watson. The evolving response of mesopelagic fishes to declining midwater oxygen concentrations in the southern and central California Current. ICES Journal of Marine Science, 76(3):626–638, May 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/3/626/5184892.

Kempf:2010:IPP

[KDH^{+}10] Alexander Kempf, Gjert Endre Dingsør, Geir Huse, Morten Vinther, Jens Floeter, and Axel Temming. The importance of predator–prey overlap: predicting North Sea cod recovery with a multispecies assessment model. ICES Journal of Marine Science, 67(9):1989–1997, December 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/9/1989/620275.

Kolling:2014:EDX

[KdS14] Juliana Almeida Kolling and Antônio Olinto Ávila da Silva. Evaluation of determinants of Xiphopenaeus kroyeri (Heller, 1862) catch abundance along a Southwest Atlantic subtropical shelf. ICES Journal of Marine Science, 71(7):1793–1804, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/71/
7/1793/663987.

Kontogianni:2014:CGF

[KE14] A. D. Kontogianni and C. J. Emmanouilides. The cost of a gelatinous future and loss of critical habitats in the Mediterranean. ICES Journal of Marine Science, 71(4): 853–866, May 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/71/4/853/669030.

Koehn:2017:TOB

[KEM^{+}17] Laura E. Koehn, Timothy E. Essington, Kristin N. Marshall, William J. Sydeman, Amber I. Szoboszlai, and Julie A. Thayer. Trade-offs between forage fish fisheries and their predators in the California Current. ICES Journal of Marine Science, 74(9):2448–2458, November 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/9/2448/3823335.

Kokkalis:2017:EUD

[KET^{+}17] Alexandros Kokkalis, Anne Maria Eikeset, Uffe H. Thygesen, Petur Steingrund, and Ken H. Andersen. Estimating uncertainty of data limited stock assessments. ICES Journal of Marine Science, 74(1):69–77, January 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/1/69/2669561.

Kordjazi:2015:ESR

[KFBG15] Ziya Kordjazi, Stewart Frusher, Colin D. Buxton, and Caleb Gardner. Estimating survival of rock lobsters from long-term tagging programmes: how survey number and interval influence estimates. ICES Journal of Marine Science, 72(S1):S244–S251, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/72/
suppl_1/i244/617934.

Konigson:2015:CPB

[KFL^{+}15] Sara J. Königson, Ronny E. Fredriksson, Sven-Gunnar Lunneryd, Patrick Strömberg, and Ulf M. Bergström. Cod pots in a Baltic fishery: are they efficient and what affects their efficiency? ICES Journal of Marine Science, 72(5):1545– 1554, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/5/1545/764414.

Kynoch:2015:STM

[KFN15] R. J. Kynoch, R. J. Fryer, and F. C. Neat. A simple technical measure to reduce bycatch and discard of skates and sharks in mixed-species bottom-trawl fisheries. ICES Journal of Marine Science, 72(6):1861–1868, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/6/1861/921176.

Kristiansen:2016:PSC

[KGH^{+}16] Inga Kristiansen, Eilif Gaard, Hjálmar Hátún, Sigrún Jónasdóttir, and A. Sofia A. Ferreira. Persistent shift of Calanus spp. in the southwestern Norwegian Sea since 2003, linked to ocean climate. ICES Journal of Marine Science, 73(5):1319–1329, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
5/1319/2296547.

Kidokoro:2010:ICR

[KGN^{+}10] Hideaki Kidokoro, Tsuneo Goto, Toru Nagasawa, Hiroshi Nishida, Tatsuro Akamine, and Yasunori Sakurai. Impact of a climate regime shift on the migration of Japanese common squid (Todarodes pacificus) in the Sea of Japan. ICES Journal of Marine Science, 67(7):1314–1322, October 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/7/1314/662573.

Krumme:2015:ASA

[KGP^{+}15] Uwe Krumme, Tommaso Giarrizzo, Rodolfo Pereira, Allan Jamesson Silva de Jesus, Christoph Schaub, and Ulrich Saint-Paul. Airborne synthetic-aperture radar (SAR) imaging to help assess impacts of stationary fishing gear on the north Brazilian mangrove coast. ICES Journal of Marine Science, 72(3):939–951, March 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/3/939/693816.

Kwasniewski:2012:ICZ

[KGW^{+}12] Slawomir Kwasniewski, Marta Gluchowska, Wojciech Walkusz, Nina J. Karnovsky, Dariusz Jakubas, Katarzyna Wojczulanis-Jakubas, Ann M. A. Harding, Ilona Goszczko, Malgorzata Cisek, Agnieszka Beszczynska-Möller, Waldemar Walczowski, Jan M. Weslawski, and Lech Stempniewicz. Interannual changes in zooplankton on the West Spitsbergen Shelf in relation to hydrography and their consequences for the diet of planktivorous seabirds. ICES Journal of Marine Science, 69(5):890–901, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/69/
5/890/653180.

Kolding:2016:BHU

[KGZH16] Jeppe Kolding, Serge M. Garcia, Shijie Zhou, and Mikko Heino. Balanced harvest: utopia, failure, or a functional strategy? ICES Journal of Marine Science, 73(6):1616– 1622, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/6/1616/2459106.

Kuparinen:2019:WPF

[KH19] Anna Kuparinen and Jeffrey A. Hutchings. When phenotypes fail to illuminate underlying genetic processes in fish and fisheries science. ICES Journal of Marine Science, 76(4):999–1006, July 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
4/999/5306594.

Kerr:2017:LLP

[KHC^{+}17] Lisa A. Kerr, Niels T. Hintzen, Steven X. Cadrin, Lotte Worsøe Clausen, Mark Dickey-Collas, Daniel R. Goethel, Emma M. C. Hatfield, Jacob P. Kritzer, and Richard D. M. Nash. Lessons learned from practical approaches to reconcile mismatches between biological population structure and stock units of marine fish. ICES Journal of Marine Science, 74(6):1708–1722, July 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/6/1708/2629217.

Kaplan:2014:FAB

[KHF14] Isaac C. Kaplan, Daniel S. Holland, and Elizabeth A. Fulton. Finding the accelerator and brake in an individual quota fishery: linking ecology, economics, and fleet dynamics of US West Coast trawl fisheries. ICES Journal of Marine Science, 71(2):308–319, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/2/308/779541.

Kaplan:2017:EII

[KHH^{+}17a] Katherine A. Kaplan, Deborah R. Hart, Karen Hopkins, Scott Gallager, Amber York, Richard Taylor, and Patrick J. Sullivan. Evaluating the interaction of the invasive tunicate Didemnum vexillum with the Atlantic sea scallop Placopecten magellanicus on open and closed fishing grounds of Georges Bank. ICES Journal of Marine Science, 74(9):2470–2479, November 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
9/2470/3831747.

Koster:2017:EBC

[KHH^{+}17b] Friedrich W. Köster, Bastian Huwer, Hans-Harald Hinrichsen, Viola Neumann, Andrei Makarchouk, Margit Eero, Burkhard V. Dewitz, Karin Hüssy, Jonna Tomkiewicz, Piotr Margonski, Axel Temming, Jens-Peter Hermann, Daniel Oesterwind, Jan Dierking, Paul Kotterba, and Maris Plikshs. Eastern Baltic cod recruitment revisited — dynamics and impacting factors. ICES Journal of Marine Science, 74 (1):3–19, January 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/74/1/3/2418194.

Kraak:2019:RLP

[KHMS19] Sarah B. M. Kraak, Stefanie Haase, Cóilín Minto, and Juan Santos. The Rosa Lee phenomenon and its consequences for fisheries advice on changes in fishing mortality or gear selectivity. ICES Journal of Marine Science, 76(7):2179– 2192, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/76/7/2179/5525343.

Kotwicki:2015:FAA

[KHPI15] Stan Kotwicki, John K. Horne, André E. Punt, and James N. Ianelli. Factors affecting the availability of walleye pollock to acoustic and bottom trawl survey gear. ICES Journal of Marine Science, 72(5):1425–1439, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/5/1425/773038.

Kaifu:2018:DWC

[KIA^{+}18] Kenzo Kaifu, Hikaru Itakura, Yosuke Amano, Kotaro Shirai, Kazuki Yokouchi, Ryoshiro Wakiya, Naoko Murakami-Sugihara, Izumi Washitani, and Takashi Yada. Discrimination of wild and cultured Japanese eels based on otolith stable isotope ratios. ICES Journal of Marine Science, 75(2): 719–726, March 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/2/719/4159462.

Kim:2012:RSO

[Kim12] Sen Tok Kim. A review of the Sea of Okhotsk ecosystem response to the climate with special emphasis on fish populations. ICES Journal of Marine Science, 69(7):1123– 1133, September 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/69/7/1123/750006.

Kotwicki:2014:CDD

[KIP14] Stan Kotwicki, James N. Ianelli, and André E. Punt. Correcting density-dependent effects in abundance estimates from bottom-trawl surveys. ICES Journal of Marine Science, 71(5):1107–1116, July 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/71/
5/1107/645300.

Knutsen:2015:DPG

[KJG^{+}15] Halvor Knutsen, Per Erik Jorde, Enrique Blanco Gonzalez, Ole Ritzau Eigaard, Ricardo T. Pereyra, Hanne Sannæs, Mikael Dahl, Carl André, and Guldborg Søvik. Does population genetic structure support present management regulations of the northern shrimp (Pandalus borealis) in Skagerrak and the North Sea? ICES Journal of Marine Science, 72(3):863–871, March 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/72/3/863/698697.

Kennedy:2015:MFL

[KJKÓ15] James Kennedy, Sigurdur T. Jónsson, Jacob M. Kasper, and Halldór G. Ólafsson. Movements of female lumpfish (Cyclopterus lumpus) around Iceland. ICES Journal of Marine Science, 72(3):880–889, March 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/72/
3/880/2835881.

Kennedy:2016:OVM

[KJÓK16] James Kennedy, Sigurdur T. Jónsson, Halldór G. Ólafsson, and Jacob M. Kasper. Observations of vertical movements and depth distribution of migrating female lumpfish (Cyclopterus lumpus) in Iceland from data storage tags and trawl surveys. ICES Journal of Marine Science, 73(4): 1160–1169, March 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/73/4/1160/2458719.

Kenny:2018:ACH

[KJW^{+}18] Andrew J. Kenny, Chris Jenkins, Daniel Wood, Stefan G. Bolam, Peter Mitchell, Callum Scougal, and Adrian Judd. Assessing cumulative human activities, pressures, and impacts on North Sea benthic habitats using a biological traits approach. ICES Journal of Marine Science, 75(3):1080– 1092, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/3/1080/4616494.

Kavousi:2018:CCC

[KK18] Javid Kavousi and Gunnar Keppel. Clarifying the concept of climate change refugia for coral reefs. ICES Journal of Marine Science, 75(1):43–49, January 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/1/43/4080230.

Kourantidou:2019:SSC

[KK19] Melina Kourantidou and Brooks A. Kaiser. Sustainable seafood certifications are inadequate to challenges of ecosystem change. ICES Journal of Marine Science, 76(4):794– 802, July 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/4/794/5288560.

Kluger:2019:CAE

[KKAV^{+}19] Lotta Clara Kluger, Sophia Kochalski, Arturo Aguirre-Velarde, Ivonne Vivar, and Matthias Wolff. Coping with abrupt environmental change: the impact of the coastal El Niño 2017 on artisanal fisheries and mariculture in North Peru. ICES Journal of Marine Science, 76(4):1122–1130, July 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/4/1122/5230896.

Kerr:2019:SLB

[KKC19] Lisa A. Kerr, Jacob P. Kritzer, and Steven X. Cadrin. Strengths and limitations of before–after–control–impact analysis for testing the effects of marine protected areas on managed populations. ICES Journal of Marine Science, 76(4):1039–1051, July 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
4/1039/5345117.

Kuparinen:2014:IED

[KKH14] Anna Kuparinen, David M. Keith, and Jeffrey A. Hutchings. Increased environmentally driven recruitment variability decreases resilience to fishing and increases uncertainty of recovery. ICES Journal of Marine Science, 71(6):1507–1514, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/71/
6/1507/623132.

Kuningas:2013:KWP

[KKLM13] Sanna Kuningas, Petter H. Kvadsheim, Frans-Peter A. Lam, and Patrick J. O. Miller. Killer whale presence in relation to naval sonar activity and prey abundance in northern Norway. ICES Journal of Marine Science, 70(7):1287–1293, November 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/70/7/1287/611145.

Kinoshita:2014:NTP

[KKMS14] Hikari Kinoshita, Yasuhiro Kamimura, Ken-Ichiro Mizuno, and Jun Shoji. Night-time predation on post-settlement Japanese black rockfish Sebastes cheni in a macroalgal bed: effect of body length on the predation rate. ICES Journal of Marine Science, 71(4):1022–1029, May 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/4/1022/663977.

Kim:2012:CFY

[KKZ^{+}12] Suam Kim, Sukyung Kang, Chang-Ik Zhang, Hyunju Seo, Minho Kang, and Jung J. Kim. Comparison of fisheries yield and oceanographic features at the southern boundaries of the western and eastern Subarctic Pacific Ocean. ICES Journal of Marine Science, 69(7):1141–1147, September 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/69/7/1141/742487.

Kaartvedt:2019:EOT

[KLA19] Stein Kaartvedt, Tom J. Langbehn, and Dag L. Aksnes. Enlightening the ocean’s twilight zone. ICES Journal of Marine Science, 76(4):803–812, July 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/4/803/5306603.

Khan:2013:FSS

[KLC13] Amina H. Khan, Elisabeth Levac, and Gail L. Chmura. Future sea surface temperatures in large marine ecosystems of the Northwest Atlantic. ICES Journal of Marine Science, 70(5):915–921, September 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/70/5/915/644252.

Koenker:2018:ETF

[KLCC18] Brittany L. Koenker, Benjamin J. Laurel, Louise A. Copeman, and Lorenzo Ciannelli. Effects of temperature and food availability on the survival and growth of larval Arctic cod (Boreogadus saida) and walleye pollock (Gadus chalcogrammus). ICES Journal of Marine Science, 75(7):2386–2402, December 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/75/
7/2386/5033020.

Kindt-Larsen:2011:FDF

[KLKD11] Lotte Kindt-Larsen, Eskild Kirkegaard, and Jørgen Dalskov. Fully documented fishery: a tool to support a catch quota management system. ICES Journal of Marine Science, 68(8):1606–1610, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/68/
8/1606/749597.

Kim:2010:MAB

[KLN^{+}10] Eunhye Kim, Hyungbeen Lee, Jungyul Na, Jee Woong Choi, and Donhyug Kang. 5-MHz acoustic-backscatter measurements of Cochlodinium polykrikoides blooms in Korean coastal waters. ICES Journal of Marine Science, 67 (8):1759–1765, November 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/67/8/1759/606728.

Kaplan:2016:GPT

[KM16] Isaac C. Kaplan and Kristin N. Marshall. A guinea pig’s tale: learning to review end-to-end marine ecosystem models for management applications. ICES Journal of Marine Science, 73(7):1715–1724, July 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
7/1715/2458746.

Karlson:2013:ADA

[KMF13] Stine Karlson, Kathrine Michalsen, and Arild Folkvord. Age determination of Atlantic halibut (Hippoglossus hippoglossus L.) along the coast of Norway: status and improvements. ICES Journal of Marine Science, 70(1):50–55, January 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/70/1/50/662683.

Katsanevakis:2010:LPP

[KMK10] Stelios Katsanevakis, Christos D. Maravelias, and Laurie T. Kell. Landings profiles and potential métiers in Greek set longliners. ICES Journal of Marine Science, 67(4):646– 656, May 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/67/4/646/681472.

Kulagin:2017:GMD

[KN17] Dmitry N. Kulagin and Tatiana V. Neretina. Genetic and morphological diversity of the cosmopolitan chaetognath Pseudosagitta maxima (Conant, 1896) in the Atlantic Ocean and its relationship with the congeneric species. ICES Journal of Marine Science, 74(7):1875–1884, September 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/7/1875/2962395.

Kang:2011:MAG

[KNH11] Myounghee Kang, Takeshi Nakamura, and Akira Hamano. A methodology for acoustic and geospatial analysis of diverse artificial-reef datasets. ICES Journal of Marine Science, 68(10):2210–2221, November 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/68/
10/2210/610677.

Kubilius:2012:TST

[KO12] Rokas Kubilius and Egil Ona. Target strength and tiltangle distribution of the lesser sandeel (Ammodytes marinus). ICES Journal of Marine Science, 69(6):1099–1107, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/69/6/1099/622527.

Kubilius:2015:MSK

[KOC15] Rokas Kubilius, Egil Ona, and Lucio Calise. Measuring in situ krill tilt orientation by stereo photogrammetry: examples for Euphausia superba and Meganyctiphanes norvegica. ICES Journal of Marine Science, 72(8):2494–2505, September 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/8/2494/2457916.

Kavanaugh:2016:SNV

[KOC^{+}16] Maria T. Kavanaugh, Matthew J. Oliver, Francisco P. Chavez, Ricardo M. Letelier, Frank E. Muller-Karger, and Scott C. Doney. Seascapes as a new vernacular for pelagic ocean monitoring, management and conservation. ICES Journal of Marine Science, 73(7):1839–1850, July 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/7/1839/2458882.

Koeller:2011:SFP

[Koe11] Peter Koeller. Satellites and fisheries: a personal view. ICES Journal of Marine Science, 68(4):642–643, March 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/4/642/650726.

Kjesbu:2014:MUJ

[KOK^{+}14] Olav Sigurd Kjesbu, Anders Frugård Opdal, Knut Korsbrekke, Jennifer A. Devine, and Jon Egil Skjæraasen. Making use of Johan Hjort’s “unknown” legacy: reconstruction of a 150-year coastal time-series on northeast Arctic cod (Gadus morhua) liver data reveals longterm trends in energy allocation patterns. ICES Journal of Marine Science, 71(8):2053–2063, October 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/8/2053/754494.

Korneliussen:2010:AIA

[Kor10] Rolf J. Korneliussen. The acoustic identification of Atlantic mackerel. ICES Journal of Marine Science, 67(8):1749– 1758, November 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/67/8/1749/603953.

Klais:2017:WSC

[KOT^{+}17] Riina Klais, Saskia A. Otto, Marilyn Teder, Mart Simm, and Henn Ojaveer. Winter–spring climate effects on smallsized copepods in the coastal Baltic Sea. ICES Journal of Marine Science, 74(7):1855–1864, September 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/7/1855/3091774.

Kim:2013:PPW

[KP13] Tae-Goun Kim and Daniel R. Petrolia. Public perceptions of wetland restoration benefits in Louisiana. ICES Journal of Marine Science, 70(5):1045–1054, September 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/5/1045/644785.

Kough:2015:MSC

[KPBB15] Andrew S. Kough, Claire B. Paris, Donald C. Behringer, and Mark J. Butler IV. Modelling the spread and connectivity of waterborne marine pathogens: the case of PaV1 in the Caribbean. ICES Journal of Marine Science, 72(S1): S139–S146, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/72/suppl_1/i139/614947.

Kram:2016:VRT

[KPD^{+}16] S. L. Kram, N. N. Price, E. M. Donham, M. D. Johnson, E. L. A. Kelly, S. L. Hamilton, and J. E. Smith. Variable responses of temperate calcified and fleshy macroalgae to elevated pCO_{2} and warming. ICES Journal of Marine Science, 73(3):693–703, February 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/3/693/2458745.

Kelly:2016:HPA

[KPGH16] Morgan W. Kelly, Jacqueline L. Padilla-Gamiño, and Gretchen E. Hofmann. High pCO_{2} affects body size, but not gene expression in larvae of the California mussel (Mytilus californianus). ICES Journal of Marine Science, 73(3):962–969, February 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
3/962/2458759.

Knights:2015:EEA

[KPJ^{+}15] Antony M. Knights, Gerjan J. Piet, Ruud H. Jongbloed, Jacqueline E. Tamis, Lydia White, Ekin Akoglu, Laura Boicenco, Tanya Churilova, Olga Kryvenko, Vivi Fleming-Lehtinen, Juha-Markku Leppanen, Bella S. Galil, Freya Goodsir, Menachem Goren, Piotr Margonski, Snejana Moncheva, Temel Oguz, K. Nadia Papadopoulou, Outi Setälä, Chris J. Smith, Kremena Stefanova, Florin Timofte, and Leonie A. Robinson. An exposure-effect approach for evaluating ecosystem-wide risks from human activities. ICES Journal of Marine Science, 72(3):1105–1115, March 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/3/1105/703182.

Kang:2014:EAT

[KPJC14] Donhyug Kang, Jusam Park, Seom-Kyu Jung, and Sungho Cho. Estimates of acoustic target strength for giant jellyfish Nemopilema nomurai Kishinouye in the coastal Northwest Pacific. ICES Journal of Marine Science, 71(3):597–603, March 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/71/3/597/639548.

Karp:2019:ASD

[KPL^{+}19] Melissa A. Karp, Jay O. Peterson, Patrick D. Lynch, Roger B. Griffis, Charles F. Adams, William S. Arnold, Lewis A. K. Barnett, Yvonne deReynier, Jane DiCosimo, Kari H. Fenske, Sarah K. Gaichas, Anne Hollowed, Kirstin Holsman, Mandy Karnauskas, Donald Kobayashi, Andrew Leising, John P. Manderson, Michelle McClure, Wendy E. Morrison, Erin Schnettler, Andrew Thompson, James T. Thorson, John F. Walter, Annie J. Yau, Richard D.

Methot, Jr., and Jason S. Link. Accounting for shifting distributions and changing productivity in the development of scientific advice for fishery management. ICES Journal of Marine Science, 76(5):1305–1315, 09- 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/5/1305/5474997.

Kovac:2016:RPP

[KPS^{+}16] Žarko Kovač, Trevor Platt, Shubha Sathyendranath, Mira Morović, and Thomas Jackson. Recovery of photosynthesis parameters from in situ profiles of phytoplankton production. ICES Journal of Marine Science, 73(2):275–285, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/2/275/2614440.

Krumsick:2012:ACG

[KR12] Kyle J. Krumsick and George A. Rose. Atlantic cod (Gadus morhua) feed during spawning off Newfoundland and Labrador. ICES Journal of Marine Science, 69(10):1701– 1709, December 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/69/10/1701/623807.

Kvile:2018:SMN

[KRD^{+}18] Kristina Øie Kvile, Giovanni Romagnoni, Knut-Frode Dagestad, Øystein Langangen, and Trond Kristiansen. Sensitivity of modelled North Sea cod larvae transport to vertical behaviour, ocean model resolution and interannual variation in ocean dynamics. ICES Journal of Marine Science, 75(7):2413–2424, December 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/7/2413/4975493.

Kraak:2012:CFM

[KRG^{+}12] Sarah B. M. Kraak, Dave G. Reid, Hans D. Gerritsen, Ciarán J. Kelly, Mike Fitzpatrick, Edward A. Codling, and Emer Rogan. 21st century fisheries management: a spatio-temporally explicit tariff-based approach combining multiple drivers and incentivising responsible fishing. ICES Journal of Marine Science, 69(4):590–601, May 2012.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/4/590/634880.

Kloser:2016:DSL

[KRKG16] Rudy J. Kloser, Tim E. Ryan, Gordon Keith, and Lisa Gershwin. Deep-scattering layer, gas-bladder density, and size estimates using a two-frequency acoustic and optical probe. ICES Journal of Marine Science, 73(8):2037– 2048, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/73/8/2037/2198218.

Kloser:2011:SMT

[KRML11] Rudy J. Kloser, Tim E. Ryan, Gavin J. Macaulay, and Mark E. Lewis. In situ measurements of target strength with optical and model verification: a case study for blue grenadier, Macruronus novaezelandiae. ICES Journal of Marine Science, 68(9):1986–1995, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/9/1986/670571.

Kuparinen:2016:SSL

[KRO^{+}16] Anna Kuparinen, Nancy E. Roney, Rebekah A. Oomen, Jeffrey A. Hutchings, and Esben M. Olsen. Smallscale life history variability suggests potential for spatial mismatches in Atlantic cod management units. ICES Journal of Marine Science, 73(2):286–292, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/2/286/2614303.

Kidd:2014:SPR

[KS14] Sue Kidd and Dave Shaw. The social and political realities of marine spatial planning: some land-based reflections. ICES Journal of Marine Science, 71(7):1535–1541, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/71/7/1535/669262.

Kienzle:2017:RTI

[KS17] Marco Kienzle and David J. Sterling. Rising temperatures increased recruitment of brown tiger prawn (Penaeus esculentus) in Moreton Bay (Australia). ICES Journal of Marine Science, 74(3):741–749, March 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/3/741/2631170.

Kerstetter:2018:BCW

[KS18a] David W. Kerstetter and Jason Schratwieser. Billfishes in a changing world. ICES Journal of Marine Science, 75(2): 840–843, March 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/2/840/4919582.

Klibansky:2018:FPP

[KS18b] Nikolai Klibansky and Frederick S. Scharf. Fecundity peaks prior to sex transition in a protogynous marine batch spawning fish, black sea bass (Centropristis striata). ICES Journal of Marine Science, 75(3):1042–1053, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/3/1042/4739745.

Kajajian:2014:LEE

[KSJ14] Antranik Kajajian, Jason J. Schaffler, and Cynthia M. Jones. Lack of equivalence in the elemental and stable isotope chemistry within the sagittal otolith pair of the summer flounder, Paralichthys dentatus. ICES Journal of Marine Science, 71(2):356–364, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/2/356/785793.

Koster:2014:GFE

[KST14] Friedrich W. Köster, Robert L. Stephenson, and Edward A. Trippel. Gadoid fisheries: the ecology and management of rebuilding. ICES Journal of Marine Science, 71(6):1311– 1316, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/71/6/1311/2835604.

Katugin:2010:MBG

[KSZ10] Oleg N. Katugin, Gennadyi A. Shevtsov, and Mikhail A. Zuev. The morphology and biology of Gonatus tinro and Gonatopsis okutanii (Teuthida: Gonatidae) indicate that they are conspecific. ICES Journal of Marine Science, 67(7):1464–1477, October 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/67/
7/1464/664844.

Kaartvedt:2018:PFF

[KT18] Stein Kaartvedt and Josefin Titelman. Planktivorous fish in a future Arctic Ocean of changing ice and unchanged photoperiod. ICES Journal of Marine Science, 75(7):2312– 2318, December 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/7/2312/4816123.

Kim:2016:CDD

[KTLB16] Tae Won Kim, Josi Taylor, Chris Lovera, and James P. Barry. CO_{2}-driven decrease in pH disrupts olfactory behaviour and increases individual variation in deep-sea hermit crabs. ICES Journal of Marine Science, 73(3):613–619, February 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/3/613/2457893.

Koutsidi:2016:FFU

[KTMV16] Martha Koutsidi, Evangelos Tzanatos, Athanassios Machias, and Vassiliki Vassilopoulou. Fishing for function: the use of biological traits to evaluate the effects of multispecies fisheries on the functioning of fisheries assemblages. ICES Journal of Marine Science, 73(4):1091–1103, March 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/4/1091/2459090.

Keinanen:2012:TDS

[KUM^{+}12] Marja Keinänen, Annika Uddström, Jaakko Mikkonen, Michele Casini, Jukka Pönni, Timo Myllylä, Eero Aro, and Pekka J. Vuorinen. The thiamine deficiency syndrome M74, a reproductive disorder of Atlantic salmon (Salmo salar) feeding in the Baltic Sea, is related to the fat and thiamine content of prey fish. ICES Journal of Marine Science, 69 (4):516–528, May 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/69/4/516/635546.

Kiorboe:2018:TBA

[KVA18] Thomas Kiørboe, André Visser, and Ken H. Andersen. A trait-based approach to ocean ecology. ICES Journal of Marine Science, 75(6):1849–1863, November 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/6/1849/5056134.

Kraak:2019:PDM

[KVFK19] S. B. M. Kraak, A. Velasco, U. Fröse, and U. Krumme. Prediction of delayed mortality using vitality scores and reflexes, as well as catch, processing, and post-release conditions: evidence from discarded flatfish in the Western Baltic trawl fishery. ICES Journal of Marine Science, 76(1):330– 341, January 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/76/1/330/5106726.

Kristiansen:2014:TEF

[KVSV14] T. Kristiansen, K. W. Vollset, S. Sundby, and F. Vikebø. Turbulence enhances feeding of larval cod at low prey densities. ICES Journal of Marine Science, 71(9):2515–2529, November 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/71/9/2515/593402.

Kirkley:2011:GEM

[KWF11] James E. Kirkley, John Walden, and Rolf Färe. A general equilibrium model for Atlantic herring (Clupea harengus) with ecosystem considerations. ICES Journal of Marine Science, 68(5):860–866, May 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/68/5/860/649292.

Kim:2012:IAE

[KYK^{+}12] Dongseon Kim, Eun Jin Yang, Kyung Hee Kim, Chang-Woong Shin, Jisoo Park, Sinjae Yoo, and Jung-Ho Hyun. Impact of an anticyclonic eddy on the summer nutrient and chlorophyll a distributions in the Ulleung Basin, East Sea (Japan Sea). ICES Journal of Marine Science, 69(1):23–29, January 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/69/1/23/672022.

Kirkman:2013:ICD

[KYK^{+}13] Stephen P. Kirkman, Dawit Yemane, John Kathena, Samuel K. Mafwila, Silvi E. Nsiangango, Toufiek Samaai, Bjorn Axelsen, and Larvika Singh. Identifying and characterizing demersal fish biodiversity hotspots in the Benguela Current large marine ecosystem: relevance in the light of global changes. ICES Journal of Marine Science, 70 (5):943–954, September 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/70/5/943/645240.

Kruse:2010:RBB

[KZS10] Gordon H. Kruse, Jie Zheng, and Diana L. Stram. Recovery of the Bristol Bay stock of red king crabs under a rebuilding plan. ICES Journal of Marine Science, 67(9):1866–1874, December 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/67/9/1866/623920.

Lorance:2011:UQQ

[LAD^{+}11] Pascal Lorance, Sveinn Agnarsson, Dimitrios Damalas, Sophie des Clers, Ivone Figueiredo, Juan Gil, and Verena M. Trenkel. Using qualitative and quantitative stakeholder knowledge: examples from European deep-water fisheries. ICES Journal of Marine Science, 68(8):1815–1824, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/68/8/1815/752229.

Lowder:2017:AOA

[LAD^{+}17] Kaitlyn B. Lowder, Michael C. Allen, James M. D. Day, Dimitri D. Deheyn, and Jennifer R. A. Taylor. Assessment of ocean acidification and warming on the growth, calcification, and biophotonics of a California grass shrimp. ICES Journal of Marine Science, 74(4):1150–1158, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/4/1150/2926064.

Lambert:2019:HSB

[LAD^{+}19] Charlotte Lambert, Matthieu Authier, Mathieu Doray, Ghislain Dorémus, Jérome Spitz, and Vincent Ridoux. Hide and seek in the Bay of Biscay — a functional investigation of marine megafauna and small pelagic fish interactions. ICES Journal of Marine Science, 76(1):113–123, January 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/1/113/5133278.

Lawrence:2016:PAP

[LAG^{+}16] Joshua M. Lawrence, Eric Armstrong, Jonathan Gordon, Susan Mærsk Lusseau, and Paul G. Fernandes. Passive and active, predator and prey: using acoustics to study interactions between cetaceans and forage fish. ICES Journal of Marine Science, 73(8):2075–2084, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/8/2075/2198273.

Leone:2019:GWS

[LÁG^{+}19] Agostino Leone, Paula Álvarez, Dorleta García, Fran Saborido-Rey, and Naiara Rodriguez-Ezpeleta. Genomewide SNP based population structure in European hake reveals the need for harmonizing biological and management units. ICES Journal of Marine Science, 76(7):2260–2266, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/7/2260/5561460.

Laptikhovsky:2011:MSS

[Lap11] Vladimir Laptikhovsky. Migrations and structure of the species range in ridge-scaled rattail Macrourus carinatus (Southwest Atlantic) and their application to fisheries management. ICES Journal of Marine Science, 68(2):309–318, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/68/2/309/616069.

Lavers:2015:PST

[Lav15] Jennifer L. Lavers. Population status and threats to fleshfooted shearwaters (Puffinus carneipes) in South and Western Australia. ICES Journal of Marine Science, 72(2):316– 327, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/72/2/316/2801467.

Linton:2011:CAA

[LB11] Brian C. Linton and James R. Bence. Catch-at-age assessment in the face of time-varying selectivity. ICES Journal of Marine Science, 68(3):611–625, March 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/3/611/656857.

Legault:2013:CSR

[LB13] Christopher M. Legault and Elizabeth N. Brooks. Can stock–recruitment points determine which spawning potential ratio is the best proxy for maximum sustainable yield reference points? ICES Journal of Marine Science, 70(6):1075–1080, September 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/70/
6/1075/639211.

Link:2014:IWL

[LB14] Jason S. Link and Howard I. Browman. Integrating what? Levels of marine ecosystem-based assessment and management. ICES Journal of Marine Science, 71(5):1170–1173, July 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/71/5/1170/647527.

Link:2017:OIE

[LB17] Jason S. Link and Howard I. Browman. Operationalizing and implementing ecosystem-based management. ICES Journal of Marine Science, 74(1):379–381, January 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/1/379/2907794.

Li:2015:EAA

[LBB15] Yang Li, James R. Bence, and Travis O. Brenden. An evaluation of alternative assessment approaches for intermixing fish populations: a case study with Great Lakes lake whitefish. ICES Journal of Marine Science, 72(1):70–81, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/1/70/824076.

Lowerre-Barbieri:2019:PFI

[LBCOJ19] Susan K. Lowerre-Barbieri, Ignacio A. Catalán, Anders Frugård Opdal, and Christian Jørgensen. Preparing for the future: integrating spatial ecology into ecosystem-based management. ICES Journal of Marine Science, 76(2):467– 476, March 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/76/2/467/5299619.

Li:2018:DPF

[LBG18] Futian Li, John Beardall, and Kunshan Gao. Diatom performance in a future ocean: interactions between nitrogen limitation, temperature, and CO_{2}-induced seawater acidification. ICES Journal of Marine Science, 75(4):1451–1464, July 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/4/1451/4788789.

Lindemann:2015:PCS

[LBJ15] Christian Lindemann, Jan O. Backhaus, and Michael A. St John. Physiological constraints on Sverdrup’s criticaldepth-hypothesis: the influences of dark respiration and sinking. ICES Journal of Marine Science, 72(6):1942–1951, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/6/1942/923220.

Lough:2017:CST

[LBK17] R. Gregory Lough, Elisabeth A. Broughton, and Trond Kristiansen. Changes in spatial and temporal variability of prey affect functional connectivity of larval and juvenile cod. ICES Journal of Marine Science, 74(6):1826–1837, July 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/6/1826/3852243.

Liu:2019:MAH

[LBK^{+}19] Chang Liu, Crista Bank, Michael Kersula, Geoffrey W. Cowles, Douglas R. Zemeckis, Steven X. Cadrin, and Christopher McGuire. Movements of Atlantic halibut in the Gulf of Maine based on geolocation. ICES Journal of Marine Science, 76(7):2020–2032, December 2019.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/7/2020/5571136.

Lowerre-Barbieri:2019:OMF

[LBKTW19] Susan K. Lowerre-Barbieri, Roland Kays, James T. Thorson, and Martin Wikelski. The ocean’s movescape: fisheries management in the bio-logging decade (2018–2028). ICES Journal of Marine Science, 76(2):477–488, March 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/2/477/5307415.

Lavery:2017:ESP

[LBLJ17] Andone C. Lavery, Christopher Bassett, Gareth L. Lawson, and J. Michael Jech. Exploiting signal processing approaches for broadband echosounders. ICES Journal of Marine Science, 74(8):2262–2275, September 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/8/2262/4096452.

Landry:2019:CSU

[LBM19] Michael R. Landry, Lynnath E. Beckley, and Barbara A. Muhling. Climate sensitivities and uncertainties in food-web pathways supporting larval bluefin tuna in subtropical oligotrophic oceans. ICES Journal of Marine Science, 76(2):359–369, March 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
2/359/5245315.

Lowerre-Barbieri:2019:ARD

[LBTS^{+}19] Susan K. Lowerre-Barbieri, Michael D. Tringali, Colin P. Shea, Sarah Walters Burnsed, Joel Bickford, Michael Murphy, and Clay Porch. Assessing red drum spawning aggregations and abundance in the Eastern Gulf of Mexico: a multidisciplinary approach. ICES Journal of Marine Science, 76(2):516–529, March 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
2/516/5218999.

Li:2018:ESN

[LCG^{+}18] Bai Li, Jie Cao, Lisha Guan, Mackenzie Mazur, Yong Chen, and Richard A. Wahle. Estimating spatial non-stationary environmental effects on the distribution of species: a case study from American lobster in the Gulf of Maine. ICES Journal of Marine Science, 75(4):1473–1482, July 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/4/1473/4939311.

Lavery:2010:MAS

[LCM10] Andone C. Lavery, Dezhang Chu, and James N. Moum. Measurements of acoustic scattering from zooplankton and oceanic microstructure using a broadband echosounder. ICES Journal of Marine Science, 67(2):379–394, March 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/2/379/694151.

Lennert-Cody:2018:RPS

[LCMR^{+}18] Cleridy E. Lennert-Cody, Gala Moreno, Victor Restrepo, Marlon H. Román, and Mark N. Maunder. Recent purseseine FAD fishing strategies in the eastern Pacific Ocean: what is the appropriate number of FADs at sea? ICES Journal of Marine Science, 75(5):1748–1757, September 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/5/1748/4976455.

Lehodey:2015:OMM

[LCS^{+}15] Patrick Lehodey, Anna Conchon, Inna Senina, Réka Domokos, Beatriz Calmettes, Julien Jouanno, Olga Hernandez, and Rudy Kloser. Optimization of a micronekton model with acoustic data. ICES Journal of Marine Science, 72(5):1399–1412, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/72/
5/1399/765328.

Laurel:2017:TDG

[LCSI17] Benjamin J. Laurel, Louise A. Copeman, Mara Spencer, and Paul Iseri. Temperature-dependent growth as a function of size and age in juvenile Arctic cod (Boreogadus saida). ICES Journal of Marine Science, 74(6):1614–1621, July 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/6/1614/3737532.

Laurel:2018:CET

[LCSI18] Benjamin J. Laurel, Louise A. Copeman, Mara Spencer, and Paul Iseri. Comparative effects of temperature on rates of development and survival of eggs and yolk-sac larvae of Arctic cod (Boreogadus saida) and walleye pollock (Gadus chalcogrammus). ICES Journal of Marine Science, 75(7):2403–2412, December 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/75/
7/2403/4975494.

Lizarraga-Cubedo:2015:SLF

[LCTB^{+}15] H. A. Lizárraga-Cubedo, I. Tuck, N. Bailey, G. J. Pierce, A. F. Zuur, and D. Bova. Scottish lobster fisheries and environmental variability. ICES Journal of Marine Science, 72(S1):S211–S224, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/72/
suppl_1/i211/617404.

Landry:2017:PMT

[LD17] Michael R. Landry and Moira R. Décima. Protistan microzooplankton and the trophic position of tuna: quantifying the trophic link between micro- and mesozooplankton in marine foodwebs. ICES Journal of Marine Science, 74(7):1885–1892, September 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
7/1885/2977679.

Link:2019:CMM

[LDCR^{+}19] Jason S. Link, Mark Dickey-Collas, Murray Rudd, Richard McLaughlin, Nicol M. Macdonald, Torsten Thiele, Johanna Ferretti, Ellen Johannesen, and Margaret Rae. Clarifying mandates for marine ecosystem-based management. ICES Journal of Marine Science, 76(1):41–44, January 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/1/41/5187407.

Large:2010:STD

[LDD^{+}10] Philip A. Large, Guzman Diez, James Drewery, Martial Laurans, Graham M. Pilling, David G. Reid, Jákup Reinert, Andrew B. South, and Vladimir I. Vinnichenko. Spatial and temporal distribution of spawning aggregations of blue ling (Molva dypterygia) west and northwest of the British Isles. ICES Journal of Marine Science, 67(3):494– 501, April 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/67/3/494/733188.

Lonergan:2011:BGS

[LDT^{+}11] Mike Lonergan, Callan D. Duck, Dave Thompson, Simon Moss, and Bernie McConnell. British grey seal (Halichoerus grypus) abundance in 2008: an assessment based on aerial counts and satellite telemetry. ICES Journal of Marine Science, 68(10):2201–2209, November 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/10/2201/616735.

Litz:2014:BAF

[LEB^{+}14] Marisa N. C. Litz, Robert L. Emmett, Paul J. Bentley, Andrew M. Claiborne, and Caren Barceló. Biotic and abiotic factors influencing forage fish and pelagic nekton community in the Columbia River plume (USA) throughout the upwelling season 1999–2009. ICES Journal of Marine Science, 71(1):5–18, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/71/1/5/643398.

Lonthair:2017:ELS

[LEE17] Joshua Lonthair, Rasmus Ern, and Andrew J. Esbaugh. The early life stages of an estuarine fish, the red drum (Sciaenops ocellatus), are tolerant to high pCO_{2}. ICES Journal of Marine Science, 74(4):1042–1050, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/4/1042/2870680.

Levesque:2011:CFN

[Lev11] Juan C. Levesque. Commercial fisheries in the northwestern Gulf of Mexico: possible implications for conservation management at the Flower Garden Banks National Marine Sanctuary. ICES Journal of Marine Science, 68(10):2175– 2190, November 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/10/2175/614895.

Levy:2015:ECD

[Lév15] Marina Lévy. Exploration of the critical depth hypothesis with a simple NPZ model. ICES Journal of Marine Science, 72(6):1916–1925, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/72/
6/1916/918121.

Leaf:2014:ABP

[LF14] Robert T. Leaf and Kevin D. Friedland. Autumn bloom phenology and magnitude influence haddock recruitment on Georges Bank. ICES Journal of Marine Science, 71(8):2017–2025, October 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/71/
8/2017/2804375.

Large:2013:DTT

[LFFL13] Scott I. Large, Gavin Fay, Kevin D. Friedland, and Jason S. Link. Defining trends and thresholds in responses of ecological indicators to fishing and environmental pressures. ICES Journal of Marine Science, 70(4):755–767, July 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/4/755/727721.

LeBris:2013:EAM

[LFGW13] Arnault Le Bris, Alain Fréchet, Peter S. Galbraith, and Joseph S. Wroblewski. Evidence for alternative migratory behaviours in the northern Gulf of St Lawrence population of Atlantic cod (Gadus morhua L.). ICES Journal of Marine Science, 70(4):793–804, July 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/4/793/728194.

Li:2019:PBR

[LFH^{+}19] Futian Li, Jiale Fan, Lili Hu, John Beardall, and Juntian Xu. Physiological and biochemical responses of Thalassiosira weissflogii (diatom) to seawater acidification and alkalization. ICES Journal of Marine Science, 76(6):1850– 1859, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/76/6/1850/5366216.

LeBris:2018:MPP

[LFM^{+}18] Arnault Le Bris, Jonathan A. D. Fisher, Hannah M. Murphy, Peter S. Galbraith, Martin Castonguay, Timothy Loher, and Dominique Robert. Migration patterns and putative spawning habitats of Atlantic halibut (Hippoglossus hippoglossus) in the Gulf of St. Lawrence revealed by geolocation of pop-up satellite archival tags. ICES Journal of Marine Science, 75(1):135–147, January 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/1/135/3884462.

Ljungstrom:2019:POCa

[LFMJ19a] Gabriella Ljungström, Tessa B. Francis, Marc Mangel, and Christian Jørgensen. Parent-offspring conflict over reproductive timing: ecological dynamics far away and at other times may explain spawning variability in Pacific herring. ICES Journal of Marine Science, 76(1):354, January 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/1/354/5262261.

Ljungstrom:2019:POCb

[LFMJ19b] Gabriella Ljungström, Tessa B. Francis, Marc Mangel, and Christian Jørgensen. Parent-offspring conflict over reproductive timing: ecological dynamics far away and at other times may explain spawning variability in Pacific herring. ICES Journal of Marine Science, 76(2):559–572, March 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/2/559/5070414.

Lomeli:2018:EBE

[LGB^{+}18] Mark J. M. Lomeli, Scott D. Groth, Matthew T. O. Blume, Bent Herrmann, and W. Waldo Wakefield. Effects on the bycatch of eulachon and juvenile groundfish by altering the level of artificial illumination along an ocean shrimp trawl fishing line. ICES Journal of Marine Science, 75(6):2224– 2234, November 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/6/2224/5073200.

Lefebvre:2011:LTH

[LGBA11] Alain Lefebvre, Natacha Guiselin, Frederique Barbet, and Felipe L. Artigas. Long-term hydrological and phytoplankton monitoring (1992–2007) of three potentially eutrophic systems in the eastern English Channel and the Southern Bight of the North Sea. ICES Journal of Marine Science, 68(10):2029–2043, November 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/10/2029/613194.

Lassalle:2012:EAA

[LGL^{+}12] Géraldine Lassalle, Didier Gascuel, François Le Loc’h, Jérémy Lobry, Graham John Pierce, Vincent Ridoux, Maria Begoña Santos, Jérôme Spitz, and Nathalie Niquil. An ecosystem approach for the assessment of fisheries impacts on marine top predators: the Bay of Biscay case study. ICES Journal of Marine Science, 69(6):925–938, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/69/6/925/617346.

Lusseau:2014:NSH

[LGR^{+}14] Susan Mærsk Lusseau, Alejandro Gallego, Jens Rasmussen, Emma M. C. Hatfield, and Mike Heath. North Sea herring (Clupea harengus L.) recruitment failure may be indicative of poor feeding success. ICES Journal of Marine Science, 71(8):2026–2041, October 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/8/2026/760161.

Lin:2019:CAL

[LGR^{+}19] Yu-Jia Lin, Edwin M. Grandcourt, Lotfi Rabaoui, Rommel H. Maneja, Mohammad A. Qurban, Khaled Al-Abdulkader, and Rubén H. Roa-Ureta. Comparative analysis of life history traits and trends of abundance in coral reefs of the orange-spotted grouper (Epinephelus coioides) from two regions of the Arabian Gulf. ICES Journal of Marine Science, 76(4):987–998, July 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/4/987/5306607.

Leopold:2014:FMS

[LGRC14] Marc Léopold, Nicolas Guillemot, Delphine Rocklin, and Cheryl Chen. A framework for mapping small-scale coastal fisheries using fishers’ knowledge. ICES Journal of Marine Science, 71(7):1781–1792, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/7/1781/662621.

Leon:2015:CLP

[LGvPH15] Rafael León, Caleb Gardner, Ingrid van Putten, and Klaas Hartmann. Changes in the lease and permanent sale quota markets of a rock lobster fishery in response to stock abundance. ICES Journal of Marine Science, 72(5):1555–1564, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/5/1555/768356.

Liu:2019:EER

[LH19] Xiaozi Liu and Mikko Heino. Evaluating effort regulation in mixed fisheries: a Monte Carlo approach. ICES Journal of Marine Science, 76(7):2114–2124, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/7/2114/5549450.

Lynam:2013:SPT

[LHH^{+}13] Christopher P. Lynam, Nicholas C. Halliday, Hannes Höffle, Peter J. Wright, Cindy J. G. van Damme, Martin Edwards, and Sophie G. Pitois. Spatial patterns and trends in abundance of larval sandeels in the North Sea: 1950– 2005. ICES Journal of Marine Science, 70(3):540–553, April 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/70/3/540/916327.

Larsen:2012:AWF

[LHHK12] Karin Margretha H. Larsen, Hjálmar Hátún, Bogi Hansen, and Regin Kristiansen. Atlantic water in the Faroe area: sources and variability. ICES Journal of Marine Science, 69 (5):802–808, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/69/5/802/649515.

Laidre:2012:SPD

[LHJ12] Kristin L. Laidre and Mads Peter Heide-Jørgensen. Spring partitioning of Disko Bay, West Greenland, by Arctic and subarctic baleen whales. ICES Journal of Marine Science, 69(7):1226–1233, September 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/69/
7/1226/748262.

Lee:2015:MCD

[LHJ15] Soxi Lee, Neil D. Hartstein, and Andrew Jeffs. Modelling carbon deposition and dissolved nitrogen discharge from sea cage aquaculture of tropical spiny lobster. ICES Journal of Marine Science, 72(S1):S260–S275, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/72/
suppl_1/i260/613696.

Lind:2018:RSF

[LHK18] Ylva Lind, Tiina Huovila, and Reijo Käkelä. A retrospective study of fatty acid composition in Baltic herring (Clupea harengus membras) caught at three locations in the Baltic Sea (1973–2009). ICES Journal of Marine Science, 75(1):330–339, January 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/1/330/4102823.

Lundstrom:2010:UDC

[LHLK10] Karl Lundström, Olle Hjerne, Sven-Gunnar Lunneryd, and Olle Karlsson. Understanding the diet composition of marine mammals: grey seals (Halichoerus grypus) in the Baltic Sea. ICES Journal of Marine Science, 67(6):1230– 1239, September 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/67/6/1230/736003.

Lamb:2019:IJY

[LHP^{+}19] Philip D. Lamb, Ewan Hunter, John K. Pinnegar, Thomas K. Doyle, Simon Creer, and Martin I. Taylor. Inclusion of jellyfish in 30+ years of Ecopath with Ecosim models. ICES Journal of Marine Science, 76(7):1941–1950, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/7/1941/5584405.

Larsen:2018:NAM

[LHS^{+}18] Roger B. Larsen, Bent Herrmann, Manu Sistiaga, Jesse Brinkhof, Ivan Tatone, and Lise Langård. New approach for modelling size selectivity in shrimp trawl fisheries. ICES Journal of Marine Science, 75(1):351–360, January 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/1/351/3952607.

Lowen:2019:AEG

[LHS^{+}19] J. Benjamin Lowen, Devorah R. Hart, Ryan R. E. Stanley, Sarah J. Lehnert, Ian R. Bradbury, and Claudio DiBacco. Assessing effects of genetic, environmental, and biotic gradients in species distribution modelling. ICES Journal of Marine Science, 76(6):1762–1775, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/6/1762/5430871.

Leonardsson:2016:OTS

[LHV^{+}16] Kjell Leonardsson, Richard Hudd, Lari Veneranta, Alpo Huhmarniemi, and Erkki Jokikokko. Optimal time and sample allocation for unicohort fish larvae, sea-spawning whitefish (Coregonus lavaretus s. l.) as a case study. ICES Journal of Marine Science, 73(2):374–383, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/2/374/2614279.

Li:2016:ASB

[LJB16] Yan Li, Yan Jiao, and Joan A. Browder. Assessment of seabird bycatch in the US Atlantic pelagic longline fishery, with an extra exploration on modeling spatial variation. ICES Journal of Marine Science, 73(10):2687–2694, November 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/10/2687/2647090.

Lambert:2012:IUA

[LJH^{+}12] Gwladys I. Lambert, Simon Jennings, Jan Geert Hiddink, Niels T. Hintzen, Hilmar Hinz, Michel J. Kaiser, and Lee G. Murray. Implications of using alternative methods of Vessel Monitoring System (VMS) data analysis to describe fishing activities and impacts. ICES Journal of Marine Science, 69 (4):682–693, May 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/69/4/682/633432.

Lough:2015:PGP

[LK15] R. G. Lough and T. Kristiansen. Potential growth of pelagic juvenile cod in relation to the 1978–2006 winter– spring zooplankton on the Northeast US continental shelf. ICES Journal of Marine Science, 72(9):2549–2568, November 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/9/2549/2458718.

Lee:2017:EPF

[LK17] Jung-Ah Lee and Tae Won Kim. Effects of potential future CO_{2} levels in seawater on emerging behaviour and respiration of Manila clams, Venerupis philippinarum. ICES Journal of Marine Science, 74(4):1013–1020, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/4/1013/2669546.

Link:2019:GNI

[LKG^{+}19] Jason S. Link, Bas Kohler, Roger Griffis, Margaret M. Peg Brady, Shin-Ichi Ito, Véronique Garçon, Anne Hollowed, Manuel Barange, Robin Brown, and Wojciech Wawrzynski. A graphic novel from the 4th International Symposium on the Effects of Climate Change on the World’s Oceans. ICES Journal of Marine Science, 76(5):1221– 1243, 09- 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/5/1221/5185677.

Levontin:2011:IBE

[LKHK11] Polina Levontin, Soile Kulmala, Päivi Haapasaari, and Sakari Kuikka. Integration of biological, economic, and sociological knowledge by Bayesian belief networks: the interdisciplinary evaluation of potential management plans for Baltic salmon. ICES Journal of Marine Science, 68(3): 632–638, March 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/3/632/661664.

Laurel:2016:GMR

[LKR16] Benjamin J. Laurel, Brian A. Knoth, and Clifford H. Ryer. Growth, mortality, and recruitment signals in age-0 gadids settling in coastal Gulf of Alaska. ICES Journal of Marine Science, 73(9):2227–2237, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/9/2227/2198209.

Lassen:2014:IAF

[LKS14a] Hans Lassen, Ciaran Kelly, and Michael Sissenwine. ICES advisory framework 1977–2012: from F_{max} to precautionary approach and beyond. ICES Journal of Marine Science, 71(2):166–172, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/71/2/166/786230.

Lechner:2014:SCA

[LKS^{+}14b] Aaron Lechner, Hubert Keckeis, Elisabeth Schludermann, Franz Loisl, Paul Humphries, Martin Glas, Michael Tritthart, and Helmut Habersack. Shoreline configurations affect dispersal patterns of fish larvae in a large river. ICES Journal of Marine Science, 71(4):930–942, May 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/4/930/666079.

Levin:2014:GII

[LKS^{+}14c] Phillip S. Levin, Christopher R. Kelble, Rebecca L. Shuford, Cameron Ainsworth, Yvonne deReynier, Rikki Dunsmore, Michael J. Fogarty, Kirstin Holsman, Evan A. Howell, Mark E. Monaco, Stephanie A. Oakes, and Francisco Werner. Guidance for implementation of integrated ecosystem assessments: a US perspective. ICES Journal of Marine Science, 71(5):1198–1204, July 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/5/1198/639833.

Leaper:2010:TPA

[LLCJ10] Russell Leaper, David M. Lavigne, Peter Corkeron, and David W. Johnston. Towards a precautionary approach to managing Canada’s commercial harp seal hunt. ICES Journal of Marine Science, 67(2):316–320, March 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/2/316/691886.

Lowe:2018:DED

[LLF18] Michael R. Lowe, Gareth L. Lawson, and Michael J. Fogarty. Drivers of euphausiid distribution and abundance in the Northeast U.S. shelf large marine ecosystem. ICES Journal of Marine Science, 75(4):1280–1295, July 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/4/1280/4846901.

LeBot:2018:TSS

[LLG18] Tangi Le Bot, Amélie Lescroël, and David Grémillet. A toolkit to study seabird–fishery interactions. ICES Journal of Marine Science, 75(5):1513–1525, September 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/5/1513/4958790.

Lan:2011:OVA

[LLL^{+}11] Kuo-Wei Lan, Ming-An Lee, Hsueh-Jung Lu, Wei-Juan Shieh, Wei-Kuan Lin, and Szu-Chia Kao. Ocean variations associated with fishing conditions for yellowfin tuna (Thunnus albacares) in the equatorial Atlantic Ocean. ICES Journal of Marine Science, 68(6):1063–1071, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/6/1063/709637.

Lassalle:2013:ESF

[LLL^{+}13] Géraldine Lassalle, Jérémy Lobry, François Le Loc’h, Steven Mackinson, Francisco Sanchez, Maciej Tomasz Tomczak, and Nathalie Niquil. Ecosystem status and functioning: searching for rules of thumb using an intersite comparison of food-web models of Northeast Atlantic continental shelves. ICES Journal of Marine Science, 70(1):135–149, January 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/70/1/135/662353.

Lacoue-Labarthe:2010:APC

[LLLB^{+}10] Thomas Lacoue-Labarthe, Estelle Le Bihan, David Borg, Noussithé Koueta, and Paco Bustamante. Acid phosphatase and cathepsin activity in cuttlefish (Sepia officinalis) eggs: the effects of Ag, Cd, and Cu exposure. ICES Journal of Marine Science, 67(7):1517–1523, October 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/7/1517/662764.

LePort:2012:CCS

[LLM12] A. Le Port, S. Lavery, and J. C. Montgomery. Conservation of coastal stingrays: seasonal abundance and population structure of the short-tailed stingray Dasyatis brevicaudata at a Marine Protected Area. ICES Journal of Marine Science, 69(8):1427–1435, September 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/8/1427/703364.

Llope:2017:EAG

[Llo17] Marcos Llope. The ecosystem approach in the Gulf of Cadiz. A perspective from the southernmost European Atlantic regional sea. ICES Journal of Marine Science, 74(1): 382–390, January 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/74/1/382/2333850.

Lopez-Lopez:2012:JAF

[LLPV^{+}12] Lucía López-López, Izaskun Preciado, Begoña Villamor, Francisco Velasco, Magdalena Iglesias, Enrique Nogueira, Jose Luis Gutierrez-Zabala, and Ignacio Olaso. Is juvenile anchovy a feeding resource for the demersal community in the Bay of Biscay? On the availability of pelagic prey to demersal predators. ICES Journal of Marine Science, 69(8):1394–1402, September 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/8/1394/703162.

Lockerbie:2018:ADT

[LLSJ18] Emma M. Lockerbie, Christopher P. Lynam, Lynne J. Shannon, and Astrid Jarre. Applying a decision tree framework in support of an ecosystem approach to fisheries: IndiSeas indicators in the North Sea. ICES Journal of Marine Science, 75(3):1009–1020, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/3/1009/4718119.

Llort:2015:OID

[LLST15] Joan Llort, Marina Lévy, Jean-Baptiste Sallée, and Alessandro Tagliabue. Onset, intensification, and decline of phytoplankton blooms in the Southern Ocean. ICES Journal of Marine Science, 72(6):1971–1984, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/6/1971/924852.

Lleonart:2010:IME

[LM10] Jordi Lleonart and Gorka Merino. Immediate maximum economic yield; a realistic fisheries economic reference point. ICES Journal of Marine Science, 67(3):577–582, April 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/67/3/577/732617.

Linnane:2014:LSP

[LMG^{+}14] Adrian Linnane, Richard McGarvey, Caleb Gardner, Terence I. Walker, Janet Matthews, Bridget Green, and André E. Punt. Large-scale patterns in puerulus settlement and links to fishery recruitment in the southern rock lobster (Jasus edwardsii), across south-eastern Australia. ICES Journal of Marine Science, 71(3):528–536, March 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/3/528/638490.

Lehuta:2010:CSU

[LMPP10] Sigrid Lehuta, Stéphanie Mahévas, Pierre Petitgas, and Dominique Pelletier. Combining sensitivity and uncertainty analysis to evaluate the impact of management measures with ISIS–fish: marine protected areas for the Bay of Biscay anchovy (Engraulis encrasicolus) fishery. ICES Journal of Marine Science, 67(5):1063–1075, July 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/5/1063/609714.

Lynch:2015:POW

[LNH^{+}15] Patrick D. Lynch, Janet A. Nye, Jonathan A. Hare, Charles A. Stock, Michael A. Alexander, James D. Scott, Kiersten L. Curti, and Katherine Drew. Projected ocean warming creates a conservation challenge for river herring populations. ICES Journal of Marine Science, 72(2):374– 387, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/72/2/374/2801343.

Linders:2018:DFT

[LNWS18] Torsten Linders, Per Nilsson, Andreas Wikström, and Mattias Sköld. Distribution and fate of trawling-induced suspension of sediments in a marine protected area. ICES Journal of Marine Science, 75(2):785–795, March 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/2/785/4683659.

Lordan:2011:UDW

[LÓGR11] Colm Lordan, Macdara Ó Cuaig, Norman Graham, and Dominic Rihan. The ups and downs of working with industry to collect fishery-dependent data: the Irish experience. ICES Journal of Marine Science, 68(8):1670–1678, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/68/8/1670/759176.

Loher:2011:AMM

[Loh11] Timothy Loher. Analysis of match–mismatch between commercial fishing periods and spawning ecology of Pacific halibut (Hippoglossus stenolepis), based on winter surveys and behavioural data from electronic archival tags. ICES Journal of Marine Science, 68(10):2240–2251, November 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/10/2240/614114.

Lennox:2018:EEA

[LØM^{+}18] Robert J. Lennox, Finn Økland, Hiromichi Mitamura, Steven J. Cooke, and Eva B. Thorstad. European eel Anguilla anguilla compromise speed for safety in the early marine spawning migration. ICES Journal of Marine Science, 75(6):1984–1991, November 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/6/1984/5070416.

Lusher:2016:MIN

[LOOO16] Amy L. Lusher, Ciaran O’Donnell, Rick Officer, and Ian O’Connor. Microplastic interactions with North Atlantic mesopelagic fish. ICES Journal of Marine Science, 73(4): 1214–1225, March 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/73/4/1214/2457960.

Lorance:2011:HDO

[Lor11] Pascal Lorance. History and dynamics of the overexploitation of the blackspot sea bream (Pagellus bogaraveo) in the Bay of Biscay. ICES Journal of Marine Science, 68(2):290– 301, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/2/290/615300.

Lassalle:2014:CQQ

[LPB^{+}14] Géraldine Lassalle, Jean-Sébastien Nelva Pasqual, Philippe Boët, Marie-Joëlle Rochet, Verena M. Trenkel, and Nathalie Niquil. Combining quantitative and qualitative models to identify functional groups for monitoring changes in the Bay of Biscay continental shelf exploited foodweb. ICES Journal of Marine Science, 71(1):105–117, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/1/105/644810.

Lazartigues:2016:DNS

[LPD^{+}16a] Angélique V. Lazartigues, Stéphane Plourde, Julian J. Dodson, Olivier Morissette, Patrick Ouellet, and Pascal Sirois. Determining natal sources of capelin in a boreal marine park using otolith microchemistry. ICES Journal of Marine Science, 73(10):2644–2652, November 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/10/2644/2647102.

Little:2016:DTO

[LPD^{+}16b] L. Richard Little, André E. Punt, Catherine M. Dichmont, Natalie Dowling, David C. Smith, Elizabeth A. Fulton, Miriana Sporcic, and Rebecca J. Gorton. Decision trade-offs for cost-constrained fisheries management. ICES Journal of Marine Science, 73(2):494–502, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/2/494/2614442.

Lassen:2013:FMA

[LPFH13] Hans Lassen, Søren Anker Pedersen, Hans Frost, and Ayoe Hoff. Fishery management advice with ecosystem considerations. ICES Journal of Marine Science, 70(2):471–479, March 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/70/2/471/798649.

LeBris:2015:MEV

[LPH^{+}15] Arnault Le Bris, Andrew J. Pershing, Christina M. Hernandez, Katherine E. Mills, and Graham D. Sherwood. Modelling the effects of variation in reproductive traits on fish population resilience. ICES Journal of Marine Science, 72(9):2590–2599, November 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/9/2590/2458809.

Law:2012:BEM

[LPK12] Richard Law, Michael J. Plank, and Jeppe Kolding. On balanced exploitation of marine ecosystems: results from dynamic size spectra. ICES Journal of Marine Science, 69 (4):602–614, May 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/69/4/602/634795.

Long:2019:EOA

[LPSF19] William Christopher Long, Paige Pruisner, Katherine M. Swiney, and Robert J. Foy. Effects of ocean acidification on the respiration and feeding of juvenile red and blue king crabs (Paralithodes camtschaticus and P. platypus). ICES Journal of Marine Science, 76(5):1335–1343, 09- 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/5/1335/5510569.

Lorance:2010:SBL

[LPT10] Pascal Lorance, Lionel Pawlowski, and Verena M. Trenkel. Standardizing blue ling landings per unit effort from industry haul-by-haul data using generalized additive models. ICES Journal of Marine Science, 67(8):1650–1658, November 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/67/8/1650/603736.

Legendre:2018:AEA

[LRJ18] Louis Legendre, Richard B. Rivkin, and Nianzhi Jiao. Advanced experimental approaches to marine water-column biogeochemical processes. ICES Journal of Marine Science, 75(1):30–42, January 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/75/1/30/4080730.

Libralato:2010:CMB

[LS10] Simone Libralato and Cosimo Solidoro. Comparing methods for building trophic spectra of ecological data. ICES Journal of Marine Science, 67(3):426–434, April 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/3/426/731157.

Ljunggren:2010:RFC

[LSB^{+}10] Lars Ljunggren, Alfred Sandström, Ulf Bergström, Johanna Mattila, Antti Lappalainen, Gustav Johansson, Göran Sundblad, Michele Casini, Olavi Kaljuste, and Britas Klemens Eriksson. Recruitment failure of coastal predatory fish in the Baltic Sea coincident with an offshore ecosystem regime shift. ICES Journal of Marine Science, 67(8):1587– 1595, November 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/67/8/1587/608411.

Lynch:2018:ATH

[LSCL18] Patrick D. Lynch, Kyle W. Shertzer, Enric Cortés, and Robert J. Latour. Abundance trends of highly migratory species in the Atlantic Ocean: accounting for water temperature profiles. ICES Journal of Marine Science, 75(4): 1427–1438, July 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/4/1427/4840572.

Lam:2011:CFA

[LSD^{+}11] Vicky W. Y. Lam, Ussif Rashid Sumaila, Andrew Dyck, Daniel Pauly, and Reg Watson. Construction and first applications of a global cost of fishing database. ICES Journal of Marine Science, 68(9):1996–2004, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/9/1996/668065.

Lewandowska:2015:IPT

[LSF^{+}15] Aleksandra M. Lewandowska, Maren Striebel, Ulrike Feudel, Helmut Hillebrand, and Ulrich Sommer. The importance of phytoplankton trait variability in spring bloom formation. ICES Journal of Marine Science, 72(6):1908– 1915, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/6/1908/925000.

Long:2016:EHP

[LSF16] W. Christopher Long, Katherine M. Swiney, and Robert J. Foy. Effects of high pCO_{2} on Tanner crab reproduction and early life history, part II: carryover effects on larvae from oogenesis and embryogenesis are stronger than direct effects. ICES Journal of Marine Science, 73(3):836–848, February 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/3/836/2459103.

Lewin:2018:EPR

[LSF^{+}18] Wolf-Christian Lewin, Harry Vincent Strehlow, Keno Ferter, Kieran Hyder, Jan Niemax, Jens-Peter Herrmann, and Marc Simon Weltersbach. Estimating post-release mortality of European sea bass based on experimental angling. ICES Journal of Marine Science, 75(4):1483–1495, July 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/4/1483/4816122.

Liu:2015:DTD

[LSI^{+}15] Yang Liu, Sei-Ichi Saitoh, Yu Ihara, Satoshi Nakada, Makoto Kanamori, Xun Zhang, Katsuhisa Baba, Yoichi Ishikawa, and Toru Hirawake. Development of a threedimensional growth prediction model for the Japanese scallop in Funka Bay, Japan, using OGCM and MODIS. ICES Journal of Marine Science, 72(9):2684–2699, November 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/9/2684/2458766.

Lee:2010:DRR

[LSJ10] Janette Lee, Andy B. South, and Simon Jennings. Developing reliable, repeatable, and accessible methods to provide high-resolution estimates of fishing-effort distributions from Vessel Monitoring System (VMS) data. ICES Journal of Marine Science, 67(6):1260–1271, September 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/6/1260/733373.

Liu:2013:IAS

[LSR^{+}13] Yang Liu, Sei-Ichi Saitoh, I. Nyoman Radiarta, Tomonori Isada, Toru Hirawake, Hiroyuki Mizuta, and Hajime Yasui. Improvement of an aquaculture site-selection model for Japanese kelp (Saccharina japonica) in southern Hokkaido, Japan: an application for the impacts of climate events. ICES Journal of Marine Science, 70(7):1460–1470, November 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/70/7/1460/609876.

Lefevre:2012:ASP

[LSWD12] Marylise A. Lefèvre, Michael J. W. Stokesbury, Frederick G. Whoriskey, and Michael J. Dadswell. Atlantic salmon post-smolt migration routes in the Gulf of St. Lawrence. ICES Journal of Marine Science, 69(6):981–990, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/6/981/622195.

Langangen:2014:EMN

[LSY^{+}14] Øystein Langangen, Leif C. Stige, Natalia A. Yaragina, Frode B. Vikebø, Bjarte Bogstad, and Yvonne Gusdal. Egg mortality of northeast Arctic cod (Gadus morhua) and haddock (Melanogrammus aeglefinus). ICES Journal of Marine Science, 71(5):1129–1136, July 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/5/1129/636948.

Lee:2018:BRI

[LTGP18] Qi Lee, James T. Thorson, Vladlena V. Gertseva, and André E. Punt. The benefits and risks of incorporating climate-driven growth variation into stock assessment models, with application to splitnose rockfish (Sebastes diploproa). ICES Journal of Marine Science, 75(1):245–256, January 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/1/245/4091482.

Leandro:2014:STS

[LTQ14] Sérgio Miguel Leandro, Peter Tiselius, and Henrique Queiroga. Spatial and temporal scales of environmental forcing of Acartia populations (Copepoda: Calanoida) in the Canal de Mira (Ria de Aveiro, Portugal). ICES Journal of Marine Science, 71(3):585–596, March 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/3/585/630584.

Link:2017:KHE

[LTS^{+}17] Jason S. Link, Olivier Thébaud, David C. Smith, Anthony D. M. Smith, Jörn Schmidt, Jake Rice, Jan Jaap Poos, Cristina Pita, Doug Lipton, Marloes Kraan, Stewart Frusher, Luc Doyen, Annie Cudennec, Keith Criddle, and Denis Bailly. Keeping humans in the ecosystem. ICES Journal of Marine Science, 74(7):1947–1956, September 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/7/1947/4210357.

Lowther:2018:RBC

[LTT^{+}18] Andrew D. Lowther, Phil Trathan, Arnaud Tarroux, Christian Lydersen, and Kit M. Kovacs. The relationship between coastal weather and foraging behaviour of chinstrap penguins, Pygoscelis antarctica. ICES Journal of Marine Science, 75(6):1940–1948, November 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/6/1940/5049103.

Lauria:2011:WIE

[LVM^{+}11] Valentina Lauria, Sandrine Vaz, Corinne S. Martin, Steve Mackinson, and André Carpentier. What influences European plaice (Pleuronectes platessa) distribution in the eastern English Channel? Using habitat modelling and GIS to predict habitat utilization. ICES Journal of Marine Science, 68(7):1500–1510, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/68/
7/1500/657970.

Loots:2010:WCS

[LVPK10] Christophe Loots, Sandrine Vaz, Benjamin Planque, and Philippe Koubbi. What controls the spatial distribution of the North Sea plaice spawning population? Confronting ecological hypotheses through a model selection framework. ICES Journal of Marine Science, 67(2):244–257, March 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/2/244/693265.

Long:2017:SGM

[LVSF17] William Christopher Long, Scott B. Van Sant, Katherine M. Swiney, and Robert J. Foy. Survival, growth, and morphology of blue king crabs: effect of ocean acidification decreases with exposure time. ICES Journal of Marine Science, 74(4):1033–1041, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
4/1033/2731237.

Lindegren:2013:TSF

[LWN^{+}13] Martin Lindegren, Staffan Waldo, P. Anders Nilsson, Henrik Svedäng, and Anders Persson. Towards sustainable fisheries of the Öresund cod (Gadus morhua) through substock-specific assessment and management recommendations. ICES Journal of Marine Science, 70(6):1140–1150, September 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/70/6/1140/635387.

Lea:2018:OPM

[LWS^{+}18] James S. E. Lea, Bradley M. Wetherbee, Lara L. Sousa, Choy Aming, Neil Burnie, Nicolas E. Humphries, Nuno Queiroz, Guy M. Harvey, David W. Sims, and Mahmood S. Shivji. Ontogenetic partial migration is associated with environmental drivers and influences fisheries interactions in a marine predator. ICES Journal of Marine Science, 75(4): 1383–1392, July 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/4/1383/4788786.

Little:2011:DEC

[LWT^{+}11] L. Richard Little, Sally E. Wayte, Geoffrey N. Tuck, Anthony D. M. Smith, Neil Klaer, Malcolm Haddon, André E. Punt, Robin Thomson, Jemery Day, and Mike Fuller. Development and evaluation of a cpue-based harvest control rule for the southern and eastern scalefish and shark fishery of Australia. ICES Journal of Marine Science, 68(8):1699– 1705, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/8/1699/746502.

Link:2010:RME

[LYS^{+}10] Jason S. Link, Dawit Yemane, Lynne J. Shannon, Marta Coll, Yunne-Jai Shin, Louize Hill, and Maria de Fatima Borges. Relating marine ecosystem indicators to fishing and environmental drivers: an elucidation of contrasting responses. ICES Journal of Marine Science, 67(4):787– 795, May 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/67/4/787/678215.

Li:2018:DID

[LYW^{+}18] Zengguang Li, Zhenjiang Ye, Rong Wan, Kisei R. Tanaka, Robert Boenish, and Yong Chen. Density-independent and density-dependent factors affecting spatio-temporal dynamics of Atlantic cod (Gadus morhua) distribution in the Gulf of Maine. ICES Journal of Marine Science, 75(4): 1329–1340, July 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/4/1329/4803072.

Martinez:2013:CRC

[MA13] Stephane Martinez and Avigdor Abelson. Coral recruitment: the critical role of early post-settlement survival. ICES Journal of Marine Science, 70(7):1294–1298, November 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/70/7/1294/606894.

Maggio:2019:HSPa

[MAA^{+}19a] Teresa Maggio, Alessandro Allegra, Franco Andaloro, João Pedro Barreiros, Pietro Battaglia, Christopher M.

Butler, Angela Cuttitta, Miguel Rodrigues Jorge Fontes, Rui Freitas, Mark Gatt, F. Saadet Karakulak, David Macias, Aldo Nicosia, Hazel A. Oxenford, Samar Saber, Nuno Vasco Rodrigues, Taner Yildiz, and Mauro Sinopoli. Historical separation and present-day structure of common dolphinfish (Coryphaena hippurus) populations in the Atlantic Ocean and Mediterranean Sea. ICES Journal of Marine Science, 76(1):352, January 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/1/352/5270872.

Maggio:2019:HSPb

[MAA^{+}19b] Teresa Maggio, Alessandro Allegra, Franco Andaloro, João Pedro Barreiros, Pietro Battaglia, Christopher M. Butler, Angela Cuttitta, Miguel Rodrigues Jorge Fontes, Rui Freitas, Mark Gatt, F. Saadet Karakulak, David Macias, Aldo Nicosia, Hazel A. Oxenford, Samar Saber, Nuno Vasco Rodrigues, Taner Yildiz, and Mauro Sinopoli. Historical separation and present-day structure of common dolphinfish (Coryphaena hippurus) populations in the Atlantic Ocean and Mediterranean Sea. ICES Journal of Marine Science, 76(4):1028–1038, July 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/4/1028/5232593.

Muxagata:2012:ATP

[MAB12] Erik Muxagata, Waldemar J. A. Amaral, and Carla N. Barbosa. Acartia tonsa production in the patos lagoon estuary, Brazil. ICES Journal of Marine Science, 69(3):475– 482, May 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/69/3/475/598875.

MacLennan:2011:RTC

[Mac11] David N. MacLennan. Real-time calibration of in situ measurements of target strength. ICES Journal of Marine Science, 68(3):626–631, March 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/68/3/626/656728.

MacCall:2012:DLM

[Mac12] Alec D. MacCall. Data-limited management reference points to avoid collapse of stocks dependent on learned migration behaviour. ICES Journal of Marine Science, 69(2): 267–270, March 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/69/2/267/706994.

MacLennan:2017:RTS

[Mac17a] David N. MacLennan. Reflections on technology and science in fishery research. ICES Journal of Marine Science, 74(8):2069–2075, September 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
8/2069/3192371.

MacLeod:2017:PIM

[Mac17b] Colin D. MacLeod. Parasitic infection: a missing piece of the ocean acidification puzzle. ICES Journal of Marine Science, 74(4):929–933, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
4/929/2669572.

Magnussen:2011:FFH

[Mag11] Eydfinn Magnussen. Food and feeding habits of cod (Gadus morhua) on the Faroe Bank. ICES Journal of Marine Science, 68(9):1909–1917, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/9/1909/665389.

Mazaris:2018:GCE

[MAGK18] Antonios D. Mazaris, Vasiliki Almpanidou, Sylvaine Giakoumi, and Stelios Katsanevakis. Gaps and challenges of the European network of protected sites in the marine realm. ICES Journal of Marine Science, 75(1):190–198, January 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/1/190/4080231.

Malkki:2012:HCD

[Mäl12] Pentti Mälkki. Half a century of decadal symposia by NAFO and ICES. ICES Journal of Marine Science, 69 (5):703–705, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/69/5/703/653608.

Mariani:2017:TIA

[MALM17] Patrizio Mariani, Ken H. Andersen, Martin Lindegren, and Brian R. MacKenzie. Trophic impact of Atlantic bluefin tuna migrations in the North Sea. ICES Journal of Marine Science, 74(6):1552–1560, July 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
6/1552/3106213.

McLellan:2015:LHT

[MAM^{+}15] William A. McLellan, Logan H. Arthur, Sarah D. Mallette, Steven W. Thornton, Ryan J. McAlarney, Andrew J. Read, and D. Ann Pabst. Longline hook testing in the mouths of pelagic odontocetes. ICES Journal of Marine Science, 72(5):1706–1713, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/5/1706/756799.

Manderson:2016:SLA

[Man16] John Pilling Manderson. Seascapes are not landscapes: an analysis performed using Bernhard Riemann’s rules. ICES Journal of Marine Science, 73(7):1831–1838, July 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/7/1831/2458762. See comment [BF17] and

response [Man17a].

Manderson:2017:RBF

[Man17a] John P. Manderson. Response to Bell and Furman (2017): Seascapes are landscapes after all; Comment on Manderson (2016): Seascapes are not landscapes: an analysis performed using Bernhard Riemann’s rules: ICES Journal of Marine Science, 73:1831–1838. ICES Journal of Marine Science, 74(8):2280–2282, September 2017.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/8/2280/3872165. See [BF17, Man16].

Mangel:2017:KYO

[Man17b] Marc Mangel. Know your organism, know your data. ICES Journal of Marine Science, 74(5):1237–1248, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/5/1237/2870686.

Makhrov:2011:CCA

[MAP^{+}11] Alexander A. Makhrov, Valentina S. Artamonova, Viktor I. Popov, Aleksei Yu. Rolskiy, and Yuri I. Bakay. Comment on: Cadrin et al. (2010) “Population structure of beaked redfish, Sebastes mentella: evidence of divergence associated with different habitats. ICES Journal of Marine Science, 67: 1617–1630”. ICES Journal of Marine Science, 68(10):2013–2015, November 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/10/2013/609582. See [CBD^{+}10a, CMP^{+}11].

Meraz:2013:RBF

[MARD13] Juan Meraz, Sergio Ancona, Cristina Rodríguez, and Hugh Drummond. Reproduction of the blue-footed booby predicts commercial fish abundance in the eastern tropical Pacific. ICES Journal of Marine Science, 70(6):1263– 1272, September 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/70/6/1263/638484.

McKeown:2019:GAR

[MAS19] Niall J. McKeown, Alexander I. Arkhipkin, and Paul W. Shaw. Genetic analysis reveals historical and contemporary population dynamics in the longfin squid Doryteuthis gahi: implications for cephalopod management and conservation. ICES Journal of Marine Science, 76(4):1019– 1027, July 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/4/1019/5307425.

Mathew:2011:FDI

[Mat11] Sebastian Mathew. Fishery-dependent information and the ecosystem approach: what role can fishers and their knowledge play in developing countries? ICES Journal of Marine Science, 68(8):1805–1808, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/8/1805/758377.

Mat:2019:CDM

[Mat19] Audrey M. Mat. Chronobiology and the design of marine biology experiments. ICES Journal of Marine Science, 76(1):60–65, January 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/76/1/60/5116189.

Maynou:2014:CAW

[May14] Francesc Maynou. Coviability analysis of Western Mediterranean fisheries under MSY scenarios for 2020. ICES Journal of Marine Science, 71(7):1563–1571, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/7/1563/2804444.

Mancini:2014:RPS

[MB14] Patrícia L. Mancini and Leandro Bugoni. Resources partitioning by seabirds and their relationship with other consumers at and around a small tropical archipelago. ICES Journal of Marine Science, 71(9):2599–2607, November 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/9/2599/2798179.

Misiuk:2019:MAC

[MBA^{+}19] Benjamin Misiuk, Trevor Bell, Alec Aitken, Craig J. Brown, and Evan N. Edinger. Mapping Arctic clam abundance using multiple datasets, models, and a spatially explicit accuracy assessment. ICES Journal of Marine Science, 76(7):2349–2361, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
7/2349/5511718.

Munoz-Benavent:2018:ABT

[MBAGVG^{+}18] Pau Muñoz-Benavent, Gabriela Andreu-García, José M. Valiente-González, Vicente Atienza-Vanacloig, Vicente Puig-Pons, and Víctor Espinosa. Automatic bluefin tuna sizing using a stereoscopic vision system. ICES Journal of Marine Science, 75(1):390–401, January 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/75/
1/390/4091483.

McGrath:2011:FTA

[MBBC11] Shane P. McGrath, Matt K. Broadhurst, Paul A. Butcher, and Stuart C. Cairns. Fate of three Australian teleosts after ingesting conventional and modified stainless- and carbonsteel hooks. ICES Journal of Marine Science, 68(10):2114– 2122, November 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/10/2114/613884.

McCay:2011:HDC

[MBC11] Bonnie J. McCay, Sylvia Brandt, and Carolyn F. Creed. Human dimensions of climate change and fisheries in a coupled system: the Atlantic surfclam case. ICES Journal of Marine Science, 68(6):1354–1367, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/6/1354/707563.

Mueter:2011:EDR

[MBIH11] Franz J. Mueter, Nicholas A. Bond, James N. Ianelli, and Anne B. Hollowed. Expected declines in recruitment of walleye pollock (Theragra chalcogramma) in the eastern Bering Sea under future climate change. ICES Journal of Marine Science, 68(6):1284–1296, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/6/1284/700865.

Muhling:2017:PFH

[MBL^{+}17] Barbara A. Muhling, Richard Brill, John T. Lamkin, Mitchell A. Roffer, Sang-Ki Lee, Yanyun Liu, and Frank Muller-Karger. Projections of future habitat use by Atlantic bluefin tuna: mechanistic vs. correlative distribution models. ICES Journal of Marine Science, 74(3):698–716, March 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/3/698/2734680.

McAllister:2015:EFR

[MBLS15] Jaime D. McAllister, Adam Barnett, Jeremy M. Lyle, and Jayson M. Semmens. Examining the functional role of current area closures used for the conservation of an overexploited and highly mobile fishery species. ICES Journal of Marine Science, 72(8):2234–2244, September 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/8/2234/2458106.

Muradian:2019:FAW

[MBP19] Melissa L. Muradian, Trevor A. Branch, and André E. Punt. A framework for assessing which sampling programmes provide the best trade-off between accuracy and cost of data in stock assessments. ICES Journal of Marine Science, 76(7):2102–2113, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/7/2102/5559547.

McGonigle:2010:IOR

[MBQ10] Christopher McGonigle, Craig J. Brown, and Rory Quinn. Insonification orientation and its relevance for imagebased classification of multibeam backscatter. ICES Journal of Marine Science, 67(5):1010–1023, July 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/5/1010/610363.

McHugh:2015:RBD

[MBS^{+}15] Matthew J. McHugh, Matt K. Broadhurst, David J. Sterling, Russell B. Millar, Greg Skilleter, and Steven J. Kennelly. Relative benthic disturbances of conventional and novel otter boards. ICES Journal of Marine Science, 72(8):2450–2456, September 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/72/
8/2450/2458772.

Monnahan:2019:OLB

[MBT^{+}19] Cole C. Monnahan, Trevor A. Branch, James T. Thorson, Ian J. Stewart, and Cody S. Szuwalski. Overcoming long Bayesian run times in integrated fisheries stock assessments. ICES Journal of Marine Science, 76(6):1477–1488, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/6/1477/5475859.

Macdonald:2014:CQD

[MCAM14] P. Macdonald, I. R. Cleasby, C. H. Angus, and C. T. Marshall. The contribution of quota to the discards problem: a case study on the complexity of common megrim Lepidorhombus whiffiagonis discarding in the northern North Sea. ICES Journal of Marine Science, 71(5):1256–1265, July 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/71/5/1256/646855.

McBride:2015:DPA

[McB15] Richard S. McBride. Diagnosis of paired age agreement: a simulation of accuracy and precision effects. ICES Journal of Marine Science, 72(7):2149–2167, October 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/7/2149/2457853.

McElhany:2017:CSE

[McE17] Paul McElhany. CO_{2} sensitivity experiments are not sufficient to show an effect of ocean acidification. ICES Journal of Marine Science, 74(4):926–928, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/4/926/2669543.

Martin:2017:SHS

[MCO^{+}17] Tyson S. H. Martin, Rod M. Connolly, Andrew D. Olds, Daniela M. Ceccarelli, Douglas E. Fenner, Thomas A. Schlacher, and Maria Beger. Subsistence harvesting by a small community does not substantially compromise coral reef fish assemblages. ICES Journal of Marine Science, 74(8):2191–2200, September 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
8/2191/3091776.

Melnychuk:2017:RGE

[MCOS17] Michael C. Melnychuk, Tyler Clavelle, Brandon Owashi, and Kent Strauss. Reconstruction of global ex-vessel prices of fished species. ICES Journal of Marine Science, 74(1): 121–133, January 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/74/1/121/2670314.

McCain:2016:LTS

[MCSL16] J. Scott P. McCain, Deneen J. Cull, David C. Schneider, and Heike K. Lotze. Long-term shift in coastal fish communities before and after the collapse of Atlantic cod (Gadus morhua). ICES Journal of Marine Science, 73(5):1415– 1426, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/5/1415/2296490.

McBride:2014:KCC

[MDD^{+}14] Margaret M. McBride, Padmini Dalpadado, Kenneth F. Drinkwater, Olav Rune Godø, Alistair J. Hobday, Anne B. Hollowed, Trond Kristiansen, Eugene J. Murphy, Patrick H. Ressler, Sam Subbey, Eileen E. Hofmann, and Harald Loeng. Krill, climate, and contrasting future scenarios for Arctic and Antarctic fisheries. ICES Journal of Marine Science, 71(7):1934–1955, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/7/1934/666832.

Miethe:2010:MRE

[MDDP10] Tanja Miethe, Calvin Dytham, Ulf Dieckmann, and Jonathan W. Pitchford. Marine reserves and the evolutionary effects of fishing on size at maturation. ICES Journal of Marine Science, 67(3):412–425, April 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/3/412/730714.

Mellon-Duval:2010:GEH

[MDdPMQ10] Capucine Mellon-Duval, Hélène de Pontual, Luisa Métral, and Loic Quemener. Growth of European hake (Merluccius merluccius) in the Gulf of Lions based on conventional tagging. ICES Journal of Marine Science, 67(1):62–70, January 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/67/1/62/594549.

Mullowney:2011:SVP

[MDMC11] Darrell R. Mullowney, Earl G. Dawe, J. Frank Morado, and Richard J. Cawthorn. Sources of variability in prevalence and distribution of bitter crab disease in snow crab (Chionoecetes opilio) along the northeast coast of Newfoundland. ICES Journal of Marine Science, 68(3):463–471, March 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/3/463/658625.

Marra:2015:PBP

[MDP^{+}15] John F. Marra, Tommy D. Dickey, Albert J. Plueddemann, Robert A. Weller, Christopher S. Kinkade, and Malgorzata Stramska. Phytoplankton bloom phenomena in the North Atlantic Ocean and Arabian Sea. ICES Journal of Marine Science, 72(6):2021–2028, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/6/2021/917902.

Menendez:2012:RSS

[MDPH12] M. Clara Menéndez, M. Sofía Dutto, M. Cintia Piccolo, and Mónica S. Hoffmeyer. The role of the seasonal and semi-diurnal tidal cycle on mesozooplankton variability in a shallow mixed estuary (Bahía Blanca, Argentina). ICES Journal of Marine Science, 69(3):389–398, May 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/3/389/600219.

McQuinn:2013:AMC

[MDS13] Ian H. McQuinn, Maxime Dion, and Jean-François St. Pierre. The acoustic multifrequency classification of two sympatric euphausiid species (Meganyctiphanes norvegica and Thysanoessa raschii), with empirical and SDWBA model validation. ICES Journal of Marine Science, 70(3): 636–649, April 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/70/3/636/915699.

Miller:2015:LOR

[MDV^{+}15] Michael J. Miller, Jeff Dubosc, Elodie Vourey, Katsumi Tsukamoto, and Valerie Allain. Low occurrence rates of ubiquitously present leptocephalus larvae in the stomach contents of predatory fish. ICES Journal of Marine Science, 72(5):1359–1369, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/72/
5/1359/779275.

Mundy:2011:ECP

[ME11] Phillip R. Mundy and Danielle F. Evenson. Environmental controls of phenology of high-latitude Chinook salmon populations of the Yukon River, North America, with application to fishery management. ICES Journal of Marine Science, 68(6):1155–1164, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/6/1155/714627.

Melvin:2016:OSA

[Mel16] Gary D. Melvin. Observations of in situ Atlantic bluefin tuna (Thunnus thynnus) with 500-kHz multibeam sonar. ICES Journal of Marine Science, 73(8):1975–1986, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/8/1975/2198427.

Merrick:2018:MSS

[Mer18] Richard Merrick. Mechanisms for science to shape US living marine resource conservation policy. ICES Journal of Marine Science, 75(7):2319–2324, December 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/7/2319/4748808.

Murray:2017:CEC

[MFB17] Christopher S. Murray, Lee A. Fuiman, and Hannes Baumann. Consequences of elevated CO_{2} exposure across multiple life stages in a coastal forage fish. ICES Journal of Marine Science, 74(4):1051–1061, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/4/1051/2670315.

MacCall:2019:HMS

[MFP^{+}19] Alec D. MacCall, Tessa B. Francis, André E. Punt, Margaret C. Siple, Derek R. Armitage, Jaclyn S. Cleary, Sherri C. Dressel, R. Russ Jones, Harvey Kitka, Lynn C. Lee, Phillip S. Levin, Jim McIsaac, Daniel K. Okamoto, Melissa Poe, Steve Reifenstuhl, Jörn O. Schmidt, Andrew O. Shelton, Jennifer J. Silver, Thomas F. Thornton, Rudi Voss, and John Woodruff. A heuristic model of socially learned migration behaviour exhibits distinctive spatial and reproductive dynamics. ICES Journal of Marine Science, 76(2):598–608, March 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
2/598/5053588.

Matis:2014:CEI

[MFS^{+}14] Paloma A. Matis, Will F. Figueira, Iain M. Suthers, Joshua Humphries, Anthony Miskiewicz, Ross A. Coleman, Brendan P. Kelaher, and Matthew D. Taylor. Cyclonic entrainment? The ichthyoplankton attributes of three major water mass types generated by the separation of the East Australian Current. ICES Journal of Marine Science, 71(7):1696–1705, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/7/1696/2804450.

Miller:2016:DPS

[MFT16] Michael J. Miller, Eric Feunteun, and Katsumi Tsukamoto. Did a “perfect storm” of oceanic changes and continental anthropogenic impacts cause northern hemisphere anguillid recruitment reductions? ICES Journal of Marine Science, 73(1):43–56, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/73/1/43/2457868.

Mackas:2012:PTS

[MG12] David L. Mackas and Moira D. Galbraith. Pteropod timeseries from the NE Pacific. ICES Journal of Marine Science, 69(3):448–459, May 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/69/3/448/598626.

Meissa:2015:OMR

[MG15] B. Meissa and D. Gascuel. Overfishing of marine resources: some lessons from the assessment of demersal stocks off Mauritania. ICES Journal of Marine Science, 72(2):414– 427, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/72/2/414/2801391.

Miethe:2014:EPM

[MGBvD14] Tanja Miethe, Tomas Gröhsler, Uwe Böttcher, and Christian von Dorrien. The effects of periodic marine inflow into the Baltic Sea on the migration patterns of Western Baltic spring-spawning herring. ICES Journal of Marine Science, 71(3):519–527, March 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/71/3/519/637417.

Mork:2012:MMP

[MGH^{+}12] Kjell Arne Mork, John Gilbey, Lars Petter Hansen, Arne J. Jensen, Jan Arge Jacobsen, Marianne Holm, Jens Christian Holst, Niall Ó Maoiléidigh, Frode Vikebø, Philip McGinnity, Webjørn Melle, Katie Thomas, Eric Verspoor, and Vidar Wennevik. Modelling the migration of post-smolt Atlantic salmon (Salmo salar) in the Northeast Atlantic. ICES Journal of Marine Science, 69(9):1616–1624, November 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/69/9/1616/639600.

Maugendre:2015:EOW

[MGL^{+}15] L. Maugendre, J.-P. Gattuso, J. Louis, A. de Kluijver, S. Marro, K. Soetaert, and F. Gazeau. Effect of ocean warming and acidification on a plankton community in the NW Mediterranean Sea. ICES Journal of Marine Science, 72(6):1744–1755, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/6/1744/917143.

Massiot-Granier:2014:ESA

[MGPC^{+}14] Félix Massiot-Granier, Etienne Prévost, Gérald Chaput, Ted Potter, Gordon Smith, Jonathan White, Samu Mäntyniemi, and Etienne Rivot. Embedding stock assessment within an integrated hierarchical Bayesian life cycle modelling framework: an application to Atlantic salmon in the Northeast Atlantic. ICES Journal of Marine Science, 71(7):1653–1670, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/71/
7/1653/666032.

Morgan:2013:CDG

[MGR^{+}13] M. J. Morgan, D. Garabana, R. M. Rideout, E. Román, A. Pérez-Rodríguez, and F. Saborido-Rey. Changes in distribution of Greenland halibut in a varying environment. ICES Journal of Marine Science, 70(2):352–361, March 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/2/352/795871.

Moritz:2015:NMD

[MGS^{+}15] C. Moritz, D. Gravel, L. Savard, C. W. McKindsey, J.-C. Brêthes, and P. Archambault. No more detectable fishing effect on Northern Gulf of St Lawrence benthic invertebrates. ICES Journal of Marine Science, 72(8):2457– 2466, September 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/72/8/2457/2459097.

Martinez-Gomez:2010:GTA

[MGVH^{+}10] C. Martínez-Gómez, A. D. Vethaak, K. Hylland, T. Burgeot, A. Köhler, B. P. Lyons, J. Thain, M. J. Gubbins, and I. M. Davies. A guide to toxicity assessment and monitoring effects at lower levels of biological organization following marine oil spills in European waters. ICES Journal of Marine Science, 67(6):1105–1118, September 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/6/1105/734325.

Matala:2016:WGD

[MHE^{+}16] Andrew P. Matala, Douglas R. Hatch, Scott Everett, Michael W. Ackerman, Brett Bowersox, Matthew Campbell, and Shawn Narum. What goes up does not come down: the stock composition and demographic characteristics of upstream migrating steelhead differ from post-spawn emigrating kelts. ICES Journal of Marine Science, 73(10):2595–2605, November 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
10/2595/2647105.

Markus:2018:DDM

[MHH^{+}18] Till Markus, Helmut Hillebrand, Anna-Katharina Hornidge, Gesche Krause, and Achim Schlüter. Disciplinary diversity in marine sciences: the urgent case for an integration of research. ICES Journal of Marine Science, 75(2):502–509, March 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/2/502/4642977.

Murray:2013:EUC

[MHHK13] Lee G. Murray, Hilmar Hinz, Natalie Hold, and Michel J. Kaiser. The effectiveness of using CPUE data derived from vessel monitoring systems and fisheries logbooks to estimate scallop biomass. ICES Journal of Marine Science, 70(7):1330–1340, November 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/70/
7/1330/608721.

Maravelias:2010:SBM

[MHHT10] Christos D. Maravelias, Richard Hillary, John Haralabous, and Efthymia V. Tsitsika. Stochastic bioeconomic modelling of alternative management measures for anchovy in the Mediterranean Sea. ICES Journal of Marine Science, 67(6):1291–1300, September 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/6/1291/734919.

McQueen:2019:AVJ

[MHK19] Kate McQueen, Josef Hrabowski, and Uwe Krumme. Age validation of juvenile cod in the Western Baltic Sea. ICES Journal of Marine Science, 76(2):430–441, March 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/2/430/5211109.

McLeod:2012:PRS

[MHLW12] David J. McLeod, Alistair J. Hobday, Jeremy M. Lyle, and Dirk C. Welsford. A prey-related shift in the abundance of small pelagic fish in eastern Tasmania? ICES Journal of Marine Science, 69(6):953–960, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/6/953/620471.

McGilliard:2011:CIM

[MHM^{+}11] Carey R. McGilliard, Ray Hilborn, Alec MacCall, André E. Punt, and John C. Field. Can information from marine protected areas be used to inform control-rule-based management of small-scale, data-poor stocks? ICES Journal of Marine Science, 68(1):201–211, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/1/201/630478.

Mytilineou:2018:MGF

[MHMP^{+}18] Chryssi Mytilineou, Bent Herrmann, Danai Mantopoulou-Palouka, Antonello Sala, and Persefoni Megalofonou. Modelling gear and fishers size selection for escapees, discards, and landings: a case study in Mediterranean trawl fisheries. ICES Journal of Marine Science, 75(5):1693–1709, September 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/5/1693/4978319.

McMillan:2019:PFM

[MHSG19] M. N. McMillan, C. Huveneers, J. M. Semmens, and B. M. Gillanders. Partial female migration and coolwater migration pathways in an overfished shark. ICES Journal of Marine Science, 76(4):1083–1093, July 2019.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/4/1083/5231958.

McMillan:2017:AVC

[MIJ^{+}17] Matthew N. McMillan, Christopher Izzo, Claudia Junge, Ole Thomas Albert, Armelle Jung, and Bronwyn M. Gillanders. Analysis of vertebral chemistry to assess stock structure in a deep-sea shark, Etmopterus spinax. ICES Journal of Marine Science, 74(3):793–803, March 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/3/793/2418183.

Millar:2010:RSS

[Mil10] Russell B. Millar. Reliability of size-selectivity estimates from paired-trawl and covered-codend experiments. ICES Journal of Marine Science, 67(3):530–536, April 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/3/530/733364.

Mahe:2019:DBA

[MIM^{+}19] Kélig Mahé, Djamila Ider, Andrea Massaro, Oussama Hamed, Alba Jurado-Ruzafa, Patrícia Gonçalves, Aikaterini Anastasopoulou, Angelique Jadaud, Chryssi Mytilineou, Romain Elleboode, Zohir Ramdane, Mahmoud Bacha, Rachid Amara, Hélène de Pontual, and Bruno Ernande. Directional bilateral asymmetry in otolith morphology may affect fish stock discrimination based on otolith shape analysis. ICES Journal of Marine Science, 76(1): 232–243, January 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/76/1/232/5184891.

Myksvoll:2014:MDE

[MJAS14] Mari S. Myksvoll, Kyung-Mi Jung, Jon Albretsen, and Svein Sundby. Modelling dispersal of eggs and quantifying connectivity among Norwegian coastal cod subpopulations. ICES Journal of Marine Science, 71(4):957–969, May 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/71/4/957/663517.

Millar:2015:MAS

[MJS^{+}15a] Colin P. Millar, Ernesto Jardim, Finlay Scott, Giacomo Chato Osio, Iago Mosqueira, and Nekane Alzorriz. Model averaging to streamline the stock assessment process. ICES Journal of Marine Science, 72(1):93–98, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/1/93/822983.

Mustamaki:2015:SSS

[MJS^{+}15b] N. Mustamäki, H. Jokinen, M. Scheinin, E. Bonsdorff, and J. Mattila. Seasonal small-scale variation in distribution among depth zones in a coastal Baltic Sea fish assemblage. ICES Journal of Marine Science, 72(8):2374–2384, September 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/8/2374/2457874.

Mustamaki:2016:SSV

[MJS^{+}16] Noora Mustamäki, Henri Jokinen, Matias Scheinin, Erik Bonsdorff, and Johanna Mattila. Seasonal shifts in the vertical distribution of fish in a shallow coastal area. ICES Journal of Marine Science, 73(9):2278–2287, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/9/2278/2198206.

Michalsen:2014:LTT

[MJSB14] Kathrine Michalsen, Torild Johansen, Sam Subbey, and Alexander Beck. Linking tagging technology and molecular genetics to gain insight in the spatial dynamics of two stocks of cod in Northeast Atlantic waters. ICES Journal of Marine Science, 71(6):1417–1432, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/6/1417/2835581.

Morgan:2018:GCR

[MKAR^{+}18] M. Joanne Morgan, Mariano Koen-Alonso, Rick M. Rideout, Alejandro D. Buren, and Dawn Maddock Parsons. Growth and condition in relation to the lack of recovery of northern cod. ICES Journal of Marine Science, 75(2):

631–641, March 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/2/631/4100551.

Maufroy:2017:MIU

[MKB^{+}17] Alexandra Maufroy, David M. Kaplan, Nicolas Bez, Alicia Delgado De Molina, Hilario Murua, Laurent Floch, and Emmanuel Chassot. Massive increase in the use of drifting fish aggregating devices (dFADs) by tropical tuna purse seine fisheries in the Atlantic and Indian oceans. ICES Journal of Marine Science, 74(1):215–225, January 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/1/215/2418180.

Mikkonen:2011:RBF

[MKC^{+}11] Jaakko Mikkonen, Marja Keinänen, Michele Casini, Jukka Pönni, and Pekka J. Vuorinen. Relationships between fish stock changes in the Baltic Sea and the M74 syndrome, a reproductive disorder of Atlantic salmon (Salmo salar). ICES Journal of Marine Science, 68(10):2134–2144, November 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/10/2134/615360.

Maina:2018:DTE

[MKD^{+}18] I. Maina, S. Kavadas, D. Damalas, M. Pantazi, and S. Katsanevakis. Dynamics of trawling effort in the Aegean Sea: investigating the potential of Vessel Monitoring System (VMS) data. ICES Journal of Marine Science, 75(6):2265– 2275, November 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/6/2265/5047862.

Marcalo:2015:QIB

[MKF^{+}15] Ana Marçalo, Isidora Katara, Diana Feijó, Helder Araújo, Isabel Oliveira, Jorge Santos, Marisa Ferreira, Sílvia Monteiro, Graham J. Pierce, Alexandra Silva, and José Vingada. Quantification of interactions between the Portuguese sardine purse-seine fishery and cetaceans. ICES Journal of Marine Science, 72(8):2438–2449, September 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/8/2438/2457905.

Melli:2018:IFB

[MKHK18] Valentina Melli, Ludvig A. Krag, Bent Herrmann, and Junita D. Karlsen. Investigating fish behavioural responses to LED lights in trawls and potential applications for bycatch reduction in the Nephrops-directed fishery. ICES Journal of Marine Science, 75(5):1682–1692, September 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/5/1682/4976456.

Marshall:2019:IEI

[MKL^{+}19] Kristin N. Marshall, Laura E. Koehn, Phillip S. Levin, Timothy E. Essington, and Olaf P. Jensen. Inclusion of ecosystem information in US fish stock assessments suggests progress toward ecosystem-based fisheries management. ICES Journal of Marine Science, 76(1):1–9, January 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/1/1/5144591.

Macaulay:2013:STS

[MKR13] Gavin J. Macaulay, Rudy J. Kloser, and Tim E. Ryan. In situ target strength estimates of visually verified orange roughy. ICES Journal of Marine Science, 70(1):215–222, January 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/70/1/215/661341.

Methratta:2012:FHF

[ML12] Elizabeth T. Methratta and Jason S. Link. Feeding hotspots for four northwest Atlantic groundfish species. ICES Journal of Marine Science, 69(10):1710–1721, December 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/69/10/1710/622034.

Mollmann:2014:IEB

[MLB^{+}14] Christian Möllmann, Martin Lindegren, Thorsten Blenckner, Lena Bergström, Michele Casini, Rabea Diekmann, Juha Flinkman, Bärbel Müller-Karulis, Stefan Neuenfeldt, Jörn O. Schmidt, Maciej Tomczak, Rüdiger Voss, and Anna Gårdmark. Implementing ecosystem-based fisheries management: from single-species to integrated ecosystem assessment and advice for Baltic Sea fish stocks. ICES Journal of Marine Science, 71(5):1187–1197, July 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/5/1187/640972.

McManus:2016:RCT

[MLC16] M. Conor McManus, Priscilla Licandro, and Steve H. Coombs. Is the Russell Cycle a true cycle? Multidecadal zooplankton and climate trends in the western English Channel. ICES Journal of Marine Science, 73(2):227–238, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/2/227/2614177.

Moore:2019:MJT

[MLC^{+}19] Bradley R. Moore, Pratiwi Lestari, Scott C. Cutmore, Craig Proctor, and Robert J. G. Lester. Movement of juvenile tuna deduced from parasite data. ICES Journal of Marine Science, 76(6):1678–1689, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/6/1678/5366209.

Muhling:2011:PEC

[MLLL11] Barbara A. Muhling, Sang-Ki Lee, John T. Lamkin, and Yanyun Liu. Predicting the effects of climate change on bluefin tuna (Thunnus thynnus) spawning habitat in the Gulf of Mexico. ICES Journal of Marine Science, 68(6): 1051–1062, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/6/1051/697909.

McGarvey:2017:DRQ

[MLMJ17] Richard McGarvey, Adrian Linnane, Janet M. Matthews, and Annabel Jones. Decision rules for quota setting to support spatial management in a lobster (Jasus edwardsii) fishery. ICES Journal of Marine Science, 74(2):588–597, March 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/2/588/2439482.

Mendo:2015:EPS

[MLMS15] T. Mendo, J. M. Lyle, N. A. Moltschaniwskyj, and J. M. Semmens. Early post-settlement mortality of the scallop Pecten fumatus and the role of algal mats as a refuge from predation. ICES Journal of Marine Science, 72(8):2322– 2331, September 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/72/8/2322/2458752.

Marshak:2017:IPI

[MLS^{+}17] Anthony R. Marshak, Jason S. Link, Rebecca Shuford, Mark E. Monaco, Ellen Johannesen, Gabriella Bianchi, M. Robin Anderson, Erik Olsen, David C. Smith, Joern O. Schmidt, and Mark Dickey-Collas. International perceptions of an integrated, multi-sectoral, ecosystem approach to management. ICES Journal of Marine Science, 74(1): 414–420, January 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/74/1/414/2660825.

Marchal:2011:QAM

[MLT11] Paul Marchal, L. Richard Little, and Olivier Thébaud. Quota allocation in mixed fisheries: a bioeconomic modelling approach applied to the Channel flatfish fisheries. ICES Journal of Marine Science, 68(7):1580–1591, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/7/1580/660873.

MacNamara:2012:SRV

[MM12] Ruairí MacNamara and T. Kieran McCarthy. Size-related variation in fecundity of European eel (Anguilla anguilla). ICES Journal of Marine Science, 69(8):1333–1337, September 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/69/8/1333/703770.

McQueen:2017:SSP

[MM17] Kate McQueen and C. Tara Marshall. Shifts in spawning phenology of cod linked to rising sea temperatures. ICES Journal of Marine Science, 74(6):1561–1573, July 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/6/1561/3065349.

Marcalo:2010:FSE

[MMA^{+}10] Ana Marçalo, Tiago A. Marques, João Araújo, Pedro Pousão-Ferreira, Karim Erzini, and Yorgos Stratoudakis. Fishing simulation experiments for predicting the effects of purse-seine capture on sardine (Sardina pilchardus). ICES Journal of Marine Science, 67(2):334–344, March 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/2/334/694741.

Murillo:2011:DDW

[MMAS11] F. J. Murillo, P. Durán Muñoz, A. Altuna, and A. Serrano. Distribution of deep-water corals of the Flemish Cap, Flemish Pass, and the Grand Banks of Newfoundland (Northwest Atlantic Ocean): interaction with fishing activities. ICES Journal of Marine Science, 68(2):319–332, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/68/2/319/614827.

Malavenda:2019:IRB

[MMC19] Sergey S. Malavenda, Svetlana V. Malavenda, and Olga Chovgan. Interspecific relationships between Palmaria palmata and three Fucus species at the Murman Coast. ICES Journal of Marine Science, 76(S1):S55–S61, December 2019. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/76/Supplement_1/i55/5481884.

Minke-Martin:2015:ODE

[MMDSP15] Vanessa Minke-Martin, J. Brian Dempson, Timothy F. Sheehan, and Michael Power. Otolith-derived estimates of marine temperature use by West Greenland Atlantic salmon (Salmo salar). ICES Journal of Marine Science, 72(7):2139–2148, October 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/72/
7/2139/2457846.

Madirolas:2017:ATS

[MMG^{+}17] Adrián Madirolas, Federico A. Membiela, Juan D. Gonzalez, Ariel G. Cabreira, Matías dell’Erba, Igor S. Prario, and Silvia Blanc. Acoustic target strength (TS) of Argentine anchovy (Engraulis anchoita): the nighttime scattering layer. ICES Journal of Marine Science, 74(5):1408–1420, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/5/1408/2742059.

Mardones:2017:TDB

[MMH17] Jorge I. Mardones, Marius N. Müller, and Gustaaf M. Hallegraeff. Toxic dinoflagellate blooms of Alexandrium catenella in Chilean fjords: a resilient winner from climate change. ICES Journal of Marine Science, 74(4):988–995, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/4/988/2670311.

Machete:2011:EFB

[MMM11] Miguel Machete, Telmo Morato, and Gui Menezes. Experimental fisheries for black scabbardfish (Aphanopus carbo) in the Azores, Northeast Atlantic. ICES Journal of Marine Science, 68(2):302–308, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/2/302/616519.

MacDiarmid:2016:MPH

[MMO16] Alison MacDiarmid, Brian MacKenzie, and Henn Ojaveer. Multidisciplinary perspectives on the history of human interactions with life in the ocean. ICES Journal of Marine Science, 73(5):1382–1385, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
5/1382/2240716.

Mamoozadeh:2018:GEP

[MMRG18] Nadya R. Mamoozadeh, Jan R. McDowell, Jay R. Rooker, and John E. Graves. Genetic evaluation of population structure in white marlin (Kajikia albida): the importance of statistical power. ICES Journal of Marine Science, 75(2):

892–902, March 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/2/892/3769293.

Mullins:2018:GAR

[MMSS18] Rachel B. Mullins, Niall J. McKeown, Warwick H. H. Sauer, and Paul W. Shaw. Genomic analysis reveals multiple mismatches between biological and management units in yellowfin tuna (Thunnus albacares). ICES Journal of Marine Science, 75(6):2145–2152, November 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/6/2145/5062948.

MacNamara:2016:FSP

[MMWC16] Ruairí MacNamara, T. Kieran McCarthy, Håkan Wickström, and Patrik D. Clevestam. Fecundity of silver-phase eels (Anguilla anguilla) from different habitat types and geographic locations. ICES Journal of Marine Science, 73(1): 135–141, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/73/1/135/2457919.

Morales-Nin:2018:DTC

[MN18] Beatriz Morales-Nin. Dealing with time: a career in fish and fisheries. ICES Journal of Marine Science, 75(2):483–493, March 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/2/483/4683660.

Morales-Nin:2015:UCT

[MNG15] Beatriz Morales-Nin and Audrey J. Geffen. The use of calcified tissues as tools to support management: the view from the 5th International Otolith Symposium. ICES Journal of Marine Science, 72(7):2073–2078, October 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/7/2073/2457895.

Morales-Nin:2017:BIB

[MNGA^{+}17] Beatriz Morales-Nin, Antoni María Grau, Juan Salvador Aguilar, María del Mar Gil, and Elena Pastor. Balearic Islands boat seine fisheries: the transparent goby fishery an example of co-management. ICES Journal of Marine Science, 74(7):2053–2058, September 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/7/2053/2998999.

Marquis:2011:IPF

[MNV^{+}11] Elise Marquis, Nathalie Niquil, Alain F. Vézina, Pierre Petitgas, and Christine Dupuy. Influence of planktonic foodweb structure on a system’s capacity to support pelagic production: an inverse analysis approach. ICES Journal of Marine Science, 68(5):803–812, May 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/5/803/650916.

Mindel:2018:SBI

[MNWB18] Beth L. Mindel, Francis C. Neat, Thomas J. Webb, and Julia L. Blanchard. Size-based indicators show depthdependent change over time in the deep sea. ICES Journal of Marine Science, 75(1):113–121, January 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/1/113/3924853.

Misund:2013:LVC

[MO13] Ole Arve Misund and Erik Olsen. Lofoten–Vesterålen: for cod and cod fisheries, but not for oil? ICES Journal of Marine Science, 70(4):722–725, July 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/4/722/728918.

MacKenzie:2018:EPE

[MO18] Brian R. MacKenzie and Henn Ojaveer. Evidence from the past: exploitation as cause of commercial extinction of autumn-spawning herring in the Gulf of Riga, Baltic Sea. ICES Journal of Marine Science, 75(7):2476–2487, December 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/7/2476/4953454.

McManus:2014:WMCa

[MOG^{+}14a] M. Conor McManus, Candace A. Oviatt, Anne E. Giblin, Jane Tucker, and Jefferson T. Turner. The Western Maine Coastal Current reduces primary production rates, zooplankton abundance and benthic nutrient fluxes in Massachusetts Bay. ICES Journal of Marine Science, 71(5): 1158–1169, July 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/71/5/1158/644017.

McManus:2014:WMCb

[MOG^{+}14b] M. Conor McManus, Candace A. Oviatt, Anne E. Giblin, Jane Tucker, and Jefferson T. Turner. The Western Maine Coastal Current reduces primary production rates, zooplankton abundance and benthic nutrient fluxes in Massachusetts Bay. ICES Journal of Marine Science, 71(7): 1987, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/71/7/1987/673386.

Marriott:2014:AIL

[MONS14] Ross J. Marriott, Michael F. O’Neill, Stephen J. Newman, and Craig L. Skepper. Abundance indices for longlived tropical snappers: estimating standardized catch rates from spatially and temporally coarse logbook data. ICES Journal of Marine Science, 71(3):618–627, March 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/3/618/637672.

Moore:2014:HWA

[Moo14] Michael J. Moore. How we all kill whales. ICES Journal of Marine Science, 71(4):760–763, May 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/4/760/670509.

Moore:2019:HWC

[Moo19] Michael J. Moore. How we can all stop killing whales: a proposal to avoid whale entanglement in fishing gear. ICES Journal of Marine Science, 76(4):781–786, July 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/4/781/5288134.

Morgan:2018:UBI

[Mor18] M. Joanne Morgan. Understanding biology to improve advice for fisheries management. ICES Journal of Marine Science, 75(3):923–931, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/75/
3/923/4772850.

Mossop:2018:RBC

[Mos18] Joanna Mossop. The relationship between the continental shelf regime and a new international instrument for protecting marine biodiversity in areas beyond national jurisdiction. ICES Journal of Marine Science, 75(1):444–450, January 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/1/444/3925335.

Morote:2010:CAE

[MOVU10] Elvira Morote, María Pilar Olivar, Fernando Villate, and Ibon Uriarte. A comparison of anchovy (Engraulis encrasicolus) and sardine (Sardina pilchardus) larvae feeding in the Northwest Mediterranean: influence of prey availability and ontogeny. ICES Journal of Marine Science, 67(5):897– 908, July 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/67/5/897/609506.

Maunder:2015:CFS

[MP15] Mark N. Maunder and Kevin R. Piner. Contemporary fisheries stock assessment: many issues still remain. ICES Journal of Marine Science, 72(1):7–18, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/1/7/819352.

Metcalf:2011:IIE

[MPB11] Sarah J. Metcalf, Matthew B. Pember, and Lynda M. Bellchambers. Identifying indicators of the effects of fishing using alternative models, uncertainty, and aggregation error. ICES Journal of Marine Science, 68(7):1417–1425, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/68/7/1417/653525.

Miller:2011:DSC

[MPBH11] Eric F. Miller, Daniel J. Pondella II, D. Shane Beck, and Kevin T. Herbinson. Decadal-scale changes in southern California sciaenids under different levels of harvesting pressure. ICES Journal of Marine Science, 68(10):2123–2133, November 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/68/10/2123/617606.

McGinty:2012:TSR

[MPJ12] Niall McGinty, Anne Marie Power, and Mark P. Johnson. Trophodynamics and stability of regional scale ecosystems in the Northeast Atlantic. ICES Journal of Marine Science, 69(5):764–775, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/69/5/764/648017.

Maps:2014:MID

[MPL^{+}14] Frédéric Maps, Stéphane Plourde, Diane Lavoie, Ian Mc-Quinn, and Joël Chassé. Modelling the influence of daytime distribution on the transport of two sympatric krill species (Thysanoessa raschii and Meganyctiphanes norvegica) in the Gulf of St Lawrence, eastern Canada. ICES Journal of Marine Science, 71(2):282–292, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/2/282/776633.

Maps:2012:GAS

[MPR12] Frédéric Maps, Andrew J. Pershing, and Nicholas R. Record. A generalized approach for simulating growth and development in diverse marine copepod species. ICES Journal of Marine Science, 69(3):370–379, May 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/3/370/599661.

Mateo:2017:HMF

[MPR17] Maria Mateo, Lionel Pawlowski, and Marianne Robert. Highly mixed fisheries: fine-scale spatial patterns in retained catches of French fisheries in the Celtic Sea. ICES Journal of Marine Science, 74(1):91–101, January 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/1/91/2669551.

Malyshev:2011:DEH

[MQ11] Andrey Malyshev and Pedro A. Quijón. Disruption of essential habitat by a coastal invader: new evidence of the effects of green crabs on eelgrass beds. ICES Journal of Marine Science, 68(9):1852–1856, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/9/1852/670391.

Mesnil:2010:CHS

[MR10] Benoit Mesnil and Marie-Joëlle Rochet. A continuous hockey stick stock–recruit model for estimating MSY reference points. ICES Journal of Marine Science, 67(8):1780– 1784, November 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/67/8/1780/604132.

Meynecke:2014:FLC

[MR14a] J.-O. Meynecke and R. G. Richards. A full life cycle and spatially explicit individual-based model for the giant mud crab (Scylla serrata): a case study from a marine protected area. ICES Journal of Marine Science, 71(3):484–498, March 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/71/3/484/638867.

Mullowney:2014:RNC

[MR14b] Darrell R. J. Mullowney and George A. Rose. Is recovery of northern cod limited by poor feeding? The capelin hypothesis revisited. ICES Journal of Marine Science, 71(4): 784–793, May 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/71/4/784/668172.

Martins:2010:EYU

[MRC^{+}10] Rodrigo S. Martins, Michael J. Roberts, Nicolette Chang, Philippe Verley, Coleen L. Moloney, and Erica A. G. Vidal. Effect of yolk utilization on the specific gravity of chokka squid (Loligo reynaudii) paralarvae: implications for dispersal on the Agulhas Bank, South Africa. ICES Journal of Marine Science, 67(7):1323–1335, October 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/7/1323/664459.

Matos:2018:HFIa

[MRC^{+}18a] Diana M. Matos, Jaime A. Ramos, Joana G. Calado, Filipe R. Ceia, Jessica Hey, and Vitor H. Paiva. How fishing intensity affects the spatial and trophic ecology of two gull species breeding in sympatry. ICES Journal of Marine Science, 75(6):1949–1964, November 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/6/1949/5061527.

Matos:2018:HFIb

[MRC^{+}18b] Diana M. Matos, Jaime A. Ramos, Joana G. Calado, Filipe R. Ceia, Jessica Hey, and Vitor H. Paiva. How fishing intensity affects the spatial and trophic ecology of two gull species breeding in sympatry. ICES Journal of Marine Science, 75(6):2288, November 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/6/2288/5124333.

Mantyniemi:2012:BPF

[MRD^{+}12] Samu Mäntyniemi, Atso Romakkaniemi, Johan Dannewitz, Stefan Palm, Tapani Pakarinen, Henni Pulkkinen, Anna Gårdmark, and Olle Karlsson. Both predation and feeding opportunities may explain changes in survival of Baltic salmon post-smolts. ICES Journal of Marine Science, 69(9):1574–1579, November 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/69/
9/1574/638745.

Miethe:2019:RPL

[MRD19] Tanja Miethe, Yves Reecht, and Helen Dobby. Reference points for the length-based indicator L_{max5%} for use in the assessment of data-limited stocks. ICES Journal of Marine Science, 76(7):2125–2139, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/7/2125/5554548.

Morton:2011:SLD

[MRMK11] Alexandra Morton, Rick Routledge, Amy McConnell, and Martin Krkošek. Sea lice dispersion and salmon survival in relation to salmon farm activity in the Broughton Archipelago. ICES Journal of Marine Science, 68(1):144– 156, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/1/144/629517.

McInnes:2016:RMU

[MRP^{+}16] Julie C. McInnes, Ben Raymond, Richard A. Phillips, Simon N. Jarman, Mary-Anne Lea, and Rachael Alderman. A review of methods used to analyse albatross diets — assessing priorities across their range. ICES Journal of Marine Science, 73(9):2125–2137, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/9/2125/2199436.

Monroy:2017:SRL

[MRSG^{+}17] Pedro Monroy, Vincent Rossi, Enrico Ser-Giacomi, Cristóbal López, and Emilio Hernández-García. Sensitivity and robustness of larval connectivity diagnostics obtained from Lagrangian flow networks. ICES Journal of Marine Science, 74(6):1763–1779, July 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
6/1763/2900533.

Miller:2010:STO

[MS10] David C. M. Miller and Peter A. Shelton. “Satisficing” and trade-offs: evaluating rebuilding strategies for Greenland halibut off the east coast of Canada. ICES Journal of Marine Science, 67(9):1896–1902, December 2010.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/9/1896/619983.

Makino:2012:ACC

[MS12] Mitsutaku Makino and Yasunori Sakurai. Adaptation to climate-change effects on fisheries in the shiretoko world natural heritage area, Japan. ICES Journal of Marine Science, 69(7):1134–1140, September 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/7/1134/749126.

McConnaughey:2014:STE

[MS14] Robert A. McConnaughey and Stephen E. Syrjala. Shortterm effects of bottom trawling and a storm event on softbottom benthos in the eastern Bering Sea. ICES Journal of Marine Science, 71(9):2469–2483, November 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/9/2469/593895.

Masuda:2015:BLM

[MS15] Michele M. Masuda and Robert P. Stone. Bayesian logistic mixed-effects modelling of transect data: relating red tree coral presence to habitat characteristics. ICES Journal of Marine Science, 72(9):2674–2683, November 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/9/2674/2458879.

Mattone:2019:IBC

[MS19] Carlo Mattone and Marcus Sheaves. The intertidal benthic community of mangrove dominated estuaries: the ecological implications of a decoupled habitat. ICES Journal of Marine Science, 76(7):2329–2337, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/7/2329/5543876.

Matthias:2019:EDI

[MSA19] Bryan G. Matthias, Colette M. St. Mary, and Robert N. M. Ahrens. Evolutionary and demographic impacts of sex change rules and size-selective exploitation on sequential hermaphrodites. ICES Journal of Marine Science, 76(7):2140–2149, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
7/2140/5533097.

Myksvoll:2014:ERR

[MSAS14] Mari S. Myksvoll, Anne D. Sandvik, Lars Asplin, and Svein Sundby. Effects of river regulations on fjord dynamics and retention of coastal cod eggs. ICES Journal of Marine Science, 71(4):943–956, May 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/71/4/943/665541.

Mace:2014:ENZ

[MSC14] Pamela M. Mace, Kevin J. Sullivan, and Martin Cryer. The evolution of New Zealand’s fisheries science and management systems under ITQs. ICES Journal of Marine Science, 71(2):204–215, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/71/
2/204/789449.

McCully:2012:LMC

[MSE12] Sophy R. McCully, Finlay Scott, and Jim R. Ellis. Lengths at maturity and conversion factors for skates (Rajidae) around the British Isles, with an analysis of data in the literature. ICES Journal of Marine Science, 69(10):1812– 1822, December 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/69/10/1812/623515.

McPherson:2011:IMF

[MSK^{+}11] Lindsay R. McPherson, Aril Slotte, Cecilie Kvamme, Sonnich Meier, and C. Tara Marshall. Inconsistencies in measurement of fish condition: a comparison of four indices of fat reserves for Atlantic herring (Clupea harengus). ICES Journal of Marine Science, 68(1):52–60, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/68/1/52/629701.

Menon:2011:AOT

[MSL^{+}11] Harilal B. Menon, Nutan Sangekar, Aneesh Lotliker, Krishnaswamy Krishna Moorthy, and Ponnumani Vethamony. Aerosol optical thickness and spatial variability along coastal and offshore waters of the eastern Arabian Sea. ICES Journal of Marine Science, 68(4):745–750, March 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/4/745/650628.

Methling:2017:RIP

[MSM17] Caroline Methling, Peter V. Skov, and Niels Madsen. Reflex impairment, physiological stress, and discard mortality of European plaice Pleuronectes platessa in an otter trawl fishery. ICES Journal of Marine Science, 74(6):1660–1671, July 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/6/1660/3059823.

McGarvey:2018:SBE

[MSMW18] Richard McGarvey, Mike A. Steer, Janet M. Matthews, and Tim M. Ward. A stage-based estimator of daily egg production. ICES Journal of Marine Science, 75(5):1638– 1646, September 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/5/1638/4947882.

Macaulay:2018:CEI

[MSOF18] Gavin J. Macaulay, Ben Scoulding, Egil Ona, and Sascha M. M. Fässler. Comparisons of echo-integration performance from two multiplexed echosounders. ICES Journal of Marine Science, 75(6):2276–2285, November 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/6/2276/5092485.

Mantzouni:2010:HMT

[MSOM10] Irene Mantzouni, Helle Sørensen, Robert B. O’Hara, and Brian R. MacKenzie. Hierarchical modelling of temperature and habitat size effects on population dynamics of North Atlantic cod. ICES Journal of Marine Science, 67(5):833– 855, July 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/67/5/833/608321.

Miller:2012:RMM

[MSR^{+}12] Alicia S. Miller, Timothy F. Sheehan, Mark D. Renkawitz, Alfred L. Meister, and Timothy J. Miller. Revisiting the marine migration of US Atlantic salmon using historical carlin tag data. ICES Journal of Marine Science, 69(9):1609–1615, November 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/69/
9/1609/635939.

Morgan:2014:EFM

[MSR14] M. J. Morgan, P. A. Shelton, and R. M. Rideout. An evaluation of fishing mortality reference points under varying levels of population productivity in three Atlantic cod (Gadus morhua) stocks. ICES Journal of Marine Science, 71(6):1407–1416, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/6/1407/2835582.

Mendo:2019:ETS

[MSR^{+}19] Tania Mendo, Sophie Smout, Tommaso Russo, Lorenzo D’Andrea, and Mark James. Effect of temporal and spatial resolution on identification of fishing activities in small-scale fisheries using pots and traps. ICES Journal of Marine Science, 76(6):1601–1609, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/6/1601/5481178.

Miranda:2019:CMV

[MSS^{+}19] T. Miranda, J. A. Smith, I. M. Suthers, D. Mazumder, D. O. Cruz, H. T. Schilling, K. Searle, and A. Vergés. Convictfish on the move: variation in growth and trophic niche space along a latitudinal gradient. ICES Journal of Marine Science, 76(7):2404–2412, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/7/2404/5525332.

Murawski:2010:WCM

[MST^{+}10] Steven A. Murawski, John H. Steele, Phillip Taylor, Michael J. Fogarty, Michael P. Sissenwine, Michael Ford, and Cynthia Suchman. Why compare marine ecosystems? ICES Journal of Marine Science, 67(1):1–9, January 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/1/1/595103.

Moriarty:2018:RBA

[MST^{+}18] M. Moriarty, A. F. Sell, V. M. Trenkel, C. P. Lynam, F. Burns, E. D. Clarke, S. P. R. Greenstreet, and C. Mc-Gonigle. Resolution of biodiversity and assemblage structure in demersal fisheries surveys: the role of tow duration. ICES Journal of Marine Science, 75(5):1672–1681, September 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/5/1672/4986954.

Melbourne-Thomas:2017:IMS

[MTCF^{+}17] Jessica Melbourne-Thomas, Andrew J. Constable, Elizabeth A. Fulton, Stuart P. Corney, Rowan Trebilco, Alistair J. Hobday, Julia L. Blanchard, Fabio Boschetti, Rodrigo H. Bustamante, Roger Cropp, Jason D. Everett, Aysha Fleming, Ben Galton-Fenzi, Simon D. Goldsworthy, Andrew Lenton, Ana Lara-Lopez, Rich Little, Martin P. Marzloff, Richard Matear, Mathieu Mongin, Eva Plagányi, Roger Proctor, James S. Risbey, Barbara J. Robson, David C. Smith, Michael D. Sumner, and E. Ingrid van Putten. Integrated modelling to support decisionmaking for marine social–ecological systems in Australia. ICES Journal of Marine Science, 74(9):2298–2308, November 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/9/2298/3855111.

Mahevas:2011:HCF

[MTDP11] Stéphanie Mahévas, Verena M. Trenkel, Mathieu Doray, and Arnaud Peyronnet. Hake catchability by the French trawler fleet in the Bay of Biscay: estimating technical and biological components. ICES Journal of Marine Science, 68(1):107–118, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/68/1/107/628722.

Methot:2014:ISB

[MTLG14] R. D. Methot, Jr., G. R. Tromble, D. M. Lambert, and K. E. Greene. Implementing a science-based system for preventing overfishing and guiding sustainable fisheries in the United States. ICES Journal of Marine Science, 71(2): 183–194, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/71/2/183/781079.

Muhling:2018:RSS

[MTO^{+}18] Barbara A. Muhling, Desiree Tommasi, Seiji Ohshimo, Michael A. Alexander, and Gerard DiNardo. Regionalscale surface temperature variability allows prediction of Pacific bluefin tuna recruitment. ICES Journal of Marine Science, 75(4):1341–1352, July 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/75/
4/1341/4917676.

MacKenzie:2012:SIR

[MTP^{+}12] Kirsteen M. MacKenzie, Clive N. Trueman, Martin R. Palmer, Andy Moore, Anton T. Ibbotson, William R. C. Beaumont, and Ian C. Davidson. Stable isotopes reveal age-dependent trophic level and spatial segregation during adult marine feeding in populations of salmon. ICES Journal of Marine Science, 69(9):1637–1645, November 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/9/1637/637398.

Mortensen:2017:RDR

[MUEO17] Lars O. Mortensen, Clara Ulrich, Søren Eliasen, and Hans Jakob Olesen. Reducing discards without reducing profit: free gear choice in a Danish result-based management trial. ICES Journal of Marine Science, 74(5):1469– 1479, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/5/1469/2870679.

Murawski:2010:RDF

[Mur10] Steven A. Murawski. Rebuilding depleted fish stocks: the good, the bad, and, mostly, the ugly. ICES Journal of Marine Science, 67(9):1830–1840, December 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/9/1830/621607.

Murawski:2011:SSP

[Mur11] Steven A. Murawski. Summing up sendai: progress integrating climate change science and fisheries. ICES Journal of Marine Science, 68(6):1368–1372, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/6/1368/715162.

Mei:2019:FHE

[MUW^{+}19] Weiping Mei, Yu Umezawa, Xin Wan, Jinghan Yuan, and Chiyuki Sassa. Feeding habits estimated from weightrelated isotope variations of mesopelagic fish larvae in the Kuroshio waters of the northeastern East China Sea. ICES Journal of Marine Science, 76(3):639–648, May 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/3/639/4934174.

Madsen:2010:USD

[MV10] Niels Madsen and Daniel Valentinsson. Use of selective devices in trawls to support recovery of the Kattegat cod stock: a review of experiments and experience. ICES Journal of Marine Science, 67(9):2042–2050, December 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/9/2042/624251.

Marchal:2013:EDF

[MV13] Paul Marchal and Youen Vermard. Evaluating deepwater fisheries management strategies using a mixedfisheries and spatially explicit modelling framework. ICES Journal of Marine Science, 70(4):768–781, July 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/4/768/728424.

Mahlum:2018:WSR

[MVB^{+}18] Shad K. Mahlum, Knut W. Vollset, Bjørn T. Barlaup, Gaute Velle, and Tore Wiers. Where the salmon roam: fjord habitat use of adult Atlantic salmon. ICES Journal of Marine Science, 75(6):2163–2171, November 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/6/2163/5046033.

Mahevas:2011:IHV

[MVH^{+}11] Stéphanie Mahévas, Youen Vermard, Trevor Hutton, Ane Iriondo, Angélique Jadaud, Christos D. Maravelias, Antonio Punzón, Jacques Sacchi, Alex Tidd, Efthymia Tsitsika, Paul Marchal, Nicolas Goascoz, Serge Mortreux, and David Roos. An investigation of human vs. technology-induced variation in catchability for a selection of European fishing fleets. ICES Journal of Marine Science, 68(10):2252–2263, November 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/68/10/2252/613445.

Maravelias:2018:PMH

[MVK18] Christos D. Maravelias, Paraskevas Vasilakopoulos, and Stefanos Kalogirou. Participatory management in a high value small-scale fishery in the Mediterranean Sea. ICES Journal of Marine Science, 75(6):2097–2106, November 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/6/2097/5098419.

Modica:2014:DLF

[MVP^{+}14] Larissa Modica, Francisco Velasco, Izaskun Preciado, Maria Soto, and Simon P. R. Greenstreet. Development of the large fish indicator and associated target for a Northeast Atlantic fish community. ICES Journal of Marine Science, 71(9):2403–2415, November 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/9/2403/2798177.

Mochizuki:2019:OHS

[MW19] Takashi Mochizuki and Masahiro Watanabe. Observed and hindcasted subdecadal variability of the tropical Pacific climate. ICES Journal of Marine Science, 76(5):1271– 1279, 09- 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/5/1271/5366945.

Marriott:2011:HCF

[MWJ11] Ross J. Marriott, Brent Wise, and Jill St John. Historical changes in fishing efficiency in the west coast demersal scalefish fishery, Western Australia: implications for assessment and management. ICES Journal of Marine Science, 68(1):76–86, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/68/1/76/631662.

Mantyniemi:2015:GSS

[MWP^{+}15] Samu H. P. Mäntyniemi, Rebecca E. Whitlock, Tommi A. Perälä, Paul A. Blomstedt, Jarno P. Vanhatalo, Margarita María Rincón, Anna K. Kuparinen, Henni P. Pulkkinen, and O. Sakari Kuikka. General state-space population dynamics model for Bayesian stock assessment. ICES Journal of Marine Science, 72(8):2209–2222, September 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/8/2209/2459021.

Munday:2016:EEC

[MWP^{+}16] Philip L. Munday, Sue-Ann Watson, Darren M. Parsons, Alicia King, Neill G. Barr, Ian M. Mcleod, Bridie J. M. Allan, and Steve M. J. Pether. Effects of elevated CO_{2} on early life history development of the yellowtail kingfish, Seriola lalandi, a large pelagic fish. ICES Journal of Marine Science, 73(3):641–649, February 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/3/641/2458924.

Madhun:2017:EPA

[MWS^{+}17] Abdullah S. Madhun, Vidar Wennevik, Ove T. Skilbrei, Egil Karlsbakk, Øystein Skaala, Ingrid U. Fiksdal, Sonnich Meier, Yongkai Tang, and Kevin A. Glover. The ecological profile of Atlantic salmon escapees entering a river throughout an entire season: diverse in escape history and genetic background, but frequently virus-infected.

ICES Journal of Marine Science, 74(5):1371–1381, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/5/1371/2930196.

Matsuno:2012:BSS

[MYI12] Kohei Matsuno, Atsushi Yamaguchi, and Ichiro Imai. Biomass size spectra of mesozooplankton in the Chukchi Sea during the summers of 1991/1992 and 2007/2008: an analysis using optical plankton counter data. ICES Journal of Marine Science, 69(7):1205–1217, September 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/7/1205/752086.

Nye:2010:CTC

[NBS^{+}10] Janet A. Nye, Alida Bundy, Nancy Shackell, Kevin D. Friedland, and Jason S. Link. Coherent trends in contiguous survey time-series of major ecological and commercial fish species in the Gulf of Maine ecosystem. ICES Journal of Marine Science, 67(1):26–40, January 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/1/26/594631.

Needle:2011:EER

[NC11] Coby L. Needle and Rui Catarino. Evaluating the effect of real-time closures on cod targeting. ICES Journal of Marine Science, 68(8):1647–1655, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/8/1647/754542.

Noblezada:2012:CAA

[NC12] Mary Mar P. Noblezada and Wilfredo L. Campos. Chaetognath assemblages along the Pacific Coast and adjacent inland waters of the Philippines: relative importance of oceanographic and biological factors. ICES Journal of Marine Science, 69(3):410–420, May 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/3/410/601052.

Norrstrom:2017:NEC

[NCH17] Niclas Norrström, Michele Casini, and Noél M. A. Holmgren. Nash equilibrium can resolve conflicting maximum sustainable yields in multi-species fisheries management. ICES Journal of Marine Science, 74(1):78–90, January 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/1/78/2669564.

Needle:2015:SSA

[NDB^{+}15] C. L. Needle, R. Dinsdale, T. B. Buch, R. M. D. Catarino, J. Drewery, and N. Butler. Scottish science applications of remote electronic monitoring. ICES Journal of Marine Science, 72(4):1214–1229, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/4/1214/801692.

Nelms:2016:PMT

[NDB^{+}16] Sarah E. Nelms, Emily M. Duncan, Annette C. Broderick, Tamara S. Galloway, Matthew H. Godfrey, Mark Hamann, Penelope K. Lindeque, and Brendan J. Godley. Plastic and marine turtles: a review and call for research. ICES Journal of Marine Science, 73(2):165–181, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/2/165/2614204.

Nikolic:2015:ESD

[NDF^{+}15] Natacha Nikolic, Joël Diméet, Spyros Fifas, Michèle Salaün, David Ravard, Laurence Fauconnet, and Marie-Joëlle Rochet. Efficacy of selective devices in reducing discards in the Nephrops trawl fishery in the Bay of Biscay. ICES Journal of Marine Science, 72(6):1869–1881, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/6/1869/920977.

Nunez:2018:DLM

[NdlPIR18] Ricardo Alberto Cavieses Núñez, Miguel Ángel Ojeda Ruiz de la Peña, Alfredo Flores Irigollen, and Manuel Rodríguez Rodríguez. Deep learning models for the prediction of small-scale fisheries catches: finfish fishery in the region of ‘Bahía Magadalena-Almejas”. ICES Journal of Marine Science, 75(6):2088–2096, November 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/6/2088/5035889.

Navarro:2016:OWE

[NDM^{+}16] Jorge M. Navarro, Cristian Duarte, Patricio H. Manríquez, Marco A. Lardies, Rodrigo Torres, Karin Acuña, Cristian A. Vargas, and Nelson A. Lagos. Ocean warming and elevated carbon dioxide: multiple stressor impacts on juvenile mussels from southern Chile. ICES Journal of Marine Science, 73(3):764–771, February 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/3/764/2459099.

Nottestad:2016:FSM

[NDP^{+}16] Leif Nøttestad, Justine Diaz, Hector Penã, Henrik Søiland, Geir Huse, and Anders Fernö. Feeding strategy of mackerel in the Norwegian Sea relative to currents, temperature, and prey. ICES Journal of Marine Science, 73(4):1127–1137, March 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/4/1127/2457913.

Needle:2015:HST

[Nee15] Coby L. Needle. Honeycomb: a spatio-temporal simulation model to evaluate management strategies and assessment methods. ICES Journal of Marine Science, 72(1):151–163, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/1/151/825964.

Nelson:2019:BCC

[Nel19] Gary A. Nelson. Bias in common catch-curve methods applied to age frequency data from fish surveys. ICES Journal of Marine Science, 76(7):2090–2101, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/7/2090/5491588.

Noonan:2016:OAA

[NF16] Sam H. C. Noonan and Katharina E. Fabricius. Ocean acidification affects productivity but not the severity of thermal bleaching in some tropical corals. ICES Journal of Marine Science, 73(3):715–726, February 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/3/715/2458717.

Nogueira:2016:CTO

[NGTT16] Adriana Nogueira, Diana González-Troncoso, and Nick Tolimieri. Changes and trends in the overexploited fish assemblages of two fishing grounds of the Northwest Atlantic. ICES Journal of Marine Science, 73(2):345–358, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/2/345/2614227.

Nguyen:2019:ELE

[NHL^{+}19] Khanh Q. Nguyen, Odd-Børre Humborstad, Svein Løkkeborg, Paul D. Winger, and Shannon M. Bayse. Effect of light-emitting diodes (LEDs) on snow crab catch rates in the Barents Sea pot fishery. ICES Journal of Marine Science, 76(6):1893–1901, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
6/1893/5475857.

Nilsson:2019:FSR

[NJFH19] Jessica A. Nilsson, Craig R. Johnson, Elizabeth A. Fulton, and Marcus Haward. Fisheries sustainability relies on biological understanding, evidence-based management, and conducive industry conditions. ICES Journal of Marine Science, 76(6):1436–1452, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/6/1436/5475863.

Neumann:2014:REB

[NKSE14] Viola Neumann, Friedrich W. Köster, Matthias Schaber, and Margit Eero. Recovery in eastern Baltic cod: is increased recruitment caused by decreased predation on early life stages? ICES Journal of Marine Science, 71(6):1382– 1392, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/71/6/1382/2835584.

Nielsen:2012:DNP

[NLB^{+}12] J. Rasmus Nielsen, Gwladys Lambert, Francois Bastardie, Henrik Sparholt, and Morten Vinther. Do Norway pout (Trisopterus esmarkii) die from spawning stress? Mortality of Norway pout in relation to growth, sexual maturity, and density in the North Sea, Skagerrak, and Kattegat. ICES Journal of Marine Science, 69(2):197–207, March 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/69/2/197/703311.

Nunes:2016:TIN

[NMF^{+}16] Joana Nunes, Sophie J. McCoy, Helen S. Findlay, Frances E. Hopkins, Vassilis Kitidis, Ana M. Queirós, Lucy Rayner, and Stephen Widdicombe. Two intertidal, non-calcifying macroalgae (Palmaria palmata and Saccharina latissima) show complex and variable responses to short-term CO_{2} acidification. ICES Journal of Marine Science, 73(3):887–896, February 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
3/887/2458302.

Noack:2017:EEF

[NMM^{+}17] Thomas Noack, Niels Madsen, Bernd Mieske, Rikke P. Frandsen, Kai Wieland, and Ludvig A. Krag. Estimating escapement of fish and invertebrates in a Danish anchor seine. ICES Journal of Marine Science, 74(9):2480–2488, November 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/9/2480/3772732.

Nicolle:2017:MLD

[NMO^{+}17] Amandine Nicolle, Roderic Moitié, Julien Ogor, Franck Dumas, Aurélie Foveau, Eric Foucher, and Eric Thiébaut. Modelling larval dispersal of Pecten maximus in the English Channel: a tool for the spatial management of the stocks. ICES Journal of Marine Science, 74(6):1812–1825, July 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/6/1812/2742036.

Niwa:2016:RSJ

[NNY16] Hiro-Sato Niwa, Kazuya Nashida, and Takashi Yanagimoto. Reproductive skew in Japanese sardine inferred from DNA sequences. ICES Journal of Marine Science, 73(9):2181–2189, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
9/2181/2198459.

Niwa:2017:AID

[NNY17] Hiro-Sato Niwa, Kazuya Nashida, and Takashi Yanagimoto. Allelic inflation in depleted fish populations with low recruitment. ICES Journal of Marine Science, 74(6): 1639–1647, July 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/74/6/1639/3069156.

Neira:2015:MDS

[NPB^{+}15] Francisco J. Neira, Robert A. Perry, Christopher P. Burridge, Jeremy M. Lyle, and John P. Keane. Molecular discrimination of shelf-spawned eggs of two co-occurring Trachurus spp. (Carangidae) in southeastern Australia: a key step to future egg-based biomass estimates. ICES Journal of Marine Science, 72(2):614–624, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/2/614/2801447.

Nogueira:2015:CED

[NPGT15] Adriana Nogueira, Xabier Paz, and Diana González-Troncoso. Changes in the exploited demersal fish assemblages in the Southern Grand Banks (NAFO Divisions 3NO): 2002–2013. ICES Journal of Marine Science, 72(3): 753–770, March 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/72/3/753/691985.

Nunez-Riboni:2019:SRP

[NRTK^{+}19] Ismael Núñez-Riboni, Marc H. Taylor, Alexander Kempf, Miriam Püts, and Moritz Mathis. Spatially resolved past and projected changes of the suitable thermal habitat of North Sea cod (Gadus morhua) under climate change. ICES Journal of Marine Science, 76(7):2389–2403, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/7/2389/5538730.

Nielsen:2017:ISF

[NS17] Julie K. Nielsen and Andrew C. Seitz. Interannual site fidelity of Pacific halibut: potential utility of protected areas for management of a migratory demersal fish. ICES Journal of Marine Science, 74(8):2120–2134, September 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/8/2120/3091775.

Nikolioudakis:2019:DSD

[NSO^{+}19] N. Nikolioudakis, H. J. Skaug, A. H. Olafsdottir, T. Jansen, J. A. Jacobsen, and K. Enberg. Drivers of the summerdistribution of Northeast Atlantic mackerel (Scomber scombrus) in the Nordic Seas from 2011 to 2017; a Bayesian hierarchical modelling approach. ICES Journal of Marine Science, 76(2):530–548, March 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
2/530/5051297.

Nieland:2015:ADE

[NSS15] Julie L. Nieland, Timothy F. Sheehan, and Rory Saunders. Assessing demographic effects of dams on diadromous fish: a case study for Atlantic salmon in the Penobscot River, Maine. ICES Journal of Marine Science, 72(8):2423– 2437, September 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/72/8/2423/2458708.

Niklitschek:2014:NSP

[NST^{+}14] Edwin J. Niklitschek, David H. Secor, Pamela Toledo, Ximena Valenzuela, Luis A. Cubillos, and Alejandro Zuleta.

Nursery systems for Patagonian grenadier off Western Patagonia: large inner sea or narrow continental shelf? ICES Journal of Marine Science, 71(2):374–390, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/2/374/783631.

Narimatsu:2010:ETC

[NUO^{+}10] Yoji Narimatsu, Yuji Ueda, Takehiro Okuda, Tsutomu Hattori, Kunihiro Fujiwara, and Masaki Ito. The effect of temporal changes in life-history traits on reproductive potential in an exploited population of Pacific cod, Gadus macrocephalus. ICES Journal of Marine Science, 67(8):1659– 1666, November 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/67/8/1659/605408.

Nottestad:2016:QCA

[NUÓ^{+}16] Leif Nøttestad, Kjell R. Utne, Gudmundur J. Óskarsson, Sigurdur Th. Jónsson, Jan Arge Jacobsen, Øyvind Tangen, Valantine Anthonypillai, Sondre Aanes, Jon Helge Vølstad, Matteo Bernasconi, Høgni Debes, Leon Smith, Sveinn Sveinbjörnsson, Jens C. Holst, Teunis Jansen, and Aril Slotte. Quantifying changes in abundance, biomass, and spatial distribution of Northeast Atlantic mackerel (Scomber scombrus) in the Nordic seas from 2007 to 2014. ICES Journal of Marine Science, 73(2):359–373, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/2/359/2614456.

Nash:2012:SLN

[NWM^{+}12] Richard D. M. Nash, Peter J. Wright, Iveta Matejusova, Stefan Petev Dimitrov, Martha O’Sullivan, Julian Augley, and Hannes Höffle. Spawning location of Norway pout (Trisopterus esmarkii Nilsson) in the North Sea. ICES Journal of Marine Science, 69(8):1338–1346, September 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/8/1338/704452.

Olsen:2010:CHS

[OAM^{+}10] Erik Olsen, Sondre Aanes, Sigbjørn Mehl, Jens Christian Holst, Asgeir Aglen, and Harald Gjøsæter. Cod, haddock, saithe, herring, and capelin in the Barents Sea and adjacent waters: a review of the biological value of the area. ICES Journal of Marine Science, 67(1):87–101, January 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/1/87/596251.

Otteraa:2012:STM

[OAT^{+}12] H. Otterå, A-L. Agnalt, A. Thorsen, O. S. Kjesbu, G. Dahle, and K. Jørstad. Is spawning time of marine fish imprinted in the genes? A two-generation experiment on local Atlantic cod (Gadus morhua L.) populations from different geographical regions. ICES Journal of Marine Science, 69(10):1722–1728, December 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/10/1722/622145.

Olsson:2012:ADC

[OBG12] Jens Olsson, Lena Bergström, and Anna Gårdmark. Abiotic drivers of coastal fish community change during four decades in the Baltic Sea. ICES Journal of Marine Science, 69(6):961–970, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/69/6/961/620764.

OBrien:2018:SCA

[OBJA^{+}18] Caitlin E. O’Brien, Cécile Bellanger, Christelle Jozet-Alves, Nawel Mezrai, Anne-Sophie Darmaillacq, and Ludovic Dickel. Stressful conditions affect reproducing cuttlefish (Sepia officinalis), reducing egg output and quality. ICES Journal of Marine Science, 75(6):2060–2069, November 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/6/2060/5092481.

Ovegaard:2012:CIA

[OBL12] Mikael Ovegård, Kim Berndt, and Sven-Gunnar Lunneryd. Condition indices of Atlantic cod (Gadus morhua) biased by capturing method. ICES Journal of Marine Science, 69(10):1781–1788, December 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/69/
10/1781/622653.

Olivar:2019:TPL

[OBLP^{+}19] M. Pilar Olivar, Antonio Bode, Cristina López-Pérez, P. Alexander Hulley, and Santiago Hernández-León. Trophic position of lanternfishes (Pisces: Myctophidae) of the tropical and equatorial Atlantic estimated using stable isotopes. ICES Journal of Marine Science, 76(3):649–661, May 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/3/649/4791959.

Obura:2018:CB

[Obu18a] David Obura. Counterpoint to basurto. ICES Journal of Marine Science, 75(3):1195–1196, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/75/
3/1195/4098831.

Obura:2018:OFW

[Obu18b] David Obura. Obura’s final word. ICES Journal of Marine Science, 75(3):1201, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/75/
3/1201/4098840.

Obura:2018:BEOa

[Obu18c] David O. Obura. On being effective, and the other 90%. ICES Journal of Marine Science, 75(3):1198–1199, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/3/1198/4098827.

Obura:2018:BEOb

[Obu18d] David O. Obura. On being effective, and the other 90%. ICES Journal of Marine Science, 75(4):1512, July 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/4/1512/4591651.

Ottersen:2014:REL

[OBY^{+}14] Geir Ottersen, Bjarte Bogstad, Natalia A. Yaragina, Leif Christian Stige, Frode B. Vikebø, and Padmini Dalpadado. A review of early life history dynamics of Barents Sea cod (Gadus morhua). ICES Journal of Marine Science, 71(8):2064–2087, October 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/8/2064/755258.

ONeill:2010:UCR

[OCBJ10] Michael F. O’Neill, Alexander B. Campbell, Ian W. Brown, and Ron Johnstone. Using catch rate data for simple cost-effective quota setting in the Australian spanner crab (Ranina ranina) fishery. ICES Journal of Marine Science, 67(8):1538–1552, November 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/8/1538/607456.

Oliveira:2010:TEE

[OCG10] Manuela M. Oliveira, Ana S. Camanho, and Miguel B. Gaspar. Technical and economic efficiency analysis of the Portuguese artisanal dredge fleet. ICES Journal of Marine Science, 67(8):1811–1821, November 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/8/1811/608459.

Oliveira:2013:ICQ

[OCG13] M. M. Oliveira, A. S. Camanho, and M. B. Gaspar. The influence of catch quotas on the productivity of the Portuguese bivalve dredge fleet. ICES Journal of Marine Science, 70(7):1378–1388, November 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/7/1378/608660.

Obradovich:2014:BLN

[OCR14] Shannon G. Obradovich, Erin H. Carruthers, and George A. Rose. Bottom-up limits to Newfoundland capelin (Mallotus villosus) rebuilding: the euphausiid hypothesis. ICES Journal of Marine Science, 71(4):775–783, May 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/71/
4/775/667471.

OKeefe:2014:EET

[OCS14] Catherine E. O’Keefe, Steven X. Cadrin, and Kevin D. E. Stokesbury. Evaluating effectiveness of time/area closures, quotas/caps, and fleet communications to reduce fisheries bycatch. ICES Journal of Marine Science, 71(5):1286– 1297, July 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/71/5/1286/638196.

Oliveira:2016:EIS

[OCWG16] Manuela M. Oliveira, Ana S. Camanho, John B. Walden, and Miguel B. Gaspar. Evaluating the influence of skipper skills in the performance of Portuguese artisanal dredge vessels. ICES Journal of Marine Science, 73(10):2721–2728, November 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/10/2721/2647101.

Oates:2017:AES

[OD17] Jennifer Oates and Lyndsey A. Dodds. An approach for effective stakeholder engagement as an essential component of the ecosystem approach. ICES Journal of Marine Science, 74(1):391–397, January 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
1/391/2967555.

ODriscoll:2012:SIS

[OdJN^{+}12] Richard L. O’Driscoll, Peter de Joux, Richard Nelson, Gavin J. Macaulay, Adam J. Dunford, Peter M. Marriott, Craig Stewart, and Brian S. Miller. Species identification in seamount fish aggregations using moored underwater video. ICES Journal of Marine Science, 69(4):648–659, May 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/69/4/648/631791.

OBoyle:2016:ISM

[ODL16] Robert O’Boyle, Micah Dean, and Christopher M. Legault. The influence of seasonal migrations on fishery selectivity. ICES Journal of Marine Science, 73(7):1774–1787, July 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/7/1774/2458751.

Olson:2019:TSF

[OdS19] Julia Olson and Patricia Pinto da Silva. Taking stock of fisheries science through oral history: voices from NOAA’s Fishery Science Centers. ICES Journal of Marine Science, 76(2):370–383, March 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
2/370/5250414.

OToole:2010:CCR

[ODSC10] Amanda C. O’Toole, Andy J. Danylchuk, Cory D. Suski, and Steven J. Cooke. Consequences of catch-and-release angling on the physiological status, injury, and immediate mortality of great barracuda (Sphyraena barracuda) in The Bahamas. ICES Journal of Marine Science, 67(8):1667– 1675, November 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/67/8/1667/606285.

Oken:2016:EES

[OE16] Kiva L. Oken and Timothy E. Essington. Evaluating the effect of a selective piscivore fishery on rockfish recovery within marine protected areas. ICES Journal of Marine Science, 73(9):2267–2277, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/9/2267/2198693.

Olafsson:2016:OAS

[OEG^{+}16] Kristinn Olafsson, Sigurdur M. Einarsson, John Gilbey, Christophe Pampoulie, Gudmundur O. Hreggvidsson, Sigridur Hjorleifsdottir, and Sigurdur Gudjonsson. Origin of Atlantic salmon (Salmo salar) at sea in Icelandic waters. ICES Journal of Marine Science, 73(6):1525–1532, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/6/1525/2458705.

Orell:2018:SSM

[OEK^{+}18] Panu Orell, Jaakko Erkinaro, Mikko Kiljunen, Jyrki Torniainen, Tapio Sutela, Mikko Jaukkuri, and Aki Mäki-Petäys. Short sea migration and precocious maturation in reared Atlantic salmon post-smolts in the northern Baltic Sea. ICES Journal of Marine Science, 75(3):1063–1070, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/3/1063/4657157.

Orio:2017:MIA

[OFB^{+}17] Alessandro Orio, Ann-Britt Florin, Ulf Bergström, Ivo Šics, Tatjana Baranova, and Michele Casini. Modelling indices of abundance and size-based indicators of cod and flounder stocks in the Baltic Sea using newly standardized trawl survey data. ICES Journal of Marine Science, 74(5):1322– 1333, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/5/1322/2972205.

Oigaard:2012:MAG

[ØFNH12] Tor Arne Øigård, Anne Kirstine Frie, Kjell Tormod Nilssen, and Mike Osborne Hammill. Modelling the abundance of grey seals (Halichoerus grypus) along the Norwegian coast. ICES Journal of Marine Science, 69(8):1436–1447, September 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/69/8/1436/701575.

Ourens:2014:IHP

[OFVF14] Rosana Ouréns, Juan Freire, Jose A. Vilar, and Luis Fernández. Influence of habitat and population density on recruitment and spatial dynamics of the sea urchin Paracentrotus lividus: implications for harvest refugia. ICES Journal of Marine Science, 71(5):1064–1072, July 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/5/1064/644671.

Ouellet:2011:OSC

[OFYS^{+}11] Patrick Ouellet, César Fuentes-Yaco, Louise Savard, Trevor Platt, Shubha Sathyendranath, Peter Koeller, David Orr, and Helle Siegstad. Ocean surface characteristics influence recruitment variability of populations of northern shrimp (Pandalus borealis) in the Northwest Atlantic. ICES Journal of Marine Science, 68(4):737–744, March 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/4/737/646723.

Olsson:2016:TRA

[OG16] Karin H. Olsson and Henrik Gislason. Testing reproductive allometry in fish. ICES Journal of Marine Science, 73(6): 1466–1473, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/73/6/1466/2458914.

Ozerov:2016:CGM

[OHD^{+}16] Mikhail Yu. Ozerov, Mikael Himberg, Paul V. Debes, Henry Hägerstrand, and Anti Vasemägi. Combining genetic markers with an adaptive meristic trait improves performance of mixed-stock analysis in Baltic whitefish. ICES Journal of Marine Science, 73(10):2529–2538, November 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/10/2529/2647117.

Olsen:2019:PMLa

[OHLK19a] Esben Moland Olsen, Kim Tallaksen Halvorsen, Torkel Larsen, and Anna Kuparinen. Potential for managing life history diversity in a commercially exploited intermediate predator, the goldsinny wrasse (Ctenolabrus rupestris). ICES Journal of Marine Science, 76(1):357, January 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/1/357/5258052.

Olsen:2019:PMLb

[OHLK19b] Esben Moland Olsen, Kim Tallaksen Halvorsen, Torkel Larsen, and Anna Kuparinen. Potential for managing life history diversity in a commercially exploited intermediate predator, the goldsinny wrasse (Ctenolabrus rupestris).

ICES Journal of Marine Science, 76(2):410–417, March 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/2/410/5224525.

Ohman:2019:STO

[Ohm19] Mark D. Ohman. A sea of tentacles: optically discernible traits resolved from planktonic organisms in situ. ICES Journal of Marine Science, 76(7):1959–1972, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/7/1959/5580550.

Oigaard:2014:PPQ

[ØHN14] Tor Arne Øigård, Tore Haug, and Kjell Tormod Nilssen. From pup production to quotas: current status of harp seals in the Greenland Sea. ICES Journal of Marine Science, 71(3):537–545, March 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/71/3/537/635422.

Okland:2010:MDM

[ØHS10] Jan-Magnus Økland, Øystein Ariansen Haaland, and Hans J. Skaug. A method for defining management units based on genetically determined close relatives. ICES Journal of Marine Science, 67(3):551–558, April 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/3/551/732928.

Osterblom:2017:TTS

[ÖHSNV17] Henrik Österblom, Jonas Hentati-Sundberg, Nea Nevonen, and Katarina Veem. Tinkering with a tanker — slow evolution of a Swedish ecosystem approach. ICES Journal of Marine Science, 74(1):443–452, January 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/1/443/2967556.

ONeill:2016:PIT

[OI16] F. G. O’Neill and A. Ivanović. The physical impact of towed demersal fishing gears on soft sediments. ICES Journal of Marine Science, 73(suppl 1):S5–S14, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/73/suppl_1/i5/2573993.

Ono:2018:IDM

[OIMP18] Kotaro Ono, James N. Ianelli, Carey R. McGilliard, and André E. Punt. Integrating data from multiple surveys and accounting for spatio-temporal correlation to index the abundance of juvenile Pacific halibut in Alaska. ICES Journal of Marine Science, 75(2):572–584, March 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/2/572/4210474.

Olsson:2019:FLS

[OJK^{+}19] Jens Olsson, Eglė Jakubavičiūtė, Olavi Kaljuste, Niklas Larsson, Ulf Bergström, Michele Casini, Massimiliano Cardinale, Joakim Hjelm, and Pär Byström. The first largescale assessment of three-spined stickleback (Gasterosteus aculeatus) biomass and spatial distribution in the Baltic Sea. ICES Journal of Marine Science, 76(6):1653–1665, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/6/1653/5480400.

Ono:2018:MAD

[OKDJ18] Kotaro Ono, Stan Kotwicki, Gjert E. Dingsør, and Espen Johnsen. Multispecies acoustic dead-zone correction and bias ratio estimates between acoustic and bottom-trawl data. ICES Journal of Marine Science, 75(1):361–373, January 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/1/361/3931851.

Ono:2015:ILA

[OLM^{+}15] Kotaro Ono, Roberto Licandeo, Melissa L. Muradian, Curry J. Cunningham, Sean C. Anderson, Felipe Hurtado-Ferro, Kelli F. Johnson, Carey R. McGilliard, Cole C. Monnahan, Cody S. Szuwalski, Juan L. Valero, Katyana A. Vert-Pre, Athol R. Whitten, and André E. Punt. The importance of length and age composition data in statistical age-structured models for marine species. ICES Journal of Marine Science, 72(1):31–43, January 2015.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/1/31/816216.

Ojaveer:2018:STC

[OLR^{+}18] Henn Ojaveer, Ain Lankov, Tiit Raid, Arno Põllumäe, and Riina Klais. Selecting for three copepods — feeding of sprat and herring in the Baltic Sea. ICES Journal of Marine Science, 75(7):2439–2449, December 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/7/2439/4836298.

ONeill:2014:LSS

[OLW^{+}14] Michael F. O’Neill, George M. Leigh, You-Gan Wang, J. Matías Braccini, and Matthew C. Ives. Linking spatial stock dynamics and economics: evaluation of indicators and fishery management for the travelling eastern king prawn (Melicertus plebejus). ICES Journal of Marine Science, 71(7):1818–1834, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/7/1818/663348.

Ommer:2018:CIG

[Omm18] Rosemary E. Ommer. Curiosity, interdisciplinarity, and giving back. ICES Journal of Marine Science, 75(5):1526– 1535, September 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/5/1526/4961487.

Ohata:2014:MTE

[OMTY14] Ryosuke Ohata, Reiji Masuda, Kohji Takahashi, and Yoh Yamashita. Moderate turbidity enhances schooling behaviour in fish larvae in coastal waters. ICES Journal of Marine Science, 71(4):925–929, May 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/4/925/663331.

Ojaveer:2018:SUB

[OND^{+}18] Henn Ojaveer, Stefan Neuenfeldt, Jan Dierking, Liina Eek, Jannica Haldin, Georg Martin, Kaire Märtin, Kaja Peterson, and Sebastian Valanko. Sustainable use of Baltic Sea resources. ICES Journal of Marine Science, 75(7):2434– 2438, December 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/7/2434/5107799.

ODriscoll:2013:STS

[OOD13] Richard L. O’Driscoll, Johannes Oeffner, and Adam J. Dunford. In situ target strength estimates of optically verified southern blue whiting (Micromesistius australis). ICES Journal of Marine Science, 70(2):431–439, March 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/2/431/795387.

Ozdemir:2019:NPL

[ÖPPM19] Nurgül Şen Özdemir, Christopher C. Parrish, Camilla Parzanini, and Annie Mercier. Neutral and polar lipid fatty acids in five families of demersal and pelagic fish from the deep Northwest Atlantic. ICES Journal of Marine Science, 76(6):1807–1815, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
6/1807/5475846.

Otteraa:2014:PIS

[OS14] Håkon Otterå and Ove T. Skilbrei. Possible influence of salmon farming on long-term resident behaviour of wild saithe (Pollachius virens L.). ICES Journal of Marine Science, 71(9):2484–2493, November 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/9/2484/2798173.

Oigaard:2015:FSS

[ØS15] Tor Arne Øigård and Hans J. Skaug. Fitting state– space models to seal populations with scarce data. ICES Journal of Marine Science, 72(5):1462–1469, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/5/1462/756991.

Ouellet:2016:CRT

[OSBG16] Patrick Ouellet, Claude Savenkoff, Hugues P. Benoît, and Peter S. Galbraith. A comparison of recent trends in demersal fish biomass and their potential drivers for three ecoregions of the Gulf of St Lawrence, Canada. ICES Journal of Marine Science, 73(2):329–344, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/2/329/2614178.

OFarrell:2017:IDS

[OSC^{+}17] Shay O’Farrell, James N. Sanchirico, Iliana Chollett, Marcy Cockrell, Steven A. Murawski, Jordan T. Watson, Alan Haynie, Andrew Strelcheck, and Larry Perruso. Improving detection of short-duration fishing behaviour in vessel tracks by feature engineering of training data. ICES Journal of Marine Science, 74(5):1428–1436, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/5/1428/2966516.

Olafsdottir:2016:CWL

[OSJ^{+}16] Anna H. Olafsdottir, Aril Slotte, Jan Arge Jacobsen, Gudmundur J. Oskarsson, Kjell R. Utne, and Leif Nøttestad. Changes in weight-at-length and size-at-age of mature Northeast Atlantic mackerel (Scomber scombrus) from 1984 to 2013: effects of mackerel stock size and herring (Clupea harengus) stock size. ICES Journal of Marine Science, 73(4):1255–1265, March 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
4/1255/2457869.

Oskarsson:2018:EPC

[Ósk18] Gudmundur J. Óskarsson. The existence and population connectivity of Icelandic spring-spawning herring over a 50-year collapse period. ICES Journal of Marine Science, 75(6):2025–2032, November 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/6/2025/5098311.

OConnell:2015:TPD

[OSL^{+}15] Victoria O’Connell, Janice Straley, Joe Liddle, Lauren Wild, Linda Behnken, Dan Falvey, and Aaron Thode. Testing a passive deterrent on longlines to reduce sperm whale depredation in the Gulf of Alaska. ICES Journal of Marine Science, 72(5):1667–1672, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/5/1667/776210.

Ouellet:2018:VDS

[OSM18] Patrick Ouellet and Bernard Sainte-Marie. Vertical distribution of snow crab (Chionoecetes opilio) pelagic stages in the Gulf of St. Lawrence (Canada) and effect of temperature on development and survival. ICES Journal of Marine Science, 75(2):773–784, March 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/2/773/4210371.

ONeill:2013:MGE

[OSP^{+}13] F. G. O’Neill, S. M. Simmons, D. R. Parsons, J. L. Best, P. J. Copland, F. Armstrong, M. Breen, and K. Summerbell. Monitoring the generation and evolution of the sediment plume behind towed fishing gears using a multibeam echosounder. ICES Journal of Marine Science, 70(4):892– 903, July 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/70/4/892/726895.

Otto:2016:MPB

[OSSL16] Saskia A. Otto, Sarah Simons, Joshua S. Stoll, and Peter Lawson. Making progress on bycatch avoidance in the ocean salmon fishery using a transdisciplinary approach. ICES Journal of Marine Science, 73(9):2380–2394, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/9/2380/2198334.

Ojaveer:2015:FOA

[OTAK15] Henn Ojaveer, Jonna Tomkiewicz, Timo Arula, and Riina Klais. Female ovarian abnormalities and reproductive failure of autumn-spawning herring (Clupea harengus membras) in the Baltic Sea. ICES Journal of Marine Science, 72(8):2332–2340, September 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/72/
8/2332/2458908.

Ovenden:2016:SSP

[OTM^{+}16] Jennifer R. Ovenden, Bree J. Tillett, Michael Macbeth, Damien Broderick, Fiona Filardo, Raewyn Street, Sean R. Tracey, and Jayson Semmens. Stirred but not shaken: population and recruitment genetics of the scallop (Pecten fumatus) in Bass Strait, Australia. ICES Journal of Marine Science, 73(9):2333–2341, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/9/2333/2198452.

Olsson:2015:TDC

[OTO^{+}15] Jens Olsson, Maciej T. Tomczak, Henn Ojaveer, Anna Gårdmark, Arno Põllumäe, Bärbel Müller-Karulis, Didzis Ustups, Grete E. Dinesen, Heikki Peltonen, Ivars Putnis, Lena Szymanek, Mart Simm, Outi Heikinheimo, Pavel Gasyukov, Philip Axe, and Lena Bergström. Temporal development of coastal ecosystems in the Baltic Sea over the past two decades. ICES Journal of Marine Science, 72(9):2539–2548, November 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/72/
9/2539/2458711.

Ottersen:2010:DTA

[Ott10] Geir Ottersen. A digital temperature atlas for the Norwegian Sea. ICES Journal of Marine Science, 67(8):1525– 1537, November 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/67/8/1525/607714.

Oresland:2013:ELS

[ØU13] Vidar Øresland and Mats Ulmestrand. European lobster subpopulations from limited adult movements and larval retention. ICES Journal of Marine Science, 70(3):532–539, April 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/70/3/532/918763.

Ovenden:2019:BMH

[Ove19] Jennifer R. Ovenden. Breaking the myths (or how to have a successful career in science). ICES Journal of Marine Science, 76(1):23–27, January 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/76/1/23/5146511.

Ozerov:2017:CMB

[OVW^{+}17] Mikhail Ozerov, Juha-Pekka Vähä, Vidar Wennevik, Eero Niemelä, Martin-A. Svenning, Sergey Prusov, Rogelio Diaz Fernandez, Laila Unneland, Anti Vasemägi, Morten Falkegård, Tiia Kalske, and Bente Christiansen. Comprehensive microsatellite baseline for genetic stock identification of Atlantic salmon (Salmo salar L.) in northernmost Europe. ICES Journal of Marine Science, 74(8):2159– 2169, September 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/74/8/2159/3609001.

Okamura:2017:RVP

[OYI17] Hiroshi Okamura, Yuuho Yamashita, and Momoko Ichinokawa. Ridge virtual population analysis to reduce the instability of fishing mortalities in the terminal year. ICES Journal of Marine Science, 74(9):2427–2436, November 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/9/2427/3855119.

Okamura:2018:CPA

[OYIN18] Hiroshi Okamura, Yuuho Yamashita, Momoko Ichinokawa, and Shota Nishijima. Comparison of the performance of age-structured models with few survey indices. ICES Journal of Marine Science, 75(6):2016–2024, November 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/6/2016/5096001.

Planque:2018:PCA

[PA18] Benjamin Planque and Per Arneberg. Principal component analyses for integrated ecosystem assessments may primarily reflect methodological artefacts. ICES Journal of Marine Science, 75(3):1021–1028, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/3/1021/4769309.

Punt:2014:FMU

[PAB^{+}14] André E. Punt, Teresa A’mar, Nicholas A. Bond, Douglas S. Butterworth, Carryn L. de Moor, José A. A. De Oliveira, Melissa A. Haltuch, Anne B. Hollowed, and Cody Szuwalski. Fisheries management under climate and environmental uncertainty: control rules and performance simulation. ICES Journal of Marine Science, 71(8):2208–2220, October 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/71/8/2208/744434.

Pendleton:2018:DEM

[PAB^{+}18] Linwood H. Pendleton, Gabby N. Ahmadia, Howard I. Browman, Ruth H. Thurstan, David M. Kaplan, and Valerio Bartolino. Debating the effectiveness of marine protected areas. ICES Journal of Marine Science, 75(3):1156– 1159, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/3/1156/4098843.

Packard:2018:TWC

[Pac18] Theodore T. Packard. From Thoreau’s woods to the Canary Islands: exploring ocean biogeochemistry through enzymology. ICES Journal of Marine Science, 75(3):912–922, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/3/912/4731708.

Paffenhofer:2018:HCC

[Paf18] Gustav-Adolf Paffenhöfer. How cooperation contributes to scientific advances. ICES Journal of Marine Science, 75(2): 494–501, March 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/2/494/4345802.

Parada:2013:CAG

[Par13] J. M. Parada. A correction to “Assessment of goose barnacle (Pollicipes pollicipes Gmelin, 1789) stocks in management plans: design of a sampling program based on the harvesters’ experience”. ICES Journal of Marine Science, 70 (1):244, January 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/70/1/244/663441.

Passantino:2014:ESF

[Pas14] Annamaria Passantino. The EU shark finning ban at the beginning of the new millennium: the legal framework. ICES Journal of Marine Science, 71(3):429–434, March 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/3/429/639895.

Pauly:2016:HSH

[Pau16] Daniel Pauly. Having to science the hell out of it. ICES Journal of Marine Science, 73(9):2156–2166, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/9/2156/2198225.

Payne:2010:MGS

[Pay10] Mark R. Payne. Mind the gaps: a state-space model for analysing the dynamics of North Sea herring spawning components. ICES Journal of Marine Science, 67(9):1939– 1947, December 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/67/9/1939/619011.

Palsson:2011:LTC

[PB11] Ólafur K. Pálsson and Höskuldur Björnsson. Longterm changes in trophic patterns of Iceland cod and linkages to main prey stock sizes. ICES Journal of Marine Science, 68(7):1488–1499, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/68/
7/1488/654725.

Plourde:2014:POZ

[PB14] Stéphane Plourde and Howard I. Browman. Parameterizing and operationalizing zooplankton population dynamic and trophic interaction models: Introduction. ICES Journal of Marine Science, 71(2):234–235, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/2/234/790963.

Payne:2016:UPC

[PBC^{+}16] Mark R. Payne, Manuel Barange, William W. L. Cheung, Brian R. MacKenzie, Harold P. Batchelder, Xochitl Cormon, Tyler D. Eddy, Jose A. Fernandes, Anne B. Hollowed, Miranda C. Jones, Jason S. Link, Philipp Neubauer, Ivonne Ortiz, Ana M. Queirós, and José Ricardo Paula. Uncertainties in projecting climate-change impacts in marine ecosystems. ICES Journal of Marine Science, 73(5): 1272–1282, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/73/5/1272/2240686.

Pendleton:2019:DDS

[PBE^{+}19] Linwood H. Pendleton, Hawthorne Beyer, Estradivari, Susan O. Grose, Ove Hoegh-Guldberg, Denis B. Karcher, Emma Kennedy, Lyndon Llewellyn, Cecile Nys, Aurélie Shapiro, Rahul Jain, Katarzyna Kuc, Terry Leatherland, Kira O’Hainnin, Guillermo Olmedo, Lynette Seow, and Mick Tarsel. Disrupting data sharing for a healthier ocean. ICES Journal of Marine Science, 76(6):1415–1423, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/6/1415/5480138.

Pinkerton:2014:CFS

[PBG14] Matthew H. Pinkerton and Janet M. Bradford-Grieve. Characterizing foodweb structure to identify potential ecosystem effects of fishing in the Ross Sea, Antarctica. ICES Journal of Marine Science, 71(7):1542–1553, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/71/7/1542/665353.

Pacariz:2014:MSL

[PBJ^{+}14] Selma Pacariz, Göran Björk, Patrik Jonsson, Patrik Börjesson, and Henrik Svedäng. A model study of the largescale transport of fish eggs in the Kattegat in relation to egg density. ICES Journal of Marine Science, 71(2):345–355, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/71/2/345/785570.

Planque:2011:UPS

[PBL11] Benjamin Planque, Edwige Bellier, and Christophe Loots. Uncertainties in projecting spatial distributions of marine populations. ICES Journal of Marine Science, 68(6):1045– 1050, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/68/6/1045/697747.

Poos:2010:IQF

[PBQ^{+}10] J. J. Poos, J. A. Bogaards, F. J. Quirijns, D. M. Gillis, and A. D. Rijnsdorp. Individual quotas, fishing effort allocation, and over-quota discarding in mixed fisheries. ICES Journal of Marine Science, 67(2):323–333, March 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/2/323/693826.

Pacariz:2014:IVT

[PBS14] Selma Pacariz, Göran Björk, and Henrik Svedäng. Interannual variability in the transport of fish eggs in the Kattegat and Öresund. ICES Journal of Marine Science, 71(7):1706– 1716, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/71/7/1706/2804371.

Padilla:2015:SIA

[PBW15] Andrew J. Padilla, Randy J. Brown, and Matthew J. Wooller. Strontium isotope analyses (^{87}Sr/^{86}Sr) of otoliths from anadromous Bering cisco (Coregonus laurettae) to determine stock composition. ICES Journal of Marine Science, 72(7):2110–2117, October 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/72/
7/2110/2457879.

Phillips:2014:STA

[PCB^{+}14] A. Jason Phillips, Lorenzo Ciannelli, Richard D. Brodeur, William G. Pearcy, and John Childers. Spatio-temporal associations of albacore CPUEs in the Northeastern Pacific with regional SST and climate environmental variables. ICES Journal of Marine Science, 71(7):1717–1727, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/71/7/1717/665880.

Pineiro-Corbeira:2018:CNF

[PCBO^{+}18] Cristina Piñeiro-Corbeira, Sara Barrientos, Mercedes Olmedo, Javier Cremades, and Rodolfo Barreiro. Bycatch in no-fed aquaculture: exploiting mussel seed persistently and extensively disturbs the accompanying assemblage. ICES Journal of Marine Science, 75(6):2213–2223, November 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/6/2213/5075182.

Pham:2013:TMF

[PCD^{+}13] Christopher K. Pham, Angela Canha, Hugo Diogo, João G. Pereira, Rui Prieto, and Telmo Morato. Total marine fishery catch for the Azores (1950–2010). ICES Journal of Marine Science, 70(3):564–577, April 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/3/564/919093.

Penn:2015:RLF

[PCdL15] J. W. Penn, N. Caputi, and S. de Lestang. A review of lobster fishery management: the Western Australian fishery for Panulirus cygnus, a case study in the development and implementation of input and output-based management systems. ICES Journal of Marine Science, 72(S1): S22–S34, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/72/suppl_1/i22/619927.

Pinchuk:2013:EAT

[PCFR13] Alexei I. Pinchuk, Kenneth O. Coyle, Edward V. Farley, and Heather M. Renner. Emergence of the Arctic Themisto libellula (Amphipoda: Hyperiidae) on the southeastern Bering Sea shelf as a result of the recent cooling, and its potential impact on the pelagic food web. ICES Journal of Marine Science, 70(6):1244–1254, September 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/6/1244/633762.

Pennino:2016:FDI

[PCLQ^{+}16] Maria Grazia Pennino, David Conesa, Antonio López-Quílez, Facundo Muñoz, Angel Fernández, and Jose Maria Bellido. Fishery-dependent and -independent data lead to consistent estimations of essential habitats. ICES Journal of Marine Science, 73(9):2302–2310, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/9/2302/2198379.

Pardo:2018:QKU

[PCRD18] Sebastián A. Pardo, Andrew B. Cooper, John D. Reynolds, and Nicholas K. Dulvy. Quantifying the known unknowns: estimating maximum intrinsic rate of population increase in the face of uncertainty. ICES Journal of Marine Science, 75 (3):953–963, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/3/953/4791419.

Przeslawski:2011:USH

[PCS^{+}11] Rachel Przeslawski, David R. Currie, Shirley J. Sorokin, Tim M. Ward, Franziska Althaus, and Alan Williams. Utility of a spatial habitat classification system as a surrogate of marine benthic community structure for the Australian margin. ICES Journal of Marine Science, 68(9):1954– 1962, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/9/1954/665764.

Petrik:2016:MCB

[PDAC^{+}16] Colleen M. Petrik, Janet T. Duffy-Anderson, Frederic Castruccio, Enrique N. Curchitser, Seth L. Danielson, Katherine Hedstrom, and Franz Mueter. Modelled connectivity between walleye Pollock (Gadus chalcogrammus) spawning and age-0 nursery areas in warm and cold years with implications for juvenile survival. ICES Journal of Marine Science, 73(7):1890–1900, July 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/7/1890/2458227.

Pinho:2014:HJB

[PDCP14] Mário Pinho, Hugo Diogo, Joana Carvalho, and João Gil Pereira. Harvesting juveniles of blackspot sea bream (Pagellus bogaraveo) in the Azores (Northeast Atlantic): biological implications, management, and life cycle considerations. ICES Journal of Marine Science, 71(9):2448–2456, November 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/71/9/2448/2798170.

Pecl:2010:TMC

[PDD^{+}10a] Gretta T. Pecl, Zoë A. Doubleday, Leonid Danyushevsky, Sarah Gilbert, and Natalie A. Moltschaniwskyj. Transgenerational marking of cephalopods with an enriched barium isotope: a promising tool for empirically estimating post-hatching movement and population connectivity. ICES Journal of Marine Science, 67(7):1372–1380, October 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/7/1372/661781.

Punt:2010:ISS

[PDD^{+}10b] André E. Punt, Roy A. Deng, Catherine M. Dichmont, Tom Kompas, William N. Venables, Shijie Zhou, Sean Pascoe, Trevor Hutton, Rob Kenyon, Tonya van der Velde, and Marco Kienzle. Integrating size-structured assessment and bioeconomic management advice in Australia’s northern prawn fishery. ICES Journal of Marine Science, 67(8):1785–1801, November 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/67/
8/1785/602838.

Prince:2011:SCE

[PDD^{+}11] Jeremy D. Prince, Natalie A. Dowling, Campbell R. Davies, Robert A. Campbell, and Dale S. Kolody. A simple costeffective and scale-less empirical approach to harvest strategies. ICES Journal of Marine Science, 68(5):947–960, May 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/5/947/651393.

Petitgas:2014:MVF

[PDH^{+}14] Pierre Petitgas, Mathieu Doray, Martin Huret, Jacques Massé, and Mathieu Woillez. Modelling the variability in fish spatial distributions over time with empirical orthogonal functions: anchovy in the Bay of Biscay. ICES Journal of Marine Science, 71(9):2379–2389, November 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/9/2379/2798181.

Petitgas:2011:SEE

[PDMG11] Pierre Petitgas, Mathieu Doray, Jacques Massé, and Patrick Grellier. Spatially explicit estimation of fish length histograms, with application to anchovy habitats in the Bay of Biscay. ICES Journal of Marine Science, 68(10):2086– 2095, November 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/10/2086/610244.

Pointin:2019:UAB

[PDR19] Fabien Pointin, Fabienne Daurès, and Marie-Joëlle Rochet. Use of avoidance behaviours to reduce the economic impacts of the EU Landing Obligation: the case study of a mixed trawl fishery. ICES Journal of Marine Science, 76(6):1554–1566, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
6/1554/5419278.

Passadore:2015:AMM

[PDS15a] Cecilia Passadore, Andrés Domingo, and Eduardo R. Secchi. Analysis of marine mammal bycatch in the Uruguayan pelagic longline fishery operating in the Southwestern Atlantic Ocean. ICES Journal of Marine Science, 72(5):1637– 1652, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/5/1637/769111.

Passadore:2015:DKW

[PDS15b] Cecilia Passadore, Andrés Domingo, and Eduardo R. Secchi. Depredation by killer whale (Orcinus orca) and false killer whale (Pseudorca crassidens) on the catch of the Uruguayan pelagic longline fishery in Southwestern Atlantic Ocean. ICES Journal of Marine Science, 72(5):1653– 1666, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/5/1653/770143.

Poole:2018:LTV

[PDT^{+}18] W. Russell Poole, Ola H. Diserud, Eva B. Thorstad, Caroline M. Durif, Conor Dolan, Odd Terje Sandlund, Knut Bergesen, Gerard Rogan, Sean Kelly, and Leif Asbjørn Vøllestad. Long-term variation in numbers and biomass of silver eels being produced in two European river systems. ICES Journal of Marine Science, 75(5):1627–1637, September 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/5/1627/4994249.

Polovina:2011:PES

[PDWH11] Jeffrey J. Polovina, John P. Dunne, Phoebe A. Woodworth, and Evan A. Howell. Projected expansion of the subtropical biome and contraction of the temperate and equatorial upwelling biomes in the North Pacific under global warming. ICES Journal of Marine Science, 68(6):986–995, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/6/986/697199.

Papetti:2013:SPC

[PDZ^{+}13] Chiara Papetti, Antonio Di Franco, Lorenzo Zane, Paolo Guidetti, Valeria De Simone, Marianna Spizzotin, Barbara Zorica, Vanja Čikeš Keč, and Carlotta Mazzoldi. Single population and common natal origin for Adriatic Scomber scombrus stocks: evidence from an integrated approach. ICES Journal of Marine Science, 70(2):387–398, March 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/2/387/798183.

Pohlot:2018:ASD

[PE18] Bruce G. Pohlot and Nelson Ehrhardt. An analysis of sailfish daily activity in the Eastern Pacific Ocean using satellite tagging and recreational fisheries data. ICES Journal of Marine Science, 75(2):871–879, March 2018.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/2/871/3845292.

Pena:2019:MFA

[Peñ19a] Marian Peña. Mesopelagic fish avoidance from the vessel dynamic positioning system. ICES Journal of Marine Science, 76(3):734–742, May 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
3/734/5181375.

Pinnegar:2019:AVA

[PEN^{+}19b] John K. Pinnegar, Georg H. Engelhard, Norman J. Norris, Derrick Theophille, and Riviere Delanco Sebastien. Assessing vulnerability and adaptive capacity of the fisheries sector in Dominica: long-term climate change and catastrophic hurricanes. ICES Journal of Marine Science, 76(5):1353– 1367, 09- 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/5/1353/5512304.

Pepin:2016:DNF

[Pep16] Pierre Pepin. Death from near and far: alternate perspectives on size-dependent mortality in larval fish. ICES Journal of Marine Science, 73(2):196–203, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/2/196/2614195.

Peterman:2019:CLT

[Pet19] Randall M. Peterman. Continuous learning, teamwork, and lessons for young scientists. ICES Journal of Marine Science, 76(1):28–40, January 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/76/1/28/5116191.

Pitcher:2016:ETS

[PEV^{+}16] C. Roland Pitcher, Nick Ellis, William N. Venables, Ted J. Wassenberg, Charis Y. Burridge, Greg P. Smith, Matthew Browne, Francis Pantus, Ian R. Poiner, Peter J. Doherty, John N. A. Hooper, and Neil Gribble. Effects of trawling on sessile megabenthos in the Great Barrier Reef and evaluation of the efficacy of management strategies. ICES Journal of Marine Science, 73(suppl 1):S115–S126, January 2016. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/73/suppl_1/i115/2573860.

Punt:2016:ELT

[PFD^{+}16] André E. Punt, Robert J. Foy, Michael G. Dalton, W. Christopher Long, and Katherine M. Swiney. Effects of long-term exposure to ocean acidification conditions on future southern Tanner crab (Chionoecetes bairdi) fisheries management. ICES Journal of Marine Science, 73(3):849– 864, February 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/73/3/849/2458912.

Pohlmann:2016:MLS

[PFH16] Jan-Dag Pohlmann, Marko Freese, and Reinhold Hanel. Minimum landing size in European eel fisheries management: limitations of simplistic management approaches in a semelparous species. ICES Journal of Marine Science, 73(10):2509–2517, November 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
10/2509/2647093.

Petursdottir:2012:TIM

[PFPG12] Hildur Petursdottir, Stig Falk-Petersen, and Astthor Gislason. Trophic interactions of meso- and macrozooplankton and fish in the Iceland Sea as evaluated by fatty acid and stable isotope analysis. ICES Journal of Marine Science, 69(7):1277–1288, September 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/7/1277/752471.

Postuma:2010:RBS

[PG10] Felippe A. Postuma and Maria A. Gasalla. On the relationship between squid and the environment: artisanal jigging for Loligo plei at São Sebastião Island (24^{o}S), southeastern Brazil. ICES Journal of Marine Science, 67(7):1353–1362, October 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/67/7/1353/665751.

Padron:2016:MED

[PG16] Mariana Padrón and Katell Guizien. Modelling the effect of demographic traits and connectivity on the genetic structuration of marine metapopulations of sedentary benthic invertebrates. ICES Journal of Marine Science, 73(7):1935– 1945, July 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/7/1935/2457906.

Palsson:2012:ESI

[PGG^{+}12] Ólafur K. Pálsson, Astthor Gislason, Hafsteinn G. Gudfinnsson, Björn Gunnarsson, Sólveig R. Ólafsdóttir, Hildur Petursdottir, Sveinn Sveinbjörnsson, Konrad Thorisson, and Hédinn Valdimarsson. Ecosystem structure in the Iceland Sea and recent changes to the capelin (Mallotus villosus) population. ICES Journal of Marine Science, 69(7):1242–1254, September 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/69/
7/1242/744363.

Parker:2017:DFU

[PGG^{+}17] Robert W. R. Parker, Caleb Gardner, Bridget S. Green, Klaas Hartmann, and Reg A. Watson. Drivers of fuel use in rock lobster fisheries. ICES Journal of Marine Science, 74(6):1681–1689, July 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
6/1681/3066286.

Priede:2011:RSE

[PGN^{+}11] Imants G. Priede, Jasmin A. Godbold, Tomasz Niedzielski, Martin A. Collins, David M. Bailey, John D. M. Gordon, and Alain F. Zuur. A review of the spatial extent of fishery effects and species vulnerability of the deepsea demersal fish assemblage of the Porcupine Seabight, Northeast Atlantic Ocean (ICES Subarea VII). ICES Journal of Marine Science, 68(2):281–289, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/2/281/613939.

Pedersen:2011:RTS

[PGOM11] Geir Pedersen, Olav Rune Godø, Egil Ona, and Gavin J. Macaulay. A revised target strength–length estimate for blue whiting (Micromesistius poutassou): implications for biomass estimates. ICES Journal of Marine Science, 68(10):2222–2228, November 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/68/
10/2222/610797.

Pernet:2019:FMRa

[PGS^{+}19a] Fabrice Pernet, Sonia Gachelin, Jean-Yves Stanisière, Bruno Petton, Elodie Fleury, and Joseph Mazurié. Farmer monitoring reveals the effect of tidal height on mortality risk of oysters during a herpesvirus outbreak. ICES Journal of Marine Science, 76(6):1816–1824, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/6/1816/5477843.

Pernet:2019:FMRb

[PGS^{+}19b] Fabrice Pernet, Sonia Gachelin, Jean-Yves Stanisière, Bruno Petton, Elodie Fleury, and Joseph Mazurié. Farmer monitoring reveals the effect of tidal height on mortality risk of oysters during a herpesvirus outbreak. ICES Journal of Marine Science, 76(6):1932–1933, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/6/1932/5511588.

Parada:2017:ECF

[PGV^{+}17] Carolina Parada, Alexandre Gretchina, Sebastián Vásquez, Ali Belmadani, Vincent Combes, Billy Ernst, Emanuele Di Lorenzo, Javier Porobic, and Aquiles Sepúlveda. Expanding the conceptual framework of the spatial population structure and life history of jack mackerel in the eastern South Pacific: an oceanic seamount region as potential spawning/nursery habitat. ICES Journal of Marine Science, 74(9):2398–2414, November 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/9/2398/3831740.

Pennington:2011:EDE

[PH11] Michael Pennington and Kristin Helle. Evaluation of the design and efficiency of the Norwegian self-sampling purse-seine reference fleet. ICES Journal of Marine Science, 68(8):1764–1768, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/68/
8/1764/746343.

Pfeiffer:2012:EDS

[PH12a] Lisa Pfeiffer and Alan C. Haynie. The effect of decreasing seasonal sea-ice cover on the winter Bering Sea pollock fishery. ICES Journal of Marine Science, 69(7):1148– 1159, September 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/69/7/1148/748562.

Piet:2012:IFP

[PH12b] G. J. Piet and N. T. Hintzen. Indicators of fishing pressure and seafloor integrity. ICES Journal of Marine Science, 69(10):1850–1858, December 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/69/
10/1850/624669.

Pauli:2013:EEE

[PH13] Beatriz Diaz Pauli and Mikko Heino. Ecological and evolutionary effects of harvesting: lessons from the candy-fish experiment. ICES Journal of Marine Science, 70(7):1281– 1286, November 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/70/7/1281/611906.

Peterson:2017:SMA

[PH17] Megan J. Peterson and Dana Hanselman. Sablefish mortality associated with whale depredation in Alaska. ICES Journal of Marine Science, 74(5):1382–1394, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/5/1382/2884303.

Punzon:2010:SOT

[PHA^{+}10] Antonio Punzón, Carmen Hernández, Esther Abad, José Castro, Nelida Pérez, and Valentín Trujillo. Spanish otter trawl fisheries in the Cantabrian Sea. ICES Journal of Marine Science, 67(8):1604–1616, November 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/8/1604/606197.

Portnoy:2016:PSG

[PHB^{+}16] David S. Portnoy, Christopher M. Hollenbeck, Dana M. Bethea, Bryan S. Frazier, Jim Gelsleichter, and John R. Gold. Population structure, gene flow, and historical demography of a small coastal shark (Carcharhinus isodon) in US waters of the Western Atlantic Ocean. ICES Journal of Marine Science, 73(9):2322–2332, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/9/2322/2199306.

Pekcan-Hekim:2016:RCF

[PHGK^{+}16] Zeynep Pekcan-Hekim, Anna Gårdmark, Agnes M. L. Karlson, Pirkko Kauppila, Mikaela Bergenius, and Lena Bergström. The role of climate and fisheries on the temporal changes in the Bothnian Bay foodweb. ICES Journal of Marine Science, 73(7):1739–1749, July 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/7/1739/2458697.

Puerta:2016:CEI

[PHH^{+}16] Patricia Puerta, Mary E. Hunsicker, Manuel Hidalgo, Patricia Reglero, Lorenzo Ciannelli, Antonio Esteban, María González, and Antoni Quetglas. Community–environment interactions explain octopus-catshark spatial overlap. ICES Journal of Marine Science, 73(7):1901–1911, July 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/7/1901/2458758.

Proud:2019:SDS

[PHK^{+}19] Roland Proud, Nils Olav Handegard, Rudy J. Kloser, Martin J. Cox, and Andrew S. Brierley. From siphonophores to deep scattering layers: uncertainty ranges for the estimation of global mesopelagic fish biomass. ICES Journal of Marine Science, 76(3):718–733, May 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/3/718/4978316.

Punt:2013:RIS

[PHM13] André E. Punt, TzuChuan Huang, and Mark N. Maunder. Review of integrated size-structured models for stock assessment of hard-to-age crustacean and mollusc species. ICES Journal of Marine Science, 70(1):16–33, January 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/1/16/663084.

Paulsen:2014:NSL

[PHM^{+}14] Matthias Paulsen, Cornelius Hammer, Arne M. Malzahn, Patrick Polte, Christian von Dorrien, and Catriona Clemmesen. Nutritional situation for larval Atlantic herring (Clupea harengus L.) in two nursery areas in the western Baltic Sea. ICES Journal of Marine Science, 71(4): 991–1000, May 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/71/4/991/666562.

Pena:2013:FHS

[PHO13] Héctor Peña, Nils Olav Handegard, and Egil Ona. Feeding herring schools do not react to seismic air gun surveys. ICES Journal of Marine Science, 70(6):1174–1180, September 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/70/6/1174/639044.

Prince:2015:RCB

[PHV^{+}15] Jeremy Prince, Adrian Hordyk, Sarah R. Valencia, Neil Loneragan, and Keith Sainsbury. Revisiting the concept of Beverton–Holt life-history invariants with the aim of informing data-poor fisheries assessment. ICES Journal of Marine Science, 72(1):194–203, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/1/194/816563.

Pinto:2017:HCB

[Pin17] Bruno Pinto. Historical connections between early marine science research and dissemination: the case study of aquarium Vasco Da Gama (Portugal) from late 19th century to mid-20th century. ICES Journal of Marine Science, 74(6): 1522–1530, July 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/74/6/1522/3074694.

Planque:2012:HVY

[PJDN12] B. Planque, E. Johannesen, K. V. Drevetnyak, and K. H. Nedreaas. Historical variations in the year-class strength of beaked redfish (Sebastes mentella) in the Barents Sea. ICES Journal of Marine Science, 69(4):547–552, May 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/4/547/632214.

Polte:2014:SBE

[PKHG14] Patrick Polte, Paul Kotterba, Cornelius Hammer, and Tomas Gröhsler. Survival bottlenecks in the early ontogenesis of Atlantic herring (Clupea harengus, L.) in coastal lagoon spawning areas of the western Baltic Sea. ICES Journal of Marine Science, 71(4):982–990, May 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/4/982/664266.

Peck:2012:TWS

[PKHM12] Myron A. Peck, Philipp Kanstinger, Linda Holste, and Meike Martin. Thermal windows supporting survival of the earliest life stages of Baltic herring (Clupea harengus). ICES Journal of Marine Science, 69(4):529–536, May 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/4/529/635278.

Probst:2013:IBS

[PKK13] Wolfgang Nikolaus Probst, Matthias Kloppmann, and Gerd Kraus. Indicator-based status assessment of commercial fish species in the North Sea according to the EU Marine Strategy Framework Directive (MSFD). ICES Journal of Marine Science, 70(3):694–706, April 2013.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/3/694/917293.

Pons:2019:PLB

[PKR^{+}19] Maite Pons, Laurence Kell, Merrill B. Rudd, Jason M. Cope, and Flávia Lucena Frédou. Performance of lengthbased data-limited methods in a multifleet context: application to small tunas, mackerels, and bonitos in the Atlantic Ocean. ICES Journal of Marine Science, 76(4):960– 973, July 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/4/960/5307953.

Planque:2016:PFS

[Pla16] Benjamin Planque. Projecting the future state of marine ecosystems, “la grande illusion”? ICES Journal of Marine Science, 73(2):204–208, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/2/204/2614187.

Plank:2017:BHB

[Pla17] Michael J. Plank. Balanced harvesting is the bioeconomic equilibrium of a size-structured Beverton–Holt model. ICES Journal of Marine Science, 74(1):112–120, January 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/1/112/2669575.

Purcell:2018:SSF

[PLCC18] Steven W. Purcell, Watisoni Lalavanua, Brian R. Cullis, and Nicole Cocks. Small-scale fishing income and fuel consumption: Fiji’s artisanal sea cucumber fishery. ICES Journal of Marine Science, 75(5):1758–1767, September 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/5/1758/4955261.

Pulver:2016:MCS

[PLSD16] Jeffrey Robert Pulver, Hui Liu, and Elizabeth Scott-Denton. Modelling community structure and species cooccurrence using fishery observer data. ICES Journal of Marine Science, 73(7):1750–1763, July 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
7/1750/2458706.

Punt:2018:CRV

[PMBM18] André E. Punt, Paula Moreno, John R. Brandon, and Michael A. Mathews. Conserving and recovering vulnerable marine species: a comprehensive evaluation of the US approach for marine mammals. ICES Journal of Marine Science, 75(5):1813–1831, September 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/5/1813/5035891.

Peterson:2013:KWO

[PMH^{+}13] Megan J. Peterson, Franz Mueter, Dana Hanselman, Chris Lunsford, Craig Matkin, and Holly Fearnbach. Killer whale (Orcinus orca) depredation effects on catch rates of six groundfish species: implications for commercial longline fisheries in Alaska. ICES Journal of Marine Science, 70(6):1220–1232, September 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/70/
6/1220/635452.

Plourde:2014:DDT

[PMM^{+}14] Stéphane Plourde, Ian H. McQuinn, Frédéric Maps, Jean-François St-Pierre, Diane Lavoie, and Pierre Joly. Daytime depth and thermal habitat of two sympatric krill species in response to surface salinity variability in the Gulf of St Lawrence, eastern Canada. ICES Journal of Marine Science, 71(2):272–281, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/2/272/777147.

Prigge:2013:MPV

[PMOH13] E. Prigge, L. Marohn, R. Oeberst, and R. Hanel. Model prediction vs. reality — testing the predictions of a European eel (Anguilla anguilla) stock dynamics model against the in situ observation of silver eel escapement in compliance with the European eel regulation. ICES Journal of Marine Science, 70(2):309–318, March 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/70/2/309/796885.

Paradinas:2016:IBF

[PMP^{+}16] Iosu Paradinas, Marcial Marín, Maria Grazia Pennino, Antonio López-Quílez, David Conesa, David Barreda, Maria Gonzalez, and José María Bellido. Identifying the best fishing-suitable areas under the new European discard ban. ICES Journal of Marine Science, 73(10):2479–2487, November 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/10/2479/2647110.

Pecuchet:2015:ILE

[PNC15] Laurène Pécuchet, J. Rasmus Nielsen, and Asbjørn Christensen. Impacts of the local environment on recruitment: a comparative study of North Sea and Baltic Sea fish stocks. ICES Journal of Marine Science, 72(5):1323–1335, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/5/1323/760685.

Peck:2014:FFI

[PNE^{+}14] Myron A. Peck, Stefan Neuenfeldt, Timothy E. Essington, Verena M. Trenkel, Akinori Takasuka, Henrik Gislason, Mark Dickey-Collas, Ken H. Andersen, Lars Ravn-Jonsen, Niels Vestergaard, Sturla F. Kvamsdal, Anna Gårdmark, Jason Link, and Jake C. Rice. Forage fish interactions: a symposium on “Creating the tools for ecosystem-based management of marine resources”. ICES Journal of Marine Science, 71(1):1–4, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/71/1/1/646783.

Paterson:2018:SES

[PNS18] Barbara Paterson, Barbara Neis, and Robert L. Stephenson. A social–ecological study of stock structure and fleet dynamics in the Newfoundland herring fishery. ICES Journal of Marine Science, 75(1):257–269, January 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/1/257/3860037.

Probst:2014:HGA

[PO14] Wolfgang Nikolaus Probst and Daniel Oesterwind. How good are alternative indicators for spawning-stock biomass (SSB) and fishing mortality (F)? ICES Journal of Marine Science, 71(5):1137–1141, July 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/71/
5/1137/645195.

Parada:2012:AGB

[POIM12] J. M. Parada, R. Outeiral, E. Iglesias, and J. Molares. Assessment of goose barnacle (Pollicipes pollicipes Gmelin, 1789) stocks in management plans: design of a sampling program based on the harvesters’ experience. ICES Journal of Marine Science, 69(10):1840–1849, December 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/10/1840/623725.

Portman:2011:MSP

[Por11] Michelle E. Portman. Marine spatial planning: achieving and evaluating integration. ICES Journal of Marine Science, 68(10):2191–2200, November 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/68/
10/2191/615795.

Powers:2014:ASN

[Pow14] Joseph E. Powers. Age-specific natural mortality rates in stock assessments: size-based vs. density-dependent. ICES Journal of Marine Science, 71(7):1629–1637, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/7/1629/664136.

Pascoe:2017:MMM

[PPD17] Sean D. Pascoe, Éva E. Plagányi, and Catherine M. Dichmont. Modelling multiple management objectives in fisheries: Australian experiences. ICES Journal of Marine Science, 74(2):464–474, March 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
2/464/2907920.

Pakhomov:2019:VDA

[PPHK19] Evgeny A. Pakhomov, Yasha Podeswa, Brian P. V. Hunt, and Lian E. Kwong. Vertical distribution and active carbon transport by pelagic decapods in the North Pacific Subtropical Gyre. ICES Journal of Marine Science, 76(3):702– 717, May 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/3/702/5107804.

Parsons:2013:VSS

[PPM13] Miles J. G. Parsons, Iain M. Parnum, and Robert D. McCauley. Visualizing samsonfish (Seriola hippos) with a Reson 7125 Seabat multibeam sonar. ICES Journal of Marine Science, 70(3):665–674, April 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/70/
3/665/916493.

Poulton:2017:CIE

[PPS17] Danielle A. Poulton, Cosima S. Porteus, and Stephen D. Simpson. Combined impacts of elevated CO_{2} and anthropogenic noise on European sea bass (Dicentrarchus labrax). ICES Journal of Marine Science, 74(4):1230–1236, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/4/1230/2907911.

Poos:2010:SSA

[PQR10] J. J. Poos, F. J. Quirijns, and A. D. Rijnsdorp. Spatial segregation among fishing vessels in a multispecies fishery. ICES Journal of Marine Science, 67(1):155–164, January 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/1/155/594920.

Pedersen:2016:YPR

[PR16] Michael I. Pedersen and Gorm H. Rasmussen. Yield per recruit from stocking two different sizes of eel (Anguilla anguilla) in the brackish Roskilde Fjord. ICES Journal of Marine Science, 73(1):158–164, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/1/158/2458789.

Pepin:2015:OUL

[PRB^{+}15] Pierre Pepin, Dominique Robert, Caroline Bouchard, John F. Dower, Marianne Falardeau, Louis Fortier, Gregory P. Jenkins, Véronique Leclerc, Keith Levesque, Joel K. Llopiz, Mark G. Meekan, Hannah M. Murphy, Marc Ringuette, Pascal Sirois, and Su Sponaugle. Once upon a larva: revisiting the relationship between feeding success and growth in fish larvae. ICES Journal of Marine Science, 72(2):359–373, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/2/359/2801506.

Pierre:2018:ACL

[PRBF18] Maud Pierre, Tristan Rouyer, Sylvain Bonhommeau, and J. M. Fromentin. Assessing causal links in fish stock– recruitment relationships. ICES Journal of Marine Science, 75(3):903–911, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/75/3/903/4590187.

Prellezo:2019:EEV

[Pre19] Raúl Prellezo. Exploring the economic viability of a mesopelagic fishery in the Bay of Biscay. ICES Journal of Marine Science, 76(3):771–779, May 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
3/771/4823616.

Pardo:2017:HFI

[PRF^{+}17] Luis M. Pardo, Marcela P. Riveros, Juan Pablo Fuentes, Ramona Pinochet, Carla Cárdenas, and Bernard Sainte-Marie. High fishing intensity reduces females’ sperm reserve and brood fecundity in a eubrachyuran crab subject to sex- and size-biased harvest. ICES Journal of Marine Science, 74(9):2459–2469, November 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/9/2459/3823336.

Perez-Rodriguez:2012:CTD

[PRKASR12] Alfonso Pérez-Rodríguez, Mariano Koen-Alonso, and Fran Saborido-Rey. Changes and trends in the demersal fish community of the Flemish Cap, Northwest Atlantic, in the period 1988–2008. ICES Journal of Marine Science, 69(5): 902–912, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/69/5/902/647250.

Pritt:2014:MDR

[PRO14] Jeremy J. Pritt, Edward F. Roseman, and Timothy P. O’Brien. Mechanisms driving recruitment variability in fish: comparisons between the Laurentian Great Lakes and marine systems. ICES Journal of Marine Science, 71(8):2252– 2267, October 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/71/8/2252/760714.

Polak:2013:EVB

[PS13] Omer Polak and Nadav Shashar. Economic value of biological attributes of artificial coral reefs. ICES Journal of Marine Science, 70(4):904–912, July 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/70/
4/904/724825.

Pascoe:2016:MEL

[PSB16] Sean D. Pascoe, Jesse A. Sharp, and Rik C. Buckworth. Modelling effort levels in a sequential fishery. ICES Journal of Marine Science, 73(2):503–511, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/2/503/2614219.

Pampoulie:2012:GSA

[PSDG12] Christophe Pampoulie, Sigurlaug Skirnisdóttir, Anna Kristín Daníelsdóttir, and Ásgeir Gunnarsson. Genetic structure of the Atlantic wolffish (Anarhichas lupus L.) at Icelandic fishing grounds: another evidence of panmixia in Iceland? ICES Journal of Marine Science, 69(4):508–515, May 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/4/508/633230.

Probst:2012:UCC

[PSF12] Wolfgang Nikolaus Probst, Vanessa Stelzenmüller, and Heino Ove Fock. Using cross-correlations to assess the relationship between time-lagged pressure and state indicators: an exemplary analysis of North Sea fish population indicators. ICES Journal of Marine Science, 69(4):670– 681, May 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/69/4/670/632477.

Punt:2012:EIB

[PSGY^{+}12] André E. Punt, M. S. M. Siddeek, Brian Garber-Yonts, Michael Dalton, Louis Rugolo, Diana Stram, Benjamin J. Turnock, and Jie Zheng. Evaluating the impact of buffers to account for scientific uncertainty when setting TACs: application to red king crab in Bristol Bay, Alaska. ICES Journal of Marine Science, 69(4):624–634, May 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/4/624/636410.

Pampoulie:2011:PGS

[PSH^{+}11] Christophe Pampoulie, Sigurlaug Skirnisdottir, Sigurbjorg Hauksdottir, Kristinn Olafsson, Hrafnkell Eiríksson, Valérie Chosson, Gudmundur O. Hreggvidsson, Gudmundur H. Gunnarsson, and Sigridur Hjorleifsdottir. A pilot genetic study reveals the absence of spatial genetic structure in Norway lobster (Nephrops norvegicus) on fishing grounds in Icelandic waters. ICES Journal of Marine Science, 68(1): 20–25, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/1/20/633073.

Pansch:2013:LDB

[PSH13] Christian Pansch, Peter Schlegel, and Jonathan Havenhand. Larval development of the barnacle Amphibalanus improvisus responds variably but robustly to near-future ocean acidification. ICES Journal of Marine Science, 70 (4):805–811, July 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/70/4/805/729185.

Parsons:2016:CDV

[PSKM^{+}16] Miles J. G. Parsons, Chandra P. Salgado-Kent, Sarah A. Marley, Alexander N. Gavrilov, and Robert D. Mc-Cauley. Characterizing diversity and variation in fish choruses in Darwin Harbour. ICES Journal of Marine Science, 73(8):2058–2074, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
8/2058/2198353.

Petitgas:2010:SCT

[PSM^{+}10] Pierre Petitgas, Dave H. Secor, Ian McQuinn, Geir Huse, and Nancy Lo. Stock collapses and their recovery: mechanisms that establish and maintain life-cycle closure in space and time. ICES Journal of Marine Science, 67(9):1841– 1848, December 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/67/9/1841/619826.

Pampoulie:2014:GSL

[PSO^{+}14] Christophe Pampoulie, Sigurlaug Skirnisdottir, Gudbjorg Olafsdottir, Sarah J. Helyar, Vilhjálmur Thorsteinsson, Sigurdur Th. Jónsson, Alain Fréchet, Caroline M. F. Durif, Sally Sherman, Magdalena Lampart-Kałużniacka, Rasmus Hedeholm, Halldór Ólafsson, Anna K. Daníelsdóttir, and Jacob M. Kasper. Genetic structure of the lumpfish Cyclopterus lumpus across the North Atlantic. ICES Journal of Marine Science, 71(9):2390–2397, November 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/9/2390/594774.

Punt:2011:ASC

[PSS11] André E. Punt, David C. Smith, and Anthony D. M. Smith. Among-stock comparisons for improving stock assessments of data-poor stocks: the “Robin Hood” approach. ICES Journal of Marine Science, 68(5):972–981, May 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/68/5/972/653125.

Punt:2014:SRA

[PSS^{+}14] André E. Punt, Anthony D. M. Smith, David C. Smith, Geoffrey N. Tuck, and Neil L. Klaer. Selecting relative abundance proxies for B_{MSY} and B_{MEY}. ICES Journal of Marine Science, 71(3):469–483, March 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/3/469/636159.

Pursche:2014:ETS

[PST14] Alexander R. Pursche, Iain M. Suthers, and Matthew D. Taylor. The effect of targeted stocking on behaviour and space utilization of a released finfish. ICES Journal of Marine Science, 71(5):1100–1106, July 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/71/
5/1100/645740.

Prokopchuk:2019:IDZ

[PT19] Irina P Prokopchuk and Alexander G Trofimov. Interannual dynamics of zooplankton in the Kola Section of the Barents Sea during the recent warming period. ICES Journal of Marine Science, 76(S1):S10–S23, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/76/Supplement_1/i10/5680377.

Poos:2013:ARB

[PTvOR13] Jan Jaap Poos, Michael N. J. Turenhout, Hans A. E. van Oostenbrugge, and Adriaan D. Rijnsdorp. Adaptive response of beam trawl fishers to rising fuel cost. ICES Journal of Marine Science, 70(3):675–684, April 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/3/675/914693.

Punt:2017:SMD

[Pun17] André E. Punt. Strategic management decision-making in a complex world: quantifying, understanding, and using trade-offs. ICES Journal of Marine Science, 74(2):499–510, March 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/2/499/2907904.

Purcell:2018:JFH

[Pur18] Jennifer E. Purcell. Of jellyfish, fish, and humans. ICES Journal of Marine Science, 75(4):1235–1244, July 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/4/1235/4794717.

Pinder:2017:CCR

[PVCB17] Adrian C. Pinder, Randolph Velterop, Steven J. Cooke, and J. Robert Britton. Consequences of catch-and-release angling for black bream Spondyliosoma cantharus, during the parental care period: implications for management. ICES Journal of Marine Science, 74(1):254–262, January 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/1/254/2669567.

Persson:2014:EFE

[PVD14] Lennart Persson, Anieke Van Leeuwen, and André M. De Roos. The ecological foundation for ecosystem-based management of fisheries: mechanistic linkages between the individual-, population-, and community-level dynamics. ICES Journal of Marine Science, 71(8):2268–2280, October 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/8/2268/751624.

Pliru:2012:SFB

[PvdKE^{+}12] Antonio Plirú, Jeroen van der Kooij, Georg H. Engelhard, Clive J. Fox, Stephen P. Milligan, and Ewan Hunter. Sprat feeding behaviour, selective predation, and impact on plaice egg mortality. ICES Journal of Marine Science, 69(6): 1019–1029, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/69/6/1019/621583.

Piet:2010:NRI

[PvOP10] G. J. Piet, H. M. J. van Overzee, and M. A. Pastoors. The necessity for response indicators in fisheries management. ICES Journal of Marine Science, 67(3):559–566, April 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/67/3/559/734210.

Papaioannou:2012:DUS

[PVQS12] Eva A. Papaioannou, Athanasios T. Vafeidis, Martin F. Quaas, and Jörn O. Schmidt. The development and use of a spatial database for the determination and characterization of the state of the German Baltic small-scale fishery sector. ICES Journal of Marine Science, 69(8):1480– 1490, September 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/69/8/1480/699568.

Page:2017:EEP

[PWCS17] Tessa M. Page, Samantha Worthington, Piero Calosi, and Jonathon H. Stillman. Effects of elevated pCO_{2} on crab survival and exoskeleton composition depend on shell function and species distribution: a comparative analysis of carapace and claw mineralogy across four porcelain crab species from different habitats. ICES Journal of Marine Science, 74(4): 1021–1032, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/74/4/1021/2631175.

Pine:2018:ISV

[PWPW18] Matthew K. Pine, Ding Wang, Lindsay Porter, and Kexiong Wang. Investigating the spatiotemporal variation of fish choruses to help identify important foraging habitat for Indo-Pacific humpback dolphins, Sousa chinensis. ICES Journal of Marine Science, 75(2):510–518, March 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/2/510/4575129.

Plaganyi:2011:AAC

[PWS^{+}11] Éva E. Plagányi, Scarla J. Weeks, Tim D. Skewes, Mark T. Gibbs, Elvira S. Poloczanska, Ana Norman-López, Laura K. Blamey, Muri Soares, and William M. L. Robinson. Assessing the adequacy of current fisheries management under changing climate: a southern synopsis. ICES Journal of Marine Science, 68(6):1305–1317, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/6/1305/710425.

Quinones:2018:SPLa

[QCA^{+}18a] Javier Quiñones, Luciano M. Chiaverano, Patricia Ayón, Grant D. Adams, Hermes W. Mianzan, and E. Marcelo Acha. Spatial patterns of large jellyfish Chrysaora plocamia blooms in the Northern Humboldt Upwelling System in relation to biological drivers and climate. ICES Journal of Marine Science, 75(4):1405–1415, July 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/4/1405/4840650.

Quinones:2018:SPLb

[QCA^{+}18b] Javier Quiñones, Luciano M. Chiaverano, Patricia Ayón, Grant D. Adams, Hermes W. Mianzan, and E. Marcelo Acha. Spatial patterns of large jellyfish Chrysaora plocamia blooms in the Northern Humboldt Upwelling System in relation to biological drivers and climate. ICES Journal of Marine Science, 75(4):1510, July 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/4/1510/4961486.

Quintino:2010:RSU

[QFM^{+}10] Victor Quintino, Rosa Freitas, Renato Mamede, Fernando Ricardo, Ana Maria Rodrigues, Jorge Mota, Ángel Pérez-Ruzafa, and Concepción Marcos. Remote sensing of underwater vegetation using single-beam acoustics. ICES Journal of Marine Science, 67(3):594–605, April 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/3/594/731557.

Quillerou:2012:WBF

[QG12] Emmanuelle Quillérou and Olivier Guyader. What is behind fleet evolution: a framework for flow analysis and application to the French Atlantic fleet. ICES Journal of Marine Science, 69(6):1069–1077, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/6/1069/619081.

Queirolo:2012:EEC

[QHG^{+}12] Dante Queirolo, Carlos F. Hurtado, Erick Gaete, Milagrosa C. Soriguer, Karim Erzini, and Juan C. Gutiérrez-Estrada. Effects of environmental conditions and fishing operations on the performance of a bottom trawl. ICES Journal of Marine Science, 69(2):293–302, March 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/2/293/702485.

Quetglas:2013:SCE

[QOH^{+}13] Antoni Quetglas, Francesc Ordines, Manuel Hidalgo, Sebastià Monserrat, Susana Ruiz, Ángel Amores, Joan Moranta, and Enric Massutí. Synchronous combined effects of fishing and climate within a demersal community. ICES Journal of Marine Science, 70(2):319–328, March 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/2/319/796397.

Quinn:2015:PEV

[QR15] Brady K. Quinn and Rémy Rochette. Potential effect of variation in water temperature on development time of American lobster larvae. ICES Journal of Marine Science, 72(S1):S79–S90, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/72/
suppl_1/i79/618195.

Quinn:2018:MBC

[Qui18] Thomas P. Quinn. From magnets to bears: is a career studying salmon narrow or broad? ICES Journal of Marine Science, 75(5):1546–1552, September 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/5/1546/4994355.

Qin:2017:SDR

[QZH^{+}17] Geng Qin, Yanhong Zhang, Adeljean L. F. C. Ho, Yuan Zhang, and Qiang Lin. Seasonal distribution and reproductive strategy of seahorses. ICES Journal of Marine Science, 74(8):2170–2179, September 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/8/2170/3192369.

Russell:2012:IFE

[RAB^{+}12] Ian C. Russell, Miran W. Aprahamian, Jon Barry, Ian C. Davidson, Peder Fiske, Anton T. Ibbotson, Richard J. Kennedy, Julian C. Maclean, Andrew Moore, Jaime Otero, Ted (E. C. E.) Potter, and Christopher D. Todd. The influence of the freshwater environment and the biological characteristics of Atlantic salmon smolts on their subsequent marine survival. ICES Journal of Marine Science, 69(9):1563–1573, November 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/69/
9/1563/634513.

Reum:2016:IDO

[RAH^{+}16] Jonathan C. P. Reum, Simone R. Alin, Chris J. Harvey, Nina Bednaršek, Wiley Evans, Richard A. Feely, Burke Hales, Noelle Lucey, Jeremy T. Mathis, Paul McElhany, Jan Newton, and Christopher L. Sabine. Interpretation and design of ocean acidification experiments in upwelling systems in the context of carbonate chemistry co-variation with temperature and oxygen. ICES Journal of Marine Science, 73(3):582–595, February 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/3/582/2457875.

Rodhouse:2010:LAP

[RB10] Paul G. Rodhouse and Peter R. Boyle. Large aggregations of pelagic squid near the ocean surface at the Antarctic Polar Front, and their capture by grey-headed albatrosses. ICES Journal of Marine Science, 67(7):1432–1435, October 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/7/1432/663941.

Rice:2014:WAR

[RB14] Jake Rice and Howard I. Browman. Where has all the recruitment research gone, long time passing? ICES Journal of Marine Science, 71(8):2293–2299, October 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/8/2293/2804458.

Reglero:2019:PHO

[RBA^{+}19] Patricia Reglero, Rosa Balbín, Franciso Javier Abascal, Antonio Medina, Diego Alvarez-Berastegui, Leif Rasmuson, Baptiste Mourre, Sámar Saber, Aurelio Ortega, Edurne Blanco, Fernando de la Gándara, Franciso Javier Alemany, G. Walter Ingram, and Manuel Hidalgo. Pelagic habitat and offspring survival in the eastern stock of Atlantic bluefin tuna. ICES Journal of Marine Science, 76(2):549– 558, March 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/76/2/549/5115365.

Raybaud:2017:FCD

[RBAB17] Virginie Raybaud, Mahmoud Bacha, Rachid Amara, and Grégory Beaugrand. Forecasting climate-driven changes in the geographical range of the European anchovy (Engraulis encrasicolus). ICES Journal of Marine Science, 74(5): 1288–1299, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/74/5/1288/2997453.

Reiss:2015:BDM

[RBB^{+}15] Henning Reiss, Silvana Birchenough, Angel Borja, Lene Buhl-Mortensen, Johan Craeymeersch, Jennifer Dannheim, Alexander Darr, Ibon Galparsoro, Mayya Gogina, Hermann Neumann, Jacques Populus, Anna M. Rengstorf, Mireia Valle, Gert van Hoey, Michael L. Zettler, and Steven Degraer. Benthos distribution modelling and its relevance for marine ecosystem management. ICES Journal of Marine Science, 72(2):297–315, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/2/297/674105.

Rijnsdorp:2016:TFQ

[RBB^{+}16] A. D. Rijnsdorp, F. Bastardie, S. G. Bolam, L. Buhl-Mortensen, O. R. Eigaard, K. G. Hamon, J. G. Hiddink, N. T. Hintzen, A. Ivanović, A. Kenny, P. Laffargue, J. R. Nielsen, F. G. O’Neill, G. J. Piet, H. Polet, A. Sala, C. Smith, P. D. van Denderen, T. van Kooten, and M. Zengin. Towards a framework for the quantitative assessment of trawling impact on the seabed and benthic ecosystem. ICES Journal of Marine Science, 73(suppl 1):S127–S138, January 2016. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/73/suppl_1/i127/2573996.

Roberts:2011:UME

[RBBC11] Lachlan W. Roberts, Paul A. Butcher, Matt K. Broadhurst, and Brian R. Cullis. Using a multi-experimental approach to assess the fate of angled-and-released yellowtail kingfish (Seriola lalandi). ICES Journal of Marine Science, 68(1): 67–75, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/1/67/630886.

Ruiz:2015:EMT

[RBC^{+}15] J. Ruiz, A. Batty, P. Chavance, H. McElderry, V. Restrepo, P. Sharples, J. Santos, and A. Urtizberea. Electronic monitoring trials on in the tropical tuna purse-seine fishery. ICES Journal of Marine Science, 72(4):1201–1213, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/4/1201/801387.

Ricard:2010:EGS

[RBCH10] Daniel Ricard, Robert M. Branton, Donald W. Clark, and Peter Hurley. Extracting groundfish survey indices from the Ocean Biogeographic Information System (OBIS): an example from Fisheries and Oceans Canada. ICES Journal of Marine Science, 67(4):638–645, May 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/4/638/680186.

Rabearisoa:2015:AIB

[RBM15] Njaratiana Rabearisoa, Pascal Bach, and Francis Marsac. Assessing interactions between dolphins and small pelagic fish on branchline to design a depredation mitigation device in pelagic longline fisheries. ICES Journal of Marine Science, 72(5):1682–1690, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/72/
5/1682/770977.

Rodriguez-Barreras:2018:ASL

[RBMAOCL18] Ruber Rodríguez-Barreras, Alfredo Montañez-Acuña, Abimarie Otaño-Cruz, and Scott D. Ling. Apparent stability of a low-density Diadema antillarum regime for Puerto Rican coral reefs. ICES Journal of Marine Science, 75(6):2193– 2201, November 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/6/2193/5055456.

Robinson:2015:QPI

[RBP15] William M. L. Robinson, Douglas S. Butterworth, and Éva E. Plagányi. Quantifying the projected impact of the South African sardine fishery on the Robben Island penguin colony. ICES Journal of Marine Science, 72(6):1822– 1833, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/6/1822/919990.

Reid:2012:SHI

[RC12] Jenny E. Reid and Gérald Chaput. Spawning history influence on fecundity, egg size, and egg survival of Atlantic salmon (Salmo salar) from the Miramichi River, New Brunswick, Canada. ICES Journal of Marine Science, 69(9):1678–1685, November 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/69/
9/1678/639366.

Richards:2015:UWS

[RC15] Russell G. Richards and Milani Chaloupka. Using a weightstructured oyster population dynamic model to explore topdown control of coastal water quality in a subtropical embayment. ICES Journal of Marine Science, 72(2):403–413, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/2/403/2801302.

Rochet:2014:BDI

[RCC14a] Marie-Joëlle Rochet, Tom Catchpole, and Steve Cadrin. Bycatch and discards: from improved knowledge to mitigation programmes. ICES Journal of Marine Science, 71(5): 1216–1218, July 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/71/5/1216/648191.

Rowley:2014:PIC

[RCC^{+}14b] Andrew F. Rowley, Maud E. Cross, Sarah C. Culloty, Sharon A. Lynch, Clara L. Mackenzie, Emer Morgan, Ruth M. O’Riordan, Peter E. Robins, Amanda L. Smith, Tara J. Thrupp, Claire L. Vogan, Emma C. Wootton, and Shelagh K. Malham. The potential impact of climate change on the infectious diseases of commercially important shellfish populations in the Irish Sea — a review. ICES Journal of Marine Science, 71(4):741–759, May 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/4/741/669736.

Ruhl:2017:LTE

[RCF^{+}17] Saskia Rühl, Piero Calosi, Sarah Faulwetter, Kleoniki Keklikoglou, Stephen Widdicombe, and Ana M. Queirós. Longterm exposure to elevated pCO_{2} more than warming modifies early-life shell growth in a temperate gastropod. ICES Journal of Marine Science, 74(4):1113–1124, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/4/1113/3059862.

Roach:2018:ETE

[RCF^{+}18] Michael Roach, Mike Cohen, Rodney Forster, Andrew S. Revill, and Magnus Johnson. The effects of temporary exclusion of activity due to wind farm construction on a lobster (Homarus gammarus) fishery suggests a potential management approach. ICES Journal of Marine Science, 75(4): 1416–1426, July 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/4/1416/4841920.

Rastrick:2018:FPM

[RCG^{+}18] Samuel P. S. Rastrick, Victoria Collier, Helen Graham, Tore Strohmeier, Nia M. Whiteley, and Øivind Strand. Feeding plasticity more than metabolic rate drives the productivity of economically important filter feeders in response to elevated CO_{2} and reduced salinity. ICES Journal of Marine Science, 75(6):2117–2128, November 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/6/2117/5059522.

Ratz:2015:ARP

[RCH^{+}15] Hans-Joachim Rätz, John Casey, Steven J. Holmes, Josep Lloret, Hendrik Dörner, Nikolaos Mitrakis, and Aymen Charef. An alternative reference point in the context of ecosystem-based fisheries management: maximum sustainable dead biomass. ICES Journal of Marine Science, 72(8):2257–2268, September 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/72/
8/2257/2458743.

Rypina:2019:ISS

[RCH^{+}19] Irina I. Rypina, Ke Chen, Christina M. Hernández, Lawrence J. Pratt, and Joel K. Llopiz. Investigating the suitability of the Slope Sea for Atlantic bluefin tuna spawning using a high-resolution ocean circulation model. ICES Journal of Marine Science, 76(6):1666–1677, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/6/1666/5487747.

Rindorf:2017:FMU

[RCS^{+}17] Anna Rindorf, Massimiliano Cardinale, Samuel Shephard, José A. A. De Oliveira, Einar Hjorleifsson, Alexander Kempf, Anna Luzenczyk, Colin Millar, David C. M. Miller, Coby L. Needle, John Simmonds, and Morten Vinther. Fishing for MSY: using “pretty good yield” ranges without impairing recruitment. ICES Journal of Marine Science, 74(2):525–534, March 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
2/525/2669544.

Rice:2014:MFF

[RD14] Jake Rice and Daniel Duplisea. Management of fisheries on forage species: the test-bed for ecosystem approaches to fisheries. ICES Journal of Marine Science, 71(1):143–152, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/71/1/143/645606.

Renfree:2016:OTI

[RD16] Josiah S. Renfree and David A. Demer. Optimizing transmit interval and logging range while avoiding aliased seabed echoes. ICES Journal of Marine Science, 73(8):1955– 1964, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/73/8/1955/2198388.

Randon:2018:QEA

[RDB^{+}18a] Marine Randon, Françoise Daverat, Gilles Bareille, Philippe Jatteau, Jean Martin, Christophe Pecheyran, and Hilaire Drouineau. Quantifying exchanges of Allis shads between river catchments by combining otolith microchemistry and abundance indices in a Bayesian model. ICES Journal of Marine Science, 75(1):9–21, January 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/1/9/4093614.

Renaud:2018:PFW

[RDB^{+}18b] Paul E. Renaud, Malin Daase, Neil S. Banas, Tove M. Gabrielsen, Janne E. Søreide, Øystein Varpe, Finlo Cottier, Stig Falk-Petersen, Claudia Halsband, Daniel Vogedes, Kristin Heggland, and Jørgen Berge. Pelagic food-webs in a changing Arctic: a trait-based perspective suggests a mode of resilience. ICES Journal of Marine Science, 75(6):1871– 1881, November 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/6/1871/5046035.

Rockmann:2011:RHE

[RDCPvH11] Christine Röckmann, Mark Dickey-Collas, Mark R. Payne, and Ralf van Hal. Realized habitats of early-stage North Sea herring: looking for signals of environmental change. ICES Journal of Marine Science, 68(3):537–546, March 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/3/537/656561.

Reiss:2010:SPI

[RDD^{+}10] Henning Reiss, Steven Degraer, Gerard C. A. Duineveld, Ingrid Kröncke, John Aldridge, Johan A. Craeymeersch, Jacqueline D. Eggleton, Hans Hillewaert, Marc S. S. Lavaleye, Andreas Moll, Thomas Pohlmann, Eike Rachor, Mike Robertson, Edward Vanden Berghe, Gert van Hoey, and Hubert L. Rees. Spatial patterns of infauna, epifauna, and demersal fish communities in the North Sea. ICES Journal of Marine Science, 67(2):278–293, March 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/2/278/695848.

Rice:2013:DFR

[RDGP13] Jake Rice, Niels Daan, Henrik Gislason, and John Pope. Does functional redundancy stabilize fish communities? ICES Journal of Marine Science, 70(4):734–742, July 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/4/734/728379.

Ryan:2015:RBD

[RDKK15] Tim E. Ryan, Ryan A. Downie, Rudy J. Kloser, and Gordon Keith. Reducing bias due to noise and attenuation in open-ocean echo integration data. ICES Journal of Marine Science, 72(8):2482–2493, September 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/8/2482/2459093.

Rindorf:2017:FTP

[RDL^{+}17] Anna Rindorf, Catherine Mary Dichmont, Phillip S. Levin, Pamela Mace, Sean Pascoe, Raul Prellezo, André E. Punt, David G. Reid, Robert Stephenson, Clara Ulrich, Morten Vinther, and Lotte Worsøe Clausen. Food for thought: pretty good multispecies yield. ICES Journal of Marine Science, 74(2):475–486, March 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/2/475/2907929.

Roberts:2012:RIS

[RDS12] Michael J. Roberts, Nicola J. Downey, and Warwick H. Sauer. The relative importance of shallow and deep shelf spawning habitats for the South African chokka squid (Loligo reynaudii). ICES Journal of Marine Science, 69 (4):563–571, May 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/69/4/563/634083.

Rindorf:2017:IEE

[RDT^{+}17] Anna Rindorf, Catherine M. Dichmont, James Thorson, Anthony Charles, Lotte Worsøe Clausen, Poul Degnbol, Dorleta Garcia, Niels T. Hintzen, Alexander Kempf, Phillip Levin, Pamela Mace, Christos Maravelias, Coilín Minto, John Mumford, Sean Pascoe, Raul Prellezo, André E. Punt, David G. Reid, Christine Röckmann, Robert L. Stephenson, Olivier Thebaud, George Tserpes, and Rüdiger Voss. Inclusion of ecological, economic, social, and institutional considerations when setting targets and limits for multispecies fisheries. ICES Journal of Marine Science, 74(2): 453–463, March 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/74/2/453/2962394.

Rosen:2012:RSA

[REFJ12] Shale Rosen, Arill Engås, Anders Fernö, and Terje Jörgensen. The reactions of shoaling adult cod to a pelagic trawl: implications for commercial trawling. ICES Journal of Marine Science, 69(2):303–312, March 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/2/303/699393.

Rey:2012:DSC

[Rey12] Francisco Rey. Declining silicate concentrations in the Norwegian and Barents Seas. ICES Journal of Marine Science, 69(2):208–212, March 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/69/2/208/706643.

Reilly:2014:IVF

[RFF^{+}14] T. O. M. Reilly, H. M. Fraser, R. J. Fryer, J. Clarke, and S. P. R. Greenstreet. Interpreting variation in fishbased food web indicators: the importance of “bottomup limitation” and “top-down control” processes. ICES Journal of Marine Science, 71(2):406–416, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/2/406/784513.

Reis-Filho:2019:ISS

[RFHG19] José Amorim Reis-Filho, Euan S. Harvey, and Tommaso Giarrizzo. Impacts of small-scale fisheries on mangrove fish assemblages. ICES Journal of Marine Science, 76(1):153– 164, January 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/76/1/153/5106384.

Rouyer:2014:CEE

[RFHS14] Tristan Rouyer, Jean-Marc Fromentin, Manuel Hidalgo, and Nils C. Stenseth. Combined effects of exploitation and temperature on fish stocks in the Northeast Atlantic. ICES Journal of Marine Science, 71(7):1554–1562, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/7/1554/673272.

Robert:2010:BSS

[RFMR10] Marianne Robert, Abdelmalek Faraj, Murdoch K. McAllister, and Etienne Rivot. Bayesian state-space modelling of the De Lury depletion model: strengths and limitations of the method, and application to the Moroccan octopus fishery. ICES Journal of Marine Science, 67(6):1272– 1290, September 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/67/6/1272/735347.

Ramajo:2019:PRJ

[RFN^{+}19] Laura Ramajo, Carolina Fernández, Yolanda Núñez, Paz Caballero, Marco A. Lardies, and María Josefina Poupin. Physiological responses of juvenile Chilean scallops (Argopecten purpuratus) to isolated and combined environmental drivers of coastal upwelling. ICES Journal of Marine Science, 76(6):1836–1849, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
6/1836/5487735.

Runge:2016:ECC

[RFT^{+}16] Jeffrey A. Runge, David M. Fields, Cameron R. S. Thompson, Steven D. Shema, Reidun M. Bjelland, Caroline M. F. Durif, Anne Berit Skiftesvik, and Howard I. Browman. End of the century CO_{2} concentrations do not have a negative effect on vital rates of Calanus finmarchicus, an ecologically critical planktonic species in North Atlantic ecosystems. ICES Journal of Marine Science, 73(3):937–950, February 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/3/937/2459142.

Rice:2011:FFS

[RG11] Jake C. Rice and Serge M. Garcia. Fisheries, food security, climate change, and biodiversity: characteristics of the sector and perspectives on emerging issues. ICES Journal of Marine Science, 68(6):1343–1353, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/6/1343/704758.

Rastrick:2018:UNA

[RGAS^{+}18] Samuel S. P. Rastrick, Helen Graham, Kumiko Azetsu-Scott, Piero Calosi, Melissa Chierici, Agneta Fransson, Haakon Hop, Jason Hall-Spencer, Marco Milazzo, Peter Thor, and Tina Kutti. Using natural analogues to investigate the effects of climate change and ocean acidification on northern ecosystems. ICES Journal of Marine Science, 75(7):2299–2311, December 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/75/
7/2299/5133274.

Richards:2019:TEM

[RGC^{+}19] Travis M. Richards, Emily E. Gipson, April Cook, Tracey T. Sutton, and R. J. David Wells. Trophic ecology of meso- and bathypelagic predatory fishes in the Gulf of Mexico. ICES Journal of Marine Science, 76(3):662– 672, May 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/3/662/5048417.

Respondek:2014:VFE

[RGFT14] G. Respondek, J. Gröger, J. Floeter, and A. Temming. Variability of fishing effort for the German brown shrimp (Crangon crangon) fishing fleet: influencing factors, and seasonal and spatial patterns. ICES Journal of Marine Science, 71(7):1805–1817, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/7/1805/671080.

Rivera:2016:ILM

[RGGFA16] Antonella Rivera, Stefan Gelcich, Lucía García-Flórez, and José Luis Acuña. Incorporating landscape metrics into invertebrate fisheries management: case study of the gooseneck barnacle in Asturias (N. Spain). ICES Journal of Marine Science, 73(6):1570–1578, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/6/1570/2459092.

Roura:2010:MAI

[RGPG10] Álvaro Roura, Ángel F. González, Santiago Pascual, and Ángel Guerra. A molecular approach to identifying the prey of cephalopod paralarvae. ICES Journal of Marine Science, 67(7):1408–1412, October 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/7/1408/663135.

Reid:2011:DBB

[RGR^{+}11] D. G. Reid, N. Graham, D. J. Rihan, E. Kelly, I. R. Gatt, F. Griffin, H. D. Gerritsen, and R. J. Kynoch. Do big boats tow big nets? ICES Journal of Marine Science, 68(8):1663– 1669, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/8/1663/761876.

Rihan:2011:PFI

[RGRL11] Dominic Rihan, Norman Graham, Dave G. Reid, and Colm Lordan. The provision of fishery information by ICES–WGFTFB to assessment working groups: use of information and lessons learned. ICES Journal of Marine Science, 68(8):1809–1814, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/68/
8/1809/756573.

Reid:2016:IBH

[RGS^{+}16] David G. Reid, Norman Graham, Petri Suuronen, Pingguo He, and Michael Pol. Implementing balanced harvesting: practical challenges and other implications. ICES Journal of Marine Science, 73(6):1690–1696, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/6/1690/2458739.

Richerson:2017:QPR

[RH17] Kate Richerson and Daniel S. Holland. Quantifying and predicting responses to a US West Coast salmon fishery closure. ICES Journal of Marine Science, 74(9):2364–2378, November 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/9/2364/3884461.

Reddin:2012:DBC

[RHB^{+}12] David G. Reddin, Lars Petter Hansen, Vegar Bakkestuen, Ian Russell, Jonathan White, E. C. E. (Ted) Potter, J. Brian Dempson, Timothy F. Sheehan, Niall Ó Maoiléidigh, Gordon W. Smith, Arni Isaksson, Jan Arge Jacobsen, Mark Fowler, Kjell Arne Mork, and Peter Amiro. Distribution and biological characteristics of Atlantic salmon (Salmo salar) at Greenland based on the analysis of historical tag recoveries. ICES Journal of Marine Science, 69(9):1589–1597, November 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/69/
9/1589/638274.

Reddin:2013:CDB

[RHB^{+}13] David G. Reddin, Lars Petter Hansen, Vegar Bakkestuen, Ian Russell, Jonathan White, E. C. E. (Ted) Potter, J. Brian Dempson, Timothy F. Sheehan, Niall Ó Maoiléidigh, Gordon W. Smith, Arni Isaksson, Jan Arge Jacobsen, Mark Fowler, Kjell Arne Mork, and Peter Amiro. A correction to “Distribution and biological characteristics of Atlantic salmon (Salmo salar) at Greenland based on the analysis of historical tag recoveries”. ICES Journal of Marine Science, 70(4):914, July 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/70/4/914/728370.

Rehberg-Haas:2012:OMA

[RHHH^{+}12] Sabine Rehberg-Haas, Cornelius Hammer, Nicola Hillgruber, Karin Hüssy, and Axel Temming. Otolith microstructure analysis to resolve seasonal patterns of hatching and settlement in western Baltic cod. ICES Journal of Marine Science, 69(8):1347–1356, September 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/8/1347/702565.

Richardson:2010:DLT

[RHOJ10] David E. Richardson, Jonathan A. Hare, William J. Overholtz, and Donna L. Johnson. Development of long-term larval indices for Atlantic herring (Clupea harengus) on the northeast US continental shelf. ICES Journal of Marine Science, 67(4):617–627, May 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/67/4/617/680402.

Riascos:2011:ESP

[RHOL11] José M. Riascos, Olaf Heilmayer, Marcelo E. Oliva, and Jürgen Laudien. Environmental stress and parasitism as drivers of population dynamics of Mesodesma donacium at its northern biogeographic range. ICES Journal of Marine Science, 68(5):823–833, May 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/68/5/823/650568.

Rogers:2012:QCD

[RHP^{+}12] Paul J. Rogers, Charlie Huveneers, Brad Page, Derek J. Hamer, Simon D. Goldsworthy, James G. Mitchell, and Laurent Seuront. A quantitative comparison of the diets of sympatric pelagic sharks in gulf and shelf ecosystems off southern Australia. ICES Journal of Marine Science, 69(8):1382–1393, September 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/69/
8/1382/700401.

Reuchlin-Hugenholtz:2016:SRP

[RHSH16] Emilie Reuchlin-Hugenholtz, Nancy L. Shackell, and Jeffrey A. Hutchings. Spatial reference points for groundfish. ICES Journal of Marine Science, 73(10):2468–2478, November 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/10/2468/2647118.

Rajasilta:2014:DSI

[RHV14] M. Rajasilta, J. Hänninen, and I. Vuorinen. Decreasing salinity improves the feeding conditions of the Baltic herring (Clupea harengus membras) during spring in the Bothnian Sea, northern Baltic. ICES Journal of Marine Science, 71(5):1148–1152, July 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/71/
5/1148/648285.

Rossi:2019:CEF

[RIBB^{+}19] Sergio Rossi, Enrique Isla, Mar Bosch-Belmar, Giovanni Galli, Andrea Gori, Michele Gristina, Gianmarco Ingrosso, Giacomo Milisenda, Stefano Piraino, Lucia Rizzo, Nadine Schubert, Marcelo Soares, Cosimo Solidoro, Ruth H. Thurstan, Núria Viladrich, Trevor J. Willis, and Patrizia Ziveri. Changes of energy fluxes in marine animal forests of the Anthropocene: factors shaping the future seascape. ICES Journal of Marine Science, 76(7):2008–2019, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/7/2008/5542774.

Rice:2011:ASF

[Ric11] Jake C. Rice. Advocacy science and fisheries decisionmaking. ICES Journal of Marine Science, 68(10):2007– 2012, November 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/10/2007/614605.

Rice:2014:EIC

[Ric14] Jake Rice. Evolution of international commitments for fisheries sustainability. ICES Journal of Marine Science, 71(2): 157–165, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/71/2/157/777867.

Rice:2019:WLS

[Ric19] Jake Rice. What a long, strange trip it’s been. ICES Journal of Marine Science, 76(7):1973–1982, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/7/1973/5528146.

Ridings:2018:RES

[Rid18] Penelope Ridings. Redefining environmental stewardship to deliver governance frameworks for marine biodiversity beyond national jurisdiction. ICES Journal of Marine Science, 75(1):435–443, January 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/75/
1/435/3953910.

Rothschild:2012:CVN

[RJ12] Brian J. Rothschild and Yue Jiao. Characterizing variation in Northwest Atlantic fish-stock abundance. ICES Journal of Marine Science, 69(5):913–922, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/5/913/646821.

Robinson:2017:EIF

[RJP^{+}17] Orin J. Robinson, Olaf P. Jensen, Mikaela M. Provost, Shuochen Huang, Nina H. Fefferman, Amira Kebir, and Julie L. Lockwood. Evaluating the impacts of fishing on sex-changing fish: a game-theoretic approach. ICES Journal of Marine Science, 74(3):652–659, March 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/3/652/2690095.

Ryan:2016:IEO

[RK16] Tim E. Ryan and Rudy J. Kloser. Improved estimates of orange roughy biomass using an acoustic-optical system in commercial trawlnets. ICES Journal of Marine Science, 73(8):2112–2124, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
8/2112/2198257.

Rudels:2012:EGC

[RKB^{+}12] Bert Rudels, Meri Korhonen, Gereon Budéus, Agnieszka Beszczynska-Möller, Ursula Schauer, Aleksi Nummelin, Detlef Quadfasel, and Hédinn Valdimarsson. The East Greenland Current and its impacts on the Nordic Seas: observed trends in the past decade. ICES Journal of Marine Science, 69(5):841–851, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/69/5/841/653841.

Rouse:2018:CFI

[RKCP18] Sally Rouse, Andronikos Kafas, Rui Catarino, and Hayes Peter. Commercial fisheries interactions with oil and gas pipelines in the North Sea: considerations for decommissioning. ICES Journal of Marine Science, 75(1):279–286, January 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/1/279/3972175.

Rothschild:2014:FEO

[RKJ14] Brian J. Rothschild, Emily F. Keiley, and Yue Jiao. Failure to eliminate overfishing and attain optimum yield in the New England groundfish fishery. ICES Journal of Marine Science, 71(2):226–233, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/2/226/780206.

Raid:2010:RDG

[RKL^{+}10] Tiit Raid, Georgs Kornilovs, Ain Lankov, Anne-Marin Nisumaa, Heli Shpilev, and Ahto Järvik. Recruitment dynamics of the Gulf of Riga herring stock: densitydependent and environmental effects. ICES Journal of Marine Science, 67(9):1914–1920, December 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/67/
9/1914/623166.

Robbins:2017:IRP

[RKW^{+}17] L. L. Robbins, P. O. Knorr, J. G. Wynn, P. Hallock, and P. J. Harries. Interpreting the role of pH on stable isotopes in large benthic foraminifera. ICES Journal of Marine Science, 74(4):955–964, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/4/955/2907923.

Rougier:2012:CAS

[RLD^{+}12] Thibaud Rougier, Patrick Lambert, Hilaire Drouineau, Michel Girardin, Gérard Castelnaud, Laurent Carry, Miran Aprahamian, Etienne Rivot, and Eric Rochard. Collapse of allis shad, Alosa alosa, in the Gironde system (southwest France): environmental change, fishing mortality, or Allee effect? ICES Journal of Marine Science, 69(10):1802–1811, December 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/69/10/1802/623126.

Rousi:2013:LTC

[RLP^{+}13] Heta Rousi, Ari O. Laine, Heikki Peltonen, Pentti Kangas, Ann-Britt Andersin, Jouko Rissanen, Eva Sandberg-Kilpi, and Erik Bonsdorff. Long-term changes in coastal zoobenthos in the northern Baltic Sea: the role of abiotic environmental factors. ICES Journal of Marine Science, 70(2): 440–451, March 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/70/2/440/797473.

Rabaoui:2015:POG

[RLQ^{+}15] Lotfi Rabaoui, Yu-Jia Lin, Mohammad A. Qurban, Rommel H. Maneja, Javier Franco, Thadickal V. Joydas, Premlal Panickan, Khaled Al-Abdulkader, and Rubén H. Roa-Ureta. Patchwork of oil and gas facilities in Saudi waters of the Arabian Gulf has the potential to enhance local fisheries production. ICES Journal of Marine Science, 72(8):2398– 2408, September 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/72/8/2398/2457887.

Rorke:2015:PAL

[RLW^{+}15] Richard ’Rorke, Shane D. Lavery, Miao Wang, Ramón Gallego, Anya M. Waite, Lynnath E. Beckley, Peter A. Thompson, and Andrew G. Jeffs. Phyllosomata associated with large gelatinous zooplankton: hitching rides and stealing bites. ICES Journal of Marine Science, 72(S1):S124–S127, July 2015. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/72/suppl_1/i124/613103.

Reum:2015:QNM

[RMF^{+}15] Jonathan C. P. Reum, P. Sean McDonald, Bridget E. Ferriss, Dara M. Farrell, Chris J. Harvey, and Phillip S.

Levin. Qualitative network models in support of ecosystem approaches to bivalve aquaculture. ICES Journal of Marine Science, 72(8):2278–2288, September 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/8/2278/2459074.

Robert:2014:PTK

[RMJF14] Dominique Robert, Hannah M. Murphy, Gregory P. Jenkins, and Louis Fortier. Poor taxonomical knowledge of larval fish prey preference is impeding our ability to assess the existence of a “critical period” driving year-class strength. ICES Journal of Marine Science, 71(8):2042–2052, October 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/8/2042/747291.

Rotherham:2011:RUA

[RMKG11] Douglas Rotherham, William G. Macbeth, Steven J. Kennelly, and Charles A. Gray. Reducing uncertainty in the assessment and management of fish resources following an environmental impact. ICES Journal of Marine Science, 68(8):1726–1733, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/68/
8/1726/752990.

Rincon:2016:EVE

[RML^{+}16] Margarita María Rincón, John D. Mumford, Polina Levontin, Adrian W. Leach, and Javier Ruiz. The economic value of environmental data: a notional insurance scheme for the European anchovy. ICES Journal of Marine Science, 73(4): 1033–1041, March 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/73/4/1033/2458923.

Reese:2011:EFD

[ROBC11] Douglas C. Reese, Robert T. O’Malley, Richard D. Brodeur, and James H. Churnside. Epipelagic fish distributions in relation to thermal fronts in a coastal upwelling system using high-resolution remote-sensing techniques. ICES Journal of Marine Science, 68(9):1865–1874, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/9/1865/666090.

Rodhouse:2010:EEV

[Rod10] Paul G. Rodhouse. Effects of environmental variability and change on cephalopod populations: an introduction to the CIAC ’09 Symposium special issue. ICES Journal of Marine Science, 67(7):1311–1313, October 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/7/1311/664413.

Roney:2018:TVO

[ROK^{+}18] Nancy E. Roney, Rebekah A. Oomen, Halvor Knutsen, Esben M. Olsen, and Jeffrey A. Hutchings. Temporal variability in offspring quality and individual reproductive output in a broadcast-spawning marine fish. ICES Journal of Marine Science, 75(4):1353–1361, July 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/4/1353/4780189.

Rothschild:2015:BDI

[Rot15] Brian J. Rothschild. On the birth and death of ideas in marine science. ICES Journal of Marine Science, 72(5): 1237–1244, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/72/5/1237/778277.

Ross:2016:TRM

[RPB16] Pauline M. Ross, Laura Parker, and Maria Byrne. Transgenerational responses of molluscs and echinoderms to changing ocean conditions. ICES Journal of Marine Science, 73(3):537–549, February 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
3/537/2459108.

Robert:2014:IGH

[RPDF14] Dominique Robert, Pierre Pepin, John F. Dower, and Louis Fortier. Individual growth history of larval Atlantic mackerel is reflected in daily condition indices. ICES Journal of Marine Science, 71(4):1001–1009, May 2014.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/4/1001/663419.

Record:2014:PPPa

[RPM14a] Nicholas R. Record, Andrew J. Pershing, and Frédéric Maps. The paradox of the “paradox of the plankton”. ICES Journal of Marine Science, 71(2):236–240, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/2/236/777680.

Record:2014:PPPb

[RPM14b] Nicholas R. Record, Andrew J. Pershing, and Frédéric Maps. Plankton post-paradox: reply to comment on “The paradox of the ‘paradox of the plankton”’ by Record et al. ICES Journal of Marine Science, 71(2):296–298, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/71/2/296/790726.

Roy:2011:MTE

[RPS11] Shovonlal Roy, Trevor Platt, and Shubha Sathyendranath. Modelling the time-evolution of phytoplankton size spectra from satellite remote sensing. ICES Journal of Marine Science, 68(4):719–728, March 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/68/4/719/647222.

Rochet:2010:CPF

[RR10] Marie-Joëlle Rochet and Jake C. Rice. Comment on “Purported flaws in management strategy evaluation: basic problems or misinterpretation?” by Butterworth et al. ICES Journal of Marine Science, 67(3):575–576, April 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/3/575/734337.

Ramirez-Rodriguez:2011:DCS

[RR11] Mauricio Ramírez-Rodríguez. Data collection on the small-scale fisheries of México. ICES Journal of Marine Science, 68(8):1611–1614, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/68/
8/1611/753193.

Ribas-Ribas:2017:SPP

[RRCT^{+}17] M. Ribas-Ribas, G. L. Cripps, M. Townend, A. J. Poulton, and T. Tyrrell. Spatial patterns of phytoplankton composition and upper-ocean biogeochemistry do not follow carbonate chemistry gradients in north-west European shelf seas. ICES Journal of Marine Science, 74(4):965– 977, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/4/965/3811715.

Robinson:2015:EEF

[RRH^{+}15] Kelly L. Robinson, James J. Ruzicka, Frank J. Hernandez, William M. Graham, Mary Beth Decker, Richard D. Brodeur, and Malinda Sutor. Evaluating energy flows through jellyfish and gulf menhaden (Brevoortia patronus) and the effects of fishing on the northern Gulf of Mexico ecosystem. ICES Journal of Marine Science, 72(8):2301– 2312, September 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/72/8/2301/2458737.

Roger:2012:CSS

[RRM^{+}12] Liza M. Roger, Anthony J. Richardson, A. David McKinnon, Brenton Knott, Richard Matear, and Cameron Scadding. Comparison of the shell structure of two tropical thecosomata (Creseis acicula and Diacavolinia longirostris) from 1963 to 2009: potential implications of declining aragonite saturation. ICES Journal of Marine Science, 69(3):465–474, May 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/69/3/465/599234.

Rosa:2010:VAP

[RS10] Rui Rosa and Brad A. Seibel. Voyage of the argonauts in the pelagic realm: physiological and behavioural ecology of the rare paper nautilus, Argonauta nouryi. ICES Journal of Marine Science, 67(7):1494–1500, October 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/7/1494/661868.

Ramos:2015:MIA

[RSB^{+}15] Jorge Ramos, Katrine Soma, Øivind Bergh, Torsten Schulze, Antje Gimpel, Vanessa Stelzenmüller, Timo Mäkinen, Gianna Fabi, Fabio Grati, and Jeremy Gault. Multiple interests across European coastal waters: the importance of a common language. ICES Journal of Marine Science, 72(2):720–731, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/72/
2/720/671711.

Radlinski:2013:STD

[RSBC13] Mary K. Radlinski, Miles A. Sundermeyer, James J. Bisagni, and Steven X. Cadrin. Spatial and temporal distribution of Atlantic mackerel (Scomber scombrus) along the northeast coast of the United States, 1985–1999. ICES Journal of Marine Science, 70(6):1151–1161, September 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/6/1151/633708.

Rambo:2017:MFC

[RSGM17] Henrike Rambo, Vanessa Stelzenmüller, Simon P. R. Greenstreet, and Christian Möllmann. Mapping fish community biodiversity for European marine policy requirements. ICES Journal of Marine Science, 74(8):2223–2238, September 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/8/2223/3782701.

Ruesink:2018:SOR

[RST18] J. L. Ruesink, A. Sarich, and A. C. Trimble. Similar oyster reproduction across estuarine regions differing in carbonate chemistry. ICES Journal of Marine Science, 75(1):340–350, January 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/1/340/4061476.

Reed:2017:EIA

[RSV^{+}17] Jodie Reed, Lynne Shannon, Laure Velez, Ekin Akoglu, Alida Bundy, Marta Coll, Caihong Fu, Elizabeth A. Fulton, Arnaud Grüss, Ghassen Halouani, Johanna J. Heymans, Jennifer E. Houle, Emma John, François Le Loc’h, Baris Salihoglu, Philippe Verley, and Yunne-Jai Shin. Ecosystem indicators — accounting for variability in species’ trophic levels. ICES Journal of Marine Science, 74(1):158–169, January 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/1/158/2669566.

Radiarta:2011:ASS

[RSY11] I. Nyoman Radiarta, Sei-Ichi Saitoh, and Hajime Yasui. Aquaculture site selection for Japanese kelp (Laminaria japonica) in southern Hokkaido, Japan, using satellite remote sensing and GIS-based models. ICES Journal of Marine Science, 68(4):773–780, March 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/4/773/645269.

Ryabov:2019:SSS

[RT19] Alexey B. Ryabov and Geraint A. Tarling. Scaling of size, shape and surface roughness in Antarctic krill swarms. ICES Journal of Marine Science, 76(4):1177–1188, July 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/4/1177/5306105.

Romero-Torres:2017:RSS

[RTAT17] Mauricio Romero-Torres, Alberto Acosta, and Eric A. Treml. The regional structure of spawning phenology and the potential consequences for connectivity of coral assemblages across the Eastern Tropical Pacific. ICES Journal of Marine Science, 74(3):613–624, March 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/3/613/2734755.

Rojbek:2014:FFQ

[RTJS14] Maria C. Røjbek, Jonna Tomkiewicz, Charlotte Jacobsen, and Josianne G. Støttrup. Forage fish quality: seasonal lipid dynamics of herring (Clupea harengus L.) and sprat (Sprattus sprattus L.) in the Baltic Sea. ICES Journal of Marine Science, 71(1):56–71, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/1/56/644463.

Rudd:2019:EMD

[RTS19] Merrill B. Rudd, James T. Thorson, and Skyler R. Sagarese. Ensemble models for data-poor assessment: accounting for uncertainty in life-history information. ICES Journal of Marine Science, 76(4):870–883, July 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/4/870/5315645.

Roa-Ureta:2012:MSP

[RU12] Rubén H. Roa-Ureta. Modelling in-season pulses of recruitment and hyperstability-hyperdepletion in the Loligo gahi fishery around the Falkland Islands with generalized depletion models. ICES Journal of Marine Science, 69(8):1403– 1415, September 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/69/8/1403/702402.

Rijnsdorp:2010:IFI

[RvDW10] Adriaan D. Rijnsdorp, Cindy J. G. van Damme, and Peter R. Witthames. Implications of fisheries-induced changes in stock structure and reproductive potential for stock recovery of a sex-dimorphic species, North Sea plaice. ICES Journal of Marine Science, 67(9):1931–1938, December 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/9/1931/619161.

Rooper:2018:VDS

[RWG^{+}18] Christopher N. Rooper, Rachel Wilborn, Pamela Goddard, Kresimir Williams, Richard Towler, and Gerald R. Hoff. Validation of deep-sea coral and sponge distribution models in the Aleutian Islands, Alaska. ICES Journal of Marine Science, 75(1):199–209, January 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/1/199/3855118.

Rosa:2011:EEV

[RYS11] A. L. Rosa, J. Yamamoto, and Y. Sakurai. Effects of environmental variability on the spawning areas, catch, and recruitment of the Japanese common squid, Todarodes pacificus (Cephalopoda: Ommastrephidae), from the 1970s to the 2000s. ICES Journal of Marine Science, 68(6):1114– 1121, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/68/6/1114/704348.

Scheld:2014:MEC

[SA14] Andrew M. Scheld and Christopher M. Anderson. Market effects of catch share management: the case of New England multispecies groundfish. ICES Journal of Marine Science, 71(7):1835–1845, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/7/1835/666705.

Sala:2016:CEO

[SAB^{+}16] M. M. Sala, F. L. Aparicio, V. Balagué, J. A. Boras, E. Borrull, C. Cardelús, L. Cros, A. Gomes, A. López-Sanz, A. Malits, R. A. Martínez, M. Mestre, J. Movilla, H. Sarmento, E. Vázquez-Domínguez, D. Vaqué, J. Pinhassi, A. Calbet, E. Calvo, J. M. Gasol, C. Pelejero, and C. Marrasé. Contrasting effects of ocean acidification on the microbial food web under different trophic conditions. ICES Journal of Marine Science, 73(3):670–679, February 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/3/670/2458721.

Skarethhamar:2018:MSL

[SAS^{+}18] Jofrid Skardhamar, Jon Albretsen, Anne D. Sandvik, Vidar S. Lien, Mari S. Myksvoll, Ingrid A. Johnsen, Lars Asplin, Bjørn Ådlandsvik, Elina Halttunen, and Pål Arne Bjørn. Modelled salmon lice dispersion and infestation patterns in a sub-Arctic fjord. ICES Journal of Marine Science, 75(5):1733–1747, September 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/5/1733/4954009.

Sassa:2019:ESB

[Sas19] Chiyuki Sassa. Estimation of the spawning biomass of myctophids based on larval production and reproductive parameters: the case study of Benthosema pterotum in the East China Sea. ICES Journal of Marine Science, 76(3): 743–754, May 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/76/3/743/4991919.

Stephenson:2017:PST

[SBB^{+}17] Robert L. Stephenson, Ashleen J. Benson, Kate Brooks, Anthony Charles, Poul Degnbol, Catherine M. Dichmont, Marloes Kraan, Sean Pascoe, Stacey D. Paul, Anna Rindorf, and Melanie Wiber. Practical steps toward integrating economic, social and institutional elements in fisheries policy and management. ICES Journal of Marine Science, 74(7):1981–1989, September 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/7/1981/3787887.

Soetaert:2019:GDU

[SBB19] Maarten Soetaert, Pim G. Boute, and William R. C. Beaumont. Guidelines for defining the use of electricity in marine electrotrawling. ICES Journal of Marine Science, 76(7):1994–2007, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
7/1994/5537344.

Shertzer:2016:RTC

[SBCF16] Kyle W. Shertzer, Nathan M. Bacheler, Lewis G. Coggins, Jr., and John Fieberg. Relating trap capture to abundance: a hierarchical state-space model applied to black sea bass (Centropristis striata). ICES Journal of Marine Science, 73(2):512–519, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/2/512/2614408.

Stanley:2015:EMT

[SBD^{+}15] Ryan R. E. Stanley, Ian R. Bradbury, Claudio DiBacco, Paul V. R. Snelgrove, Simon R. Thorrold, and Shaun S. Killen. Environmentally mediated trends in otolith composition of juvenile Atlantic cod (Gadus morhua). ICES Journal of Marine Science, 72(8):2350–2363, September 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/8/2350/2457883.

Schweigert:2010:RFL

[SBFC10] Jacob F. Schweigert, Jennifer L. Boldt, Linnea Flostrand, and Jaclyn S. Cleary. A review of factors limiting recovery of Pacific herring stocks in Canada. ICES Journal of Marine Science, 67(9):1903–1913, December 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/9/1903/623650.

Saillant:2010:GVS

[SBG10] Eric Saillant, S. Coleen Bradfield, and John R. Gold. Genetic variation and spatial autocorrelation among youngof-the-year red snapper (Lutjanus campechanus) in the northern Gulf of Mexico. ICES Journal of Marine Science, 67(6):1240–1250, September 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/67/
6/1240/733403.

Segvic-Bubic:2017:GCW

[ŠBGT^{+}17] Tanja Šegvić-Bubić, Leon Grubišić, Željka Trumbić, Rino Stanić, Jelena Ljubković, Jasna Maršić-Lučić, and Ivan Katavić. Genetic characterization of wild and farmed European seabass in the Adriatic sea: assessment of farmed escapees using a Bayesian approach. ICES Journal of Marine Science, 74(1):369–378, January 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/1/369/2669571.

Simons:2014:ISA

[SBH^{+}14] Sarah Laura Simons, Heleen Bartelings, Katell Gaelle Hamon, Alexander Johannes Kempf, Ralf Döring, and Axel Temming. Integrating stochastic age-structured population dynamics into complex fisheries economic models for management evaluations: the North Sea saithe fishery as a case study. ICES Journal of Marine Science, 71(7):1638– 1652, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/71/7/1638/2804379.

Stenson:2016:ICC

[SBKA16] Garry B. Stenson, Alejandro D. Buren, and Mariano Koen-Alonso. The impact of changing climate and abundance on reproduction in an ice-dependent species, the Northwest Atlantic harp seal, Pagophilus groenlandicus. ICES Journal of Marine Science, 73(2):250–262, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/2/250/2614432.

Simpson:2017:SLL

[SBL^{+}17] Cara A. Simpson, Hongsheng Bi, Dong Liang, Michael J. Wilberg, Amy M. Schueller, Geneviève M. Nesslage, and Harvey J. Walsh. Spawning locations and larval dispersal of Atlantic menhaden during 1977–2013. ICES Journal of Marine Science, 74(6):1574–1586, July 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/6/1574/3609043.

Skaret:2015:MPH

[SBLS15] Georg Skaret, Eneko Bachiller, Herdis Langøy, and Erling K. Stenevik. Mackerel predation on herring larvae during summer feeding in the Norwegian Sea. ICES Journal of Marine Science, 72(8):2313–2321, September 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/8/2313/2458733.

Shackell:2012:CLS

[SBNL12] Nancy L. Shackell, Alida Bundy, Janet A. Nye, and Jason S. Link. Common large-scale responses to climate and fishing across Northwest Atlantic ecosystems. ICES Journal of Marine Science, 69(2):151–162, March 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/2/151/698337.

Stenson:2011:UBD

[SBR11] Garry B. Stenson, Steven Benjamins, and David G. Reddin. Using bycatch data to understand habitat use of small cetaceans: lessons from an experimental driftnet fishery. ICES Journal of Marine Science, 68(5):937–946, May 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/5/937/653328.

Shin:2010:CSU

[SBS^{+}10] Yunne-Jai Shin, Alida Bundy, Lynne J. Shannon, Monique Simier, Marta Coll, Elizabeth A. Fulton, Jason S. Link, Didier Jouffre, Henn Ojaveer, Steven Mackinson, Johanna J. Heymans, and Tiit Raid. Can simple be useful and reliable? Using ecological indicators to represent and compare the states of marine ecosystems. ICES Journal of Marine Science, 67(4):717–731, May 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/67/4/717/683004.

Steele:2019:CMTa

[SBS19a] John H. Steele, Kenneth H. Brink, and Beth E. Scott. Comparison of marine and terrestrial ecosystems: suggestions of an evolutionary perspective influenced by environmental variation. ICES Journal of Marine Science, 76(1):50–59, January 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/1/50/5161207.

Steele:2019:CMTb

[SBS19b] John H. Steele, Kenneth H. Brink, and Beth E. Scott. Comparison of marine and terrestrial ecosystems: suggestions of an evolutionary perspective influenced by environmental variation. ICES Journal of Marine Science, 76(1):355, January 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/1/355/5259213.

Svedang:2019:LCG

[SBS^{+}19c] Henrik Svedäng, Julia M. I. Barth, Anders Svenson, Patrik Jonsson, Sissel Jentoft, Halvor Knutsen, and Carl André. Local cod (Gadus morhua) revealed by egg surveys and population genetic analysis after longstanding depletion on the Swedish Skagerrak coast. ICES Journal of Marine Science, 76(2):418–429, March 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/2/418/5224779. See comments [CMH19].

Sundblad:2014:NHA

[SBSE14] Göran Sundblad, Ulf Bergström, Alfred Sandström, and Peter Eklöv. Nursery habitat availability limits adult stock sizes of predatory coastal fish. ICES Journal of Marine Science, 71(3):672–680, March 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/71/3/672/631404.

Svedang:2016:CSA

[SC16] Henrik Svedäng and Massimiliano Cardinale. Comment on stock assessment of eels in the Baltic by westerberg and Wickström (2015): do we need more unknowns? ICES Journal of Marine Science, 73(6):1610–1612, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/6/1610/2458736.

Scanlon:2018:AUP

[Sca18] Zoe Scanlon. The art of “not undermining”: possibilities within existing architecture to improve environmental protections in areas beyond national jurisdiction. ICES Journal of Marine Science, 75(1):405–416, January 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/1/405/4657156.

Spitz:2013:PPA

[SCC^{+}13] J. Spitz, T. Chouvelon, M. Cardinaud, C. Kostecki, and P. Lorance. Prey preferences of adult sea bass Dicentrarchus labrax in the northeastern Atlantic: implications for bycatch of common dolphin Delphinus delphis. ICES Journal of Marine Science, 70(2):452–461, March 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/2/452/797965.

Soetaert:2015:DSR

[SCD^{+}15] Maarten Soetaert, Koen Chiers, Luc Duchateau, Hans Polet, Bart Verschueren, and Annemie Decostere. Determining the safety range of electrical pulses for two benthic invertebrates: brown shrimp (Crangon crangon L.) and ragworm (Alitta virens S.). ICES Journal of Marine Science, 72(3):973–980, March 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/72/3/973/691591.

Schwach:2014:SCJ

[Sch14] Vera Schwach. A sea change: Johan hjort and the natural fluctuations in the fish stocks. ICES Journal of Marine Science, 71(8):1993–1999, October 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/8/1993/2804434.

Stanton:2010:NBM

[SCJI10] Timothy K. Stanton, Dezhang Chu, J. Michael Jech, and James D. Irish. New broadband methods for resonance classification and high-resolution imagery of fish with swimbladders using a modified commercial broadband echosounder. ICES Journal of Marine Science, 67(2):365– 378, March 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/67/2/365/696001.

Strong:2019:RIM

[SCL^{+}19] James Asa Strong, Annika Clements, Helen Lillis, Ibon Galparsoro, Tim Bildstein, and Roland Pesch. A review of the influence of marine habitat classification schemes on mapping studies: inherent assumptions, influence on end products, and suggestions for future developments. ICES Journal of Marine Science, 76(1):10–22, January 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/1/10/5183306.

Stirnimann:2019:DRS

[SCM19] Luca Stirnimann, Alessandra Conversi, and Simone Marini. Detection of regime shifts in the environment: testing “STARS” using synthetic and observed time series. ICES Journal of Marine Science, 76(7):2286–2296, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/7/2286/5542081.

Simmonds:2011:DSR

[SCS^{+}11] E. John Simmonds, Andrew Campbell, Dankert Skagen, Beatriz A. Roel, and Ciaran Kelly. Development of a stock– recruit model for simulating stock dynamics for uncertain situations: the example of Northeast Atlantic mackerel (Scomber scombrus). ICES Journal of Marine Science, 68 (5):848–859, May 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/5/848/646854.

Shannon:2010:CDB

[SCY^{+}10] Lynne J. Shannon, Marta Coll, Dawit Yemane, Didier Jouffre, Sergio Neira, Arnaud Bertrand, Erich Diaz, and Yunne-Jai Shin. Comparing data-based indicators across upwelling and comparable systems for communicating ecosystem states and trends. ICES Journal of Marine Science, 67 (4):807–832, May 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/67/4/807/678850.

Silbiger:2017:CBO

[SD17] Nyssa J. Silbiger and Thomas M. DeCarlo. Comment on “Bioerosion: the other ocean acidification problem”: on field studies and mechanisms. ICES Journal of Marine Science, 74(9):2489–2493, November 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/9/2489/4317440. See reply [STF^{+}17].

Siddon:2018:SIS

[SDAM^{+}18] Elizabeth C. Siddon, Janet T. Duffy-Anderson, Kathryn L. Mier, Morgan S. Busby, and Lisa B. Eisner. Seasonal, interannual, and spatial patterns of community composition over the eastern Bering Sea shelf in cold years. part II: ichthyoplankton and juvenile fish. ICES Journal of Marine Science, 75(1):87–101, January 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/1/87/4082286.

Skiftesvik:2015:DHP

[SDBB15] Anne Berit Skiftesvik, Caroline M. F. Durif, Reidun M. Bjelland, and Howard I. Browman. Distribution and habitat preferences of five species of wrasse (family Labridae) in a Norwegian fjord. ICES Journal of Marine Science, 72(3): 890–899, March 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/72/3/890/701274.

Simmonds:2011:RFD

[SDDA11] E. J. Simmonds, R. Döring, P. Daniel, and V. Angot. The role of fisheries data in the development evaluation and impact assessment in support of European fisheries plans. ICES Journal of Marine Science, 68(8):1689–1698, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/68/8/1689/750780.

Sethi:2012:MHR

[SDH12] Suresh Andrew Sethi, Michael Dalton, and Ray Hilborn. Managing harvest risk with catch-pooling cooperatives. ICES Journal of Marine Science, 69(6):1038–1044, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/6/1038/619625.

Schulte:2018:ECM

[SDH^{+}18] Katharina Friederike Schulte, Andreas Dänhardt, Marc Hufnagl, Volker Siegel, Werner Wosniok, and Axel Temming. Not easy to catch: multiple covariates influence catch rates of brown shrimp (Crangon crangon L.), potentially affecting inferences drawn from catch and landings data. ICES Journal of Marine Science, 75(4):1318–1328, July 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/4/1318/4855829.

Schmiing:2015:MCM

[SDSA15] Mara Schmiing, Hugo Diogo, Ricardo Serrão Santos, and Pedro Afonso. Marine conservation of multispecies and multi-use areas with various conservation objectives and targets. ICES Journal of Marine Science, 72(3):851–862, March 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/3/851/2835888.

Subbey:2014:MFS

[SDSN14] Sam Subbey, Jennifer A. Devine, Ute Schaarschmidt, and Richard D. M. Nash. Modelling and forecasting stock– recruitment: current and future perspectives. ICES Journal of Marine Science, 71(8):2307–2322, October 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/8/2307/2804451.

Simons:2015:MFR

[SDT15] Sarah Laura Simons, Ralf Döring, and Axel Temming. Modelling fishers’ response to discard prevention strategies: the case of the North Sea saithe fishery. ICES Journal of Marine Science, 72(5):1530–1544, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/5/1530/763146.

Samuelsen:2019:ROC

[SDW19] Annette Samuelsen, Ute Daewel, and Cecilie Wettre. Risk of oil contamination of fish eggs and larvae under different oceanic and weather conditions. ICES Journal of Marine Science, 76(6):1902–1916, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/6/1902/5418568.

Stige:2019:DIE

[SEDO19] Leif Christian Stige, Elena Eriksen, Padmini Dalpadado, and Kotaro Ono. Direct and indirect effects of sea ice cover on major zooplankton groups and planktivorous fishes in the Barents Sea. ICES Journal of Marine Science, 76(S1):S24–S36, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/76/
Supplement_1/i24/5475871.

Seitz:2014:VCH

[Sei14] Rochelle D. Seitz. Value of coastal habitats for exploited species: introduction to a theme set of articles. ICES Journal of Marine Science, 71(3):636–637, March 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/3/636/638818.

Smith:2017:IME

[SFA^{+}17] David C. Smith, Elizabeth A. Fulton, Petrina Apfel, Ian D. Cresswell, Bronwyn M. Gillanders, Marcus Haward, Keith J. Sainsbury, Anthony D. M. Smith, Joanna Vince, and Tim M. Ward. Implementing marine ecosystembased management: lessons from Australia. ICES Journal of Marine Science, 74(7):1990–2003, September 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/7/1990/3958176.

Schonberg:2017:BOO

[SFCS^{+}17] Christine H. L. Schönberg, James K. H. Fang, Marina Carreiro-Silva, Aline Tribollet, and Max Wisshak. Bioerosion: the other ocean acidification problem. ICES Journal of Marine Science, 74(4):895–925, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/4/895/3064242.

Smith:2015:IFI

[SFD15] Michael D. Smith, Elizabeth A. Fulton, and Robert W. Day. An investigation into fisheries interaction effects using Atlantis. ICES Journal of Marine Science, 72(1):275–283, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/1/275/2804300.

Strahl:2016:BRO

[SFD^{+}16] J. Strahl, D. S. Francis, J. Doyle, C. Humphrey, and K. E. Fabricius. Biochemical responses to ocean acidification contrast between tropical corals with high and low abundances at volcanic carbon dioxide seeps. ICES Journal of Marine Science, 73(3):897–909, February 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/3/897/2458787.

Stelzenmuller:2015:QER

[SFG^{+}15] V. Stelzenmüller, H. O. Fock, A. Gimpel, H. Rambo, R. Diekmann, W. N. Probst, U. Callies, F. Bockelmann, H. Neumann, and I. Kröncke. Quantitative environmental risk assessments in the context of marine spatial management: current approaches and some perspectives. ICES Journal of Marine Science, 72(3):1022–1042, March 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/3/1022/699064.

Shephard:2012:SSF

[SFH^{+}12] Samuel Shephard, Tak Fung, Jennifer E. Houle, Keith D. Farnsworth, David G. Reid, and Axel G. Rossberg. Sizeselective fishing drives species composition in the Celtic Sea. ICES Journal of Marine Science, 69(2):223–234, March 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/2/223/700491.

Svenning:2019:CMP

[SFN^{+}19] Martin-A. Svenning, Morten Falkegård, Eero Niemelä, Juha-Pekka Vähä, Vidar Wennevik, Mikhail Ozerov, Sergey Prusov, J. Brian Dempson, Michael Power, and Per Fauchald. Coastal migration patterns of the four largest Barents Sea Atlantic salmon stocks inferred using genetic stock identification methods. ICES Journal of Marine Science, 76(6):1379–1389, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
6/1379/5522970.

Shackell:2016:TDD

[SFNdH16] Nancy L. Shackell, Kenneth T. Frank, Janet A. Nye, and Cornelia E. den Heyer. A transboundary dilemma: dichotomous designations of Atlantic halibut status in the Northwest Atlantic. ICES Journal of Marine Science, 73(7): 1798–1805, July 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/73/7/1798/2458730.

Sinerchia:2012:UIB

[SFW^{+}12] Matteo Sinerchia, Anthony J. Field, John D. Woods, Silvana Vallerga, and Wes R. Hinsley. Using an individualbased model with four trophic levels to model the effect of predation and competition on squid recruitment. ICES Journal of Marine Science, 69(3):439–447, May 2012.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/3/439/599768.

Sherwood:2010:ELH

[SG10] Graham D. Sherwood and Jonathan H. Grabowski. Exploring the life-history implications of colour variation in offshore Gulf of Maine cod (Gadus morhua). ICES Journal of Marine Science, 67(8):1640–1649, November 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/8/1640/607076.

Sherwood:2016:CCL

[SG16a] Graham D. Sherwood and Jonathan H. Grabowski. A comparison of cod life-history parameters inside and outside of four year-round groundfish closed areas in New England, USA. ICES Journal of Marine Science, 73(2):316–328, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/2/316/2614455.

Silva:2016:IEF

[SG16b] Catarina N. S. Silva and Jonathan P. A. Gardner. Identifying environmental factors associated with the genetic structure of the New Zealand scallop: linking seascape genetics and ecophysiological tolerance. ICES Journal of Marine Science, 73(7):1925–1934, July 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/7/1925/2457917.

Sakinan:2017:SDB

[SG17] Serdar Sakınan and Ali Cemal Gücü. Spatial distribution of the Black Sea copepod, Calanus euxinus, estimated using multi-frequency acoustic backscatter. ICES Journal of Marine Science, 74(3):832–846, March 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/3/832/2452356.

Sala:2018:CH

[SG18a] Enric Sala and Sylvaine Giakoumi. Counterpoint to hilborn. ICES Journal of Marine Science, 75(3):1163–1164, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/3/1163/4098835.

Sala:2018:NTM

[SG18b] Enric Sala and Sylvaine Giakoumi. No-take marine reserves are the most effective protected areas in the ocean. ICES Journal of Marine Science, 75(3):1166–1168, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/3/1166/4098821.

Sala:2018:SGF

[SG18c] Enric Sala and Sylvaine Giakoumi. Sala and Giakoumi’s final word. ICES Journal of Marine Science, 75(3):1171, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/3/1171/4098842.

Shephard:2011:FED

[SGK^{+}11] Samuel Shephard, Hans D. Gerritsen, Michel J. Kaiser, Holly S. Truszkowska, and David G. Reid. Fishing and environment drive spatial heterogeneity in Celtic Sea fish community size structure. ICES Journal of Marine Science, 68(10):2106–2113, November 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/68/
10/2106/610062.

Scoulding:2017:EVM

[SGM^{+}17] Ben Scoulding, Sven Gastauer, David N. MacLennan, Sascha M. M. Fässler, Phillip Copland, and Paul G. Fernandes. Effects of variable mean target strength on estimates of abundance: the case of Atlantic mackerel (Scomber scombrus). ICES Journal of Marine Science, 74(3):822–831, March 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/3/822/2734660.

Shephard:2015:SIT

[SGP^{+}15] Samuel Shephard, Simon P. R. Greenstreet, GerJan J. Piet, Anna Rindorf, and Mark Dickey-Collas. Surveillance indicators and their use in implementation of the marine strategy framework directive. ICES Journal of Marine Science, 72(8):2269–2277, September 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/8/2269/2459102.

Santos:2012:CTR

[SGQR^{+}12] M. Begoña Santos, Rafael González-Quirós, Isabel Riveiro, José M. Cabanas, Carmela Porteiro, and Graham J. Pierce. Cycles, trends, and residual variation in the Iberian sardine (Sardina pilchardus) recruitment series and their relationship with the environment. ICES Journal of Marine Science, 69(5):739–750, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/69/5/739/646010.

Selgrath:2018:ISD

[SGV18] Jennifer C. Selgrath, Sarah E. Gergel, and Amanda C. J. Vincent. Incorporating spatial dynamics greatly increases estimates of long-term fishing effort: a participatory mapping approach. ICES Journal of Marine Science, 75(1): 210–220, January 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/1/210/4091481.

Smith:2014:ISD

[SH14a] Stephen J. Smith and Brad Hubley. Impact of survey design changes on stock assessment advice: sea scallops. ICES Journal of Marine Science, 71(2):320–327, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/2/320/779846.

Stenson:2014:CIB

[SH14b] G. B. Stenson and M. O. Hammill. Can ice breeding seals adapt to habitat loss in a time of climate change? ICES Journal of Marine Science, 71(7):1977–1986, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/7/1977/2804452.

Svedang:2015:WFB

[SH15] Henrik Svedäng and Sara Hornborg. Waiting for a flourishing Baltic cod (Gadus morhua) fishery that never comes: old truths and new perspectives. ICES Journal of Marine Science, 72(8):2197–2208, September 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/8/2197/2458929.

Szuwalski:2016:CCN

[SH16] Cody S. Szuwalski and Anne B. Hollowed. Climate change and non-stationary population processes in fisheries management. ICES Journal of Marine Science, 73(5):1297– 1305, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/5/1297/2240681.

Skilbrei:2010:HMS

[SHAM10] Ove T. Skilbrei, Jens Christian Holst, Lars Asplin, and Stein Mortensen. Horizontal movements of simulated escaped farmed Atlantic salmon (Salmo salar) in a western Norwegian fjord. ICES Journal of Marine Science, 67(6):1206–1215, September 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/67/
6/1206/736803.

Sherman:2015:SWL

[She15] Kenneth Sherman. Sustaining the world’s large marine ecosystems. ICES Journal of Marine Science, 72(9):2521– 2531, November 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/72/9/2521/2457877.

Sistiaga:2018:CSG

[SHG^{+}18] Manu Sistiaga, Bent Herrmann, Eduardo Grimaldo, Roger B. Larsen, Leonore Olsen, Jesse Brinkhof, and Ivan Tatone. Combination of a sorting grid and a square mesh panel to optimize size selection in the North-East Arctic cod (Gadus morhua) and redfish (Sebastes spp.) trawl fisheries. ICES Journal of Marine Science, 75(3):1105–1116, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/3/1105/4780188.

Saha:2017:CSN

[SHH^{+}17] Atal Saha, Lorenz Hauser, Rasmus Hedeholm, Benjamin Planque, Svein-Erik Fevolden, Jesper Boje, and Torild Johansen. Cryptic Sebastes norvegicus species in Greenland waters revealed by microsatellites. ICES Journal of Marine Science, 74(8):2148–2158, September 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/8/2148/3782698.

Schimel:2010:QEC

[SHJI10] Alexandre C. G. Schimel, Terry R. Healy, David Johnson, and Dirk Immenga. Quantitative experimental comparison of single-beam, sidescan, and multibeam benthic habitat maps. ICES Journal of Marine Science, 67(8):1766–1779, November 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/67/8/1766/607895.

Saha:2015:SGS

[SHK^{+}15] Atal Saha, Lorenz Hauser, Matthew Kent, Benjamin Planque, Francis Neat, Tina Graceline Kirubakaran, Irene Huse, Eydna Í. Homrum, Svein-Erik Fevolden, Sigbjørn Lien, and Torild Johansen. Seascape genetics of saithe (Pollachius virens) across the North Atlantic using single nucleotide polymorphisms. ICES Journal of Marine Science, 72(9):2732–2741, November 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/72/
9/2732/2457957.

Secor:2017:ERP

[SHK17] David H. Secor, Edward D. Houde, and Loren L. Kellogg. Estuarine retention and production of striped bass larvae: a mark-recapture experiment. ICES Journal of Marine Science, 74(6):1735–1748, July 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
6/1735/2918632.

Soule:2010:ETS

[SHL10] Michael A. Soule, Ian Hampton, and Marek R. Lipiński. Estimating the target strength of live, free-swimming chokka squid Loligo reynaudii at 38 and 120 kHz. ICES Journal of Marine Science, 67(7):1381–1391, October 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/7/1381/663454.

Samhouri:2014:LLD

[SHL^{+}14] Jameal F. Samhouri, Alison J. Haupt, Phillip S. Levin, Jason S. Link, and Rebecca Shuford. Lessons learned from developing integrated ecosystem assessments to inform marine ecosystem-based management in the USA. ICES Journal of Marine Science, 71(5):1205–1215, July 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/5/1205/641719.

Singh:2011:RFA

[SHS11] Warsha Singh, Einar Hjorleifsson, and Gunnar Stefansson. Robustness of fish assemblages derived from three hierarchical agglomerative clustering algorithms performed on Icelandic groundfish survey data. ICES Journal of Marine Science, 68(1):189–200, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/68/
1/189/629176.

Skilbrei:2015:USE

[SHS15] Ove T. Skilbrei, Mikko Heino, and Terje Svåsand. Using simulated escape events to assess the annual numbers and destinies of escaped farmed Atlantic salmon of different life stages from farm sites in Norway. ICES Journal of Marine Science, 72(2):670–685, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/2/670/2801341.

Shafait:2017:TAU

[SHS^{+}17] Faisal Shafait, Euan S. Harvey, Mark R. Shortis, Ajmal Mian, Mehdi Ravanbakhsh, James W. Seager, Philip F. Culverhouse, Danelle E. Cline, and Duane R. Edgington. Towards automating underwater measurement of fish length: a comparison of semi-automatic and manual stereo–video measurements. ICES Journal of Marine Science, 74(6):1690–1701, July 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
6/1690/3056949.

Shepperson:2018:CVA

[SHS^{+}18] Jennifer L. Shepperson, Niels T. Hintzen, Claire L. Szostek, Ewen Bell, Lee G. Murray, and Michel J. Kaiser. A comparison of VMS and AIS data: the effect of data coverage and vessel position recording frequency on estimates of fishing footprints. ICES Journal of Marine Science, 75(3):988– 998, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/3/988/4775005.

Skogen:2018:FEC

[SHST18] Morten D. Skogen, Solfrid S. Hjøllo, Anne Britt Sandø, and Jerry Tjiputra. Future ecosystem changes in the Northeast Atlantic: a comparison between a global and a regional model system. ICES Journal of Marine Science, 75(7):2355–2369, December 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/75/
7/2355/5056135.

Shelton:2015:MAE

[SHW^{+}15] Andrew Olaf Shelton, Jeffrey A. Hutchings, Robin S. Waples, David M. Keith, H. Resit Akçakaya, and Nicholas K. Dulvy. Maternal age effects on Atlantic cod recruitment and implications for future population trajectories. ICES Journal of Marine Science, 72(6):1769–1778, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/6/1769/924938.

Shelton:2018:STM

[SHW^{+}18] Andrew O. Shelton, Mary E. Hunsicker, Eric J. Ward, Blake E. Feist, Rachael Blake, Colette L. Ward, Benjamin C. Williams, Janet T. Duffy-Anderson, Anne B. Hollowed, and Alan C. nHaynie. Spatio-temporal models reveal subtle changes to demersal communities following the Exxon Valdez oil spill. ICES Journal of Marine Science, 75(1):287–297, January 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/75/
1/287/3852237.

Spencer:2016:MSD

[SHZ^{+}16] Paul D. Spencer, Kirstin K. Holsman, Stephani Zador, Nicholas A. Bond, Franz J. Mueter, Anne B. Hollowed, and James N. Ianelli. Modelling spatially dependent predation mortality of eastern Bering Sea walleye pollock, and its implications for stock dynamics under future climate scenarios. ICES Journal of Marine Science, 73(5):1330– 1342, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/5/1330/2296608.

Stram:2015:EES

[SI15] Diana L. Stram and James N. Ianelli. Evaluating the efficacy of salmon bycatch measures using fishery-dependent data. ICES Journal of Marine Science, 72(4):1173–1180, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/4/1173/799876.

Szuwalski:2018:RRP

[SIP18] Cody S. Szuwalski, James N. Ianelli, and André E. Punt. Reducing retrospective patterns in stock assessment and impacts on management performance. ICES Journal of Marine Science, 75(2):596–609, March 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/2/596/4106929.

Shoji:2011:PEG

[SiTiM^{+}11] Jun Shoji, Shun ichi Toshito, Ken ichiro Mizuno, Yasuhiro Kamimura, Masakazu Hori, and Koji Hirakawa. Possible effects of global warming on fish recruitment: shifts in spawning season and latitudinal distribution can alter growth of fish early life stages through changes in daylength. ICES Journal of Marine Science, 68(6):1165– 1169, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/68/6/1165/711833.

Smith:2015:VMC

[SJ15] Walker O. Smith, Jr. and Randolph M. Jones. Vertical mixing, critical depths, and phytoplankton growth in the Ross Sea. ICES Journal of Marine Science, 72(6):1952– 1960, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/6/1952/917810.

Sundin:2016:DEE

[SJ16] Josefin Sundin and Fredrik Jutfelt. 9–28 d of exposure to elevated pCO_{2} reduces avoidance of predator odour but had no effect on behavioural lateralization or swimming activity in a temperate wrasse (Ctenolabrus rupestris). ICES Journal of Marine Science, 73(3):620–632, February 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/3/620/2458707.

Smith:2018:MAL

[SJ18] Danielle Smith and Julia Jabour. MPAs in ABNJ: lessons from two high seas regimes. ICES Journal of Marine Science, 75(1):417–425, January 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/75/
1/417/4430996.

Sathyendranath:2015:RSC

[SJB15] Shubha Sathyendranath, Rubao Ji, and Howard I. Browman. Revisiting Sverdrup’s critical depth hypothesis. ICES Journal of Marine Science, 72(6):1892–1896, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/6/1892/927499.

Solmundsson:2018:CAO

[SJRB18] Jón Sólmundsson, Ingibjörg G. Jónsdóttir, Stefán Á. Ragnarsson, and Björn Björnsson. Connectivity among offshore feeding areas and nearshore spawning grounds; implications for management of migratory fish. ICES Journal of Marine Science, 75(1):148–157, January 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/1/148/3884463.

Skaar:2011:AVD

[SJUE11] K. L. Skaar, T. Jørgensen, B. K. H. Ulvestad, and A. Engås. Accuracy of VMS data from Norwegian demersal stern trawlers for estimating trawled areas in the Barents Sea. ICES Journal of Marine Science, 68(8):1615–1620, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/68/8/1615/754372.

Sivle:2012:INS

[SKA^{+}12] Lise Doksæter Sivle, Petter Helgevold Kvadsheim, Michael A. Ainslie, Andrew Solow, Nils Olav Handegard, Nina Nordlund, and Frans-Peter A. Lam. Impact of naval sonar signals on Atlantic herring (Clupea harengus) during summer feeding. ICES Journal of Marine Science, 69(6):1078–1085, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/69/6/1078/621196.

Sivle:2015:PPL

[SKA15] L. D. Sivle, P. H. Kvadsheim, and M. A. Ainslie. Potential for population-level disturbance by active sonar in herring. ICES Journal of Marine Science, 72(2):558–567, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/2/558/2801451. See erratum [SKA22].

Sivle:2022:EPP

[SKA22] L. D. Sivle, P. H. Kvadsheim, and M. A. Ainslie. Erratum to: Potential for population-level disturbance by active sonar in herring. ICES Journal of Marine Science, 79 (1):244, January 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/79/1/244/6459379. See [SKA15].

Smale:2012:RSB

[SKH^{+}12] Dan A. Smale, Gary A. Kendrick, Euan S. Harvey, Timothy J. Langlois, Renae K. Hovey, Kimberly P. Van Niel, Kris I. Waddington, Lynda M. Bellchambers, Matthew B. Pember, Russ C. Babcock, Mathew A. Vanderklift, Damian P. Thomson, Michael V. Jakuba, Oscar Pizarro, and Stefan B. Williams. Regional-scale benthic monitoring for ecosystem-based fisheries management (EBFM) using an autonomous underwater vehicle (AUV). ICES Journal of Marine Science, 69(6):1108–1118, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/6/1108/621843.

Stanley:2015:DIE

[SKKM15] Richard D. Stanley, Tameezan Karim, John Koolman, and Howard McElderry. Design and implementation of electronic monitoring in the British Columbia groundfish hook and line fishery: a retrospective view of the ingredients of success. ICES Journal of Marine Science, 72(4):1230– 1236, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/4/1230/801122.

Savina:2018:DTC

[SKM18] Esther Savina, Ludvig Ahm Krag, and Niels Madsen. Developing and testing a computer vision method to quantify 3D movements of bottom-set gillnets on the seabed. ICES Journal of Marine Science, 75(2):814–824, March 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/2/814/4616492.

Sigwart:2016:EPD

[SLF^{+}16a] Julia D. Sigwart, Gillian Lyons, Artur Fink, Magdalena A. Gutowska, Darren Murray, Frank Melzner, Jonathan D. R. Houghton, and Marian Yong an Hu. Elevated pCO_{2} drives lower growth and yet increased calcification in the early life history of the cuttlefish Sepia officinalis (Mollusca: Cephalopoda). ICES Journal of Marine Science, 73(3):970–980, February 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
3/970/2458765.

Swiney:2016:EHP

[SLF16b] Katherine M. Swiney, William Christopher Long, and Robert J. Foy. Effects of high pCO_{2} on Tanner crab reproduction and early life history — part i: long-term exposure reduces hatching success and female calcification, and alters embryonic development. ICES Journal of Marine Science, 73(3):825–835, February 2016.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/3/825/2458897.

Swiney:2017:DPI

[SLF17] Katherine M. Swiney, W. Christopher Long, and Robert J. Foy. Decreased pH and increased temperatures affect young-of-the-year red king crab (Paralithodes camtschaticus). ICES Journal of Marine Science, 74(4):1191–1200, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/4/1191/3739849.

Spanier:2015:CRL

[SLG^{+}15] Ehud Spanier, Kari L. Lavalli, Jason S. Goldstein, Johan C. Groeneveld, Gareth L. Jordaan, Clive M. Jones, Bruce F. Phillips, Marco L. Bianchini, Rebecca D. Kibler, David Díaz, Sandra Mallol, Raquel Goñi, Gro I. van Der Meeren, Ann-Lisbeth Agnalt, Donald C. Behringer, William F. Keegan, and Andrew Jeffs. A concise review of lobster utilization by worldwide human populations from prehistory to the modern era. ICES Journal of Marine Science, 72(S1): S7–S21, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/72/suppl_1/i7/620442.

Shiao:2017:CSA

[SLH^{+}17] Jen-Chieh Shiao, Han-Bo Lu, Jhen Hsu, Hui-Yu Wang, Shui-Kai Chang, Min-Yu Huang, and Taiki Ishihara. Changes in size, age, and sex ratio composition of Pacific bluefin tuna (Thunnus orientalis) on the northwestern Pacific Ocean spawning grounds. ICES Journal of Marine Science, 74(1):204–214, January 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/1/204/2669558.

Sheehan:2010:PBG

[SLKS10] Timothy F. Sheehan, Christopher M. Legault, Timothy L. King, and Adrian P. Spidle. Probabilistic-based genetic assignment model: assignments to subcontinent of origin of the West Greenland Atlantic salmon harvest. ICES Journal of Marine Science, 67(3):537–550, April 2010.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/3/537/730129.

Santos:2011:DSA

[SLM^{+}11] Miguel N. Santos, Francisco Leitão, Ana Moura, Marco Cerqueira, and Carlos C. Monteiro. Diplodus spp. on artificial reefs of different ages: influence of the associated macrobenthic community. ICES Journal of Marine Science, 68(1):87–97, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/68/1/87/628472.

Stehfest:2015:UAA

[SLS15] Kilian M. Stehfest, Jeremy M. Lyle, and Jayson M. Semmens. The use of acoustic accelerometer tags to determine seasonal changes in activity and catchability of a recreationally caught marine teleost. ICES Journal of Marine Science, 72(8):2512–2520, September 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/8/2512/2458984.

Schweitzer:2018:IMT

[SLS18] Cara C. Schweitzer, Romuald N. Lipcius, and Bradley G. Stevens. Impacts of a multi-trap line on benthic habitat containing emergent epifauna within the Mid-Atlantic Bight. ICES Journal of Marine Science, 75(6):2202–2212, November 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/6/2202/5096002.

Sammarco:2014:CCA

[SLT^{+}14] P. W. Sammarco, A. Lirette, Y. F. Tung, G. S. Boland, M. Genazzio, and J. Sinclair. Coral communities on artificial reefs in the Gulf of Mexico: standing vs. toppled oil platforms. ICES Journal of Marine Science, 71(2):417–426, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/71/2/417/785062.

Sun:2017:BEE

[SLT^{+}17] Peng Sun, Xiaozi Liu, Yanli Tang, Wenzhi Cheng, Runlong Sun, Xinxin Wang, Rong Wan, and Mikko Heino. The bio-economic effects of artificial reefs: mixed evidence from Shandong, China. ICES Journal of Marine Science, 74(8):2239–2248, September 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
8/2239/3852234.

Sallee:2015:CDB

[SLTL15] Jean-Baptiste Sallée, J. Llort, A. Tagliabue, and M. Lévy. Characterization of distinct bloom phenology regimes in the Southern Ocean. ICES Journal of Marine Science, 72(6): 1985–1998, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/72/6/1985/925140.

Soetaert:2016:RBB

[SLV16] Maarten Soetaert, Heleen Lenoir, and Bart Verschueren. Reducing bycatch in beam trawls and electrotrawls with (electrified) benthos release panels. ICES Journal of Marine Science, 73(9):2370–2379, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
9/2370/2199034.

Skagseth:2012:HCN

[SM12] Øystein Skagseth and Kjell Arne Mork. Heat content in the Norwegian Sea, 1995–2010. ICES Journal of Marine Science, 69(5):826–832, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/69/5/826/648254.

Stewart:2015:RSA

[SM15] Ian J. Stewart and Steven J. D. Martell. Reconciling stock assessment paradigms to better inform fisheries management. ICES Journal of Marine Science, 72(8):2187– 2196, September 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/72/8/2187/2457865.

Small:2016:TFS

[SMB^{+}16a] Daniel P. Small, Marco Milazzo, Camilla Bertolini, Helen Graham, Chris Hauton, Jason M. Hall-Spencer, and Samuel P. S. Rastrick. Temporal fluctuations in seawater pCO_{2} may be as important as mean differences when determining physiological sensitivity in natural systems. ICES Journal of Marine Science, 73(3):604–612, February 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/3/604/2459051.

Szostek:2016:NVF

[SMB^{+}16b] Claire L. Szostek, Lee G. Murray, Ewen Bell, Gemma Rayner, and Michel J. Kaiser. Natural vs. fishing disturbance: drivers of community composition on traditional king scallop, Pecten maximus, fishing grounds. ICES Journal of Marine Science, 73(suppl 1):S70–S83, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/73/suppl_1/i70/2573994.

Sbragaglia:2018:SMF

[SMB^{+}18] Valerio Sbragaglia, Lorenzo Morroni, Lorenzo Bramanti, Boris Weitzmann, Robert Arlinghaus, and Ernesto Azzurro. Spearfishing modulates flight initiation distance of fishes: the effects of protection, individual size, and bearing a speargun. ICES Journal of Marine Science, 75(5):1779– 1789, September 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/5/1779/5032467.

Steen:2016:RAK

[SMBH16] Henning Steen, Frithjof E. Moy, Torjan Bodvin, and Vivian Husa. Regrowth after kelp harvesting in Nord–Trøndelag, Norway. ICES Journal of Marine Science, 73(10):2708– 2720, November 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/73/10/2708/2647119.

Stierhoff:2016:IET

[SMD^{+}16] Kevin L. Stierhoff, David W. Murfin, David A. Demer, Scott A. Mau, and Deanna R. Pinkard-Meier. Improving the estimations of transect length and width for underwater visual surveys of targets on or near the seabed. ICES Journal of Marine Science, 73(10):2729–2736, November 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/10/2729/2647107.

Silva:2019:OCP

[SMH^{+}19] C. N. S. Silva, H. S. Macdonald, M. G. Hadfield, M. Cryer, and J. P. A. Gardner. Ocean currents predict fine-scale genetic structure and source-sink dynamics in a marine invertebrate coastal fishery. ICES Journal of Marine Science, 76(4):1007–1018, July 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/4/1007/5303700.

Skjaeraasen:2011:ESS

[SMK^{+}11] Jon Egil Skjæraasen, Justin J. Meager, Ørjan Karlsen, Jeffrey A. Hutchings, and Anders Fernö. Extreme spawning-site fidelity in Atlantic cod. ICES Journal of Marine Science, 68(7):1472–1477, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/68/
7/1472/654316.

Stone:2015:AEI

[SMK15] Robert P. Stone, Michele M. Masuda, and John F. Karinen. Assessing the ecological importance of red tree coral thickets in the eastern Gulf of Alaska. ICES Journal of Marine Science, 72(3):900–915, March 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/3/900/694278.

Sissenwine:2014:POE

[SML14] Michael M. Sissenwine, Pamela M. Mace, and Hans J. Lassen. Preventing overfishing: evolving approaches and emerging challenges. ICES Journal of Marine Science, 71(2):153–156, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/71/2/153/791087.

Stanley:2011:AAC

[SMMK11] Richard D. Stanley, Howard McElderry, Tameezan Mawani, and John Koolman. The advantages of an audit over a census approach to the review of video imagery in fishery monitoring. ICES Journal of Marine Science, 68(8):1621– 1627, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/8/1621/749488.

Seitz:2014:EFS

[SMNE14] Andrew C. Seitz, Kathrine Michalsen, Jennifer L. Nielsen, and Mark D. Evans. Evidence of fjord spawning by southern Norwegian Atlantic halibut (Hippoglossus hippoglossus). ICES Journal of Marine Science, 71(5):1142–1147, July 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/71/5/1142/646806.

Skern-Mauritzen:2018:OAE

[SMOH18] Mette Skern-Mauritzen, Erik Olsen, and Geir Huse. Opportunities for advancing ecosystem-based management in a rapidly changing, high latitude ecosystem. ICES Journal of Marine Science, 75(7):2425–2433, December 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/7/2425/5144590.

Steingrund:2010:TSS

[SMR^{+}10] Petur Steingrund, Rógvi Mouritsen, Jákup Reinert, Eilif Gaard, and Hjálmar Hátún. Total stock size and cannibalism regulate recruitment in cod (Gadus morhua) on the Faroe Plateau. ICES Journal of Marine Science, 67(1):111– 124, January 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/67/1/111/596456.

Saitoh:2011:SOU

[SMR^{+}11] Sei-Ichi Saitoh, Robinson Mugo, I. Nyoman Radiarta, Shinsuke Asaga, Fumihiro Takahashi, Toru Hirawake, Yoichi Ishikawa, Toshiyuki Awaji, Teiji In, and Shigeki Shima. Some operational uses of satellite remote sensing and marine GIS for sustainable fisheries and aquaculture. ICES Journal of Marine Science, 68(4):687–695, March 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/4/687/649951.

Sabates:2012:CLH

[SMR12] Ana Sabatés, Paloma Martín, and Vanesa Raya. Changes in life-history traits in relation to climate change: bluefish (Pomatomus saltatrix) in the northwestern Mediterranean. ICES Journal of Marine Science, 69(6):1000–1009, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/69/6/1000/618385.

Shafait:2016:FIV

[SMS^{+}16] Faisal Shafait, Ajmal Mian, Mark Shortis, Bernard Ghanem, Phil F. Culverhouse, Duane Edgington, Danelle Cline, Mehdi Ravanbakhsh, James Seager, and Euan S. Harvey. Fish identification from videos captured in uncontrolled underwater environments. ICES Journal of Marine Science, 73(10):2737–2746, November 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
10/2737/2647103.

Stephenson:2017:EPI

[SMS^{+}17] Fabrice Stephenson, Aileen C. Mill, Catherine L. Scott, Nicholas V. C. Polunin, and Clare Fitzsimmons. Experimental potting impacts on common UK reef habitats in areas of high and low fishing pressure. ICES Journal of Marine Science, 74(6):1648–1659, July 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/6/1648/3052250.

Spitz:2010:PCE

[SMSR10] Jérôme Spitz, Emeline Mourocq, Valérie Schoen, and Vincent Ridoux. Proximate composition and energy content of forage species from the Bay of Biscay: high- or lowquality food? ICES Journal of Marine Science, 67(5):909– 915, July 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/67/5/909/610336.

Shephard:2014:STS

[SMZ^{+}14] Samuel Shephard, Cóilín Minto, Melanie Zölck, Simon Jennings, Deirdre Brophy, and David Reid. Scavenging on trawled seabeds can modify trophic size structure of bottom-dwelling fish. ICES Journal of Marine Science, 71(2):398–405, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/71/2/398/784277.

Steingrund:2010:DDD

[SO10] Petur Steingrund and Lise H. Ofstad. Density-dependent distribution of Atlantic cod (Gadus morhua) into deep waters on the Faroe Plateau. ICES Journal of Marine Science, 67(1):102–110, January 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/67/1/102/596360.

Skilbrei:2016:VDS

[SO16] O. T. Skilbrei and H. Otterå. Vertical distribution of saithe (Pollachius virens) aggregating around fish farms. ICES Journal of Marine Science, 73(4):1186–1195, March 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/4/1186/2458753.

Sherman:2011:ASR

[SOB^{+}11] Kenneth Sherman, John O’Reilly, Igor M. Belkin, Christopher Melrose, and Kevin D. Friedland. The application of satellite remote sensing for assessing productivity in relation to fisheries yields of the world’s large marine ecosystems. ICES Journal of Marine Science, 68(4):667–676, March 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/68/4/667/647498.

Scharff-Olsen:2019:DSB

[SOGT^{+}19] Camilla Hjorth Scharff-Olsen, Anders Galatius, Jonas Teilmann, Rune Dietz, Signe May Andersen, Simon Jarnit, Anne-Mette Kroner, Amanda Bolt Botnen, Karl Lundström, Peter Rask Møller, and Morten Tange Olsen. Diet of seals in the Baltic Sea region: a synthesis of published and new data from 1968 to 2013. ICES Journal of Marine Science, 76(1):284–297, January 2019.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/1/284/5184893.

Saunders:2012:UKA

[SOK^{+}12] Ryan A. Saunders, Ciaran O’Donnell, Rolf J. Korneliussen, Sascha M. M. Fässler, Maurice W. Clarke, Afra Egan, and Dave Reid. Utility of 18-kHz acoustic data for abundance estimation of Atlantic herring (Clupea harengus). ICES Journal of Marine Science, 69(6):1086–1098, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/6/1086/618867.

Straley:2015:SAS

[SOL^{+}15] Janice Straley, Victoria O’Connell, Joe Liddle, Aaron Thode, Lauren Wild, Linda Behnken, Dan Falvey, and Chris Lunsford. Southeast Alaska sperm whale avoidance project (SEASWAP): a successful collaboration among scientists and industry to study depredation in Alaskan waters. ICES Journal of Marine Science, 72(5):1598–1609, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/5/1598/780563.

Szuwalski:2013:FMR

[SP13] Cody S. Szuwalski and André E. Punt. Fisheries management for regime-based ecosystems: a management strategy evaluation for the snow crab fishery in the eastern Bering Sea. ICES Journal of Marine Science, 70(5):955– 967, September 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/70/5/955/642524.

Staaterman:2014:MLF

[SP14] Erica Staaterman and Claire B. Paris. Modelling larval fish navigation: the way forward. ICES Journal of Marine Science, 71(4):918–924, May 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/71/4/918/665210.

Sarmento:2017:IPC

[SPdJFMS17] Visnu Cunha Sarmento, Bárbara Ramos Pinheiro, Manuel de Jesus Flores Montes, and Paulo Jorge Parreira Santos. Impact of predicted climate change scenarios on a coral reef meiofauna community. ICES Journal of Marine Science, 74(4):1170–1179, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
4/1170/2870688.

Sonderblohm:2014:EFD

[SPE14] Carlos P. Sonderblohm, João Pereira, and Karim Erzini. Environmental and fishery-driven dynamics of the common octopus (Octopus vulgaris) based on time-series analyses from leeward Algarve, southern Portugal. ICES Journal of Marine Science, 71(8):2231–2241, October 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/8/2231/746913.

Serra-Pereira:2010:MRS

[SPFM^{+}10] Bárbara Serra-Pereira, Inês Farias, Teresa Moura, Leonel Serrano Gordo, Miguel Santos, and Ivone Figueiredo. Morphometric ratios of six commercially landed species of skate from the Portuguese continental shelf, and their utility for identification. ICES Journal of Marine Science, 67(8):1596–1603, November 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/67/
8/1596/604204.

Sampedro:2017:SSE

[SPG^{+}17] Paz Sampedro, Raúl Prellezo, Dorleta García, José María Da-Rocha, Santiago Cerviño, Julia Torralba, Julia Touza, Javier García-Cutrín, and María José Gutiérrez. To shape or to be shaped: engaging stakeholders in fishery management advice. ICES Journal of Marine Science, 74(2): 487–498, March 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/74/2/487/2670307.

Stephenson:2017:STC

[SPM^{+}17] Fabrice Stephenson, Nicholas V. C. Polunin, Aileen C. Mill, Catherine Scott, Paula Lightfoot, and Clare Fitzsimmons. Spatial and temporal changes in pot-fishing effort and habitat use. ICES Journal of Marine Science, 74(8):2201– 2212, September 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/74/8/2201/3778475.

Sharma:2018:FRD

[SPM^{+}18] Rishi Sharma, Maite Pons, Sarah Martin, Laurie Kell, John Walter, Matthew Lauretta, and Michael Schirripa. Factors related to the decline and rebuilding of billfish stocks in the Atlantic and Indian oceans. ICES Journal of Marine Science, 75(2):880–891, March 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/2/880/3836955.

Stephenson:2016:IFK

[SPP^{+}16] Robert L. Stephenson, Stacey Paul, Martin A. Pastoors, Marloes Kraan, Petter Holm, Melanie Wiber, Steven Mackinson, Dorothy J. Dankel, Kate Brooks, and Ashleen Benson. Integrating fishers’ knowledge research in science and management. ICES Journal of Marine Science, 73(6):1459– 1465, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/6/1459/2459055.

Sato:2018:EPP

[SPRL18] Kirk N. Sato, Jackson Powell, Dave Rudie, and Lisa A. Levin. Evaluating the promise and pitfalls of a potential climate change–tolerant sea urchin fishery in southern California. ICES Journal of Marine Science, 75(3):1029–1041, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/3/1029/4745800.

Simmonds:2010:IAB

[SPS^{+}10] E. John Simmonds, Enrique Portilla, Dankert Skagen, Doug Beare, and Dave G. Reid. Investigating agreement between different data sources using Bayesian statespace models: an application to estimating NE Atlantic mackerel catch and stock abundance. ICES Journal of Marine Science, 67(6):1138–1153, September 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/6/1138/733749.

Stuart:2011:FFS

[SPS11] Venetia Stuart, Trevor Platt, and Shubha Sathyendranath. The future of fisheries science in management: a remotesensing perspective. ICES Journal of Marine Science, 68(4):644–650, March 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/68/4/644/651758.

Siemann:2015:SCH

[SPS15] Liese A. Siemann, Christopher J. Parkins, and Ronald J. Smolowitz. Scallops caught in the headlights: swimming escape behaviour of the Atlantic sea scallop (Placopecten magellanicus) reduced by artificial light. ICES Journal of Marine Science, 72(9):2700–2706, November 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/9/2700/2458906.

Stuart:2011:RSF

[SPSP11] Venetia Stuart, Trevor Platt, Shubha Sathyendranath, and P. Pravin. Remote sensing and fisheries: an introduction. ICES Journal of Marine Science, 68(4):639–641, March 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/4/639/650725.

Sys:2016:CIB

[SPV^{+}16] Klaas Sys, Jan Jaap Poos, Jef Van Meensel, Hans Polet, and Jeroen Buysse. Competitive interactions between two fishing fleets in the North Sea. ICES Journal of Marine Science, 73(6):1485–1493, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/6/1485/2458780.

Sorochan:2014:HDD

[SQ14] Kevin A. Sorochan and Pedro A. Quijón. Horizontal distributions of Dungeness crab (Cancer magister) and red rock crab (Cancer productus) larvae in the Strait of Georgia, British Columbia. ICES Journal of Marine Science, 71(9):2564–2577, November 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/9/2564/2798167.

Sullivan:2016:QAS

[SR16] Patrick J. Sullivan and Lars G. Rudstam. Quantifying acoustic survey uncertainty using Bayesian hierarchical modeling with an application to assessing Mysis relicta population densities in Lake Ontario. ICES Journal of Marine Science, 73(8):2104–2111, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/8/2104/2198462.

Smolinski:2017:SPD

[SR17] Szymon Smoliński and Krzysztof Radtke. Spatial prediction of demersal fish diversity in the Baltic Sea: comparison of machine learning and regression-based techniques. ICES Journal of Marine Science, 74(1):102–111, January 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/1/102/2669554.

Soma:2014:MAE

[SRB^{+}14] Katrine Soma, Jorge Ramos, Øivind Bergh, Torsten Schulze, Hans van Oostenbrugge, Arie Pieter van Duijn, Kathrin Kopke, Vanessa Stelzenmüller, Fabio Grati, Timo Mäkinen, Claus Stenberg, and Erik Buisman. The “mapping out” approach: effectiveness of marine spatial management options in European coastal waters. ICES Journal of Marine Science, 71(9):2630–2642, November 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/9/2630/592994.

Saunders:2011:WMD

[SRC11] Ryan A. Saunders, François Royer, and Maurice W. Clarke. Winter migration and diving behaviour of porbeagle shark, Lamna nasus, in the Northeast Atlantic. ICES Journal of Marine Science, 68(1):166–174, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/1/166/629325.

Sheehan:2012:SNA

[SRCR12] Timothy F. Sheehan, David G. Reddin, Gérald Chaput, and Mark D. Renkawitz. SALSEA North America: a pelagic ecosystem survey targeting Atlantic salmon in the Northwest Atlantic. ICES Journal of Marine Science, 69(9):1580–1588, November 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/69/
9/1580/636388.

Shephard:2014:ASP

[SRDC^{+}14] Samuel Shephard, Anna Rindorf, Mark Dickey-Collas, Niels T. Hintzen, Keith Farnsworth, and David G. Reid. Assessing the state of pelagic fish communities within an ecosystem approach and the European marine strategy framework directive. ICES Journal of Marine Science, 71(7):1572–1585, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/71/
7/1572/668673.

Shephard:2011:ILF

[SRG11] Samuel Shephard, David G. Reid, and Simon P. R. Greenstreet. Interpreting the large fish indicator for the Celtic Sea. ICES Journal of Marine Science, 68(9):1963–1972, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/9/1963/667144.

Shephard:2015:EBF

[SRGF15] Samuel Shephard, David G. Reid, Hans D. Gerritsen, and Keith D. Farnsworth. Estimating biomass, fishing mortality, and “total allowable discards” for surveyed non-target fish. ICES Journal of Marine Science, 72(2):458–466, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/2/458/2801434.

Smout:2014:MPC

[SRH^{+}14] Sophie Smout, Anna Rindorf, Philip S. Hammond, John Harwood, and Jason Matthiopoulos. Modelling prey consumption and switching by UK grey seals. ICES Journal of Marine Science, 71(1):81–89, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/1/81/645258.

Sherwin:2012:ICN

[SRHJ12] Toby J. Sherwin, Jane F. Read, N. Penny Holliday, and Clare Johnson. The impact of changes in North Atlantic Gyre distribution on water mass characteristics in the Rockall Trough. ICES Journal of Marine Science, 69(5): 751–757, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/69/5/751/645870.

Syrjanen:2015:RBS

[SRKV15] Jukka Tapani Syrjänen, Timo Juhani Ruokonen, Tarmo Ketola, and Pentti Valkeajärvi. The relationship between stocking eggs in boreal spawning rivers and the abundance of brown trout parr. ICES Journal of Marine Science, 72(5):1389–1398, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/72/
5/1389/776651.

Simonsen:2016:STD

[SRRZ16] Kirsten A. Simonsen, Patrick H. Ressler, Christopher N. Rooper, and Stephani G. Zador. Spatio-temporal distribution of euphausiids: an important component to understanding ecosystem processes in the Gulf of Alaska and eastern Bering Sea. ICES Journal of Marine Science, 73(8):2020–2036, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
8/2020/2198237.

Santora:2011:SOK

[SRS11] Jarrod A. Santora, Stephen Ralston, and William J. Sydeman. Spatial organization of krill and seabirds in the central California Current. ICES Journal of Marine Science, 68(7): 1391–1402, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/7/1391/653365.

Smith:2013:DWB

[SRW13] Joy N. Smith, Patrick H. Ressler, and Joseph D. Warren. A distorted wave Born approximation target strength model for Bering Sea euphausiids. ICES Journal of Marine Science, 70(1):204–214, January 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/70/
1/204/661080.

Shin:2010:UIEa

[SS10a] Yunne-Jai Shin and Lynne J. Shannon. Using indicators for evaluating, comparing, and communicating the ecological status of exploited marine ecosystems. 1. The IndiSeas project. ICES Journal of Marine Science, 67(4):686– 691, May 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/67/4/686/679710.

Sjare:2010:CRP

[SS10b] Becky Sjare and Garry B. Stenson. Changes in the reproductive parameters of female harp seals (Pagophilus groenlandicus) in the Northwest Atlantic. ICES Journal of Marine Science, 67(2):304–315, March 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/67/
2/304/696120.

Smith:2012:MPL

[SS12] Woollcott K. Smith and Andrew R. Solow. Missing and presumed lost: extinction in the ocean and its inference. ICES Journal of Marine Science, 69(1):89–94, January 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/1/89/671867.

Skaret:2017:HSD

[SS17] Georg Skaret and Aril Slotte. Herring submesoscale dynamics through a major spawning wave: duration, abundance fluctuation, distribution, and schooling. ICES Journal of Marine Science, 74(3):717–727, March 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/3/717/2444574.

Stoehr:2018:FSS

[SSA^{+}18] Ashley Stoehr, Joshua St. Martin, Scott Aalbers, Chugey Sepulveda, and Diego Bernal. Free-swimming swordfish, Xiphias gladius, alter the rate of whole body heat transfer: morphological and physiological specializations for thermoregulation. ICES Journal of Marine Science, 75(2): 858–870, March 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/2/858/4111230.

Shin:2010:UIEb

[SSB^{+}10a] Yunne-Jai Shin, Lynne J. Shannon, Alida Bundy, Marta Coll, Kerim Aydin, Nicolas Bez, Julia L. Blanchard, Maria de Fatima Borges, Ibrahima Diallo, Erich Diaz, Johanna J. Heymans, Louize Hill, Edda Johannesen, Didier Jouffre, Souad Kifani, Pierre Labrosse, Jason S. Link, Steven Mackinson, Hicham Masski, Christian Möllmann, Sergio Neira, Henn Ojaveer, Khairdine ould Mohammed Abdallahi, Ian Perry, Djiga Thiao, Dawit Yemane, and Philippe M. Cury. Using indicators for evaluating, comparing, and communicating the ecological status of exploited marine ecosystems.

2. setting the scene. ICES Journal of Marine Science, 67 (4):692–716, May 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/67/4/692/683270.

Solari:2010:DSP

[SSB^{+}10b] A. P. Solari, M. T. G. Santamaría, M. F. Borges, A. M. P. Santos, H. Mendes, E. Balguerías, J. A. Díaz Cordero, J. J. Castro, and C. Bas. On the dynamics of Sardina pilchardus: orbits of stability and environmental forcing. ICES Journal of Marine Science, 67(8):1565–1573, November 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/67/8/1565/608088.

Sonntag:2012:SSN

[SSF^{+}12] Nicole Sonntag, Henriette Schwemmer, Heino Ove Fock, Jochen Bellebaum, and Stefan Garthe. Seabirds, setnets, and conservation management: assessment of conflict potential and vulnerability of birds to bycatch in gillnets. ICES Journal of Marine Science, 69(4):578–589, May 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/4/578/634399.

Siple:2018:CAM

[SSF^{+}18a] Margaret C. Siple, Andrew O. Shelton, Tessa B. Francis, Dayv Lowry, Adam P. Lindquist, and Timothy E. Essington. Contributions of adult mortality to declines of Puget Sound Pacific herring. ICES Journal of Marine Science, 75(1):319–329, January 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/1/319/3862121.

Solvang:2018:CDB

[SSF18b] Hiroko K. Solvang, Sam Subbey, and Anna S. J. Frank. Causal drivers of Barents Sea capelin (Mallotus villosus) population dynamics on different time scales. ICES Journal of Marine Science, 75(2):621–630, March 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/2/621/4345801.

Staby:2018:STD

[SSGH18] Arved Staby, Jon Egil Skjæraasen, Audrey J. Geffen, and Daniel Howell. Spatial and temporal dynamics of European hake (Merluccius merluccius) in the North Sea. ICES Journal of Marine Science, 75(6):2033–2044, November 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/6/2033/5071917.

Smith:2014:IHS

[SSH^{+}14] Anthony D. M. Smith, David C. Smith, Malcolm Haddon, Ian A. Knuckey, Keith J. Sainsbury, and Sean R. Sloan. Implementing harvest strategies in Australia: 5 years on. ICES Journal of Marine Science, 71(2):195–203, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/2/195/788673.

Sarmento:2017:EEC

[SSH^{+}17] Visnu Cunha Sarmento, Paulo Jorge Parreira Santos, Rachel Hale, Jeroen Ingels, and Stephen Widdicombe.

Effects of elevated CO_{2} and temperature on an intertidal harpacticoid copepod community. ICES Journal of Marine Science, 74(4):1159–1169, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
4/1159/2667172.

Schram:2016:TAR

[SSM^{+}16] Julie B. Schram, Kathryn M. Schoenrock, James B. Mc-Clintock, Charles D. Amsler, and Robert A. Angus. Testing Antarctic resilience: the effects of elevated seawater temperature and decreased pH on two gastropod species. ICES Journal of Marine Science, 73(3):739–752, February 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/3/739/2459083.

Siddiqui:2018:AFS

[SSM^{+}18] Shoaib Ahmed Siddiqui, Ahmad Salman, Muhammad Imran Malik, Faisal Shafait, Ajmal Mian, Mark R. Shortis, and Euan S. Harvey. Automatic fish species classification in underwater videos: exploiting pre-trained deep neural network models to compensate for limited labelled data. ICES Journal of Marine Science, 75(1):374–389, January 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/1/374/3924506.

Skagen:2013:SFE

[SSMD^{+}13] Dankert W. Skagen, Mette Skern-Mauritzen, Dorothy Dankel, Katja Enberg, Olav S. Kjesbu, and Richard D. M. Nash. A simulation framework for evaluating fisheries management decisions using environmental information. ICES Journal of Marine Science, 70(4):743–754, July 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/4/743/726035.

Su:2011:MIE

[SSP^{+}11] Nan-Jay Su, Chi-Lu Sun, André E. Punt, Su-Zan Yeh, and Gerard DiNardo. Modelling the impacts of environmental variation on the distribution of blue marlin, Makaira nigricans, in the Pacific Ocean. ICES Journal of Marine Science, 68(6):1072–1080, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/6/1072/701442.

Sparrevohn:2012:UIB

[SSP12] Claus Reedtz Sparrevohn and Marie Storr-Paulsen. Using interview-based recall surveys to estimate cod Gadus morhua and eel Anguilla anguilla harvest in Danish recreational fishing. ICES Journal of Marine Science, 69(2): 323–330, March 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/69/2/323/705466.

Su:2013:EAP

[SSP^{+}13] Nan-Jay Su, Chi-Lu Sun, André E. Punt, Su-Zan Yeh, Gerard DiNardo, and Yi-Jay Chang. An ensemble analysis to predict future habitats of striped marlin (Kajikia audax) in the North Pacific Ocean. ICES Journal of Marine Science, 70(5):1013–1022, September 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/5/1013/643037.

Sagebiel:2016:EVB

[SSR^{+}16] Julian Sagebiel, Carmen Schwartz, Mounaim Rhozyel, Sandra Rajmis, and Jesko Hirschfeld. Economic valuation of Baltic marine ecosystem services: blind spots and limited consistency. ICES Journal of Marine Science, 73(4):991– 1003, March 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/73/4/991/2458767.

Schopka:2010:UTE

[SSRT10] Sigfus A. Schopka, Jon Solmundsson, Stefan Aki Ragnarsson, and Vilhjalmur Thorsteinsson. Using tagging experiments to evaluate the potential of closed areas in protecting migratory Atlantic cod (Gadus morhua). ICES Journal of Marine Science, 67(5):1024–1035, July 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/5/1024/607467.

Santora:2012:KSC

[SSS^{+}12] Jarrod A. Santora, William J. Sydeman, Isaac D. Schroeder, Christian S. Reiss, Brian K. Wells, John C. Field, Anthony M. Cossio, and Valerie J. Loeb. Krill space: a comparative assessment of mesoscale structuring in polar and temperate marine ecosystems. ICES Journal of Marine Science, 69(7):1317–1327, September 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/69/
7/1317/742083.

Stortini:2015:AMS

[SSTB15] Christine H. Stortini, Nancy L. Shackell, Peter Tyedmers, and Karen Beazley. Assessing marine species vulnerability to projected warming on the Scotian Shelf, Canada. ICES Journal of Marine Science, 72(6):1731–1743, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/6/1731/918246.

Strehlow:2012:CCT

[SSZH12] Harry V. Strehlow, Norbert Schultz, Christopher Zimmermann, and Cornelius Hammer. Cod catches taken by the German recreational fishery in the western Baltic Sea, 2005–2010: implications for stock assessment and management. ICES Journal of Marine Science, 69(10):1769–1780, December 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/69/10/1769/623611.

Sveegaard:2011:ASC

[STB^{+}11] Signe Sveegaard, Jonas Teilmann, Per Berggren, Kim N. Mouritsen, Douglas Gillespie, and Jakob Tougaard. Acoustic surveys confirm the high-density areas of harbour porpoises found by satellite tracking. ICES Journal of Marine Science, 68(5):929–936, May 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/68/5/929/650225.

Strom:2017:OMP

[STC^{+}17] John Fredrik Strøm, Eva B. Thorstad, Graham Chafe, Sigrunn H. Sørbye, David Righton, Audun H. Rikardsen, and Jonathan Carr. Ocean migration of pop-up satellite archival tagged Atlantic salmon from the Miramichi River in Canada. ICES Journal of Marine Science, 74(5):1356– 1370, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/5/1356/2901808.

Stevenson:2011:FBI

[STD11] Todd C. Stevenson, Brian N. Tissot, and Jan Dierking. Fisher behaviour influences catch productivity and selectivity in West Hawaii’s aquarium fishery. ICES Journal of Marine Science, 68(5):813–822, May 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/5/813/648880.

Silk:2016:ECA

[STF^{+}16] Janet R. D. Silk, Sally E. Thorpe, Sophie Fielding, Eugene J. Murphy, Philip N. Trathan, Jonathan L. Watkins, and Simeon L. Hill. Environmental correlates of Antarctic krill distribution in the Scotia Sea and southern Drake Passage. ICES Journal of Marine Science, 73(9):2288– 2301, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/73/9/2288/2199119.

Schonberg:2017:VBR

[STF^{+}17] Christine H. L. Schönberg, Aline Tribollet, James K.-H. Fang, Marina Carreiro-Silva, and Max Wisshak. Viewpoints in bioerosion research — are we really disagreeing? A reply to the comment by Silbiger and DeCarlo (2017). ICES Journal of Marine Science, 74(9):2494–2500, November 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/9/2494/4317441. See [SD17].

Sassa:2016:IVD

[STKT16] Chiyuki Sassa, Motomitsu Takahashi, Yoshinobu Konishi, and Youichi Tsukamoto. Interannual variations in distribution and abundance of Japanese jack mackerel Trachurus japonicus larvae in the East China Sea. ICES Journal of Marine Science, 73(4):1170–1185, March 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/4/1170/2458926.

Sgardeli:2019:OEA

[STM19] Vasiliki Sgardeli, George Tserpes, and Christos D. Maravelias. Optimizing effort allocation in data poor mixed fisheries. ICES Journal of Marine Science, 76(6):1505– 1514, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/76/6/1505/5421313.

Strehlow:2010:MMP

[Str10] Harry V. Strehlow. The multiannual management plan for cod in the Baltic Sea: reactions and sentiments in two German fishing communities. ICES Journal of Marine Science, 67(9):1963–1971, December 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/67/
9/1963/623140.

Sigurdsson:2016:PAL

[STR16] Gudjon Mar Sigurdsson, Michael John Tremblay, and Rémy Rochette. Patchiness in American lobster benthic recruitment at a hierarchy of spatial scales. ICES Journal of Marine Science, 73(2):394–404, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/2/394/2614245.

Steenbergen:2017:CTN

[STSP17] Josien Steenbergen, Brita K. Trapman, Nathalie A. Steins, and Jan Jaap Poos. The commons tragedy in the North Sea brown shrimp fishery: how horizontal institutional interactions inhibit a self-governance structure. ICES Journal of Marine Science, 74(7):2004–2011, September 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/7/2004/3744774.

Subbey:2018:PESa

[Sub18a] Sam Subbey. Parameter estimation in stock assessment modelling: caveats with gradient-based algorithms. ICES Journal of Marine Science, 75(4):1511, July 2018.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/4/1511/4996293.

Subbey:2018:PESb

[Sub18b] Sam Subbey. Parameter estimation in stock assessment modelling: caveats with gradient-based algorithms. ICES Journal of Marine Science, 75(5):1553–1559, September 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/5/1553/4973625.

Sumaila:2013:HMP

[Sum13] U. Rashid Sumaila. How to make progress in disciplining overfishing subsidies. ICES Journal of Marine Science, 70(2):251–258, March 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/70/2/251/795042.

Shea:2010:OCD

[SV10] Elizabeth K. Shea and Michael Vecchione. Ontogenic changes in diel vertical migration patterns compared with known allometric changes in three mesopelagic squid species suggest an expanded definition of a paralarva. ICES Journal of Marine Science, 67(7):1436–1443, October 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/7/1436/665415.

Siskey:2016:FYF

[SWA^{+}16] M. R. Siskey, M. J. Wilberg, R. J. Allman, B. K. Barnett, and D. H. Secor. Forty years of fishing: changes in age structure and stock mixing in northwestern Atlantic bluefin tuna (Thunnus thynnus) associated with size-selective and long-term exploitation. ICES Journal of Marine Science, 73(10):2518–2528, November 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/10/2518/2647111.

Seitz:2014:EVC

[SWB^{+}14] Rochelle D. Seitz, Håkan Wennhage, Ulf Bergström, Romuald N. Lipcius, and Tom Ysebaert. Ecological value of coastal habitats for commercially and ecologically important species. ICES Journal of Marine Science, 71(3): 648–665, March 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/71/3/648/634683.

Stockwell:2013:UOS

[SWBJ13] Jason D. Stockwell, Thomas C. Weber, Adam J. Baukus, and J. Michael Jech. On the use of omnidirectional sonars and downwards-looking echosounders to assess pelagic fish distributions during and after midwater trawling. ICES Journal of Marine Science, 70(1):196–203, January 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/1/196/660983.

Szalaj:2018:GBF

[SWRC^{+}18] D. Szalaj, L. Wise, S. Rodríguez-Climent, M. M. Angélico, V. Marques, C. Chaves, A. Silva, and H. Cabral. A GISbased framework for addressing conflicting objectives in the context of an ecosystem approach to fisheries management — a case study of the Portuguese sardine fishery. ICES Journal of Marine Science, 75(6):2070–2087, November 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/6/2070/5062949.

Sundelof:2013:RHO

[SWS13] Andreas Sundelöf, Håkan Wennhage, and Henrik Svedäng. A red herring from the Öresund (ICES40G2): the apparent recovery of the Large Fish Indicator (LFI) in the North Sea hides a non-trawled area. ICES Journal of Marine Science, 70(6):1081–1084, September 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/6/1081/639321.

Stock:2019:USM

[SWT^{+}19] Brian C. Stock, Eric J. Ward, James T. Thorson, Jason E. Jannot, and Brice X. Semmens. The utility of spatial model-based estimators of unobserved bycatch. ICES Journal of Marine Science, 76(1):255–267, January 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/1/255/5144592.

Somerton:2011:UAE

[SWvSR11] David A. Somerton, Kresimir Williams, Paul G. von Szalay, and Craig S. Rose. Using acoustics to estimate the fish-length selectivity of trawl mesh. ICES Journal of Marine Science, 68(7):1558–1565, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/7/1558/658269.

Sugie:2016:EHC

[SY16] Koji Sugie and Takeshi Yoshimura. Effects of high CO_{2} levels on the ecophysiology of the diatom Thalassiosira weissflogii differ depending on the iron nutritional status. ICES Journal of Marine Science, 73(3):680–692, February 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/3/680/2459144.

Siddeek:2010:EBC

[SZM^{+}10] M. Shareef M. Siddeek, Jie Zheng, Joseph F. Morado, Gordon H. Kruse, and William R. Bechtol. Effect of bitter crab disease on rebuilding in Alaska Tanner crab stocks. ICES Journal of Marine Science, 67(9):2027–2032, December 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/67/9/2027/619756.

Secor:2019:ODF

[SZOL19] David H. Secor, Fan Zhang, Michael H. P. O’Brien, and Ming Li. Ocean destratification and fish evacuation caused by a mid-Atlantic tropical storm. ICES Journal of Marine Science, 76(2):573–584, March 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/2/573/4791423.

Trzcinski:2016:RAH

[TB16] M. Kurtis Trzcinski and W. Don Bowen. The recovery of Atlantic halibut: a large, long-lived, and exploited marine predator. ICES Journal of Marine Science, 73(4):1104– 1114, March 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/73/4/1104/2458915.

Thorson:2017:CEA

[TB17] James T. Thorson and Lewis A. K. Barnett. Comparing estimates of abundance trends and distribution shifts using single- and multispecies models of fishes and biogenic habitat. ICES Journal of Marine Science, 74(5):1311–1321, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/5/1311/2907795.

Toft:2014:MSM

[TBC^{+}14] J. E. Toft, J. L. Burke, M. P. Carey, C. K. Kim, M. Marsik, D. A. Sutherland, K. K. Arkema, A. D. Guerry, P. S. Levin, T. J. Minello, M. Plummer, M. H. Ruckelshaus, and H. M. Townsend. From mountains to sound: modelling the sensitivity of Dungeness crab and Pacific oyster to land–sea interactions in Hood Canal, WA. ICES Journal of Marine Science, 71(3):725–738, March 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/3/725/631931.

Thompson:2015:IPA

[TBC15] Paul M. Thompson, Kate L. Brookes, and Line S. Cordes. Integrating passive acoustic and visual data to model spatial patterns of occurrence in coastal dolphins. ICES Journal of Marine Science, 72(2):651–660, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/2/651/675254.

Taylor:2018:TCR

[TBD^{+}18] S. M. Taylor, S. J. Blight, C. J. Desfosses, A. S. Steffe, K. L. Ryan, A. M. Denham, and B. S. Wise. Thermographic cameras reveal high levels of crepuscular and nocturnal shore-based recreational fishing effort in an Australian estuary. ICES Journal of Marine Science, 75(6):2107–2116, November 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/6/2107/5039627.

Talwar:2016:APR

[TBG16] Brendan Talwar, Edward J. Brooks, and R. Dean Grubbs. An assessment of post-release mortality for a commonly discarded deep-sea isopod (Bathynomus giganteus) using reflex impairment. ICES Journal of Marine Science, 73(9):2356–2363, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
9/2356/2198931.

Turner:2017:DTS

[TBHK17] Joseph A. Turner, Russell C. Babcock, Renae Hovey, and Gary A. Kendrick. Deep thinking: a systematic review of mesophotic coral ecosystems. ICES Journal of Marine Science, 74(9):2309–2320, November 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/9/2309/3855117.

Tamdrari:2010:DID

[TCBD10] Hacène Tamdrari, Martin Castonguay, Jean-Claude Brêthes, and Daniel Duplisea. Density-independent and -dependent habitat selection of Atlantic cod (Gadus morhua) based on geostatistical aggregation curves in the northern Gulf of St Lawrence. ICES Journal of Marine Science, 67(8):1676– 1686, November 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/67/8/1676/608265.

Tremblay:2016:AUD

[TCD^{+}16] V. Tremblay, C. Cossette, J-D. Dutil, G. Verreault, and P. Dumont. Assessment of upstream and downstream passability for eel at dams. ICES Journal of Marine Science, 73(1):22–32, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/73/1/22/2458709.

Tzanatos:2013:MSI

[TCF^{+}13] Evangelos Tzanatos, Jose Castro, Aitor Forcada, Sanja Matić-Skoko, Miguel Gaspar, and Constantin Koutsikopoulos. A Métier–Sustainability–Index (MSI25) to evaluate fisheries components: assessment of cases from data-poor fisheries from southern Europe. ICES Journal of Marine Science, 70(1):78–98, January 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/70/1/78/
661623.

Tulp:2016:ABS

[TCH^{+}16] Ingrid Tulp, Chun Chen, Holger Haslob, Katharina Schulte, Volker Siegel, Josien Steenbergen, Axel Temming, and Marc Hufnagl. Annual brown shrimp (Crangon crangon) biomass production in Northwestern Europe contrasted to annual landings. ICES Journal of Marine Science, 73(10):2539–2551, November 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
10/2539/2647123.

Tian:2015:MSN

[TCQ^{+}15] Rucheng Tian, Changsheng Chen, Jianhua Qi, Rubao Ji, Robert C. Beardsley, and Cabell Davis. Model study of nutrient and phytoplankton dynamics in the Gulf of Maine: patterns and drivers for seasonal and interannual variability. ICES Journal of Marine Science, 72(2):388–402, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/2/388/671414.

Tanaka:2019:MBA

[TCS^{+}19] Kisei R. Tanaka, Jie Cao, Burton V. Shank, Samuel B. Truesdell, Mackenzie D. Mazur, Luoliang Xu, and Yong Chen. A model-based approach to incorporate environmental variability into assessment of a commercial fishery: a case study with the American lobster fishery in the Gulf of Maine and Georges Bank. ICES Journal of Marine Science, 76(4):884–896, July 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/4/884/5363241.

Thorpe:2019:CCF

[TD19] Robert B. Thorpe and José A. A. De Oliveira. Comparing conceptual frameworks for a fish community MSY (FCMSY) using management strategy evaluation — an example from the North Sea. ICES Journal of Marine Science, 76(4):813–823, July 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
4/813/5359482.

Trijoulet:2018:BMG

[TDHC18] Vanessa Trijoulet, Helen Dobby, Steven J. Holmes, and Robin M. Cook. Bioeconomic modelling of grey seal predation impacts on the West of Scotland demersal fisheries. ICES Journal of Marine Science, 75(4):1374–1382, July 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/4/1374/4803069.

Toledo:2019:PME

[TDN^{+}19] Pamela Toledo, Audrey M. Darnaude, Edwin J. Niklitschek, Vilma Ojeda, Raphaël Voué, Félix P. Leiva, Maylis Labonne, and Cristian B. Canales-Aguirre. Partial migration and early size of southern hake Merluccius australis: a journey between estuarine and oceanic habitats off Northwest Patagonia. ICES Journal of Marine Science, 76(4): 1094–1106, July 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/76/4/1094/5232582.

Thorpe:2016:AFE

[TDR^{+}16] Robert Brian Thorpe, Paul James Dolder, Stuart Reeves, Peter Robinson, and Simon Jennings. Assessing fishery and ecological consequences of alternate management options for multispecies fisheries. ICES Journal of Marine Science, 73(6):1503–1512, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
6/1503/2459085.

Thrush:2016:IFI

[TED16] Simon F. Thrush, Kari E. Ellingsen, and Kathryn Davis. Implications of fisheries impacts to seabed biodiversity and ecosystem-based management. ICES Journal of Marine Science, 73(suppl 1):S44–S50, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/73/
suppl_1/i44/2573991.

Trijoulet:2019:PMAa

[TFC^{+}19a] Vanessa Trijoulet, Gavin Fay, Kiersten L. Curti, Brian Smith, and Timothy J. Miller. Performance of multispecies assessment models: insights on the influence of diet data. ICES Journal of Marine Science, 76(6):1464–1476, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/6/1464/5475848.

Trijoulet:2019:PMAb

[TFC^{+}19b] Vanessa Trijoulet, Gavin Fay, Kiersten L. Curti, Brian Smith, and Timothy J. Miller. Performance of multispecies assessment models: insights on the influence of diet data. ICES Journal of Marine Science, 76(6):1938, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/6/1938/5511587.

Todd:2012:PPC

[TFM^{+}12] Christopher D. Todd, Kevin D. Friedland, Julian C. MacLean, Bryce D. Whyte, Ian C. Russell, Michael E. Lonergan, and Michael B. Morrissey. Phenological and phenotypic changes in Atlantic salmon populations in response to a changing climate. ICES Journal of Marine Science, 69(9):1686–1698, November 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/69/
9/1686/640751.

Tixier:2015:HAH

[TGDG15] Paul Tixier, Nicolas Gasco, Guy Duhamel, and Christophe Guinet. Habituation to an acoustic harassment device (AHD) by killer whales depredating demersal longlines. ICES Journal of Marine Science, 72(5):1673–1681, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/5/1673/756429.

Tixier:2015:MKW

[TGG^{+}15] Paul Tixier, Jade Vacquie Garcia, Nicolas Gasco, Guy Duhamel, and Christophe Guinet. Mitigating killer whale depredation on demersal longline fisheries by changing fishing practices. ICES Journal of Marine Science, 72(5):1610– 1620, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/5/1610/755146.

Torrissen:2012:GEP

[TGH^{+}12] Ole Torrissen, Kevin Alan Glover, Tore Haug, Ole Arve Misund, Hans Julius Skaug, and Matthias Kaiser. Good ethics or political and cultural censoring in science? ICES Journal of Marine Science, 69(4):493–497, May 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/4/493/632714.

Temming:2015:DPL

[TH15] Axel Temming and Marc Hufnagl. Decreasing predation levels and increasing landings challenge the paradigm of non-management of North Sea brown shrimp (Crangon crangon). ICES Journal of Marine Science, 72(3):804–823, March 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/3/804/697167.

Truesdell:2016:ESH

[THC16] Samuel B. Truesdell, Deborah R. Hart, and Yong Chen. Effects of spatial heterogeneity in growth and fishing effort on yield-per-recruit models: an application to the US Atlantic sea scallop fishery. ICES Journal of Marine Science, 73(4): 1062–1073, March 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/73/4/1062/2457902.

Then:2018:EFN

[THH18] Amy Y. Then, John M. Hoenig, and Quang C. Huynh. Estimating fishing and natural mortality rates, and catchability coefficient, from a series of observations on mean length and fishing effort. ICES Journal of Marine Science, 75(2): 610–620, March 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/2/610/4161439.

Then:2015:EPP

[THHH15] Amy Y. Then, John M. Hoenig, Norman G. Hall, and David A. Hewitt. Evaluating the predictive performance of empirical estimators of natural mortality rate using information on over 200 fish species. ICES Journal of Marine Science, 72(1):82–92, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/72/1/82/
2804320.

Then:2018:EPP

[THHH18] Amy Y. Then, John M. Hoenig, Norman G. Hall, and David A. Hewitt. Evaluating the predictive performance of empirical estimators of natural mortality rate using information on over 200 fish species. ICES Journal of Marine Science, 75(4):1509, July 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/75/
4/1509/4591652.

Tidd:2011:EEF

[THKP11] Alex N. Tidd, Trevor Hutton, Laurence T. Kell, and Gurpreet Padda. Exit and entry of fishing vessels: an evaluation of factors affecting investment decisions in the North Sea English beam trawl fleet. ICES Journal of Marine Science, 68(5):961–971, May 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/68/5/961/647587.

Thorson:2015:REE

[THM15] James T. Thorson, Allan C. Hicks, and Richard D. Methot, Jr. Random effect estimation of time-varying factors in stock synthesis. ICES Journal of Marine Science, 72(1): 178–185, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/72/1/178/814992.

Thorson:2011:AFA

[Tho11] James T. Thorson. Auxiliary and focal assessment models: a proof-of-concept involving time-varying catchability and fishery stock-status evaluation. ICES Journal of Marine Science, 68(10):2264–2276, November 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/10/2264/616466.

Thorson:2014:SCD

[Tho14] James T. Thorson. Standardizing compositional data for stock assessment. ICES Journal of Marine Science, 71(5): 1117–1128, July 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/71/5/1117/646485.

terHofstede:2011:CDF

[tHR11] Remment ter Hofstede and Adriaan D. Rijnsdorp. Comparing demersal fish assemblages between periods of contrasting climate and fishing pressure. ICES Journal of Marine Science, 68(6):1189–1198, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/68/
6/1189/711056.

Twist:2016:DNZ

[THR16] Brenton A. Twist, Christopher D. Hepburn, and William J. Rayment. Distribution of the New Zealand scallop (Pecten novaezealandiae) within and surrounding a customary fisheries area. ICES Journal of Marine Science, 73(2):384–393, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/2/384/2614470.

Tang:2018:STP

[THSM^{+}18] Feng Tang, Marthe Larsen Haarr, Bernard Sainte-Marie, Michel Comeau, M. John Tremblay, Julien Gaudette, and Rémy Rochette. Spatio-temporal patterns and reproductive costs of abnormal clutches of female American lobster, Homarus americanus, in eastern Canada. ICES Journal of Marine Science, 75(6):2045–2059, November 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/6/2045/5042979.

Trenkel:2016:OOI

[THW16] Verena M. Trenkel, Nils Olav Handegard, and Thomas C. Weber. Observing the ocean interior in support of integrated management. ICES Journal of Marine Science, 73 (8):1947–1954, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/73/8/1947/2198679.

Tugores:2010:LID

[TID^{+}10] M. Pilar Tugores, Magdalena Iglesias, Núria Díaz, Dolores Oñate, Joan Miquel, and Ana Giráldez. Latitudinal and interannual distribution of the European anchovy (Engraulis encrasicolus) and sardine (Sardina pilchardus) in the western Mediterranean, and sampling uncertainty in abundance estimates. ICES Journal of Marine Science, 67(8):1574– 1586, November 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/67/8/1574/604252.

Tang:2019:DHC

[TIS^{+}19] K. W. Tang, J. A. Ivory, S. Shimode, Y. Nishibe, and K. Takahashi. Dead heat: copepod carcass occurrence along the Japanese coasts and implications for a warming ocean. ICES Journal of Marine Science, 76(6):1825–1835, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/6/1825/5368121.

Thorpe:2017:RBC

[TJD17] Robert B. Thorpe, Simon Jennings, and Paul J. Dolder. Risks and benefits of catching pretty good yield in multispecies mixed fisheries. ICES Journal of Marine Science, 74(8):2097–2106, September 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
8/2097/3787892.

Thorson:2015:PSD

[TJH15] James T. Thorson, Olaf P. Jensen, and Ray Hilborn. Probability of stochastic depletion: an easily interpreted diagnostic for stock assessment modelling and fisheries management. ICES Journal of Marine Science, 72(2):428–435, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/2/428/2801337.

Thorisson:2011:UOC

[TJMC11] Konrad Thorisson, Ingibjörg G. Jónsdóttir, Gudrun Marteinsdottir, and Steven E. Campana. The use of otolith chemistry to determine the juvenile source of spawning cod in Icelandic waters. ICES Journal of Marine Science, 68(1): 98–106, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/1/98/628232.

Thorarinsdottir:2010:CES

[TJR^{+}10] Gudrun G. Thorarinsdóttir, Larry Jacobson, Stefan Áki Ragnarsson, Elena Guijarro Garcia, and Karl Gunnarsson. Capture efficiency and size selectivity of hydraulic clam dredges used in fishing for ocean quahogs (Arctica islandica): simultaneous estimation in the SELECT model. ICES Journal of Marine Science, 67(2):345–354, March 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/2/345/692769.

Trygonis:2018:CFS

[TK18] Vasilis Trygonis and Zacharias Kapelonis. Corrections of fish school area and mean volume backscattering strength by simulation of an omnidirectional multi-beam sonar. ICES Journal of Marine Science, 75(4):1496–1508, July 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/4/1496/4924926.

Tomiyasu:2016:RBB

[TKA^{+}16] Makoto Tomiyasu, Wan-Yu Kao, Koki Abe, Kenji Minami, Taro Hirose, Michio Ogawa, and Kazushi Miyashita. The relationship between body angle and target strength of ribbonfish (Trichiurus japonicus) displaying a vertical swimming motion. ICES Journal of Marine Science, 73(8):2049– 2057, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/73/8/2049/2198578.

Taranger:2015:RAE

[TKB^{+}15] Geir Lasse Taranger, Ørjan Karlsen, Raymond John Bannister, Kevin Alan Glover, Vivian Husa, Egil Karlsbakk, Bjørn Olav Kvamme, Karin Kroon Boxaspen, Pål Arne Bjørn, Bengt Finstad, Abdullah Sami Madhun, H. Craig Morton, and Terje Svåsand. Risk assessment of the environmental impact of Norwegian Atlantic salmon farming. ICES Journal of Marine Science, 72(3):997–1021, March 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/3/997/686282.

Thygesen:2019:ISM

[TKJB19] Uffe Høgsbro Thygesen, Kasper Kristensen, Teunis Jansen, and Jan E. Beyer. Intercalibration of survey methods using paired fishing operations and log-Gaussian Cox processes. ICES Journal of Marine Science, 76(4):1189–1199, July 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/4/1189/5280811.

Thanassekos:2016:IBA

[TLF16] Stéphane Thanassekos, Robert J. Latour, and Mary C. Fabrizio. An individual-based approach to year-class strength estimation. ICES Journal of Marine Science, 73(9):2252– 2266, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/73/9/2252/2198441.

Thebaud:2017:MMS

[TLK^{+}17] Olivier Thébaud, Jason S. Link, Bas Kohler, Marloes Kraan, Romain López, Jan Jaap Poos, Jörn O. Schmidt, and David C. Smith. Managing marine socio-ecological systems: picturing the future. ICES Journal of Marine Science, 74(7):1965–1980, September 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
7/1965/3058980.

Tsai:2015:APB

[TLPS15] Wen-Pei Tsai, Kwang-Ming Liu, André E. Punt, and Chi-Lu Sun. Assessing the potential biases of ignoring sexual dimorphism and mating mechanism in using a single-sex demographic model: the shortfin mako shark as a case study. ICES Journal of Marine Science, 72(3):793–803, March 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/3/793/700838.

Tam:2017:TEB

[TLR^{+}17] Jamie C. Tam, Jason S. Link, Axel G. Rossberg, Stuart I. Rogers, Philip S. Levin, Marie-Joëlle Rochet, Alida Bundy, Andrea Belgrano, Simone Libralato, Maciej Tomczak, Karen van de Wolfshaar, Fabio Pranovi, Elena Gorokhova, Scott I. Large, Nathalie Niquil, Simon P. R. Greenstreet, Jean-Noel Druon, Jurate Lesutiene, Marie Johansen, Izaskun Preciado, Joana Patricio, Andreas Palialexis, Paul Tett, Geir O. Johansen, Jennifer Houle, and Anna Rindorf. Towards ecosystem-based management: identifying operational food-web indicators for marine ecosystems. ICES Journal of Marine Science, 74(7):2040–2052, September 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/7/2040/2970046.

Thorson:2015:MEU

[TM15] James T. Thorson and Cóilín Minto. Mixed effects: a unifying framework for statistical modelling in fisheries biology. ICES Journal of Marine Science, 72(5):1245–1256, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/5/1245/758333.

Tull:2016:ESI

[TMG16] M. Tull, S. J. Metcalf, and H. Gray. The economic and social impacts of environmental change on fishing towns and coastal communities: a historical case study of Geraldton, Western Australia. ICES Journal of Marine Science, 73(5): 1437–1446, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/73/5/1437/2296432.

Trueman:2012:SIR

[TMP12] Clive N. Trueman, Kirsteen M. MacKenzie, and Martin R. Palmer. Stable isotopes reveal linkages between ocean climate, plankton community dynamics, and survival of two populations of Atlantic salmon (Salmo salar). ICES Journal of Marine Science, 69(5):784–794, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/5/784/652467.

Trifonova:2017:PER

[TMP^{+}17] Neda Trifonova, David Maxwell, John Pinnegar, Andrew Kenny, and Allan Tucker. Predicting ecosystem responses to changes in fisheries catch, temperature, and primary productivity with a dynamic Bayesian network model. ICES Journal of Marine Science, 74(5):1334–1343, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/5/1334/2798349.

Turner:2016:UHA

[TMR^{+}16] Sara M. Turner, John P. Manderson, David E. Richardson, John J. Hoey, and Jonathan A. Hare. Using habitat association models to predict alewife and blueback herring marine distributions and overlap with Atlantic herring and Atlantic mackerel: can incidental catches be reduced? ICES Journal of Marine Science, 73(7):1912–1924, July 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/7/1912/2457910.

Tenningen:2017:BAH

[TMR^{+}17] Maria Tenningen, Gavin J. Macaulay, Guillaume Rieucau, Héctor Peña, and Rolf J. Korneliussen. Behaviours of Atlantic herring and mackerel in a purse-seine net, observed using multibeam sonar. ICES Journal of Marine Science, 74(1):359–368, January 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
1/359/2670305.

Thode:2015:CCD

[TMS^{+}15] Aaron Thode, Delphine Mathias, Janice Straley, Victoria O’Connell, Linda Behnken, Dan Falvey, Lauren Wild, John Calambokidis, Gregory Schorr, Russell Andrews, Joseph Liddle, and Phillip Lestenkof. Cues, creaks, and decoys: using passive acoustic monitoring as a tool for studying sperm whale depredation. ICES Journal of Marine Science, 72(5): 1621–1636, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/72/5/1621/777470.

Tokunaga:2019:ECM

[TMW19] Kanae Tokunaga, Tsutom Miyata, and Hiroki Wakamatsu. Evolution of a co-management arrangement in Japanese offshore fisheries: current policy and legal framework and a case study of the tiger puffer fishery in Ise–Mikawa Bay. ICES Journal of Marine Science, 76(6):1567–1580, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/6/1567/5475455.

Takahashi:2016:ELT

[TNFC16] M. Takahashi, S. H. C. Noonan, K. E. Fabricius, and C. J. Collier. The effects of long-term in situ CO_{2} enrichment on tropical seagrass communities at volcanic vents. ICES Journal of Marine Science, 73(3):876–886, February 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/3/876/2458728.

Therkildsen:2010:DDE

[TNH^{+}10] Nina Overgaard Therkildsen, Einar Eg Nielsen, Karin Hüssy, Dorte Meldrup, and Audrey J. Geffen. Does DNA extraction affect the physical and chemical composition of historical cod (Gadus morhua) otoliths? ICES Journal of Marine Science, 67(6):1251–1259, September 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/6/1251/734177.

Tian:2013:SAT

[TNS13] Yongjun Tian, Kazuya Nashida, and Hideo Sakaji. Synchrony in the abundance trend of spear squid Loligo bleekeri in the Japan Sea and the Pacific Ocean with special reference to the latitudinal differences in response to the climate regime shift. ICES Journal of Marine Science, 70(5):968– 979, September 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/70/5/968/644467.

Tam:2013:TCF

[TNY^{+}13] Yin-Ki Tam, I-Hsun Ni, Cynthia Yau, Man-Yee Yan, Wai-Shan Chan, Sze-Man Chan, and Hsueh-Jung Lu. Tracking the changes of a fish community following a megascale reclamation and ensuing mitigation measures. ICES Journal of Marine Science, 70(6):1206–1219, September 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/6/1206/635684.

Tsujii:2016:MFW

[TOA^{+}16] Koki Tsujii, Mayuko Otsuki, Tomonari Akamatsu, Ikuo Matsuo, Kazuo Amakasu, Minoru Kitamura, Takashi Kikuchi, Kazushi Miyashita, and Yoko Mitani. The migration of fin whales into the southern Chukchi Sea as monitored with passive acoustics. ICES Journal of Marine Science, 73(8):2085–2092, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/8/2085/2198231.

Townsend:2014:CCW

[Tow14] Howard Townsend. Comparing and coupling a water quality and a fisheries ecosystem model of the Chesapeake Bay for the exploratory assessment of resource management strategies. ICES Journal of Marine Science, 71(3):703–712, March 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/71/3/703/631733.

Taylor:2017:FLH

[TPBL17] Matthew D. Taylor, Nicholas L. Payne, Alistair Becker, and Michael B. Lowry. Feels like home: homing of mature largebodied fish following translocation from a power-station canal. ICES Journal of Marine Science, 74(1):301–310, January 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/1/301/2670313.

Tuck:2015:SMC

[TPHC15] Ian D. Tuck, Darren M. Parsons, Bruce W. Hartill, and Stephen M. Chiswell. Scampi (Metanephrops challengeri) emergence patterns and catchability. ICES Journal of Marine Science, 72(S1):S199–S210, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/72/
suppl_1/i199/616941.

Tenningen:2019:EPS

[TPHdJ19] Maria Tenningen, Armin Pobitzer, Nils Olav Handegard, and Karen de Jong. Estimating purse seine volume during capture: implications for fish densities and survival of released unwanted catches. ICES Journal of Marine Science, 76(7):2481–2488, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/7/2481/5528147.

Toft:2011:MEE

[TPL11] Jodie E. Toft, André E. Punt, and L. Richard Little. Modelling the economic and ecological impacts of the transition to individual transferable quotas in the multispecies US west coast groundfish trawl fleet. ICES Journal of Marine Science, 68(7):1566–1579, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/7/1566/659841.

Tuck:2011:ASF

[TPS^{+}11] G. N. Tuck, R. A. Phillips, C. Small, R. B. Thomson, N. L. Klaer, F. Taylor, R. M. Wanless, and H. Arrizabalaga. An assessment of seabird–fishery interactions in the Atlantic Ocean. ICES Journal of Marine Science, 68(8):1628– 1637, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/8/1628/760591.

Tsagarakis:2014:MFD

[TPV14] K. Tsagarakis, A. Palialexis, and V. Vassilopoulou. Mediterranean fishery discards: review of the existing knowledge. ICES Journal of Marine Science, 71(5):1219– 1234, July 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/71/5/1219/638845.

Teilmann:2010:OSD

[TRH10] Jonas Teilmann, Frank Rigét, and Tero Harkonen. Optimizing survey design for Scandinavian harbour seals: population trend as an ecological quality element. ICES Journal of Marine Science, 67(5):952–958, July 2010.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/5/952/609604.

Teniswood:2013:QAM

[TRHB13] Clara M. H. Teniswood, Donna Roberts, William R. Howard, and Jodie E. Bradby. A quantitative assessment of the mechanical strength of the polar pteropod Limacina helicina antarctica shell. ICES Journal of Marine Science, 70(7):1499–1505, November 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/7/1499/608824.

Tidd:2016:EPT

[TRPH16] Alex N. Tidd, Chris Reid, Graham M. Pilling, and Shelton J. Harley. Estimating productivity, technical and efficiency changes in the Western Pacific purse-seine fleets. ICES Journal of Marine Science, 73(4):1226–1234, March 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/4/1226/2458757.

Thomalla:2015:HRV

[TRSM15] Sandy J. Thomalla, Marie-Fanny Racault, Sebastiaan Swart, and Pedro M. S. Monteiro. High-resolution view of the spring bloom initiation and net community production in the Subantarctic Southern Ocean using glider data. ICES Journal of Marine Science, 72(6):1999–2020, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/6/1999/926541.

Tyson:2015:ISF

[TSBV^{+}15] Chris Tyson, Judy Shamoun-Baranes, E. Emiel Van Loon, Kees (C. J.) Camphuysen, and Niels T. Hintzen. Individual specialization on fishery discards by lesser black-backed gulls (Larus fuscus). ICES Journal of Marine Science, 72(6):1882–1891, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/72/
6/1882/918142.

Thorson:2012:DAA

[TSP12] James T. Thorson, Ian J. Stewart, and André E. Punt. Development and application of an agent-based model to evaluate methods for estimating relative abundance indices for shoaling fish such as Pacific rockfish (Sebastes spp.). ICES Journal of Marine Science, 69(4):635–647, May 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/69/4/635/631390.

Tsikopoulou:2019:FGB

[TSP^{+}19] Irini Tsikopoulou, Chris J. Smith, Nadia K. Papadopoulou, Elena Eleftheriadou, and Ioannis Karakassis. A fishing ground benthic ecosystem improved during the economic crisis. ICES Journal of Marine Science, 76(2):402–409, March 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/2/402/5203445.

Tsai:2014:DAS

[TSPL14] Wen-Pei Tsai, Chi-Lu Sun, André E. Punt, and Kwang-Ming Liu. Demographic analysis of the shortfin mako shark, Isurus oxyrinchus, in the Northwest Pacific using a two-sex stage-based matrix model. ICES Journal of Marine Science, 71(7):1604–1618, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/7/1604/2804400.

Tseng:2013:STV

[TSS^{+}13] Chen-Te Tseng, Nan-Jay Su, Chi-Lu Sun, André E. Punt, Su-Zan Yeh, Don-Chung Liu, and Wei-Cheng Su. Spatial and temporal variability of the Pacific saury (Cololabis saira) distribution in the northwestern Pacific Ocean. ICES Journal of Marine Science, 70(5):991–999, September 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/5/991/643838.

Thorson:2015:GDG

[TSWS15] James T. Thorson, Andrew O. Shelton, Eric J. Ward, and Hans J. Skaug. Geostatistical delta-generalized linear mixed models improve precision for estimated abundance indices for West Coast groundfishes. ICES Journal of Marine Science, 72(5):1297–1310, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/72/
5/1297/767661.

Tseng:2011:ICD

[TSY^{+}11] Chen-Te Tseng, Chi-Lu Sun, Su-Zan Yeh, Shih-Chin Chen, Wei-Cheng Su, and Don-Chung Liu. Influence of climate-driven sea surface temperature increase on potential habitats of the Pacific saury (Cololabis saira). ICES Journal of Marine Science, 68(6):1105–1113, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/6/1105/713320.

Todd:2015:RIM

[TTG^{+}15] Victoria L. G. Todd, Ian B. Todd, Jane C. Gardiner, Erica C. N. Morrin, Nicola A. MacPherson, Nancy A. Di-Marzio, and Frank Thomsen. A review of impacts of marine dredging activities on marine mammals. ICES Journal of Marine Science, 72(2):328–340, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/2/328/676320.

Townhill:2018:HAB

[TTJ^{+}18] Bryony L. Townhill, Jonathan Tinker, Miranda Jones, Sophie Pitois, Veronique Creach, Stephen D. Simpson, Stephen Dye, Elizabeth Bear, and John K. Pinnegar. Harmful algal blooms and climate change: exploring future distribution changes. ICES Journal of Marine Science, 75(6):1882–1893, November 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/75/
6/1882/5094977.

Townhill:2019:HAB

[TTJ^{+}19] Bryony L. Townhill, Jonathan Tinker, Miranda Jones, Sophie Pitois, Veronique Creach, Stephen D. Simpson, Stephen Dye, Elizabeth Bear, and John K. Pinnegar. Harmful algal blooms and climate change: exploring future distribution changes. ICES Journal of Marine Science, 76 (1):353, January 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/76/1/353/5270873.

Towers:2019:MDB

[TTR^{+}19] Jared R. Towers, Paul Tixier, Katherine A. Ross, John Bennett, John P. Y. Arnould, Robert L. Pitman, and John W. Durban. Movements and dive behaviour of a toothfish-depredating killer and sperm whale. ICES Journal of Marine Science, 76(1):298–311, January 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/1/298/5103434.

Turrell:2019:MSWa

[Tur19a] W. R. Turrell. Marine science within a net-zero emission statutory framework. ICES Journal of Marine Science, 76(7):1983–1993, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
7/1983/5569821.

Turrell:2019:MSWb

[Tur19b]  W. R. Turrell.  Marine science within a net-zero

emission statutory framework. ICES Journal of Marine Science, 76(7):2500, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
7/2500/5614225.

Tian:2014:CFF

[TUUC14] Yongjun Tian, Kazuhisa Uchikawa, Yuji Ueda, and Jiahua Cheng. Comparison of fluctuations in fish communities and trophic structures of ecosystems from three currents around Japan: synchronies and differences. ICES Journal of Marine Science, 71(1):19–34, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/1/19/646293.

Torniainen:2014:MCT

[TVJ^{+}14] Jyrki Torniainen, Pekka J. Vuorinen, Roger I. Jones, Marja Keinänen, Stefan Palm, Kristiina A. M. Vuori, and Mikko Kiljunen. Migratory connectivity of two Baltic Sea salmon populations: retrospective analysis using stable isotopes of scales. ICES Journal of Marine Science, 71(2):336–344, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/71/2/336/788099.

Tenningen:2012:RHH

[TVO12] Maria Tenningen, Aud Vold, and Rolf Erik Olsen. The response of herring to high crowding densities in purseseines: survival and stress reaction. ICES Journal of Marine Science, 69(8):1523–1531, September 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/8/1523/702815.

Tiano:2019:AIB

[TWB^{+}19] Justin C. Tiano, Rob Witbaard, Magda J. N. Bergman, Pieter van Rijswijk, Anton Tramper, Dick van Oevelen, and Karline Soetaert. Acute impacts of bottom trawl gears on benthic metabolism and nutrient cycling. ICES Journal of Marine Science, 76(6):1917–1930, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/6/1917/5475867.

Talbot:2019:ADE

[TWHB19] S. Elizabeth Talbot, Stephen Widdicombe, Chris Hauton, and Jorn Bruggeman. Adapting the dynamic energy budget (DEB) approach to include non-continuous growth (moulting) and provide better predictions of biological performance in crustaceans. ICES Journal of Marine Science, 76(1):192–205, January 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
1/192/5200553.

Uglem:2011:DWF

[UBV^{+}11] Ingebrigt Uglem, Marius Berg, Rebecca Varne, Rune Nilsen, Jarle Mork, and Pål Arne Bjørn. Discrimination of wild and farmed Atlantic cod (Gadus morhua) based on morphology and scale-circuli pattern. ICES Journal of Marine Science, 68(9):1928–1936, September 2011.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/9/1928/667832.

Uhlmann:2011:MDP

[UBvH11] Sebastian S. Uhlmann, Stijn M. Bierman, and Aloysius T. M. van Helmond. A method of detecting patterns in mean lengths of samples of discarded fish, applied to the self-sampling programme of the Dutch bottom-trawl fishery. ICES Journal of Marine Science, 68(8):1712–1718, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/8/1712/750498.

Ulvan:2018:RMA

[UFJ^{+}18] Eva Marita Ulvan, Anders Foldvik, Arne Johan Jensen, Bengt Finstad, Eva Bonsak Thorstad, Audun Håvard Rikardsen, and Tor Fredrik Næsje. Return migration of adult Atlantic salmon (Salmo salar L.) to northern Norway. ICES Journal of Marine Science, 75(2):653–661, March 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/2/653/4161440.

Uusitalo:2013:NAE

[UFLH^{+}13] Laura Uusitalo, Vivi Fleming-Lehtinen, Heidi Hällfors, Andres Jaanus, Seija Hällfors, and Lauri London. A novel approach for estimating phytoplankton biodiversity. ICES Journal of Marine Science, 70(2):408–417, March 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/2/408/797699.

Urmy:2012:MVD

[UHB12] Samuel S. Urmy, John K. Horne, and David H. Barbee. Measuring the vertical distributional variability of pelagic fauna in Monterey Bay. ICES Journal of Marine Science, 69(2):184–196, March 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/69/2/184/701699.

Ulrich:2015:DCD

[UOB^{+}15] Clara Ulrich, Hans Jakob Olesen, Heidrikur Bergsson, Josefine Egekvist, Kirsten Birch Håkansson, Jørgen Dalskov, Lotte Kindt-Larsen, and Marie Storr-Paulsen. Discarding of cod in the Danish fully documented fisheries trials. ICES Journal of Marine Science, 72(6):1848–1860, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/6/1848/918592.

Urquhart:2010:SPM

[UPF^{+}10] Katy Urquhart, Campbell C. Pert, Rob J. Fryer, Paul Cook, Sarah Weir, Rachel Kilburn, Una McCarthy, Judy Simons, Sonia J. McBeath, Iveta Matejusova, and Ian R. Bricknell. A survey of pathogens and metazoan parasites on wild sea trout (Salmo trutta) in Scottish waters. ICES Journal of Marine Science, 67(3):444–453, April 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/3/444/733826.

Urban:2015:DMR

[Urb15] J. Daniel Urban. Discard mortality rates in the Bering Sea snow crab, Chionoecetes opilio, fishery. ICES Journal of Marine Science, 72(5):1525–1529, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/72/
5/1525/772588.

Underwood:2018:SSR

[URE^{+}18] Melanie J. Underwood, Shale Rosen, Arill Engås, Terje Jørgensen, and Anders Fernö. Species-specific residence times in the aft part of a pelagic survey trawl: implications for inference of pre-capture spatial distribution using the deep vision system. ICES Journal of Marine Science, 75(4): 1393–1404, July 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/4/1393/4829680.

Ulrich:2011:RSS

[URV^{+}11] Clara Ulrich, Stuart A. Reeves, Youen Vermard, Steven J. Holmes, and Willy Vanhee. Reconciling single-species TACs in the North Sea demersal fisheries using the Fcube mixed-fisheries advice framework. ICES Journal of Marine Science, 68(7):1535–1547, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/68/
7/1535/655273.

Uhlmann:2016:IRI

[UTA^{+}16] Sebastian S. Uhlmann, Ruben Theunynck, Bart Ampe, Marieke Desender, Maarten Soetaert, and Jochen Depestele. Injury, reflex impairment, and survival of beamtrawled flatfish. ICES Journal of Marine Science, 73(4): 1244–1254, March 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/73/4/1244/2458731.

Ulrich:2017:AMS

[UVD^{+}17] Clara Ulrich, Youen Vermard, Paul J. Dolder, Thomas Brunel, Ernesto Jardim, Steven J. Holmes, Alexander Kempf, Lars O. Mortensen, Jan-Jaap Poos, and Anna Rindorf. Achieving maximum sustainable yield in mixed fisheries: a management approach for the North Sea demersal fisheries. ICES Journal of Marine Science, 74(2): 566–575, March 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/74/2/566/2669548.

Uhlmann:2014:DFE

[UvHS^{+}14] Sebastian S. Uhlmann, Aloysius T. M. van Helmond, Elísabet Kemp Stefánsdóttir, Sigrídur Sigurdardóttir, John Haralabous, Jose Maria Bellido, A. Carbonell, Tom Catchpole, Dimitrios Damalas, Laurence Fauconnet, Jordan Feekings, Teresa Garcia, Niels Madsen, Sandra Mallold, Sveinn Margeirsson, Andreas Palialexis, Lisa Readdy, Julio Valeiras, Vassiliki Vassilopoulou, and Marie-Joëlle Rochet. Discarded fish in European waters: general patterns and contrasts. ICES Journal of Marine Science, 71(5):1235– 1245, July 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/71/5/1235/637441.

VanDover:2011:MSM

[Van11] Cindy Lee Van Dover. Mining seafloor massive sulphides and biodiversity: what is at risk? ICES Journal of Marine Science, 68(2):341–348, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/2/341/616315.

Valdimarsson:2012:HVI

[VAP12] Hédinn Valdimarsson, Olafur S. Astthorsson, and Jonbjorn Palsson. Hydrographic variability in Icelandic waters during recent decades and related changes in distribution of some fish species. ICES Journal of Marine Science, 69(5): 816–825, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/69/5/816/648504.

Velho:2010:DND

[VBA10] Filomena Vaz Velho, Pedro Barros, and Bjørn Erik Axelsen. Day–night differences in cunene horse mackerel (Trachurus trecae) acoustic relative densities off Angola. ICES Journal of Marine Science, 67(5):1004–1009, July 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/5/1004/610010.

Verisimo:2011:CBC

[VBGG^{+}11] Patricia Verísimo, Cristina Bernárdez, Eduardo González-Gurriarán, Juan Freire, Ramón Muiño, and Luis Fernández. Changes between consecutive broods in the fecundity of the spider crab, Maja brachydactyla. ICES Journal of Marine Science, 68(3):472–478, March 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/68/
3/472/655364.

Volkenandt:2015:PHD

[VBO^{+}15] Mareike Volkenandt, Simon Berrow, Ian O’Connor, Jean-Marc Guarini, and Ciaran O’Donnell. Prespawning herring distribution in the Irish Celtic Sea between 2005 and 2012. ICES Journal of Marine Science, 72(2):498–507, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/2/498/2801379.

Vaudo:2018:HVM

[VBW^{+}18] J. J. Vaudo, M. E. Byrne, B. M. Wetherbee, G. M. Harvey, A. Mendillo, and M. S. Shivji. Horizontal and vertical movements of white marlin, Kajikia albida, tagged off the Yucatán Peninsula. ICES Journal of Marine Science, 75(2): 844–857, March 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/2/844/4259720.

Verspoor:2012:RMS

[VCF^{+}12] Eric Verspoor, Sonia Consuegra, Olafur Fridjonsson, Sigridur Hjorleifsdottir, David Knox, Kristinn Olafsson, Scott Tompsett, Vidar Wennevik, and Carlos Garcia de Leaniz. Regional mtDNA SNP differentiation in European Atlantic salmon (Salmo salar): an assessment of potential utility for determination of natal origin. ICES Journal of Marine Science, 69(9):1625–1636, November 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/9/1625/635799.

Vestfals:2016:CHU

[VCH16] Cathleen D. Vestfals, Lorenzo Ciannelli, and Gerald R. Hoff. Changes in habitat utilization of slope-spawning flatfish across a bathymetric gradient. ICES Journal of Marine Science, 73(7):1875–1889, July 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
7/1875/2458913.

Vasquez:2013:IOP

[VCRPS13] Sebastián Vásquez, Marco Correa-Ramírez, Carolina Parada, and Aquiles Sepúlveda. The influence of oceanographic processes on jack mackerel (Trachurus murphyi) larval distribution and population structure in the southeastern Pacific Ocean. ICES Journal of Marine Science, 70(6):1097–1107, September 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/70/
6/1097/637731.

vonderHeyden:2010:MMD

[vdHBSM10] Sophie von der Heyden, Jaco Barendse, Anthony J. Seebregts, and Conrad A. Matthee. Misleading the masses: detection of mislabelled and substituted frozen fish products in South Africa. ICES Journal of Marine Science, 67(1):176–185, January 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/67/1/176/595224.

vanderHammen:2016:ECM

[vdHdGL16] Tessa van der Hammen, Martin de Graaf, and Jeremy M. Lyle. Estimating catches of marine and freshwater recreational fisheries in the Netherlands using an online panel survey. ICES Journal of Marine Science, 73(2):441–450, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/2/441/2614378.

vanDenderen:2015:TAB

[vDHvKR15] P. Daniël van Denderen, Niels T. Hintzen, Tobias van Kooten, and Adriaan D. Rijnsdorp. Temporal aggregation of bottom trawling and its implication for the impact on the benthic ecosystem. ICES Journal of Marine Science, 72(3): 952–961, March 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/72/3/952/2835891.

Varpe:2015:FEV

[VDK15] Øystein Varpe, Malin Daase, and Trond Kristiansen. A fish-eye view on the new Arctic lightscape. ICES Journal of Marine Science, 72(9):2532–2538, November 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/9/2532/2457864.

Vollset:2018:DRS

[VDK^{+}18] Knut Wiik Vollset, Ian Dohoo, Ørjan Karlsen, Elina Halttunen, Bjørn Olav Kvamme, Bengt Finstad, Vidar Wennevik, Ola H. Diserud, Andrew Bateman, Kevin D. Friedland, Shad Mahlum, Christian Jørgensen, Lars Qviller, Martin Krkošek, Åse Åtland, and Bjørn Torgeir Barlaup. Disentangling the role of sea lice on the marine survival of Atlantic salmon. ICES Journal of Marine Science, 75(1): 50–60, January 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/1/50/3867702.

vanderKooij:2016:ORA

[vdKFS^{+}16] Jeroen van der Kooij, Sascha M. M. Fässler, David Stephens, Lisa Readdy, Beth E. Scott, and Beatriz A. Roel. Opportunistically recorded acoustic data support Northeast Atlantic mackerel expansion theory. ICES Journal of Marine Science, 73(4):1115–1126, March 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/4/1115/2458700.

vanderKooij:2011:DHD

[vdKKS11] Jeroen van der Kooij, Sven Kupschus, and Beth E. Scott. Delineating the habitat of demersal fish assemblages with acoustic seabed technologies. ICES Journal of Marine Science, 68(9):1973–1985, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/9/1973/669292.

vanDeurs:2014:DCS

[vDKR14] Mikael van Deurs, Marja Koski, and Anna Rindorf. Does copepod size determine food consumption of particulate feeding fish? ICES Journal of Marine Science, 71(1):35–43, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/71/1/35/644277.

vanderReijden:2017:SUP

[vdRMC^{+}17] K. J. van der Reijden, P. Molenaar, C. Chen, S. S. Uhlmann, P. C. Goudswaard, and B. van Marlen. Survival of undersized plaice (Pleuronectes platessa), sole (Solea solea), and dab (Limanda limanda) in North Sea pulsetrawl fisheries. ICES Journal of Marine Science, 74(6): 1672–1680, July 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/74/6/1672/3059374.

vanderSleen:2018:OIE

[vdSSM^{+}18] Peter van der Sleen, Christoph Stransky, John R. Morrongiello, Holger Haslob, Melita Peharda, and Bryan A. Black. Otolith increments in European plaice (Pleuronectes platessa) reveal temperature and density-dependent effects on growth. ICES Journal of Marine Science, 75(5):1655– 1663, September 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/5/1655/4850985.

vanDamme:2014:FRH

[vDTF^{+}14] Cindy J. G. van Damme, Anders Thorsen, Merete Fonn, Paula Alvarez, Dolores Garabana, Brendan O’Hea, José R. Perez, and Mark Dickey-Collas. Fecundity regulation in horse mackerel. ICES Journal of Marine Science, 71(3): 546–558, March 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/71/3/546/635500.

Vinther:2013:QRF

[VE13] Morten Vinther and Margit Eero. Quantifying relative fishing impact on fish populations based on spatiotemporal overlap of fishing effort and stock density. ICES Journal of Marine Science, 70(3):618–627, April 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/3/618/915321.

Vetemaa:2010:CFS

[VEA^{+}10] Markus Vetemaa, Redik Eschbaum, Anu Albert, Lauri Saks, Aare Verliin, Kristiina Jürgens, Martin Kesler, Kalvi Hubel, Rögnvaldur Hannesson, and Toomas Saat. Changes in fish stocks in an Estonian estuary: overfishing by cormorants? ICES Journal of Marine Science, 67(9):1972– 1979, December 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/67/9/1972/620105.

Vasconcelos:2014:PPH

[VELT14] Rita P. Vasconcelos, David B. Eggleston, Olivier Le Pape, and Ingrid Tulp. Patterns and processes of habitat-specific demographic variability in exploited marine species. ICES Journal of Marine Science, 71(3):638–647, March 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/3/638/633232.

Valderrama:2017:FES

[VF17] Diego Valderrama and KathrynAnn H. Fields. Flawed evidence supporting the metabolic theory of ecology may undermine goals of ecosystem-based fishery management: the case of invasive Indo-Pacific lionfish in the western Atlantic. ICES Journal of Marine Science, 74(5):1256–1267, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/5/1256/2742003.

VanBeveren:2016:FHS

[VFR^{+}16] Elisabeth Van Beveren, Jean-Marc Fromentin, Tristan Rouyer, Sylvain Bonhommeau, Pablo Brosset, and Claire Saraux. The fisheries history of small pelagics in the Northern Mediterranean. ICES Journal of Marine Science, 73(6): 1474–1484, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/73/6/1474/2458965.

Vishnyakova:2015:SSB

[VG15] Karina Vishnyakova and Pavel Gol’din. Seasonality of strandings and bycatch of harbour porpoises in the Sea of Azov: the effects of fisheries, weather conditions, and life history. ICES Journal of Marine Science, 72(3):981–991, March 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/3/981/695484.

vanGemert:2018:CFA

[vGA18a] Rob van Gemert and Ken H. Andersen. Challenges to fisheries advice and management due to stock recovery. ICES Journal of Marine Science, 75(6):1864–1870, November 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/6/1864/5068220.

vanGemert:2018:ILL

[vGA18b] Rob van Gemert and Ken H. Andersen. Implications of late-in-life density-dependent growth for fishery sizeat-entry leading to maximum sustainable yield. ICES Journal of Marine Science, 75(4):1296–1305, July 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/4/1296/4817093.

Visconti:2017:MRL

[VGB^{+}17] Giulia Visconti, Fabrizio Gianguzza, Emanuela Butera, Valentina Costa, Salvatrice Vizzini, Maria Byrne, and Paola Gianguzza. Morphological response of the larvae of Arbacia lixula to near-future ocean warming and acidification. ICES Journal of Marine Science, 74(4):1180–1190, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/4/1180/3744779.

Viana:2011:FDI

[VGWJ11] Mafalda Viana, Norman Graham, James G. Wilson, and Andrew L. Jackson. Fishery discards in the Irish Sea exhibit temporal oscillations and trends reflecting underlying processes at an annual scale. ICES Journal of Marine Science, 68(1):221–227, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/68/1/221/632386.

vanHoof:2010:CMA

[vH10] Luc van Hoof. Co-management: an alternative to enforcement? ICES Journal of Marine Science, 67(2):395–401, March 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/67/2/395/693675.

vanHoof:2013:DPE

[vH13] Luc van Hoof. Design or pragmatic evolution: applying ITQs in EU fisheries management. ICES Journal of Marine Science, 70(2):462–470, March 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/70/
2/462/797111.

vanHelmond:2015:HEE

[vHCP15] Aloysius T. M. van Helmond, Chun Chen, and Jan Jaap Poos. How effective is electronic monitoring in mixed bottom-trawl fisheries? ICES Journal of Marine Science, 72(4):1192–1200, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/72/4/1192/800914.

vanHelmond:2017:UEM

[vHCP17] Aloysius T. M. van Helmond, Chun Chen, and Jan Jaap Poos. Using electronic monitoring to record catches of sole (Solea solea) in a bottom trawl fishery. ICES Journal of Marine Science, 74(5):1421–1427, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/5/1421/2884306.

Vollset:2017:SLI

[VHF^{+}17] Knut Wiik Vollset, Elina Halttunen, Bengt Finstad, Ørjan Karlsen, Pål Arne Bjørn, and Ian Dohoo. Salmon lice infestations on sea trout predicts infestations on migrating salmon post-smolts. ICES Journal of Marine Science, 74(9):2354–2363, November 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
9/2354/3860036.

Viddi:2010:SSV

[VHGTFR10] Francisco A. Viddi, Rodrigo Hucke-Gaete, Juan P. Torres-Florez, and Sandra Ribeiro. Spatial and seasonal variability in cetacean distribution in the fjords of northern Patagonia, Chile. ICES Journal of Marine Science, 67(5):959– 970, July 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/67/5/959/607974.

Vidal:2010:DPS

[VHH10] Erica A. G. Vidal, Manuel Haimovici, and Vivian C. S. Hackbart. Distribution of paralarvae and small juvenile cephalopods in relation to primary production in an upwelling area off southern Brazil. ICES Journal of Marine Science, 67(7):1346–1352, October 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/7/1346/663980.

Voss:2011:TCB

[VHQ^{+}11] Rüdiger Voss, Hans-Harald Hinrichsen, Martin F. Quaas, Jörn O. Schmidt, and Olli Tahvonen. Temperature change and Baltic sprat: from observations to ecological–economic modelling. ICES Journal of Marine Science, 68(6):1244– 1256, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/68/6/1244/711976.

Vikebo:2010:EHD

[VHS^{+}10] Frode B. Vikebø, Åse Husebø, Aril Slotte, Erling Kåre Stenevik, and Vidar S. Lien. Effect of hatching date, vertical distribution, and interannual variation in physical forcing on northward displacement and temperature conditions of Norwegian spring-spawning herring larvae. ICES Journal of Marine Science, 67(9):1948–1956, December 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/67/9/1948/620033.

Voss:2011:EMP

[VHS^{+}11] Rüdiger Voss, Hans-Harald Hinrichsen, Daniel Stepputtis, Matthias Bernreuther, Bastian Huwer, Viola Neumann, and Jörn O. Schmidt. Egg mortality: predation and hydrography in the central Baltic. ICES Journal of Marine Science, 68(7):1379–1390, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/7/1379/656150.

Volstad:2011:PBS

[VKN^{+}11] J. H. Vølstad, K. Korsbrekke, K. H. Nedreaas, M. Nilsen, G. N. Nilsson, M. Pennington, S. Subbey, and R. Wienerroither. Probability-based surveying using self-sampling to estimate catch and effort in Norway’s coastal tourist fishery. ICES Journal of Marine Science, 68(8):1785–1791, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/8/1785/752535.

Vikebo:2012:STO

[VKS^{+}12] Frode B. Vikebø, Anton Korosov, Erling Kåre Stenevik, Åse Husebø, and Aril Slotte. Spatio-temporal overlap of hatching in Norwegian spring-spawning herring and the spring phytoplankton bloom at available spawning substrata. ICES Journal of Marine Science, 69(7):1298–1302, September 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/69/7/1298/745061.

Vehmaa:2012:CRS

[VKT^{+}12] Anu Vehmaa, Anke Kremp, Timo Tamminen, Hedvig Hogfors, Kristian Spilling, and Jonna Engström-Öst. Copepod reproductive success in spring-bloom communities with modified diatom and dinoflagellate dominance. ICES Journal of Marine Science, 69(3):351–357, May 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/3/351/597867.

Vasconcelos:2011:SSF

[VMAMAG11] Paulo Vasconcelos, António Morgado-André, Carlos Morgado-André, and Miguel B. Gaspar. Shell strength and fishing damage to the smooth clam (Callista chione): simulating impacts caused by bivalve dredging. ICES Journal of Marine Science, 68(1):32–42, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/68/
1/32/629897.

Verissimo:2011:PSD

[VMG11] Ana Veríssimo, Jan R. McDowell, and John E. Graves. Population structure of a deep-water squaloid shark, the Portuguese dogfish (Centroscymnus coelolepis). ICES Journal of Marine Science, 68(3):555–563, March 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/3/555/661444.

Vasilakopoulos:2011:MYD

[VOM11] Paraskevas Vasilakopoulos, Finbarr G. O’Neill, and C. Tara Marshall. Misspent youth: does catching immature fish affect fisheries sustainability? ICES Journal of Marine Science, 68(7):1525–1534, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/68/
7/1525/657754.

Vilata:2010:PFS

[VOS10] Juan Vilata, Doris Oliva, and Maritza Sepúlveda. The predation of farmed salmon by South American sea lions (Otaria flavescens) in southern Chile. ICES Journal of Marine Science, 67(3):475–482, April 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/67/
3/475/731368.

Vollestad:2019:ULA

[VP19] L. Asbjørn Vøllestad and Craig R. Primmer. Understanding local adaptation in a freshwater salmonid fish: evolution of a research programme. ICES Journal of Marine Science, 76(6):1404–1414, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/6/1404/5384518.

vanPutten:2019:PSI

[vPBLR19] Ingrid van Putten, Fabio Boschetti, Scott Ling, and Shane A. Richards. Perceptions of system-identity and regime shift for marine ecosystems. ICES Journal of Marine Science, 76(6):1736–1747, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/76/
6/1736/5475850.

vanPutten:2016:EED

[vPFF^{+}16] Ingrid E. van Putten, Stewart Frusher, Elizabeth A. Fulton, Alistair J. Hobday, Sarah M. Jennings, Sarah Metcalf, and Gretta T. Pecl. Empirical evidence for different cognitive effects in explaining the attribution of marine range shifts to climate change. ICES Journal of Marine Science, 73(5): 1306–1318, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/73/5/1306/2240649.

vanPutten:2013:RBF

[vPGFT13] Ingrid E. van Putten, Rebecca J. Gorton, Elizabeth A. Fulton, and Olivier Thebaud. The role of behavioural flexibility in a whole of ecosystem model. ICES Journal of Marine Science, 70(1):150–163, January 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/1/150/662711.

Vatnehol:2018:MAD

[VPH18] Sindre Vatnehol, Hector Peña, and Nils Olav Handegard. A method to automatically detect fish aggregations using horizontally scanning sonar. ICES Journal of Marine Science, 75(5):1803–1812, September 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/5/1803/4984479.

Vatnehol:2017:EVF

[VPO17] Sindre Vatnehol, Hector Peña, and Egil Ona. Estimating the volumes of fish schools from observations with multibeam sonars. ICES Journal of Marine Science, 74(3): 813–821, March 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/74/3/813/2709601.

Vinas:2010:GDB

[VPSV^{+}10] Jordi Viñas, Alexandra Pérez-Serra, Oriol Vidal, Jaime R. Alvarado Bremer, and Carles Pla. Genetic differentiation between eastern and western Mediterranean swordfish revealed by phylogeographic analysis of the mitochondrial DNA control region. ICES Journal of Marine Science, 67(6):1222–1229, September 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/67/
6/1222/737624.

VanIseghem:2011:ERE

[VQB^{+}11] Sylvie Van Iseghem, Emmanuelle Quillérou, Cécile Brigaudeau, Claire Macher, Olivier Guyader, and Fabienne Daurès. Ensuring representative economic data: survey data-collection methods in France for implementing the common fisheries policy. ICES Journal of Marine Science, 68(8):1792– 1799, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/68/8/1792/757663.

Valenzuela-Quinonez:2014:IPD

[VQGDAMdL14] Fausto Valenzuela-Quiñonez, John Carlos Garza, Juan A. De-Anda-Montañez, and Francisco J. García de León. Inferring past demographic changes in a critically endangered marine fish after fishery collapse. ICES Journal of Marine Science, 71(7):1619–1628, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/7/1619/2804417.

Vincent:2016:FBP

[VRF^{+}16] Cecile Vincent, Vincent Ridoux, Mike A. Fedak, Bernie J. McConnell, Carole E. Sparling, Jean-Pierre Leaute, Joffrey Jouma’a, and Jerome Spitz. Foraging behaviour and prey consumption by grey seals (Halichoerus grypus) — spatial and trophic overlaps with fisheries in a marine protected area. ICES Journal of Marine Science, 73(10):2653–2665, November 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/10/2653/2647099.

Vikebo:2014:STO

[VRL^{+}14] Frode B. Vikebø, Petter Rønningen, Vidar S. Lien, Sonnich Meier, Mark Reed, Bjørn Ådlandsvik, and Trond Kristiansen. Spatio-temporal overlap of oil spills and early life stages of fish. ICES Journal of Marine Science, 71(4):970– 981, May 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/71/4/970/665925.

Vicente:2016:IHP

[VSB^{+}16] Jan Vicente, Nyssa J. Silbiger, Billie A. Beckley, Charles W. Raczkowski, and Russell T. Hill. Impact of high pCO_{2} and warmer temperatures on the process of silica biomineralization in the sponge Mycale grandis. ICES Journal of Marine Science, 73(3):704–714, February 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/3/704/2459091.

Vinas:2014:GPS

[VSP^{+}14] Jordi Viñas, Núria Sanz, Luis Peñarrubia, Rosa-Maria Araguas, José-Luis García-Marín, Maria-Inés Roldán, and Carles Pla. Genetic population structure of European anchovy in the Mediterranean Sea and the Northeast Atlantic Ocean using sequence analysis of the mitochondrial DNA control region. ICES Journal of Marine Science, 71(2):391– 397, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/71/2/391/784190.

VanGinderdeuren:2014:SFP

[VVP^{+}14] Karl Van Ginderdeuren, Sofie Vandendriessche, Yves Prössler, Hakimu Matola, Magda Vincx, and Kris Hostens. Selective feeding by pelagic fish in the Belgian part of the North Sea. ICES Journal of Marine Science, 71(4):808– 820, May 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/71/4/808/667002.

Vanalderweireldt:2019:EGM

[VWMS19] Lucie Vanalderweireldt, Gesche Winkler, Marc Mingelbier, and Pascal Sirois. Early growth, mortality, and partial migration of striped bass (Morone saxatilis) larvae and juveniles in the St. Lawrence Estuary, Canada. ICES Journal of Marine Science, 76(7):2235–2246, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/7/2235/5528149.

Welsh:2016:SSE

[WABZ16] Stuart A. Welsh, Joni L. Aldinger, Melissa A. Braham, and Jennifer L. Zimmerman. Synergistic and singular effects of river discharge and lunar illumination on dam passage of upstream migrant yellow-phase American eels. ICES Journal of Marine Science, 73(1):33–42, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/1/33/2457863.

Wolfe:2016:IAL

[WAH^{+}16] A. Megan Wolfe, Susan E. Allen, Michal Hodal, Rich Pawlowicz, Brian P. V. Hunt, and Desiree Tommasi. Impact of advection loss due to wind and estuarine circulation on the timing of the spring phytoplankton bloom in a fjord. ICES Journal of Marine Science, 73(6):1589–1609, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/6/1589/2457904.

Walmsley:2018:CQG

[WBW^{+}18] Sarah Walmsley, Julie Bremner, Alan Walker, Jon Barry, and David Maxwell. Challenges to quantifying glass eel abundance from large and dynamic estuaries. ICES Journal of Marine Science, 75(2):727–737, March 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/2/727/4210157.

Wiebe:2010:APS

[WCK^{+}10] Peter H. Wiebe, Dezhang Chu, Stein Kaartvedt, Anna Hundt, Webjorn Melle, Egil Ona, and Paola Batta-Lona. The acoustic properties of Salpa thompsoni. ICES Journal of Marine Science, 67(3):583–593, April 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/3/583/733009.

Wright:2019:ISP

[WCR^{+}19] P. J. Wright, A. Christensen, T. Régnier, A. Rindorf, and M. van Deurs. Integrating the scale of population processes into fisheries management, as illustrated in the sandeel, Ammodytes marinus. ICES Journal of Marine Science, 76(6):1453–1463, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/6/1453/5365500.

Waggitt:2016:PHC

[WCT^{+}16] James J. Waggitt, Pierre W. Cazenave, Ricardo Torres, Benjamin J. Williamson, and Beth E. Scott. Predictable hydrodynamic conditions explain temporal variations in the density of benthic foraging seabirds in a tidal stream environment. ICES Journal of Marine Science, 73(10):2677– 2686, November 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/73/10/2677/2647098.

Waggitt:2018:RSP

[WDE^{+}18] James J. Waggitt, Holly K. Dunn, Peter G. H. Evans, Jan Geert Hiddink, Laura J. Holmes, Emma Keen, Ben D. Murcott, Marco Piano, P. E. Robins, Beth E. Scott, Jenny Whitmore, and Gemma Veneruso. Regional-scale patterns in harbour porpoise occupancy of tidal stream environments. ICES Journal of Marine Science, 75(2):701–710, March 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/2/701/4101444.

Whyte:2017:TCM

[WDG^{+}17] Callum Whyte, Keith Davidson, Linda Gilpin, Elaine Mitchell, Grigorios Moschonas, Sharon McNeill, and Paul Tett. Tracking changes to a microplankton community in a North Atlantic sea loch using the microplankton index PI(mp). ICES Journal of Marine Science, 74(1):311–325, January 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/1/311/2669547.

Wahle:2015:ALN

[WDOJ15] Richard A. Wahle, Lanny Dellinger, Scott Olszewski, and Phoebe Jekielek. American lobster nurseries of southern New England receding in the face of climate change. ICES Journal of Marine Science, 72(S1):S69–S78, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/72/suppl_1/i69/622623.

Welcomme:2011:OGC

[Wel11] Robin L. Welcomme. An overview of global catch statistics for inland fish. ICES Journal of Marine Science, 68(8):1751–1756, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/68/
8/1751/747359.

Woods:2018:EER

[WESS18] Pamela J. Woods, Bjarki Tór Elvarsson, Thorsteinn Sigurdsson, and Gunnar Stefánsson. Evaluating the effectiveness of real-time closures for reducing susceptibility of small fish to capture. ICES Journal of Marine Science, 75(1): 298–308, January 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/1/298/4093371.

Wise:2015:KBM

[WFM^{+}15] Laura Wise, Paulo Fonseca, Alberto G. Murta, Cristina Silva, Hugo Mendes, João P. Carvalho, Maria de Fátima Borges, and Aida Campos. A knowledge-based model for evaluating the impact of gear-based management measures under Europe’s new common fisheries policy. ICES Journal of Marine Science, 72(4):1140–1151, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/72/
4/1140/801998.

Williamson:2011:BAE

[WFS^{+}11] Robert Williamson, John G. Field, Frank A. Shillington, Astrid Jarre, and Anet Potgieter. A Bayesian approach for estimating vertical chlorophyll profiles from satellite remote sensing: proof-of-concept. ICES Journal of Marine Science, 68(4):792–799, March 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/68/4/792/646407.

Walker:2019:DLA

[WGCL^{+}19] Nicola D. Walker, Bernardo García-Carreras, Will J. F. Le Quesne, David L. Maxwell, and Simon Jennings. A datalimited approach for estimating fishing mortality rates and exploitation status of diverse target and non-target fish species impacted by mixed multispecies fisheries. ICES Journal of Marine Science, 76(4):824–836, July 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/4/824/5303699.

Walton:2018:TSM

[WHAA^{+}18] Mark E. M. Walton, Jamie Hayes, Mohsin Al-Ansi, Mohamed Abdallah, Ibrahim Al Maslamani, Mohammed Al-Mohannadi, Ismail Al-Shaikh, Tim D’Urban Jackson, Claire Szostek, Jack Egerton, Michel J. Kaiser, and Lewis Le Vay. Towards spatial management of fisheries in the Gulf: benthic diversity, habitat and fish distributions from Qatari waters. ICES Journal of Marine Science, 75(1):178– 189, January 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/1/178/4049573.

Waldbusser:2016:CCS

[WHH16] George G. Waldbusser, Burke Hales, and Brian A. Haley. Calcium carbonate saturation state: on myths and this or that stories. ICES Journal of Marine Science, 73(3):563– 568, February 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/73/3/563/2458755.

Wang:2017:RLC

[WHL^{+}17] Kang Wang, Brian P. V. Hunt, Cui Liang, Daniel Pauly, and Evgeny A. Pakhomov. Reassessment of the life cycle of the pteropod Limacina helicina from a high resolution interannual time series in the temperate North Pacific. ICES Journal of Marine Science, 74(7):1906–1920, September 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/7/1906/3589818.

Woods:2015:HCQ

[WHMP15] Pamela J. Woods, Daniel S. Holland, Gudrún Marteinsdóttir, and André E. Punt. How a catch–quota balancing system can go wrong: an evaluation of the species quota transformation provisions in the Icelandic multispecies demersal fishery. ICES Journal of Marine Science, 72(5):1257–1277, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/5/1257/772020.

Windsland:2014:DIR

[WHNS14] Kristin Windsland, Carsten Hvingel, Einar M. Nilssen, and Jan H. Sundet. Dispersal of the introduced red king crab (Paralithodes camtschaticus) in Norwegian waters: a tag-recapture study. ICES Journal of Marine Science, 71(7):1966–1976, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/71/
7/1966/666428.

Woods:2016:EBR

[WHP16] Pamela J. Woods, Daniel S. Holland, and André E. Punt. Evaluating the benefits and risks of species-transformation provisions in multispecies IFQ fisheries with joint production. ICES Journal of Marine Science, 73(7):1764–1773, July 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/7/1764/2458588.

Wilson:2011:RRR

[Wil11] Cara Wilson. The rocky road from research to operations for satellite ocean-colour data in fishery management. ICES Journal of Marine Science, 68(4):677–686, March 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/4/677/645604.

Windsland:2015:TNM

[Win15] Kristin Windsland. Total and natural mortality of red king crab (Paralithodes camtschaticus) in Norwegian waters: catch–curve analysis and indirect estimation methods. ICES Journal of Marine Science, 72(2):642–650, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/2/642/2801345.

Wall:2016:IAA

[WJM16] Carrie C. Wall, J. Michael Jech, and Susan J. McLean. Increasing the accessibility of acoustic data through global access and imagery. ICES Journal of Marine Science, 73(8):2093–2103, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
8/2093/2198304.

Wells:2015:NTR

[WKK^{+}15] R. J. David Wells, Michael J. Kinney, Suzanne Kohin, Heidi Dewar, Jay R. Rooker, and Owyn E. Snodgrass. Natural tracers reveal population structure of albacore (Thunnus alalunga) in the eastern North Pacific. ICES Journal of Marine Science, 72(7):2118–2127, October 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/7/2118/2457862.

Wigand:2013:PGA

[WKL13] Laura A. Wigand, Terrie Klinger, and Miles G. Logsdon. Patterns in groundfish abundance along the Eastern Bering Sea outer continental margin. ICES Journal of Marine Science, 70(6):1181–1197, September 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/6/1181/636413.

Whitlock:2017:MRE

[WKP^{+}17] Rebecca E. Whitlock, Juho Kopra, Tapani Pakarinen, Eero Jutila, Adrian W. Leach, Polina Levontin, Sakari Kuikka, and Atso Romakkaniemi. Mark-recapture estimation of mortality and migration rates for sea trout (Salmo trutta) in the northern Baltic Sea. ICES Journal of Marine Science, 74(1):286–300, January 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/1/286/2669568.

Watts:2010:TPS

[WKW^{+}10] Phillip C. Watts, Suzanne M. Kay, Drew Wolfenden, Clive J. Fox, Audrey J. Geffen, Stephen J. Kemp, and Richard D. M. Nash. Temporal patterns of spatial genetic structure and effective population size in European plaice (Pleuronectes platessa) along the west coast of Scotland and in the Irish Sea. ICES Journal of Marine Science, 67(4): 607–616, May 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/67/4/607/679843.

Williams:2011:BIN

[WLB11] Ashley J. Williams, L. Richard Little, and Gavin A. Begg. Balancing indigenous and non-indigenous commercial objectives in a coral reef finfish fishery. ICES Journal of Marine Science, 68(5):834–847, May 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/5/834/652226.

Wiebe:2013:IAN

[WLL^{+}13] Peter H. Wiebe, Gareth L. Lawson, Andone C. Lavery, Nancy J. Copley, Erich Horgan, and Albert Bradley. Improved agreement of net and acoustical methods for surveying euphausiids by mitigating avoidance using a net-based LED strobe light system. ICES Journal of Marine Science, 70(3):650–664, April 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/70/3/650/916210.

Williams:2013:CDA

[WLN^{+}13] Ashley J. Williams, Bruno M. Leroy, Simon J. Nicol, Jessica H. Farley, Naomi P. Clear, Kyne Krusic-Golub, and Campbell R. Davies. Comparison of daily- and annual- increment counts in otoliths of bigeye (Thunnus obesus), yellowfin (T. albacares), southern bluefin (T. maccoyii) and albacore (T. alalunga) tuna. ICES Journal of Marine Science, 70(7):1439–1450, November 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/7/1439/607820.

Walther:2014:BIE

[WM14] Yvonne M. Walther and Christian Möllmann. Bringing integrated ecosystem assessments to real life: a scientific framework for ICES. ICES Journal of Marine Science, 71(5):1183–1186, July 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/71/
5/1183/642891.

Wright:2011:IDM

[WMG11] Peter J. Wright, Colin P. Millar, and Fiona M. Gibb. Intrastock differences in maturation schedules of Atlantic cod, Gadus morhua. ICES Journal of Marine Science, 68(9):1918–1927, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/68/
9/1918/666918.

White:2017:INV

[WMG^{+}17] Jonathan White, Niall O’ Maoiléidigh, Paddy Gargan, Elvira de Eyto, Gerald Chaput, Willie Roche, Phil McGinnity, Walter W. Crozier, Paddy Boylan, Dennis Doherty, Kealan O’Higgins, Brian Kennedy, Ian Lawler, David Lyons, and Ferdia Marnell. Incorporating natural variability in biological reference points and population dynamics into management of Atlantic salmon (Salmo salar L.) stocks returning to home waters. ICES Journal of Marine Science, 74(3):888, March 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
3/888/2960057.

Wilson:2013:ERF

[WMJ13] Matthew T. Wilson, Kathryn L. Mier, and Christina M. Jump. Effect of region on the food-related benefits to age-0 walleye pollock (Theragra chalcogramma) in association with midwater habitat characteristics in the Gulf of Alaska. ICES Journal of Marine Science, 70(7):1396–1407, November 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/70/7/1396/611587.

Walker:2017:EES

[WMLJ17] Nicola D. Walker, David L. Maxwell, Will J. F. Le Quesne, and Simon Jennings. Estimating efficiency of survey and commercial trawl gears from comparisons of catch-ratios. ICES Journal of Marine Science, 74(5):1448–1457, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/5/1448/2938485.

White:2011:FEA

[WMN^{+}11] Emma White, Cóilín Minto, Conor P. Nolan, Erna King, Eugene Mullins, and Maurice Clarke. First estimates of age, growth, and maturity of boarfish (Capros aper): a species newly exploited in the Northeast Atlantic. ICES Journal of Marine Science, 68(1):61–66, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/1/61/630114.

Wakefield:2010:ABD

[WNM10] Corey B. Wakefield, Stephen J. Newman, and Brett W. Molony. Age-based demography and reproduction of hapuku, Polyprion oxygeneios, from the south coast of Western Australia: implications for management. ICES Journal of Marine Science, 67(6):1164–1174, September 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/6/1164/735784.

Wakefield:2013:CLH

[WNM^{+}13] Corey B. Wakefield, Stephen J. Newman, Ross J. Marriott, Dion K. Boddington, and David V. Fairclough. Contrasting life history characteristics of the eightbar grouper Hyporthodus octofasciatus (Pisces: Epinephelidae) over a large latitudinal range reveals spawning omission at higher latitudes. ICES Journal of Marine Science, 70(3):485–497, April 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/70/3/485/918902.

Werner:2015:MBD

[WNPY15] Timothy B. Werner, Simon Northridge, Kate McClellan Press, and Nina Young. Mitigating bycatch and depredation of marine mammals in longline fisheries. ICES Journal of Marine Science, 72(5):1576–1586, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/5/1576/781679.

Williams:2015:EPO

[WNW^{+}15] Ashley J. Williams, Stephen J. Newman, Corey B. Wakefield, Melanie Bunel, Tuikolongahau Halafihi, Jeremie Kaltavara, and Simon J. Nicol. Evaluating the performance of otolith morphometrics in deriving age compositions and mortality rates for assessment of data-poor tropical fisheries. ICES Journal of Marine Science, 72(7):2098–2109, October 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/7/2098/2457850.

White:2016:INV

[WÓG^{+}16] Jonathan White, Niall Ó Maoiléidigh, Paddy Gargan, Elvira de Eyto, Gerald Chaput, Willie Roche, Phil McGinnity, Walter W. Crozier, Paddy Boylan, Dennis Doherty, Kealan O’Higgins, Brian Kennedy, Ian Lawler, David Lyons, and Ferdia Marnell. Incorporating natural variability in biological reference points and population dynamics into management of Atlantic salmon (Salmo salar L.) stocks returning to home waters. ICES Journal of Marine Science, 73(6):1513–1524, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/6/1513/2458792.

Wakefield:2017:ABP

[WOW^{+}17] Corey B. Wakefield, Joseph M. O’Malley, Ashley J. Williams, Brett M. Taylor, Ryan S. Nichols, Tuikolongahau Halafihi, Robert L. Humphreys, Jeremie Kaltavara, Simon J. Nicol, and Stephen J. Newman. Ageing bias and precision for deep-water snappers: evaluating nascent otolith preparation methods using novel multivariate comparisons among readers and growth parameter estimates. ICES Journal of Marine Science, 74(1):193–203, January 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/1/193/2669576.

Waldo:2013:ISD

[WP13] Staffan Waldo and Anton Paulrud. ITQs in Swedish demersal fisheries. ICES Journal of Marine Science, 70(1):68–77, January 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/70/1/68/661197.

Wetzel:2016:IAR

[WPeEA16] Chantel R. Wetzel, André E. Punt, and Handling editor: Emory Anderson. The impact of alternative rebuilding strategies to rebuild overfished stocks. ICES Journal of Marine Science, 73(9):2190–2207, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/9/2190/2198555.

Walczowski:2012:CAW

[WPGW12] Waldemar Walczowski, Jan Piechura, Ilona Goszczko, and Piotr Wieczorek. Changes in Atlantic water properties: an important factor in the European Arctic marine climate. ICES Journal of Marine Science, 69(5):864–869, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/69/5/864/652808.

Wakefield:2015:MVR

[WPH^{+}15] Corey B. Wakefield, Ian C. Potter, Norman G. Hall, Rodney C. J. Lenanton, and Sybrand A. Hesp. Marked variations in reproductive characteristics of snapper (Chrysophrys auratus, Sparidae) and their relationship with temperature over a wide latitudinal range. ICES Journal of Marine Science, 72(8):2341–2349, September 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/8/2341/2458918.

Wakefield:2017:TGZ

[WPH^{+}17] Corey B. Wakefield, Ian C. Potter, Norman G. Hall, Rodney C. J. Lenanton, and Sybrand A. Hesp. Timing of growth zone formations in otoliths of the snapper, Chrysophrys auratus, in subtropical and temperate waters differ and growth follows a parabolic relationship with latitude. ICES Journal of Marine Science, 74(1):180–192, January 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/1/180/2669555.

Walter:2015:IET

[WPvBS15] B. Walter, J. Peters, J. E. E. van Beusekom, and M. A. St. John. Interactive effects of temperature and light during deep convection: a case study on growth and condition of the diatom Thalassiosira weissflogii. ICES Journal of Marine Science, 72(6):2061–2071, July 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/6/2061/917661.

Williams:2011:LSR

[WPWH11] Kresimir Williams, André E. Punt, Christopher D. Wilson, and John K. Horne. Length-selective retention of walleye pollock, Theragra chalcogramma, by midwater trawls. ICES Journal of Marine Science, 68(1):119–129, January 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/1/119/631245.

Wiebe:2012:IVN

[WRC^{+}12] Peter H. Wiebe, Bert Rudels, Steven X. Cadrin, Ken F. Drinkwater, and Alicia Lavin. Introduction to variability of the North Atlantic and its marine ecosystems, 2000–2009, the proceedings of an ICES/NAFO symposium held in Santander, Spain, 10–12 May 2011. ICES Journal of Marine Science, 69(5):697–702, July 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/69/5/697/654163.

Windle:2010:ESN

[WRDF10] Matthew J. S. Windle, George A. Rose, Rodolphe Devillers, and Marie-Josée Fortin. Exploring spatial non-stationarity of fisheries survey data using geographically weighted regression (GWR): an example from the Northwest Atlantic. ICES Journal of Marine Science, 67(1):145–154, January 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/1/145/595559.

Wright:2014:TUL

[Wri14] Peter J. Wright. Are there useful life history indicators of stock recovery rate in gadoids? ICES Journal of Marine Science, 71(6):1393–1406, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/6/1393/2835583.

Wilhelm:2013:DSS

[WRMJ13] Margit R. Wilhelm, Jean-Paul Roux, Coleen L. Moloney, and Astrid Jarre. Data from fur seal scats reveal when Namibian Merluccius capensis are hatched and how fast they grow. ICES Journal of Marine Science, 70(7):1429– 1438, November 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/70/7/1429/608936.

Weltersbach:2013:DAE

[WS13] Marc Simon Weltersbach and Harry V. Strehlow. Dead or alive — estimating post-release mortality of Atlantic cod in the recreational fishery. ICES Journal of Marine Science, 70(4):864–872, July 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/70/4/864/725642.

Westgaard:2017:LFS

[WSG^{+}17] Jon-Ivar Westgaard, Arved Staby, Jane Aanestad Godiksen, Audrey J. Geffen, Anders Svensson, Gregory Charrier, Henrik Svedäng, and Carl André. Large and fine scale population structure in European hake (Merluccius merluccius) in the Northeast Atlantic. ICES Journal of Marine Science, 74(5):1300–1310, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/5/1300/2999000.

Wakefield:2017:PBR

[WSGD^{+}17] Corey B. Wakefield, Julia Santana-Garcon, Stacey R. Dorman, Stuart Blight, Ainslie Denham, John Wakeford, Brett W. Molony, and Stephen J. Newman. Performance of bycatch reduction devices varies for chondrichthyan, reptile, and cetacean mitigation in demersal fish trawls: assimilating subsurface interactions and unaccounted mortality. ICES Journal of Marine Science, 74(1):343–358, January 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/1/343/2669559.

Wilson:2019:MGP

[WSW^{+}19] Robert E. Wilson, George K. Sage, Kate Wedemeyer, Sarah A. Sonsthagen, Damian M. Menning, Megan C. Gravley, Matthew G. Sexson, R. John Nelson, and Sandra L. Talbot. Micro-geographic population genetic structure within Arctic cod (Boreogadus saida) in Beaufort Sea of Alaska. ICES Journal of Marine Science, 76(6):1713– 1721, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/76/6/1713/5430868.

Williams:2013:EDD

[WVG^{+}13] Rob Williams, Gisli A. Vikingsson, Astthor Gislason, Christina Lockyer, Leslie New, Len Thomas, and Philip S. Hammond. Evidence for density-dependent changes in body condition and pregnancy rate of North Atlantic fin whales over four decades of varying environmental conditions. ICES Journal of Marine Science, 70(6):1273– 1280, September 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/70/6/1273/637166.

Westerberg:2016:RCS

[WW16a] Håkan Westerberg and Håkan Wickström. Response to comments by Svedäng and Cardinale. ICES Journal of Marine Science, 73(6):1613–1615, May 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
6/1613/2458927.

Westerberg:2016:SAE

[WW16b] Håkan Westerberg and Håkan Wickström. Stock assessment of eels in the Baltic: reconciling survey estimates to achieve quantitative analysis. ICES Journal of Marine Science, 73(1):75–83, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/73/1/75/2457858.

Wang:2015:DOF

[WWCO15] Na Wang, You-Gan Wang, Anthony J. Courtney, and Michael F. O’Neill. Deriving optimal fishing effort for managing Australia’s Moreton Bay multispecies trawl fishery with aggregated effort data. ICES Journal of Marine Science, 72(5):1278–1284, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/72/
5/1278/759976.

Woillez:2016:ETU

[WWI^{+}16] Mathieu Woillez, Paul D. Walline, James N. Ianelli, Martin W. Dorn, Christopher D. Wilson, and André E. Punt. Evaluating total uncertainty for biomass- and abundance-at-age estimates from eastern Bering Sea walleye pollock acoustic-trawl surveys. ICES Journal of Marine Science, 73(9):2208–2226, September 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
9/2208/2198276.

Waller:2017:LRP

[WWMF17] Jesica D. Waller, Richard A. Wahle, Halley McVeigh, and David M. Fields. Linking rising pCO_{2} and temperature to the larval development and physiology of the American lobster (Homarus americanus). ICES Journal of Marine Science, 74(4):1210–1219, May 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/4/1210/2669570.

Wang:2016:BCR

[WZZ^{+}16] Yu Wang, Rui Zhang, Qiang Zheng, Ye Deng, Joy D. Van Nostrand, Jizhong Zhou, and Nianzhi Jiao. Bacterioplankton community resilience to ocean acidification: evidence from microbial network analysis. ICES Journal of Marine Science, 73(3):865–875, February 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/73/3/865/2458761.

Xavier:2013:HDC

[XCRP13] José C. Xavier, Yves Cherel, Jim Roberts, and Uwe Piatkowski. How do cephalopods become available to seabirds: can fish gut contents from tuna fishing vessels be a major food source of deep-dwelling cephalopods? ICES Journal of Marine Science, 70(1):46–49, January 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/1/46/662137.

Xie:2012:CFN

[XMM12] Yunbo Xie, Catherine G. J. Michielsens, and Fiona J. Martens. Classification of fish and non-fish acoustic tracks using discriminant function analysis. ICES Journal of Marine Science, 69(2):313–322, March 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/69/
2/313/699098.

Xavier:2011:CMP

[XPC11] José C. Xavier, Richard A. Phillips, and Yves Cherel. Cephalopods in marine predator diet assessments: why identifying upper and lower beaks is important. ICES Journal of Marine Science, 68(9):1857–1864, September 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/9/1857/664851.

Xu:2012:ECM

[XS12] Cailin Xu and David C. Schneider. Efficacy of conservation measures for the American lobster: reproductive value as a criterion. ICES Journal of Marine Science, 69(10):1831– 1839, December 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/69/10/1831/622515.

Yaragina:2010:BPI

[Yar10] Natalia A. Yaragina. Biological parameters of immature, ripening, and non-reproductive, mature northeast Arctic cod in 1984–2006. ICES Journal of Marine Science, 67(9):2033–2041, December 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/67/
9/2033/619451.

Yatsu:2013:CFK

[YCY^{+}13] Akihiko Yatsu, Sanae Chiba, Yasuhiro Yamanaka, Shin ichi Ito, Yugo Shimizu, Masahide Kaeriyama, and Yoshioro Watanabe. Climate forcing and the Kuroshio/Oyashio ecosystem. ICES Journal of Marine Science, 70(5):922– 933, September 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/70/5/922/646538.

Yu:2019:HSM

[YCZY19] Wei Yu, Xinjun Chen, Yang Zhang, and Qian Yi. Habitat suitability modelling revealing environmental-driven abundance variability and geographical distribution shift of winter–spring cohort of neon flying squid Ommastrephes bartramii in the northwest Pacific Ocean. ICES Journal of Marine Science, 76(6):1722–1735, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/6/1722/5456841.

Yasue:2014:TST

[YDT14] Naotaka Yasue, Ryu Doiuchi, and Akinori Takasuka. Trophodynamic similarities of three sympatric clupeoid species throughout their life histories in the Kii Channel as revealed by stable isotope approach. ICES Journal of Marine Science, 71(1):44–55, January 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/71/
1/44/645325.

Yesson:2017:ITE

[YFG^{+}17] Chris Yesson, Jess Fisher, Taylor Gorham, Chris J. Turner, Nanette Hammeken Arboe, Martin E. Blicher, and Kirsty M. Kemp. The impact of trawling on the epibenthic megafauna of the west Greenland shelf. ICES Journal of Marine Science, 74(3):866–876, March 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/3/866/2730113.

Young:2019:ASC

[YFP^{+}19] Talia Young, Emma C. Fuller, Mikaela M. Provost, Kaycee E. Coleman, Kevin St. Martin, Bonnie J. Mc-Cay, and Malin L. Pinsky. Adaptation strategies of coastal fishing communities as species shift poleward. ICES Journal of Marine Science, 76(1):93–103, January 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/1/93/5199399.

Yellapu:2017:PST

[YJBL17] Bhargavi Yellapu, Andrew Jeffs, Stephen Battaglene, and Shane D. Lavery. Population subdivision in the tropical spiny lobster Panulirus ornatus throughout its Indo-West Pacific distribution. ICES Journal of Marine Science, 74(3):759–768, March 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/74/
3/759/2447882.

Yang:2018:STD

[YLLG18] Qing Yang, Hui Liu, Guize Liu, and Yanbin Gu. Spatiotemporal distribution pattern of Calanus sinicus and its relationship with climate variability in the northern Yellow Sea. ICES Journal of Marine Science, 75(2):764–772, March 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/75/2/764/4084298.

Young:2018:IDP

[YPN^{+}18] Talia Young, Jennifer Pincin, Philipp Neubauer, Sofía Ortega-García, and Olaf P. Jensen. Investigating diet patterns of highly mobile marine predators using stomach contents, stable isotope, and fatty acid analyses. ICES Journal of Marine Science, 75(5):1583–1590, September 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/5/1583/4955965.

Yates:2015:ISA

[YS15] K. L. Yates and D. S. Schoeman. Incorporating the spatial access priorities of fishers into strategic conservation planning and marine protected area design: reducing cost and increasing transparency. ICES Journal of Marine Science, 72(2):587–594, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/2/587/2801319.

Yasuma:2010:SCT

[YST^{+}10] Hiroki Yasuma, Kouichi Sawada, Yoshimi Takao, Kazushi Miyashita, and Ichiro Aoki. Swimbladder condition and target strength of myctophid fish in the temperate zone of the Northwest Pacific. ICES Journal of Marine Science, 67(1): 135–144, January 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/67/1/135/594800.

Yu:2016:MEC

[YYCC16] Wei Yu, Qian Yi, Xinjun Chen, and Yong Chen. Modelling the effects of climate variability on habitat suitability of jumbo flying squid, Dosidicus gigas, in the Southeast Pacific Ocean off Peru. ICES Journal of Marine Science, 73(2):239–249, January 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
2/239/2614468.

Zhou:2015:GAI

[ZBE^{+}15] Shijie Zhou, Rik C. Buckworth, Nick Ellis, Roy A. Deng, and Sean Pascoe. Getting all information out of logbooks: estimating banana prawn fishable biomass, catchability, and fishing power increase, with a focus on natural mortality. ICES Journal of Marine Science, 72(1):54–61, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/1/54/816867.

Zhou:2019:CPU

[ZCH19] Shijie Zhou, Robert A. Campbell, and Simon D. Hoyle. Catch per unit effort standardization using spatio-temporal models for Australia’s Eastern tuna and billfish fishery. ICES Journal of Marine Science, 76(6):1489–1504, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/6/1489/5374756.

Zhang:2015:AUM

[ZCR15] Chongliang Zhang, Yong Chen, and Yiping Ren. Assessing uncertainty of a multispecies size-spectrum model resulting from process and observation errors. ICES Journal of Marine Science, 72(8):2223–2233, September 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/8/2223/2458725.

Zhang:2011:DEH

[ZCW11] Yuying Zhang, Yong Chen, and Carl Wilson. Developing and evaluating harvest control rules with different biological reference points for the American lobster (Homarus americanus) fishery in the Gulf of Maine. ICES Journal of Marine Science, 68(7):1511–1524, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/7/1511/656815.

Zwolinski:2013:MNM

[ZD13] Juan P. Zwolinski and David A. Demer. Measurements of natural mortality for Pacific sardine (Sardinops sagax). ICES Journal of Marine Science, 70(7):1408–1415, November 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/70/7/1408/610035.

Zwolinski:2014:EPC

[ZD14] Juan P. Zwolinski and David A. Demer. Environmental and parental control of Pacific sardine (Sardinops sagax) recruitment. ICES Journal of Marine Science, 71(8):2198– 2207, October 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/71/8/2198/745945.

Zemeckis:2019:IDA

[ZDD^{+}19] Douglas R. Zemeckis, Micah J. Dean, Annamaria I. DeAngelis, Sofie M. Van Parijs, William S. Hoffman, Mark F. Baumgartner, Leila T. Hatch, Steven X. Cadrin, and Christopher H. McGuire. Identifying the distribution of Atlantic cod spawning using multiple fixed and glider-mounted acoustic technologies. ICES Journal of Marine Science, 76(6):1610–1625, November 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/6/1610/5475874.

Zhou:2019:DLM

[ZDF^{+}19] Shijie Zhou, Ross M. Daley, Michael Fuller, Cathy M. Bulman, and Alistair J. Hobday. A data-limited method for assessing cumulative fishing risk on bycatch. ICES Journal of Marine Science, 76(4):837–847, July 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/76/4/837/5303697.

Zimmermann:2017:CLM

[ZE17] Fabian Zimmermann and Katja Enberg. Can less be more? Effects of reduced frequency of surveys and stock assessments. ICES Journal of Marine Science, 74(1):56–68, January 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/74/1/56/2669553.

Zwolinski:2011:PHO

[ZED11] Juan P. Zwolinski, Robert L. Emmett, and David A. Demer. Predicting habitat to optimize sampling of Pacific sardine (Sardinops sagax). ICES Journal of Marine Science, 68(5):867–879, May 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/68/5/867/652826.

Zador:2017:LEP

[ZGK^{+}17] Stephani G. Zador, Sarah K. Gaichas, Stephen Kasperski, Colette L. Ward, Rachael E. Blake, Natalie C. Ban, Amber Himes-Cornell, and J. Zachary Koehn. Linking ecosystem processes to communities of practice through commercially fished species in the Gulf of Alaska. ICES Journal of Marine Science, 74(7):2024–2033, September 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/7/2024/3746139.

Zischke:2013:RGW

[ZGT13] Mitchell T. Zischke, Shane P. Griffiths, and Ian R. Tibbetts. Rapid growth of wahoo (Acanthocybium solandri) in the Coral Sea, based on length-at-age estimates using annual and daily increments on sagittal otoliths. ICES Journal of Marine Science, 70(6):1128–1139, September 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/6/1128/634994.

Zischke:2013:SIW

[ZGTL13] Mitchell T. Zischke, Shane P. Griffiths, Ian R. Tibbetts, and Robert J. G. Lester. Stock identification of wahoo (Acanthocybium solandri) in the Pacific and Indian Oceans using morphometrics and parasites. ICES Journal of Marine Science, 70(1):164–172, January 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/70/1/164/661815.

Zimmermann:2013:SDP

[ZH13] Fabian Zimmermann and Mikko Heino. Is size-dependent pricing prevalent in fisheries? The case of Norwegian demersal and pelagic fisheries. ICES Journal of Marine Science, 70(7):1389–1395, November 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/70/
7/1389/610595.

Zador:2017:ECA

[ZHAG17] Stephani G. Zador, Kirstin K. Holsman, Kerim Y. Aydin, and Sarah K. Gaichas. Ecosystem considerations in Alaska: the value of qualitative assessments. ICES Journal of Marine Science, 74(1):421–430, January 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/1/421/2669560.

Zavalaga:2010:MHU

[ZHD10] Carlos B. Zavalaga, Joanne Halls, and Giacomo Dell’Omo. Marine habitat use of Peruvian boobies: a geographic and oceanographic comparison between inshore and offshore islands. ICES Journal of Marine Science, 67(5):940–951, July 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/67/5/940/609008.

Zemeckis:2014:SSF

[ZHD^{+}14] Douglas R. Zemeckis, William S. Hoffman, Micah J. Dean, Michael P. Armstrong, and Steven X. Cadrin. Spawning site fidelity by Atlantic cod (Gadus morhua) in the Gulf of Maine: implications for population structure and rebuilding. ICES Journal of Marine Science, 71(6):1356– 1365, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/71/6/1356/2835586.

Zhong:2019:PCR

[ZHL^{+}19] Yanping Zhong, Jun Hu, Edward A. Laws, Xin Liu, Jixin Chen, and Bangqin Huang. Plankton community responses to pulsed upwelling events in the southern Taiwan Strait. ICES Journal of Marine Science, 76(7):2374–2388, December 2019. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/76/7/2374/5542622.

Zhang:2011:IAA

[ZHLK11] Chang Ik Zhang, Anne Babcock Hollowed, Jae-Bong Lee, and Do-Hoon Kim. An IFRAME approach for assessing impacts of climate change on fisheries. ICES Journal of Marine Science, 68(6):1318–1328, July 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/68/6/1318/713532.

Ziegler:2016:SSD

[ZHV^{+}16] Friederike Ziegler, Sara Hornborg, Daniel Valentinsson, Erik Skontorp Hognes, Guldborg Søvik, and Ole Ritzau Eigaard. Same stock, different management: quantifying the sustainability of three shrimp fisheries in the Skagerrak from a product perspective. ICES Journal of Marine Science, 73(7):1806–1814, July 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/73/
7/1806/2458710.

Zhou:2014:MMF

[ZKD^{+}14] Shijie Zhou, Neil L. Klaer, Ross M. Daley, Zhengyuan Zhu, Michael Fuller, and Anthony D. M. Smith. Modelling multiple fishing gear efficiencies and abundance for aggregated populations using fishery or survey data. ICES Journal of Marine Science, 71(9):2436–2447, November 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/71/9/2436/594507.

Zemeckis:2017:SMC

[ZLC^{+}17] Douglas R. Zemeckis, Chang Liu, Geoffrey W. Cowles, Micah J. Dean, William S. Hoffman, David Martins, and Steven X. Cadrin. Seasonal movements and connectivity of an Atlantic cod (Gadus morhua) spawning component in the western Gulf of Maine. ICES Journal of Marine Science, 74(6):1780–1796, July 2017. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/74/6/1780/2907792.

Zhou:2012:IRA

[ZMF12] Shijie Zhou, David A. Milton, and Gary C. Fry. Integrated risk analysis for rare marine species impacted by fishing: sustainability assessment and population trend modelling. ICES Journal of Marine Science, 69(2):271–280, March 2012. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/69/2/271/707384.

Zemeckis:2014:SIA

[ZMKC14] Douglas R. Zemeckis, David Martins, Lisa A. Kerr, and Steven X. Cadrin. Stock identification of Atlantic cod (Gadus morhua) in US waters: an interdisciplinary approach. ICES Journal of Marine Science, 71(6):1490– 1506, September 2014. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/71/6/1490/623280.

Zwolinski:2010:SPH

[ZOQS10] Juan P. Zwolinski, Paulo B. Oliveira, Victor Quintino, and Yorgos Stratoudakis. Sardine potential habitat and environmental forcing off western Portugal. ICES Journal of Marine Science, 67(8):1553–1564, November 2010. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/67/8/1553/604811.

Zhou:2018:OCO

[ZPS^{+}18] Shijie Zhou, André E. Punt, Anthony D. M. Smith, Yimin Ye, Malcolm Haddon, Cathy M. Dichmont, and David C. Smith. An optimized catch-only assessment method for data poor fisheries. ICES Journal of Marine Science, 75 (3):964–976, May 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/75/3/964/4772849.

Zhai:2011:PPS

[ZPT^{+}11] Li Zhai, Trevor Platt, Charles Tang, Shubha Sathyendranath, and Rafael Hernández Walls. Phytoplankton phenology on the Scotian Shelf. ICES Journal of Marine Science, 68(4):781–791, March 2011. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/68/4/781/646884.

Zhang:2018:ECD

[ZRN18] Fan Zhang, Kevin B. Reid, and Thomas D. Nudds. Ecosystem change and decadal variation in stock–recruitment relationships of Lake Erie yellow perch (Perca flavescens). ICES Journal of Marine Science, 75(2):531–540, March 2018. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/75/2/531/4364860.

Zarauz:2015:CDS

[ZRU^{+}15] L. Zarauz, J. Ruiz, A. Urtizberea, E. Andonegi, E. Mugerza, and I. Artetxe. Comparing different survey methods to estimate European sea bass recreational catches in the Basque country. ICES Journal of Marine Science, 72(4):1181–1191, May 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/72/4/1181/802575.

Zhang:2016:PRS

[ZSC16] Haoyu Zhang, Paul K. S. Shin, and Siu Gin Cheung. Physiological responses and scope for growth in a marine scavenging gastropod, Nassarius festivus (Powys, 1835), are affected by salinity and temperature but not by ocean acidification. ICES Journal of Marine Science, 73(3):814–824, February 2016. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/73/3/814/2458919.